A Complete Bibliography of Lecture Notes in Mathematics (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.06

Title word cross-reference

(7, 1) [Met91l]. (d > 1) [LO98b]. (k − 1) [Küh95a]. (n + 1, 1) [Met91a, Met91d]. (π_{N}^{S}, 32 ≤ N ≤ 45) [Koc90a]. (π_{N}^{S}, 46 ≤ N ≤ 64) [Koc90g]. (π_{N}, 9 ≤ N ≤ 31) [Koc90e]. ∗ [SZ90]. + [DP91c, Pat90]. 0 [Mar96f]. 1 [DRSŻ91b, Hub95b, Hub95g, Lea90, Mar91, Oda90a, Paz91, Rut97n, Rut97o, Yap96b]. 1/4 [Ham91]. 16^{th} [Den92]. 2 [BL92a, BR90f, Cos90a, DR91, Dub96, DRSŻ91b, Dũn92, Fre94, Ger92, Han90, HYL91a, HA90a, Huc90, KW94, Kim94, Kin94, Mar90b, Møl90, NZ99f, Oda90b, Räb91, Yap96b]. 256 [Sho92r, Sho92s]. 27 [Hun96a]. 2^{n} [GC92]. 2k [Küh95a]. 2l > n [Kar97c]. 2n [Cas92]. 3 [BP97c, BF93g, CG97e, CS92b, DRSŻ91a, Gra91a, GDT91, HM90, Joh95a, Joh95c, Joh95f, Kai97a, Kai97b, Kai97d, Kai97e, Kai97h, Kna95h, Küh95b, LP90a, McC90c, Wir98a]. 3n + 1 [Wir98g, Wir98f]. 4 [Aud90, HH92, Kna95e]. 4l > n + 1 [Kar97e]. 5 [AE90]. 7 [Rou90]. 9 [AE90]. [0, 1] [CV93]. [0, T] [Der95]. [X, Y] [Ark90b]. _{p}^{sq} [Li91]. A(5) [Met91l]. A(6) [Met91l]. A(t, B_{t}) [Wal93]. a∗_{p} [BZ92]. A_{3} [Koz91]. A_{5} [BK94, KW94, Kin94]. a_{f} [Mas95l]. A_{∞} [HSS91]. α [Thi91]. Aut(X(q)) [BOERT96a]. B_{dR} [Kat93a]. ¯∂ [Cha91n]. βN [Tod97d]. modP [Gre90]. C [Bek91a, Che90b, CHT92, deL94c]. c(a, p, q) [GE98f]. C(K, X) [CM97a]. C∗ [Pis96j, Tro91]. C[a, b] [Kit94c]. C^{∗} [NZ99a]. C^{+} − C^{−} [Pir94]. C^{0} [Bot92, Pil94e]. C^{1,α} [Cha91g]. C^{2l,l} [Mog92]. C^{∞} [Col92g, Rou90]. C^{n} [Che90a]. c_{0} [CM97d, GJ91]. C_{0}(Q) [Kit94c]. c_{1} [Xi94m]. c_{1}^{2} = 10 [DR91]. c_{2} = 3 [DR91]. C_{n−1} [Cha91b]. C_{n−1}(Q̄) [Cha91l]. C_{R}(X) [VM92]. CW_{A} [Far96g]. D [Ber90a, Ber90c, Bui92a, Bui92c, Bui92f, SS93, vdG93]. d(a, p, q) [GE98f]. d[X, X]_{t} = dt − X_{t−}^{+}dX_{t} [AR94]. D_{G}(X) [BL94a]. D_{MR} [Hub95a]. Δ [Mil94d]. δ_{3,r}(S, H) [MO93d]. divμ = ϵ [vW90]. E [Sło91]. e^{z} [VC90]. E_{2} [Lin90]. E_{8} [Shi91b]. E_{n} [Bia90a, Bia90b]. ℓ [Jan90c]. ℓ^{1} [Leu91]. ℓ^{∞} [Sch90h]. ℓ_{1} [CM97d]. ℓ_{p} [KO91]. ϵ [Eke90, Ger90, Sch91j]. ϵ(X × Y) [BH90b]. ϵ^{∗}(X) [Rut97b, Rut97a]. ϵ_{#}(X) [iM90a]. ϵ_{B}^{B}(X ×_{B} Y) [BH90b]. f [SM99u, SM99n]. F [a, b] [Kit94e]. F_{q} [Abr96a]. F_{q}[t] [Abr96a]. G [Bar93g, BV93, Bou97i, HL92, KV92, Kum97, Møl90, Mor90b, Sch94m, Sch94o, Sch94j, Tam99c, Win95b, Win95c]. G(T) [IK99a]. G/K [Mor90b]. G^{d+2} [JL91]. G_{T} (Θ) [MS94a]. Γ [Nie91d, Nie91g, Nie91j]. Γ_{0}(N) [Wan94]. Γ_{0}H, ⟨, ⟩ [Nie91a]. Γ_{2} [Bak91]. γ_{3}(S) [MO93d]. GL(2, p^{k}) [Sho92l, Sho92o]. GL(4, 2) [Sho92m]. GL(4, 3) [Sho92m]. GL(4, p^{k}) [Sho92t]. GL(6, 2) [Sho92p]. GL(n) [Hen90b, Oda90a]. GL(q, p^{k}) [Sho92u]. GL_{2} [Coa99]. GL_{n} [Har90b]. Gor(T) [IK99a, IK99e]. H [BL91a, Neu96a, Neu96c, Rut97h, SM99a, SM99h]. H ∗ (B(Z/p)^{r}) [HS92]. H ∗ V [BZ92]. H^{1} [Bea92, JcC91, Sar92]. H^{2}(C^{K}) [YSA92]. H^{∞} [Foi91, Gra99]. H^{P} [LS93, DK94, Han91, pL91, Suf90]. H_{∞} [Fra91, Kwa91, MP91]. Hilb^{3}P^{3} [RL90]. Hilb^{4}P^{2} [AV90]. hom_{A}(A, W) [Vas92]. I [Sło91]. ℑ [CP92]. ℑ_{n} [HP92]. IP_{4} [DPlPS90]. J [Koc90c, Muk90]. J_{0}(N) [Rib99]. J_{3,0} [Ura93f]. Jac(W_{d}^{1}(C)) [CHT92]. K [BS90a, CT93, Far96a, Gil92, Gub90, HKR92, IP92, Ina90a, Ina90b, Jan90b, Jan90d, Jan90c, Kan90, LPLG^{+}92, Lod92, Mur94, Sou92, Wah95i, Wür91, Bal90b, Bor92, Küh95h, MM90]. K/J [Mil94a]. K3 [Ura93e]. KK [Pus96a]. L [CP91e, CP91g, Fuj90a, HWW93a, Har90c, Kat93a, iM90b, Pan91h, Sch90x, Kin94]. L(n, d) [Met91o]. L[R] [Bek91d]. L^{1} [Pra92, Pea91]. L^{2} [Cha91c, Cha91f, Cha91b, Dem96, Lam92, Mit94a]. L^{2}(Γ \ G) [Kon90, Wal90c, Wan90]. L^{2}(F_{′}) + G_{T} (×) [MS94a]. L^{∞} [Gra99]. L^{p} [Dor93, Ebe99f, Ebe99a, GGS93, Iwa92a, Iwa92b, Nil94f, Röc99, RS91b, Tai91a, Tai91b, Üst95a, Zuc91]. L^{q} [FS92, RS91b]. L_{+}^{0} (Ω, F, P) [BS99]. L_{1}(G) [Ouy91]. L_{2} [Jaj90, Jaj91a, Jaj91f, Jaj91h, Wah95c, Wah95d]. L_{f} [Far96g]. L_{∞}(μ, X) [CM97h]. L_{P} [LS93, Wie90]. L_{p}(μ, X) [CM97d, CM97b]. L_{p}^{n} [SS91]. ≤ 0′ [HS90a]. M [HWW93b, HWW93c, HWW93d, HWW93e, HWW93f, HWW93g, MT91b]. M(∂_{t}) [KM97f]. M(∂_{t}) − ω(t) [KM97d]. M(∂_{t}) − ω_{0} [KM97e]. M(∂_{t} − ω(t)) [KM97g, KM97i]. C^{n} [Aiz94, Kul91]. Cp^{n} [BOWV91]. P^{1} [BFS98f]. P^{n} [BS92]. PV_{s}(u, v; r) [IK99b]. Q [KW94]. R [Arn99, Fou95]. R^{6} [Mul90c]. R^{d} [BIOS90]. R^{n} [Tur90a]. Z [Phi90]. Z[i] [Phi90]. Z_{2} [Kan90]. Z_{p} [vH90]. B [HMN91j]. FS [BFS98g]. P [vdPS97c]. H [Nie91d, Nie91g, Nie91j]. ∫↕_{\} [Ser92o]. μ [vW90]. N [HP92, Joh95c, Kit93, Mey99i, Son90, BT93, Day92, Met91n, Met91o]. N(p) [Acc94b]. n^{2} + n + 1 [Met91g]. n^{2} + n + 2 [Met91h]. ν [MO93g]. O′′′ [Dow90, BP97k]. O(3) [Bar93a, CA91]. O^{∗} [Ino98a, Ino98e]. O′ [Ars90]. o_{K} [Lan96a]. ∮ Δ^{0}(a) ∮ [Abr96c]. Ω [Far96g]. ω(t) [KM97g]. ω_{0} [KM97e]. ⊗ [Hub95a]. P [Mar96f, AS90a, And90, BBD90, Cri90, DG90, DNQ99, DN99d, DN99b, Far96e, Gre92e, Ish92, KLGP97h, KLGP97g, KLGP97o, Oda90a, Ogu90, Sch90h, Zuo99c]. p = 1 [CHT92]. p > 1 [Iwa92b]. p^{−1} ≥ R [Gre92d]. P^{2} [EP90]. P_{d}(X) [CDP90j]. p_{g} = 1 [MO93b, MO93c]. PG(t, q) [HH92]. π_{1} [CT92]. π: R × M [Kru97g]. PS(s, j; r) [IK99f]. PSL_{2}(1) [Kin94]. Q [Kim94, BF93i, Pra91, vdPS97j]. q = 3 [HH92]. q > 2 [Sho92u]. Q_{2,0} [Ura93f]. q_{s} [Xi94a]. R [Ani90, Kru97g, SS93, Sto92h, Yap96c]. r = 3 [IK99d]. r > (n − 2) [Ale91]. R^{3} [Bou91d, Bor92]. R^{N} [BL91b, Nil94e]. rotμ = γ [vW90]. Rp^{2} [BOWV91]. S [Abr96e, BCDS92, Kun92, Sch91g, vdPS97k, Met91q, Pra91]. S^{2n} [DM90a]. S^{4} [Huc90]. S^{n−1} [Lou91a]. S_{g,1} [LO98a, LO98e]. S_{g,d} [LO98b, LO98f]. σ [HFS94a]. σ(X, X^{⊙}) [vN92a]. Σ_{2} [GM90]. Σ_{∞} ∫ X [Dũn92]. Sl(2) [HN93k]. Sl(2, R) [Kni91]. SP [CG97e]. Sp(2) [DM90a]. Sp(3) [DM90a]. SU(2) [Bia92]. SU_{q}(2) [NM92]. ∑_{n≤N{nα+β}} [Sch90l]. T (b) [Dav90]. Θ [SV91b]. p̃ [KLGP97h].

P̃− [FP98a]. Q̃− [FP98a]. W̃^{1,2} [Cha91m]. T × C^{∗} [Xi94d]. t → ∞ [Dix97f, HS91]. U [Uri90, Gin97a]. U_{q}^{red}(sℓ_{2}) [FK93h]. V [DP91b]. V_{1} [Far92, Coh90, DM90a, Far96a]. φ [DNQ99]. W [LO98i]. G̃_{2}, B̃_{2}, Ã_{2} [Xi94c]. X [Pus96c, Pus96g, Pus96k, Sch94r]. X = G/H [Win95b, Win95c]. X[\sqrt{−1}] [Sch94r]. X^{α} [CV93]. X_{b} [Sch94c, Sch94b]. Y [DG90, KZ92d]. y′ [KZ92d]. y′′ [KZ92d]. y∂/∂x + (±x^{3} ± x^{6}y)∂/∂y [Rou90]. Z_{1,0} [Ura93f]. Z_{2} [Eyr98a, Eyr98h, Eyr98i]. Z_{p} [IR92]. |X| [Sta93].

* [LM90, Nie91j]. *-algebraic [LM90].

- [Dam91a, Pra91, Sło91]. -action [IR92]. -Adic [AS90a, And90, BBD90, Cri90, Hub95b, Hub95g, Jan90c, Oda90a, Ogu90, Sch90h, Wir98a, Zuo99c]. -adically [KLGP97h]. -algebra [Pis96j]. -algebras [BZ92, Ino98a, Ino98e, NZ99a]. -Analogue [Xi94a]. -analysis [Röc99]. -Arithmetic [Abr96e]. -balls [Ger90]. -body [LP90a, Mey99i, Kit93]. -Boundary [Cha91f, Cha91g, Cha91b]. -bundles [Kum97]. -categories [Mar90b, Sło91]. -cells [Lea90]. -closed [SM99a, SM99h]. -Cohomology [Zuc91]. -complex [Pus96c, Pus96g]. -complexes [HA90a, Huc90, Pus96k, Tro91]. -connected [Küh95a, Rut97n, Rut97o]. -constructible [SS93]. -Control [MP91]. -convex [Day92, MM90]. -cover [Sto92h]. -D [GDT91]. -designs [Met91a, Met91d, Met91l]. -Difference [vdPS97j]. -differentials [Lea90]. -dimensional [Fre94, Kai97a, KW94, Kim94, Kin94, Kna95e, Kna95h, NZ99f]. -divisors [SV91b]. -entropie [Thi91]. -equivalent [Rou90]. -equivariant [Bar93a]. -estimate [Cha91b]. -estimates [KO91, Räb91]. -extensions [Gre92e]. -fields [BK94]. -filtered [HFS94a]. -Fold [Met91q]. -forms [Paz91]. -functions [Li91, CP91g, Pan91h, CP91e, DP91b, Fuj90a, Har90c, Kat93a, iM90b, Sch90x]. -graded [Eyr98a, Eyr98h, Eyr98i, Kan90]. -graphs [BP97k]. -group [Bui92a, Lan96a]. -Groups [KLGP97g, SZ90, KLGP97h, KLGP97o, Bui92c, Bui92f, CP92]. -homomorphism [Muk90]. -Ideals [HWW93f, HWW93b, HWW93c, HWW93d, HWW93e, HWW93g]. -induction [GM90]. -invariant [BV93, Tam99c]. -jet [Rou90]. -local [Ani90]. -manifold [HM90]. -Manifolds [Joh95f, BP97c, CS92b, Joh95a, Joh95c, Kai97b, Kai97d, Kai97e, Kai97h, Küh95a, Küh95b, McC90c]. -map [MO93g]. -matrix [BCDS92]. -method [Cha91c]. -modules [Ber90a, Ber90c, SS93]. -multiplicity [Iwa92b]. -multipliers [Üst95a]. -Neumann [Cha91n]. -opération [HL92]. -operator [Wah95i]. -optimal [Kwa91, Pea91]. -partite [Yap96c]. -periodic [DM90a, Far96a, Coh90, Far92]. -pinched [Ham91]. -point [AE90]. -point/ [AE90]. -polynomials [Pra91]. -processes [Gin97a]. -product [Hub95a]. -projections [Wah95c, Wah95d]. -prospaces [HP92]. -Régularité [Bek91a]. -relation [Joh95c]. -resolvents [deL94c]. -rings [SM99u, SM99n]. -sets [Bou97i]. -Spaces [Jaj91h, Bar93g, FS92, Jaj90, Jaj91a, KV92, Møl90, Rut97h]. -spanned [Bal90b]. -spannedness [BS90a]. -sphere [Han90]. -spheres [BR90f]. -stage [Møl90]. -statistics [Gin97a]. -structures [Mar96f]. -summands [HWW93a]. -term [Lin90]. -th [Cas92]. -theorem [BL91a]. -théorie [CT93]. -theories [HKR92, Wür91]. -theory [Ina90a, Jan90d, LPLG^{+}92, Mur94, Far96a, Gil92, Gub90, IP92, Ina90b, Iwa92a, Jan90b, Jan90c, Kan90, Lod92, Pus96a, Sou92, Uri90]. -tightness [Küh95h]. -Topology [Pil94e, vN92a]. -topos [Sch94o]. -toposes [Sch94m, Sch94j]. -tori [Ish92]. -torsion [Far96e, Kin94]. -type [KW94, Kin94]. -types [HP92]. -unit [Sch91g]. -units [Kun92]. -Variation [DN99d, DNQ99, DN99b]. -variational [Eke90]. -widths [BT93].

1 [HOR91a]. 17ème [GC92]. 1987 [XD90]. 1988 [AvW90, Alb90, AYM90, BS90c, BLM^{+}90, CDZ91, FG90, GT90, Hen90a, HMRS90, KÜ90b, LW90b, Lat90, Mim90, NF90, Odi90, Pic90, SBL90, TBR91, Vil90]. 1988/89 [AYM90]. 1989 [ASMS90, AK90, BBD90, BDF91, BPSH91, CR91, Cel90, FR90c, FIK90, JOP91, JLP90, Kov90, LS90a, NO91, RSSV90]. 1990 [ACC92, ACE91, BCC92b, Ben91b, BM91, CF92, CPR91, FPSW91, Hen92, HLN91, HPSS92, KRW92, Kul92, MP91, Vuo92, YS92, dBT91]. 1991 [Bal93, CC92, CTKV93a, CT93, CMR92, FZ93, GS93, Gra94, Hen93, HMRS92, KZ93, Kom93a, ST92c]. 1992 [BGM94, BV93, CP93b, DM93b, ZD93]. 1993 [BA94, Ber95, FG96, Vis94]. 1994 [Ber96, CC96, Joh95g, Rug96]. 1995 [Ber98b, DL97, Nar97, TT96]. 1996 [GGSC97, HS99, Run97, TPZ98]. 1997 [Ber99a, Cap99, Qua98, Vio99]. 1998 [Da 99]. 1st [Cap99, Cel90, DM93b, FG96, FZ93, Odi90, Rug96, TT96, TPZ98, dBT91].

21-29 [BA94]. 25th [NO91]. 2nd [Bal93, BA94, CR91, Da 99, DL97, FG90, HRS91, HS99, Joh95g, MP91, Qua98, ZD93].

3rd [CF92, CC96, CP93b, Gra94, Nar97, Run97, Vio99, Vis94].

40ème [Mal91].

88 [Ina90a]. 89 [HT90a, OR91].

90 [Gre92b, LM91].

A. [DN99a, Tik90]. Abel [GV91c, GV91a, GV91b, GV91f, GV91h, GV91i, GV91j, GV91k, Rei90c]. Abelian [BCC92b, AH90, AR96b, BL92c, Bro90b, CG91, Dat90, Eva92d, FS90d, FPS92, Hin98, LO98o, Zuo99c, AH92d, AR96e, DRSŻ91a, DRSŻ91i, EE93, GH90, Hui93, Hul90, LO98h, Van96h, vG94]. abeliennes [Deb92]. above [LQ95c]. Abramov [BC92m]. absolute [Jan90e]. absolutely [Bun92]. Abstract [Cor91, Del91b, Kat93b, Mar96a, Mar96i, Ric99a, XL98d, Col92a, Dor93, Fla91, Mil94k, Rob90]. abundant [Pel92]. Acceleration [Bal94a]. accéléro [Eca90]. accéléro-sommation [Eca90]. Acknowledgement [AH92a, BFS98a, Pan91b]. actifs [Ans94]. acting [Abr96b, BMMN98d]. Action [MF94a, CP93c, IR92, Tur90a]. actions [CW91a, KB96b, KB96a, Sad91, Sch94g, Sch94l]. acyclic [Huc90]. adapted [DS95]. Adaptive [Joh98]. Addendum [Oba94a]. addition [AR96g, Mol93j]. additional [Par91b, Vät97d, Zim98e]. Additive [AB91, Ban91a, AS98h, Fuk93, Jaw91]. Adèlic [Sho92a]. Adiabatic [Bor98a]. Adic [AS90a, And90, BBD90, Cri90, Hub95b, Hub95g, Jan90c, Oda90a, Ogu90, Sch90h, Wir98a, Zuo99c]. adically [KLGP97h]. Adjoint [Bou97a, vN92c, vN92e, Fuj92, Pal90, vN92b, vN92d]. Adjoints [Neu97j]. Adjunction [Bou97b, BSW92, Dem96]. Admissible [CP91e]. Advanced [Qua98]. adverse [BR97]. aerosol [Oud91]. Affine [AM91, BCDS92, Bui92a, Cal90, Dek96b, Xi94b, Xi94l, Bou91e, Koz91, MM91a, Sch94d, Van96f, Xi94c, Xi94e, Xi94g, Xi94h]. affinely [MR91a]. affinely-rotation-invariant [MR91a]. after [Oor93b]. Alain [LW90b]. Aldous [Pra99]. aléatoires [Fer91, Le 97]. Algebra [BS90c, Her94g, Ina90a, Ser92q, VM92, Bab92, Bou97c, Bro90b, Dũn92, Erd90i, Jaj91f, KV92, Lin90, Mit94a, MZ90, Nie91a, Ouy91, Pir90, Pis96j, Rei90a, Ric99d, Sch92a, SG97]. Algebraic [ACC92, Bal93, Bar92, Bou98b, Bui92d, CC96, CMR92, Dek96a, Ehr92, GT90, Gro97a, Har90a, HPSS92, Jan90b, Jan90d, LPLG^{+}92, Mur94, Nis96a, Par92, Pus96b, Sch90j, SY91, Sou92, Spr93a, ST92b, ST92c, Van96a, Van96g, Voi94, Zie96a, Zie96b, Zuo99e, AK92, Aub92, BK91, BLR90, BIOS90, BK90, Bru92a, Bru92b, Bui92b, Bui92c, Bui92f, BCG90, Cen91, CTKV93a, Deg92, FR92, Gil92, GC90, GCR90, Gub90, IP92, JT91a, JR91b, Kaw92, Kun92, LM90, Lod92, McC90c, MO93h, Nab90, Nar90, Oko91, Pol92, Pom91, Pyt99f, RS90e, SM99r, SM99q, Shu92, Sil92, Spr93f, Tat96, Tia96, Tyu92, Voj93, ACC92, BA94, BDF91, JOP91, SBL90]. Algebraic-Geometric [Ehr92]. Algebraically [AFF90, Fre91i, DM96, Pil98b]. Algebras [Erd90a, Erd90d, Erd90b, Erd90c, Erd90e, HWW93f, HS94, IK99g, Ino98f, Koc91, Koe99f, Nie91c, Ric99e, Ser92f, Ser92g, Ser92l, Ser92k, Ser92n, Ser92p, Tam99b, Xi94l, AFSTS92, App90, BZ92, Bru90, CRV90, CK90, DH94, Erd90i, Erd90k, FK93b, FK93i, Gub90, GR92, HWW93e, HFS94a, HFS94d, HFS94g, Hel92, Hen90e, IK99a, IK99i, Ino98a, Ino98e, Iva93, KS92, Koe99a, Koe99d, Koe99e, Koe99h, McK90, Mit94c, Nie91b, NZ99a, OS90, Ol’92, Ric99a, Ric99c, Ric99f, Ric99g, Sid92, Sim90, Šip92, Sta92, Tam99g, Vas90, Xi94b, Xi94f, Xi94g, Xi94h, Xi94k, Zar91a]. Algèbre [SG97]. Algèbres [RS94, JS92]. algébrique [Bun90, CT93, Wal90a]. algébriques [CT93, Cos90a, Dur90c, MNM90]. Algebro [Pra91]. Algebro-geometric [Pra91]. algebroid [HM91]. Algorithm [CP92, KLGP97a, AH92c, CL99d, FS91, Sto92g]. Algorithms [KLGP97b, vdPS97a, Ben99, Cat99, Egg91, HO91, Vai93a]. Aligned [Mas95a]. alloys [HMS94]. Almost [Dek96b, Dek96f, Git98, Jaj90, Jaj91a, Mol93b, XL98a, Dek96a, HMN91a, Kar91, Mao91, RB92]. Almost-Bieberbach [Dek96b, Dek96f]. almost-crystallographic [Dek96a]. along [Les92]. Alternative [RG96a, BL92a, Pra99]. Amalgamation [JR92a]. amidst [BHS96i]. Ampère [Sch90q, Sch90t, Sch90u]. Ampèrean [Wer94]. Ample [dJ93, Fuj92, Pal90]. amplitudes [LR90]. analog [Hol90a, OK91]. analogue [Ogu90, Xi94a]. analogues [GE98d]. Analyse [Bon91, FM90, Lem90a]. analysing [Die99]. Analysis [AE90, BBD90, BGV90, BGV92, CR91, Cel90, CDZ91, ELL99b, FPSW91, GS93, GV91a, HN94a, HN94b, HT90a, JL93a, Kom93a, KÜ90b, Lem90c, LM91, Mil94f, Nua98, OR91, Sch95a, Ste90a, SW92, TPZ98, Üst95e, Vil90, Wei94b, Wir98b, YS92, Bar90b, Bel94, Bre94c, Chi90a, Chu93, FS91, Gil94, Gin98, Gli92, KP90, Kom91, Led96, Moo96a, NSS92, Nik94b, Oba94d, OSK93, Pro90a, Ran90a, Rob90, Roc90, Röc99, Sim98a, Sjö91, SOA91, mS91b, Tak91, Tal96, Tib90a, Woy98d]. Analytic [Bal94e, GT90, Per91, Ser92a, Ser92b, Ser92c, Tum98, AB90a, CP91b, Ebe98, Fro91, JL93c, Maĭ93a, Mon91a, Mor90c, Neu97k, Ore90, Pan91f, Pus96c, Sim98c, FIK90, NF90]. Analyticity [Ant92, Kur90, XL98b, Lie93g, OU91]. analytique [Dur90a, Dur90b]. analytiques [KPR90]. Anderson [Har93a]. Andersons [Har93a]. Ando [Pis96l]. André [McG92]. Andreev [MR90]. angle [BK92a, Krö90, Mat93]. angle-doubling [BK92a]. angles [Pri93]. annealing [Cat99]. anneau [Kro91]. Année [Mal91]. Annihilating [IK99e]. annotated [Len93]. Annulation [Bea92]. anomalous [BL90]. Anosov [NZ99b]. Ans [LW90b]. anticipatif [Gro90a]. anticipating [Nua98, Par90]. Anticipative [NV90, JS90]. anticipatives [BRS94, Pra94]. antisymplectic [Wan91a]. Anwendungen [Hla90]. any [LS90b, Wah95g]. apartment [Lan96b]. Appendice [Arn99]. Appendices [BHS96a, ELL99a]. Appendix [Bek91b, Mum99b]. appetizer [Wil95]. Application [Alb98a, Bal94e, CJ90, ELL99b, KM97a, Mey91a, OL90d, Tai91c, Tak91, CTR96, CDP90i, Eca90, Met91p, Ric99b, Yan90, ZR92, Pra92]. Applications [Abr96e, AvW90, Ber99b, BMMN98a, BGV90, BGV92, Bor98b, Bui92b, BR90j, CR91, Cha91c, DL96a, Dix97a, DL97, DT97a, DT97d, FP98b, GV91c, GV91a, GV91b, GE98a, HN93a, HN93f, Koe99f, LL91, Mie91c, Mil94f, Mit94b, Mol93c, MM91b, MN90, Nil94a, Nis96b, Oka90, Par90, RG96b, RS91a, Sna92, Tat96, Tij93a, Tot94a, Voj93, Wei94b, Xu96a, Yuk98a, AP94b, AP91, And91, Bar97, Bel99, BSW92, BKS93, Ber98a, Bul90, Bur91, CP91a, Che91, CS96, Cri90, Del91g, DGT91, Erd90h, ERHG91, FS91, Fri90a, GDT91, Gru91a, HH92, HOR91a, HFS94f, Hla90, Ino98c, Isa92g, KO93, Kos95d, Kos95f, KB96f, Lar90, MY99, MY07, Mal92, Mil91, Mil94d, Mil94g, Mol93h, Mot91, Mur91, Ole90, Pil99a, Pra91, Röc93, Sch92a, Sho92d, Sin92, Son97, Sou97b]. applications [Üst95i, Üst95k, Vät97f, Zak91, SvW92, Wat90]. applied [Gus90, SV91b]. Approach [AP94b, BEGG90c, Koc90h, Ric99e, Zie96a, Alp90, AS98d, Bas90, Bor90c, Bra98d, FS92, Fur90b, Gra96, Her90b, Kat93a, Kwa91, Pan91a, Tai91a, Zip91]. approximants [Maĭ93a]. Approximate [Tad98]. Approximating [Kit94f, Kit94c]. Approximation [Beu93, Bra98a, BVH94a, CG97a, CDP90a, EE93, GS93, HS93, KR92a, KR92e, Kit94a, Kit94b, Qua98, Tij93a, Tot94b, Tot94f, WY92, Wie92, BMO98, Ben99, Bra98b, Bra98d, Bra98f, CV93, Fis90, Fre90, Kar91, Lar90, LS93, LL91, Maĭ93a, Sch91d, She93, Sta93, Woy98f, Woy98i, Hla90]. Approximations [Bre94d, Bre94a, Dur90c, Sch91e, Arn95a, Hou90, Pyt99a, SM99a, Voj93, Yur95b]. April [HPSS92]. Aquila [SBL90]. arbitrage [Jou97]. Arbitrary [Bre94b, deL94a, Bia90a, Tan91, Tho90b, Wal91a]. arbres [CDP90a]. Archimedean [CP91b, CP91g, Pan91f, Pan91g, Pan91h, Pan91i]. area [Wer94]. areas [Shi94]. arf [Koc90b]. argument [dM99]. arguments [CW91b, DJS90, LL90]. arising [BFS98i, ELLV90, Ken94, Mor91, RS90d]. Arithmetic [Abr96e, CTKV93a, Edi93, Hun96b, Hun96e, LS90a, AR96a, AR96d, Ber90d, Har90b, Now90, RS90c, Roh90, Sch90x, Voj93, Bal93, Vio99]. Arithmetical [CP91g, CP91d]. Arities [SM99a]. arrangements [BFS98d, Mas95i, MT91b]. Array [DJS90]. Arthur [Sho92g]. Articulés [Hau91]. Artin [Fre94, KW94, Kim94]. Artinian [FZH90, Khu90, Xue92a, Xue92b]. Arveson [Mey93a]. Askey [NM92]. Aspects [LM91, MM91b, Mur94, TBR91, AD90, BK91, Dur90a, Dur90b, Ebe99f, Gin97b, Hod90, Khi92, Mar90d, Ore90, Ver94]. aspherical [FJ91]. Asplund [Phe93a, Phe93i, Phe93j]. assembly [Ros91]. Assisted [Koc90h, OSK93]. associated [Aub92, Bel90b, Bou97c, Bou97d, Coh92, ELL99e, Lun92, MR91b, Raj90, Sad91]. Association [Zie96a]. associative [GR92]. Associées [Par91a, HL92, JS92]. associés [Maz97]. assumptions [KM97g]. asymmetric [Sap90]. Asymptotic [Bal94b, BW94a, Bla91, Bre94c, Bre94d, HS91, Jaj91b, Kit93, Krö90, Mél96, Mor93a, Pus96e, Pus96d, Sim91a, BV90, BL91a, Gol91a, IK90, JN96, JL93b, Pus96g, Pus96f, Pus96a, Sch90l, Sch92b, Sim91g, Tsf92, Wat90, Woy98d, Yan96, ERHG91]. asymptotically [Wir98c]. Asymptotics [KM97b, VC90, Yuk98b, Ike90, Jen90, Rak93, Wan91b, Yur95a, Yur95e]. asymptotique [BW94b, Mic98]. asymptotiques [Arn93]. Athens [JLP90]. attached [Xi94m]. attention [Dow90]. attraction [GS90a]. attractivité [Mai94]. Attractors [PR97c, BV90, Bot92, LQ95f, Pil94d]. Aug [BS90c, Kom93a, Pic90]. Augmented [Far96b, Lan96g]. August [BLM^{+}90, Da 99, HMRS92, Mim90, NO91, YS92]. Auslander [Erd90d, Erd90g, Erd90j]. Austin [OR91]. Australian [Kov90]. auto [Le 97]. auto-évitantes [Le 97]. Automates [CDP90b]. automatic [deL94s]. Automorphic [LS90a, Sch90x, Sho92f]. Automorphism [BOERT96c, BEGG90c, BEGG90a, BEGG90b]. Automorphisms [Acc94a, Acc94b, Eva92a, HM91, Acc94h, BOERT96e, BEGG90d, BEGG90f, Gol90, McC90c, Rut97h]. autonomous [HMN91f]. Autour [BÉK^{+}98]. aux [Bon91, CJ90, JS92, Par91a, Pra92]. averages [Wir98a]. averaging [Wah95i]. Axioms [Bek91g, Bek91c, Hub95d]. Axler [Iva93]. Azéma [CF95, Sch91l]. Azumaya [Weh90, Xue92a].

B [Hit91, Kin94]. Back [BS97a, BHS96b, CTKV93b, Del92a, FR90a, Hru98a, KP96a, Liu96a, MF94b, Mey93d, Nus93a, Oor93a, Pan91c, Phe93b, Pis96a, Rau95a, Rub99a, Sce90a, Ser92d, Shu98a, Smi99a, Sou97a, Ste99a, Str93a, Val90a, Van96b, Var93a, XL98c]. Background [DS90f, Mil94b, Moe95a, Sho92j]. Backscattered [Fri91]. backward [EQ97, MY99, MY07]. Bade [Ric99b]. Baire [HOR91a, Pia90]. Baire-1 [HOR91a]. balanced [PsT90]. balancing [FK93a]. ball [Hol90b]. balls [BL91b, Ger90, SS91]. Banach [AMM90, CG97b, CG97h, CG97g, CM97c, CNZ91, DH94, Day92, DGT91, GJ91, HW91, HWW93b, HWW93a, HWW93e, HWW93f, HOR91a, Kan90, Kis94, KB96g, Leu91, LT96, OL90b, OL90d, OL90f, Phe93d, Pis96d, Räb91, Ric99h, Sch90f, Sch90h, Sim98k, Sza91, WY92, Woj91, deL94i]. banachiques [Sch94u]. band [GKW91, Kai97d]. band-operations [Kai97d]. Barcelona [ACC92, CC92]. Bari [CF92]. barotropic [ST92a]. Barratt [Bau91]. barrelled [Nar93]. Barsotti [Cri90]. Barycentre [ESH99, EM91]. Barycentres [Arn95a]. Base [Sch94a, Chr91, Hen90b, Her94k, JT91a]. Based [Sho92g, DX93, DK93a, NV90, Pyt99f, Xi94c]. bases [Her94b, JJS92, JR92b, KS90b, Nik94a, RS90b, SZ90]. Basic [AS98a, GV91d, HWW93c, Her94a, Hub95c, IK99d, JL93a, Kön98a, KM97c, Kru97a, Lor92a, Met91a, Sch93a, Sto92a, Vät97a, Yap96a, Zie96c, Acc94d, AR92, DS90f, Ebe99e, Küh95d, KJ90, Mas95c, Met91b, Mie91k, Sch94n, SSS92, Wil97e, Yap96g]. basin [GS90a]. basis [Jaw91, SDS90]. basse [Mic98]. Baxter [Ge92, GR92, Mor90b]. Bayreuth [HPSS92]. beam [Gra91a]. beams [Ani91]. Bébé [Mey91a]. Behavior [Dix97d, Har98a, IS90, Sim91a, Alb98e, BM90d, Mor93a, Sch90p, Tan91]. behaviors [HS91]. Behaviour [DH90, BV90, Gol91a, Mél96, Sch92b, van90]. behaviours [Shi94]. behind [KLGP97n]. Beilinson [Hod90]. Beispiele [Har93b]. Belavkin [Mey93c]. below [Ars90, LQ95d]. Bénard [GP91a, Nag90a]. Bendixson [NZ99j]. Berechnung [Hla90]. Berechungen [Hla90]. Berlin [FPSW91]. Bernoulli [DS92]. Bernstein [BL93a, CS92a, Fre90, Hod90]. Besov [LP91]. Bessel [AY90, Ber90b, DWY98, Shi96, Xue90]. best [CV93, Heb96a, LS93, Maĭ93a]. Betti [Kai97a, Kai97e, And90, Göt94b, Kai97d]. Betti-trivial [Kai97a, Kai97e, Kai97d]. between [DS90e, Ebe98, FF97, GV91j, MS90a, McK90, Räb91, Riv90, SOA91, Woj91, Wri90]. betweenness [BMMN98m]. Beĭlinson [Har93c]. Bi [Lie93a]. Bi-symplectic [Lie93a]. Bialgebras [Sch93g]. Biane [Leu96]. Bias [HALM90, Oll93]. Bibliographies [DNQ99]. Bibliography [Arb99, Len93]. bicommutant [Ric99g]. bidimensional [PR97e]. biduals [HWW93b, HWW93a]. Bieberbach [Dek96b, Dek96f, Pri93]. bifurcating [GS90b, Sch90a]. Bifurcation [Ast91, Bar93a, Chi91, Per90, Rei91d, RS90d, Ama91, Ben91c, BF93a, BF93b, BF93e, BF93i, BF93k, CDD90, Chi90b, CvdM90, DW91, DR90, ERHG91, Fit93, Fru91, LP90c, Oka90, Sap90, ZR92]. Bifurcations [Ben91b, DRSŻ91b, DRSŻ91i, GHS97a, RS91a, Bae91, Bor90b, Dam91b, FS91, Fur91, Lau91, Lob91, Pet91, FR90c]. Bilinear [Mil94c, Sna92]. bin [Pro90a]. bin-packing [Pro90a]. Binaires [Rod92]. binaries [MS90b]. binary [Bax92, Sch90v]. binding [PR97b, PR97a]. biological [Had99]. Biology [Cap99, RS90d]. biomagnetic [Lou91b]. biorthogonal [RS93]. bipolar [BS99]. birationnelle [Sil90]. Birch [Rub99b]. Birkhoff [Lor92a, Lor92e, Lor92f]. biserial [Erd90i]. Bismut [Léa96]. bistability [DW91]. Blaschke [Par93]. Bloch [Bab92, Hub95d, Min90]. block [Cen91]. Blocking [GE98c, GE98b]. Blocks [Erd90e, Erd90f, Erd90j, Rou98, Rut97c]. Blow [MS90b, OT90, MO93a]. Blow-up [OT90, MO93a]. BMO [DK94, Gra99, JcC91, Kaz94a, Kaz94b, Kaz94d, Li91, Shi91a]. BMO-martingales [Kaz94a]. BO [I90]. Bochner [Ban91b]. bodies [Ika90]. Body [Mey99i, Kit93, LP90a, MS90b, Sap90]. Boltzmann [BL91a, DGDMN91, Hen91, Lan91b, Mél96, SOA91]. Book [HN94a, HN94b, Mum99a, Mum99g]. Boolean [HFS94a, HS94, HFS94d, Ric99a, Ric99c, Ric99d, Ric99f]. bootstrap [Gin97b]. Bord [CDP90c, Arn93, CDP90m]. Bordered [BEGG90c]. bordism [Kai97b, Kai97c, Ver91]. Borel [Bal94f, Del91d, Gre90]. Borsuk [IR92]. Bose [Nie91c, Nie91a, Nie91b, Sch93e, Sch93h]. Boson [AL90a, Fag90, FS90b]. bound [Lus90, Met91m, Tam90, XL98f]. boundaries [Kan91, Law91a]. Boundary [Alb98g, Cha91f, CGG^{+}91, Kaz90, Pom98a, RS90c, Sch90d, Sch95b, Sch95d, Tai91d, deL94b, Alb98e, Ant92, Bax92, BP97g, BBG91, BEGG90a, BEGG90b, BM92d, Cha91g, Cha91i, Cha91b, CMN95, CP93c, CP93d, CP93e, CP93k, Fis90, Fur90a, Gal92, Gru91a, Küh95c, Lie93h, Lio93a, Mar93b, Maw93, Mog92, Neu97f, Neu97g, OT90, Pil90, Riv90, Soc92, ST92a, Son90, Tai91a, Tai91c, Tib90b, Ver94, Wah95g, Zvy90a, vW90]. Boundary-Domain [Pom98a]. boundary-layer [Fis90]. Boundary-value [Kaz90, Gal92]. Bounded [Gol90, Pis96k, Ban92, Bra98c, Bra98g, Cha91d, Gil90, HWW93g, Kan91, Kar97a, Kos95c, Pis96b, Pis96c, Pis96d, Pis96h, Räb91, Woj91, deL94a, deL94i, deL94j, deL94n]. Boundedly [JJS92]. Boundedness [Mit94a, Bas90, Han91]. Bounds [Sim91b, BLR90, Del92e, Fab93b, GKKL91, Met91k, Yap96h]. Bourgain [SW93, Sch91j]. Boussinesq [SvW92]. Boussinesq-equations [SvW92]. brackets [Koc90i, VK90]. Braid [MT91b, DLS90, FK93d, FK93e]. braided [Maj92]. Branched [BP97a, BP97b, BP97c, Acc94e]. Branches [Sch90d, Sch90a]. Branching [Git98, War97, JN96]. Branges [YSA92]. Bras [Hau91]. Brauer [Heg92]. Brazil [BS90c]. breaking [Nag90a]. Bressanone [Run97]. bridge [Pit99]. brief [Wal91b]. Brill [FP98c]. Bros [Del92c]. Brown [TY90]. Brownian [BW94a, Bro93, BM92d, Bur93, Don98, DÉY93, ÉS99a, KT93, Kho97, Le 92, Le 94, Mou92a, Neu96a, Neu96d, Pit99, Shi94, War97, War99, Wer94, Wer95, Zhe96]. brownien [Arn99, Ber93, Mey94, dM99]. brownienne [BW94b]. Browniennes [BÉK^{+}98, ABÉH95]. browniens [ÉY98, JY90]. Bruhat [Lan96e]. BU [Koz90a]. Bucy [Bou91a]. Building [Lan96e, Rut97c, Lan96c, Lan96d]. Buildings [Abr96e, Abr96b]. bundle [App95, Hub95e, Pal90]. Bundles [Brî96d, Nar97, dJ93, AR96b, AR96d, Bel90b, BEL91, BIOS90, BK90, BS92, Brî96a, Brî96b, Brî96e, Brî96f, CT92, Dam92, Dem96, DR91, DM96, Fuj91, Fuj92, FP98e, Göt94a, Isk91, JP91b, KO90b, Kob90, Küh95b, Kum97, MO93b, Nog91, PSW92, Sai91a, TO97g, Tyu92, Zuo99a]. Burgers [Woy98a, Woy98c, Woy98b, Woy98g]. Burkhardt [Hun96c]. Burkholder [Mon94]. Burnside [Wal90b].

C [Ste90b]. C.I.M.E [Bal93, BA94, CF92, Cap99, CC96, CR91, Cel90, CP93b, CT93, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. cône [KPR90]. càdlàg [Fer91, PE92]. Calcul [Bia95a, Pic91, Gro90a]. calculation [Hla90]. Calculations [Gre92a, Mil94d, Don98, Rut97d, Rut97o]. Calculi [deL94h, Gra91b]. Calculus [Eyr98g, Mey93m, Oba94e, AS98d, BK92b, BOERT96d, Bel90a, BP97h, DZ93j, Eyr98f, Eyr98h, Eyr98i, Fag90, HS90b, JS90, Kom93b, Mey95h, Nor92, NV90, Nua98, Par90, Raj96, Vät97c, deL94j, HS99, Mar90a]. Calderón [Han91]. calibrations [Hon90]. California [BM91]. Cameron [LN97]. Canada [Pic90]. canard [CD91]. Canberra [Kov90]. cancellation [BL92d, Gol90]. Canonical [CP91a, Dek96c, Col92b, DPlPS90, KT91, Rei90b, Wil97a, vdPS97b]. canonique [ESH99]. capacities [BVH94a]. Capillarity [Mie91a]. Caractère [EP90, dM99]. caractérisation [Pra92]. caractéristiques [MY91]. carcinogenesis [KY93]. Cardinals [Bek91g]. Carleman [DP91j, Sid92, Sin92]. Carlen [Hu93]. Carleson [Li91]. Cartan [FZH90]. cas [Eme91, Ser94b]. cascade [Bae91]. Case [Bus92, Oor93b, Ale91, Bae91, BGR91, Col92c, Fru91, Fur92, GN97, GU92, IK99d, Kar97c, Kar97e, Kit93, KP96c, Lan96d, Lan96a, Mil94c, Neu97b, Neu97c, Neu97d, Neu97o, Pet92, Sch94c, Sch94b, Ser94b, Wan90, Wol92, Yuk98b, Yuk98l]. cases [Dix97e, Küh95e]. Castelnuovo [Acc94c]. Catalans [CC92]. catalecticant [IK99c, IK99f, IK99j]. Categorical [HP92]. Categories [FK93g, Hub95q, AR91, BFS98i, Bjö91, Bou91e, Dat90, Fre91i, FK93a, FK93f, Hub95i, Hub95j, Hub95k, Jib90, Kan90, KV91, Mar90b, Pir90, Ric98, Sło91, Uri90]. Category [Bar93c, Bar93b, Bau91, BL94a, BFS98g, Bou97d, CW91a, Law91b, Pia90, Pus96f, Rou98, CPR91]. Cauchy [Bal94c, Del92h, Kaj91, KN91, Nis91, Wie90, XL98d, deL94d]. causal [Kan91]. Caustics [Del92b, Mon91b, JR91a]. Cayley [DV91]. celestial [Mey99d]. cell [Xi94e, Xi94i]. cells [Lea90, Sim91m, Xi94c]. Cellular [Far96c, Far96d, Rut97e, Rut97d, Far96e]. Cellular/homology [Rut97e]. Center [Mie91c, Box91, DvG91, Mie91d, Mie91k]. centered [Kun92]. centers [Chi90b, Shp90]. Central [Mey99a, DRSŻ91c, GVV90, Shp90, Wac92]. centre [LP90c]. Centres [Bou97e, Erd90f]. Centro [Bal93, BA94, CF92, Cap99, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. centroid [MR91a]. Certain [MO93b, Aki90, BSÉ99, BS92, EY93, FGL91, HOR91a, Law91a, Lei91, Met91l, Pan91a, Räb91, Sch99, Son90, Tak91, Tho90a, Wat99, Zip91]. Certaines [Rod92, Dix91, JS92]. certains [ÉY98, Mor90b]. Cetraro [CC96, Da 99, HS99, Nar97, Qua98, Vio99]. CG [BCMS90, van90]. CG-S [van90]. chaînes [Mic97]. chain [Fur92, Jac90, Rut97i, Wir98c]. Chainable [GC90]. Chains [Har98e, Cat99, Har98a, Har98b, Har98d, SC97, Sim91i]. Champernowne [DK92]. champs [Mou90]. change [AS98h, Her94k, JT91a, Sch94a, ÜZ97]. change-of-base [JT91a]. changement [Hen90b]. changes [ÉS99a]. changing [Żoł90]. channels [Mor93b]. Chaos [BHS96i, Fra95, Gra94, Rei91d, Ble91, Jou98, Led90, Sin91, Szn91, Woy98e]. Chaotic [BC92b, Eyr98b]. Chaoticity [Der95]. chaotique [Eme91]. Chapter [Wah95a, Wah95f]. Character [Pot95d, BV93]. characteristic [FP98d, Hru98b, Kai97d, LS90b, Mar91, Sri90, WJ90, vdPS97e]. characteristics [Bor90c, BS90b, Col92g, JL91]. Characterization [ITO93b, Li91, LQ95a, MR91a, Pus93, Üst95b, HH92, Met91o, OL90b, PsT90, Spr90].

Characterizations [DX93, Kit94c, Dek96a, DK93a, JcC91, Mas95j]. characterized [HSS91]. characterizing [Muk90, Pri90]. characters [HKR92]. Chebyshev [Fre90]. Cheponkus [Mar90c]. Chern [BV93, Hub95e, Hub95n, PSW92, Zuo99f]. Chevalley [KLGP97c]. Chicago [BDF91, Koc90a]. Chile [RSSV90]. China [CDZ91]. Chorin [Ran92]. Chow [Göt94a, RL90, Vis92a]. Chowla [Ogu90]. chromatic [Yap96h, Yap96i]. CIME [BV93]. Cinquante [LW90b]. circle [BK92a, BC92e, FW91, Ger93]. Claremont [BM91]. Class [Lan96h, Bot92, DK92, FGL91, Fuh91, Gus90, HM90, HH90, Hub95e, JL91, KPV90, KÜ90a, LO98m, LP90c, Oda90a, Par91c, Spr93b, cSyW91]. class- [Oda90a]. classe [Val94]. Classes [Ene95a, HL92, Bar92, FP98c, HOR91a, Hub95n, Jan90b, Kom91, MO93g, PSW92, Räb91, Res92, Rud92]. Classical [CG97b, Gra94, Gub90, LT96, RS93, Spr93c, Yuk98g, Abr96a, Ble91, Bor98a, Col92i, Daf96, EGI94, FP98e, FP98a, KLGP97j, Lun92, Ol’92, Pot95b, Wil97f]. Classification [Alb98b, AS98b, BMMN98b, BF93a, BF93b, Brî96a, CDP90d, Dek96d, Far96e, JR91a, Maz97, MO93c, Pet91, Sil90, Win95a, Yap96b, vdPS97b, CS92b, Pol92, Win95b, Win95c, Xi94h, BCC92b]. classifier [Koc91]. Classifying [Hub95f, Ish92, Moe95b, Moe95c, Moe95d, HKR92, JT91b, Pri90]. Clifford [Mit94a, Mit94c, Mit94e, Mit94d]. Climb [Per99]. Close [Con94, Shk93]. Closed [BOERT96a, Éme97, Mol93g, Ban92, BP97d, BP97h, CMM91, Del98, McC90a, Mol93b, Mol93d, Mol93f, Paz91, Pil98b, Ric99c, Ric99d, SM99a, SM99h, SM99k, SM99s, Woo98]. Closedness [GK98]. closure [Fur92, SM99l]. Closures [IK99a]. clustering [Kit93]. CM [HH90]. CM-fields [HH90]. co [Law91a, Ura93a]. co-Heyting [Law91a]. co-root [Ura93a]. coadjointes [Don90]. coalescences [Lor92h]. coarse [GVV90]. Coaugmented [Far96f]. Cobordism [Mor90a]. cocompact [HN93h]. cocycles [Kni91]. code [Neu97l]. codegree [KO90b]. Codes [Ehr92, Pot95a, Aub92, GL92, HH92, Han92, KTV92, Pel92, Sko92, Sto92o, Vos92, YK92]. codimension [Cos90a, DW91, DRSŻ91a, DRSŻ91b, OL90e]. codimension-four [DW91]. coding [OK91, ST92b, ST92c]. Coefficients [Jib90, Bor90a, DS90g, Fre91h, Ger93, Hla90, JPA97, KM97a, KM97c, Wei94a, MNM90]. coercive [BM92b]. Coexistence [PR97c]. cofibrations [Far96g]. coherent [Nie91d]. Cohn [HA90b]. Cohomological [Lan96i, Sch94c, Sch94b]. Cohomologie [Ber90a, Léa96, Ser94b, Ser94c, Ser94d, Ser94a]. cohomologies [LZ92, Mor90c, VK90]. Cohomology [ACC92, Dek96e, Har93d, Jan90a, Lan96p, Mur94, Pir90, Pus96d, Sch94d, Sch94p, Sch90x, TY90, Zuc91, AH90, Ade92, Bar92, BP97f, BL94b, BL94c, BFS98d, CG91, CP92, Dũn92, Gek90, Har90b, Har93e, Hub95b, Hub95h, Hub95t, Jan90c, Jaw91, Jib90, Lin90, Lom92, Pus96e, Pus96h, Pus96a, Ser94b, Ser94c, Ser94d, Sna92, Tat96, ZL90, LS90a]. Coifman [Dav90]. Coin [BSÉ99, BM90a]. Coin-Tossing [BSÉ99]. coincidence [CMZ91, Col90b]. Cole [Woy98d]. Coleraine [BLM^{+}90]. colimits [Far96h, Jay91, Sło91]. collapses [Bog90]. collapsing [MS90b]. Collatz [Wir98b]. collection [Vuo92]. Collège [SG97, SG97]. collineation [Pot95f]. Collision [BF93c, BF93d]. collocation [Vai93f]. Colloque [Alb90]. Colourings [Yap96j]. Comb [Sto92b]. Combed [BP97d]. Combinatorial [BEGG90c, Lat90, Ver90]. Combinatorics [Joh95f]. combings [BP97h]. Comet [Mey99b]. Commutant [Foi91]. commutatif [Bia95a, Ser94b]. Commutation [Far96g]. Commutative [Bui92c, Gre92c, Mey93j, Bel94, Gre92g, Man97d, Mey95e, NSS92, Ser94b, Ser94d, Xu96a, Xu96e, BS90c, JLP90]. commutativity [Dug92]. commutators [Mil94e]. Commuting [Ric99e, deL94j]. Como [CPR91, MP91]. compacité [Fer91]. Compact [Brî96d, BEGG90c, Ham91, Tod97a, Tod97b, Blü90, Brî96c, BEGG90a, BEGG90b, BEGG90d, Cas92, DG90, Ish92, Kna95e, Kna95f, LN97, Lem90a, NZ99i]. compacte [Arn99]. compactes [Aud90, DM90b]. Compactification [Lan96e, Lan96c, Soa94f]. compactifications [Zuc91]. compactly [Tod97a]. compactness [Sar92]. compactum [Sza91]. Comparaison [CTR96]. Comparing [Gil92, SM99b]. Comparison [Hub95g, Moe95e, Neu97a, Sch94e, Mor93b]. competition [Smi99b]. complement [Huc90, Kul91, Ros90]. Complementarity [Isa92k, Egg91, Isa92d, Isa92f, Isa92h, Isa92i, Isa92j]. complementation [Sch90h]. Complements [Sim91c, Sch94h]. complet [Fra95]. Complete [Man97a, Man97g, Ser92e, Yap96c, App95, BK91, DL96b, Fre91i, JJS92, Kus91, Nor92, Pel91, Mul90c]. Completely [DTW90, Pis96b, Pis96c, Pis96d, Pis96k, DM96, Van96a]. complètement [DM90b]. completeness [AP90, AM91, Kit93]. Complex [AP94a, Brî96d, GS93, HN94a, HN94b, Koe99a, MO93c, Nie91c, NO91, OU91, TPZ98, Win95a, Bar90b, BKS93, Brî96c, Brî96f, BCG90, Coa99, Dan91, Ebe98, Gin98, JL93c, Kaj91, Kap91, Kna95g, Koe99h, Kom91, MT91b, Mor90a, Mor91, Nie91e, Ohs91, Pus96c, Pus96g, Rub99b, Siu91, Tak91, Win95b, Win95c, HPSS92, Vil90]. Complex-analyticity [OU91]. complex-valued [Kaj91]. complexes [Col90b, Gra90b, HP92, HA90a, Huc90, Pus96b, Pus96k, Rut90, Rut97e, Rut97d, Rut97i, Tro91, Yam90, Zim98c, Zim98e]. Complexity [Gro97b, RS90e, Kel91a, MS90a, Ver90]. Complicated [BHS96c]. component [BEGG90a]. components [BLR90, Chr91, Erd90j, IK99b, Mar96f, Sch91l, Sch93d, Tyu92, Ura93d]. Comportement [BW94b, DW97, Mic98]. Composite [Vai93a, Zhe96]. Composition [Bou97f, DN99a, JN96, Lam92, LS91, Sar92]. compressible [ST92a, Val90c]. Compression [HN93b]. computability [Coo90]. computable [MS90a]. Computation [Göt94b, Rei91d, CP92, Fuh91, Lus90, RS90e, Sma92]. Computational [Sto92c, Mus92, Ore90, Wah95b, RSSV90]. Computations [FJ91, BP97f, Coo90, Koc90f, Soa94g]. computed [De 91]. Computer [Koc90h, Odi90, Neu97l, OSK93, Rei90a]. Computer-Assisted [Koc90h, OSK93]. computers [Pla90]. Computing [CW91a, Cou93, Foi91, Kem91b]. concave [MR91a]. Concentration [Led99, How90, Tal91, Yuk98e]. Concept [Ura93a]. concepts [Acc94d, AS98a, DT97c, Haz93, Sch93a]. concerning [Coh90, MM97, Pra99, Pro90a, Rus90, dF90]. Concise [Lun92]. Concluding [GHS97b, Yap96d]. Conclusion [Lan91a, Sho92k]. Conclusions [BHS96d, DRSŻ91d]. Concretely [Mac91]. condensation [SOA91]. condensed [SOA91]. Condition [MS91a, Gre92d, HSS91, MM97, Mas95l, Mog92, OT90, Wol92]. Conditional [EY93, Jaj91c, Pus93, Tay92, Wes93]. conditionally [Sch93d]. conditioning [BD94, NÜZ90]. conditionnellement [Mey93c]. conditionnelles [BM92c]. Conditions [Kit94f, LP90c, Alb98b, CGG^{+}91, GK90, Gru92, HW91, Neu97f, Neu97g, Pel91, Ura93f, Zvy90a, vW90]. conducting [Wol92]. cone [Gra91a, KR92h]. Cones [HN93e, KR92e, Isa92f, KR92c, KR92d]. Conference [ACC92, ASMS90, ACE91, AK90, BBD90, BCC92b, Ben91b, BLM^{+}90, BM91, CPR91, CMR92, FPSW91, FIK90, GT90, HLN91, HMRS90, HMRS92, HPSS92, JOP91, JLP90, Kom93a, KRW92, LS90a, LW90b, Mim90, NO91, Pic90, RSSV90, SBL90, XD90, YS92]. Configuration [BFS98b, DL96a, BFS98i]. configurations [Bor92, Mey99a]. Conformal [AVV92, Hou90, Mat90b, Sch90m, AG90, BBG91, PS90]. conformally [FGL91]. conic [Dul91, Isk91]. Conical [Lie93c, Lie93b]. conics [MT91b]. Conjecture [Fre94, Met91c, Oor93b, Top93, vG94, Apt93, Blo92, Bou98b, Bou98c, Clo90, DLS90, Hin98, Hom90, KW94, Kim94, Lin90, Lup90, Pil98a, Rub99b, Sai91b]. conjectures [AMS90, Bar90b, Bou91c, Har93c, LW90a]. conjugacy [BHS96e, Hig90]. Conjugate [AK90, Gus90, WS90]. Conley [Bar93f, Mis95]. Connected [Küh95c, BK92a, BLR90, Küh95a, Mar96f, Neu96e, Rut97n, Rut97o, Yam90]. Connectedness [IK99b]. connection [Gek90, KT91]. connections [ELL99c, GS90b, PR97i]. connective [Koz90a]. connectivity [Sto92n]. Connexions [Arn92]. conséquence [Hen90b]. consequences [Kuk93c]. conservation [Daf96, Joh98, Jou98, Neu97r, Shu98b, Tad98]. considerations [Wah95g]. Constancy [Mas95b]. constant [AP94c, CS92c, KM97c, KM97e, Kus91, MM91a, Mat93, RS90a, Vai93f]. constantes [Fon98]. constants [Heb96a, Min90, Vel90]. Constitutive [Alb98c, Alb98g, Alb98d, Alb98j, Col92c]. Constrained [Elf91]. constraining [Bor98d]. Constraints [Pyt99e, BL90, CGG^{+}91, Cvi97, Sto92n]. construct [Box91]. Constructible [Sch94f, Mar96d, SS93]. Constructing [AK92, SM99c]. Construction [AS98c, Bou97g, DL96b, ELL99c, Kna95i, Sim91d, Wil97e, BMMN98f, FS90a, GM90, Har93e, HK92, Kel98, Kin94, Mic93, Pat90, RG96a, Wil97a]. Constructions [Mit94e, Bou97a, CG97f, Jay91, OU91, Sch94l, Tuc90]. contact [Krö90, Soc92]. contacts [AMS90]. containing [ABW92, GJ91]. content [Pil98a, Sab91]. contents [Pis96g]. context [LR90, Neu97i]. contexts [Ebe99e, LS90c]. contingents [Ans94]. continu [Arn93].

Continuation [Mar93b, SS92, BHS96f, CMZ91, CPS90]. continued [BC92h, Jac90, Thr90]. continues [Arn95a, AW97, Fer91]. Continuity [HM92, Ban92, BK92b, SSS92, Sch90o, XL98e]. Continuous [AS98e, AS98d, Jam90, Kaz94b, Neu97b, Neu97c, vW96, AS98c, AS98g, CW91b, DÉY93, DN99a, Éme97, Ene95c, KR92h, RY90, Sch90f, Tak99, deL94a, deL94j, deL94n, vH90]. contractible [Huc90]. Contractions [CGST92, Jaj90, Jaj91d]. Contractive [PR97d]. contractivity [Gro93]. contre [De 96, Mic98, Pra94]. contre-exemple [De 96, Mic98]. contre-exemples [Pra94]. contribution [Col90b]. contributions [RS90c]. Control [MP91, Pyt99e, Tib90c, Bar97, DZ93d, DZ93e, DZ93i, Fra91, Pea91, Pyt99f]. Controlled [Son97]. controllers [Foi91]. convection [Coc98, Mor93a, Nag90a]. convection-dominated [Coc98]. conventions [Bui92g, Dek96h]. Convergence [Clu92, CJ90, Fer91, Har98b, Jaj91c, Jaj91d, Jaj91e, BP90b, BRS94, Jaj90, Jaj91a, Kar91, KP96b, KP96c, LS93, MR94, Mol93b, PG90, Pra92, Pra94, Pra99, Pri93, Sho92e, Thr90, Vai93e, Xia92, Yan93, Yee90, Yuk98h, van90, CJ90]. convergences [DZ93h, Fuk93]. convering [BL91b]. Convex [MM91a, Phe93c, Phe93d, Ban92, Bou91b, CMM91, Day92, Ika90, JJS92, KR92c, Kit94b, Mar90a, Mat93, MM90, Mou92a, Phe93h, Sch90r, Sim98a, XL98i]. convexes [Arn95a, CDP90k]. convexity [CS92a, MT90]. Convolution [MT91a, gDH91, LP90b]. convolutions [Kal93, Pan91g]. Cooke [BM91]. coordinates [CS92c]. Copies [CM97d]. corank [Mar91]. Corestriction [Gre92b]. corps [Dum91]. correct [deL94n]. correcting [HH92]. Corrections [BRS94]. corrélés [ÉY98]. correspondences [Hub95s]. Corrigenda [Bur90]. Cosenza [Nar97]. coset [Fin90]. cotangent [Gir90]. cotorsion [Xu96c]. Coulombian [LP90a]. counter [MM97]. counter-example [MM97]. counterexamples [LW90a]. counting [KP90, Wir98a]. couplage [Mai94]. couples [Her94e]. coupling [AFL90, Bor98a, Wan91b]. courbure [ESH99]. cours [SG97, CT93]. courses [SG97]. cover [Sto92h]. covering [Mil90a, Sto92o, Tuc90]. Coverings [Che90a, Che90b, Acc94e, Erd90d]. Covers [Xu96d, DM96, Xu96b, Xu96c, Xu96e]. CR [KR90, Tum98]. Cramér [Gao97]. critère [Bak91, Sch90i]. Criteria [Amg98, LS90b, OL90c, Sch90o]. criterion [Pro90b]. Critical [Che91, PR97e, Bar93c, Bor90b, iM90b, Mou92a]. crochet [Gro90a]. crochets [Rog90]. croissants [AJK^{+}96]. Cross [Ste90b]. crossed [HP92]. Crossing [Bre94e]. crossings [Raj90]. crystal [Kar97a]. crystallographic [CK91, Dek96a]. crystals [Bré93]. CS [Smi90]. CS-modules [Smi90]. cubature [Tem93, Vai93a]. Cube [Mic93]. Cubic [Pro98a, BB90, Hun96a, LP90c, Pol93a]. cubics [MD90]. Cup [Lan96j]. Current [Ene95g, Yau92]. Currents [Soa94a, BL92a]. curvature [AP94c, Eva97, GS90b, Haj91, Ham91, Kus91, MM91a, MORT91, Pet96, Ros91, Ste99b]. curve [CMM91, Sch91k]. curved [JcC91]. Curves [Dav91d, Mum99b, Mum99c, Nar97, STV92, Vio99, vdG93, BEL91, BCG90, Cal90, Cas91, Coa99, Col90a, Dav91b, GCR90, Gre99, Joh90, Kin94, KMM92, Mat93, MT91a, MT91b, Pel92, Pol92, Pre96, RS90e, Rub99b, Shu92, Van96e]. cusp [Wan94, Wol91]. cuspidal [MD90, MT91b]. cutting [Sto92g]. CW [Rut90]. Cwikel [Mil94c]. cycle [Sho92q, Sim91j]. Cycles [BA94, CT93, Mas95h, Mur94, Voi94, Blo92, Chi90b, CA91, Eca90, GH90, Jan90b, Jan90e, Kap91, LP90c, Mas95j, Mon91a, Per90, Sie91, Son90]. cycles-limites [Eca90]. Cyclic [BOERT96b, BOERT96c, BEGG90d, Gre92e, Gre92c, Gre92d, Pus96d, Pus96h, Chr91, Gre92h, HA90a, Lod92, Pis96j, Pus96e, Rou98, Ste91]. Cyclicité [Mou90]. Cyclotomic [Eva92b, Gre92f]. cylindrical [Mie91b, RS90e].

D [Sch91j, BL92a, Dub96, Ger92, Gra91a, GDT91, HYL91a, ZD93]. D-modules [ZD93]. D. [Top93]. d’Algèbre [Mal91]. Damek [BTV95a, BTV95c]. d’Analyse [Lem90c]. dans [Arn92, Arn93, Arn95a, Arn99, CDP90l, Dix91, EM91, Fou90, Gir90, Mai94, Mul90c, Pra92]. d’approximations [Bun90]. d’Après [BSÉ99, Dav90, Hu93, Mey93c, Mey93b, Mey93a, Sch99]. Darbouxian [GS90b]. dare [WS90]. d’arrêt [AJKY93]. Data [Cha91f, Eva92c, WS90, Cha91b, Fra91, Göt94e, Lio93a]. d’atteinte [DW97]. Davis [Mon94]. Dawson [GRCW90]. Decay [Wie90, IG90, SV90]. December [BCC92b]. decidability [FR92]. Decoding [Ehr92, KTV92, Pel92]. decomposing [Pol92]. Decomposition [ELL99d, Lor92b, NZ99c, SvW92, Sch95d, Sto92d, Zie96d, AS93a, AS98g, Blo92, Erd90f, Hol90a, Mas95f, MT99, PS90, RS90e, Sch95c, SY98, VM92, Yan90, CS96]. Decompositions [Mil94f, Maa91, Mil94i]. Decoupling [Gin97a, Sob90]. defect [Ric92, Rou98]. deficiency [Lus90]. defined [gDH91]. defining [Sri90]. definite [Not90]. Definition [Met91b, Ser92l, KJ90, LO98l]. Definitions [DRSŻ91e, Mas95c, NZ99d, Pil94a, Bal94h, Ebe99e, Isa92i, Kön98a, Vät97a]. Deformation [DG90, Pom91, VK90, GR92, Rei90a, Shi91b, Ste91, Taf92]. Deformations [Sch91k, Rog90, Băd90, Beh91, Wan91a]. Degeneracy [FP98i, FP98d, FP98f, FP98h]. degenerate [BF93i, Oka90]. degenerated [Kos95c]. degenerations [Kaw92, MD90]. Degree [KB96b, KB96a, KB96d, Maĭ93a, Sho92s, vdG93, BEL91, CMZ91, FS90a, Gel92, GM90, GZ91, KLGP97e, Lor92l, MS90a, Man92, Maw93, Met91n, Met91o, Sho92q, Sho92r, Son90, Yap96e, Yap96f, ZR92]. degrees [Ars90, DM90c, HS90a]. Del [Fuj90b]. Delay [HMN91c, DvG91, DD91, FS91, Fru91, Żoł90, BM91]. delay-difference [FS91]. delays [Hal91]. Delhi [YS92]. Deligne [Han92, Har93c]. Deligne-Vermutungen [Har93c]. démonstration [Dav90, Leu96]. Denjoy [Yee90]. dense [How90, Pus96e]. d’ensemble [AJKY93]. densité [KPR90]. Density [McC99a, AL90a, Dow90, Oud91, Wir98e]. Dependence [Wal90b, Har90a, Per91, Sch91j]. dependent [BV90, FMSW91, Fro91, JPA97]. d’équations [MS91a]. d’Équivalences [Did90]. Dérivation [AY90]. derivations [Pis96b]. derivative [Gra91a, Ino92, Kac91, Üst95c]. Derivatives [KZ92c, Cha91h, EY93, Wah95b]. Derived [BL94a, Bjö91, Hub95i, Hub95q, KZ98, Hub95j, Rou98]. derivée [Dur90d]. Derivées [Par91a, Bon91]. descent [Gre92f, Neu97b, Neu97c, Neu97r]. description [Gli92, Woy98i, Pis96g]. Design [Sto92e, Hel91, Sto92n]. designs [Har93e, Met91a, Met91d, Met91l]. Désiré [McG92]. d’Estudis [CC92]. desuspensions [Cru90]. d’Été [BGM94, Ber95, Ber96, Ber98b, Ber99a, GGSC97, Hen90a, Hen92, Hen93, BPSH91]. detecting [Fit93]. determinant [Kra90]. Determinantal [IK99g, FP98e, IK99b, IK99f]. determinants [Lor92n]. Determination [GV91c, Oud91, Met91l]. Deux [CS96, Pra94]. development [LL93, Nov92, Wic91d, Wic91e, Zor92]. Developments [dBT91, Sch92a]. deviation [JS91b]. deviations [Bal90a, Fre92, Kif91, Led90, MWNPA92, Yur95a, Yur95e]. device [BMMP91]. d’exemples [Val94]. DG [BL94b, Kap91]. DG-modules [BL94b, Kap91]. d’Heisenberg [Ech96]. diagram [Ros90]. diagrammatic [Bog90]. diameter [BL91b]. Dichotomy [vN92f, DU92]. Dickson [JS92]. diédrales [Lab90]. Dieudonné [LO98g, LO98j, LO98i]. diffemorphisms [Gli92]. difféomorphismes [Don90]. diffeomorphisms [LQ95b, LQ95i, LQ95h]. Difference [Nil94b, Pot95b, vdPS97c, vdPS97f, FS91, KZ92a, Neu97d, Nil94e, Pot95g, Wah95k, Wet92, vdPS97a, vdPS97g, vdPS97h, vdPS97l, vdPS97m, vdPS97j]. Differences [KZ92c, Bar90c]. different [FPS92, Haz93, Hel92, SOA91]. Differentiability [DN99a, DN99d, Kos95a, Phe93c, XL98e, Day92, DN99c, Hu95, Phe93f]. Differentiable [Rei91a, Rei91b, FF97, GI91, HT90b]. Differential [Bui92d, BM91, Dim91, FPSW91, FIK90, FZ93, HMN91c, MO93c, Neu97n, Pia90, Sim91a, Sta92, TT96, Van91, XL98d, AT98, AS98e, BMO98, Bar97, Bax91, Bia90c, BH90c, Bui92b, Cal90, CMZ91, Car91b, Coh95, CW91b, CP91d, Dor93, Egg91, EQ97, ÉS99b, Gil90, Gru91a, Har90c, HS91, HS93, HMN91b, Hu96, JL91, Kac91, Kal99, KY91, KM97a, KP96b, KP96c, Liu96b, MY99, MY07, MT90, Mar92, Maw93, Mor91, NSS92, Neu97q, Nil94a, Nor92, Par90, Pom98b, Pom98d, Pyt99a, Pyt99f, Rau95b, Sch90p, Sch95a, Sch95b, Sim91k, Tai91b, Tal96, Tay92, Yau92, ZL90, Zvy90b, ZD92]. Differentially [Woo98]. differentials [Lea90, Mat90b]. différentielles [Cép95, JP91a, JY90, MS91a]. différentiels [Ber90c, dM93a]. Diffraction [CL99a, GDT91, JS91a]. Diffuse [Grü91b]. Diffusion [Ebe99g, ELL99h, Git98, Ama91, Blü90, Cap98, CML90, DP91i, DW97, FMSW91, KT93, Kus93, Maz97, MSS94, Nag93, Rai96, Röc99, Wat99]. Diffusions [Bar98, App90, Bal90a, Hec98, Kry99, McG92, Woy98e]. digit [Gra90a]. digits [AS90b, Kir90]. digression [Sim98a]. dihedral [Erd90a]. dilations [HS90b]. Dimension [AS98d, BC92c, Bui92d, Tur90a, AB90a, Aud90, Dat90, FK93f, Kna95f, Lus90, McC90b, Rut97m, Sch94c, Sch94b, cSyW91, Wal91a, Wri90, Zuo99f]. Dimensional [GHS97e, Göt94d, Kuk93a, Win95a, Bar93b, Bou91b, Dar91, Fis90, Fre94, Hen90c, Hey93, HS90b, IK99d, IK99e, Kai97a, KW94, Kim94, Kin94, Kna95e, Kna95h, Koc90f, KZ91, Kry99, KP96c, KB96b, KB96a, KB96g, Lan90a, Mat90a, Miz92, Mot91, Mul90a, NZ99f, OSK93, Ol’92, Ore90, Pal91, Rei90c, RG96g, Röc99, Tam99e, Val90c, Wah95h, Win95b, Win95c]. Dimensions [Da 99, Sch90u, Ber90b, Ebe99b, Ebe99a, FH90, FK93a, Sch91a, Wah95c, Wah95g, Yur95f]. diminishing [CS92a]. d’immersions [Gir90]. d’intégrales [BRS94, Pra94]. d’interpolation [DDD90, Phi90]. d’invariance [Cad97]. Diophantine [Beu93, EE93, Fal92, Lar90, MN91, Sch91d, Sch91e, Sch91f, Spr93c, Spr93h, Tij93a, Voj93]. direct [Apt93, CRW90, FS90a, HM91, VM92]. Directions [Nar97, Eva92h, PR97d]. Dirichlet [BH90c, Bou91c, Cha91d, Cha91f, Cha91e, Cha91b, DM93b, Fuj90a, Fuk93, Kur90, Röc93, Sch90a, Sch90e]. d’irreductibilité [Sch90i]. discontinuities [Rei91b]. Discontinuity [Bra98a, Kni91]. Discontinuous [Sim98b, Coc98]. Discrepancies [DT97d]. Discrepancy [DT97b]. Discrete [Mar99, Car91b, Day92, DS92, Fru91, Hen91, MN90, Vai93e, GN97]. discrètes [Lép92]. discretization [AL90b, ELLV90, Hu96]. discriminant [Chr91]. discriminants [Dam91a]. discs [Tum98]. discussion [CFJ^{+}92, DRSŻ91d]. disentanglement [Mar91]. Disentanglements [dJvS91]. d’Ising [Mai94, Mey91a]. disk [Ros90]. disordered [Gri97]. Dispersive [Dix97d, KPV90]. dissipative [GS90a, Liu96b]. distance [Bar90c, Che91, YK92]. distances [Bar90c, Woj91]. distant [Nag90a]. Distinguished [Her94b, DG90]. Distributed [Tib90c, DTW90]. distribution [Bou91b, DX93, DK93a, Drm90, DT97c, FT93, Fuj90a, GK90, Ika90, Mat90c, Pit99, Pra99, Win90]. Distributions [Col92h, Mol93d, Col92e, Col92i, CP91b, KÜ90a, MY91, Mol93i, Pan91f, Pom98f, Pus93, Shk93, Üst95a, Üst95i, Üst95k, Wes93, Yur95b]. distributive [KV91]. Divergence [Pfe90, Ken93, Üst95c]. Divergent [Bal94e, Éme97, Kos95c]. Diviseurs [FT90]. divisibility [Mol93e]. divisible [Wes93]. division [HS92]. Divisors [Now90, AMSV91, Ale91, SV91b]. does [FP91, Shi96]. Dold [Far96h]. Domain [Pom98a, Cha91d, Hel91, Hyl91b, PS90, Sto92o]. Domains [Cha91g, Koe99b, Väi92, BM90d, BM92b, Gal92, GKKL91, Han91, Kan91, Koe99e, Koe99g, Kom91, KB96c, Mie91b, MN90, Pil94c, Pri93, Rau95b, SM99t, SV90, Wie92, YR90]. dominated [Coc98]. domination [Del91g]. donné [Bor90a]. Doob [Pra90b]. d’opérateurs [Mey93b]. d’ordre [Jac91]. Doss [RS94]. Double [DL96c, PG90, AMS90, Tam90]. doublepoints [GCR90]. doubling [Bae91, BK92a]. Dowling [Met91c]. Dq [BF93a]. Dq-equivariant [BF93a]. drift [HW97, QH95]. drifted [KT93]. Drinfeld [BS97b, Gek90, Lau97]. driven [Kar91]. droites [Bea92]. dual [Bia92, Cas91, DGT91, Gau91, KR92h, Maj92, Zak91]. Dualité [Bel99]. Duality [Chi90a, FK93a, JcC91, LM90, LZ92, She93, Xue92f, Ban91e, Ben92, Mil91, Rut97k, Xue92b, Xue92c, Xue92d, Xue92e, vN92g]. Dubins [Sch99].

Dubreil [Mal91]. ducks [GI91]. due [Mey94]. d’un [BÉK^{+}98, Bor90a, CTR96, CDP90c, CDP90k, CDP90m, Dur90c, Dur90d, Eis97, LW90c]. d’une [CS98, DW97, EM91, Gra90b, Leu96, PE92, Rai96, Tur90a, Val94]. Dunford [Räb91]. Durand [LW90b]. DV [Pom98b]. dyadic [Skv90]. Dyer [Rub99b]. Dynamic [Lob91, Bae91, Ben91c, Krö90, Soc92, Ben91b]. Dynamical [BHS96i, BM91, Hor90, LQ95g, Pil94e, Pil99b, Rei91d, Wir98g, AC91, Arn95b, Car91a, CP93d, Cra91, ELL99g, HS90b, Ken94, Mat90a, Miz92, Pil94b, Rei91a, Rei91b, Wal91b, Joh95g]. Dynamics [Ben99, BC92c, CP93j, Hal91, Lau91, Mar99, Rau95b, Smi99b, TBR91, BK92a, BL94f, BC92g, BC92h, CP93i, DP91k, DGT91, Fie91, Jag93, Kru97f, McC99e, RS91a]. Dynamique [DM90b]. dynamiques [RS94]. Dynkin [Ura93b].

E. [Bou98b, DN99a]. é.d.s. [Mey91b]. each [WJ90]. échelle [Fla90]. échelles [Lem90a]. Ecole [BPSH91, BGM94, Ber95, Ber96, Ber98b, Ber99a, GGSC97, Hen90a, Hen92, Hen93]. Edoardo [Vil90]. effect [Bae91, Oza90, Tam90, Wan91b]. Effective [Kem91a, Hol90b, PSW92, Spr90]. effectively [Nis91]. effects [IG90, Nag90a]. effectués [BW94b]. efficiency [CRW90]. Efimov [Tam90]. Eichler [Sai91a]. Eigenvalue [Lei91, Fis90, LN91]. Einstein [Tia96]. Eisenstein [Har90b, Har93e]. Eisensteinkohomologie [Har93e]. Elastic [CL99a]. electric [Oza90]. electron [Ani91, Fro91]. Element [Wah95l, ELLV90, Gus90, Joh98, Mau90]. élémentaire [Leu96]. Elemental [McC99b]. Elementary [Col92d, DRSŻ91f, Eva92d, KLGP97d, Mas95d, deL94k]. Elements [Tal96, Tib90a, Koc90b, Wah95m]. elimination [Nis96b]. ellipses [Fre90]. Elliptic [Cha91f, Kos95e, Mie91b, Mie91c, MO93c, Pom98a, Rub99b, Sch90u, Spr93d, Tam99a, Tam99b, Vio99, Beh91, BS97b, Coa99, Eva97, ELLV90, Gre99, Ken93, Kin94, Kos95a, Kos95c, Kos95f, Mat90a, McI90, Mey99c, Mic93, MO93b, Pom98d, Sch90n, Sch90s, Sch90v, SY91, SS92, Tai91a, Tam99e, Ura93e, Wah95g, Wil97d, Zvy90b, Zvy90a, ZD92]. elliptique [CF90]. Embedded [Ion90, Col90b, Kus91]. Embedding [Leu91, Met91d, Wan90, HFS94f, Met91m, Mil94a]. Embeddings [Huc90, Heb96c, Hun96h, Met91s]. end [Amg98]. endomorphism [Azu90, Khu90]. endomorphisms [BL92c, CHT92]. energies [Bré93]. energy [BM90d, Jen90, Soa94a]. engineering [Alb98a, Col92h, Hel91]. enriched [Gra91b]. Enriques [DR91]. ensemble [DI98, Mic92]. ensembles [Cos90a]. Enskog [BL91a]. entière [Léa96]. entières [Gra90b]. entiers [FT90]. Entire [GLO94, deL94e]. entrance [Raj96]. entre [KS96]. entropie [Mic97, Thi91]. Entropy [BC92i, Daf96, LQ95b, Var93b, Iwa92b, Lan91b, LQ95a, LQ95c, LQ95d]. enumerable [FS90a]. Enumeration [Coo90, Kai97b, Kra90, Kum90]. enumerative [Pel91, XD90]. Envelopes [Xu96b, Xu96c]. environment [KT93]. epi [DZ93h]. epi- [DZ93h]. Epimorphic [SM99d]. Epimorphisms [SM99e]. epireflectors [SM99h]. Equation [KM97h, AFL90, AD90, BL91a, CD91, DP91j, DGDMN91, Ehr92, ÉS99b, GN97, GV91c, GV91b, GR92, HW91, Hla90, Hum90, Ino92, Kal99, Kat90, KM97c, Lax96, Mat90a, Mat90b, McG92, Meu90, Mic93, Mor93a, OT90, Sch91i, Sjö91, SOA91, Spr93j, Spr93i, Tur90b, Vai93d, deL94d, deL94r, AR94, Mor90b]. Equational [JR92b]. Equations [Alb98g, Beu93, BM91, Cha91f, Da 99, Dix97d, FIK90, FZ93, GV91a, HMRS90, HMRS92, HMN91c, Kru97c, Kru97k, Mey99d, Neu97n, Qua98, Sch91e, Sch90u, Sim91a, Spr93c, Spr93e, TT96, Vai93c, XL98d, deL94h, vdPS97f, Alb98c, Alb98d, Alb98j, AT98, AS98e, BMO98, Bax91, BL91a, Bel90b, Bet99, Bia90c, BM90d, BH90c, BT93, Caf91, Car91b, Coh95, Col92c, CNZ91, CW91b, Cor91, DP91i, DvG91, DGT91, Dor93, Egg91, EQ97, Eva97, Fab93b, FS92, FS91, Fur90a, Fur90b, Fur92, GH90, Ge92, Gil90, GV91f, GV91h, GV91i, Gru92, Gus90, HS91, Hel98, Hey90, HMN91b, HMN91i, Hu96, HP99, Ino92, Kac91, Kat93b, KY91, Kaz90, KPV90, Ken93, KF93, KM97a, Kry99, KP96b, KP96c, Liu96b, MY99, MY07, Mar92, Maw93, Mey99f, Mil94k, Mis92, Moo96a]. equations [Mor91, Mur93, NSS92, Neu97r, Neu97s, Neu97q, Oka90, OSK93, Ôta93, Oza90, Par90, Pil90, Pom98b, Pom98d, Pom98e, Pom98h, Pom98g, Pro90b, Puk90, Pyt99a, Pyt99f, Ran90a, Ran92, Rau95b, Riv90, Sal92, Sch91f, Sch91g, SvW92, Sch92b, Sch90n, Sch90q, Sch90s, Sch90t, Sim91k, Soc92, SV90, ST92a, Spi91, Spr93d, Spr93h, SS92, SW92, Tal96, Tay92, Tib90d, Vai93i, Val90c, Var92, Wie90, Wie92, Yau92, Zab99, ZL90, vdPS97a, vdPS97j, vdPS97d, vdPS97g, vdPS97h, vdPS97l, vdPS97m, vW90, Bon91, JP91a, JY90, Cép95, Par91a]. Equidimensional [Che90a, Lan90a]. Equidistribution [EGI94]. equilibria [Bar93i]. equilibrium [DP91a, DU92, Jou97, LP90a]. equivalence [Bou97l, Bou91e, Dam91a, HALM90, JR91a, Tho90b, Xi94d]. Equivalences [Isa92a, KZ98, Pic90, Rut97r, Ark90a, Boo90, Kah90a, Kah90b, Kah90c, Kel98, Lin98, Mar90b, May90, Møl90, Rut90, Yam90]. Equivalent [Boo90, Bal94h, BSÉ99, Did90, ÉS99a, Rou90, Sch99]. Equivariant [And91, BL94c, BL94d, BF93d, BF93j, Dam91b, DS90a, KB96d, Møl90, Sad91, Sch94g, Vis92a, Bar93a, BV93, BL94b, BF93a, BF93b, BF93h, BF93k, DS90b, DS90c, DS90e, Fie91, GZ91, KB96b, KB96a, KB96c, KB96g]. Ergodic [Led92, Les92, LQ95g, McC99b, NZ99e, Hof91, Jaj91f, McC99f, Sch91a, KRW92]. Errata [Gro97c, Isa92b]. error [Far96i, HH92, KTV92]. error-correcting [HH92]. errors [van90]. Escape [DP91a, LP90a]. Escape-equilibrium [LP90a]. escompté [Mic98]. Escott [RS90a]. espace [CDP90c, CDP90k, CDP90m, Deb92, ESH99, Gir90, MY91, Mey93a]. Espaces [CDP90e, Did90, dlR93, CDP90h, CDP90l]. Essential [SM99f]. Essentially [Shu98b]. estimate [Cha91b]. Estimates [Cha91h, DP91b, FMSW91, Ger90, Mil94e, Mog92, Pyt99a, AS90a, BM92b, FS92, GGS93, Kal93, KR92f, KO91, LN91, Ôta93, Räb91, Sch90q, Sch90r, Tai91f, Wah95g]. Estimation [LQ95c, LQ95d, Hen90e]. estimations [RS94]. Estivo [Bal93, BA94, CF92, Cap99, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. étale [Sch94k, Sch94r, Sch94q, Sch94p]. Étude [Val94]. Euclidean [AM91, FGL91, GKKL91, Lum91, Yuk98b, Yuk98g, Yuk98i, Yuk98k, Yur95c]. Euler [Kul92, FP98d, Kar91, Mar91]. Evaluation [HO91, Kra90, Yan96]. Evaporation [SOA91]. every [Met91f]. everywhere [Fuk93, Mou92b]. évitantes [Le 97]. Evolution [HS99, ST92a, deL94h, HP99, Kat93b, Ore90, Oza90, Mic97]. evolutions [Fri90b, Par91b, Par91c]. Exact [FS90c, Hen91, LS93, MR94, Pus96i, CG91, Hub95i, Sch94i, Uri90, Xue92a]. Exactly [Kaw93]. example [Chi95, MM97, Pom98c, Wah95a]. Examples [Ber99b, BMMN98c, Bra98b, CK91, Ene95b, HN93c, KW94, Kna95a, LO98a, LO98b, Lor92c, Pot95c, Pus96j, Sch94h, Ser92l, vdPS97d, Ark93, Bia95b, BL92d, BHS96h, Kön98a, Lan91h, Mas95d, NZ99d, Ros90, Sch94m, Sim91c, Soa94g, TO97a, Vis92b, Yuk98a, deL94k, deL94l, Har93b]. Exceptional [Cha91i, Acc94f, Acc94h, Bur93, Nog91]. exchange [Bab92]. exclusion [PR97f]. Excursion [Bur93, vdPS97e]. excursions [BM92c]. exemple [De 96, Mic98]. exemples [Pra94]. Existence [Alb98d, Brî96b, Fru91, GV91e, HMN91a, Isa92c, Lan96k, OL90d, Pot95c, Tam90, deL94g, deL94h, deL94f, AS93b, Haj91, Liv90, Met91l, Mot91, Sal92, Wat99, deL94e]. exit [Bar93d]. exit-length [Bar93d]. Exotic [Ban91c]. Expanding [Bot92, Kif91]. Expansion [Dix97b, BP90a, CD91, JL93b, Sch90l, Sho92b, vH90]. expansions [Bla91, Ble91, Dix97c, Dix97f, Eyr98b, GI91, Gra90a, Krö90, Mer94, Sim91g]. expectations [EY93, Jaj91c]. experimental [Kim94]. Explanation [Lan91b, Lan91c]. Explicit [JL94, JLG94, Gol94]. Explorations [Bur91]. explosion [Wer94]. exponent [BC95, Dar91, Hof91, Kni91]. Exponential [Kaz94b, Kaz94d, Kaz94c, Le 94, XL98f, AS90a, BD94, DK93a, HYL91a, Hen91, Mao91, Pus93]. Exponentially [deL94i]. Exponentielles [Rod92, DL90]. Exponents [ACE91, Cra91, Fla91, Gol91a, HH90, Kel91a, Lei91, LQ95c, LQ95d, LQ95b, Per91]. exposants [Bou91a]. exposé [Arn99]. Exposed [Ban92]. Expressions [Kro91]. exptally [ASHY90]. Extended [Man97b, Her90a, Lan91b, Lan91c, Lem90d, Spi91]. extendibility [Tum98]. extending [Rei90a]. extension [GKW91, GC92, GDT91, Jan90b, KB96c, Le 91, MS90a, RS94, SDS90, Zip91]. Extensions [Ber90c, Eyr98c, FS91, Gre92c, CT92, CF95, Dat90, FPS92, Gre92e, Gre92h, Jan90a, Lan96l, Ole90, Opo91, SM99d, Shp90, Wil97e, Wil97f, Wil97d, Xue92d]. exterior [Gal92, Sal92, SV90]. Extrapolation [Mil94f, Mil94g, Mil94c, Mil94b, Mil94j, Mil94k, Rus90, vN92g]. Extremal [Par93, Saf90, Suf90, Tan91].

F [Ste90b]. F.B.I. [Del92b]. facteur [FT90]. factor [GJ91]. Factoring [BLP93, Pol93a]. Factorisation [BR90a]. Factorizable [BFS98c]. factorization [AE90, Bea90, Kel91b, Zuo99c]. Facts [Brî96c, Göt94c, Mie91k, Sch94n]. Fading [HMN91b]. faibles [Bak91]. Fall [Kul92, Lat90]. Faltings [Edi93, Fab93a, Koo93, Oor93b]. Families [Brö92, BHS96i, deL94h, BHS96c, BEGG90f, Fit93, XL98i, deL94g, deL94f, deL94e]. famille [CR92, Maz97]. family [BD94, Muk91, PR97j]. Fano [Ale91, Ara91, KMM92]. Fans [Mar96b]. fast [Del91a]. Feb [KZ93]. Feldman [Sch99]. Feliu [ACC92]. Fermat [Col90a]. Fermeture [MS94a]. fermions [Hu93]. Few [Met91j, Met91i, Met91f, Sab91]. Feynman [Fuj93, Ich90]. fiabilité [CTR96]. fibered [Kru97f]. fibering [Koz90a]. Fibonacci [CS92b, Tem93]. fibrations [Far96g, Far96i, Far96j, Mas95b, Per92]. fibre [Bel90b, Mas95f]. fibres [Sie91, Bea92]. Field [FK93g, AG90, BL93b, BCDS92, BLP93, Cou93, Dub96, LL93, Len93, LLMP93, Mat90b, Ped92, Rou90, Sch94c, Sch94b]. Fields [AH92d, CC92, DRSŻ91i, FR90c, KLGP97e, Ser92e, vdG93, BK94, Boi96, Bou98c, Box91, BK92c, Col92b, Del91a, Del98, Dob96, DRSŻ91a, FPS92, GS90a, Gem90, GC90, Gre92e, Gre92h, HS93, HH90, Iwa90, Kin94, KLGP97b, KB96g, Mar96e, NZ99l, Now90, Oda90a, Oza90, Pfe90, Pil98b, SvW92, SM99o, SM99s, STV92, Son90, Tha90, Woo98, Yan90]. Fifty [LW90b]. Filtered [FZH90, HFS94d, Hub95j, Ser92f, HFS94a, HFS94g, McC90b]. filtering [ELL99d]. Filtrage [Par91a]. Filtration [JY90, Arn99, KLGP97a, Rai96, Wat99]. filtrations [ABÉH95, BP90a, BÉK^{+}98, ÉS99a, Her94c, Her94j, JS94]. Filtre [Bou91a, Jac91]. filtres [Mey90]. fin [AJKY93]. finance [Son97]. Financial [Run97]. find [Sto92f]. finding [KLGP97a]. Fine [Yur95a]. fini [DI98, Mic92]. finie [Mou90]. finies [Mic97]. Finite [AH92d, Bui92d, Ene95c, HM90, KLGP97g, Neu97d, Pot95d, Pra90a, Sch91c, Soa94b, Wah95l, Wet92, Ade92, Alp90, Ben92, CP93c, Dam91b, Ebe99a, ELLV90, Far96e, GI91, Gus90, HKR92, JS94, JPA97, Joh98, KLGP97o, KS90b, Kry99, KB96b, KB96a, Mau90, Ric98, Röc99, SC97, Sch90g, Seg90, SDS90, STV92, Šip92, Soa94a, Sri90, TY90, WJ90]. finite-dimensional [KB96b, KB96a]. Finitely [BK92c, GI91, CP93d, Gre92h, Neu90, Rut97b, Rut97a]. Finiteness [Abr96a, Sim91e, And91, Hol90b, Sch94f, Son90]. Finitude [Eca90]. Finsler [AP94a, AP94b, AP94d]. First [Bui92e, Ebe99b, Hyl91b, Pyt99b, Raj96, Ari91, Del92c, GV91f, GV91i, Gra91a, MZ90, OS90, Sch93a, Wah95b]. five [DLS90]. Fixed [IT90, Vig90, Wit92, Acc94h, AB91, Bru92a, Gel90, Kuč90, Nus93b, Zvy90a]. Fixpoint [Jay91]. FKG [BM92a]. Flag [BR90b, FP98e, LO98c, FP98g, Taf92]. Flat [Xu96c, Xu96e, Xu96d, FGL91, Fuj91, Her94e, Zuo99a]. Flatifying [Her94c]. Flatness [Sch90g, Her94i, Her94k]. Flensted [Wan90]. Floquet [Lan91c]. Flour [BGM94, Ber95, Ber96, Ber98b, Ber99a, BPSH91, GGSC97, Hen90a, Hen92, Hen93, BPSH91]. Flow [Neu97e, Ant92, Fla91, Hey90, Koz90b, Krö90, Puk90, SvW92]. Flows [ELL99h, GHS97e, Mie91c, NZ99f, AH90, Ari91, EY93, Fis90, LS91, LQ95i, Mie91d, NZ99a, NZ99c, NZ99g, NZ99l, Raj91, Sim91d, Woy98g]. fluctuante [Ber90b]. Fluid [TBR91, Gli92]. fluids [Raj91, Riv90, Val90c, Wol92]. focal [BW91]. Fock [Mey91a, Bel90a, Mey93e, Mey93f, Mey93m, Mey93a, Mey95a, Mey95b, Mey95h, Mor90a, NV90, Oba94e, Par91b, Par91c, Sch93e, Sch93h]. focus [PR97i, Sch90j]. focusing [Bun92]. Fold [Met91q]. Folding [BW91]. Foliage [Lat90]. Foliations [Neu97f, Bot92, Led92, MORT91]. fonction [Fra95, HT90b]. fonctionnelles [Lim90]. Fonctions [Gra90b, Fer91, KPR90, MY91, Mey93c]. Footnotes [Cat90]. Forced [Woy98b, FP91]. Forcing [Bek91c, Bek91g, HRS91]. form [Ant92, Col92f, DRSŻ91h, Gor91, Ken93, Kos95c, Lum91, Maj92]. Formal [Bal94d, CD91, AFF90, Har90a]. formalism [Col92b, Nor92]. formation [Caf91]. formations [Lan96h]. forme [Mou90]. Formes [Gir90]. Forms [CP91g, DM93b, Fre91e, IK99c, IK99d, IK99e, LS90a, Nil94b, Pan91h, Pro98a, BH90c, BK92c, CP91d, CP91e, CP91c, Dim91, Fre91d, Fre91f, Fuk93, Har90c, HS93, IK99i, Lyu92, Man92, Mer94, Oda90b, Pan91g, Pan91i, Pan91j, Paz91, Röc93, Sch95a, Sch95b, Sch90x, Sna92, Spr93f, Wan94, vdPS97b]. Formula [Sho92g, AR96g, AS98f, BC92m, DL96c, Eyr98j, FR92, Gol94, JL93b, LQ95a, MO93a, Ogu90, Sho92b, Sho92d, Sho92h, Tak91, Tha90, ÜZ97, Woy98d]. formulae [Jaj91b]. Formulas [JLG94, Col90b, Col92f, DS90g, FP98e, FP98a, GV91d, Hla90, IK90, JL94, Kle90, LN97, Mas95e, Pal91, Tem93]. formulation [FP98g, LR90, MR90]. formules [Le 90]. fortement [DW97]. Forward [MY99, MY07]. Forward-backward [MY99, MY07]. Foundations [Kna95b, Kna95i, Dob93, Spr93a]. four [DW91, Moo96d, Tyu96]. four-manifolds [Moo96d]. Fourier [Del92c, Hou90, Kaj91, Lie93d, Lou91a, Mer94, Sin92, Wei94a, Wei94b]. Fourth [AvW90, Man92]. Fox [Lus90]. Fractal [FH90, Woy98c, Wic91g]. Fractals [DK93c, Wic91a, Bar98, Kus93]. fractional [Var92]. fractions [Dum91, Jac90, Thr90]. Fragments [Coa99]. Fraïssé [Clo90]. Framed [BP97e]. frames [App95]. framework [Gra91a]. Franca [Cap99]. France [Alb90, Ben91b, CMR92, FR90c, LS90a, LW90b, SG97, ST92c, SG97]. Fréchet [Day92]. Fredholm [Fit93, ZR92]. Fredholmian [Khi92]. Free [Bra98a, Ser92g, Tib90b, Ant92, Ben91a, CW91a, CP93e, Cru90, Gol90, HA90a, IT90, IR92, Kar97a, Kuč90, LN97, Neu90, Pil90, Shp90, Soc92, ST92a, Ver94, Wol92]. Free-Discontinuity [Bra98a]. French [NF90, Anc90, AS90c, AS93a, AS93b, Ans94, Arn92, Arn93, Arn95a, Arn99, AAG90, ABÉH95, Aud90, AW97, AY90, AJKY93, AR94, AJK^{+}96, ARY96, Bak91, BM92a, Bak94, BE96, BÉK^{+}98, Bea92, Bek91a, Bel99, Ber90a, Ber90b, Ber93, BW94b, Ber90c, Bia92, Bia95a, BRS94, Bon91, Bor90a, Bou91a, Bun90, BPSH91, Cad97, CDD90, Cap98, CF90, Cép95, Cha96, CS96, CS98, CTR96, CT93, CDP90b, CDP90a, CDP90c, CDP90d, CDP90e, CDP90g, CDP90i, CDP90h, CDP90j, CDP90k, CDP90l, CDP90m, CJ90, Cos90a, CR92, Dav90, Daz90, De 96, DW97, Deb92, Del90a, Del90b, Del91g, DI98, DL90, DDD90, DM90b, Did90, Dix91, Don90, Dum91, Dur90c, Dur90a, Dur90b, Dur90d, Eca90, Ech96, Eis97, EP90, EM91, Eme91, ÉY98, ESH99, Fer91]. French [Fla90, Fon98, Fou95, FT90, Fou90, Fra95, Fra90, FM90, Gir90, GC92, GRCW90, Gra90b, Gro90a, Hau91, HL92, Hen90a, Hen92, Hen93, Hen90b, HT90b, Hu93, JP91a, Jac91, JS92, JY90, JY93, Jou98, KS96, KS90a, Kro91, KPR90, Lab90, Lan90c, Le 90, Le 97, LW90c, Léa96, Lem90a, Lem90b, Lem90c, Lép92, Leu96, Lic90, Lim90, Mai94, Mai93b, Man97f, Mar90c, Mat98, Maz97, MNM90, MS91a, Mey90, Mey91a, MY91, Mey91b, Mey93c, Mey93b, Mey93a, Mey94, Mic92, Mic97, Mic98, Mig90, MS94a, Mon94, Mor90b, Mou90, Mul90c, Par91a, Phi90, Pic91, PE92, Pra90b, Pra92, Pra94, RS94, Rai96, Rod92, Rog90, Sch90i, SN90, SW93, Sch94u, Ser94b, Ser94c, Ser94d, Ser94a, SG97, Thi91, Tur90a, Val94, Wal90a, dM99, dM93a]. French [dlR93]. fréquence [Fla90]. Frequency [Hel91]. Freud [Tot94c]. FRG [AvW90, ASMS90, HMRS90]. Friedman [Fur92]. friends [I90]. Frobenius [Col90a, OS90]. Front [AGADH90, ALMV93, AJMR95, BCE^{+}97, BEF^{+}94, BHMS99, BDLL96, BHS96g, CRGR99, CSJT98, CTKV93c, Del92d, DDH^{+}99, DK93b, EMM^{+}90, FFG^{+}93, FMMN93, GKM^{+}96, HNP^{+}90, KLS97, Mey93g, Pan91d, Phe93e, Pis96e, Ser92h, Ser92i, Ser92j, Sou97b, Van96c, XL98g, Lie93e]. Frontière [Bia92]. ^{≫} [Kru97g]. FTQs [LO98k]. Fuchs [GH90]. Fuchsian [Bur90, Bol92, Bur74]. Fully [Nar95a, Eva97]. Function [Neu97o, RSSV90, Sch90n, SM99q, Wir98g, AP94b, AS90b, Bor90b, Bou98c, CL99g, Dur94, Fuj90a, Hit91, Kaj91, Kir90, KM97d, LS93, Meu90, Mou92a, Ped92, Rei97d, Tan91, Tha90, Tod97b]. Functional [Fro91, Hey93, HMN91c, LM91, NV94, Sim98c, deL94h, deL94j, Bal90a, CMZ91, HMN91b, Ino92, Kho93, Meu90, Res92, Üst95h, Wat90, FIK90, Kom93a, OR91, YS92]. Functional-analytic [Fro91, FIK90]. functionals [AS98h, CMN95, Fuk93, KS92, Oba94f, Ric99b, Sch93d, Üst95b, Xue90, Yuk98b, Yuk98i, Yuk98k]. Functions [Bal94e, Bia90d, Bra98c, CM97c, Che90b, CP91g, DN99d, Ene95g, Kit94f, Nis96c, Pan91h, Phe93c, Pro98a, Rob90, SM99g, Ser92a, AVV92, AB90a, BL93a, BGR91, Bek91g, Bia90a, Bia90b, Bra98g, CMZ91, Cas92, Che91, CPS90, Col92a, Col92e, Coo90, CP91e, CP91b, Day92, DP91b, DN99a, Ene95a, Ene95c, FS90c, Fuj90a, GK94, GLO94, Har90c, HOR91a, Hu95, Jam90, JS91b, JL93b, Kat90, Kat93a, Kit94a, Kit94b, Kuč90, KZ92e, Li91, Maĭ93a, Mao91, Mat90c, MO90, Mit94b, Mit94f, iM90b, Mol93h, Mol93a, Mot91, Mou92b, Nab90, Nis96a, Nis96d, Nis96e, Nis96f, NM92, Oda90a, Ôta93, Pan91f, Pan91i, Phe93d, Phe93h, Pom98d, Pom98e, Pom98h, RS93, Rus90, SM99r, Sch90x, Sen90, Shi90, Sho92e, Suf90, Tho90a, Tum98, UT92, Wir98a, vH90]. functor [Bou97c, Bou97d, JT91a, Kap91]. functorial [Mac91]. Functors [BL94d, Bou97i, HFS94b, BL94a, Bou97b, Bou97g, Bou97h, Bou97k, Far96b, Far96f, HS92, Hub95g]. Fundamental [Göt94c, HMN91d, KB96c, Xu96f, Zuo99e, AR96a, Ara91, Fab93b, Fre91a, HA90a, JR91b, Kön98b, Kul91, Lum91, Sch94i, Wil97c, Zuo99f]. Fundamentals [Ino98a]. Further [Eva92e, GP91a, Pro90a, Lea90]. fusion [FK93i]. future [Law91b, Shi96].

G [Hit91]. G. [Ste90b]. Gâteaux [Day92, Phe93f]. GAFA [LM91].

Galerkin [Coc98, Fis90, SW92, Wah95l]. Galois [Ber90c, Gre92c, Gre92d, Gre92g, Jan91, KW94, Kim94, Kin94, Lan96m, Rib99, Ser94b, Ser94c, Ser94d, Tat96, vdPS97d, vdPS97f]. galoisième [Ser94b]. galoisienne [Ser94d, Ser94c]. Galtchouk [CS96]. games [Bar97]. gaps [GL92]. Garnier [Van96d]. gas [Kaz90, SOA91, Spi91]. gases [Fro91]. gasket [Hec98]. gauches [Dum91]. gauge [BL90]. Gauss [AMSV91, Gek90]. Gaussian [BEL91, Fab93b, Led96, Wes93, Yur95c, Yur95b, Yur95f, Yur95g]. gaussienne [LW90c]. Gelfand [McC90b]. gem [Hun96d, Far96i]. gemischte [Har93a]. gemischter [Har93e]. Genera [Tam99b, Tam99a, Tam99e]. General [Bui92f, DT97c, Far96i, HMN91e, JR92c, Nil94c, Oor93b, Röc93, Ros94, Sch90b, Sim98d, BL93b, Bra98d, CHT92, Lan91d, McK90, Nil94f, Sim98k, Tot94b, Tyu92]. Généralisation [Thi91]. Generalised [McG92, Dug92, Mul90a]. généralisée [Cha96]. generalization [Bom91, Phe93g, Pis96l, Tyu91]. Generalizations [Sch90c, Hom90, Kuč90, vH90]. Generalized [AB90b, BTV95b, BTV95c, BP91, Bia90a, Bia90b, Bia90c, Bia90d, DV91, Gel92, Joh95a, Mas95e, XL98h, Yee90, deL94u, Alb98c, BF93g, Col92a, Col92e, gDH91, HK92, HKR92, JS90, Kow92, Koz90a, Li91, Muk90, WS90, Xi94e]. Generated [Ebe99g, Ric99e, Wir98g, AR91, Gre92h, Nag90b, Neu90, Nie90, Pom98h, Ric99d, Rut97b, Tod97a]. generation [Mar96d]. generator [FK93f]. Generators [Rut97f, ELL99e, Zie96f, deL94a, deL94j, deL94n]. generatrices [Gir90]. Generic [BF93e, DZ93a, HU90, Miz92, Mot91, Pil94b, BGR91, BV91, Pil99e]. Genetic [Lan91g]. Genus [Per92, Bar93c, Bar93b, DG90, Mat90b, Tho90a]. geodesics [McC90a]. géodésiques [CDP90l]. Geometric [Del92e, Dur94, Ehr92, Fab93a, Gre92h, HWW93d, Heb96b, HS99, HP99, Jon95, Kru97b, KB96d, Moe95f, MM91b, Tam99d, Yuk98c, AP94b, And91, Bou98b, HSS91, Kal93, Kru97a, Par92, Pom91, Pra91, Sho92b, Sho92c, Suz90, LM91, dBT91]. Geometrical [AD90, CG97c, Kin94]. Géométrie [Don90, Lan90c, Alb90, CDP90f]. geometries [Mar98, Fra90]. géométrique [Gra90b]. Geometry [AR96a, Bal93, BDF91, Bou98b, BL90, CC96, CMR92, Dub96, ELL99h, FPSW91, FG90, GT90, HN93d, Hun96e, Kru97c, NO91, Pot95d, SBL90, ST92c, TPZ98, Van96g, Vil90, XD90, AP94a, AP94d, Ade92, Ber90d, BBG91, BR90d, Cal90, Che91, CTKV93a, Con92, Fal92, FP98j, Gin98, GC90, Hit90, JP91b, JY91, Leh90, Lie93a, MM90, MN91, MT91b, Moo96d, Muk91, Mus92, Nab90, Par92, Pel91, Sch92a, Sil92, ST92b, Sza91, Tat96, Tyu92, Van96h, Voj93, Yau92, Zak91]. Georgia [Ina90a]. germ [Mar91]. German [Har93a, Har93c, Har93b, Har93e, Har93d, Hla90]. Germany [HLN91, HMRS92, HPSS92, KRW92]. Gevrey [CD91, Kom91]. Gibbs [Str93b]. Gibbsian [Dob96]. Ginzburg [Bet99, Neu97s]. Given [Joh95g, Bal93, BA94, BM90c, CF92, Cap99, CC92, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, HK92, Lio93a, MP91, Nar97, Odi90, Qua98, Rug96, Run97, Ser92k, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. GKS [BE96]. Glauber [DP91c, Mar99].

Gleichungssystemen [Hla90]. Global [AP94b, BGV90, BGV92, FPSW91, Lio93b, Moo96a, Sch90d, Siu91, CDD90, FPS92, Ruí90, Wah95d, Wah95g, Zip91, Zor92, ZR92, FPSW91, FG90]. Globally [Gil90]. Glueing [Sch94k]. Gluing [Hub95k]. Godement [Hub95l]. gonality [Pel92]. good [Ale91]. Goppa [GL92]. Gordon [Bal91]. Gorenstein [IK99g, IK99i, Oli90]. Göttingen [Woy98a]. graded [CRV90, Eyr98a, Eyr98h, Eyr98i, IK99a, Kan90, Sta92]. Gradient [AK90, BF93a, BF93b, BF93k, Gus90, SOA91, WS90]. Gradients [Neu97n, Neu97a, Neu97i, Neu97m, Neu97p, Neu97o]. graining [GVV90]. grand [FT90]. Grande [Alb90]. graph [Blo92, Cas91, Met91p, SDS90, Tuc90, Wir98b]. Graphes [Anc90]. Graphs [Eva92g, Ura93b, Yap96e, Yap96f, Yap96j, BP97k, Joh95e, Mur91, SZ90, Spe93, Yap96c, Yap96b, Yap96h, Yap96i]. Grassmannians [FP98g, Jaw91, Sko92, Cra91, Kum97]. gravity [Ger92]. Green [Bou97h, Bou97c, Bou97b, Bou97d, Bou97g, Bou97i, KM97d]. Gromov [CDP90f, CP93g]. Gross [Tha90, Üst95c]. Grothendieck [Hub95m, MNM90, Pis96i, Vis92a]. Group [ACC92, Bur74, KZ98, Kov90, Lan96l, Lat90, Leh90, Neu96d, Rut97g, Sch94l, Ark90a, BMMN98n, BH90b, Bui92a, BEGG90a, BEGG90b, CP93c, CP93d, CP93e, Dri92, DLS90, FF97, FK93d, Ge92, Ger92, Haj91, Han92, Hig90, HA90a, Hol90a, JL91, Kah90b, Kho93, Kul91, Lan96a, LO98n, McK90, MT91b, Ped92, Pis96h, Pri90, Sna92, Win95b, Win95c, WJ90, Xi94e, Zim98c, Bur90]. groupe [Dix91]. Groupes [CDP90g, EP90, KS90a, Lic90, Maz97, Arn92, Ber90c, CDP90b, CDP90f, Don90, Ser94a]. Groupoídes [Daz90]. Groups [Abr96b, Abr96e, AR96f, BOERT96c, BTV95c, BC92a, BFS98c, BMMN98d, BMMN98j, Bui92d, BEGG90c, CP93j, Dek96b, Eva92g, FG96, FK93g, KLGP97g, Koc90h, Kul92, LS90a, Lan96m, Lan96p, New90, SZ90, Ser92b, Ser92f, Ser92k, Sho92s, Tho90a, ZD93, Zuo99e, Ade92, AR96b, AFSTS92, AP90, AP91, Alp90, Ans90, Ara91, Ban91c, Ban91d, Ben92, BTV95b, BFS98b, BMMN98c, BMMN98g, BMMN98h, BMMN98i, Bia95b, BL92b, Boo90, Bui92b, Bui92c, Bui92f, BEGG90d, BEGG90f, BEGG90e, BEGG90g, BR90a, CRW90, CS92b, CGST92, CP92, Coh92, CK91, Cri90, Dam91b, DP93, Dek96a, Dek96e, Dek96f, Eva92c, Eva92d, FF96, FH90, FK93b, FK93e, FP98e, FP98a, Gli92, Gro90b, HM90, Haz93, HN93h, HA90a, HKR92, HH90, IT90, Ish92, JR91b, Kah90a, Kin94, KLGP97h, KLGP97c]. groups [KLGP97j, KLGP97l, KLGP97o, Kol91, Kom93b, Las98, pL91, Lun92, Lus90, Lyu92, Maj92, McC90b, Neu96e, Neu90, Nil94f, Pat90, Pot95e, Pot95f, Pra90a, Ric98, RS90c, Roh90, Rou98, Rut97f, Rut97j, Sch94g, Sch91c, Sch90x, Seg90, SDS90, Sho92r, Shp90, Sri90, mS91b, Tam99d, TY90, Tod97a, Vis92a, Wil97c, Xi94b, Xi94c, Zuo99f, deL94a, deL94j, deL94n, vdPS97d, Kov90, Pic90]. growth [BC95, XL98f, Yuk98l]. Guide [Mum99c, Neu97f, Sce90b]. Guíxols [ACC92]. Gundy [Mon94]. Güstrow [KRW92].

Haar [Kit94f, Kit94e, Mül91]. Haar-like [Kit94f, Kit94e]. Hadamard [DZ93b, DZ93c, DZ93d]. Half [DPlPS90, Koe99c, Cha91e, Cha92, Koz90b]. Half-canonical [DPlPS90]. half-space [Cha91e, Koz90b]. Half-spaces [Koe99c]. Halperin [Lup90]. Hamiltonian [Bar93i, Col92b, CvdM90, DM96, FS90c, FS90d, Fuh91, Kru97d, Kuk93a, Kuk93d, LR90, MF94a, Mey99e, Mie91d, Mie91c, Mie91e, Mon91b, Van96a, Van96f, Van96e]. hamiltoniens [DM90b, Aud90]. Hampshire [Lat90]. handle [Mas95f]. Handlebodies [Joh95b, Joh95d]. Hankel [Fuh91, Pee94]. Hankelian [Pis96f]. hard [SOA91]. hard-sphere [SOA91]. Hardy [Mit94f, GE98c, LL91, Mit94d, MN90, Shi91a, Wei94b, Wei94c, Wei94g]. Harmonic [BTV95c, BR90c, BR90f, BR90j, Cha91j, JY91, Lem90c, Mey93h, mS91b, Zuo99a, BL93a, BP91, Bré93, BR90a, HS93, Mey95c, Mit94b, MN91, Oba94d, Ohs91, Sch90k, CDZ91]. Harmonique [Lem90c]. Harnack [Mur93]. Harvard [Ser92k]. Hasse [Kat93a]. Hausdorff [DU92, cSyW91, Wic91d]. Heat [Wol92, Wat90, Yan96]. Heat-conducting [Wol92]. Hecke [Aki90, Oda90a, Wan94, Xi94b, Xi94f, Xi94g, Xi94h, Xi94l]. Heegaard [Joh95e]. Heidelberg [AvW90]. Height [MS94b]. Heights [Hui93, Sch91h]. Heiles [Van96d]. Heine [Bal94c]. Heinz [Sch90w]. Heisenberg [AR96b, AR96f, BTV95b, BTV95c, Ech96, KV92, Neu96d]. Held [CC96, AvW90, ACC92, ASMS90, ACE91, AK90, BBD90, BCC92b, Bal93, BA94, Ben91b, BDF91, BS90c, BLM^{+}90, BM91, CF92, Cap99, CPR91, CR91, Cel90, CP93b, CMR92, Da 99, DM93b, DL97, FPSW91, FG90, FG96, FR90c, FIK90, FZ93, GT90, Gra94, HLN91, HMRS90, HMRS92, HS99, HPSS92, Ina90a, JOP91, JLP90, Joh95g, KZ93, Kom93a, KÜ90b, KRW92, Kul92, LS90a, LW90b, Lat90, Mim90, MP91, NF90, Nar97, NO91, Odi90, Pic90, Qua98, Rug96, Run97, RSSV90, SBL90, ST92c, TT96, TPZ98, TBR91, Vil90, Vio99, Vis94, XD90, YS92, ZD93, dBT91]. Helgason [Bom91]. helices [Nog91]. Hénon [Van96d]. Henri [LW90b]. Henstock [BLM^{+}90, MO90, Mul90b, Sen90, dF90]. Hermite [Lor92k, Lor92l]. Hermitian [Dam92, mS91b, YK92]. Herz [Pis96f]. heteroclinic [CA91]. heuristic [Lan91d]. heuristics [Yuk98h]. Heyting [Law91a]. Hidden [AFSTS92]. hierarchical [Mau90]. High [Jen90, Bou91b, GM90, Yap96e]. Higher [Day92, Hub95n, KM97h, Lie93b, Lie93e, LPLG^{+}92, Sai91a, Vai93b, XL98d, Eyr98j, Göt94e, Kna95f, Lan90a, Lie93g, Lum91, Mat90b, Rei90c, Sta92]. Higher-Order [KM97h]. highest [Kac91]. highly [Yam90]. Hilbert [Bol92, Den92, DL96a, GC92, Göt94b, Göt94a, Göt94d, Gre92b, Har90c, Her94d, Her94h, Kle90, Kuk93d, Lab90, Mad91, Neu97b, Neu97c, Oda90b, OL90b, Pan91g, Pan91h, Rus90, Zab99]. Hill [Mey99f]. Hille [deL94c]. Historical [Zim98a]. Hitting [BM92d]. Hochschild [Léa96]. Hodge [BA94, BK91, Gre94, Hub95o, Jan90e, Oda90b, Sai91b, Sch95c, Sch95d, Zar91a, Zuo99c, vG94]. Hölder [SSS92, Sch90o, ZD92]. Holomorphic [BL94f, Brî96d, Ebe98, GK94, Zai90, AP94c, Brî96b, Cal90, CS92c, CP91d, CP91e, CP91c, MN91, MR91b, Pan91j, Sar92, Vig90, deL94b]. holonomy [SSS92]. Hom [Hub95a]. homeomorphism [Zor92]. homeomorphisms [Mil90a]. homoclinic [BC92l]. Homogeneous [BMMN98e, BR90d, Gro97a, Hec98, Win95a, HN93i, KT91, pL91, Lun92, Spi91, Win95b, Win95c, Wir98c]. Homogenization [Bor98d, Bor98c]. Homological [CG97d, Dat90, Far96j, Ina90a, Bro90b, Hel92, Her90b, Lus90]. Homology [BP97f, Del91c, Del91d, Gre90, Gro90b, Hel92, Koz90a, Lod92, Rud92, Rut97d, Rut97q, Sim91i, iY90]. homomorphism [Muk90]. homomorphisms [IT90, Pis96b, Pis96j]. homotheties [Res92]. homotopical [Bro90b, Pir90]. Homotopie [Did90]. homotopiques [HL92]. Homotopy [Abr96c, ACC92, Far92, Far96d, Koc90h, Mar90b, McC90a, Pus96k, Rut97h, Rut97i, Rut97j, Rut97r, SV91a, Sło91, TO97d, Ani90, Ark90a, Ark90b, Bau91, BL92d, Boo90, Coh90, Coh92, DM90a, Far96b, Far96f, HALM90, HA90a, Kah90a, Kah90c, Lun92, Møl90, Pus96f, Rut90, Rut97q, Rut97u, Sta92, Yam90, Did90, Mim90]. homotopy-self-equivalences [Kah90c]. Honest [ASHY90]. honor [BM91]. honour [LW90b, Vil90]. Hopf [Ama91, Bru92a, CvdM90, FS91, FK93b, Ino92, Sta92, Wil95, Woy98d]. horizontal [App95, PE92]. horloge [Ber90b]. Hörmander [LL91]. Hrushovski [BMMN98f, Bou98b]. hull [Mou92a]. Hurwitz [AS90b]. Hutchinson [Wic91b]. Huxley [MN90]. Hydrodynamic [DP91d, DP91e, DP91h, Woy98e, Var93b]. Hydrodynamics [DP91f]. Hyperbolic [CP93j, KN91, MM90, Qua98, Woy98f, AB91, Boi96, Caf91, CP93c, CP93d, CP93e, CP93g, CP93k, Daf96, Drm90, Hör91b, Kaj91, LQ95f, Liu96b, Nis91, SSS92, Shu98b]. hyperbolic-dissipative [Liu96b]. hyperbolicity [Lie93c, Zuo99f]. hyperbolique [CDP90c, CDP90k, CDP90m]. hyperboliques [CDP90b, CDP90e, CDP90g, CDP90f, CDP90h, CDP90l, Mou90]. Hypercontractivité [Hu93, Bak94, Mic97]. Hypercontractivity [Üst95d, HW97, QH95]. hyperelliptic [BEGG90b, Sch91g, Spr93d, Spr93e]. hypergroups [Hey93, Kum92]. Hyperkähler [Ati90, Fuj91]. hyperplane [Băd90, Bia90e, CvdG92, Mas95i, Rei90a]. Hyperspaces [Tod97c, Wic91a]. Hypersurface [Mas95h, Dim91, Kul91]. hypersurfaces [FGL91, HB91, HP99]. hypoelliptic [CML90]. hypothetical [Met91e]. Hysteresis [Bro94, Vis94, Ver94].

I. [Cat90]. Iagolnitzer [Del92c]. Ideal [Vät97b, Vät97d, Vät97c, HH90, Iva93, Sch90g, Ver91]. ideal-transforms [Sch90g]. Ideals [HWW93f, CG97e, HWW93b, HWW93c, HWW93d, HWW93e, HWW93g, HALM90, Lus90, Sid92, Sim90, TT90, Wri90]. ideas [Wir98f]. idempotent [Far96b, Far96f]. identical [Mor93b]. Identification [MO93d]. identité [Leu96]. identity [BH90a]. ie771 [Gre92e]. ie771--extensions [Gre92e]. Igusa [Meu90]. II [HMRS92, KÜ90b, RS91a, Tai91j, Acc94b, Aik96, Bor90c, Dix91, Dur90b, Fuj90a, HN94b, IK99e, Kit93, Kle90, KP96c, MO90, MT91b, MS90c, Mum99e, Pri90, Rei97c, Sai91b, Sch94c, Shi90, TY90, VC90, Wil97f, Xue92a].

III [KRW92, Ion90, Mum99f, Wil97d]. Ik [Ura93f]. Ik-conditions [Ura93f]. ill [Fur90a]. ill-posed [Fur90a]. Illposedness [GV91f]. Image [Mad91, Koc90c, Mar91, Mey90, MO93g]. images [Dam91a]. imaginary [Lyu92, Now90]. imaging [Gem90, GDT91, Lou91b]. imbedding [Mil94j]. Immersed [CL99a]. immersions [BOWV91]. imperfections [Jou97]. Implementation [Pyt99c, Sto92g, BL93b]. Implicit [Kac91, Isa92d]. implies [Iwa92b]. implique [Mat98]. imprimitive [Sho92l, Sho92m]. improper [Koz91]. Incidence [Pot95e]. inclusions [Pia90]. incomplete [Bea90]. incompletely [Lio93a]. incompressible [Gli92, Hey90, Ran92, Riv90]. increasing [Kit94b]. increment [Sch91l]. Increments [Mey93i, Mey95d]. indecomposable [Hub90]. indefinite [ELLV90]. Indépendance [Bun90, Wal90a, De 96]. independence [Nis96a, NÜZ90, Üst95b]. Independent [Mey93i, Mey95d, Alb98j, AFF90, DP91d, FS90b, JPA97, KK93, Sch91l, Yur95d]. Index [SS93, Ale91, Bar93f, BV93, Brö90a, FR92, Mis95, Nus93b, Wat90, ZR92]. indexé [Arn99]. indexés [Fou95]. India [YS92]. Indice [dM93a, LW90c]. indistinguishability [Kal99]. Individual [Jaj91f, Sim91l]. Induced [BL92b, Xi94j]. induction [GM90, Hla90]. inductive [PR97g]. Induktionskoeffizienten [Hla90]. Industrial [CF92, Bus92, Kui92]. industry [For92]. inégalité [Cap98, Mat98]. Inégalités [Bak91, CDP90i, CDP90h, BM92a, BE96, Mon94]. inelastic [Alb98e]. Inequalities [Her94e, KZ92c, Wei94a, Yur95d, Far96c, FS90c, Fre90, Gro93, GE98a, Hit91, Led99, Lou91a, MT90, Mil94a, Pan91a, Sch90p, Sto92a, Sto92b, Sto92f, Sto92h, Sto92l, Sto92m, Str93b, Tib90e, Üst95g, Üst95h, Üst95a, Vel90, Yuk98e]. Inequality [GE98c, Acc94c, DS95, Hof91, Hu91, JS91b, Lar90, LN91, Mur93, Pis96l, Pis96i, Rez91, Tal91]. infimum [Shi96]. Infinitary [McC99c]. Infinite [BMMN98j, Da 99, HMN91c, Kuk93a, Kum97, KP96c, Mol93e, Mul90a, Soa94d, Tam99e, Ark90b, Bar93b, Dar91, Ebe99b, Hen90c, Iwa92b, KLGP97o, KB96g, Mat90a, Mot91, Ol’92, Röc99, Sch91a, Yur95f]. Infinite-Dimensional [Kuk93a, Mul90a, Bar93b, Dar91, Hen90c, Ol’92]. infinitely [Mil94j, Wes93]. Infinitesimal [Băd90, Gre94, Pir94, Rei90a, Beh91, ELL99e]. infinitesimally [Wan91a]. infinity [KM97b, Kus91]. inflated [Met91q]. information [CL96]. informative [Sab91]. Infra [Dek96f, Dek96g]. Infra-nilmanifolds [Dek96f, Dek96g]. inhomogeneous [Fri90b]. Initial [Alb98e, Alb98g, Gru92, HS90a, KPV90, Lio93a, Tai91c]. Initial-Boundary [Alb98g, Alb98e, Tai91c]. injectif [AJK^{+}96]. injective [Sch91b]. injectivity [AHJLP90]. inner [Dix97b, NV91, Neu97d, Neu97e, Neu97p]. input [Bel90a]. Inspired [Cap99]. instabilities [Jag93]. Instanton [Oko91]. Institut [CC92, LW90b]. Institute [CDZ91, Ina90a, Kul92]. integers [BLP93, Pol93a, Sou92]. intégrabilité [Lép92, Val94]. integrability [MM97, OS90, Tak99]. Integrable [Kuk93a, Van96f, Van96e, Van96g, DM96, FS90d, MO90, Nov92, Sen90, Van96a, FG96]. intégrables [DM90b]. Integral [GV91a, GE98d, Lie93d, Pom98a, Ric99d, Sch90o, Vai93c, Ari91, BI92, Bas90, Dav91a, Gin98, GV91c, GV91b, GV91f, GV91h, GV91i, GV91j, HW91, Ich90, Kaj91, KJ90, MWNPA92, MT91a, Mit94a, Muk91, Nar95c, Nil94a, Pom98h, Pom98g, Sob90, Tak91, TPDZ91, Vai93d, Vai93i, Xue90, TPZ98]. integral-valued [Nar95c]. intégrale [Cha96]. Integralen [Hla90]. Intégrales [Fra90]. Integrality [iM90b]. Integrals [Bre94d, BLM^{+}90, Dav91d, DRSŻ91i, Ene95d, FF96, Mit94d, Bre94a, Bre94b, Bre94h, Bre94i, Dav91b, DN99b, DRSŻ91a, DK94, FS90c, Fuj93, GH90, GK98, Hen90d, HMN91j, Hla90, Hu91, JS90, KP96b, KP96c, Mar90a, Mey99l, MR98, Mul90a, OS90, PG90, Sai91a, Sch90j, Yee90, dF90]. integrated [deL94p]. Integration [Hen90c, LN97, Sim91j, Wat90, Yan91b]. intense [Ani91]. interacting [Col92b, Mél96]. interaction [Kar97a]. interactions [McK90]. Interest [Bjö97]. interface [DP91k]. interference [Nag93]. intergration [Kru97b]. interior [Alb98i, Pri93]. Interludium [Van96h]. Internal [Alb98g, Hub95a, Jan90a]. International [Alb90, BBD90, CPR91, KZ93, Kom93a, KRW92, Kul92, Mim90, NO91, SBL90, ST92c, TBR91, Vil90]. Internationale [BA94]. Internazionale [Bal93, CF92, Cap99, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. interpolants [IS90]. Interpolation [BKS93, Lor92f, Nik94a, vN92g, BGR91, DDD90, Lor92a, Lor92e, Lor92i, Lor92k, Lor92l, Lor92m, Mil94e, Wei94e]. interpretation [AFF90, Gol94]. Intersection [BFS98d, Bur74, dJ93, Bur90, Del91f, Gor91, Le 94, SS91]. Intersections [Man97g, HU90, Man97a, Neu90]. Intertwining [Bia95b]. interval [Der95, Hof91, Kel91a, KM97f]. intervalles [Fou90]. intervenant [SN90]. Intrinsic [Law91a, Oll93]. Introducing [Neu97g]. Introduction [Abr96d, AH92b, AR96c, Alb98f, Bar93e, BTV95d, BL94e, BFS98e, BC92d, Bor98e, Bou97j, Bou98a, Bra98e, Bre94f, BF93f, BHS96h, BR90e, CM97e, Cha91k, Col92e, CP93f, CP91f, CL99b, DP91g, Del92f, DL96d, DS90b, DRSŻ91g, Ebe99c, ELL99f, Eva92f, Eyr98d, Fre91b, FK93c, FP98f, GHS97c, GV91g, GE98e, Gro97d, Har98c, Har98d, HFS94c, Hen90d, Hin98, Hun96f, IP92, Ino98b, Isa92e, Kar97b, KR92b, KLGP97f, Kna95c, Koc90d, Kru97e, Küh95d, KB96e, KZ92b, Lan96f, Lan91e, Lan91g, Lem90b, LO98d, Lie93f, Lod92, Lor92d, Man97c, Mar96c, Mas95g, McC99d, Mey99g, Mie91f, Mil94h, Moe95g, MO93e, NSS92, Neu97p, Neu96b, Nie91f, Nil94d, OL90a, Pan91e, Pis96g, PR97h, Pyt99d, Rei97a, Rei91c, RG96c, Sch93b, SM99i, Sch95e, Sho92n, Sim98e, Sim91f, Suz90, Tai91e, Tam99f, Tot94d, Ura93c, Üst95e, Üst95f, Van96i, Vät97e, Wah95c]. Introduction [Wic91c, Wil97b, Wir98d, Xu96g, Xue92c, Yuk98d, Zie98, Zim98b, Zuo99b, vG94, Bou98b, CP93a, Coc98, Col92d, GP91b, Lan91b, Lan91c, Pla90, ST92b, Yap96a]. Introductory [Per99]. Intuition [deL94k]. intuitionistic [Ner90]. Invariance [Wac92]. Invariant [AR96d, Bax91, Gro97a, HN93e, Mal91, Nil94b, Pir94, AP90, Bal90a, BV93, Fit93, GCR90, Hit91, Hof91, Kem91a, Koc90b, LQ95h, MR91a, Nar95a, Pil99c, Res92, Sch91c, Sho92d, Tam99c, Wah95a, Wah95k, Wic91b]. invariantes [Lic90]. Invariants [GHS97d, Gro97e, Moo96b, Mur91, NZ99g, AVV92, BP97a, Dek96g, Dix91, Har90c, Ion90, Kai97e, Kem91b, Oko91, Van91]. Inverse [For92, Sch94m, BL92a, Gem90, Gro96, KZ91, vdPS97g, vW96]. Inversion [Pal91]. investigation [Wah95b]. inviscid [Woy98f]. involutions [Cru90, Wan91a]. involutive [App90]. Iomdine [Mas95e]. ionized [Fro91]. IR [Maz97]. Ireland [BLM^{+}90]. irreducible [Sho92o, Sho92p, Sho92u, Zar91a]. Irregular [BCC92b]. irregularities [Drm90]. isochore [Fra90]. Isogenous [Xi94g]. isolated [Gor91, NR90, Ohs91, Sch91k]. Isometries [OL90b, Soa94e, CDP90d, CDP90l]. isomorphism [FF97, Hom90, LM90]. isoparametric [Wah95a]. isoperimetric [Rez91, Tal91]. isopérimétriques [CDP90i, CDP90h]. Isoperimetry [Led96, Yuk98e]. Isoptics [CMM91]. isosceles [LP90a]. Isotone [Isa92f]. Isotopies [Nab90]. isotropy [Mar96f]. Israel [LM91]. Itô [Eyr98e, Eyr98j, Hol90a, MWNPA92]. Italy [BBD90, BCC92b, Bal93, BA94, CF92, Cap99, CPR91, CC96, CR91, Cel90, CP93b, CTKV93a, Da 99, DM93b, DL97, FG90, FG96, FZ93, GT90, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, SBL90, TT96, TPZ98, TBR91, Vil90, Vio99, Vis94, ZD93, dBT91]. iterated [Kry95, LL90]. iterates [Jaj90, Jaj91d]. itératif [DDD90]. Iteration [AL90b, BB90, Mil94i, Neu97k]. iterations [Wir98f]. iterative [De 91, Kos95g, Neu97q, Not90, DDD90]. Itinerary [BOERT96d]. ITPACKV [KO90a]. IV [BGV90, Mic93]. Iwasawa [Coa99, Gre99, Kat93a].

J [Ste90b]. J. [Sch91j]. Jacobi [Fon98, Hou90, Rei90c]. Jacobian [CS92c, CvdG92]. Jacobians [Mum99b, Mum99c]. James [KS92, Lun92, Sho92g]. Jan [BM91, KZ93]. Japan [FIK90, Kom93a, Mim90, NF90]. Japanese [NF90, Koc90e]. Jensen [Wan90]. jet [Rou90]. jets [Cha92]. Jeulin [Mey94]. Johnson [Hit91]. joining [War99]. Jones [Mil90b]. Jordan [BMMN98c, BMMN98g, Koe99a, Koe99d, Koe99e, Koe99f, Koe99h]. juin [CT93]. Julia [BK92a]. July [Bal93, BDF91, CPR91, CC96, CR91, CDZ91, FG90, FIK90, FZ93, Gra94, Kom93a, KÜ90b, NO91, Run97, Vio99, Vis94]. Jump [Col92f, Xia92]. Jumping [HRS91, JS94]. jumps [Coh95]. June [ACC92, ACE91, AK90, BBD90, Bal93, BA94, BDF91, Cap99, CC92, Cel90, CP93b, CMR92, DM93b, DL97, FPSW91, FG96, FZ93, HLN91, HS99, JOP91, Joh95g, MP91, Nar97, Odi90, Qua98, SBL90, ST92c, TPZ98, XD90, ZD93, dBT91]. just [KLGP97o].

Kôsaku [Kom93a]. Kac [AFSTS92]. Kadomtsev [AD90]. Kähler [JY91, Tam99e, Tia96, TO97f, TO97d]. Kähleriennes [Lic90]. Kähleriens [Mor90b]. Kahn [Mil90b]. Kalman [Bou91a]. Katata [NO91]. Katenoid [Pom98c]. Kawasaki [DP91c]. Kazhdan [Xi94h]. KDV [I90, Lax96, Mat90b]. Keller [Fur92]. Kenneth [BM91]. Kergin [Lor92e]. Kernel [Sho92e, Bou91e, Kna95g, Yan96]. kernels [Wat90, vW96]. key [Bor90b, Ehr92]. Khinchin [Kry95]. kind [GV91f, GV91h, GV91i]. Kinematic [Kna95d]. Kinetic [Mar93a, Pul96, Ark93, BM90d, DP91e, Lio93b, Spi91]. Kinosaki [Mim90]. Kirchhoff [Soa94c]. Kirillov [McC90b]. Klein [BEGG90a, BEGG90c, BEGG90b, BEGG90d, BEGG90f, BEGG90e]. Kleinian [FH90]. Kloosterman [CPS90]. Knight [War97]. Knödel [Pro90a]. knot [Mur91]. Knots [GHS97e, Jon92, Cru90]. known [Lan91f]. KO [Ros91, iY90, KO90b]. KO-assembly [Ros91]. KO-codegree [KO90b]. KO-homology [iY90]. Koblitz [Tha90]. Kodaira [Her94f, Her94g, Zuo99f]. Kohomologie [Har93d]. Kolmogorov [Da 99, Kry99, OL90c, Tik90]. Konstruktion [Har93e]. KPZ [Woy98a, Woy98c]. Krull [Wri90]. Kubota [Ste90b]. Kullback [CL96]. Kunita [CS96, AS93a]. Künneth [Hub95p]. Kurzweil [Sen90]. Kutta [Pyt99f]. Kyoto [Kom93a, NO91].

L [BM90a, Ste90b, Rut97k]. L-R [Rut97k]. L. [Mey91b]. Lê [Mas95h, Mas95e, Mas95i, Mas95j]. lack [CK91, MT90]. lacunary [cSyW91]. Lagrange [CvdM90, IS90, SW92]. Lagrangean [Kru97f, Kru97g, Kru97h, Kru97i, Kru97j]. Lagrangian [AMSV91, Gli92, Mie91c, Mie91g]. Lagrangienne [Mul90c]. lagrangiennes [Gir90]. Laguerre [Fon98]. lambda [Gra91b]. Landau [Bet99, Neu97s]. Lang [Bou98b, Bou98c, Hin98, Oor93b, Pil98a]. Langevin [AFL90]. language [AFF90]. Lannes [HS92]. Laplace [Bal94f, BC95, Bre94h, Dix97c, Sim91g, XL98i, deL94d]. Laplacian [BBG91]. Laplacians [Neu97j, mS91b]. Large [Bal90a, Bek91g, Dix97d, Kif91, MWNPA92, ABW92, BEL91, BM90d, Fre92, Fur92, Gra96, Hen90e, Jaj91e, Jib90, Led90, Woy98h, Yur95e]. Large-time [Dix97d]. larger [deL94u]. last [Kne92, Raj96, Sch92a]. Lattice [MN90, Bab92, JR92a, Pol93b, Van96d]. Lattices [JR92g, And90, GJ91, GP91b, Hol90b, Räb91, Shi91b]. Laurent [Dum91]. law [Hen90e, Kry95, Xue90]. Laws [Ast91, Bru90, Bur93, Daf96, Jaj91e, Joh98, KS90b, Shu98b, Soa94c, Tad98]. layer [Fis90]. leading [Vai93i]. Leary [SS90]. least [Elf91]. Leaving [Tho92]. Lebesgue [Bek91g, dlR93]. Lecture [Kus93, Cha91n]. Lectures [Bal93, BA94, CF92, Cap99, CC92, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, Fra91, FZ93, Gil94, Gin97b, Gra94, Gro96, HS99, Joh95g, Kat93a, Mar99, Mol94, Moo96b, MP91, Nar97, Odi90, Qua98, Rug96, Run97, SC97, Ser92k, Sho92g, TT96, TPZ98, Vio99, Vis94, Woy98a, ZD93, dBT91, Dur95, Ner90, Spe93, Tyu96, BGM94, BV93, Ber95, Ber96, Ber98b, Ber99a, GGSC97]. Lefschetz [RS90c, Roh90, Tro91]. left [AP90]. left-invariant [AP90]. Leibniz [Law91a]. Lemma [Koo93, Fre91a, RS90e, Sch91h, Sim91h]. lemmas [Her94a, Sim91e]. lemme [Mig90, Pra90b]. length [Bar93d, Bar93f, Bar93h, Bar93g, FZH90, Tan91]. Leningrad [GS93, Kul92]. lente [Rai96]. Lenz [Kna95g]. less [Bar90c, FK93f, Sho92r, Sho92s]. level [Box91]. Levi [Nar90, Pom98d, Pom98e, Pom98h]. Lévy [App95, Bal90a, BK92b, Chi95, DS92, DÉY93, Fag90, MR94, Shi94]. Lévy-transformation [DS92]. Lewis [Bal91]. Lewy [Sch90w]. Liapunov [Kat90]. Lie [Ser92k, AP90, Alp90, Arn92, BR90a, Daz90, Eyr98g, Eyr98h, GR92, HN93a, HN93d, HN93g, HN93f, Hol90a, Ish92, JL91, Kho93, KS90a, Lic90, Lun92, Naz92, Neu96e, OS90, Ol’92, RS94, Ser92f, Ser92g, Ser92l, Ser92k, Ser92m, Ser92n, Ser92p, Ser92q, Sri90, Sta92, mS91b, Wal90b, Win95b, Win95c, Zar91a]. Lieb [Hu93]. lien [KS96]. Liénard [AV91]. lift [Böc90, Heg92]. Lifted [Sto92h]. Lifting [Nie91g, Sto92i, Foi91]. liftings [Kim94]. lifts [BR90h]. light [DvG91]. like [BTV95e, KS92, Kit94f, Kit94e]. likelihoods [CJ90]. Liminf [Shi94]. Limit [GH90, Jaj91h, KK93, Mol93f, Mol93g, Neu96d, Pil94c, Sch94n, Sch93c, Wal90c, AL90a, AFL90, Chi90b, ERHG91, FH90, Fre92, Gin97a, GVV90, Hey93, Kho93, LP90c, Mal92, Mol93c, Mol93h, Mol93j, Neu96c, Per90, Wac92, Woy98f]. limite [GRCW90]. Limited [Sch91b]. limites [Eca90, Eis97]. limiting [HMN91i, Mil94c, Mil94a]. Limits [DP91h, Brö92, DP91d, DP91e, Iva93, Pul96, Sch94m, Woy98e]. Lindel [Man97h]. Lindenstrauss [Sch91j]. Line [Brî96e, dJ93, AR96b, BEL91, Bor92, Col90b, Dam92, Dem96, Gor91, Hub95e, KO90b, Met91f, Pal90, Soc92]. Linéaire [Par91a, Daz90, Del90a, Del91g]. linéaires [Bon91, JY90]. Linear [Ben91c, BMMN98h, Cha91f, Col92g, Dix97d, HN94a, HN94b, HMN91f, HMN91g, KLGP97g, LS91, Met91i, Met91j, Met91h, Met91g, Met91f, Neu97b, Nil94b, Rus90, Sko92, Spr93f, Zar91a, vN92c, ABW92, Bal94d, Boi96, Bor90c, CT92, EQ97, Fis90, Fre92, GV91f, Gra90a, HS91, HMN91e, HMN91j, IK99i, Kat93b, KTV92, Khi92, KS92, Met91b, Met91m, Met91o, Met91p, Mie91h, Not90, OS90, Phà91, RB92, Sch90s, Sch93d, Sim98b, Wil95, WJ90, Zvy90a]. linearization [Shi90]. linearized [She93]. Lines [Met91j, BSW92, GS90b, Hun96a, Met91k, Met91i, Met91h, Met91f, MT91b]. Link [Kai97c, Kai97d, Kai97e, Kai97f, Kai97b, Kos91]. Linking [Kai97g]. Links [GHS97e, Wri90, Kai97h]. L’interversion [GRCW90]. Liouville [Bab92, KM97h, Sim91n, Wal90a]. Lipschitz [MT91a]. Lipschitzian [BH90c, KO93]. liquid [Ant92, Bré93, ST92a]. List [Kah90a]. lists [GCR90]. Literature [Mum99c]. Littlewood [MN90]. Ljapunov [Hof91]. Local [BM90b, Bog90, Fie91, FK93d, GT91, Her94g, Joh90, NZ99h, Sch90p, SG97, Wah95d, Ani90, BK92b, BL92a, Bra98f, CDD90, Del91a, ELY97, Erd90i, Erd90k, FK93a, Her94e, Her94h, Le 94, MR94, Meu90, NZ99k, Pit99, Sid92, Wah95j, Wah95g, Zim98d, Mum99f]. locale [HT90b, SG97]. Localisation [SN90]. Localization [BFS98f, Far96d, Rut97l, Far92, Far96f, Far96j, MZ90, SS92]. localizations [BC92a, TT90]. Localizing [iM90a]. Locally [Azu90, Del91c, FS90b, KR92c, Kna95e, Kna95f, AR91, BK92a, Del91b, Del91e, HB91, JJS92, KO93, KT91, Sch90r, XL98i]. location [MR91a, Yuk98c]. locaux [Eis97, ÉY98]. Loci [FP98i, IK99g, FP98c, FP98d, FP98f, FP98h, IK99f]. locus [IK99b, IK99f]. Loewy [FZH90]. log [Ale91, Mat98, MR91a]. log-concave [MR91a]. log-Sobolev [Mat98]. logarithm [Kry95]. Logarithmic [Gro93, Str93b, Led99]. logarithmique [Cap98]. logarithmiques [Fon98]. logarithms [Spr93f]. logic [Ner90, Odi90]. loi [AS93b, Fer91, MS91a]. Lois [BM92c, CF90, CTR96, Fer91, Jou98]. long [CG91, Kit93, Met91f, Sch94i, Shi96]. long-range [Kit93]. loop [DR90, Hub90, Jay91, Wer94]. loss [Żoł90]. Lösungen [Hla90]. Lotka [Egg91]. Low [Koc90f, AL90a, Cha92, FZH90, IG90, Yap96f]. Lower [Met91k, Phe93h, Räb91, DK92, Lus90]. lowest [Xi94i]. LSS [Woy98h]. Luminy [Ben91b, FR90c, LS90a, ST92c]. Luminy/Marseille [LS90a]. Lunar [Mey99h, Mey99f]. Lusztig [Han92, Xi94h]. Lyapounov [Bou91a]. Lyapunov [ACE91, Bor90b, Cra91, Dar91, Fla91, Gol91a, Kel91a, Kni91, Lei91, LQ95c, LQ95d, LQ95b, Mao91, Ole90, Ôta93, Per91].

M [Eve93, Mal91, Ste90b]. M. [Blo92]. M.-P [Mal91]. Maassen [vW96]. MacDonald [UT92]. Machlup [CMN95]. Mackey [Bou97k]. Macroscopic [vdB91]. magnetic [GP91a, Iwa90, Sjö91]. Mahler [Nis96a, Nis96c, Nis96d, Nis96e, Nis96f, Spr93i, vH90]. Mai [Alb90]. Main [LO98e, LO98f, AH92c, AH92h, Kuk93b, Kuk93c, Sim91h, Wah95g]. mainly [Wah95i]. major [Sch92a]. Making [Pla90]. Malliavin [Ste90a]. manifold [AB90a, App95, Chi91, DvG91, Fuj93, HM90, HB91, LN97]. Manifolds [AP94c, BP97g, Heb96f, Joh95f, Kai97f, Mie91c, NZ99f, Pet96, Rut97m, Ser92c, TO97f, Win95a, ABW92, AT98, Ati90, BK91, BP97c, BP97d, BP97e, BTV95e, BBG91, BC92l, Blü90, Box91, Brî96f, BR90b, CS92b, FJ91, Gau91, Ham91, JcC91, Joh95a, Joh95c, Kai97b, Kai97c, Kai97d, Kai97e, Kai97h, KR90, KT91, Kru97f, Küh95a, Küh95c, Küh95b, KB96b, KB96a, LQ95h, McC90c, Mie91d, Mie91k, Moo96d, Mot91, Nar90, Nor92, Sob90, Tam99d, Tam99e, Tia96, TO97d, Tyu96, Win95b, Win95c]. Manin [Gek90, Hru98b]. Many [CP93b, BSW92, CP93a, KLGP97o, Mil94j, STV92]. Many-Particle [CP93b, CP93a]. map [BF93g, GZ91, Heg92, Hof91, Koc91, Mar91, MR91b, MO93g, PR97e, Ros91]. mapping [CS92c, HM90, Hou90, PS90]. Mappings [Nar95b, Vuo92, BS90b, Ebe98, Geh92, Gel90, Gel92, Iwa92a, KR90, Mar92, MN91, Nar95a, RB92, Ric92, Sch90k, Sch90m, Sre92, Vig90, Wah95a, Wit92, Zor92, dPW91]. Maps [BC92e, BF93j, BR90j, Her94g, KB96d, Pis96k, AMSV91, AVV92, Ara91, BEL91, BC92b, Blo92, Bré93, BF93c, BF93d, BF93h, BF93i, BW91, Bru92a, BR90c, DU92, Fit93, GU92, Han90, Her94f, JY91, Kai97g, Kel91a, Kif91, Kos91, KB96b, KB96a, KB96c, Lie93a, Mon91a, OU91, Oli90, Pet92, Phe93g, Pis96c, Pis96d, Sch90b, Sch90e, SSS92, Väi92, Zuo99f, ZR92]. March [ASMS90, Ben91b, CDZ91, RSSV90]. Marches [Le 97]. Marcinkiewicz [Bul90]. Marden [Min90]. marges [Bel99]. marginals [HK92]. Market [Jou97, BR97]. Markov [AS98c, AS98e, AS98d, AS98g, AS98h, Bar90a, Bax92, Bia95b, Cat99, CL96, CTR96, Coh95, Daw93, Ebe99d, Ebe99f, Fou95, Har98a, Har98b, Har98d, Har98e, Lei91, Mai94, Mai93b, Mal92, Mic97, Par91c, SC97, Son97, Tai91d, Wir98c]. Markovian [CP93e, CP93h]. markoviennes [BM92c, JP91a]. Martin [Bia92, LN97]. Martina [Cap99]. Martingale [Hit91, Wei94b, Wei94c, Wei94g, Bur91, CF95, CS98, Éme97, Hob98, Kry95, PE92, Wat99]. Martingales [Kaz94b, Arn92, Arn93, Arn95a, AW97, ARY96, DÉY93, ELY97, JS94, JPA97, Jaj91c, Kaz94a, Kaz94c, Lép92, Mey93b, MT99, Pra99, Raj90, Sch91l, Sch94u, Tak99, Val94, Wei94f]. Maschke [Weh90]. Maslov [NSS92]. master [Van96j]. matching [Dix97b]. Matematico [Bal93, BA94, CF92, Cap99, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. material [DS90f]. Materials [Alb98g, Alb98c]. Mathematical [Bre94g, DP91h, FG90, GS93, Gra91a, Ina90a, Kui92, Kul92, Lou91b, Mai99, Rug96, Sin91, Dob93, DZ93c, OL90e, Son97, CF92, CC92, HLN91, TBR91]. Mathematics [CDZ91, Run97, Bus92, Col92h, For92, Jon92, Nov92, Tho92, Tik90, Cap99]. matrices [Col90a, Dar91, DTW90, FZH90, Gol91a, Har98f, IK99c, IK99f, KS90b, Man97i, Mau90, Nag90b, deL94a, deL94n]. matrix [BGR91, BCDS92, GE98a, Ika90, SV91a]. Matter [AGADH90, ALMV93, AJMR95, BCE^{+}97, BEF^{+}94, BHMS99, BS97a, BDLL96, BHS96b, BHS96g, CRGR99, CSJT98, CTKV93b, CTKV93c, Del92a, Del92d, DDH^{+}99, DK93b, EMM^{+}90, FFG^{+}93, FR90a, FMMN93, GKM^{+}96, HNP^{+}90, Hru98a, KLS97, KP96a, Liu96a, MF94b, Mey93d, Mey93g, Nus93a, Oor93a, Pan91c, Pan91d, Phe93b, Phe93e, Pis96a, Pis96e, Rau95a, Rub99a, Sce90a, Ser92d, Ser92h, Ser92i, Ser92j, Shu98a, Smi99a, Sou97a, Ste99a, Str93a, Val90a, Van96b, Van96c, Var93a, XL98c, XL98g]. Mawhin [Ste90b]. Maximal [DD91, HN93h, Sid92, Gro97e, HSS91, Iva93, Pan91a, Sim98i, YR90]. maximum [Del91g, Hob98, KT93, Kus91]. May [ACE91, BBD90, GS93, LS90a, LW90b, Rug96, SBL90, TT96, Vil90, Lin90]. Mazur [Sza91, Woj91]. McDuff [TO97a]. McKean [Mél96]. Mean [GE98c, BB90, Eva97, Fuj90a, HSS91, Kus91, MM91a, SvW92, Ste99b]. méandre [Ber93]. means [Kin94]. Measurability [Bek91d, Bek91g]. Measure [Daw93, Ric99e, Cha91j, HK92, Hof91, Led99, Li91, Oll93, Sch99, Tal91]. Measure-valued [Daw93]. measurements [Sab91]. measures [Bax91, BM90c, CP91e, CP91b, DU92, LQ95a, MR91a, Pan91f, Ric99f, Ric99h, Val90b, Vät97c, Woj91, Yur95c]. Mécanique [Alb90]. Mechanical [Bor98c, Bor98d]. Mechanics [DK93c, Gra94, Dob93, Ich90, Mey99d]. Medalists [CC92]. media [Gol91b, Mol94]. medial [BH90a]. median [Shk93]. medium [Col92g, Zhe96]. Meeting [FR90c]. meets [Met91f]. mehrfachen [Hla90]. Mehta [Tyu91]. Memory [Alb98g, Kom93a, YS92, HMN91b, HMS94]. Meromorphic [Che90b, Cas92, CPS90]. meshes [Wah95a, Wah95k]. mesurabilité [Pra90b]. mesures [DI98, Le 97]. metals [Alb98e]. metamathematics [Clo90]. Metaplectic [Pro98a]. Method [Pom98a, Sch95d, AE90, AL90b, Bek91e, BL92a, Bor90b, Cha91c, Coc98, Dow90, Fab93b, GKW91, Hen90e, KY91, Mil94d, MN90, NSS92, Neu97i, Neu97k, PS90, Pia90, Pyt99b, Pyt99g, Ran92, Sak90, Sim91g, SW92, Var92, Yan90]. méthode [Mey91b]. méthodes [Bun90, Fla90]. Methods [AK90, Bar93j, CG97e, CC96, DP91h, FIK90, FZ93, Gre94, HLN91, HMRS90, HMRS92, KB96d, McC99b, Pyt99e, Rug96, RSSV90, Voi94, Wah95l, Ark93, Bel94, Ber90d, Bet99, BH90c, Col92f, De 91, Dob96, Gus90, Ike90, Joh98, Kem91a, Kru97b, Neu97q, Not90, Par92, Riv90, Rod94, Rut97e, Rut97n, SY91, Tal96, Vai93b, Vai93f, Var93b, WS90, Wet92, Cel90, GS93]. metric [Sch90m, Wic91d]. Metrics [AP94b, AP90, Dam92, Ger90, Haj91, Mol93i, Tia96, Zuo99a]. métrique [CDP90m, ESH99]. métriques [CDP90e]. Meyer [Üst95g, Üst95a, Zhe96]. Microlocal [Kom91, Sjö91, CR91]. microlocale [Bon91]. Microlocalization [Del92b, Lie93b, Del92c, Del92g, Lie93g]. microstructure [Mül99]. Mild [vdPS97h]. Milman [Sch91j]. Milnor [Mas95b, Mas95f, Sch91k]. minihypers [HH92]. Minimal [BCC92b, BOWV91, Mar96d, Neu97h, OL90d, OL90e, Yuk98f, Bek91e, Bor90b, FS90a, HB91, Hon90, OL90c, OL90f]. minimale [AS93b, Tur90a]. minimally [Beh91]. Minimax [Sim98f]. Minimization [CL96]. minimizing [MF94a]. minimum [Bar90c, YK92]. Minkowski [Mol93j]. Minnesota [Koe99f]. Minoration [Rod92]. minorations [Fon98]. miroirs [Mey90]. Mirror [Bor92]. mirrors [Bun92]. Miscellaneous [BL94g, Jaj91g, Küh95e]. mise [Mou90]. Misiurewicz [Blo92]. Mixed [Hub95q, Jan90e, Jan90d, Wil97c, Bor92, ELLV90, Har93a, Har93e, Hub95r, Sai91b, Wil97a, Wil97e, Wil97f, Wil97d]. Mixing [Wir98e, Kow92]. mixtures [Spi91]. mobiles [JY93]. Model [Bou98b, Pil98b, Bou98a, Ech96, Grü91b, KY93, Pil98a, Pom98b]. model-theoretic [Pil98a]. modèle [Ech96, Jac91, KS96, Mai94, Mey91a]. Modèles [Del90a]. Modeling [Ani91, Die99]. Modelling [CF92, DGDMN91, Had99, Kui92]. Models [BCC92b, Isa92g, KZ93, Mar99, TT96, AL90a, Alb98a, Car91a, DP91e, DP91j, DP91i, Dur99, Hen91, HP92, Kaw93, Lio93b, Mai99, Mar93a, Mél96, Mül99, NV94, Nov92, Sto92k, Woy98c, vG92]. Moderate [Kaw92, Yur95a]. Modern [FP98g, Bor90c]. Modified [Bea90, DS92, Kar91]. modulaires [Lab90]. Modular [BFS98g, CP91g, Fre91e, JR92d, Man92, Pan91h, Ree90, Sri90, AB90b, CP91d, CP91e, CP91c, Fre91d, Fre91f, Har90c, Hun96d, Hun96h, Ino98d, Mer94, Oda90b, Pan91g, Pan91i, Pan91j, Ric98, Rut97f, Zim98d, vG92]. modulators [HYL91a]. module [Vas92]. Modules [Khu90, Man97d, Man97g, Xu96d, AHJLP90, Azu90, BL94b, Ber90a, Ber90c, BS97b, Bou97n, Bru90, Dau90, Deb92, Erd90g, Gek90, Jan90a, Kap91, LO98g, LO98j, LO98i, Man97b, Man97e, MS90c, RY90, SS93, Smi90, Ura93a, Wan90, Xi94h, Xi94j, Xi94m, YSA92, vdPS97c, ZD93]. Moduli [AR96e, Hun96g, LO98g, LO98h, MN91, Sil92, CHT92, DM96, Fuj91, Hit90, JP91b, Kob90, Kum97, MO93b, Sai91a, Tyu92]. modulo [Dũn92]. modulus [Bal90a]. molecules [SOA91]. Moment [HW91, Üst95h, Fur92]. moments [BVH94b, Le 94, Pus93]. momentumless [Puk90]. monad [Koc91]. Monge [Sch90q, Sch90t, Sch90u]. Monodromy [Sim91a, Bol92]. monogenic [Mit94f]. monoid [Gub90]. monoids [Ben91a]. monopole [Hor90]. monoreflectors [SM99f]. Monotone [Ari91, Phe93c, Phe93i, Alb98c, Alb98d, Egg91, GK90, Kit94b, Phe93g, Sim98d, Sim98i]. monotonic [Hof91]. Monotonicity [Ene95e, Sim98f, Kat90]. Montecatini [CR91, CP93b, DL97, FG90, FG96, FZ93, Gra94, Joh95g, Odi90, Rug96, TT96, Vis94, dBT91]. Montgomery [Bou91c]. Montreal [Pic90]. Moody [AFSTS92]. Moore [Bau91, Del91d]. Moore-homology [Del91d]. Morava [HKR92, Wür91]. Morayne [Ste90b]. Mordell [Top93, Bou98b, Bou98c, Hin98, Shi91b]. Morita [Bou97l, Bou97m, Lin98, LS90c, Xue92b, Xue92c, Xue92d, Xue92f]. Morley [Zie98]. morphism [Hub95p]. morphisms [BK90]. morphogenesis [Mai99]. Morse [FS90c, Mot91, Paz91]. morsifications [Dam91b]. Mosco [DZ93h]. Moser [Mil94i]. Motion [Neu96d, BW94a, BM92d, Don98, Eva97, FF96, Le 92, Le 94, Shi94, ST92a, War97, War99, Wer95, Zhe96]. Motions [BHS96i, DÉY93, Neu96a]. Motivation [Ebe99e, Sto92j]. Motivations [BP97i, Fri90a]. Motive [Har93a, Har93e]. Motives [Hub95q, Jan90d, Har93a, Hub95m, Jan90e, Sai91b]. Motte [Alb90]. mountain [FW91]. mouvement [Arn99, Ber93, Mey94]. mouvements [ÉY98]. Moving [Sim91j, Sim91i, Cle90]. Moyennes [JY93, KPR90]. Muller [Aub92]. multi [Gel92, JN96, Lem90a]. multi-échelles [Lem90a]. multi-type [JN96]. multi-valued [Gel92]. Multidimensional [Vai93c, Sap90, Vai93i]. multiéchelle [FM90]. Multifunctions [Sim98g, Mar90d, Sim98d, Sim98i]. multihomogeneous [Lie93a]. multiloop [LR90]. multiparameter [BF93i, JS90]. Multiple [Bar93i, Kle90, Mey93f, Mey95b, Oda90a, DJS90, Hla90, JL91, JS90, MWNPA92, Muk91, Wat99]. Multiple-point [Kle90]. multipler [Vas92]. multiples [Bor90a, Cle90]. multiplicatif [Fra95]. Multiplication [Col92h, Nil94e, Coa99, Rub99b, STV92]. Multiplications [Col92i]. Multiplicative [De 91, Sch91a, Egg91, Hol90a, KS92]. multiplicités [SG97]. multiplicities [Joh90, Kon90, SG97, Sim91b, Wal90c]. multiplicity [Chi97, Iwa92b, Xi94a]. multiplier [Bou91d, LL91]. Multipliers [Ouy91, BF93c, BF93d, Nik94a, Pis96f, Pis96i, Üst95a]. multisécantes [Le 90]. multisummability [Bal94h]. Multisummable [Bal94e, Bal94g]. Multivalued [Mol93h, Gel90, Mol93a]. Multivariate [Bre94h, Lor92f, Wes93]. multivoques [Cép95]. Mumford [Top93]. mutual [Kni93].

N [Tik90, YS92]. Nachbin [KR92d]. Nagell [Tur90b]. Näherungsformeln [Hla90]. Naissance [Mai93b]. Nankai [CDZ91, Cha91n]. Nash [Fab93b, Mil94i]. National [Kov90]. natural [Bor98d, Wat99]. naturelle [Arn99]. Navier [HMRS90, HMRS92, BM90d, Cha92, Fur90a, Fur90b, Gru92, Hey90, Kat90, Kaz90, Kos95b, Koz90b, OSK93, Pil90, Pro90b, Puk90, Ran90a, Ran92, Riv90, Sal92, Sch92b, Soc92, SW92, Val90c, Var92, Wie90, Wie92]. near [Bar93i, Fri91, Lau91, Maĭ93a, Pil99c]. Nearly [HS94, Kuk93a, CP91d, CP91e, Far96j]. nebentype [Wan94]. NEC [BEGG90e, BEGG90g]. necessary [vW90]. Necking [Mie91i]. negative [Băd90, Gau91, Ham91, SOA91, ZR92, ESH99]. Nehari [BGR91, GKW91]. nested [Kus93]. Netherlands [AK90]. nets [Eva92a]. Network [Sto92k, Sto92n]. Networks [Soa94d, Sto92e, Gra96, Soa94b]. Neumann [CK90, Cha91n, Gru92, Hen90e, Jaj91f, Pis96l, Sch90e, SM99u]. neurones [KS96]. Nevanlinna [Che90a, Che90b, Lan90b, LC90]. Newton [DL90, KY91, Neu97i]. Newtonian [Raj91]. nice [TT90]. Nijmegen [AK90]. Nilmanifolds [TO97b, Dek96f, Dek96g]. Nilpotent [DRSŻ91i, Ser92n, Dek96e, JL91, Kho93, Neu96e, Rut97a, mS91b]. Nilpotentes [RS94]. Nine [Spe93]. niveau [GC92]. no [TO97f]. Node [Sto92l, Sto92k]. Noether [FP98c]. Noetherian [Brö90a, Khu90, Tho90b]. Noise [Bae91, Oba94e, Sch93g, AB91, Ben91c, ELL99d, Heg90, Oba94f, Oba94g, Sch93e, Sch93h]. nombre [BW94b, Dur90c]. Nombres [Fou90, Wal90a]. Non [Bel94, BL92a, Boi96, Bra98f, CP91g, Ebe99g, EQ97, Mar90a, Mey93j, Mey95e, NZ99i, Pan91h, Pan91i, Par91a, Pis96h, Rut97n, Rut97o, Wil95, Zuo99c, AH90, Bal94d, BÉK^{+}98, BK91, Bia95a, BC92b, BL92d, Bon91, Bor90c, Bro90b, CG91, Daz90, Del90a, Del91g, gDH91, Gau91, Gro90a, HT90b, Ken93, Kos95c, Mat93, Met91l, Mor93b, Nab90, NSS92, Neu97e, Neu97p, Pet92, Pot95c, Raj91, Sch91k, Ser94d, Shu98b, Sri90, Ste90a, Suf90, Zvy90a, ZR92, dM99, CP91b, JLP90, Pan91f, Pan91g]. Non- [Rut97n, Rut97o]. Non-abelian [Zuo99c, Bro90b, CG91]. Non-Archimedean [CP91g, Pan91h, Pan91i, CP91b, Pan91f, Pan91g]. non-brownien [dM99]. non-cancellation [BL92d]. non-chaotic [BC92b]. non-commutatif [Bia95a]. Non-Commutative [Mey93j, Mey95e, NSS92, JLP90]. Non-compact [NZ99i]. non-complete [BK91]. non-convex [Mat93]. non-convolution [gDH91]. non-defining [Sri90]. non-différentiable [HT90b]. non-divergence [Ken93]. non-existence [Met91l, Pot95c]. non-identical [Mor93b]. non-inner [Neu97e, Neu97p]. non-isolated [Sch91k]. Non-linear [EQ97, Wil95, Bal94d, Bor90c, Zvy90a]. Non-local [BL92a, Bra98f]. non-negative [Gau91, ZR92]. non-Newtonian [Raj91]. non-oscillatory [Shu98b]. non-recursive [Nab90]. non-rigid [Pet92]. non-symétrique [Gro90a]. Non-Uniqueness [Ebe99g]. Non-unitarizable [Pis96h]. non-vanishing [Suf90]. nonadditive [Wac92]. nonarchimedean [AMM90, Her90a]. Nonautonomous [Mie91j]. Noncommutative [Con92, Jaj91h, Jaj90, Jaj91a, Ser94d]. Noncompact [Pil90, Wan90]. nonconservative [Col92f]. Nonconvex [Cel90]. nondeformability [Siu91]. nondemolition [Bel90a]. nondeterministic [Coo90]. Nonequilibrium [CP93a, CP93b]. Nonexpansive [RB92, Wit92]. nongeneric [FS91]. nonholonomic [Ger90]. nonhomogeneous [Gol91b]. Nonlinear [Bia90d, GV91h, Liu96b, Neu97c, Qua98, Rei91d, Rug96, TT96, Vai93d, Bax91, Bia90c, BT93, CP91a, Eva97, Fit93, Fri90b, GP91a, Hel91, KPV90, Ken94, Mat90a, Maw93, Mis92, OT90, Ôta93, Sob90, Tad98, Tib90a, Wah95e, Woy98e, Yau92, ZD92]. nonlinearities [Kos95c]. nonlinearity [HRS91]. nonlinearly [ITO93b]. Nonlocal [Sob90]. Nonmeasurable [Bek91f]. Nonmodular [JR92e]. nonnegative [BD94]. nonrecursive [DJS90]. nonreflexive [KF93]. nonsingular [Bor92]. Nonsmooth [DN99d, Roc90, Tib90c]. Nonstandard [Wic91d, Wic91e]. nonstationary [Mog92, Ran90a]. nontangential [Iva93]. nonzero [Kon90]. Norm [BL92c, KZ92c, Dug92, Hu91, Lou91a, Pan91a, XL98e]. Norm-endomorphisms [BL92c]. Normal [BEGG90g, BOERT96a, Bre94a, BK92c, Con94, DRSŻ91h, Oli90, PsT90, Sak90]. normale [Mou90]. normaliser [Sho92q]. normality [BVH94b]. Normalization [Bou91e]. Normalized [Lom92, MR98, Mol93f]. Normed [Bal91, Ina90b, RB92]. norms [KZ92d]. Northern [BLM^{+}90]. Notation [CL99c]. notational [Dek96h]. Notations [Mie91k, DRSŻ91e, Wei94d]. Note [Lup90, AB90a, Coh90, Dug92, Gao97, Heg92, Kel91b, Led90, Tur90b]. Notes [BMMN98j, Dow90, Koe99f, Nog91, Pea91, Yan91a, Neu97l]. notions [Hub95c, Küh95d, Wic91g]. nouvelle [Dav90]. NSA [Ark93]. Nualart [Léa96]. Nuclear [Ban91d]. Null [Far96d, KO91]. nullité [ÉY98]. nullity [Far96e]. Nullstellensätze [LW90a]. Number [NF90, Sho92f, vdG93, BLR90, BL93b, BLP93, Cha92, Cou93, DK92, Gre92e, HYL91a, Kai97g, Kin94, KB96g, LL93, Len93, LLMP93, Met91k, Now90, Sch91k, Spr93b, Tsf92, Wah95g, Yap96h, Yap96i, HT90a]. Number-Theoretic [HT90a]. numbers [BMMN98l, Erd90f, Fur92, Göt94b, Hen90e, Jaj91e, LO98m, Lun92, Mas95i, Nie90, RS90c, Roh90, Spr93f, Tro91]. Numerical [CL99d, CL99e, HMRS90, HMRS92, Hum90, Pil99a, Pyt99e, Qua98, Rei91d, Riv90, SOA91, Ver94, BMMP91, CV93, Col92f, Col92h, GV91i, PS90, Pom98g, Ran90a, Tal96]. Numerically [PSW92]. numérique [Mey90].

Oberwolfach [ASMS90, ACE91, HLN91, HMRS90, HMRS92]. objects [RG96d]. obliquity [KLGP97a]. Observable [Gro97f]. obstacle [KZ91]. Obstruction [Ura93d, And91, Cle90, Haj91]. obstructions [HALM90]. obtain [CG97e]. Occupation [AS98f]. occurring [DW91]. October [GT90, KRW92, NF90]. Odd [Fre94, Coh90, KW94, Kim94]. ODE [Bal94d, HRS91]. Ogawa [MM97]. Ogus [Hub95d]. Ohio [JLP90]. old [Bau91, dF90]. Ol’shanskii [HN93e]. Omega [Koe99g, Las98, Koe99e]. Omega-stable [Las98]. Ondelettes [Lem90c, Mey90, AAG90, FM90, Lem90a, Lem90b]. One [AS98d, BC92c, Wei94c, AB90a, Bor92, BEGG90a, Fit93, Hey93, HS90b, Joh95a, Koc90b, Lan90b, Miz92, Nis96e, OL90e, Pal90, Pet92, Val90c, Wah95h, Xue90]. one-dimensional [Hey93, HS90b, Miz92, Val90c, Wah95h]. One-parameter [Wei94c]. one-relator [Joh95a]. ones [Shk93]. Onesided [Zim98c]. only [Met91f]. Onsager [CMN95]. onto [OL90e]. Open [Alb98h, BCC92a, Bar90b, Hey90, Sim98h, BMMN98a, Bia90b, Bru92b, CM97g, FP98b, Heg90, KLGP97k, Lan91f, Mar90d, Zor92]. opérateur [LW90c]. opérateurs [Fon98, Sch94u]. opération [HL92]. Operational [Kom93b, NSS92]. Operations [Hub95s, Kai97d, Lom92]. Operator [Kar97f, Oba94b, Ric99e, Tam99b, YS92, CS92a, CP91c, Gau91, GK94, Hen90e, HSS91, Hu95, Kar97c, Kar97e, Kar97d, KS92, KM97a, KM97d, KM97e, KM97f, KM97i, KZ92a, LR90, Nag90b, Pan91j, Ric99c, Ros94, Tam99c, Tam99g, Üst95c, Wah95i, XL98i, Zip91, deL94s, deL94t]. Operators [DN99d, Ebe99g, ELL99h, GJ91, GE98c, Ino98f, Lan91g, Lie93d, Phe93c, Vät97f, vN92c, Aki90, AS92, Bal94a, Bas90, BV93, BL92b, Bor90c, Bro94, CMZ91, Con94, CP91d, Dav91a, DN99a, DN99c, ELL99e, Fuh91, GV91k, GE98a, GKKL91, Han91, HWW93g, Har90c, HMN91j, Hör91b, Iwa90, JL91, Kaj91, KR92h, Kit93, KO93, Kom93b, Lam92, LS91, Lie93i, LN91, Mal92, McI90, MT91a, Mil94j, Mit94a, MR91b, Nie91b, Nie91g, Nie91i, Nil94a, Pan91a, Pee94, Phe93g, Phe93i, Räb91, Ric99b, Röc99, RS91b, Sar92, Sch90f, Sen90, Sim98b, Tai91b, Van91, Wan91b, Wan94, Zvy90b, ZD92]. optical [DW91, JS91a]. Optimal [Cvi97, Mil94f, Mil94i, Pyt99e, Tib90c, Bar97, DZ93d, DZ93e, DZ93i, Foi91, Kwa91, LS93, Met91m, Pea91, Pyt99f, Vel90]. Optimization [DZ93g, Yuk98g, Phe93a, Roc90, Ver90]. optional [SY98]. orbifold [Kaw93]. orbit [BEGG90e]. orbitals [BMMN98o]. orbites [Don90, Tur90a]. orbits [AV91, BC92f, BC92k, EGI94, MF94a, Mey99b, Mey99h, Mey99k]. Order [BHS96i, Gra91b, KM97h, XL98d, Bel90b, Clo90, DX93, ELLV90, Eyr98j, GH90, Göt94e, Gra90a, HRS91, Hu96, Isa92h, Kos95a, Kos95c, Le 91, Lie93e, Lum91, New90, Pyt99b, Pyt99g, Räb91, Raj96, Sch90f, Sch94o, Vai93b, Wah95g, Wes93]. Order-enriched [Gra91b]. Ordered [HN93i, KR92e, SM99q, HN93a, SM99j, SM99o, SM99t, SM99s]. ordering [Nie91k]. Orderings [Mar96e, Mar96i, Mar96h]. Ordinary [FZ93, Kru97c, Sim91k, Car91b, HS91, KM97a]. ordres [GC92]. Ore [MZ90, TT90]. orientations [Deg92]. Oriented [ZR92]. origin [JL93b]. original [GV91c]. Origins [Spr93g]. Ornstein [BW94a, HW97, QH95, Üst95c]. Orsay [Lem90c]. Orthogonal [Jac90, Neu97j, Nev94, Ger93, Jaj91e, Khr93, Rak93]. Orthogonalité [Lép92, Val94]. orthogonaux [Maz97]. Orthomorphism [Eva92g]. orthomorphisms [Eva92b]. orthonormal [App95]. Oscillation [AW97]. oscillations [FP91]. Oscillator [Mey93h, Mey95c, vW96]. oscillators [BP91]. Oscillatory [TPDZ91, Fuj93, Hu91, Shu98b]. osculating [PsT90]. Oseen [Gal92]. Oseledets [Le 91]. Other [Mas95j, Neu96c, Xue92e, Acc94h, Far96h, FP98e, FP98a, GV91j, Man97e, Sch94n]. outer [Dix97c]. Outline [Lax96]. Outlines [Vai93e]. output [Bel90a]. Overstability [Wal91a]. Overview [AH92c, BFS98h, NZ99f, EGI94, LPLG^{+}92].

P [Mal91, Mey93c]. P. [Ste90a]. packing [Pro90a]. packings [Kun92]. Pairings [MS94b]. Paley [LP91]. paper [Bur90, Hit91]. Papers [Ste90b, Kah90a]. Parabolic [Gru91a, Kos95e, Zab99, Bel90b, DH90, DU92, Fab93b, Fla91, Kos95d, Mil94k, Mur93, Ôta93, Sob90, Ver94, Woy98i, Zvy90b, deL94q]. parabolicity [XL98b]. parallel [BCMS90, Lum91, SOA91]. parallelization [KO90a]. Parameter [Tib90c, Wei94g, AV91, AS92, BV90, DH90, Fit93, IK99a, Pet91, Wei94c, Żoł90]. Parameters [Khr93, Pot95b, PR97f]. Parametric [Ste99b, Roc90]. parametrized [IR92]. paraproduct [gDH91]. Pardoux [Léa96]. Pareto [DX93]. Paris [LW90b, Lem90c]. Paris-Sud [Lem90c]. parity [Fit93]. Parseval [JL93b]. Part [Edi93, Fab93a, MM91b, Pro98b, Pro98c, Pro98d, RS91a, Tai91i, Tai91j, Aik96, Den92, Ess96, Lio93a, Pla90, HN94a, HN94b, Kat93a, Rei97b, Rei97c, Wil97e, Wil97f, Wil97d]. Parthasarathy [Mey93b]. Partial [FIK90, Lie93g, Mar92, TT96, Bia90c, Coo90, CJ90, Koc91, Liu96b, MT90, Pom98d, Rau95b, Tal96, VC90, Woj91]. Partially [SM99q, SM99j]. participants [Zar91b]. Particle [CP93b, DP91j, DP91i, Gol91b, CP93a, Dur95, Mél96, Pul96]. particles [DP91d]. particules [KS96]. Partielles [Par91a, Bon91, CJ90]. Parties [CDP90k]. partite [Yap96c]. Partition [Sto92m, Fra95, Sto92l]. partitions [Kra90]. parts [BCG90, LN97]. passage [ERHG91]. Passive [Woy98g]. Patching [Man97e]. Path [FK93e, BI92, Blü90, FF97, Fuj93, Git98, Ich90, LN97]. paths [Bur93, ELL99b]. Pathwise [BMO98, BP90a]. patterns [Mau90]. PDE [Da 99, Fre92]. PDEs [Fla91, Ken94]. Peano [Kar91]. Peculiar [Šip92]. PEL [Hun96g]. [AE90]. 89 [AYM90]. homology [Rut97e]. Marseille [LS90a]. pencils [Col90b]. pendulum [FP91]. Percolation [Gri97]. perfect [Azu90, Coh92, Khu90, Pat90]. performance [PS90]. Period [BF93g, Har90c, Bae91, BF93i, Oda90b, Pet92, PR97a, Sch90e]. Period- [BF93g, BF93i]. period-doubling [Bae91]. périodes [Fra90]. Periodic [AV91, BC92f, BF93h, BHS96i, Cru90, GS90b, Kar97f, Mey99i, Rei91d, Bar93i, BC92k, BF93e, CMZ91, Coh90, DH90, DM90a, EGI94, Far92, Far96a, HMN91a, HMN91g, Kar97d, Mey99m, OT90, Sjö91]. periodic-parabolic [DH90]. Periodicity [DS90a, DS90c, KLGP97i, DS90d, XL98a]. périodiques [Aud90]. periods [Blo92]. permitting [DJS90]. Permutation [BMMN98j, BH90a, Sho92s, Pra90a, Sho92r]. Permutations [Mül91, Hig90]. Perron [Yee90]. persistance [GRCW90]. Persistence [PR97c, Wal91b]. perspectives [BP97j]. Perturbation [Bir94, Dob96, Kar97c, Kar97e, Kar97d, Kar97f, vN92h, BCDS92, Jon95]. Perturbations [Pil94d, deL94m, DZ93e, LQ95f, RS94]. Perturbed [Bor98c, DWY98, Sim91a, Don98, HMN91i, ITO93b, Phe93a, Wer95]. Pesin [Thi91]. Peterson [TY90]. Petites [RS94, Ber90b]. petits [Fou90]. Petrovsky [deL94n]. Pettis [Räb91]. Petviashvili [AD90]. Pezzo [Fuj90b]. Pfeffer [Ste90b]. PFTQs [LO98g, LO98i]. Phase [DP91k, HMN91h, Jen90, Kaj91, Ken94, Mül99, Sim91g, TPDZ91, Vis94]. phases [SOA91]. phenomenon [SOA91, Tal91]. Philippe [Dwo90]. philosophy [Tho92]. photons [Fri91]. Physical [Fri90a, Ino98c]. Physics [FG90, GS93, Col92i, Col92h, Gra96, Jon92]. physiologically [Die99]. Picard [GH90, Ric92, vG92, vdPS97i]. picture [Lan91d]. Piecewise [Shi90, Vai93f, Col92g, CW91b, Hof91]. pinched [Ham91]. Pisa [Vil90]. plan [BP97i, Ber93, Mou90]. Planar [DRSŻ91i, FR90c, Sch90k, AV91, BW94a, Col90b, DRSŻ91a, GS90a, Le 92, Le 94, MS90b, Mou92a, MN90, Rus90, Shi94, Wer94, Yap96i]. plane [Ant92, BW94b, Drm90, GT91, GCR90, Gor91, Kra90, MD90, Pol92, Sch91k, Shu92, SOA91, Sto92g, Wah95b, Wat99]. Planes [Kna95g, Kna95i, Kna95e, Kna95f, Met91d, Met91q, Met91m, Pot95f]. Plasma [TBR91]. plats [Bea92]. plongée [Mul90c]. pluriadjoint [Oli90]. pluriharmonic [OU91]. plusieurs [Gra90b]. poids [DL90]. Poincaré [LW90b, Ben92, NZ99j, Rud92, Yam90]. Point [Sch90d, AB91, AE90, Bar90c, Bar93c, BK92c, Bru92a, BM92d, Fur90a, How90, Kle90, Kuč90, Nus93b, PR97b, Sam91, TO97d, Wit92]. point/ [AE90]. Points [Met91n, Met91o, vdG93, Acc94f, Acc94h, AS98b, Ban92, Ber93, BC92l, Bor90b, Bor92, BF93c, BF93e, BF93h, BF93g, BF93i, Bur93, Che91, DG90, DL96c, EP90, ERHG91, Gel90, Kin94, Met91g, MN90, PR97e, Rib99, STV92, Tsf92, Vig90]. Poisson [Fri90a, KS90a, LP90b, NV90, Van96f, VK90]. Poissonxo [Rog90]. Pol [CD91]. Poland [JOP91]. polar [CML90, Kho97]. polarisées [Deb92]. polarized [Oli90]. poles [Ika90]. poloidal [SvW92]. Pólya [Kah92]. polycycles [Mou90]. polyèdre [CDP90j]. Polyèdres [DL90]. polygogarithm [Wil97d]. Polygogarithmic [Wil97d]. polyharmonic [Kar97c, Kar97e]. Polyhedral [Küh95f, Küh95g]. polyhedrons [Zai90]. polylogarithm [Wil97f]. Polylogarithmic [Wil97e, Wil97f]. Polylogarithms [Wil97g]. polymère [Le 97]. polymorphic [Sce90b]. polynôme [Bor90a, Dur90d]. Polynômes [Bor90a, LW90b, Mig90, Phi90, Dur90a, Dur90b, Lan90c, Maz97, Sch90i, SN90]. Polynomial [Dek96b, Nar95b, AV91, ASHY90, CV93, Fre91h, FP98a, GU92, Kwa91, Maĭ93a, Man97b, Nar95a, RS90d, Son90, TPDZ91, Zai90]. Polynomially [NR90]. Polynomials [FP98h, Ger93, LW90b, Bou91b, Clu92, FP98f, FP98j, Jac90, Khr93, MO93h, Nar95c, Nev94, NM92, Pra91, Pri93, Rak93, Saf90, Spr93h, UT92, Woj91, Yan93]. Polytopes [RG96f, RG96g]. Pontryagin [Ban91e]. ponts [JY90]. population [DGT91, Jag93]. populations [Die99, JN96]. porings [SM99d, SM99e, SM99g]. Posed [DZ93g, Fur90a]. posedness [DZ93a, DZ93b, DZ93c, DZ93d, DZ93f, DZ93j, DZ93i, DZ93h, deL94s, deL94u]. positif [Mey93c]. positifs [Bor90a]. position [EP90]. Positive [CM90, Iwa92b, Üst95i, Dam92, Dem96, HRS91, Haj91, Hru98b, Nag90b, Not90, Pet96, Ros91, Sch93d, Sim98b, vN92b, vdPS97e]. positively [JcC91]. positivity [Kni91, Koe99b]. Post [CM90]. Postnikov [Rut97n]. Potential [Aik96, AE96, Ess96, Kar97f, Ken93, Soa94d, Bor98d, Jen90, Kar97d, OT90, Suz90]. Potentialen [Hla90]. potentials [BM90c, GV91c, Hla90, Soa94a, Tam90, Van96d]. Potentiel [Anc90, Del90a, Del91g, Mic92]. pour [Bea92, Ber90b, Cap98, Eis97, ESH99, Fon98, Fou95, Hen90b, Hu93, Jac91, Jou98, Le 90, Mey91b, Mic97, Thi91]. Power [Bal94e, IK99g, AFF90, Bal94b, Bal94g, Bal94i, Gre90, Har90a, New90, WY92]. powers [CRW90, Cas92, CS92c, Her90b, IK99f, IK99i, Kur90]. Poznań [JOP91]. PR [CDZ91]. pre [Mic93]. pre-wavelets [Mic93]. Precategories [Jan91]. précédent [Arn99]. preconditioned [BCMS90, Gus90, Not90, van90, AK90]. Preconditioning [ELLV90, WS90]. predecessor [Wir98e]. Predictable [Ran90b, DS95, Ove95]. preduals [CK90]. Preframe [JV91]. Préliminaires [Man97f]. Preliminaries [Ban91f, CM97f, Dek96h, Ene95f, Isa92i, JR92f, Kit94d, LQ95e, Moo96c, Pot95e, Rut97p, Üst95j, Vai93g, Wei94d, Wic91f, Zuo99d, Bre94g, Heb96b, Man97e, Sim98c, Man97f]. Preliminary [BEGG90h, Pil94a, Wah95g]. premier [FT90]. premiers [Fou90]. prenant [Gra90b]. Preordered [SM99j, SM99m]. preparation [IG90]. preparatory [AS98a, Nil94c]. Prerequisites [GHS97f, LO98j, Oba94c, LO98n, Zie96b]. Preschemes [Mum99e]. prescribed [BEGG90f, Ste99b]. presence [Heb96e, van90]. present [JV91, Rud92]. presentation [CP93c, HM90, Wah95g]. presentations [Fin90, JV91]. presented [CP93d, Rut97a]. preservation [Far96g]. preserves [Far96j]. preserving [HK92, Rut90]. presque [AW97, Fou90]. preuve [Ber93, Cad97]. prévisible [AJKY93]. pricing [EQ97]. Priddy [Mil90b]. Primality [AH92d]. prime [Kur90, New90, Sch94o, Sho92q, Wri90]. prime-power [New90]. primer [Cra97]. Primitive [Sho92s, TT90, Pra90a, Sho92r, Sho92t]. primitives [Skv90]. Primitivity [BMMN98k]. Principal [Xi94j, Bel90b, Isa92i, Ver91]. principalement [Deb92]. principe [Cad97, Del91g]. Principle [Nag93, Eke90, Gro97g, Kum92, Kus91, Mur93, She93]. Principles [Kna95i, Mar93b, Phe93j, Wac92]. priori [Ôta93, Sch90q, Sch90r, Tai91f]. Priority [LL90, Dow90]. prix [Ans94]. Pro [KLGP97g, HP92, KLGP97o]. Pro- [KLGP97g, KLGP97o]. pro-crossed [HP92]. Probabilistic [Ebe99f, Ike90, Kah94, Tal96, Bas90, KP90, TT96]. Probabilists [Mey93k, Mey95f]. probabilité [EM91]. Probabilités [AYM90, AYM91, AYM92, AYM93, AYM94, AEMY95, AYE96, AYE97, AYÉL98, AÉLY99, BGM94, Ber95, Ber96, Ber98b, GGSC97, Hen90a, Hen92, Hen93, BPSH91]. Probabilities [Neu96d, BSÉ99, Haz93, Per91, Sch99, Tay92]. Probability [AvW90, BGM94, Ber95, Ber96, Ber98b, Ber99a, Bre94d, BPSH91, GGSC97, Mey93j, Mey93k, Mey95f, Mol93i, Neu96e, Per99, Yuk98g, Bar90a, BP91, Bre94b, DN99c, Itō93a, KTV92, Mey95e]. Probailités [Ber99a]. Problem [Arb99, Cha91f, HN94a, HN94b, KN91, Kos95e, Mey99i, OL90f, XL98d, BM90a, BGR91, Ble91, Bol92, BM92b, Bor90b, Cha91d, Cha91e, Cha91b, Cha91n, Cha92, CM90, DH90, Del92h, Den92, Fre90, Fur92, GP91a, Gal92, Heg90, Hon90, Isa92d, Isa92h, Isk91, Iva93, Kaj91, KPV90, Kni91, LP90a, Lou91b, MS90b, McI90, Mey99c, Mey99j, Mie91l, Mog92, Muk91, Nar90, Nar93, NZ99b, Nis91, OL90b, OL90e, Oka90, Pis96j, Pro90a, Puk90, RS90a, Rod94, Rud92, Sal92, Sim98k, Sim98l, Sim91n, Soc92, ST92a, Spr93b, Wie90, YR90, Zvy90a, deL94d, vdPS97g]. problème [Bel99, GC92]. Problems [Alb98g, Bar93j, BP97j, Bra98a, CG97g, CP93b, DZ93g, HS99, Isa92k, KZ93, Mie91c, OL90d, Pis96k, Pyt99e, RG96e, Sch90d, Sch95d, Ste90c, Tai91d, Yuk98g, Alb98e, Alb98h, BCC92a, Bar90b, Bar93a, BL94g, Bou91d, BMMP91, BF93a, BF93b, BF93k, Bro90b, CM97g, CP93a, CMN95, Coc98, Dri92, DRSŻ91d, Ebe99e, Egg91, Elf91, ELLV90, Fis90, For92, FP98b, Fur90a, Gel90, Gem90, GKW91, Gro96, Gru91a, Gru92, Heb96a, Hey90, IK99h, Isa92i, Isa92j, Kah94, Kah90c, Kaz90, Khi92, KLGP97k, Lan91f, Lau91, Lie93j, MT90, Mar93b, Maw93, Mie91b, Mus92, Neu97e, Par93, Pil90, Saf90, Sch90e, Sch95b, Sil92, Sim98h, Sin91, Sko92, Suf90, Tai91a, Tai91c, Tib90b, Vai93i, Ver94, Wah95e, Wah95g, Wat90, Yuk98c, Zip91, Zor92, deL94q]. procedure [Pyt99f]. procedures [AL90b]. Proceedings [AvW90, ACC92, ASMS90, ACE91, AK90, BBD90, BCC92b, Ben91b, BDF91, BS90c, BLM^{+}90, BM91, CPR91, CDZ91, CMR92, FPSW91, FR90c, FIK90, GT90, HLN91, HMRS90, HMRS92, HPSS92, JOP91, JLP90, KZ93, Kom93a, KÜ90b, KRW92, Kul92, LS90a, LW90b, Lat90, Mim90, NF90, NO91, OR91, Pic90, RSSV90, SBL90, ST92c, TBR91, Vil90, XD90, YS92]. process [App95, BW94a, DP91c, DP91f, DS95, KT93, Kos95g, NV90, PR97g, Shi96, Xue90]. Processes [AS98d, CF92, Git98, Tai91d, AS98c, AS98h, BI92, BK92b, Bax92, Bel90a, Blü90, Bre94e, CL96, CML90, Chi97, Daw93, DWY98, Fri90a, Gin97a, Hol90a, Kry95, Kui92, Kus93, Lei91, MR94, MWNPA92, MSS94, Nag93, Par91c, Ran90b, Sch91l, Son97, Wac92, War97, Xia92]. Processus [Del90b, Mai93b, AY90, AJK^{+}96, Ber90b, CTR96, Eis97, Fou95, GRCW90, Mai94, dM99]. Product [DN99b, vdP93, CS92a, DS90g, Dũn92, Eyr98j, Gol91a, Han91, HA90a, Hub95a, LS91, Lor92k, iM90b, Neu97d, Neu97e, Neu97p, Vät97c, Wah95m]. Production [Del90b]. Products [DS90d, JL93a, JLG94, Pus96l, BMMN98q, Dar91, Far96h, Gol90, HM91, Hul90, JL93c, JL94, Kow92, Lan96g, Lan96j, Leu91, Par93, RS91b, Rut97t, Sch91b, Tay92, Van96e, vN92i]. Produits [Mor90b, Mey93a]. Professor [Kom93a]. profinis [Ser94a]. Program [Kov90, CDZ91]. Programming [Sac90, DZ93c, Mac91, OL90e]. programs [Pla90]. progress [Fal92, Tik90]. progression [Gra90b]. progressions [Now90]. Projecting [Mil90a]. Projection [Rai96, CP91c, DS95, HYL91a, Isa92f, Pan91j, Ran92]. Projections [OL90d, Ric99e, Neu97j, OL90c, OL90e, OL90f, Ric99c, Ric99d, Ric99f, Wah95c, Wah95d]. Projective [ABW92, HS94, Hun96h, Man97g, Pot95f, vG92, Aub92, Bal90b, BS90a, Göt94a, HFS94g, Hul90, Ion90, Kim94, Kna95h, Met91d, Met91q, Met91m, Siu91, Van91, WJ90, iY90, Zak91]. projectivity [Bou97l]. Prolog [Sac90]. Proof [AH92e, AH92f, AH92g, AMS90, Bou98c, Edi93, Fab93a, Hru98b, Kuk93b, Tai91i, Tai91j, Tai91g, Tai91h, AH92c, Apt93, Bom91, Bou98b, HA90b, Kit93, Kry95, MR90, MSS94, MT99, MO93f, Muk90, Pra99, Ste90a]. Proofs [CL99f, LO98k, Pla90]. Propagation [Jou98, Lie93h, Rug96, Col92g, Sou97b, Szn91, Woy98e]. propagators [CP91a, Yaj90]. Proper [GU92]. Properness [ZD92, Zvy90b]. Properties [CG97a, CG97c, CG97d, CG97h, DS90a, KM97g, Mum99f, Abr96a, Abr96c, BI92, Blü90, Bui92e, CS92a, Coh95, DK93a, DRSŻ91f, Git98, GV91k, Gro93, HWW93d, HS93, Hel92, Jan93, JL93c, KLGP97d, Kos95f, KM97i, KJ90, Le 92, Led92, Lor92a, Mas95c, Met91a, Met91b, Miz92, Pil94b, Pus96k, Rut97g, Sch90b, Shk93, Skv90, SV90, Sre92, TO97d, Vät97a, Vät97d, Wah95d, Wil97e, Zak91]. property [Bal91, Bor92, Ove95, Sch90h, SDS90, Win90, WJ90, Yaj90]. proposition [AH92e, AH92f, AH92g]. propositions [AH92h, LO98k]. propriété [ARY96, Fou95]. Propriétés [Arn93, Fer91]. prospaces [HP92]. prospects [Yau92, NO91]. proximinality [RB92]. Prym [FP98c]. pseudo [AP90, Gru91a, Tai91b]. pseudo-differential [Gru91a, Tai91b]. pseudo-Riemannian [AP90]. Pseudodifferential [Lie93i]. pseudoelasticity [HMS94]. pseudoequivalence [DS90d]. Pseudohomogeneous [Pom98f]. pseudoisotopy [FJ91]. pseudomanifolds [Küh95e]. pseudonanalytic [Iva93]. Pseudoorthogonality [CS92c]. Pseudorandom [Nie90]. PU [KJ90]. PU-integral [KJ90]. pures [ARY96].

QFT [BL92a]. Quadratic [Hör91b, Sch93d, ELY97, Now90, Sch90j]. quadrature [LS93, Mey90]. Quadrics [Rei90b]. Quadrilateral [Ura93b, Ura93e]. quadruples [DL96b]. qualitative [Kum92, Sam91]. Quand [Mat98]. quantile [DN99a]. Quantitative [KR92f]. Quantum [AFL90, AH90, App90, BFS98c, DP93, Eyr98f, Eyr98g, Fag90, FG96, Fri90a, FK93g, GGS92, Ger92, GVV90, Gra94, KV92, McC90b, Mey93k, Mey95f, Taf92, ZD93, AFSTS92, Bar90a, Bel90a, BP97a, BL92a, BFS98b, BP91, BL92b, Ble91, Bor98a, CGST92, CF95, Col92b, Dri92, EGI94, Eyr98h, Eyr98i, FF96, Fri90b, FK93a, FK93b, FK93d, FK93e, FK93f, Ge92, Heg90, HS90b, Ich90, IG90, Lyu92, Maj92, Sch91l, Sin91, UT92, AvW90, Kul92]. quantum-classical [Bor98a]. quartic [Hun96c, OT90, RS90d]. Quasi [Ble91, BHS96i, CDP90l, Mou92b, Ant92, CTR96, CDP90k, Fuk93, Lan96d, Lan96a, McG92, Sta92, iY90]. Quasi-classical [Ble91]. quasi-convexes [CDP90k]. quasi-diffusions [McG92]. Quasi-everywhere [Mou92b]. Quasi-géodésiques [CDP90l]. quasi-Hopf [Sta92]. quasi-isométries [CDP90l]. Quasi-Periodic [BHS96i]. quasi-split [Lan96d, Lan96a]. quasi-stationnaires [CTR96]. quasi-steady [Ant92]. Quasiconformal [KR90, AVV92, Geh92, Zor92, Vuo92]. quasifibrations [May90]. quasihomogeneity [Zai90]. Quasilinear [Kos95e, Sch90u, Ama91, Kat93b, Kos95c, Kos95f, Sch90s]. quasimartingales [Pra92]. Quasiplasticity [HMS94]. quasiregular [Iwa92a, Mar92, Pot95f, Ric92, Sre92]. quasitriangulaires [KS90a]. quaternion [Erd90b]. quaternionic [Rei97d]. quelconque [Jac91]. Quelques [AS90c, Dur90a, Dur90b, Eme91, Fla90, Le 90, ABÉH95, Del91g]. Questions [IK99h, BMMN98a, Lio93b, Lou91b, Sjö91, Ver90]. queueing [Mor93b]. quick [CP93g]. Quillen [Pat90]. quivers [Erd90d, Erd90g]. Quotient [Sch94o, SM99k, Xi94k, Chr91, Nar93, Ste91]. Quotient-closed [SM99k]. Quotients [Hun96e, CW91a, Hun96b, LO98c, Wah95k].

R [Rut97k]. R. [BM90a]. Rôle [Nov92]. Rademacher [JS91b]. Radial [GKKL91, HYL91a]. radiation [KY93]. radical [BH90a, HN93h]. radii [NV91]. Radon [Hel98, OK91, TPZ98, Bom91, Gra91a, HYL91a, Maa91, Opo91]. Ralph [Mul90b]. Ramanathan [Tyu91]. Ramanujan [IK90, Tur90b]. Ramer [Üst95f]. Ramsey [McC99a, McC99b, McC99c, McC99e]. Random [AC91, Arn95b, Gem90, Lan91g, LQ95f, LQ95g, Mol93g, BD94, Cra91, Dar91, Ger93, Gol91a, Hey93, JPA97, Jan93, Kif91, KK93, LQ95b, LQ95h, Mol93b, Mol93d, Mol93e, Mol93f, Mol93h, Mol93i, Mol93j, Mol94, Spe93, Yan93, Yur95d]. Randomness [Kuč90]. Randon [Ber98a]. range [DP91f, Kit93, YR90]. Ranges [Ric99f]. Rank [Bur74, Maj92, Bur90, DR91, Lus90, Pet92, Zie98, FT93]. Rank-Sum [FT93]. rapport [AY90]. rare [Cat99]. rarefied [SOA91]. Rate [Wer94, Alb98j, BC95, Bjö97]. Rates [Yuk98h, Bre94e, LS93, MR94, Yuk98l]. Rational [BMMN98l, KMM92, Tho90b, BGR91, BF93c, DU92, LS93, PsT90, RS93, She93, Sta93, Tho90a]. rationality [Isk91]. rationnelles [Dum91, Kro91, Sil90]. ray [Pal91, War97]. Rayleigh [Nag90a]. rays [Muk91, Wit92]. Razmadze [Ina90a]. rc [HFS94d]. rc-Filtered [HFS94d]. reach [Shi96]. reachable [Phà91]. Reaction [Had99, Ama91, DP91i, FMSW91]. reaction-diffusion [Ama91, DP91i]. Real [AP94d, BCG90, Ene95g, FS90d, Gin98, Koe99h, Lyu92, Mar96i, Nie91c, Sch92a, Sch94p, Sch94r, Sch94q, SM99m, Wei94e, AK92, AB90a, BLR90, BFS98d, BK90, Deg92, Ebe98, FR92, GU92, GC90, GCR90, Jaw91, KO90b, LL91, Mar96a, Mar96g, Mat90b, Mil94e, Nab90, Nie91h, Par92, Phe93d, Rob90, RS90e, Sch94k, SM99l, SM99s, Shu92, Sil92, Ver90, Win95c, iY90, CMR92, GT90]. Realisation [Par91c, JS92, Kin94]. Realization [Hub95q, RG96f, Cri90, GCR90, Hub95g, Hub95o, Hub95r, Mar96j, Moe95f, Rud92]. Realizations [Wil97g]. realm [Van96g]. reals [Bek91f]. rearrangement [Hit91]. reconstruction [Cen91, Gra91a, HO91, Mad91]. reconstructions [KZ91]. Recuit [Mic92]. recurrence [Ger93, Jac90]. recurring [Gra90a]. Recursion [ASMS90]. Recursive [Kum90, AE90, Hig90, Nab90]. recursively [FS90a]. Red [Mum99a, Mum99g]. Reduced [SM99n, MT99, Ros90, SM99d, SM99e]. Reducibility [Spr93h, Bog90, Coo90]. reducible [Bol92, IK99e]. Reduction [AH92h, Lor92g, Sch94b, Hru98b, Mey99l, Tam99d, WS90]. reductions [ASHY90]. Ree [Ped92]. Reed [Aub92]. réelles [Sil90]. réels [Cos90a]. Rees [Sim90]. References [CP91h, Mum99c, Mum99g]. refined [Grü91b]. refinement [Mic93]. reflected [BM92d]. reflecting [Wer95]. Reflections [SM99o, Muk91]. Reflective [SM99p]. reflector [SM99l]. Reflectors [SM99q, SM99a, SM99b, SM99c, SM99k]. reflexional [GT91]. reflexive [Ouy91, Sim98l]. reflexivity [Ric99g]. Refraction [Lie93b, Lie93c]. Regeneration [Tho93, Sve93, Mai93b]. regime [Sap90]. region [Dix97b]. regions [Dix97c, MM90, Pan91a]. register [Nie90]. Regular [Kru97h, Nis96d, Bru92b, Khu90, Kos95f, SM99u, Tay92, vdPS97l]. Régularité [Bek91a, HT90b, Jac91]. Regularity [Eva97, Kos95e, Kos95c, Kos95d, Sch90s, Sch90t, Sch90r, Sch90u, Sim91l, Amg98, Dor93, Lie93h, Sal92]. Regularized [JL93a, JLG94, deL94o, Elf91, JL93c, JL94, Var92, deL94l]. regularly [HFS94g, Mol93h, Mol93a]. regulated [DN99a]. regulation [Kwa91]. Reider [Cat90, Sak90]. Reimannian [App95]. Reiten [Erd90d, Erd90g, Erd90j]. reiteration [Mil94g]. Related [Erd90e, Kom93a, KÜ90b, KRW92, Mim90, Pic90, Alb98h, BMMN98m, CMZ91, CS92b, CM97g, Fuk93, Her94e, Lio93b, Neu97k, Neu97q, OL90e, Ped92, Ros90, Sjö91, Soa94h, Vai93f, Woy98g, Yur95b, Zor92]. relatif [ÉY98]. Relation [Dur90d, Her94i, DN99c, Joh95c, Oda90b, Ric92, Xi94d]. Relations [BMMN98m, DS90e, Fre91e, GV91j, Lan96n, Sch94s, AR96a, Jac90, Kel91b, Mil91, Nie91k]. Relationship [deL94p, MS90a]. Relative [BS90b, Hon90, Joh95d, AHJLP90, Bou97l, Coo90, Göt94a, Kel91b, Pot95g, Res92, Sim91i]. relative-injectivity [AHJLP90]. relativistic [Ani91, Ich90]. relator [Ans90, Joh95a]. relators [Wal90b]. Relaxation [DZ93d]. relevant [Kah90a]. Relévement [PE92]. remaining [DRSŻ91d]. Remark [MS94b, Sab91, Top93, ÉS99b, Hit91, Par91b, SS91, Tay92]. remarkable [BB90]. Remarks [Bou91c, Coh92, JS91b, Lio93a, Wah95f, Żoł90, Ben91a, GHS97b, Gra99, Kni93, Mil94k, NÜZ90, SY98, Tak99, Tsf92, Wal93, Wer95, Woj91, Yan91b, Yap96d, Zim98a]. remarque [JP91a, SW93]. Remarques [Ans94, Cha96, Mic97, Mon94, AS90c]. reminders [Sch94t]. renewal [DGT91]. Rennes [CMR92]. renormalisation [AAG90, GRCW90]. Renormalization [BOERT96c, BOERT96b, BOERT96d, Yan91a]. renormalized [Le 94]. rentrante [DW97]. repetition [Kum90]. representable [SM99g]. Representation [Erd90e, FK93h, HN93j, ZD93, Zie96e, Mey93c, Mey93b, Bar93b, Bar93h, BV91, Erd90d, GV91d, Lei91, Mar96b, Mor90a, Nie91e, Nie91h, Ohs91, Ove95, Ree90, Ric98, Ric99d, Vas92, Zim98d, Eme91, LW90c, Lim90]. Representations [Eyr98g, Eyr98h, Fre94, Kah90b, Nie91c, Rut97q, Ser92o, Xi94l, Zuo99e, BL92b, Dek96c, Ene95c, FK93e, KW94, Kim94, Kin94, Kon90, Leh90, Nie91j, OS90, Pis96h, Rib99, Sho92f, Sri90, Xi94j, Zuo99c, Zuo99f, Dix91, Lab90]. Representatives [MO93g, Fin90]. repulsive [LP90a]. rescaling [DRSŻ91c].

Research [CC92, Eva92h, Kah90c]. réseaux [KS96]. residual [HU90]. Residually [Seg90]. residue [SS90]. resolution [BM90b]. Resolutions [Ben92, Hub95l]. resolvent [FS92]. resolvents [deL94c]. respect [Gra90a, Jan93, Mol93e, Sch90m]. rest [BK92c]. restricted [Clu92, Mey99j, Mur93]. restriction [Bou91d, Lou91a]. restrictions [Sch94t]. result [Bau91, Met91p]. Results [BSW92, Nil94f, Alb98h, AS98a, BM90c, Bor98a, BH90c, BEGG90h, BDT90, CV93, CM97i, CL99c, CL99e, CL99f, DK92, Dix97e, DS90f, Dow90, EGI94, Eva92e, Eyr98i, Fre91c, FK93c, GP91a, Har90b, IK99d, Jaj91g, JR92c, Kra90, Lan91f, LO98e, LO98f, Lie93j, Mar96j, Met91a, Nil94c, Pil94a, Pro90a, Raj91, Sch92a, Sch92b, Sch93a, Sto92c, Tai91e, Tam99f, Thr90, Wah95c, Wah95h, Yap96g, Zie96c, dF90]. Retard [CDD90]. rétinotopie [KS96]. retour [Eca90]. retournement [Mai93b]. return [Les92]. Reversibility [BF93i, deL94q]. reversible [Fur91, Mon91b]. Review [Acc94d, BP97k, CP93g]. revisited [BW94a, Eke90, HW97, Kat93b, Kum90, Per92]. Rewriting [Gro90b]. Reynolds [Cha92, Fur92]. Rham [MNM90, Dan91, Gek90, Hub95h, Kap91, Pus96b]. Rho [Bek91g]. Rho-functions [Bek91g]. ribbon [Ros90]. ribbon-disk [Ros90]. Ricci [BTV95a, BTV95c, Gau91]. Rickard [Kön98b]. Riemann [Acc94a, Acc94b, Acc94d, Acc94e, Acc94f, Acc94g, Bol92, BR90j, Cha96, DG90, Fuj90a, HT90b, JP91b, Lan90c, Mul90a, Oda90b, Sai91a, Shi91a, Sim91a, Sim91k, Tan91, Wol91]. Riemannian [AP90, BTV95e, BR90h, BR90i, BR90j, Haj91, Heb96f, KT91, LN97, MORT91, Mot91, Sch90m]. Right [dPW91]. Right-symmetry [dPW91]. rigid [Chi90a, LO98g, Mor90c, Pet92, Sap90, Zuo99c]. rigide [Ber90a]. Rigidity [GP91b, LP90b, CK91, Zai90]. Rigorous [Spi91, BL91a]. RIMS [Kom93a]. Ring [Her94j, Hod90, JLP90, Göt94a, Mar96g, Mor90a, RL90, SV91a, Wri90, Xue92d, vdPS97k]. Ring-theoretical [Hod90]. Rings [Gre92c, KZ98, Nar95c, SM99q, Xue92f, AHJLP90, Azu90, BH90a, BV91, Brö90a, Dat90, Dau90, FZH90, Gre92g, Her94e, Her94h, Ina90b, Khu90, Man97b, Man97d, McC90b, MS90c, SV91a, SM99u, SM99j, SM99n, SM99m, Ver91, Weh90, Xi94c, Xu96a, Xu96e, Xue92a, Xue92b, Xue92e, Zim98c, vdPS97i]. Risk [BR97]. Rivers [Sam91, Bla91, vdB91]. Rivest [BM90a]. RNP [Phe93a, vN92e]. Robba [Dwo90]. Rogers [Ste90b]. Rokhlin [BC92m]. Role [KS92]. root [Cou93, McI90, Ura93a]. Rooted [BOERT96c, BOERT96e]. roots [FK93b, FK93e, IS90]. Rotated [Yan90]. rotation [MR91a]. rotational [GT91]. rotations [Sap90]. Roth [McC99f, Tij93b]. Rough [Soa94e, Yur95e]. Round [CFJ^{+}92, FS90c]. Round-table [CFJ^{+}92]. rounding [van90]. Royden [Soa94f]. rule [FK93i, Law91a]. ruled [Bia90e]. rules [Far96g]. Runge [Pyt99f]. Russia [KZ93].

S [ACC92, van90]. S. [Oor93b]. S.D.E. [Kar91]. sûre [AW97]. sa [Rai96]. Saddle [PR97i, DR90]. Saddle-focus [PR97i]. safe [Pla90]. Saint [BGM94, Ber95, Ber96, Ber98b, Ber99a, BPSH91, GGSC97, Hen90a, Hen92, Hen93, BPSH91, Mie91l]. Saint-Flour [BGM94, Ber95, Ber96, Ber98b, Ber99a, BPSH91, GGSC97, Hen90a, Hen92, Hen93, BPSH91]. Saint-Venant [Mie91l]. Salice [TBR91]. Salvador [BS90c]. same [BL91b]. Sample [Blü90, BI92, CP91a, Jan93, Neu97l]. sampled [Fra91]. sampled-data [Fra91]. sampler [Tuc90]. sans [FT90, Mic92]. Satake [Zuc91]. Satellite [Mey99k]. satisfying [GH90, KW94, LQ95a]. sauts [Pic91]. scalaire [Jou98]. scalar [Haj91, Ric99b, Ros91]. scalar-type [Ric99b]. scale [Woy98h]. scales [Kuk93d, Mil94i]. Scaling [Ast91, Var93b]. scattering [BL92a, Bir94, Hor90, Ika90, Jen90, KZ91]. Schachermayer [BSÉ99]. Schatten [Dug92]. Schechtman [Hit91]. Schéma [DDD90]. scheme [DDD90, Her94h, Kle90, LO98l]. scheme-theoretic [LO98l]. Schemes [DL96a, Göt94d, Mum99f, Mum99a, Mum99g, Zie96a, BCMS90, Bui92a, Göt94b, Göt94a, Hu96, IK99d, IK99e, IK99j, Lan96a, LO98n, Lor92i, Shu98b, Tho90b]. Schiffer [Hum90]. Schmidt [Bor90b, Eve93]. School [BPSH91]. Schottky [Arb99]. Schrödinger [BM90c, GKKL91, Iwa90, Kar97d, Kar97f, Lan91g, LN91, MR91b, OT90, Oza90, Sjö91, Wan91b, Yaj90, deL94r]. Schubert [AMS90, Col90b, FP98f, FP98g, FP98i, MD90]. Schur [Pis96f, Pis96i, Pra91]. Schwartz [Mey91b]. Science [Odi90]. scrolls [PsT90]. Second [Bel90b, Del92b, Del92g, KÜ90b, Pyt99g, Wah95g, Den92, ELLV90, GV91h, Göt94e, Gra90a, Kos95a, Kos95c, Le 91, Wes93]. Second-order [Bel90b, ELLV90]. section [CvdG92, MR91a, Rei90a]. section- [Rei90a]. section-centroid [MR91a]. sections [Băd90, Bia90e, Dam91a]. sectors [FK93i]. Segal [Ouy91]. segments [HS90a]. Segre [Sko92]. Seiberg [Moo96a, Moo96b]. Selberg [CPS90, Ogu90, Sho92g, Sho92h]. Selected [AE96]. selection [BR97]. Self [Gau91, Pic90, Rut97r, Yam90, Ark90a, Boo90, Han90, Kah90a, Kah90b, Kah90c, Le 94, Møl90, Rut90]. Self-dual [Gau91]. Self-Equivalences [Pic90, Rut97r, Boo90, Kah90b, Rut90]. Self-homotopy [Yam90, Ark90a, Kah90a, Møl90]. self-intersection [Le 94]. Semi [CP93h, Fre92, Maz97, NZ99k, Sch94u, SM99r, SM99q, vdPS97l, Ari91, Arn93, Bia95b, Bru92a, Bru92b, CP93e, Kar97a, Kom93b, Not90, Pet96, PE92, Raj90, Rei97d, Van91]. Semi-algebraic [SM99r, SM99q, Bru92a, Bru92b]. semi-bounded [Kar97a]. semi-flows [Ari91]. Semi-Groupes [Maz97]. semi-groups [Bia95b, Kom93b]. semi-invariants [Van91]. Semi-linear [Fre92]. Semi-local [NZ99k]. Semi-Markovian [CP93h, CP93e]. semi-martingale [PE92]. Semi-martingales [Sch94u, Arn93, Raj90]. semi-positive [Pet96]. Semi-regular [vdPS97l]. semi-simple [Rei97d]. Semialgebraic [Del91c, Del91d, Kne92, AR92, Brö92, Del91b, Del91e]. semianalytic [Ruí90]. semiaxis [KM97f]. semiclassical [Bel94]. Semiconductor [DGDMN91, BMMP91]. semiconductors [Mar93a]. semicontinuous [Phe93h]. Semicontractible [Kos91]. semidihedral [Erd90c, Erd90i]. Semigroup [Tai91k, vN92c, AS92, CF95, Tod97d, Yan91a, vN92b, vN92d]. semigroupes [Bak94]. Semigroups [BV90, DGT91, Ebe99g, HN93f, Itō93a, KO93, Bar90a, BM90c, Gro93, HW97, HN93a, HN93b, HN93d, HN93e, HN93g, HN93h, HS90b, ITO93b, LS91, Nag90b, QH95, deL94b, deL94l, deL94p, deL94o, vN92e]. Semilinear [Del92b, Del92h, Tib90d, DH90, Sch90e, Tai91c]. Semimartingale [AS98g, Kar91]. semimartingales [AS98e, AS98g, BP90b, JY93]. Séminaire [Lem90c, Mal91, AYM90, AYM91, AYM92, AYM93, AYM94, AEMY95, AYE96, AYE97, AYÉL98, AÉLY99]. Seminar [HT90a, Ina90a, KZ93, Lem90c, LM91, OR91]. Seminars [Lat90]. Seminorms [Yur95f, HM92, Yur95d]. Semiprime [Dau90]. Semiregular [Pot95g]. Semisimple [Ser92p]. Semmes [Dav90]. sense [Acc94h, Sve93]. separable [Nar93]. Separably [Del98]. separate [Thr90]. Séparation [CS98, DP91k, Mur93]. Sept [AvW90, CF92, FR90c, HMRS90, JLP90]. September [Da 99, TBR91]. Sequence [LT96, CG91, GE98f, Les92, Lin90, Sch94i]. Sequences [DT97d, DTW90, DT97b, Erd90g, GK90, Gra90a, Jac90, KO91, Lan96o, Nie90, Nis96d, Pot95a, Pus96i, Räb91, vdPS97k]. serial [MS90c, Wri90]. Series [Bal94e, JL93a, JLG94, AFF90, Bal94b, Bal94g, Bal94i, Fre91f, Fre91h, Har90a, Her94d, Jaj91e, JL93c, JL94, Kur90, Oda90a, PG90, cSyW91, WY92, Xi94j, Dum91]. Serrano [Sak90]. server [DK93a]. service [DK93a]. ses [Bel99]. Session [Bal93, BA94, CF92, Cap99, CC96, CR91, Cel90, CP93b, Da 99, DM93b, DL97, FG90, FG96, FZ93, Gra94, HS99, Joh95g, MP91, Nar97, Odi90, Qua98, Rug96, Run97, TT96, TPZ98, Vio99, Vis94, ZD93, dBT91]. Set [Bek91g, Har98e, Ark90b, BLR90, BL91b, Cas92, CNZ91, Har98a, Har98b, Har98d, MS90a, Pil99c, Ran90b, Zvy90a]. Set-Chains [Har98e, Har98a, Har98b, Har98d]. set-valued [CNZ91, Ran90b]. Sets [Mol93g, AK92, ASHY90, AR92, BK92a, Ban92, Bar90c, Bek91f, BMMN98d, Bou97i, Bou91b, Brö92, BW91, Bru92b, Cha91i, CML90, DJS90, Éme97, FH90, FR92, Kho97, Lie93e, Mar96d, Mol93b, Mol93d, Mol93e, Mol93f, Mol93h, Mol93i, Mol93j, Nar95a, NR90, Pil94c, Pot95b, Pot95g, Ran90b, Ruí90, Sch91b, SSS92, Tod97b, Wic91b]. Setting [HFS94e, Neu97d, Pis96d]. Seven [CC92]. Seventy [Pre96]. Several [Neu97m, BKS93, Col92c, Ika90, Jam90, Nis96f, Wah95c]. Severi [Lor92j, Acc94c]. Shadowing [Pil99b, Pil99c, Pil99a]. Shafarevich [Kol91, Zuo99f]. Shape [GCR90, HMS94]. sharing [BR97]. Sheaf [Del91e, BV91]. Sheaves [BL94d, BFS98c, BS97b, Lie93g, SS93, Sch94f, Sch94g, Wil97a]. sheet [Bro93, Kho97, Ore90]. sheets [Caf91]. shell [Pom98c, Pom98e]. Shields [Iva93]. shift [Nie90]. shifted [AP91]. Shifts [Lor92h, Lam92]. Shimura [Ber93, Har93d, Rei97d, Wil97a, Wil97e, Wil97f, Wil97d]. Shimura-Varietäten [Har93d]. Shock [Woy98h]. Short [Wat90, CRV90, HA90b, Kit93, MT99, Yan96]. short-range [Kit93]. shot [Heg90]. showing [Wah95g]. shtukas [Lau97]. side [Kal93]. sided [Xi94e, Xi94i]. sidewalls [Nag90a]. Siegel [CP91d, CP91e, CP91g, CP91c, Fre91d, Koo93, Mig90, Pan91h, Pan91i, Pan91j, Sch91h]. Sierpiński [Hec98]. sieve [BL93b, BLP93, Cou93, LL93, Len93, LLMP93, Pol93b]. signal [Fla90, SN90]. Signature [Lab90]. signed [BM90c]. signum [LS93]. Silivri [KÜ90b]. Silov [Kan91]. Similarity [Pis96k, Pis96j]. Simple [Xi94m, Ade92, Ber93, Bou97n, Fra95, HALM90, Ish92, KLGP97h, Lan91h, MSS94, Rei97d, Rut97u, SY91, Xi94h]. simple-elliptic [SY91]. simple-homotopy [HALM90]. simples [Del90a]. simplices [Rez91]. simplification [dM99]. simply [Neu96e]. Simulated [Cat99]. simulation [BMMP91, Tho93]. simulé [Mic92]. Singer [Pot95e, Sho92q]. Singular [Dam92, Dav91b, Dav91a, Dav91d, DZ93e, DK94, Ebe99g, Fre91f, Fre91e, Fre91g, Hub95t, Kru97i, Lor92i, Maa91, Mat90a, Mit94d, Vai93c, Bas90, DRSŻ91a, Fuh91, Fuj90b, Hu91, Hub95g, Jon95, JL93b, MT91a, Mit94a, Nil94a, RY90, Vai93i]. Singularities [DRSŻ91i, Mas95h, RS91a, Ura93b, AMSV91, Beh91, BM90b, Bor90b, Chr91, Dim91, FS90d, JS91a, Mas95a, Mas95k, Ohs91, Pfe90, Sch91k, SY91, Shi91b, SS92, Ste91, Ura93e, MM91b]. Singularity [BF93k, BF93j, Caf91, SV91b, Gor91, Rou90, MM91b, RS91a]. Singularly [Bor98c, Sim91a]. Sinha [Mey93b]. site [Sch94k, Sch94r, Sch94q]. Sitges [XD90]. Six [DN99d, Tyu96, DN99c]. size [Jan93]. Sizes [Sim91m]. skeleton [Rut90]. sketches [Gra91b]. Skew [Tay92, Kow92, LS91]. skew-product [LS91]. Skorohod [BI92]. sleeping [Sap90]. Slices [Rob90]. slight [Tyu91]. slip [Wol92]. slip-condition [Wol92]. slit [YR90]. Slow [ERHG91, Del91a]. slow-fast [Del91a]. slowly [BC95]. slowness [Lie93c]. Sluggish [IG90]. small [AV91, DH90, Eva92c, Ion90, KLGP97e, PSW92, Pir90]. Smooth [Acc94e, Göt94d, KB96a, LQ95g, MO93c, Phe93j, BM90c, BHS96c, BK92c, Sad91]. Smoothing [GV91k, Oza90, Yaj90]. Smoothness [Vai93h]. Smorodinsky [Sch99]. Sobolev [Bak91, BT93, Cap98, Cha91m, Fon98, Fuj93, Gro93, Heb96c, Heb96d, Heb96e, Heb96f, I90, Led99, LN91, Mat98, Mil94j, Neu97e, Neu97i, Neu97n, Neu97p, Neu97o, Str93b, Üst95c]. solenoidal [SvW92]. Soluble [Sho92s, Sho92l, Sho92o, Sho92m, Sho92p, Sho92r, Sho92t, Sho92u]. Solution [Sch90d, AB91, Cha91b, Ino92, Kal99, Kar91, Koz90b, Mog92, Pom98g, Pro90b, Spi91, Vai93h, deL94i, deL94s]. Solutions [Dix97d, DL97, Mey99i, Rei91d, Alb98d, AS98e, BMO98, Bal94d, Bar97, Bar93i, Bia90c, BM90d, CD91, CMN95, CM90, Cra97, Daf96, Fab93b, Ge92, Gil90, HS91, Hen91, HMN91a, Hla90, Hum90, JP91a, Kos95a, Kos95b, Kos95c, Kos95d, Kos95f, Kos95g, KM97b, LP90a, Lio93b, Mat90a, MS91a, Mey99m, Mis92, Mor93a, Neu97o, OT90, Pyt99a, Sch92b, Sch90p, Sch90v, Sch90w, Son97, SS92, Tad98, Wie90, Wie92, Zvy90a, ZR92, deL94u]. solvability [KM97i, XL98j, vW90]. Solvable [Ser92n, Kac91, Kaw93]. Solving [Not90, Sch95d, Ehr92]. Solvmanifolds [TO97c]. Some [Acc94f, AP91, Ark93, Bal94h, Bar90a, Bar97, BK91, Ben91a, BMMN98n, BH90c, Bro90b, Bul90, BDT90, CV93, CM97g, Coh95, Del91f, DK93a, Don98, Erd90h, Eyr98i, Fuj93, GK90, Gra99, Har90b, Hub90, IK90, Jan93, JL93c, Kah90c, Kho97, Kit94e, KLGP97j, KLGP97k, KLGP97l, Kni93, Kos95f, Koz91, KB96f, Lan91f, Law91b, Le 92, LO98m, LO98n, Lie93j, Mal92, MT90, McK90, Mil91, Mil94k, Mor91, Ner90, Neu97q, Nie91i, NÜZ90, Opo91, Raj91, Sch92a, Sch94t, Sch92b, Shk93, Sho92u, Skv90, Soa94g, SY98, Tak99, Thr90, Tsf92, Tuc90, Üst95k, Vai93i, Wah95h, Wal93, Wer95, Wir98f, Woj91, Yan91b, Yap96g, Yap96h, Zak91, Zim98d, Acc94d, Acc94h, Bar92, BI92, Bia95b, BM92b, BCMS90, BMMP91, CRW90, CMZ91, Car91a, CL96, Cri90, Dri92, FH90, Gel90]. some [Gin97b, GK98, GDT91, HH92, Hun96e, Iva93, Kön98a, Kos95d, Kra90, KJ90, Liv90, Lun92, Mar90d, Mél96, Mus92, Ore90, Ôta93, Shp90, Üst95b, Vel90, Wah95c, Wah95f]. sommation [Eca90]. Sommes [Rod92, DL90]. son [Bak94]. Sous [Cos90a, DI98, Dum91, KPR90, Mou90]. sous-analytiques [KPR90]. Sous-corps [Dum91]. Sous-ensembles [Cos90a]. Sous-mesures [DI98]. sousmartingale [CS98]. Space [Mey93e, Mey93f, Mey93m, NZ99l, Oba94e, Pil94e, Üst95e, Vuo92, AS98b, Bel90a, BR90i, Cas91, CG97f, CG97g, CM97h, Cha91e, Cha91l, Cha91m, Cha92, Chr91, CP93e, Ebe98, FF97, Fuj91, HW91, Han90, HOR91a, Her90a, IK99a, Ino92, Iva93, Joh95e, Joh90, Koz90b, LN97, Mad91, Mey95a, Mey95b, Mey95h, MO90, Mol93i, Neu97b, Neu97c, Neu97o, NV90, NÜZ90, Nua98, Oba94g, Par91b, Par91c, Phà91, Pis96d, Pri90, Sch93e, Sch93h, Sen90, Siu91, Tuc90, Tyu92, ÜZ97, Wah95c, Wah95g, Woj91, Yur95c, Zor92, deL94s, deL94u]. space-time [Ino92]. Spaces [Ban91a, BTV95c, BR90j, CG97h, CM97c, Del91c, DL96a, Far96d, Gro97a, HWW93f, Heb96f, HMN91h, JJS92, Jaj91h, Kit94f, LT96, Mar96h, Mar96i, Met91j, Mey93f, Mit94d, Moe95c, Nil94b, OL90d, RG96f, Rut97r, Vät97b, Wei94b, Bal91, Bar93g, Bau91, BTV95a, BC92a, BFS98b, BFS98i, BEGG90e, BR90c, BR90g, BR90h, BT93, CMZ91, CG97b, CM97a, CM97b, CNZ91, CP93g, CP93h, Day92, Del91b, Del91e, Did90, DGT91, Dul91, ELL99b, Far92, Far96e, Far96a, FJ91, FS92, GK98, GE98f, HWW93b, HWW93a, HWW93g, HM91, Heb96d, Heb96e, Hen90c, HN93a, HN93i, HMN91b, Hit90, Hit91, HKR92, Hub90, Hub95f, Hul90, Hun96g, IK99j, I90, Ish92, Jaj90, Jaj91a, JP91b, JT91b, KS92, Kis94, Kit94c, Kit94e, Kna95d, Kna95h, Koe99c, KF93, KÜ90a, KV92, Kuk93d]. spaces [Kum97, KB96g, Leu91, LP91, LL91, pL91, Lum91, Lun92, Met91b, Met91i, Met91h, Met91g, Met91f, Met91m, Met91o, Met91p, Mey95b, MR98, Mil90a, Mil94k, Mit94f, MN91, Moe95b, Møl90, MO93b, Mor90c, NV91, Nar93, Neu97e, Neu97p, Nil94e, Nil94f, OL90b, OL90f, Ohs91, Phe93a, Phe93d, Phe93f, Phe93i, Phe93j, PsT90, RB92, Ric99a, Ric99h, Rut97h, Rut97o, Sch94l, Sch90h, SV91a, Shi91a, Sim98k, Sim98l, Spr90, Tam99d, Tod97b, TO97g, UT92, Van91, Vät97d, Vät97c, WY92, Wei94c, Wei94g, Woj91, XL98i, iY90, Zab99, ZD92, deL94i, dlR93]. Spain [ACC92, CC92, XD90]. spanned [Bal90b, Pal90]. spannedness [BS90a]. sparsity [Mau90]. spatial [Dur99]. spatio [FMSW91]. spatio-temporally [FMSW91]. Special [Bra98g, Dix97e, Erd90i, Sim98i, AVV92, Bar92, CDZ91, Dow90, Hun96e, Ina90b, Nie91i, Pil99d, Sie91, Sim90, Tyu92, WJ90]. specific [Oud91]. Spectra [KTV92, Mar96i, Mar96a, SM99m]. Spectral [Blo92, BT93, DM96, How90, Lan91g, Lan96o, Nik94b, BBG91, CL99g, Gol94, Lin90, Mat98, Pre96, Ric99b, Ric99f, Sho92i, Ste90a, SW92, mS91b]. spectraux [Mic98]. spectrum [AB90a, How90, Mar96g, Sch94r, Sch94q, Wol91]. Spencer [Her94f, Her94g, LZ92, ZL90]. Sphere [Kun92, Lan90c, Han90, Küh95b, SOA91, Mul90c]. Spheres [Koc90h, Bar93b, Bar93h, BR90a, BR90f, Coh90, Cru90, Koz91]. spherical [FP91, Lau91, NM92, Ros90, UT92]. Spherically [BOERT96e, MM90]. spider [Wat99]. Spin [Mar99, Mey93l, Mey95g, Moo96d, BP97e]. Spines [BP97c, BP97a, BP97b, BP97k, CS92b]. spinned [Sai91a]. splines [Mic93]. split [Azu90, Lan96d, Lan96a]. splittings [Sad91]. Spreads [Kna95h]. square [Cou93, McI90]. squares [Elf91]. stabilité [MS91a]. Stabilities [HMN91i, Jag93]. Stability [AT98, ELL99g, Rei91d, BS92, Brö90a, Car91b, Cha91a, Daf96, GP91a, Haz93, Hon90, Mao91, Mol93e, Mol93j, Nag90b, Pil94f, Wie90, XL98f, Zie98, Żoł90, KZ93]. stability-concepts [Haz93]. stabilization [GV91f]. Stable [BR90f, Erd90j, Koc90h, LQ95h, Bot92, CA91, Eis97, ÉS99a, FJ91, GRCW90, Jan93, JP91b, Kah90b, Las98, Led92, Lin98, LZ92, Pil99e, Shk93]. stably [Hub90]. stacks [JT91b]. stage [HFS94e, Møl90]. stalks [Sch94n]. Standard [AR96f, BP97c, Ino98d, Ino98e, Alb98c, BP97b, BP97k, BF93g, BL90, Cad97, CP91e, Pan91i]. star [Mor90b]. Stark [Wan91b]. stars [DvG91]. starting [TO97d]. State [Pyt99e, AS98b, Lau91, Yau92]. Statement [Kuk93c]. statements [BP97i]. states [DU92, Str93b, Tam90]. stationarity [Tho93]. Stationary [Mis92, AB91, Pro90b, Sal92, Sap90, Sim91g, Soc92, Wac92]. stationnaires [CTR96]. statistic [FT93]. Statistical [DK93c, Woy98i, Dob93, Fur90b, Gra96]. Statistics [Ber96, Ber98b, Ber99a, GGSC97, Bro93, DX93, DN99c, FK93d, Gin97a]. statistique [Jac91]. stay [BD94]. steady [Ant92, Cha92, Lau91]. Steenrod [Lin90]. Steepest [Neu97r, Neu97b, Neu97c]. Stefan [Rod94]. Steinberg [Ree90]. Steinhaus [Sch90f]. stem [Koc90a]. stems [Koc90e, Koc90g]. step [Ino92, Kit94a, Var92]. Stephen [Lup90]. steps [Ebe99b, Hyl91b]. stéréo [FM90]. sticky [War97, War99]. Stiefel [Deg92]. Stieltjes [HMN91j]. Stochastic [BK92b, Car91a, CK90, Dur99, ELL99h, Eyr98g, Fla91, Gli92, Har98f, HS90b, KZ93, LQ95i, Mey93m, Mey95h, AH90, AT98, AS98e, AS98d, BMO98, Bax91, Bel90a, Ben99, BH90c, Box91, Bre94e, CMN95, Chi97, Coh95, Der95, EQ97, EY93, ELL99g, ÉS99b, Eyr98f, Eyr98h, Eyr98i, Fag90, GK98, Hu96, JS90, Kal99, KY91, KY93, KP96b, KP96c, MY99, MY07, MR98, Mis92, Nor92, Nua98, Par90, Pul96, Shi94, Tay92, Yan91b, Da 99, KÜ90b]. stochastique [Bia95a, Gro90a, Pic91]. Stochastiques [Par91a, Cép95, JP91a, JY90, MS91a]. Stokes [HMRS90, HMRS92, BM90d, BM92b, Cha92, FS92, Fur90a, Fur90b, GGS93, Gru92, Hey90, Kat90, Kaz90, Kos95b, Koz90b, Mog92, OSK93, Pil90, Pro90b, Puk90, Ran90a, Ran92, Riv90, Sal92, Sch92b, Soc92, SV90, SW92, Val90c, Var92, Wie90, Wie92]. Stoll [Cad97]. Stone [Ric99a, AMM90]. Straightening [Bru90]. Strange [PR97c, AK92, Naz92]. strata [Sko92]. strategies [Bea90, LL90]. stratification [Deb92]. strengthened [Rez91]. strictly [CMM91, ELY97]. string [AG90, DLS90, LR90]. strips [Mie91i]. Strong [AS98d, Hu96, Jaj91h, JT91b, Koz90b, Rak93, Win90, AS98c, AS98e, AS98g, AS98h, Bor98d, Hen90e, Ino92, Jaj91e, Kal99, Kos95b, Ohs91, Wie90]. Strongly [BIOS90, MT99, Ric99d, deL94a, deL94j, deL94n]. Stroock [RS94]. Structurally [CA91, Pil99e]. Structure [CDP90m, Mar96j, TO97f, Ars90, AR94, BFS98g, BR97, Fie91, FK93i, Fuh91, Fuj91, Kol91, Met91e, Mul90c, Pil99d, Tam99d, Wes93, Woy98h, Zim98e, Zvy90a]. structured [Die99]. Structures [Dek96b, AR96g, BMMN98e, Eyr98a, Ge92, Hun96g, Kai97h, Kan91, Kuk93d, Lic90, Mar96f, Pom91, Pot95e, TO97g, Wil97c, Zar91a]. Studies [BGV90, BGV92, Bus92]. Study [Voi94, Del91a, Fro91, Ike90, Lea90, SY91]. Sturm [KM97h, Sim91n]. sub [BZ92]. sub- [BZ92]. Subadditive [Yuk98i]. Subadditivity [Yuk98j]. Subalgebra [RS90b]. subalgebras [Pus96e, Tam99c]. Subanaliticity [Den92]. subanalytic [Shi90]. subcategories [SM99p]. Subcones [KR92g]. Subdifferentials [Sim98j, Phe93i]. subfactors [Kaw93]. Subgroup [Fin90]. Subgroups [Ban91a, Bur74, AP91, BOERT96a, BEGG90g, Bur90, Gro97e, Gro97f, Har90b, Lan96n, Neu90, Sho92l, Sho92o, Sho92m, Sho92p, Sho92t, Sho92u]. subharmonic [GLO94]. subharmonics [Fur91]. subject [Ant92]. Submanifolds [Küh95f, Lum91, Ebe98, Kai97a]. submeasures [Šip92]. submodules [Azu90]. Suborbits [BMMN98o]. Subordinators [Ber99b, BC95]. Subschemes [Göt94d]. subsequences [DK92]. subset [Bia90a, Bia90b]. subsets [IT90]. Subspace [Eve93]. subspaces [ABW92, OL90e]. subvarieties [BL92c, Cle90]. Sud [Lem90c]. Sufficient [Bro93, vW90]. sum [Aki90, Kir90, Sim98k, Sim98l, VM92, FT93]. summability [BDT90]. Summable [Bal94i]. summands [HWW93a, HS92, Pri90]. Summary [DS90f, Lem90d, Tam99f]. Summer [BPSH91]. Sums [AS90b, Cas92, IK99g, IK99i, Yur95d, Yur95g, AS90a, Bou91c, IK99f, JPA97, KK93, Küh95c, VC90, Woj91, Yur95b]. suns [DvG91]. Super [Tam99b, Tam99c, Tam99g]. Super-Algebras [Tam99b]. superadditive [Yuk98i]. superadditivity [Yuk98j]. superalgebra [Eyr98e]. Superalgebras [Eyr98g, Eyr98h, Naz92]. Superbranching [DM90c]. superconductivity [Neu97s]. Supercongruences [Cos90b]. Superconvergence [Wah95i, Wah95k, Wah95l, Wah95j, Wah95m, Wah95b, Wah95c, Wah95e, Wah95h, Wah95g, Wah95a]. supercritical [JN96]. supermartingales [CJ90]. Supermoduli [FR90b]. superposition [Nag93, Sen90]. superprocesses [Ove95]. Superselection [FK93i]. Supersingular [LO98h, LO98o]. supersingularity [LO98l]. superstrings [FR90b]. supersymmetric [BL90]. Supplementary [Arb99]. supplements [Kna95a]. support [Bom91, Lem90a, MSS94]. supported [FP98h]. supporting [Éme97]. supremum [ELY97]. sur-et [CS98]. sure [Git98, Jaj90, Jaj91a, Kar91, Mao91, Mol93b]. Surface [Kai97h, Sim91a, Ant92, CvdG92, Fri91, Hun96a, Mie91a, Mog92, NZ99a, NZ99g, Rei90b, Rut97f, Sim91k, Tan91, Wol92, Wol91]. surface-near [Fri91]. Surfaces [Acc94g, Brî96d, BEGG90c, BR90j, Dav91d, MO93c, Rut97s, Acc94a, Acc94b, Acc94d, Acc94e, Acc94f, Bal90b, BS90a, Bia90e, Bor90b, Brî96c, Brî96e, BEGG90a, BEGG90b, BEGG90d, BEGG90f, BEGG90e, Dav91b, DPlPS90, DG90, DV91, DR91, Hon90, Hul90, JP91b, Kaw92, Küh95g, Kus91, Le 90, Lie93c, Liv90, MM91a, MO93b, MO93h, Neu97h, NZ99i, Oko91, Oli90, Sai91a, Sak90, Sch90r, Shi91a, Sil90, Ste99b, Tyu92, Ura93e]. Surgery [DS90a, DS90b, DS90d, DS90c, DS90e, DS90g]. Surjectivity [BEL91]. surmartingales [CJ90]. Survey [Mol93j, Rut97r, TO97e, Win95d, And91, Ark90a, CW91b, Dek96d, DN99c, FK93c, McI90, Pra90a, Sto92n, Wal91b, Wür91]. surveys [Vuo92]. survivability [Sto92k]. Survivable [Sto92e, Sto92n]. survival [Gil94]. Suspending [Mas95k]. suspension [Far96e]. suspensions [Kos91]. Suzdal [KZ93]. swarms [Fro91]. Swinnerton [Rub99b]. Swinnerton-Dyer [Rub99b]. swirling [Raj91]. Symbolic [BK92a, CP93i, CP93j, Her90b]. symbols [AB90b, Ree90]. symétrie [CF90]. symétrique [Dix91, Eis97, Gro90a]. symétriques [DI98]. Symmetric [AMSV91, BTV95e, BR90g, BR90j, FP98j, Sch93e, Vas90, Bar93i, BMMN98i, Bru90, BR90h, BR90i, Erd90k, Far96h, HB91, JR91a, Kan91, MR94, Phà91, Rou90, Shk93, UT92, Van96e, Yan93]. Symmetric-like [BTV95e]. Symmetries [Bar93j, Kru97j, Mey99l, Bar93c, Heb96e, Hor90, Mie91e, Tam99e]. Symmetrization [Sch93f]. Symmetry [Nag90a, CA91, CGG^{+}91, GGS92, Ger92, GT91, Lau91, Wah95j, dPW91]. Symmetry-breaking [Nag90a]. Symplectic [BF93j, JS91a, Kuk93d, TO97f, TO97g, BF93c, BF93h, JR91a, Lie93a, Wan91a]. Symplectique [Alb90, Fra90, Mul90c, Tur90a]. symplectiques [Aud90, Daz90, Lic90]. Symposium [BDF91, FIK90, NF90, NO91, Vil90]. synthesis [Nik94b, Ste90a]. synthetic [Hyl91b]. System [Wir98g, CA91, CP93d, FS90d, GGS93, HS91, Kel91b, Kos95b, Mat90a, Mog92, Mor93b, Mül91, Pom98g, RS90d]. systèmes [DM90b, KS96, Mey93a, RS94, dM93a]. systèmes-produits [Mey93a]. Systems [Bor98c, BHS96i, BM91, CP93b, CNZ91, FG96, Joh95g, Ken94, Kos95e, Kuk93a, LQ95g, Pil94e, Pil99b, Pil99d, Pom98a, Pom98h, Rei91d, Sch90u, Tib90c, Van96g, AV91, Ama91, AC91, Arn95b, Bar93i, Bol92, Bor98d, CP91a, Car91a, CT92, CP93a, Col92f, Cra91, Daf96, DM96, Dur95, ELL99g, ELLV90, FS90b, FMSW91, FS90c, Fra91, Fur91, Gri97, Gro90b, Had99, Hel91, HMN91e, HMN91f, HMN91g, HMN91i, Hla90, IG90, Ino98d, Kit93, Kos95a, Kos95c, Kos95d, Kos95f, Kru97d, Kru97g, Kru97h, Kru97i, Kru97j, Kuk93d, Lio93a, LP90c, MF94a, Mél96, Mey99e, Mie91e, Mie91g, Mie91j, Miz92, Mon91b, Not90, OS90, Phà91, Pil94b, Pil99e, Pom98d, Pul96, Rei91a, Rei91b, Rut97i, Sch90j, Sch90s, Sch90v, Sch90w, Sob90, Van96a, Van96f, Van96e, Van96j, Wal91b, Żoł90]. Szego [Dur90d]. Szemerédi [McC99g].

table [BK94, CFJ^{+}92, FS90a]. Tables [KLGP97m, KLGP97n, Lun92]. Tabulation [CM97i]. tails [ELY97]. take [Shi96]. Takesaki [Ino98f]. Talagrand [JS91b]. Talk [Ada90]. Tame [Erd90e, Erd90k, Erd90j]. Tangent [IK99j, HN93g, KPR90]. Tangential [Her94k, Joh90]. Taniguchi [NO91]. Tarry [RS90a]. Tate [Cri90, Hub95u, Kol91]. tautness [MORT91]. Taylor [Oka90]. Tbilisi [Ina90a]. technical [Tay92, Wah95g]. Technique [GE98c, GE98b, MT91b]. techniques [Cen91, Foi91, Fro91, Man97h, RG96a]. Teichmüller [Her90a, NV91]. telecommunications [Mor91]. temperature [IG90, SOA91, Mic98]. temperature-gradient [SOA91]. temperatures [SOA91]. Template [GHS97g]. Templates [GHS97h]. temporally [FMSW91]. temps [DW97, Eis97, ÉY98, Fla90, GRCW90]. temps-échelle [Fla90]. temps-fréquence [Fla90]. Ten [Dur95, Kne92, Sch92a]. tension [Ant92, Mog92]. Tensor [BFS98i, vN92i, BL92b, CS92a, Leu91, Lor92k, RS91b, Sch91b, Wah95m]. tensor-product [Lor92k, Wah95m]. term [Jac90, Lin90]. Terme [CR91, DL97, FG90, FG96, FZ93, Gra94, Joh95g, Odi90, Rug96, TT96, TBR91, Vis94, dBT91]. Terminology [Bui92g, Yap96a]. terms [Far96i, PsT90, Sim91l]. Testing [AH92d]. Texas [OR91]. th [Cas92, Son90]. Their [DS90a, Hub95q, Koe99f, BV90, BSW92, Bru90, Dek96g, HWW93b, HWW93a, Har98f, HN93g, HN93f, JcC91, Kos91, MY99, MY07, Mal92, Mol93h, Mum99b, Mum99c, Mur91, Nis96a, OL90d, OU91, PsT90, Pyt99a, SV91a, Wei94b, Zak91]. Theorem [Eve93, Gre92b, Lan91b, Lan91c, MO93f, Tai91g, Tai91h, Tij93b, vdP93, AH92c, AH92h, AMM90, Ban91b, BL91a, Bom91, BS99, Bru92a, Bul90, Cat90, Fag90, FW91, Gao97, Hod90, HA90b, Hol90b, Hru98b, IR92, Kah92, Kho93, Kön98b, Kuk93b, Kuk93c, Le 91, Les92, LL91, Lor92j, MR90, Mar96b, Mar96f, Met91r, Mil90b, MSS94, Mil94c, Mor93b, Muk90, Ohs91, OS90, Opo91, Pfe90, Pra99, RG96h, Ric99g, Ric92, Sch90f, Sch91j, Sho92e, SY98, Tyu91, Üst95f, Üst95a, War97, Zor92, Ole90, Tai91i, Tai91j, Weh90]. théorème [AS90c, AJKY93, BÉK^{+}98, Ber93, Dav90, Daz90, ÉY98, Mar90c, SW93, Sch94u, Thi91]. Théorèmes [Eis97]. Theorems [Ale91, HYL91a, Jaj91h, Mol93g, Neu96d, Ura93f, BL93a, BKS93, BMMN98b, CMZ91, Dem96, DS90c, Dug92, Fre92, Fuj90a, Gin97a, GV91e, Hey93, HMN91d, Isa92c, Jaj91f, KK93, Lor92b, Mal92, McC99f, McC99g, Mil94a, Mil94j, Moe95e, Mol93c, Mol93f, Mol93h, Mol93j, Neu96c, Nus93b, SS93, Sch91a, Sch94a, Sch94f, Sch94e, Sch94n, Sch93c, Sto92i, Wac92, Wat90, Xu96f, Yee90, Yuk98k, Zuo99c, deL94c, vN92f]. theoretic [LO98l, Pil98a, Rut97g, Sch94l, HT90a]. theoretical [Col92h, Hod90]. Théorie [Anc90, Del90b, Del91g, Bak94, Ber90a, CT93, CDP90f, Del90a, Dur90a, Dur90b, Hen90b, Lem90b, SN90]. Theories [DS90a, Boo90, BL90, DS90e, Dub96, FK93a, HKR92, Wür91]. Theory [Acc94g, AE96, ASMS90, BGM94, Bal94e, BA94, Bek91g, Ber95, Ber96, Ber98b, Ber99a, Bia90d, Bou98b, BF93j, BR90j, CP91a, CPR91, CG97g, Che90a, Che90b, FK93g, GGSC97, GS93, Gre94, Gro97a, HMRS90, HMRS92, Ino98f, JLP90, Jan90d, Kai97f, Kar97f, Kov90, KM97h, KRW92, KB96d, LC90, Lan91g, Lat90, LQ95g, LPLG^{+}92, Mal91, McC99b, Mey99m, Mim90, MM91b, MP91, Mur94, NF90, Pot95d, RS91a, RSSV90, Sch90u, Ser92m, Sma92, Soa94d, Sou97b, ST92c, Vio99, YS92, Yuk98g, ZD93, Zie96f, dJ93, vdPS97f, AP94b, Acc94d, AR96d, Aik96, Alb98a, AFF90, AG90, Ani90, Ark93, Bab92, Bar93c, Bar93f, BK91, BL91a, BSW92, BMMN98n, Bir94, Bjö97, Bor90c, Bor98a, Bou97m, Bou98a, BCDS92, BF93k, BHS96e, BHS96f, BDT90, Bur91, BT93, Coa99, Coh92]. theory [Col92a, Del91f, Del91e, Dem96, Dob96, Dri92, Dur94, EQ97, Erd90d, ERHG91, Ess96, Far96i, Far96a, FK93a, FK93d, FK93h, Fur90a, Gel90, GHS97g, Gil92, GV91d, Gre99, Gre92d, Gre92g, Gre92h, Gub90, HWW93c, HOR91a, Hey90, HN93j, HN93k, Hyl91b, IP92, Ina90a, Ina90b, Ino92, Itō93a, Iwa92a, Jan91, Jan90b, Jan90c, Jen90, Jon95, Kan90, Kar97c, Kar97e, Kar97d, Kat93a, Kem91a, Ken93, KLGP97n, KB96b, KB96a, Lan90a, Lan90b, Lan91b, Lan91c, Law91b, Lax96, Lod92, LR90, Man97i, Mat90b, McC90a, McC99a, McC99c, McC99e, McK90, Met91p, Mie91h, Mil94b, Mis95, Mit94b, Moe95a, Mur91, Neu96e, NZ99e, NZ99h, NZ99j, NZ99k, Nis96b, Nis96e, Nis96f, Oba94b, Paz91, Pil98b, Pom91, Pus96a, Rei90a, Rei90c, Ric98, Röc93, Ros94, Sam91, Sch91c]. theory [Sch90n, Sho92a, Sho92c, Sho92f, Sho92i, Sho92j, Sin91, SV91b, Sou92, SS90, ST92b, mS91b, Suz90, Tai91k, Tai91b, Tho90b, Tuc90, Uri90, Vai93e, Ver90, Wat90, Wil95, Yan91b, Zie98, Zie96d, Zie96e, Zim98d, Zuo99c, vN92h]. Theta [AR96g, Fre91e, Fre91h, Pro98a, AMSV91, Böc90, Fre91f]. thin [KLGP97l, Rau95b]. third [GH90]. Thom [Far96h, Mas95l, RS90e]. thoughts [Law91b]. Three [CG97f, CG97g, GHS97e, KZ91, McC99f, RG96g, Win95a, vH90, AH92h, Fis90, Jac90, MS90b, Pal91, Rut97m, Win95b, Win95c]. Three-Dimensional [GHS97e, Win95a, RG96g, Fis90, Pal91, Win95b, Win95c]. Three-space [CG97f, CG97g]. three-term [Jac90]. threefolds [Bar92]. Thue [Sch91i, Spr93j, Spr93i]. Thurston [MR90]. Tianjin [CDZ91]. Tight [Küh95g, Küh95f]. Tightness [Küh95h]. Tilting [Zim98e, Zim98c]. Time [Bor98c, Fri90b, GKW91, AS98f, AS98h, BM90d, DK93a, Dix97d, ÉS99a, Fro91, GN97, Hu96, Ino92, JPA97, Le 94, Les92, Mon91b, Sch90b, Wat90, Yan96]. time-changes [ÉS99a]. time-dependent [Fro91]. Time-inhomogeneous [Fri90b]. times [BK92b, MR94, Pit99, Raj96]. tiny [Ino92]. Titchmarsh [FT90]. Tits [DLS90, Lan96e]. Toda [Ada90, BCDS92, Koc90i, Van96d]. Today [CC92]. Toeplitz [Pee94]. Tokyo [FIK90, NF90]. Tomita [Ino98f]. Tomography [HLN91, De 91, Fri91, GDT91, Grü91b, Oud91]. Tomorrow [CC92]. tool [Col92h, Del91a]. tools [Kru97a, RG96d, Wah95f, Wah95g]. top [CvdM90, Sap90]. Topics [Acc94g, AE96, AG90, Bek91g, Ene95g, Geh92, Isa92j, Kom93a, KÜ90b, KRW92, LC90, Lan96p, Mim90, Pic90, Sim90, Szn91, Tod97e, Fuk93, Hel91, Kit94e, Soa94h, Mal91]. Topoi [Moe95c, Moe95b, Moe95d]. Topological [Ban91a, Bar93j, BC92g, BC92h, BC92i, CG97h, KB96b, Maw93, McC90c, Pil94f, Sch94l, Sre92, Bor90c, BS90b, Dek96g, Dub96, Gil92, Hel92, HA90b, McC99e, MR98, Sch94g, Spr90, Tuc90, Ver90, FZ93]. Topologically [Pil99e]. Topologie [Hau91]. topologies [Cor91]. topologique [Bek91a, JS92]. topologists [Tuc90]. Topology [ACC92, JOP91, Joh95f, Lat90, MO93c, Pil94e, Shu92, Tod97e, Zhe96, dBT91, BK92a, HN93d, Hit90, Kne92, Kus91, Leh90, Mus92, Phà91, RS90e, Ruí90, Sch94s, Wic91e, vN92a, Lat90]. topos [Ben91a, Moe95a, Sch94l, Sch94o]. topos-theoretic [Sch94l]. toposes [Law91a, Sch94m, Sch94j]. toral [Sad91]. tori [Ish92, Kun92, Van91]. toric [BL94c]. Torino [BA94]. toroidal [Dan91, SvW92]. Torsion [Rib99, CT93, CP93e, Far96e, Kin94, Pom98c, Rut97u, Ver91]. torsion-free [CP93e]. torus [YSA92]. Tossing [BSÉ99, BM90a]. Total [Yap96j, Yap96i, BM90d, Jen90, Lor92l, TO97g, Yap96h]. Totally [SM99t, SM99s, SM99o]. totten [Met91r]. touchant [De 96]. tours [BW94b]. Trace [Sho92g, Sho92b, Sho92d, Sho92h, Vos92]. tracer [Woy98g]. traces [Aki90]. trading [Cvi97]. traitement [Fla90, Fra95, Lim90, Mey90]. trajectoire [BW94b]. trajectorielle [Jou98]. tranches [KPR90]. transcendant [Dur90c]. Transcendence [Nis96c, Nis96e, Nis96f]. Transcendental [Ber90d, Voi94, CC96]. transfer [Gro97g]. transform [Bal94c, GV91j, Gra91a, HYL91a, LM90, Lou91a, OK91, Pal91, Sim91g, Sin92]. Transformation [Alb98i, Alb98j, AAG90, Del92b, DRSŻ91h, AP91, Del92c, DS92, DÉY93, Mey94]. transformations [ABÉH95, HK92, Wan91a]. transformed [GV91b]. Transforms [TPZ98, Bal94f, Ber98a, Bom91, GV91j, Hel98, Maa91, McG92, Sch90g, XL98i]. transient [Shi96]. Transition [Rei91d, Per91, Gra94]. Transitions [Vis94, Cat99, CK90, Ken94, Mül99]. transitive [BOERT96e]. Transitivity [BMMN98p]. Translation [Kna95i, Eva92a, Kna95e, Kna95f, Wah95a, Wah95k]. transmission [Khi92, Lie93j]. transport [Gol91b, Had99, Woy98g]. Transversal [MORT91]. Transversality [LS90b]. Transversals [Hig90]. treatment [GV91i]. Tree [Wei94f, Amg98, Bax92]. Trees [BOERT96c, CP93k, KV91, Per99, BOERT96e, LL90]. Trefethen [Apt93]. Trento [BBD90, BCC92b, Bal93, GT90, CTKV93a, Vis92b]. triangle [Kel98]. triangular [KS90b]. Triangulated [Ric98]. Triangulations [Küh95f, Yuk98f]. trigonometric [cSyW91]. Trilinear [Har90c]. triple [DL96c, iM90b]. Triptych [NZ99m]. trisecants [Joh90]. trivial [BC92a, Kai97a, Kai97d, Kai97e]. trivialité [Bek91a]. Trivialités [CR92]. triviality [Hel92, Lan96i]. trois [Sch94u]. troisième [Daz90]. trou [Mat98]. Trous [Mic98]. true [YK92]. truth [FS90a]. truth-table [FS90a]. Tsirelson [BÉK^{+}98, ÉY98, ÉS99b, Sch99, dM99]. Turbulence [BC92j, Woy98a, Woy98b]. Turkey [KÜ90b]. turns [AB91]. twin [Abr96b, Abr96e]. twist [Hub95u]. Twisted [Cha91a, DS90g, Ol’92, HFS94f, Küh95b]. Twistor [BR90h, BR90j, BR90c, BR90i]. Two [BLR90, Fuk93, Hel91, Kal93, Lan91h, McC99g, Sch90d, Sch94m, Sch90u, Wei94g, Ans90, AS92, CS92c, Fis90, FK93f, Neu97a, OSK93, Ore90, Per92, Pet91, Pet92, Riv90, SS91, Sch91f, SOA91, Xi94e, Xi94i]. Two-Parameter [Wei94g, AS92, Pet91]. Two-side [Kal93]. two-sided [Xi94e, Xi94i]. twosided [Zim98e]. Tykhonov [DZ93b, DZ93f]. Type [Erd90e, deL94t, Alb98c, Alb98d, AJK^{+}96, Bre94h, Clo90, Dek96c, Erd90a, Erd90d, Erd90b, Erd90c, Fla91, FS90c, Fre90, GKW91, Hen91, HA90a, JN96, KW94, Kin94, Kna95g, Lan96m, LO98c, LP91, Lin98, Lor92k, Lor92l, Mey93c, OL90c, Ric99b, Rut97j, Sch90g, Sch90w, Shi91b, Spr93e, Sri90, Tyu92, Xi94c, Yap96b, deL94c]. typed [Gra91b]. Types [BC92k, Bor90b, Cha91a, Far96e, HP92, Sce90b, Sto92k, Xue92e, iY90].

U [YS92]. Uhlenbeck [BW94a, HW97, QH95, Üst95c]. Ulam [IR92]. ultracontractivity [Mur93]. ultrasound [GDT91]. umbilical [GS90b]. Umbrella [Yuk98k]. Unbounded [Ino98f, BM90d, BM92b, Wie92, Zuo99c]. uncertainity [Kum92]. Uncountably [KLGP97o]. Undecidability [HS90a]. unfolding [Rou90]. unfoldings [DRSŻ91a]. ungraded [Eyr98j]. Unicité [AS93b]. unidimensionnelle [Cap98]. unifié [Del91g, Lim90]. Uniform [Fuj90a, Lor92k, Lor92l, MS90c, Mur93, Sta93, CV93, Cha91a, DT97c, GK90, Tak99, Win90, Yuk98b]. uniforme [EP90, Lép92, Val94]. Uniformly [Dix97f, KR92h, DTW90, Pis96h, Ric99c, Woj91]. unimodal [PR97j]. Unions [Mol93g, Rut97t, Mol93b, Mol93e, Mol93f, Mol93h]. unipotent [Gro97e]. unique [Kal99]. Uniqueness [BVH94b, Ebe99e, Ebe99g, KM97i, Met91s, OL90d, Soa94h, BMO98, BP90a, Ebe99d, Ebe99f, Ebe99a, GV91e, Lan96k, OL90f, Pro90b, deL94f]. Unit [KZ92e, Ger93, Sch91g]. unitarizable [Pis96h]. Unitary [Kon90, Par91b, Par91c]. united [Mic93]. units [Gre92a, Kun92]. unity [FK93b, FK93e, IS90]. Univalent [Sch90v, Sch90w]. Univariate [Bre94i, Lor92m]. Universal [Mer94, Rou90, Ser92q, Bab92, Kos95g]. Universality [Tem93, RG96h]. universe [Hun96d, Woy98h]. universelles [Rog90]. Université [Lem90c]. University [Kom93a, Kov90, OR91, RG96b, Lem90c, Ser92k]. unknown [Riv90]. unreliable [DK93a]. unsolved [Dri92]. Unstable [BC92l, MO93h, DP91a]. Upper [DK92, Del92e, Fab93b, KO91, Kni91, Mou92b, Räb91, Yap96h]. Usco [Phe93g]. used [Wah95f]. useful [Fri91, FP98j]. Using [Mic93, HS90b, LL90, RS90e, Sto92k]. USSR [BDF91, Ina90a, Zar91b]. UT [MS91a]. utilisation [Bak94, Mai94].

V [AvW90, BGV92, Kna95g]. V. [Mey93c, Sch91j]. valeurs [Arn93, Gra90b]. valid [Dix97f, Sto92f]. Valparaíso [RSSV90]. Value [Alb98g, Mat90c, Sch90d, Sch95d, Tai91d, Alb98e, CMN95, DZ93d, Fuj90a, Fur90a, Gal92, Gru92, Kaz90, KPV90, Maa91, Mar93b, Maw93, Mog92, Sch95b, Spr93b, Tai91a, Tai91c, Wah95g]. Value-distribution [Mat90c]. Valued [CM97c, CNZ91, Daw93, Gel92, Kaj91, Kit94b, Mit94a, Nar95c, Ran90b, Rus90]. values [Fuh91, Hol90a, iM90b, Nis96a, deL94b]. Vandermonde [Lor92n]. Vanishing [Mon91a, Sie91, Dem96, Kap91, Mor90c, Suf90, vW90]. Varenna [Cel90, DM93b]. variable [Ant92, Cen91, Jam90, Lan90b, Nis96e]. Variables [Alb98g, Alb98i, BKS93, CS92c, Gra90b, JPA97, Jan93, KK93, Nis96f, Sch91f, ÜZ97]. variance [Kir90]. variant [GKW91]. variants [Har98f]. Variation [CS92a, DN99d, Bra98c, Bra98g, DNQ99, DN99b, ELY97, JS94]. Variational [Bar93j, Bet99, Kru97c, Kru97k, Mie91c, Mül99, Rod94, Tib90e, DZ93h, Eke90, MT90, Mie91b, Phe93j]. Variations [HS99, DZ93j, Mar90a]. Varietäten [Har93d]. variété [Arn99, EM91, Pic91]. Variétés [Anc90, Arn93, Arn95a, Aud90, Deb92, DM90b, Lab90, MNM90]. Varieties [AH92d, AR96e, BCC92b, EE93, FP98i, Göt94d, Har93d, Hui93, HPSS92, JR92g, LO98h, Mum99a, Mum99g, Zuo99e, vG94, AR96b, Ale91, Aub92, BSW92, BL94c, FS90d, Fuj90b, FP98c, FP98g, Göt94e, Han92, Hin98, Hun96h, Ion90, JR92a, JR92d, JR92e, JR91b, KMM92, LO98o, Rei90c, Rei97d, Sch94d, Van96h, Van96f, Wal90b, Wil97a, Wil97e, Wil97f, Wil97d, Zak91, Zuo99f, vG92, Mum99d]. variety [Dan91, Deg92, DL96b, IK99a, Taf92]. various [Ebe99e, JJS92, KM97g]. Varying [Tot94e, Tot94f, BC95, Mol93h, Mol93a]. vecteurs [Mou90]. Vector [Ban91a, Brî96d, Brî96f, CT92, CM97c, DRSŻ91i, FR90c, Ric99e, Ric99h, AR96d, Bel90b, BIOS90, BK90, BS92, Box91, Brî96a, Brî96b, BK92c, Del91a, DR91, DRSŻ91a, Fuj92, GS90a, HS93, JP91b, Kit94b, Kob90, KB96g, MR98, NZ99l, Nog91, PSW92, Pfe90, Rou90, Sai91a, Son90, Tam99d, Tyu92, Yan90, Zuo99a, Nar97]. Vector-Valued [CM97c, Kit94b]. vectorization [KO90a]. Vectors [Yur95g, Nie91d, Yur95d, Yur95f, deL94e]. Venant [Mie91l]. Venezia [ZD93]. Venice [TPZ98]. Venise [BV93]. verification [Kim94]. Vermutungen [Har93c]. Vers [Deb92]. Versal [Wan91a, Chr91, Ste91]. Version [Koo93]. versus [Hub95g, Sch94l]. Vertex [Tam99b, Tam99g, Tam99c]. Vesentini [Vil90]. Vessiot [vdPS97i]. via [BW94a, DGDMN91, Gra91a, Kat93a, Lor92e, Lus90, Ôta93]. vicinity [BK92c]. Vienna [HT90a]. Vietoris [Wic91e]. view [Fur90a, Rei90a, Sam91]. Viewpoints [CC92]. Views [Wic91g]. Vilenkin [Wei94a]. virtually [Dek96e]. Viscosity [Cra97, Bar97, Son97, DL97]. viscous [Cha92, Gli92, Hey90, Kaz90, Riv90, ST92a]. vision [FM90]. Vlasov [Mél96]. Volterra [Cor91, Egg91, HW91]. volume [SS91]. vortex [Caf91, Ore90]. vorticity [Wet92]. vraisemblances [CJ90]. VS [EGI94, BCDS92].

W [Hit91, Ste90b]. W. [Eve93, Mey93a]. W.-C [Ste90b]. Wald [GN97]. walks [Bek91e, BD94, Hey93]. walls [Wol92]. Walsh [dM99]. Warwick [MM91b, RS91a]. Watanabe [AS93a, CS96, GRCW90]. water [Oka90]. Wave [KF93, Nie91c, Nie91j, Rug96, Col92g, Hel98, Kar97a, Kit93, Lie93e, Nie91e, Nie91h, Oka90]. wavelet [Chu93]. Wavelets [Dav91c, Dav91d, Lem90d, Mit94d, Ber98a, Mic93, Mit94e, Lem90c]. waves [Bab92, Boi96, Mie91a, Woy98h]. way [Dob93]. Weak [Kos95g, KP96b, KP96c, May90, Sar92, Wan91b, Xia92, AFL90, AHJLP90, Bal91, BM90d, Cor91, Hu96, pL91, Mor93a, Neu97o, Phe93j, Sch90f, Sch90j, Smi90, Wie92]. weak-Gordon [Bal91]. Weakly [AS98h, HFS94g, KO91, Vai93c, Fro91, Vai93i]. Wedge [CL99a]. wedges [HN93g]. Week [ASMS90]. Weierstrass [GL92]. Weight [Tot94f, KZ92e, Mat90b, Oda90b, Pet92, Xi94a]. Weighted [Cha91m, GE98c, Lam92, LN91, Lou91a, Pan91a, Tot94f, Fre90, Hu91, I90, LS91, Shu98b, Van91, WY92]. Weights [KZ92f, Băd90, Fre91g, Ino98e, Tot94b, Tot94c, Tot94e, Vos92]. Weil [Gol94, Kat93a, NZ99b, Sch94t, Shi91b]. Well [DZ93j, DZ93i, DZ93g, DZ93h, deL94u, DZ93a, DZ93b, DZ93c, DZ93d, DZ93f, Tam90, deL94s]. Well-Posed [DZ93g]. Well-posedness [DZ93j, DZ93i, DZ93h, deL94u, DZ93a, DZ93b, DZ93c, DZ93d, DZ93f, deL94s]. Wellposedness [XL98j, XL98h]. Weyl [Lan91c, MZ90, Nie91k, Win90, Xi94b, Xi94c, Xi94e]. where [Win95b, Win95c]. which [Ari91, GJ91, HWW93b, HWW93a, JPA97, Jan93, Kac91, Met91f]. White [Oba94e, Oba94f, Oba94g, Sch93h, Sch93g, Sch93e]. Whitehead [Rut97u]. Whitney [BHS96c]. Whitney-smooth [BHS96c]. Whittaker [Oda90a]. Wick [Nie91k]. wide [Sve93]. wide-sense [Sve93]. Width [KLGP97g, KLGP97d, KLGP97o]. widths [BT93, Par93]. Wiener [MY91, FF97, KÜ90a, Kry95, Led90, Lim90, LP91, MWNPA92, NÜZ90, Nua98, Sch99, Üst95b, Üst95e, Üst95h, ÜZ97, Wat90, Wil95, Yan91a]. Wilcoxon [FT93]. Wild [vdPS97m]. Wilson [Met91c, NM92]. winding [KB96g]. windings [BW94a, Kni93, Shi94]. without [Coa99, Fin90, Gre92d, Kum90]. Witt [Cri90]. Witten [Moo96a, Moo96b]. Work [Dwo90, Ada90]. Workshop [AvW90, BS90c, KÜ90b, ST92c, TBR91]. Workshops [Kul92]. world [Pla90]. Worst [Yuk98l]. Wreath [BMMN98q, Dũn92]. wtt [HS90a]. wtt-degrees [HS90a].

XIX [BPSH91]. XVIII [Hen90a]. XX [Hen92]. XXI [Hen93]. XXII [BGM94]. XXIII [Ber95]. XXIV [AYM90, Ber96]. XXIX [AEMY95]. XXV [Ber98b, AYM91]. XXVI [GGSC97, AYM92]. XXVII [Ber99a, AYM93]. XXVIII [AYM94]. XXX [AYE96]. XXXI [AYE97]. XXXII [AYÉL98]. XXXIII [AÉLY99].

Yan [AS90c]. Yang [Ste90b, Ge92, GR92, Mor90b]. Yangians [Naz92, Ol’92]. Years [LW90b, Kne92, Pre96, Sch92a]. Yor [Leu96, Mey94]. Yoshida [Böc90]. Yosida [Kom93a, deL94c]. Young [Val90b, DN99b].

Zariski [Mar98, Sch94s]. Zero [Göt94d, Cas92, DP91f, IK99d, IK99e, Pri93, Sch90a, Tho90a, Xue90]. Zero-Dimensional [Göt94d, IK99d, IK99e]. zeroes [Fit93, GH90]. zeros [Clu92, Fuj90a, HYL91a, VC90, SN90]. zeta [AS90b, CPS90, Fuj90a, Mat90c, Meu90, Pan91i, Rei97d]. zeta-functions [Mat90c]. Zonal [UT92]. zoom [Del91a]. Zq [BF93b]. Zq-equivariant [BF93b]. zur [Hla90]. Zygmund [Han91].

References

Arneodo:1990:TOR

[AAG90] A. Arneodo, F. Argoul, and G. Grasseau. Transformation en ondelettes et renormalisation. (French) []. Lecture Notes in Mathematics, 1438:125–191, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083519/.

Andradas:1990:NRS

[AB90a] C. Andradas and E. Becker. A note on the real spectrum of analytic functions on an analytic manifold of dimension one. Lecture Notes in Mathematics, 1420:1–21, 1990. CODEN LN-MAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083907/.

Ash:1990:GMS

[AB90b] Avner Ash and Armand Borel. Generalized modular symbols. Lecture Notes in Mathematics, 1447:57–75, 1990. CODEN LN-MAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085726/.

Arnold:1991:ANT

[AB91] Ludwig Arnold and Petra Boxler. Additive noise turns a hyperbolic fixed point into a stationary solution. Lecture Notes in Mathematics, 1486:159–164, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086665/.

Attal:1995:QFT

[ABÉH95] S. Attal, K. Burdzy, M. Émery, and Y. Hu. Sur quelques filtrations et transformations browniennes. (French) []. Lecture Notes in Mathematics, 1613:56–69, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094200/.

Abramenko:1996:FPC

[Abr96a] Peter Abramenko. Finiteness properties of classical F_{q} over F_{q}[t]. Lecture Notes in Mathematics, 1641:107–114, 1996. CODEN LNMAA2. ISBN 3-540-61973-9 (print), 3-540-49570-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094083/.

Abramenko:1996:GAT

[Abr96b] Peter Abramenko. Groups acting on twin buildings. Lecture Notes in Mathematics, 1641:11–55, 1996. CODEN LNMAA2. ISBN 3-540-61973-9 (print), 3-540-49570-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094081/.

Abramenko:1996:HP

[Abr96c] Peter Abramenko. Homotopy properties of ∮ Δ^{0}(a) ∮. Lecture Notes in Mathematics, 1641:56–106, 1996. CODEN LN-MAA2. ISBN 3-540-61973-9 (print), 3-540-49570-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094082/.

Abramenko:1996:I

[Abr96d] Peter Abramenko. Introduction. Lecture Notes in Mathematics, 1641:1–10, 1996. CODEN LNMAA2. ISBN 3-540-61973-9 (print), 3-540-49570-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094080/.

Abramenko:1996:TBA

[Abr96e] Peter Abramenko. Twin Buildings and Applications to S-Arithmetic Groups, volume 1641 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61973-9 (print), 3-540-49570-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 123 pp. LCCN QA3 .L28 no. 1641.

URL http://link.springer.com/book/10.1007/BFb0094079;
 http://www.springerlink.com/content/978-3-540-49570-

Andreatta:1992:PMC

[ABW92] M. Andreatta, E. Ballico, and J. Wiśniewski. Projective manifolds containing large linear subspaces. Lecture Notes in Mathematics, 1515:1–11, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098333/.

Arnold:1991:RDS

[AC91] Ludwig Arnold and Hans Crauel. Random dynamical systems. Lecture Notes in Mathematics, 1486:1–22, 1991. CODEN LN-MAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086654/.

Aguade:1992:ATH

[ACC92] Jaume Aguadé, Manuel Castellet, and Frederick Ronald Cohen, editors. Algebraic Topology Homotopy and Group Cohomology: Proceedings of the 1990 Barcelona Conference on Algebraic Topology, held in S. Feliu de Guíxols, Spain, June 6–12, 1990, volume 1509 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1509. URL http:/

/link.springer.com/book/10.1007/BFb0087495;   http://
www.springerlink.com/content/978-3-540-46772-4.

Accola:1994:ARSa

[Acc94a] Robert D. M. Accola. Automorphisms of Riemann surfaces, I. Lecture Notes in Mathematics, 1595:42–51, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073580/.

Accola:1994:ARSb

[Acc94b] Robert D. M. Accola. Automorphisms of Riemann surfaces, II; N(p). Lecture Notes in Mathematics, 1595:74–98, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073582/.

Accola:1994:ICS

[Acc94c] Robert D. M. Accola. The inequality of Castelnuovo–Severi. Lecture Notes in Mathematics, 1595:20–29, 1994. CODEN LN-MAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073578/.

Accola:1994:RSB

[Acc94d] Robert D. M. Accola. Review of some basic concepts in the theory of Riemann surfaces. Lecture Notes in Mathematics, 1595: 1–12, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073576/.

Accola:1994:SBC

[Acc94e] Robert D. M. Accola. Smooth and branched coverings of Riemann surfaces. Lecture Notes in Mathematics, 1595:30–41, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073579/.

Accola:1994:SEP

[Acc94f] Robert D. M. Accola. Some exceptional points on Riemann surfaces. Lecture Notes in Mathematics, 1595:13–19, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073577/.

Accola:1994:TTR

[Acc94g] Robert D. M. Accola. Topics in the Theory of Riemann Surfaces, volume 1595 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 105 pp. LCCN QA3 .L28 no. 1595. URL http:

//link.springer.com/book/10.1007/BFb0073575;   http:/
/www.springerlink.com/content/978-3-540-49056-2.

Accola:1994:WFP

[Acc94h] Robert D. M. Accola. When are fixed points of automorphisms exceptional in some other sense? Lecture Notes in Mathematics, 1595:52–73, 1994. CODEN LNMAA2. ISBN 3-540-58721-7 (print), 3-540-49056-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073581/.

Arnold:1991:LEP

[ACE91] Ludwig Arnold, Hans Crauel, and Jean-Pierre Eckmann, editors. Lyapunov Exponents: Proceedings of a Conference held in Oberwolfach, May 28–June 2, 1990, volume 1486 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1486.

URL http://link.springer.com/book/10.1007/BFb0086653;
 http://www.springerlink.com/content/978-3-540-46431-
0.

Arbarello:1990:GAK

[AD90] Enrico Arbarello and Corrado De Concini. Geometrical aspects of the Kadomtsev–Petviashvili equation. Lecture Notes in Mathematics, 1451:95–137, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085066/.

Adams:1990:TTW

[Ada90] J. F. Adams. Talk on Toda’s work. Lecture Notes in Mathematics, 1418:7–14, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0083688/;   http://link.springer.com/chapter/
10.1007/BFb0083689/.

Adem:1992:GCF

[Ade92] Alejandro Adem. On the geometry and cohomology of finite simple groups. Lecture Notes in Mathematics, 1509:1–9, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087496/.

Axelsson:1990:ARP

[AE90] O. Axelsson and V. Eijkhout. Analysis of a recursive 5-point/9-point factorization method. Lecture Notes in Mathematics, 1457: 154–173, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090907/.

Aikawa:1996:PTS

[AE96] Hiroaki Aikawa and Matts Essén. Potential Theory — Selected Topics, volume 1633 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61583-0 (print), 3-540-69991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 200 pp. LCCN QA3 .L28 no. 1633. URL http:

//link.springer.com/book/10.1007/BFb0093410;   http:/
/www.springerlink.com/content/978-3-540-69991-0.

Azema:1999:SPX

[AÉLY99] Jacques Azéma, Michel Émery, Michel Ledoux, and Marc Yor, editors. Séminaire de Probabilités XXXIII, volume 1709 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1709.

URL http://link.springer.com/book/10.1007/BFb0096508;
 http://www.springerlink.com/content/978-3-540-48407-
3.

Azema:1995:SPX

[AEMY95] Jacques Azéma, Michel Emery, Paul André Meyer, and Marc Yor, editors. Séminaire de Probabilités XXIX, volume 1613 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1613.

URL http://link.springer.com/book/10.1007/BFb0094193;
 http://www.springerlink.com/content/978-3-540-44744-
3.

Allouche:1990:AIF

[AFF90] J.-P. Allouche, P. Flajolet, and M. Mendes France. Algebraically independent formal power series: a language theory interpretation. Lecture Notes in Mathematics, 1434:11–18, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097121/.

Accardi:1990:QLE

[AFL90] Luigi Accardi, Alberto Frigerio, and Yun-Gang Lu. Quantum Langevin equation in the weak coupling limit. Lecture Notes in Mathematics, 1442:1–16, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085499/.

Alekseev:1992:HQG

[AFSTS92] A. Alekseev, L. Faddeev, and M. Semenov-Tian-Shansky. Hidden quantum groups inside Kac–Moody algebras. Lecture Notes in Mathematics, 1510:148–158, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101187/.

Alvarez-Gaume:1990:TCF

[AG90] L. Alvarez-Gaumé. Topics in conformal field theory and string theory. Lecture Notes in Mathematics, 1451:49–94, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085065/.

Alvarez-Gaume:1990:FM

[AGADH90] Luis Alvarez-Gaumé, Enrico Arbarello, Corrado De Concini, and Nigel J. Hitchin. Front matter. Lecture Notes in Mathematics, 1451:??, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46813-4/1.pdf.

Accardi:1990:QSF

[AH90] L. Accardi and R. L. Hudson. Quantum stochastic flows and non abelian cohomology. Lecture Notes in Mathematics, 1442: 54–69, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085501/.

Adleman:1992:A

[AH92a] Leonard M. Adleman and Ming-Deh A. Huang. Acknowledgement. Lecture Notes in Mathematics, 1512:4, 1992. CODEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090187.

Adleman:1992:I

[AH92b] Leonard M. Adleman and Ming-Deh A. Huang. Introduction. Lecture Notes in Mathematics, 1512:1–3, 1992. CODEN LN-MAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090186/.

Adleman:1992:OAP

[AH92c] Leonard M. Adleman and Ming-Deh A. Huang. Overview of the algorithm and the proof of the main theorem. Lecture Notes in Mathematics, 1512:5–14, 1992. CODEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090188/.

Adleman:1992:PTA

[AH92d] Leonard M. Adleman and Ming-Deh A. Huang. Primality Testing and Abelian Varieties Over Finite Fields, volume 1512 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 142 pp. LCCN QA3 .L28 no. 1512. URL http:/

/link.springer.com/book/10.1007/BFb0090185;   http://
www.springerlink.com/content/978-3-540-47021-2.

Adleman:1992:PPa

[AH92e] Leonard M. Adleman and Ming-Deh A. Huang. Proof of proposition 1. Lecture Notes in Mathematics, 1512:21–109, 1992. CODEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090190/.

Adleman:1992:PPb

[AH92f] Leonard M. Adleman and Ming-Deh A. Huang. Proof of proposition 2. Lecture Notes in Mathematics, 1512:110–125, 1992. CO-DEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090191/.

Adleman:1992:PPc

[AH92g] Leonard M. Adleman and Ming-Deh A. Huang. Proof of proposition 3. Lecture Notes in Mathematics, 1512:126–136, 1992. CO-DEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090192/.

Adleman:1992:RMT

[AH92h] Leonard M. Adleman and Ming-Deh A. Huang. Reduction of main theorem to three propositions. Lecture Notes in Mathematics, 1512:15–21, 1992. CODEN LNMAA2. ISBN 3-540-55308-8 (print), 3-540-47021-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090189/.

Al-Huzali:1990:WRI

[AHJLP90] Abdullah Al-Huzali, S. K. Jain, and S. R. López-Permouth. On the weak relative-injectivity of rings and modules. Lecture Notes in Mathematics, 1448:93–98, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091254/.

Aikawa:1996:PTP

[Aik96] Hiroaki Aikawa. Potential theory part II. Lecture Notes in Mathematics, 1633:103–200, 1996. CODEN LNMAA2. ISBN 3-540-61583-0 (print), 3-540-69991-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093412/.

Aizenberg:1994:C

[Aiz94] L. A. Aizenberg. C^{n}. Lecture Notes in Mathematics, 1574:339– 382, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101067/.

Azema:1996:PCT

[AJK^{+}96] J. Azéma, T. Jeulin, F. Knight, G. Mokobodzki, and M. Yor. Sur les processus croissants de type injectif. (French) []. Lecture Notes in Mathematics, 1626:312–343, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094656/.

Azema:1993:TAF

[AJKY93] J. Azéma, T. Jeulin, F. Knight, and M. Yor. Le théorème d’arrêt en une fin d’ensemble prévisible. (French) []. Lecture Notes in Mathematics, 1557:133–158, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087971/.

Arnold:1995:FM

[AJMR95] Ludwig Arnold, Christopher K. R. T. Jones, Konstantin Mischaikow, and Geneviève Raugel. Front matter. Lecture Notes in Mathematics, 1609:??, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-49415-7/1.pdf.

Axelsson:1990:PCG

[AK90] Owe Axelsson and Lily Yu. Kolotilina, editors. Preconditioned Conjugate Gradient Methods: Proceedings of a Conference held in Nijmegen, The Netherlands, June 19–21, 1989, volume 1457 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1457. URL http:/

/link.springer.com/book/10.1007/BFb0090897;   http://
www.springerlink.com/content/978-3-540-46746-5.

Akbulut:1992:CSR

[AK92] S. Akbulut and H. King. Constructing strange real algebraic sets. Lecture Notes in Mathematics, 1524:120–127, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084612/.

Akiyama:1990:CST

[Aki90] Shigeki Akiyama. On a certain sum of traces of Hecke operators. Lecture Notes in Mathematics, 1434:1–10, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097120/.

Accardi:1990:LDL

[AL90a] L. Accardi and Yun Gang Lu. On the low density limit of boson models. Lecture Notes in Mathematics, 1442:17–53, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085500/.

Axelsson:1990:IMD

[AL90b] O. Axelsson and W. Layton. Iteration method as discretization procedures. Lecture Notes in Mathematics, 1457:174–193, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090908/.

Albert:1990:GSM

[Alb90] Claude Albert, editor. Géométrie Symplectique et Mécanique: Colloque International La Grande Motte, France, 23–28 Mai, 1988, volume 1416 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1416. URL http:/

/link.springer.com/book/10.1007/BFb0097461;   http://
www.springerlink.com/content/978-3-540-46920-9.

Alber:1998:ATE

[Alb98a] Hans-Dieter Alber. Application of the theory to engineering models. Lecture Notes in Mathematics, 1682:117–136, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096281/.

Alber:1998:CC

[Alb98b] Hans-Dieter Alber. Classification conditions. Lecture Notes in Mathematics, 1682:75–97, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096279/.

Alber:1998:CEM

[Alb98c] Hans-Dieter Alber. Constitutive equations of monotone type and generalized standard materials. Lecture Notes in Mathematics, 1682:23–44, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096276/.

Alber:1998:ESC

[Alb98d] Hans-Dieter Alber. Existence of solutions for constitutive equations of monotone type. Lecture Notes in Mathematics, 1682: 45–56, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096277/.

Alber:1998:IBV

[Alb98e] Hans-Dieter Alber. Initial-boundary value problems for the inelastic behavior of metals. Lecture Notes in Mathematics, 1682: 7–22, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096275/.

Alber:1998:I

[Alb98f] Hans-Dieter Alber. Introduction. Lecture Notes in Mathematics, 1682:1–5, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096274/.

Alber:1998:MMI

[Alb98g] Hans-Dieter Alber. Materials with Memory: Initial-Boundary Value Problems for Constitutive Equations with Internal Variables, volume 1682 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 166 pp. LCCN QA3 .L28 no. 1682. URL http:

//link.springer.com/book/10.1007/BFb0096273;   http:/
/www.springerlink.com/content/978-3-540-69689-6.

Alber:1998:OPR

[Alb98h] Hans-Dieter Alber. Open problems and related results. Lecture Notes in Mathematics, 1682:137–142, 1998. CODEN LN-MAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096282/.

Alber:1998:TIV

[Alb98i] Hans-Dieter Alber. Transformation of interior variables. Lecture Notes in Mathematics, 1682:57–73, 1998. CODEN LN-MAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096278/.

Alber:1998:TRI

[Alb98j] Hans-Dieter Alber. Transformation of rate independent constitutive equations. Lecture Notes in Mathematics, 1682:99–116, 1998. CODEN LNMAA2. ISBN 3-540-64066-5 (print), 3-540-69689-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096280/.

Alexeev:1991:TAG

[Ale91] V. A. Alexeev. Theorems about good divisors on log Fano varieties (case of index r > (n − 2)). Lecture Notes in Mathematics, 1479:1–9, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086258/.

Arkeryd:1993:FM

[ALMV93] Leif Arkeryd, Pierre-Louis Lions, Peter A. Markowich, and Srinivasa R. S. Varadhan. Front matter. Lecture Notes in Mathematics, 1551:??, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-47832-4/1.pdf.

Alperin:1990:LAF

[Alp90] J. L. Alperin. A Lie approach to finite groups. Lecture Notes in Mathematics, 1456:1–9, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100726/.

An-Min:1991:ACE

[AM91] Li An-Min. Affine completeness and Euclidean completeness. Lecture Notes in Mathematics, 1481:115–125, 1991. CODEN LN-MAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083635/.

Amann:1991:HBQ

[Ama91] Herbert Amann. Hopf bifurcation in quasilinear reaction-diffusion systems. Lecture Notes in Mathematics, 1475:53–63, 1991. CO-DEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083479/.

Amghibech:1998:CRE

[Amg98] S. Amghibech. Criteria of regularity at the end of a tree. Lecture Notes in Mathematics, 1686:128–136, 1998. CODEN LN-MAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101755/.

Araujo:1990:NBS

[AMM90] Jesús Araujo and J. Martinez-Maurica. The nonarchimedean Banach–Stone theorem. Lecture Notes in Mathematics, 1454: 64–79, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091134/.

Arrondo:1990:PSC

[AMS90] Enrique Arrondo, Raquel Mallavibarrena, and Ignacio Sols. Proof of Schubert’s conjectures on double contacts. Lecture Notes in Mathematics, 1436:1–29, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084039/.

Adams:1991:SLS

[AMSV91] Malcolm R. Adams, Clint McCrory, Theodore Shifrin, and Robert Varley. Symmetric Lagrangian singularities and Gauss maps of theta divisors. Lecture Notes in Mathematics, 1462:1–26, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086371/.

Ancona:1990:TDP

[Anc90] A. Ancona. Théorie du potentiel sur les graphes et les variétés. (French) []. Lecture Notes in Mathematics, 1427:1–112, 1990. CODEN LNMAA2. ISBN 3-540-53508-X (print), 3-540-46718-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103041/.

Andre:1990:ABL

[And90] Yves André. p-adic Betti lattices. Lecture Notes in Mathematics, 1454:23–63, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091133/.

Andrzejewski:1991:EFO

[And91] Pawel Andrzejewski. Equivariant finiteness obstruction and its geometric applications — a survey. Lecture Notes in Mathematics, 1474:20–37, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084735/.

Anick:1990:LHT

[Ani90] David J. Anick. R-local homotopy theory. Lecture Notes in Mathematics, 1418:78–85, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083694/.

Anile:1991:MIR

[Ani91] A. M. Anile. Modeling intense relativistic electron beams. Lecture Notes in Mathematics, 1460:1–14, 1991. CODEN LN-MAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091357/.

Anshel:1990:TRG

[Ans90] Iris Lee Anshel. On two relator groups. Lecture Notes in Mathematics, 1440:1–21, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084445/.

Ansel:1994:RPA

[Ans94] Jean-Pascal Ansel. Remarques sur le prix des actifs contingents. (French) []. Lecture Notes in Mathematics, 1583:181–188, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073846/.

Antanovskii:1992:AFB

[Ant92] Leonid K. Antanovskii. Analyticity of a free boundary in plane quasi-steady flow of a liquid form subject to variable surface tension. Lecture Notes in Mathematics, 1530:1–16, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090330/.

Alekseevskii:1990:CLI

[AP90] D. V. Alekseevskii and B. A. Putko. On the completeness of left-invariant pseudo-Riemannian metrics on Lie groups. Lecture Notes in Mathematics, 1453:171–185, 1990. CODEN LN-MAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085954/.

Allday:1991:SAS

[AP91] C. Allday and V. Puppe. Some applications of shifted subgroups in transformation groups. Lecture Notes in Mathematics, 1474: 1–19, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084734/.

Abate:1994:CFG

[AP94a] Marco Abate and Giorgio Patrizio. Complex Finsler geometry. Lecture Notes in Mathematics, 1591:63–125, 1994. CODEN LN-MAA2. ISBN 3-540-58465-X (print), 3-540-48812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073982/.

Abate:1994:FMG

[AP94b] Marco Abate and Giorgio Patrizio. Finsler Metrics — a Global Approach: with applications to geometric function theory, volume 1591 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58465-X (print), 3-540-48812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 180 pp. LCCN QA3 .L28 no. 1591. URL http:/

/link.springer.com/book/10.1007/BFb0073980;   http://
www.springerlink.com/content/978-3-540-48812-5.

Abate:1994:MCH

[AP94c] Marco Abate and Giorgio Patrizio. Manifolds with constant holomorphic curvature. Lecture Notes in Mathematics, 1591:127– 170, 1994. CODEN LNMAA2. ISBN 3-540-58465-X (print), 3-540-48812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073983/.

Abate:1994:RFG

[AP94d] Marco Abate and Giorgio Patrizio. Real Finsler geometry. Lecture Notes in Mathematics, 1591:1–62, 1994. CODEN LN-MAA2. ISBN 3-540-58465-X (print), 3-540-48812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073981/.

Applebaum:1990:QDI

[App90] David Applebaum. Quantum diffusions on involutive algebras. Lecture Notes in Mathematics, 1442:70–85, 1990. CODEN LN-MAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085502/.

Applebaum:1995:HLP

[App95] David Applebaum. A horizontal Lévy process on the bundle of orthonormal frames over a complete Reimannian manifold. Lecture Notes in Mathematics, 1613:166–180, 1995. CODEN LN-MAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094209/.

Aptekarev:1993:DPT

[Apt93] A. I. Aptekarev. A direct proof for Trefethen’s conjecture. Lecture Notes in Mathematics, 1550:147–148, 1993. CODEN LN-MAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117479.

Adamek:1991:WLG

[AR91] Jiří Adámek and Jiří Rosický. What are locally generated categories? Lecture Notes in Mathematics, 1488:14–19, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084209/.

Andradas:1992:MBS

[AR92] Carlos Andradas and Jesús M. Ruíz. More on basic semialgebraic sets. Lecture Notes in Mathematics, 1524:128–139, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084613/.

Azema:1994:ESF

[AR94] Jacques Azéma and Catherine Rainer. Sur l’équation de structure d[X, X]_{t} = dt − X_{t−}^{+}dX_{t}. (French) []. Lecture Notes in Mathematics, 1583:236–255, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073849/.

Adler:1996:GAF

[AR96a] Allan Adler and Sundararaman Ramanan. Geometry and arithmetic of the fundamental relations. Lecture Notes in Mathematics, 1644:52–76, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093664/.

Adler:1996:HGL

[AR96b] Allan Adler and Sundararaman Ramanan. Heisenberg groups of line bundles on abelian varieties. Lecture Notes in Mathematics, 1644:18–30, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093662/.

Adler:1996:I

[AR96c] Allan Adler and Sundararaman Ramanan. Introduction. Lecture Notes in Mathematics, 1644:1–7, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093660/.

Adler:1996:ITA

[AR96d] Allan Adler and Sundararaman Ramanan. Invariant theory, arithmetic and vector bundles. Lecture Notes in Mathematics, 1644: 77–106, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093665/.

Adler:1996:MAV

[AR96e] Allan Adler and Sundararaman Ramanan. Moduli of Abelian Varieties, volume 1644 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 196 pp. LCCN QA3 .L28 no. 1644. URL http:/

/link.springer.com/book/10.1007/BFb0093659;   http://
www.springerlink.com/content/978-3-540-49609-0.

Adler:1996:SHG

[AR96f] Allan Adler and Sundararaman Ramanan. Standard Heisenberg groups. Lecture Notes in Mathematics, 1644:8–17, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093661/.

Adler:1996:TSA

[AR96g] Allan Adler and Sundararaman Ramanan. Theta structures and the addition formula. Lecture Notes in Mathematics, 1644:31– 51, 1996. CODEN LNMAA2. ISBN 3-540-62023-0 (print), 3-540-49609-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093663/.

Arapura:1991:FMF

[Ara91] Donu Arapura. Fano maps and fundamental groups. Lecture Notes in Mathematics, 1479:10–14, 1991. CODEN LN-MAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086259/.

Arbarello:1999:SBS

[Arb99] Enrico Arbarello. Supplementary bibliography on the Schottky problem. Lecture Notes in Mathematics, 1358:301–304, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46021-3_7/.

Arino:1991:MSF

[Ari91] Ovide Arino. Monotone semi-flows which have a monotone first integral. Lecture Notes in Mathematics, 1475:64–75, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083480/.

Arkowitz:1990:GSH

[Ark90a] Martin Arkowitz. The group of self-homotopy equivalences — a survey. Lecture Notes in Mathematics, 1425:170–203, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083840/.

Arkowitz:1990:WHS

[Ark90b] Martin Arkowitz. When is the homotopy set [X, Y] infinite? Lecture Notes in Mathematics, 1440:22–26, 1990. CODEN LN-MAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084446/.

Arkeryd:1993:SEN

[Ark93] L. Arkeryd. Some examples of NSA methods in kinetic theory. Lecture Notes in Mathematics, 1551:14–57, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090928/.

Arnaudon:1992:CMG

[Arn92] Marc Arnaudon. Connexions et martingales dans les groupes de Lie. (French) []. Lecture Notes in Mathematics, 1526:146– 156, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084317/.

Arnaudon:1993:PAS

[Arn93] Marc Arnaudon. Propriétés asymptotiques des semi-martingales à valeurs dans des variétés à bord continu. (French) []. Lecture Notes in Mathematics, 1557:182–201, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087975/.

Arnaudon:1995:BCA

[Arn95a] Marc Arnaudon. Barycentres convexes et approximations des martingales continues dans les variétés. (French) []. Lecture Notes in Mathematics, 1613:70–85, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094201/.

Arnold:1995:RDS

[Arn95b] Ludwig Arnold. Random dynamical systems. Lecture Notes in Mathematics, 1609:1–43, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095238/.

Arnaudon:1999:AEP

[Arn99] M. Arnaudon. Appendice à l’exposé précédent: La filtration naturelle du mouvement brownien indexé par R dans une variété compacte. (French) []. Lecture Notes in Mathematics, 1709:304– 314, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096519/.

Arslanov:1990:SDB

[Ars90] Marat M. Arslanov. On the structure of degrees below O′. Lecture Notes in Mathematics, 1432:23–32, 1990. CODEN LN-MAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086111/.

Azema:1996:PMP

[ARY96] J. Azéma, C. Rainer, and M. Yor. Une propriété des martingales pures. (French) []. Lecture Notes in Mathematics, 1626: 243–254, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094652/.

Adolphson:1990:AEE

[AS90a] Alan Adolphson and Steven Sperber. p-adic estimates for exponential sums. Lecture Notes in Mathematics, 1454:11–22, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091132/.

Allouche:1990:SDH

[AS90b] Jean-Paul Allouche and Jeffrey Shallit. Sums of digits and the Hurwitz zeta function. Lecture Notes in Mathematics, 1434:19– 30, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097122/.

Ansel:1990:QRT

[AS90c] Jean-Pascal Ansel and Christophe Stricker. Quelques remarques sur un théorème de Yan. (French) []. Lecture Notes in Mathematics, 1426:266–274, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083769/.

Arora:1992:TPS

[AS92] S. C. Arora and Sharda Sharma. On two-parameter semigroup of operators. Lecture Notes in Mathematics, 1511:147–153, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093808/.

Ansel:1993:DKW

[AS93a] J. P. Ansel and C. Stricker. Décomposition de Kunita–Watanabe. (French) []. Lecture Notes in Mathematics, 1557:30–32, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087960/.

Ansel:1993:UEL

[AS93b] J. P. Ansel and C. Stricker. Unicité et existence de la loi minimale. (French) []. Lecture Notes in Mathematics, 1557:22–29, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087959/.

Assing:1998:BCP

[AS98a] Sigurd Assing and Wolfgang M. Schmidt. Basic concepts and preparatory results. Lecture Notes in Mathematics, 1688:1– 13, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096152/.

Assing:1998:CPS

[AS98b] Sigurd Assing and Wolfgang M. Schmidt. Classification of the points of the state space. Lecture Notes in Mathematics, 1688: 15–25, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096153/.

Assing:1998:CCS

[AS98c] Sigurd Assing and Wolfgang M. Schmidt. Construction of continuous strong Markov processes. Lecture Notes in Mathematics, 1688:79–102, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096157/.

Assing:1998:CSMb

[AS98d] Sigurd Assing and Wolfgang M. Schmidt. Continuous Strong Markov Processes in Dimension One: A stochastic calculus approach, volume 1688 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 137 pp. LCCN QA3 .L28 no. 1688. URL http:/

/link.springer.com/book/10.1007/BFb0096151;   http://
www.springerlink.com/content/978-3-540-69786-2.

Assing:1998:CSMa

[AS98e] Sigurd Assing and Wolfgang M. Schmidt. Continuous strong Markov semimartingales as solutions of stochastic differential equations. Lecture Notes in Mathematics, 1688:103–118, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096158/.

Assing:1998:OTF

[AS98f] Sigurd Assing and Wolfgang M. Schmidt. Occupation time formula. Lecture Notes in Mathematics, 1688:53–77, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096156/.

Assing:1998:SDC

[AS98g] Sigurd Assing and Wolfgang M. Schmidt. Semimartingale decomposition of continuous strong Markov semimartingales. Lecture Notes in Mathematics, 1688:33–52, 1998. CODEN LN-MAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096155/.

Assing:1998:WAF

[AS98h] Sigurd Assing and Wolfgang M. Schmidt. Weakly additive functionals and time change of strong Markov processes. Lecture Notes in Mathematics, 1688:27–32, 1998. CODEN LNMAA2. ISBN 3-540-64465-2 (print), 3-540-69786-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096154/.

Ambos-Spies:1990:HPR

[ASHY90] Klaus Ambos-Spies, Steven Homer, and Dongping Yang. Honest polynomial reductions and exptally sets. Lecture Notes in Mathematics, 1432:1–22, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086110/.

Ambos-Spies:1990:RTW

[ASMS90] Klaus Ambos-Spies, Gert H. Müller, and Gerald E. Sacks, editors. Recursion Theory Week: Proceedings of a Conference held in Oberwolfach, FRG, March 19–25, 1989, volume 1432 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1432.

URL http://link.springer.com/book/10.1007/BFb0086109;
 http://www.springerlink.com/content/978-3-540-47142-
4.

Aston:1991:SLB

[Ast91] P. J. Aston. Scaling laws and bifurcation. Lecture Notes in Mathematics, 1463:1–21, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085423/.

Arnaudon:1998:SSD

[AT98] Marc Arnaudon and Anton Thalmaier. Stability of stochastic differential equations in manifolds. Lecture Notes in Mathematics, 1686:188–214, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101758/.

Atiyah:1990:HMG

[Ati90] Michael Atiyah. Hyperkähler manifolds. Lecture Notes in Mathematics, 1422:1–13, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089400/.

Aubry:1992:RMC

[Aub92] Yves Aubry. Reed–Muller codes associated to projective algebraic varieties. Lecture Notes in Mathematics, 1518:4–17, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087988/.

Audin:1990:HPV

[Aud90] Michèle Audin. Hamiltoniens périodiques sur les variétés symplectiques compactes de dimension 4. (French) []. Lecture Notes in Mathematics, 1416:1–25, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097462/.

Avritzer:1990:HP

[AV90] D. Avritzer and I. Vainsencher. Hilb^{4}P^{2}. Lecture Notes in Mathematics, 1436:30–59, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084040/.

Albrecht:1991:POP

[AV91] Felix Albrecht and Gabriele Villari. Periodic orbits of planar polynomial Liénard systems with a small parameter. Lecture Notes in Mathematics, 1475:41–52, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083478/.

Anderson:1992:CIQ

[AVV92] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Conformal invariants, quasiconformal maps, and special functions. Lecture Notes in Mathematics, 1508:1–19, 1992. CODEN LN-MAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094235/.

Accardi:1990:QPA

[AvW90] Luigi Accardi and Wilhelm von Waldenfels, editors. Quantum Probability and Applications V: Proceedings of the Fourth Workshop, held in Heidelberg, FRG, Sept. 26–30, 1988, volume 1442 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1442. URL http:/

/link.springer.com/book/10.1007/BFb0085497;   http://
www.springerlink.com/content/978-3-540-46311-5.

Azais:1997:OPS

[AW97] Jean-Marc Azaïs and Mario Wschebor. Oscillation presque sûre de martingales continues. (French) []. Lecture Notes in Mathematics, 1655:69–76, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0119293/.

Azema:1990:DPR

[AY90] J. Azéma and M. Yor. Dérivation par rapport au processus de Bessel. (French) []. Lecture Notes in Mathematics, 1426:210– 226, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083767/.

Azema:1996:SPX

[AYE96] Jacques Azéma, Marc Yor, and Michel Emery, editors. Séminaire de Probabilités XXX, volume 1626 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1626. URL http:/

/link.springer.com/book/10.1007/BFb0094636;   http://
www.springerlink.com/content/978-3-540-68463-3.

Azema:1997:SPX

[AYE97] Jacques Azéma, Marc Yor, and Michel Emery, editors. Séminaire de Probabilités XXXI, volume 1655 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1655. URL http:/

/link.springer.com/book/10.1007/BFb0119286;   http://
www.springerlink.com/content/978-3-540-68352-0.

Azema:1998:SPX

[AYÉL98] Jacques Azéma, Marc Yor, Michel Émery, and Michel Ledoux, editors. Séminaire de Probabilités XXXII, volume 1686 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1686.

URL http://link.springer.com/book/10.1007/BFb0101744;
 http://www.springerlink.com/content/978-3-540-69762-
6.

Azema:1990:SPX

[AYM90] Jacques Azéma, Marc Yor, and Paul André Meyer, editors. Séminaire de Probabilités XXIV 1988/89, volume 1426 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1426.

URL http://link.springer.com/book/10.1007/BFb0083752;
 http://www.springerlink.com/content/978-3-540-47098-
4.

Azema:1991:SPX

[AYM91] Jaques Azéma, Marc Yor, and Paul André Meyer, editors. Séminaire de Probabilités XXV, volume 1485 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1485.

URL http://link.springer.com/book/10.1007/BFb0100839;
 http://www.springerlink.com/content/978-3-540-38496-
0.

Azema:1992:SPX

[AYM92] Jacques Azéma, Marc Yor, and Paul André Meyer, editors. Séminaire de Probabilités XXVI, volume 1526 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1526.

URL http://link.springer.com/book/10.1007/BFb0084305;
 http://www.springerlink.com/content/978-3-540-47342-
8.

Azema:1993:SPX

[AYM93] Jacques Azéma, Marc Yor, and Paul André Meyer, editors. Séminaire de Probabilités XXVII, volume 1557 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1557.

URL http://link.springer.com/book/10.1007/BFb0087956;
 http://www.springerlink.com/content/978-3-540-48034-
1.

Azema:1994:SPX

[AYM94] Jacques Azéma, Marc Yor, and Paul André Meyer, editors. Séminaire de Probabilités XXVIII, volume 1583 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1583.

URL http://link.springer.com/book/10.1007/BFb0073829;
 http://www.springerlink.com/content/978-3-540-48656-
5.

Azumaya:1990:LSS

[Azu90] Goro Azumaya. Locally split submodules and modules with perfect endomorphism rings. Lecture Notes in Mathematics, 1448: 1–6, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091245/.

Bardelli:1994:ACH

[BA94] Fabio Bardelli and Alberto Albano, editors. Algebraic Cycles and Hodge Theory: Lectures given at the 2nd Session of the Centro Internationale Matematico Estivo (C.I.M.E.) held in Torino, Italy, June 21-29, 1993, volume 1594 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1594. URL

http://link.springer.com/book/10.1007/b52495;   http:
//www.springerlink.com/content/978-3-540-49046-3.

Babelon:1992:LTL

[Bab92] O. Babelon. Liouville theory on the lattice and universal exchange algebra for Bloch waves. Lecture Notes in Mathematics, 1510:159–175, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101188/.

Badescu:1990:IDN

[Băd90] Lucian Bădescu. Infinitesimal deformations of negative weights and hyperplane sections. Lecture Notes in Mathematics, 1417: 1–22, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083329/.

Baesens:1991:NED

[Bae91] Claude Baesens. Noise effect on dynamic bifurcations: The case of a period-doubling cascade. Lecture Notes in Mathematics, 1493: 107–130, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085027/.

Bakry:1991:ISF

[Bak91] Dominique Bakry. Inégalités de Sobolev faibles: un critère Γ_{2}. (French) []. Lecture Notes in Mathematics, 1485:234–261, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100859/.

Bakry:1994:LSU

[Bak94] Dominique Bakry. L’hypercontractivité et son utilisation en théorie des semigroupes. (French) []. Lecture Notes in Mathematics, 1581:1–114, 1994. CODEN LNMAA2. ISBN 3-540-58208-8 (print), 3-540-48568-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073872/.

Baldi:1990:LDF

[Bal90a] Paolo Baldi. Large deviations and the functional Lévy’s modulus for invariant diffusions. Lecture Notes in Mathematics, 1444: 193–203, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083616/.

Ballico:1990:SPS

[Bal90b] Edoardo Ballico. On k-spanned projective surfaces. Lecture Notes in Mathematics, 1417:23, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083330.

Ball:1991:NSW

[Bal91] Keith Ball. Normed spaces with a weak-Gordon–Lewis property. Lecture Notes in Mathematics, 1470:36–47, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090210/.

Ballico:1993:AAG

[Bal93] Edoardo Ballico, editor. Arithmetic Algebraic Geometry: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Trento, Italy, June 24– July 2, 1991, volume 1553 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1553. URL http:/

/link.springer.com/book/10.1007/BFb0084727;   http://
www.springerlink.com/content/978-3-540-47909-3.

Balser:1994:AO

[Bal94a] Werner Balser. Acceleration operators. Lecture Notes in Mathematics, 1582:41–52, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073569/.

Balser:1994:APS

[Bal94b] Werner Balser. Asymptotic power series. Lecture Notes in Mathematics, 1582:1–12, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073565/.

Balser:1994:CHT

[Bal94c] Werner Balser. Cauchy–Heine transform. Lecture Notes in Mathematics, 1582:33–40, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073568/.

Balser:1994:FSN

[Bal94d] Werner Balser. Formal solutions to non-linear ODE. Lecture Notes in Mathematics, 1582:83–101, 1994. CODEN LN-MAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073572/.

Balser:1994:DPS

[Bal94e] Werner Balser. From Divergent Power Series to Analytic Functions: Theory and Application of Multisummable Power Series, volume 1582 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 108 pp. LCCN QA3 .L28 no. 1582. URL http:

//link.springer.com/book/10.1007/BFb0073564;   http:/
/www.springerlink.com/content/978-3-540-48594-0.

Balser:1994:LBT

[Bal94f] Werner Balser. Laplace and Borel transforms. Lecture Notes in Mathematics, 1582:13–22, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073566/.

Balser:1994:MPS

[Bal94g] Werner Balser. Multisummable power series. Lecture Notes in Mathematics, 1582:53–74, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073570/.

Balser:1994:SED

[Bal94h] Werner Balser. Some equivalent definitions of multisummability. Lecture Notes in Mathematics, 1582:75–81, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073571/.

Balser:1994:SPS

[Bal94i] Werner Balser. Summable power series. Lecture Notes in Mathematics, 1582:23–32, 1994. CODEN LNMAA2. ISBN 3-540-58268-1 (print), 3-540-48594-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073567/.

Banaszczyk:1991:AST

[Ban91a] Wojciech Banaszczyk. Additive Subgroups of Topological Vector Spaces, volume 1466 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 178 pp. LCCN QA3 .L28 no. 1466. URL http:/

/link.springer.com/book/10.1007/BFb0089147;   http://
www.springerlink.com/content/978-3-540-46396-2.

Banaszczyk:1991:BT

[Ban91b] Wojciech Banaszczyk. The Bochner theorem. Lecture Notes in Mathematics, 1466:110–131, 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089151/.

Banaszczyk:1991:EG

[Ban91c] Wojciech Banaszczyk. Exotic groups. Lecture Notes in Mathematics, 1466:45–71, 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089149/.

Banaszczyk:1991:NG

[Ban91d] Wojciech Banaszczyk. Nuclear groups. Lecture Notes in Mathematics, 1466:72–109, 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089150/.

Banaszczyk:1991:PD

[Ban91e] Wojciech Banaszczyk. Pontryagin duality. Lecture Notes in Mathematics, 1466:132–167, 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089152/.

Banaszczyk:1991:P

[Ban91f] Wojciech Banaszczyk. Preliminaries. Lecture Notes in Mathematics, 1466:1–44, 1991. CODEN LNMAA2. ISBN 3-540-53917-4 (print), 3-540-46396-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089148/.

Bandyopadhyaya:1992:EPP

[Ban92] Pradipta Bandyopadhyaya. Exposed points and points of continuity in closed bounded convex sets. Lecture Notes in Mathematics, 1511:120–127, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093805/.

Barchielli:1990:SMS

[Bar90a] Alberto Barchielli. Some Markov semigroups in quantum probability. Lecture Notes in Mathematics, 1442:86–98, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085503/.

Barnard:1990:OPC

[Bar90b] Roger W. Barnard. Open problems and conjectures in complex analysis. Lecture Notes in Mathematics, 1435:1–26, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087893/.

Baron:1990:PSD

[Bar90c] G. Baron. On point sets with differences of distances not less than the minimum distance. Lecture Notes in Mathematics, 1452: 1–5, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096976/.

Bardelli:1992:ACC

[Bar92] Fabio Bardelli. Algebraic cohomology classes on some special threefolds. Lecture Notes in Mathematics, 1515:12–20, 1992. CO-DEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098334/.

Bartsch:1993:BEP

[Bar93a] Thomas Bartsch. Bifurcation for O(3)-equivariant problems. Lecture Notes in Mathematics, 1560:113–126, 1993. CODEN LN-MAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073867/.

Bartsch:1993:CGI

[Bar93b] Thomas Bartsch. Category and genus of infinite-dimensional representation spheres. Lecture Notes in Mathematics, 1560:30– 52, 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073862/.

Bartsch:1993:CGC

[Bar93c] Thomas Bartsch. Category, genus and critical point theory with symmetries. Lecture Notes in Mathematics, 1560:8–29, 1993. CO-DEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073861/.

Bartsch:1993:EL

[Bar93d] Thomas Bartsch. The exit-length. Lecture Notes in Mathematics, 1560:96–112, 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073866/.

Bartsch:1993:I

[Bar93e] Thomas Bartsch. Introduction. Lecture Notes in Mathematics, 1560:1–7, 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073860/.

Bartsch:1993:LCI

[Bar93f] Thomas Bartsch. The length and Conley index theory. Lecture Notes in Mathematics, 1560:86–95, 1993. CODEN LN-MAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073865/.

Bartsch:1993:LS

[Bar93g] Thomas Bartsch. The length of G-spaces. Lecture Notes in Mathematics, 1560:53–71, 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073863/.

Bartsch:1993:LRS

[Bar93h] Thomas Bartsch. The length of representation spheres. Lecture Notes in Mathematics, 1560:72–85, 1993. CODEN LN-MAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073864/.

Bartsch:1993:MPS

[Bar93i] Thomas Bartsch. Multiple periodic solutions near equilibria of symmetric Hamiltonian systems. Lecture Notes in Mathematics, 1560:127–141, 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073868/.

Bartsch:1993:TMV

[Bar93j] Thomas Bartsch. Topological Methods for Variational Problems with Symmetries, volume 1560 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57378-X (print), 3-540-48099-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 152 pp. LCCN QA3 .L28 no. 1560.

URL http://link.springer.com/book/10.1007/BFb0073859;
 http://www.springerlink.com/content/978-3-540-48099-
0.

Bardi:1997:SAV

[Bar97] Martino Bardi. Some applications of viscosity solutions to optimal control and differential games. Lecture Notes in Mathematics, 1660:44–97, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094295/.

Barlow:1998:DF

[Bar98] Martin T. Barlow. Diffusions on fractals. Lecture Notes in Mathematics, 1690:1–121, 1998. CODEN LNMAA2. ISBN 3-540-64620-5 (print), 3-540-69228-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092537/.

Bass:1990:PAB

[Bas90] Richard F. Bass. A probabilistic approach to the boundedness of singular integral operators. Lecture Notes in Mathematics, 1426: 15–40, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083754/.

Baues:1991:HCM

[Bau91] Hans Joachim Baues. On the homotopy category of Moore spaces and an old result of Barratt. Lecture Notes in Mathematics, 1474: 207–230, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084748/.

Baxendale:1991:IMN

[Bax91] Peter H. Baxendale. Invariant measures for nonlinear stochastic differential equations. Lecture Notes in Mathematics, 1486: 123–140, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086663/.

Baxter:1992:MPB

[Bax92] Martin Baxter. Markov processes on the boundary of the binary tree. Lecture Notes in Mathematics, 1526:210–224, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084323/.

Borwein:1990:RCM

[BB90] J. M. Borwein and P. B. Borwein. A remarkable cubic mean iteration. Lecture Notes in Mathematics, 1435:27–31, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087894/.

Baldassarri:1990:AAP

[BBD90] Francesco Baldassarri, Siegfried Bosch, and Bernard Dwork, editors. p-adic Analysis: Proceedings of the International Conference held in Trento, Italy, May 29–June 2, 1989, volume 1454 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1454. URL http:/

/link.springer.com/book/10.1007/BFb0091130;   http://
www.springerlink.com/content/978-3-540-46906-3.

Blazic:1991:SGL

[BBG91] Novica Blažić, Neda Bokan, and Peter Gilkey. The spectral geometry of the Laplacian and the conformal Laplacian for manifolds with boundary. Lecture Notes in Mathematics, 1481:5–17, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083623/.

Berrick:1992:GSA

[BC92a] A. J. Berrick and Carles Casacuberta. Groups and spaces with all localizations trivial. Lecture Notes in Mathematics, 1509:20– 29, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087498/.

Block:1992:CNC

[BC92b] Louis Stuart Block and William Andrew Coppel. Chaotic and non-chaotic maps. Lecture Notes in Mathematics, 1513:121– 166, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084769/.

Block:1992:DOD

[BC92c] Louis Stuart Block and William Andrew Coppel. Dynamics in One Dimension, volume 1513 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 234 (est.) pp. LCCN QA3 .L28 no. 1513.

URL http://link.springer.com/book/10.1007/BFb0084762;
 http://www.springerlink.com/content/978-3-540-47023-
6.

Block:1992:I

[BC92d] Louis Stuart Block and William Andrew Coppel. Introduction. Lecture Notes in Mathematics, 1513:1–3, 1992. CODEN LN-MAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084763/.

Block:1992:MC

[BC92e] Louis Stuart Block and William Andrew Coppel. Maps of the circle. Lecture Notes in Mathematics, 1513:219–234, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084772/.

Block:1992:PO

[BC92f] Louis Stuart Block and William Andrew Coppel. Periodic orbits. Lecture Notes in Mathematics, 1513:5–23, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084764/.

Block:1992:TD

[BC92g] Louis Stuart Block and William Andrew Coppel. Topological dynamics. Lecture Notes in Mathematics, 1513:69–89, 1992. CO-DEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084767/.

Block:1992:TDC

[BC92h] Louis Stuart Block and William Andrew Coppel. Topological dynamics (continued). Lecture Notes in Mathematics, 1513:91– 119, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084768/.

Block:1992:TE

[BC92i] Louis Stuart Block and William Andrew Coppel. Topological entropy. Lecture Notes in Mathematics, 1513:189–218, 1992. CO-DEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084771/.

Block:1992:T

[BC92j] Louis Stuart Block and William Andrew Coppel. Turbulence. Lecture Notes in Mathematics, 1513:25–46, 1992. CODEN LN-MAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084765/.

Block:1992:TPO

[BC92k] Louis Stuart Block and William Andrew Coppel. Types of periodic orbits. Lecture Notes in Mathematics, 1513:167–188, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084770/.

Block:1992:UMH

[BC92l] Louis Stuart Block and William Andrew Coppel. Unstable manifolds and homoclinic points. Lecture Notes in Mathematics, 1513: 47–67, 1992. CODEN LNMAA2. ISBN 3-540-55309-6 (print), 3-540-47023-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084766/.

Bogenschutz:1992:ARF

[BC92m] Thomas Bogenschütz and Hans Crauel. The Abramov–Rokhlin formula. Lecture Notes in Mathematics, 1514:32–35, 1992. CO-DEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097526/.

Bertoin:1995:RGS

[BC95] Jean Bertoin and Ma.-Emilia Caballero. On the rate of growth of subordinators with slowly varying Laplace exponent. Lecture Notes in Mathematics, 1613:125–132, 1995. CODEN LN-MAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094205/.

Ballico:1992:OP

[BCC92a] E. Ballico, C. Ciliberto, and F. Catanese. Open problems. Lecture Notes in Mathematics, 1515:140–146, 1992. CODEN LN-MAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098343/.

Ballico:1992:CIV

[BCC92b] Edoardo Ballico, Fabrizio Catanese, and Ciro Ciliberto, editors. Classification of Irregular Varieties: Minimal Models and Abelian Varieties. Proceedings of a Conference held in Trento, Italy, 17–21 December, 1990, volume 1515 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1515. URL http:/

/link.springer.com/book/10.1007/BFb0098332;   http://
www.springerlink.com/content/978-3-540-47016-8.

Braden:1992:ATF

[BCDS92] H. W. Braden, E. Corrigan, P. E. Dorey, and R. Sasaki. Affine Toda field theory: S-matrix vs perturbation. Lecture Notes in Mathematics, 1510:210–220, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101191/.

Bardi:1997:FM

[BCE^{+}97] Martino Bardi, Michael G. Crandall, Lawrence C. Evans, Halil Mete Soner, and Panagiotis E. Souganidis. Front matter. Lecture Notes in Mathematics, 1660:??, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-69043-6/1.pdf.

Bujalance:1990:RPC

[BCG90] Emilio Bujalance, Antonio F. Costa, and J. M. Gamboa. Real parts of complex algebraic curves. Lecture Notes in Mathematics, 1420:81–110, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083913/.

Bramley:1990:SPP

[BCMS90] R. Bramley, H.-C. Chen, U. Meier, and A. Sameh. On some parallel preconditioned CG schemes. Lecture Notes in Mathematics, 1457:17–27, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090899/.

Bertoin:1994:CRW

[BD94] J. Bertoin and R. A. Doney. On conditioning random walks in an exponential family to stay nonnegative. Lecture Notes in Mathematics, 1583:116–121, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073840/.

Bloch:1991:AGP

[BDF91] Spencer Bloch, Igor. V. Dolgachev, and William Fulton, editors. Algebraic Geometry: Proceedings of the US–USSR Symposium held in Chicago, June 20–July 14, 1989, volume 1479 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1479.

URL http://link.springer.com/book/10.1007/BFb0086257;
 http://www.springerlink.com/content/978-3-540-38388-
8.

Boillat:1996:FM

[BDLL96] Guy Boillat, Constantin M. Dafermos, Peter D. Lax, and Tai-Ping Liu. Front matter. Lecture Notes in Mathematics, 1640: ??, 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-49565-9/1.pdf.

Burg:1990:SNR

[BDT90] Thomas B. Burg, Michael Drmota, and Robert F. Tichy. Some new results in summability theory. Lecture Notes in Mathematics, 1452:20–30, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096978/.

Bakry:1996:IGF

[BE96] Dominique Bakry and Mireille Echerbault. Sur les inégalités GKS. (French) []. Lecture Notes in Mathematics, 1626:178–206, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094648/.

Beauwens:1990:MIF

[Bea90] R. Beauwens. Modified incomplete factorization strategies. Lecture Notes in Mathematics, 1457:1–16, 1990. CODEN LN-MAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090898/.

Beauville:1992:ADP

[Bea92] Arnaud Beauville. Annulation du H^{1} pour les fibrés en droites plats. (French) []. Lecture Notes in Mathematics, 1507:1–15, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094507/.

Bellissard:1994:FM

[BEF^{+}94] Jean Bellissard, Mirko Degli Esposti, Giovanni Forni, Sandro Graffi, Stefano Isola, and John N. Mather. Front matter. Lecture Notes in Mathematics, 1589:??, 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-48782-1/1.pdf.

Bujalance:1990:AGCa

[BEGG90a] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. The automorphism group of compact Klein surfaces with one boundary component. Lecture Notes in Mathematics, 1439:138–152, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084983/.

Bujalance:1990:AGH

[BEGG90b] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. The automorphism group of hyperelliptic compact Klein surfaces with boundary. Lecture Notes in Mathematics, 1439:153–164, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084984/.

Bujalance:1990:AGCb

[BEGG90c] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Automorphism Groups of Compact Bordered Klein Surfaces: A Combinatorial Approach, volume 1439 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 201 pp. LCCN QA3 .L28 no. 1439. URL http:/

/link.springer.com/book/10.1007/BFb0084977;   http://
www.springerlink.com/content/978-3-540-47180-6.

Bujalance:1990:CGA

[BEGG90d] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Cyclic groups of automorphisms of compact Klein surfaces. Lecture Notes in Mathematics, 1439:60–97, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084981/.

Bujalance:1990:KSO

[BEGG90e] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Klein surfaces as orbit spaces of NEC groups. Lecture Notes in Mathematics, 1439:21–37, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084979/.

Bujalance:1990:KSG

[BEGG90f] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Klein surfaces with groups of automorphisms in prescribed families. Lecture Notes in Mathematics, 1439:98–137, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084982/.

Bujalance:1990:NNS

[BEGG90g] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Normal NEC subgroups of NEC groups. Lecture Notes in Mathematics, 1439:38–59, 1990. CODEN LN-MAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084980/.

Bujalance:1990:PR

[BEGG90h] Emilio Bujalance, José Javier Etayo, José Manuel Gamboa, and Grzegorz Gromadzki. Preliminary results. Lecture Notes in Mathematics, 1439:1–20, 1990. CODEN LNMAA2. ISBN 3-540-52941-1 (print), 3-540-47180-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084978/.

Behnke:1991:IDM

[Beh91] Kurt Behnke. On infinitesimal deformations of minimally elliptic singularities. Lecture Notes in Mathematics, 1462:27–41, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086372/.

Bekka:1991:RTT

[Bek91a] K. Bekka. C-régularité et trivialité topologique. (French) []. Lecture Notes in Mathematics, 1462:42–62, 1991. CODEN LN-MAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086373/.

Bekkali:1991:A

[Bek91b] Mohamed Bekkali. Appendix. Lecture Notes in Mathematics, 1476:105–114, 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098403/.

Bekkali:1991:FA

[Bek91c] Mohamed Bekkali. Forcing axioms. Lecture Notes in Mathematics, 1476:23–60, 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098401/.

Bekkali:1991:M

[Bek91d] Mohamed Bekkali. Measurability in L[R]. Lecture Notes in Mathematics, 1476:15–22, 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098400/.

Bekkali:1991:MMW

[Bek91e] Mohamed Bekkali. The method of minimal walks. Lecture Notes in Mathematics, 1476:61–104, 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098402/.

Bekkali:1991:NSR

[Bek91f] Mohamed Bekkali. Nonmeasurable sets of reals. Lecture Notes in Mathematics, 1476:1–13, 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098399/.

Bekkali:1991:TST

[Bek91g] Mohamed Bekkali. Topics in Set Theory: Lebesgue Measurability, Large Cardinals, Forcing Axioms, Rho-functions, volume 1476 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54121-7 (print), 3-540-47422-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 120 pp. LCCN QA3 .L28 no. 1476. URL http:/

/link.springer.com/book/10.1007/BFb0098398;   http://
www.springerlink.com/content/978-3-540-47422-7.

Barlow:1998:ATT

[BÉK^{+}98] M. T. Barlow, M. Émery, F. B. Knight, S. Song, and M. Yor. Autour d’un théorème de Tsirelson sur des filtrations browniennes et non browniennes. (French) []. Lecture Notes in Mathematics, 1686:264–305, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101763/.

Belavkin:1990:QSC

[Bel90a] Viacheslav Belavkin. A quantum stochastic calculus in Fock space of input and output nondemolition processes. Lecture Notes in Mathematics, 1442:99–125, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085504/.

Belopolskaya:1990:SOP

[Bel90b] Ya. I. Belopol’skaya. Second-order parabolic equations in principal fibre bundles and associated vector bundles. Lecture Notes in Mathematics, 1453:187–200, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085955/.

Bertram:1991:SGM

[BEL91] A. Bertram, Lawrence Ein, and Robert Lazarsfeld. Surjectivity of Gaussian maps for line bundles of large degree on curves. Lecture Notes in Mathematics, 1479:15–25, 1991. CODEN LN-MAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086260/.

Bellissard:1994:NCM

[Bel94] Jean Bellissard. Non commutative methods in semiclassical analysis. Lecture Notes in Mathematics, 1589:1–64, 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074074/.

Belili:1999:DDP

[Bel99] Nacereddine Belili. Dualité du problème des marges et ses applications. (French) []. Lecture Notes in Mathematics, 1709:371– 387, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096527/.

Benabou:1991:SRF

[Ben91a] Jean Benabou. Some remarks on free monoids in a topos. Lecture Notes in Mathematics, 1488:20–29, 1991. CODEN LN-MAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084210/.

Benoit:1991:DBP

[Ben91b] Eric Benoît, editor. Dynamic Bifurcations: Proceedings of a Conference held in Luminy, France, March 5–10, 1990, volume 1493 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1493. URL http:/

/link.springer.com/book/10.1007/BFb0085019;   http://
www.springerlink.com/content/978-3-540-46471-6.

Benoit:1991:LDB

[Ben91c] Eric Benoit. Linear dynamic bifurcation with noise. Lecture Notes in Mathematics, 1493:131–150, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085028/.

Benson:1992:RPD

[Ben92] D. J. Benson. Resolutions and Poincaré duality for finite groups. Lecture Notes in Mathematics, 1509:10–19, 1992. CODEN LN-MAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087497/.

Benaim:1999:DSA

[Ben99] Michel Benaïm. Dynamics of stochastic approximation algorithms. Lecture Notes in Mathematics, 1709:1–68, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096509/.

Berthelot:1990:CRT

[Ber90a] Pierre Berthelot. Cohomologie rigide et théorie des D-modules. (French) []. Lecture Notes in Mathematics, 1454:80–124, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091135/.

Bertoin:1990:HFP

[Ber90b] Jean Bertoin. Sur une horloge fluctuante pour les processus de Bessel de petites dimensions. (French) []. Lecture Notes in Mathematics, 1426:117–136, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083761/.

Bertrand:1990:EMG

[Ber90c] D. Bertrand. Extensions de D-modules et groupes de Galois différentiels. (French) []. Lecture Notes in Mathematics, 1454: 125–141, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091136/.

Bertrand:1990:TMA

[Ber90d] Daniel Bertrand. Transcendental methods in arithmetic geometry. Lecture Notes in Mathematics, 1434:31–44, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097123/.

Bertoin:1993:PSD

[Ber93] Jean Bertoin. Une preuve simple du théorème de Shimura sur les points méandre du mouvement brownien plan. (French) []. Lecture Notes in Mathematics, 1557:33–35, 1993. CODEN LN-MAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087961/.

Bernard:1995:LPT

[Ber95] Pierre Bernard, editor. Lectures on Probability Theory: École d’Été de Probabilités de Saint-Flour XXIII — 1993, volume 1608 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60015-9 (print), 3-540-49402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1608. URL http:/

/link.springer.com/book/10.1007/BFb0095745;   http://
www.springerlink.com/content/978-3-540-49402-7.

Bernard:1996:LPT

[Ber96] Pierre Bernard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXIV — 1994, volume 1648 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62055-9 (print), 3-540-49635-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1648. URL http:/

/link.springer.com/book/10.1007/BFb0095673;   http://
www.springerlink.com/content/978-3-540-49635-9.

Berenstein:1998:RTW

[Ber98a] Carlos Berenstein. Randon transforms, wavelets, and applications. Lecture Notes in Mathematics, 1684:1–33, 1998. CODEN LNMAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096089/.

Bernard:1998:LPT

[Ber98b] Pierre Bernard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXV — 1995, volume 1690 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64620-5 (print), 3-540-69228-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1690. URL http:/

/link.springer.com/book/10.1007/BFb0092536;   http://
www.springerlink.com/content/978-3-540-69228-7.

Bernard:1999:LPT

[Ber99a] Pierre Bernard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probailités de Saint-Flour XXVII — 1997, volume 1717 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66593-5 (print), 3-540-48115-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1717. URL http:/

/link.springer.com/book/10.1007/b72002;   http://www.
springerlink.com/content/978-3-540-48115-7.

Bertoin:1999:SEA

[Ber99b] Jean Bertoin. Subordinators: Examples and applications. Lecture Notes in Mathematics, 1717:1–91, 1999. CODEN LNMAA2.

ISBN 3-540-66593-5 (print), 3-540-48115-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48115-7_1/.

Bethuel:1999:VMG

[Bet99] F. Bethuel. Variational methods for Ginzburg–Landau equations. Lecture Notes in Mathematics, 1713:1–43, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092668/.

Beukers:1993:DEA

[Beu93] Frits Beukers. Diophantine equations and approximation. Lecture Notes in Mathematics, 1566:1–11, 1993. CODEN LN-MAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48208-6_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
540-48208-6/1.pdf.

Bridges:1993:CDE

[BF93a] Thomas J. Bridges and Jacques E. Furter. Classification of dq-equivariant gradient bifurcation problems. Lecture Notes in Mathematics, 1558:89–99, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073477/.

Bridges:1993:CZE

[BF93b] Thomas J. Bridges and Jacques E. Furter. Classification of zq-equivariant gradient bifurcation problems. Lecture Notes in Mathematics, 1558:63–83, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073475/.

Bridges:1993:CMR

[BF93c] Thomas J. Bridges and Jacques E. Furter. Collision of multipliers at rational points for symplectic maps. Lecture Notes in Mathematics, 1558:149–174, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073480/.

Bridges:1993:EMC

[BF93d] Thomas J. Bridges and Jacques E. Furter. Equivariant maps and the collision of multipliers. Lecture Notes in Mathematics, 1558:175–183, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073481/.

Bridges:1993:GBP

[BF93e] Thomas J. Bridges and Jacques E. Furter. Generic bifurcation of periodic points. Lecture Notes in Mathematics, 1558:9–31, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073473/.

Bridges:1993:I

[BF93f] Thomas J. Bridges and Jacques E. Furter. Introduction. Lecture Notes in Mathematics, 1558:1–8, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073472/.

Bridges:1993:PPG

[BF93g] Thomas J. Bridges and Jacques E. Furter. Period-3 points of the generalized standard map. Lecture Notes in Mathematics, 1558: 85–88, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073476/.

Bridges:1993:PPE

[BF93h] Thomas J. Bridges and Jacques E. Furter. Periodic points of equivariant symplectic maps. Lecture Notes in Mathematics, 1558: 119–147, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073479/.

Bridges:1993:RDB

[BF93i] Thomas J. Bridges and Jacques E. Furter. Reversibility and degenerate bifurcation of period-q points of multiparameter maps. Lecture Notes in Mathematics, 1558:101–118, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073478/.

Bridges:1993:STEb

[BF93j] Thomas J. Bridges and Jacques E. Furter. Singularity Theory and Equivariant Symplectic Maps, volume 1558 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 226 pp. LCCN QA3 .L28 no. 1558. URL http:/

/link.springer.com/book/10.1007/BFb0073471;   http://
www.springerlink.com/content/978-3-540-48040-2.

Bridges:1993:STEa

[BF93k] Thomas J. Bridges and Jacques E. Furter. Singularity theory for equivariant gradient bifurcation problems. Lecture Notes in Mathematics, 1558:33–62, 1993. CODEN LNMAA2. ISBN 3-540-57296-1 (print), 3-540-48040-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073474/.

Bezrukavnikov:1998:A

[BFS98a] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Acknowledgement. Lecture Notes in Mathematics, 1691: 5, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092345.

Bezrukavnikov:1998:CSQ

[BFS98b] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Configuration spaces and quantum groups. Lecture Notes in Mathematics, 1691:71–121, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092348/.

Bezrukavnikov:1998:FSQ

[BFS98c] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Factorizable Sheaves and Quantum Groups, volume 1691 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 287 pp. LCCN QA3 .L28 no. 1691. URL http:/

/link.springer.com/book/10.1007/BFb0092343;   http://
www.springerlink.com/content/978-3-540-69231-7.

Bezrukavnikov:1998:ICR

[BFS98d] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Intersection cohomology of real arrangements. Lecture Notes in Mathematics, 1691:50–70, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092347/.

Bezrukavnikov:1998:I

[BFS98e] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Introduction. Lecture Notes in Mathematics, 1691:1–4, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092344/.

Bezrukavnikov:1998:L

[BFS98f] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Localization on P^{1}. Lecture Notes in Mathematics, 1691: 173–196, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092350/.

Bezrukavnikov:1998:MSC

[BFS98g] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Modular structure on the category FS. Lecture Notes in Mathematics, 1691:197–251, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092351/.

Bezrukavnikov:1998:O

[BFS98h] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Overview. Lecture Notes in Mathematics, 1691:6–49, 1998. CODEN LNMAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092346/.

Bezrukavnikov:1998:TCA

[BFS98i] Roman Bezrukavnikov, Michael Finkelberg, and Vadim Schechtman. Tensor categories arising from configuration spaces. Lecture Notes in Mathematics, 1691:122–172, 1998. CODEN LN-MAA2. ISBN 3-540-64619-1 (print), 3-540-69231-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092349/.

Bakry:1994:LPT

[BGM94] Dominique Bakry, Richard D. Gill, and Stanislav A. Molchanov, editors. Lectures on Probability Theory: École d’Été de Probabilités de Saint-Flour XXII — 1992, volume 1581 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58208-8 (print), 3-540-48568-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1581.

URL http://link.springer.com/book/10.1007/BFb0073871;
 http://www.springerlink.com/content/978-3-540-48568-
1.

Ball:1991:NIP

[BGR91] Joseph A. Ball, Israel Gohberg, and Leiba Rodman. Nehari interpolation problem for rational matrix functions: The generic case. Lecture Notes in Mathematics, 1496:277–308, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084472/.

Borisovich:1990:GAS

[BGV90] Yurii G. Borisovich, Yurii E. Gliklikh, and A. M. Vershik. Global Analysis — Studies and Applications IV, volume 1453 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 320 pp. LCCN QA3 .L28 no. 1453. URL http:/

/link.springer.com/book/10.1007/BFb0085943;   http://
www.springerlink.com/content/978-3-540-46861-5.

Borisovich:1992:GAS

[BGV92] Yuri G. Borisovich, Yuri E. Gliklikh, and A. M. Vershik. Global Analysis — Studies and Applications V, volume 1520 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 284 pp. LCCN QA3 .L28 no. 1520. URL http:/

/link.springer.com/book/10.1007/BFb0084711;   http://
www.springerlink.com/content/978-3-540-47223-0.

Birkenmeier:1990:PIR

[BH90a] Gary Birkenmeier and Henry Heatherly. Permutation identity rings and the medial radical. Lecture Notes in Mathematics, 1448: 125–138, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091257/.

Booth:1990:G

[BH90b] P. I. Booth and P. R. Heath. On the group ϵ(X ×Y) and ϵ_{B}^{B}(X ×_{B} Y). Lecture Notes in Mathematics, 1425:17–31, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083827/.

Bouleau:1990:SRL

[BH90c] Nicolas Bouleau and Francis Hirsch. Some results on Lipschitzian stochastic differential equations by Dirichlet forms methods. Lecture Notes in Mathematics, 1444:128–140, 1990. CODEN LN-MAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083613/.

Bethuel:1999:FM

[BHMS99] Fabrice Bethuel, Gerhard Huisken, Stefan Müller, and Klaus Steffen. Front matter. Lecture Notes in Mathematics, 1713: ??, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-48813-2/1.pdf.

Broer:1996:A

[BHS96a] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Appendices. Lecture Notes in Mathematics, 1645:141–167, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49613-7_6/.

Broer:1996:BM

[BHS96b] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Back matter. Lecture Notes in Mathematics, 1645:169–195, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-49613-7/1.pdf.

Broer:1996:CWS

[BHS96c] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Complicated Whitney-smooth families. Lecture Notes in Mathematics, 1645:83–121, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-49613-7_4/.

Broer:1996:C

[BHS96d] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Conclusions. Lecture Notes in Mathematics, 1645:123–139, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49613-7_5/.

Broer:1996:CTa

[BHS96e] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. The conjugacy theory. Lecture Notes in Mathematics, 1645:41– 75, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49613-7_2/.

Broer:1996:CTb

[BHS96f] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. The continuation theory. Lecture Notes in Mathematics, 1645:

77–82, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49613-7_3/.

Broer:1996:FM

[BHS96g] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Front matter. Lecture Notes in Mathematics, 1645: i–xi, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-49613-7/1.pdf.

Broer:1996:IE

[BHS96h] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Introduction and examples. Lecture Notes in Mathematics, 1645: 1–40, 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49613-7_1/.

Broer:1996:QPM

[BHS96i] Hendrik W. Broer, George B. Huitema, and Mikhail B. Sevryuk. Quasi-Periodic Motions in Families of Dynamical Systems: Order amidst Chaos, volume 1645 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62025-7 (print), 3-540-49613-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 196 pp. LCCN QA3 .L28 no.

1645. URL http://link.springer.com/book/10.1007/978-
3-540-49613-7;   http://www.springerlink.com/content/
978-3-540-49613-7.

Barlow:1992:SSP

[BI92] Martin T. Barlow and Peter Imkeller. On some sample path properties of Skorohod integral processes. Lecture Notes in Mathematics, 1526:70–80, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084311/.

Biagioni:1990:GFA

[Bia90a] Hebe Azevedo Biagioni. Generalized functions on an arbitrary subset of E_{n}. Lecture Notes in Mathematics, 1421:69–82, 1990.

CODEN LNMAA2. ISBN 0-387-52408-8, 3-540-52408-8, 3-540-46981-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089554/.

Biagioni:1990:GFO

[Bia90b] Hebe Azevedo Biagioni. Generalized functions on an open subset of E_{n}. Lecture Notes in Mathematics, 1421:1–68, 1990. CODEN LNMAA2. ISBN 0-387-52408-8, 3-540-52408-8, 3-540-46981-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089553/.

Biagioni:1990:GSN

[Bia90c] Hebe Azevedo Biagioni. Generalized solutions of nonlinear partial differential equations. Lecture Notes in Mathematics, 1421: 83–141, 1990. CODEN LNMAA2. ISBN 0-387-52408-8, 3-540-52408-8, 3-540-46981-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089555/.

Biagioni:1990:NTG

[Bia90d] Hebe Azevedo Biagioni. A Nonlinear Theory of Generalized Functions, volume 1421 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1990. CODEN LNMAA2. ISBN 0-387-52408-8, 3-540-52408-8, 3-540-46981-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 214 pp. LCCN QA3 .L28 no. 1421; QA3 .L28 no. 1421; QA299.6-433; QA324 .B53 1990. URL http:

//link.springer.com/book/10.1007/BFb0089552.

Biancofiore:1990:HSR

[Bia90e] Aldo Biancofiore. On the hyperplane sections of ruled surfaces. Lecture Notes in Mathematics, 1417:52–66, 1990. CODEN LN-MAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083332/.

Biane:1992:FMD

[Bia92] Philippe Biane. Frontière de Martin du dual de SU(2). (French) []. Lecture Notes in Mathematics, 1526:225–233, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084324/.

Biane:1995:CSN

[Bia95a] Philippe Biane. Calcul stochastique non-commutatif. (French) []. Lecture Notes in Mathematics, 1608:1–96, 1995. CODEN LNMAA2. ISBN 3-540-60015-9 (print), 3-540-49402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095746/.

Biane:1995:IMS

[Bia95b] Philippe Biane. Intertwining of Markov semi-groups, some examples. Lecture Notes in Mathematics, 1613:30–36, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094197/.

Bhatwadekar:1990:SAV

[BIOS90] Shrikant M. Bhatwadekar, Friedrich Ischebeck, Manuel Ojanguren, and Gerhard Schabhüser. Strongly algebraic vector bundles over R^{d}. Lecture Notes in Mathematics, 1420:36–41, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083909/.

Birman:1994:PTS

[Bir94] M. Sh. Birman. Perturbation theory scattering theory. Lecture Notes in Mathematics, 1573:259–292, 1994. CODEN LN-MAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100207/.

Bjork:1991:DC

[Bjö91] Jan-Erik Björk. Derived categories. Lecture Notes in Mathematics, 1478:67–129, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083502/.

Bjork:1997:IRT

[Bjö97] Tomas Björk. Interest rate theory. Lecture Notes in Mathematics, 1656:53–122, 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091999/.

Bochnak:1990:VBR

[BK90] J. Bochnak and W. Kucharz. On vector bundles and real algebraic morphisms. Lecture Notes in Mathematics, 1420:65–71, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083911/.

Bauer:1991:SAH

[BK91] Ingrid Bauer and Siegmund Kosarew. Some aspects of Hodge theory on non-complete algebraic manifolds. Lecture Notes in Mathematics, 1468:281–316, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086198/.

Bandt:1992:SDA

[BK92a] Christoph Bandt and Karsten Keller. Symbolic dynamics for angle-doubling on the circle I. The topology of locally connected Julia sets. Lecture Notes in Mathematics, 1514:1–23, 1992. CO-DEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097524/.

Bass:1992:SCC

[BK92b] Richard Bass and Davar Khoshnevisan. Stochastic calculus and the continuity of local times of Lévy processes. Lecture Notes in Mathematics, 1526:1–10, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084306/.

Bronstein:1992:FSN

[BK92c] I. U. Bronsteĭn and A. Ya. Kopanskiĭ. Finitely smooth normal forms of vector fields in the vicinity of a rest point. Lecture Notes in Mathematics, 1520:157–172, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084720/.

Basmaji:1994:TF

[BK94] Jacques Basmaji and Ian Kiming. A table of A_{5}-fields. Lecture Notes in Mathematics, 1585:37–46, 1994. CODEN LN-MAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074108/.

Berenstein:1993:ITS

[BKS93] C. A. Berenstein, T. Kawai, and D. C. Struppa. Interpolation theorems in several complex variables and applications. Lecture Notes in Mathematics, 1540:1–9, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085470/.

Bruzzo:1990:GSC

[BL90] Ugo Bruzzo and Giovanni Landi. Geometry of standard constraints and anomalous supersymmetric gauge theories. Lecture Notes in Mathematics, 1451:138–147, 1990. CODEN LN-MAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085067/.

Bellomo:1991:ATB

[BL91a] N. Bellomo and M. Lachowicz. On the asymptotic theory of the Boltzmann and Enskog equations a rigorous H-theorem for the Enskog equation. Lecture Notes in Mathematics, 1460:15– 30, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091358/.

Bourgain:1991:CSB

[BL91b] J. Bourgain and J. Lindenstrauss. On convering a set in R^{N} by balls of the same diameter. Lecture Notes in Mathematics, 1469:138–144, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089220/.

Bernard:1992:NLC

[BL92a] Denis Bernard and André Leclair. Non-local currents in 2D QFT: an alternative to the quantum inverse scattering method. Lecture Notes in Mathematics, 1510:176–196, 1992. CODEN LN-MAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101189/.

Biedenharn:1992:IRT

[BL92b] L. C. Biedenharn and M. A. Lohe. Induced representations and tensor operators for quantum groups. Lecture Notes in Mathematics, 1510:197–209, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101190/.

Birkenhake:1992:NEA

[BL92c] Ch. Birkenhake and H. Lange. Norm-endomorphisms of abelian subvarieties. Lecture Notes in Mathematics, 1515:21–32, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098335/.

Bokor:1992:MEN

[BL92d] Imre Bokor and Irene Llerena. More examples of non-cancellation in homotopy. Lecture Notes in Mathematics, 1509:30–34, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087499/.

Bagby:1993:BTH

[BL93a] Thomas Bagby and Norman Levenberg. Bernstein theorems for harmonic functions. Lecture Notes in Mathematics, 1550:7– 18, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0117470/.

Bernstein:1993:GNF

[BL93b] Daniel J. Bernstein and A. K. Lenstra. A general number field sieve implementation. Lecture Notes in Mathematics, 1554:103– 126, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091541/.

Bernstein:1994:DCF

[BL94a] Joseph Bernstein and Valery Lunts. Derived category D_{G}(X) and functors. Lecture Notes in Mathematics, 1578:2–67, 1994. CODEN LNMAA2. ISBN 3-540-58071-9 (print), 3-540-48430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073551/.

Bernstein:1994:DME

[BL94b] Joseph Bernstein and Valery Lunts. Dg-modules and equivariant cohomology. Lecture Notes in Mathematics, 1578:68–125, 1994. CODEN LNMAA2. ISBN 3-540-58071-9 (print), 3-540-48430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073552/.

Bernstein:1994:ECT

[BL94c] Joseph Bernstein and Valery Lunts. Equivariant cohomology of toric varieties. Lecture Notes in Mathematics, 1578:126–132, 1994. CODEN LNMAA2. ISBN 3-540-58071-9 (print), 3-540-48430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073553/.

Bernstein:1994:ESF

[BL94d] Joseph Bernstein and Valery Lunts. Equivariant Sheaves and Functors, volume 1578 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58071-9 (print), 3-540-48430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 139 pp. LCCN QA3 .L28 no. 1578. URL http:/

/link.springer.com/book/10.1007/BFb0073549;   http://
www.springerlink.com/content/978-3-540-48430-1.

Bernstein:1994:I

[BL94e] Joseph Bernstein and Valery Lunts. Introduction. Lecture Notes in Mathematics, 1578:1, 1994. CODEN LNMAA2. ISBN 3-540-58071-9 (print), 3-540-48430-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073550.

Bielefeld:1994:HD

[BL94f] B. Bielefeld and M. Lyubich. Holomorphic dynamics. Lecture Notes in Mathematics, 1574:423–462, 1994. CODEN LN-MAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101069/.

Bielefeld:1994:MP

[BL94g] B. Bielefeld and M. Lyubich. Miscellaneous problems. Lecture Notes in Mathematics, 1574:463–482, 1994. CODEN LN-MAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101070/.

Blais:1991:AER

[Bla91] François Blais. Asymptotic expansions of rivers. Lecture Notes in Mathematics, 1493:181–189, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085031/.

Bleher:1991:QCE

[Ble91] P. M. Bleher. Quasi-classical expansions and the problem of quantum chaos. Lecture Notes in Mathematics, 1469:60–89, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089215/.

Bullen:1990:NIP

[BLM^{+}90] Peter S. Bullen, Peng Yee Lee, Jean L. Mawhin, Patrick Muldowney, and Washek F. Pfeffer, editors. New Integrals: Proceedings of the Henstock Conference held in Coleraine, Northern Ireland, August 9–12, 1988, volume 1419 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1419. URL http:/

/link.springer.com/book/10.1007/BFb0083092;   http://
www.springerlink.com/content/978-3-540-46955-1.

Blokh:1992:SDP

[Blo92] A. M. Blokh. Spectral decomposition, periods of cycles and a conjecture of M. Misiurewicz for graph maps. Lecture Notes in Mathematics, 1514:24–31, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097525/.

Buhler:1993:FIN

[BLP93] J. P. Buhler, H. W. Lenstra, Jr., and Carl Pomerance. Factoring integers with the number field sieve. Lecture Notes in Mathematics, 1554:50–94, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091539/.

Benedetti:1990:TBN

[BLR90] Riccardo Benedetti, François Loeser, and Jean-Jacques Risler. Two bounds for the number of connected components of a real algebraic set. Lecture Notes in Mathematics, 1420:22–35, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083908/.

Blumlinger:1990:SPP

[Blü90] Martin Blümlinger. Sample path properties of diffusion processes on compact manifolds. Lecture Notes in Mathematics, 1452:6– 19, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096977/.

Bachelis:1990:CTP

[BM90a] Gregory F. Bachelis and Frank J. Massey III. A coin tossing problem of R. L. Rivest. Lecture Notes in Mathematics, 1450: 36–52, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084897/.

Bierstone:1990:LRS

[BM90b] Edward Bierstone and Pierre D. Milman. Local resolution of singularities. Lecture Notes in Mathematics, 1420:42–64, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083910/.

Blanchard:1990:NRS

[BM90c] Ph. Blanchard and Zhiming Ma. New results on the Schrödinger semigroups with potentials given by signed smooth measures.

Lecture Notes in Mathematics, 1444:213–243, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083618/.

Borchers:1990:LTB

[BM90d] Wolfgang Borchers and Tetsuro Miyakawa. On large time behavior of the total kinetic energy for weak solutions of the Navier– Stokes equations in unbounded domains. Lecture Notes in Mathematics, 1431:73–83, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086056/.

Busenberg:1991:DDE

[BM91] Stavros Busenberg and Mario Martelli, editors. Delay Differential Equations and Dynamical Systems: Proceedings of a Conference in honor of Kenneth Cooke held in Claremont, California, Jan. 13–16, 1990, volume 1475 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1475. URL http:/

/link.springer.com/book/10.1007/BFb0083474;   http://
www.springerlink.com/content/978-3-540-47418-0.

Bakry:1992:IFF

[BM92a] Dominique Bakry and Dominique Michel. Sur les inégalités FKG. (French) []. Lecture Notes in Mathematics, 1526:170–188, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084321/.

Borchers:1992:SCE

[BM92b] Wolfgang Borchers and Tetsuro Miyakawa. On some coercive estimates for the Stokes problem in unbounded domains. Lecture Notes in Mathematics, 1530:71–84, 1992. CODEN LN-MAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090334/.

Boutabia:1992:LCE

[BM92c] Hacène Boutabia and Bernard Maisonneuve. Lois conditionnelles des excursions markoviennes. Lecture Notes in Mathematics, 1526: 162–166, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084319/.

Burdzy:1992:HBP

[BM92d] Krzysztof Burdzy and Donald Marshall. Hitting a boundary point with reflected Brownian motion. Lecture Notes in Mathematics, 1526:81–94, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084312/.

Bhattacharjee:1998:AOQ

[BMMN98a] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Applications and open questions. Lecture Notes in Mathematics, 1698:171–180, 1998. CO-DEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092566/.

Bhattacharjee:1998:CT

[BMMN98b] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Classification theorems. Lecture Notes in Mathematics, 1698:131–142, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092563/.

Bhattacharjee:1998:EJG

[BMMN98c] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Examples of Jordan groups. Lecture Notes in Mathematics, 1698:99–113, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092561/.

Bhattacharjee:1998:GAS

[BMMN98d] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Groups acting on sets. Lecture Notes in Mathematics, 1698:9–18, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092552/.

Bhattacharjee:1998:HS

[BMMN98e] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Homogeneous structures. Lecture Notes in Mathematics, 1698:143–158, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092564/.

Bhattacharjee:1998:HC

[BMMN98f] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. The Hrushovski construction. Lecture Notes in Mathematics, 1698:159–170, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092565/.

Bhattacharjee:1998:JG

[BMMN98g] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Jordan groups. Lecture Notes in Mathematics, 1698:87–97, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092560/.

Bhattacharjee:1998:LG

[BMMN98h] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Linear groups. Lecture Notes in Mathematics, 1698:57–66, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092557/.

Bhattacharjee:1998:MAS

[BMMN98i] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. More about symmetric groups. Lecture Notes in Mathematics, 1698:49–56, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092556/.

Bhattacharjee:1998:NIP

[BMMN98j] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Notes on Infinite Permutation Groups, volume 1698 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 180 (est.) pp. LCCN QA3 .L28 no. 1698. URL http:/

/link.springer.com/book/10.1007/BFb0092550;   http://
www.springerlink.com/content/978-3-540-49813-1.

Bhattacharjee:1998:P

[BMMN98k] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Primitivity. Lecture Notes in Mathematics, 1698:31–38, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092554/.

Bhattacharjee:1998:RN

[BMMN98l] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Rational numbers. Lecture Notes in Mathematics, 1698:77–86, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092559/.

Bhattacharjee:1998:RRB

[BMMN98m] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Relations related to betweenness. Lecture Notes in Mathematics, 1698:115–129, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092562/.

Bhattacharjee:1998:SGT

[BMMN98n] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Some group theory. Lecture Notes in Mathematics, 1698:1–8, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092551/.

Bhattacharjee:1998:SO

[BMMN98o] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Suborbits and orbitals. Lecture Notes in Mathematics, 1698:39–48, 1998. CODEN LN-MAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092555/.

Bhattacharjee:1998:T

[BMMN98p] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Transitivity. Lecture Notes in Mathematics, 1698:19–30, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092553/.

Bhattacharjee:1998:WP

[BMMN98q] Meenaxi Bhattacharjee, Dugald Macpherson, Rögnvaldur G. Möller, and Peter M. Neumann. Wreath products. Lecture Notes in Mathematics, 1698:67–76, 1998. CODEN LNMAA2. ISBN 3-540-64965-4 (print), 3-540-49813-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092558/.

Brezzi:1991:SNP

[BMMP91] F. Brezzi, L. D. Marini, P. Markowich, and P. Pietra. On some numerical problems in semiconductor device simulation. Lecture Notes in Mathematics, 1460:31–42, 1991. CODEN LN-MAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091359/.

Bahlali:1998:PUA

[BMO98] Khaled Bahlali, Brahim Mezerdi, and Youssef Ouknine. Pathwise uniqueness and approximation of solutions of stochastic differential equations. Lecture Notes in Mathematics, 1686:166–187, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101757/.

Bocherer:1990:YTL

[Böc90] Siegfried Böcherer. On Yoshida’s theta lift. Lecture Notes in Mathematics, 1447:77–83, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085727/.

Bass:1996:CNS

[BOERT96a] Hyman Bass, Maria Victoria Otero-Espinar, Daniel Rockmore, and Charles Tresser. Closed normal subgroups of Aut(X(q)). Lecture Notes in Mathematics, 1621:135–156, 1996. CODEN LNMAA2. ISBN 3-540-60595-9 (print), 3-540-47812-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096325/.

Bass:1996:CR

[BOERT96b] Hyman Bass, Maria Victoria Otero-Espinar, Daniel Rockmore, and Charles Tresser. Cyclic renormalization. Lecture Notes in Mathematics, 1621:1–51, 1996. CODEN LNMAA2. ISBN 3-540-60595-9 (print), 3-540-47812-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096322/.

Bass:1996:CRA

[BOERT96c] Hyman Bass, Maria Victoria Otero-Espinar, Daniel Rockmore, and Charles Tresser. Cyclic Renormalization and Automorphism Groups of Rooted Trees, volume 1621 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-60595-9 (print), 3-540-47812-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 163 pp. LCCN QA3 .L28 no. 1621.

URL http://link.springer.com/book/10.1007/BFb0096321;
 http://www.springerlink.com/content/978-3-540-47812-
6.

Bass:1996:ICR

[BOERT96d] Hyman Bass, Maria Victoria Otero-Espinar, Daniel Rockmore, and Charles Tresser. Itinerary calculus and renormalization. Lecture Notes in Mathematics, 1621:53–101, 1996. CODEN LN-MAA2. ISBN 3-540-60595-9 (print), 3-540-47812-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096323/.

Bass:1996:STA

[BOERT96e] Hyman Bass, Maria Victoria Otero-Espinar, Daniel Rockmore, and Charles Tresser. Spherically transitive automorphisms of rooted trees. Lecture Notes in Mathematics, 1621:103–133, 1996. CODEN LNMAA2. ISBN 3-540-60595-9 (print), 3-540-47812-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096324/.

Bogley:1990:LCD

[Bog90] William A. Bogley. Local collapses for diagrammatic reducibility. Lecture Notes in Mathematics, 1440:27–38, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084447/.

Boillat:1996:NLH

[Boi96] Guy Boillat. Non linear hyperbolic fields and waves. Lecture Notes in Mathematics, 1640:1–47, 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093705/.

Bolibruch:1992:FSR

[Bol92] A. A. Bolibruch. Fuchsian systems with reducible monodromy and the Riemann–Hilbert problem. Lecture Notes in Mathematics, 1520:139–155, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084719/.

Boman:1991:HST

[Bom91] Jan Boman. Helgason’s support theorem for Radon transforms — a new proof and a generalization. Lecture Notes in Mathematics, 1497:1–5, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084503/.

Bony:1991:AME

[Bon91] Jean-Michel Bony. Analyse microlocale des équations aux dérivées partielles non linéaires. (French) []. Lecture Notes in Mathematics, 1495:1–45, 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085121/.

Booth:1990:EHT

[Boo90] Peter Booth. Equivalent homotopy theories and groups of selfequivalences. Lecture Notes in Mathematics, 1425:1–16, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083826/.

Borel:1990:PCP

[Bor90a] Jean-Pierre Borel. Polynômes à coefficients positifs multiples d’un polynôme donné. (French) []. Lecture Notes in Mathematics, 1415: 97–115, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084881/.

Borisovich:1990:LSM

[Bor90b] A. Yu. Borisovich. Lyapunov–Schmidt method and types of singularities of critical points of key function in the problem of bifurcations of minimal surfaces. Lecture Notes in Mathematics, 1453:201–210, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085956/.

Borisovich:1990:MAT

[Bor90c] Yu. G. Borisovich. A modern approach to the theory of topological characteristics of non-linear operators. II. Lecture Notes in Mathematics, 1453:21–49, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085945/.

Borobia:1992:MPN

[Bor92] Alberto Borobia. Mirror property for nonsingular mixed configurations of one line and k points in R^{3}. Lecture Notes in Mathematics, 1524:140–144, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084614/.

Bornemann:1998:ARQ

[Bor98a] Folkmar Bornemann. Adiabatic results in quantum theory and quantum-classical coupling. Lecture Notes in Mathematics, 1687: 89–114, 1998. CODEN LNMAA2. ISBN 3-540-64447-4 (print), 3-540-69784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092095/.

Bornemann:1998:A

[Bor98b] Folkmar Bornemann. Applications. Lecture Notes in Mathematics, 1687:73–88, 1998. CODEN LNMAA2. ISBN 3-540-64447-4 (print), 3-540-69784-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092094/.

Bornemann:1998:HTS

[Bor98c] Folkmar Bornemann. Homogenization in Time of Singularly Perturbed Mechanical Systems, volume 1687 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CO-DEN LNMAA2. ISBN 3-540-64447-4 (print), 3-540-69784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 156 pp. LCCN QA3 .L28 no. 1687. URL http:/

/link.springer.com/book/10.1007/BFb0092091;   http://
www.springerlink.com/content/978-3-540-69784-8.

Bornemann:1998:HNM

[Bor98d] Folkmar Bornemann. Homogenization of natural mechanical systems with a strong constraining potential. Lecture Notes in Mathematics, 1687:17–72, 1998. CODEN LNMAA2. ISBN 3-540-64447-4 (print), 3-540-69784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092093/.

Bornemann:1998:I

[Bor98e] Folkmar Bornemann. Introduction. Lecture Notes in Mathematics, 1687:1–16, 1998. CODEN LNMAA2. ISBN 3-540-64447-4 (print), 3-540-69784-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092092/.

Bothe:1992:EAS

[Bot92] H. G. Bothe. Expanding attractors with stable foliations of class C^{0}. Lecture Notes in Mathematics, 1514:36–61, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097527/.

Bougerol:1991:FKB

[Bou91a] Philippe Bougerol. Filtre de Kalman Bucy et exposants de Lyapounov. (French) []. Lecture Notes in Mathematics, 1486:112– 122, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086662/.

Bourgain:1991:DPH

[Bou91b] J. Bourgain. On the distribution of polynomials on high dimensional convex sets. Lecture Notes in Mathematics, 1469:127– 137, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089219/.

Bourgain:1991:RMC

[Bou91c] J. Bourgain. Remarks on Montgomery’s conjectures on Dirichlet sums. Lecture Notes in Mathematics, 1469:153–165, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089222/.

Bourgain:1991:RMP

[Bou91d] Jean Bourgain. On the restriction and multiplier problems in R^{3}. Lecture Notes in Mathematics, 1469:179–191, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089225/.

Bourn:1991:NEK

[Bou91e] Dominique Bourn. Normalization equivalence, kernel equivalence and affine categories. Lecture Notes in Mathematics, 1488:43– 62, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084212/.

Bouc:1997:AC

[Bou97a] Serge Bouc. Adjoint constructions. Lecture Notes in Mathematics, 1671:183–222, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095831/.

Bouc:1997:AGF

[Bou97b] Serge Bouc. Adjunction and Green functors. Lecture Notes in Mathematics, 1671:223–274, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095832/.

Bouc:1997:AAG

[Bou97c] Serge Bouc. The algebra associated to a Green functor. Lecture Notes in Mathematics, 1671:81–97, 1997. CODEN LN-MAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095826/.

Bouc:1997:CAG

[Bou97d] Serge Bouc. The category associated to a Green functor. Lecture Notes in Mathematics, 1671:61–80, 1997. CODEN LN-MAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095825/.

Bouc:1997:Cb

[Bou97e] Serge Bouc. Centres. Lecture Notes in Mathematics, 1671:305– 336, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095834/.

Bouc:1997:Ca

[Bou97f] Serge Bouc. Composition. Lecture Notes in Mathematics, 1671: 167–182, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095830/.

Bouc:1997:CGF

[Bou97g] Serge Bouc. Construction of Green functors. Lecture Notes in Mathematics, 1671:123–152, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095828/.

Bouc:1997:GF

[Bou97h] Serge Bouc. Green functors. Lecture Notes in Mathematics, 1671: 41–60, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095824/.

Bouc:1997:GFS

[Bou97i] Serge Bouc. Green Functors and G-sets, volume 1671 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 342 pp. LCCN QA3 .L28 no. 1671. URL http:/

/link.springer.com/book/10.1007/BFb0095821;   http://
www.springerlink.com/content/978-3-540-69596-7.

Bouc:1997:I

[Bou97j] Serge Bouc. Introduction. Lecture Notes in Mathematics, 1671: 1–3, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095822/.

Bouc:1997:MF

[Bou97k] Serge Bouc. Mackey functors. Lecture Notes in Mathematics, 1671:5–39, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095823/.

Bouc:1997:MER

[Bou97l] Serge Bouc. Morita equivalence and relative projectivity. Lecture Notes in Mathematics, 1671:99–121, 1997. CODEN LN-MAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095827/.

Bouc:1997:MT

[Bou97m] Serge Bouc. A Morita theory. Lecture Notes in Mathematics, 1671:153–165, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095829/.

Bouc:1997:SM

[Bou97n] Serge Bouc. The simple modules. Lecture Notes in Mathematics, 1671:275–304, 1997. CODEN LNMAA2. ISBN 3-540-63550-5 (print), 3-540-69596-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095833/.

Bouscaren:1998:IMT

[Bou98a] Elisabeth Bouscaren. Introduction to model theory. Lecture Notes in Mathematics, 1696:1–18, 1998. CODEN LN-MAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-68521-0_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
540-68521-0/1.pdf.

Bouscaren:1998:MTA

[Bou98b] Elisabeth Bouscaren. Model Theory and Algebraic Geometry: An introduction to E. Hrushovski’s proof of the geometric Mordell– Lang conjecture, volume 1696 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 211 pp. LCCN QA3 .L28 no. 1696.

URL http://link.springer.com/book/10.1007/978-3-540-
68521-0;   http://www.springerlink.com/content/978-3-
540-68521-0.

Bouscaren:1998:PML

[Bou98c] Elisabeth Bouscaren. Proof of the Mordell–Lang conjecture for function fields. Lecture Notes in Mathematics, 1696:177– 196, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-68521-0_10/.

Bolton:1991:MI

[BOWV91] J. Bolton, W. M. Oxbury, L. M. Woodward, and L. Vrancken. Minimal immersions of Rp^{2} into Cp^{n}. Lecture Notes in Mathematics, 1481:18–27, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083624/.

Boxler:1991:HCS

[Box91] Petra Boxler. How to construct stochastic center manifolds on the level of vector fields. Lecture Notes in Mathematics, 1486: 141–158, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086664/.

Barlow:1990:PUE

[BP90a] Martin T. Barlow and Edwin A. Perkins. On pathwise uniqueness and expansion of filtrations. Lecture Notes in Mathematics, 1426:194–209, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083766/.

Barlow:1990:CS

[BP90b] Martin T. Barlow and Philip Protter. On convergence of semimartingales. Lecture Notes in Mathematics, 1426:188–193, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083765/.

Bhat:1991:GHO

[BP91] B. V. Rajarama Bhat and K. R. Parthasarathy. Generalized harmonic oscillators in quantum probability. Lecture Notes in Mathematics, 1485:39–51, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100845/.

Benedetti:1997:BSQ

[BP97a] Riccardo Benedetti and Carlo Petronio. Branched spines and quantum invariants. Lecture Notes in Mathematics, 1653:98– 107, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093628/.

Benedetti:1997:BSSa

[BP97b] Riccardo Benedetti and Carlo Petronio. Branched standard spines. Lecture Notes in Mathematics, 1653:23–39, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093623/.

Benedetti:1997:BSSb

[BP97c] Riccardo Benedetti and Carlo Petronio. Branched Standard Spines of 3-manifolds, volume 1653 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 132 pp. LCCN QA3 .L28 no. 1653.

URL http://link.springer.com/book/10.1007/BFb0093620;
 http://www.springerlink.com/content/978-3-540-68345-
2.

Benedetti:1997:CCM

[BP97d] Riccardo Benedetti and Carlo Petronio. Combed closed manifolds. Lecture Notes in Mathematics, 1653:64–72, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093625/.

Benedetti:1997:FSM

[BP97e] Riccardo Benedetti and Carlo Petronio. Framed and spin manifolds. Lecture Notes in Mathematics, 1653:85–97, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093627/.

Benedetti:1997:HCC

[BP97f] Riccardo Benedetti and Carlo Petronio. Homology and cohomology computations. Lecture Notes in Mathematics, 1653:121– 126, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093630/.

Benedetti:1997:MB

[BP97g] Riccardo Benedetti and Carlo Petronio. Manifolds with boundary. Lecture Notes in Mathematics, 1653:40–63, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093624/.

Benedetti:1997:MCC

[BP97h] Riccardo Benedetti and Carlo Petronio. More on combings, and the closed calculus. Lecture Notes in Mathematics, 1653:73– 84, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093626/.

Benedetti:1997:MPS

[BP97i] Riccardo Benedetti and Carlo Petronio. Motivations, plan and statements. Lecture Notes in Mathematics, 1653:1–12, 1997. CO-DEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093621/.

Benedetti:1997:PP

[BP97j] Riccardo Benedetti and Carlo Petronio. Problems and perspectives. Lecture Notes in Mathematics, 1653:108–120, 1997. CO-DEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093629/.

Benedetti:1997:RSS

[BP97k] Riccardo Benedetti and Carlo Petronio. A review on standard spines and o-graphs. Lecture Notes in Mathematics, 1653:13– 22, 1997. CODEN LNMAA2. ISBN 3-540-62627-1 (print), 3-540-68345-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093622/.

Burkholder:1991:EEP

[BPSH91] D. L. (Donald L.) Burkholder, E. (Etienne) Pardoux, Alain-Sol Sznitman, and Paul Louis Hennequin, editors. Ecole d’été de probabilités de Saint-Flour XIX, 1989. (French) [Saint-Flour Summer School on Probability XIX, 1989], volume 1464 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNMAA2. ISBN 0-387-53841-0, 3-540-53841-0 (print), 3-540-46319-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1464. URL http:/

/link.springer.com/book/10.1007/BFb0085166;   http://
www.springerlink.com/content/978-3-540-46319-1.

Burstall:1990:FHS

[BR90a] Francis E. Burstall and John H. Rawnsley. Factorisation of harmonic spheres in Lie groups. Lecture Notes in Mathematics, 1424: 90–105, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095570/.

Burstall:1990:FM

[BR90b] Francis E. Burstall and John H. Rawnsley. Flag manifolds. Lecture Notes in Mathematics, 1424:39–62, 1990. CODEN LN-MAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095566/.

Burstall:1990:HMT

[BR90c] Francis E. Burstall and John H. Rawnsley. Harmonic maps and twistor spaces. Lecture Notes in Mathematics, 1424:15–21, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095564/.

Burstall:1990:HG

[BR90d] Francis E. Burstall and John H. Rawnsley. Homogeneous geometry. Lecture Notes in Mathematics, 1424:6–14, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095563/.

Burstall:1990:I

[BR90e] Francis E. Burstall and John H. Rawnsley. Introduction. Lecture Notes in Mathematics, 1424:1–5, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095562/.

Burstall:1990:SHS

[BR90f] Francis E. Burstall and John H. Rawnsley. Stable harmonic 2spheres. Lecture Notes in Mathematics, 1424:81–89, 1990. CO-DEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095569/.

Burstall:1990:SS

[BR90g] Francis E. Burstall and John H. Rawnsley. Symmetric spaces. Lecture Notes in Mathematics, 1424:22–38, 1990. CODEN LN-MAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095565/.

Burstall:1990:TLR

[BR90h] Francis E. Burstall and John H. Rawnsley. Twistor lifts over Riemannian symmetric spaces. Lecture Notes in Mathematics, 1424: 71–80, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095568/.

Burstall:1990:TSR

[BR90i] Francis E. Burstall and John H. Rawnsley. The twistor space of a Riemannian symmetric space. Lecture Notes in Mathematics, 1424:63–70, 1990. CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095567/.

Burstall:1990:TTR

[BR90j] Francis E. Burstall and John H. Rawnsley. Twistor Theory for Riemannian Symmetric Spaces: With Applications to Harmonic Maps of Riemann Surfaces, volume 1424 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

CODEN LNMAA2. ISBN 3-540-52602-1 (print), 3-540-47052-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 112 pp. LCCN QA3 .L28 no. 1424. URL http:/

/link.springer.com/book/10.1007/BFb0095561;   http://
www.springerlink.com/content/978-3-540-47052-6.

Biais:1997:RSA

[BR97] Bruno Biais and Jean Charles Rochet. Risk sharing, adverse selection and market structure. Lecture Notes in Mathematics, 1656: 1–51, 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091998/.

Braides:1998:AFD

[Bra98a] Andrea Braides. Approximation of Free-Discontinuity Problems, volume 1694 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 149 pp. LCCN QA3 .L28 no. 1694. URL http:/

/link.springer.com/book/10.1007/BFb0097344;   http://
www.springerlink.com/content/978-3-540-68714-6.

Braides:1998:EA

[Bra98b] Andrea Braides. Examples of approximation. Lecture Notes in Mathematics, 1694:39–86, 1998. CODEN LNMAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097348/.

Braides:1998:FBV

[Bra98c] Andrea Braides. Functions of bounded variation. Lecture Notes in Mathematics, 1694:7–26, 1998. CODEN LNMAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097346/.

Braides:1998:GAA

[Bra98d] Andrea Braides. A general approach to approximation. Lecture Notes in Mathematics, 1694:87–102, 1998. CODEN LN-MAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097349/.

Braides:1998:I

[Bra98e] Andrea Braides. Introduction. Lecture Notes in Mathematics, 1694:1–6, 1998. CODEN LNMAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097345/.

Braides:1998:NLA

[Bra98f] Andrea Braides. Non-local approximation. Lecture Notes in Mathematics, 1694:103–130, 1998. CODEN LNMAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097350/.

Braides:1998:SFB

[Bra98g] Andrea Braides. Special functions of bounded variation. Lecture Notes in Mathematics, 1694:27–38, 1998. CODEN LN-MAA2. ISBN 3-540-64771-6 (print), 3-540-68714-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097347/.

Brezis:1993:NEH

[Bré93] Haïm Brézis. New energies for harmonic maps and liquid crystals. Lecture Notes in Mathematics, 1540:11–24, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085471/.

Breitung:1994:ANI

[Bre94a] Karl Wilhelm Breitung. Approximations for normal integrals. Lecture Notes in Mathematics, 1592:85–105, 1994. CODEN LN-MAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073544/.

Breitung:1994:API

[Bre94b] Karl Wilhelm Breitung. Arbitrary probability integrals. Lecture Notes in Mathematics, 1592:106–120, 1994. CODEN LN-MAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073545/.

Breitung:1994:AA

[Bre94c] Karl Wilhelm Breitung. Asymptotic analysis. Lecture Notes in Mathematics, 1592:34–44, 1994. CODEN LNMAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073541/.

Breitung:1994:AAP

[Bre94d] Karl Wilhelm Breitung. Asymptotic Approximations for Probability Integrals, volume 1592 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 146 pp. LCCN QA3 .L28 no. 1592.

URL http://link.springer.com/book/10.1007/BFb0073538;
 http://www.springerlink.com/content/978-3-540-49033-
3.

Breitung:1994:CRS

[Bre94e] Karl Wilhelm Breitung. Crossing rates of stochastic processes. Lecture Notes in Mathematics, 1592:121–134, 1994. CODEN LNMAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073546/.

Breitung:1994:I

[Bre94f] Karl Wilhelm Breitung. Introduction. Lecture Notes in Mathematics, 1592:1–8, 1994. CODEN LNMAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073539/.

Breitung:1994:MP

[Bre94g] Karl Wilhelm Breitung. Mathematical preliminaries. Lecture Notes in Mathematics, 1592:9–33, 1994. CODEN LN-MAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073540/.

Breitung:1994:MLT

[Bre94h] Karl Wilhelm Breitung. Multivariate Laplace type integrals. Lecture Notes in Mathematics, 1592:51–84, 1994. CODEN LN-MAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073543/.

Breitung:1994:UI

[Bre94i] Karl Wilhelm Breitung. Univariate integrals. Lecture Notes in Mathematics, 1592:45–50, 1994. CODEN LNMAA2. ISBN 3-540-58617-2 (print), 3-540-49033-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073542/.

Brinzanescu:1996:CVB

[Brî96a] Vasile Brînzănescu. Classification of vector bundles. Lecture Notes in Mathematics, 1624:119–155, 1996. CODEN LNMAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093701/.

Brinzanescu:1996:EHV

[Brî96b] Vasile Brînzănescu. Existence of holomorphic vector bundles. Lecture Notes in Mathematics, 1624:85–117, 1996. CODEN LN-MAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093700/.

Brinzanescu:1996:FCC

[Brî96c] Vasile Brînzănescu. Facts on compact complex surfaces. Lecture Notes in Mathematics, 1624:29–52, 1996. CODEN LN-MAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093698/.

Brinzanescu:1996:HVB

[Brî96d] Vasile Brînzănescu. Holomorphic Vector Bundles over Compact Complex Surfaces, volume 1624 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 170 pp. LCCN QA3 .L28 no. 1624.

URL http://link.springer.com/book/10.1007/BFb0093696;
 http://www.springerlink.com/content/978-3-540-49845-
2.

Brinzanescu:1996:LBS

[Brî96e] Vasile Brînzănescu. Line bundles over surfaces. Lecture Notes in Mathematics, 1624:53–83, 1996. CODEN LNMAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093699/.

Brinzanescu:1996:VBC

[Brî96f] Vasile Brînzănescu. Vector bundles over complex manifolds. Lecture Notes in Mathematics, 1624:1–27, 1996. CODEN LN-MAA2. ISBN 3-540-61018-9 (print), 3-540-49845-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093697/.

Brocker:1990:SIN

[Brö90a] Ludwig Bröcker. On the stability index of Noetherian rings. Lecture Notes in Mathematics, 1420:72–80, 1990. CODEN LN-MAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083912/.

Brown:1990:SPN

[Bro90b] Ronald Brown. Some problems in non-abelian homotopical and homological algebra. Lecture Notes in Mathematics, 1418:105– 129, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083697/.

Brocker:1992:FSS

[Brö92] Ludwig Bröcker. Families of semialgebraic sets and limits. Lecture Notes in Mathematics, 1524:145–162, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084615/.

Brockhaus:1993:SSB

[Bro93] Oliver Brockhaus. Sufficient statistics for the Brownian sheet. Lecture Notes in Mathematics, 1557:44–52, 1993. CODEN LN-MAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087963/.

Brokate:1994:HO

[Bro94] Martin Brokate. Hysteresis operators. Lecture Notes in Mathematics, 1584:1–38, 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073394/.

Bobadilla:1994:CCI

[BRS94] Gladys Bobadilla, Rolando Rebolledo, and Eugenio Saavedra. Corrections à: “Sur la convergence d’intégrales anticipatives”. (French) []. Lecture Notes in Mathematics, 1583:113–115, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073839/.

Bruns:1990:SLM

[Bru90] Winfried Bruns. Straightening laws on modules and their symmetric algebras. Lecture Notes in Mathematics, 1430:1–20, 1990. CODEN LNMAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085533/.

Brumfiel:1992:HFP

[Bru92a] G. W. Brumfiel. A Hopf fixed point theorem for semi-algebraic maps. Lecture Notes in Mathematics, 1524:163–169, 1992. CO-DEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084616/.

Brumfiel:1992:ROS

[Bru92b] G. W. Brumfiel. On regular open semi-algebraic sets. Lecture Notes in Mathematics, 1524:170–173, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084617/.

Beltrametti:1990:SPS

[BS90a] Mauro Beltrametti and Andrew J. Sommese. On K-spannedness for projective surfaces. Lecture Notes in Mathematics, 1417:24– 51, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083331/.

Borisovich:1990:RTC

[BS90b] Yu. G. Borisovich and M. I. Shpilberg. Relative topological characteristics of mappings. Lecture Notes in Mathematics, 1453:211– 225, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085957/.

Bruns:1990:CAP

[BS90c] Winfried Bruns and Aron Simis, editors. Commutative Algebra: Proceedings of a Workshop held in Salvador, Brazil, Aug. 8–17, 1988, volume 1430 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1430. URL http:/

/link.springer.com/book/10.1007/BFb0085532;   http://
www.springerlink.com/content/978-3-540-47136-3.

Bohnhorst:1992:SCV

[BS92] Guntram Bohnhorst and Heinz Spindler. The stability of certain vector bundles on P^{n}. Lecture Notes in Mathematics, 1507: 39–50, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094509/.

Blum:1997:BM

[BS97a] A. Blum and U. Stuhler. Back matter. Lecture Notes in Mathematics, 1649:??, 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-49701-1/1.pdf.

Blum:1997:DME

[BS97b] A. Blum and U. Stuhler. Drinfeld modules and elliptic sheaves. Lecture Notes in Mathematics, 1649:110–188, 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094426/.

Brannath:1999:BT

[BS99] W. Brannath and W. Schachermayer. A bipolar theorem for L_{+}^{0} (Ω, F, P). Lecture Notes in Mathematics, 1709:349–354, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096525/.

Beghdadi-Sakrani:1999:CPE

[BSÉ99] S. Beghdadi-Sakrani and M. Émery. On certain probabilities equivalent to coin-tossing, d’après Schachermayer. Lecture Notes in Mathematics, 1709:240–256, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096514/.

Beltrametti:1992:RVM

[BSW92] Mauro C. Beltrametti, Andrew J. Sommese, and Jarosław A. Wiśniewski. Results on varieties with many lines and their applications to adjunction theory. Lecture Notes in Mathematics, 1507: 16–38, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094508/.

Buslaev:1993:STN

[BT93] A. P. Buslaev and V. M. Tikhomirov. Spectral theory of nonlinear equations and n-widths of Sobolev spaces. Lecture Notes in Mathematics, 1550:19–30, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0117471/.

Berndt:1995:DRS

[BTV95a] Jürgen Berndt, Franco Tricerri, and Lieven Vanhecke. Damek– Ricci spaces. Lecture Notes in Mathematics, 1598:78–114, 1995. CODEN LNMAA2. ISBN 3-540-59001-3 (print), 3-540-49171-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0076906/.

Berndt:1995:GHGa

[BTV95b] Jürgen Berndt, Franco Tricerri, and Lieven Vanhecke. Generalized Heisenberg groups. Lecture Notes in Mathematics, 1598:

21–77, 1995. CODEN LNMAA2. ISBN 3-540-59001-3 (print), 3-540-49171-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076905/.

Berndt:1995:GHGb

[BTV95c] Jürgen Berndt, Franco Tricerri, and Lieven Vanhecke. Generalized Heisenberg Groups and Damek–Ricci Harmonic Spaces, volume 1598 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-59001-3 (print), 3-540-49171-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 125 pp. LCCN QA3 .L28 no. 1598. URL http:/

/link.springer.com/book/10.1007/BFb0076902;   http://
www.springerlink.com/content/978-3-540-49171-2.

Berndt:1995:I

[BTV95d] Jürgen Berndt, Franco Tricerri, and Lieven Vanhecke. Introduction. Lecture Notes in Mathematics, 1598:1–3, 1995. CODEN LNMAA2. ISBN 3-540-59001-3 (print), 3-540-49171-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076903/.

Berndt:1995:SLR

[BTV95e] Jürgen Berndt, Franco Tricerri, and Lieven Vanhecke. Symmetriclike Riemannian manifolds. Lecture Notes in Mathematics, 1598: 4–20, 1995. CODEN LNMAA2. ISBN 3-540-59001-3 (print), 3-540-49171-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076904/.

Buium:1992:AGS

[Bui92a] Alexandru Buium. Affine D-group schemes. Lecture Notes in Mathematics, 1506:30–59, 1992. CODEN LNMAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087238/.

Buium:1992:ADA

[Bui92b] Alexandru Buium. Applications to differential algebraic groups. Lecture Notes in Mathematics, 1506:99–136, 1992. CODEN LN-MAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087241/.

Buium:1992:CAG

[Bui92c] Alexandru Buium. Commutative algebraic D-groups. Lecture Notes in Mathematics, 1506:60–86, 1992. CODEN LN-MAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087239/.

Buium:1992:DAG

[Bui92d] Alexandru Buium. Differential Algebraic Groups of Finite Dimension, volume 1506 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 145 pp. LCCN QA3 .L28 no. 1506. URL http:/

/link.springer.com/book/10.1007/BFb0087235;   http://
www.springerlink.com/content/978-3-540-46764-9.

Buium:1992:FP

[Bui92e] Alexandru Buium. First properties. Lecture Notes in Mathematics, 1506:7–29, 1992. CODEN LNMAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087237/.

Buium:1992:GAG

[Bui92f] Alexandru Buium. General algebraic D-groups. Lecture Notes in Mathematics, 1506:87–98, 1992. CODEN LNMAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087240/.

Buium:1992:TC

[Bui92g] Alexandru Buium. Terminology and conventions. Lecture Notes in Mathematics, 1506:1–6, 1992. CODEN LNMAA2. ISBN 3-540-55181-6 (print), 3-540-46764-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087236/.

Bullen:1990:SAT

[Bul90] P. S. Bullen. Some applications of a theorem of Marcinkiewicz. Lecture Notes in Mathematics, 1419:10–18, 1990. CODEN LN-MAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083095/.

Bundschuh:1990:IAP

[Bun90] P. Bundschuh. Indépendance algébrique par des méthodes d’approximations. (French) []. Lecture Notes in Mathematics, 1415:116–122, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084882/.

Bunimovich:1992:AFM

[Bun92] L. A. Bunimovich. On absolutely focusing mirrors. Lecture Notes in Mathematics, 1514:62–82, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097528/.

Burns:1974:RIS

[Bur74] R. G. Burns. On the rank of the intersection of subgroups of a Fuchsian group. Lecture Notes in Mathematics, 372: 165–187, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21571-5_14/;
http://link.springer.com/chapter/10.1007/BFb0065168/.

See corrigenda [Bur90].

Burns:1990:CPR

[Bur90] R. G. Burns. Corrigenda to the paper ‘On the rank of the intersection of subgroups of a Fuchsian group’. Lecture Notes in Mathematics, 1456:198, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100742. See [Bur74].

Burkholder:1991:EMT

[Bur91] Donald L. Burkholder. Explorations in martingale theory and its applications. Lecture Notes in Mathematics, 1464:1–66, 1991.

CODEN LNMAA2. ISBN 3-540-53841-0 (print), 3-540-46319-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085167/.

Burdzy:1993:ELE

[Bur93] Krzysztof Burdzy. Excursion laws and exceptional points on Brownian paths. Lecture Notes in Mathematics, 1557:177–181, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087974/.

Busenberg:1992:CSI

[Bus92] Stavros Busenberg. Case studies in industrial mathematics. Lecture Notes in Mathematics, 1521:111–153, 1992. CODEN LN-MAA2. ISBN 3-540-55595-1 (print), 3-540-47247-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098367/.

Babin:1990:SDP

[BV90] A. V. Babin and M. I. Vishik. Semigroups dependent on a parameter, their attractors and asymptotic behaviour. Lecture Notes in Mathematics, 1453:1–19, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085944/.

Borceux:1991:GSR

[BV91] Francis Borceux and Gilberte Van den Bossche. A generic sheaf representation for rings. Lecture Notes in Mathematics, 1488: 30–42, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084211/.

Berline:1993:ECC

[BV93] Nicole Berline and Michèle Vergne. The equivariant Chern character and index of G-invariant operators. Lectures at CIME, Venise 1992. Lecture Notes in Mathematics, 1565:157–200, 1993. CO-DEN LNMAA2. ISBN 3-540-57498-0 (print), 3-540-48195-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073468/.

Brennan:1994:AC

[BVH94a] J. Brennan, A. Volberg, and V. P. Havin. Approximation and capacities. Lecture Notes in Mathematics, 1574:73–176, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101062/.

Brennan:1994:UMN

[BVH94b] J. Brennan, A. Volberg, and V. P. Havin. Uniqueness, moments, normality. Lecture Notes in Mathematics, 1574:207–258, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101064/.

Bruce:1991:FMF

[BW91] J. W. Bruce and T. C. Wilkinson. Folding maps and focal sets. Lecture Notes in Mathematics, 1462:63–72, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086374/.

Bertoin:1994:AWP

[BW94a] Jean Bertoin and Wendelin Werner. Asymptotic windings of planar Brownian motion revisited via the Ornstein–Uhlenbeck process. Lecture Notes in Mathematics, 1583:138–152, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073842/.

Bertoin:1994:CAD

[BW94b] Jean Bertoin and Wendelin Werner. Comportement asymptotique du nombre de tours effectués par la trajectoire brownienne plane. (French) []. Lecture Notes in Mathematics, 1583:164–171, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073844/.

Broto:1992:SA

[BZ92] Carlos Broto and Saïd Zarati. On sub-a∗_{p}-algebras of H ∗ V. Lecture Notes in Mathematics, 1509:35–49, 1992. CODEN LN-MAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087500/.

Chossat:1991:SSH

[CA91] P. Chossat and D. Armbruster. Structurally stable heteroclinic cycles in a system with O(3) symmetry. Lecture Notes in Mathematics, 1463:38–62, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085425/.

Cadre:1997:PSD

[Cad97] B. Cadre. Une preuve standard du principe d’invariance de Stoll. (French) []. Lecture Notes in Mathematics, 1655:85–102, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0119296/.

Caflisch:1991:SFV

[Caf91] Russel E. Caflisch. Singularity formation for vortex sheets and hyperbolic equations. Lecture Notes in Mathematics, 1460:57– 69, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091361/.

Calabi:1990:ADG

[Cal90] Eugenio Calabi. Affine differential geometry and holomorphic curves. Lecture Notes in Mathematics, 1422:15–21, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089401/.

Capitaine:1998:ISL

[Cap98] Mireille Capitaine. Sur une inégalité de Sobolev logarithmique pour une diffusion unidimensionnelle. (French) []. Lecture Notes in Mathematics, 1686:6–13, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101746/.

Capasso:1999:MIB

[Cap99] Vincenzo Capasso, editor. Mathematics Inspired by Biology: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Martina Franca, Italy, June 13–20, 1997, volume 1714 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1714. URL http:/

/link.springer.com/book/10.1007/BFb0092373;   http://
www.springerlink.com/content/978-3-540-48170-6.

Carleson:1991:SMS

[Car91a] Lennart Carleson. Stochastic models of some dynamical systems. Lecture Notes in Mathematics, 1469:1–12, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089212/.

Carvalho:1991:SDE

[Car91b] L. A. V. Carvalho. On the stability of discrete equations and ordinary differential equations. Lecture Notes in Mathematics, 1475: 88–97, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083482/.

Castellanos:1991:DGS

[Cas91] Julio Castellanos. The dual graph for space curves. Lecture Notes in Mathematics, 1462:73–80, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086375/.

Castilla:1992:SPM

[Cas92] Ana Castilla. Sums of 2n-th powers meromorphic functions with compact zero set. Lecture Notes in Mathematics, 1524:174– 177, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084618/.

Catanese:1990:FTR

[Cat90] Fabrizio Catanese. Footnotes to a theorem of I. Reider. Lecture Notes in Mathematics, 1417:67–74, 1990. CODEN LN-MAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083333/.

Catoni:1999:SAA

[Cat99] Olivier Catoni. Simulated annealing algorithms and Markov chains with rare transitions. Lecture Notes in Mathematics, 1709: 69–119, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096510/.

Casacuberta:1992:MRT

[CC92] Carles Casacuberta and Manuel Castellet, editors. Mathematical Research Today and Tomorrow: Viewpoints of Seven Fields Medalists Lectures given at the Institut d’Estudis Catalans, Barcelona, Spain, June 1991, volume 1525 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1525.

URL http://link.springer.com/book/10.1007/BFb0089200;
 http://www.springerlink.com/content/978-3-540-47341-
1.

Catanese:1996:TMA

[CC96] Fabrizio Catanese and Ciro Ciliberto, editors. Transcendental Methods in Algebraic Geometry: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) Held in Cetraro, Italy, July 4–12, 1994, volume 1646 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62038-9 (print), 3-540-49632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1646.

URL http://link.springer.com/book/10.1007/BFb0094301;
 http://www.springerlink.com/content/978-3-540-49632-
8.

Canalis-Durand:1991:FEV

[CD91] Mireille Canalis-Durand. Formal expansion of van der Pol equation canard solutions are Gevrey. Lecture Notes in Mathematics, 1493:29–39, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085022/.

Candelpergher:1990:RBD

[CDD90] Bernard Candelpergher, Francine Diener, and Marc Diener. Retard à la bifurcation: du local au global. (French) []. Lecture Notes in Mathematics, 1455:1–19, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085388/.

Coornaert:1990:APA

[CDP90a] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Approximation par des arbres. (French) []. Lecture Notes in Mathematics, 1441:90–96, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084921/.

Coornaert:1990:AGH

[CDP90b] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Automates et groupes hyperboliques. (French) []. Lecture Notes in Mathematics, 1441:140–158, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084925/.

Coornaert:1990:BEH

[CDP90c] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Bord d’un espace hyperbolique. (French) []. Lecture Notes in Mathematics, 1441:16–23, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084915/.

Coornaert:1990:CIF

[CDP90d] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Classification des isométries. (French) []. Lecture Notes in Mathematics, 1441:97–105, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084922/.

Coornaert:1990:EMH

[CDP90e] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Espaces métriques hyperboliques. (French) []. Lecture Notes in Mathematics, 1441:1–15, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084914/.

Coornaert:1990:GTG

[CDP90f] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Géométrie et théorie des groupes: Les groupes hyperboliques de Gromov, volume 1441 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 165 pp. LCCN QA3 .L28 no. 1441. URL http:

//link.springer.com/book/10.1007/BFb0084913;   http:/
/www.springerlink.com/content/978-3-540-46294-1.

Coornaert:1990:GHF

[CDP90g] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Groupes hyperboliques. (French) []. Lecture Notes in Mathematics, 1441:43–56, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084917/.

Coornaert:1990:IIE

[CDP90h] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Inégalités isopérimétriques et espaces hyperboliques. (French) []. Lecture Notes in Mathematics, 1441:65–80, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084919/.

Coornaert:1990:IIA

[CDP90i] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Inégalités isopérimétriques: une application. (French) []. Lecture Notes in Mathematics, 1441:81–89, 1990. CODEN LN-MAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084920/.

Coornaert:1990:PF

[CDP90j] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Le polyèdre P_{d}(X). (French) []. Lecture Notes in Mathematics, 1441:57–64, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084918/.

Coornaert:1990:PQC

[CDP90k] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Parties quasi-convexes d’un espace hyperbolique. (French) []. Lecture Notes in Mathematics, 1441:106–123, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084923/.

Coornaert:1990:QGQ

[CDP90l] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Quasi-géodésiques et quasi-isométries dans les espaces hyperboliques. (French) []. Lecture Notes in Mathematics, 1441: 24–42, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084916/.

Coornaert:1990:SMB

[CDP90m] Michel Coornaert, Thomas Delzant, and Athanase Papadopoulos. Structure métrique sur le bord d’un espace hyperbolique. (French) []. Lecture Notes in Mathematics, 1441:124–139, 1990. CODEN LNMAA2. ISBN 3-540-52977-2 (print), 3-540-46294-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084924/.

Cheng:1991:HAP

[CDZ91] Min-Teh Cheng, Dong-Gao Deng, and Xing-Wei Zhou, editors. Harmonic Analysis: Proceedings of the special program at the Nankai Institute of Mathematics Tianjin, PR China, March–July, 1988, volume 1494 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1494. URL http:/

/link.springer.com/book/10.1007/BFb0087751;   http://
www.springerlink.com/content/978-3-540-46474-7.

Cellina:1990:MNA

[Cel90] Arrigo Cellina, editor. Methods of Nonconvex Analysis: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Varenna, Italy, June 15–23, 1989, volume 1446 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1446. URL http:/

/link.springer.com/book/10.1007/BFb0084928;   http://
www.springerlink.com/content/978-3-540-46715-1.

Censor:1991:VBA

[Cen91] Yair Censor. On variable block algebraic reconstruction techniques. Lecture Notes in Mathematics, 1497:133–140, 1991. CO-DEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084514/.

Cepa:1995:EDS

[Cép95] Emmanuel Cépa. Équations différentielles stochastiques multivoques. (French) []. Lecture Notes in Mathematics, 1613:86– 107, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094202/.

Cellier:1990:LSE

[CF90] Dominique Cellier and Dominique Fourdrinier. Sur les lois à symétrie elliptique. (French) []. Lecture Notes in Mathematics, 1426:300–328, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083772/.

Capasso:1992:MMI

[CF92] Vincenzo Capasso and Antonio Fasano, editors. Mathematical Modelling of Industrial Processes: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Bari, Italy, Sept. 24–29, 1990, volume 1521 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2.

ISBN 3-540-55595-1 (print), 3-540-47247-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1521.

URL http://link.springer.com/book/10.1007/BFb0098364;
 http://www.springerlink.com/content/978-3-540-47247-
6.

Chebotarev:1995:QEA

[CF95] A. M. Chebotarev and F. Fagnola. On quantum extensions of the Azéma martingale semigroup. Lecture Notes in Mathematics, 1613:1–16, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094194/.

Connes:1992:RTD

[CFJ^{+}92] Alain Connes, Gerd Faltings, Vaughan Jones, Stephen Smale, René Thom, and Jorge Wagensberg. Round-table discussion. Lecture Notes in Mathematics, 1525:87–108, 1992. CODEN LN-MAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089208/.

Cegarra:1991:LES

[CG91] Antonio M. Cegarra and Antonio R. Garzón. A long exact sequence in non-abelian cohomology. Lecture Notes in Mathematics, 1488:79–94, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084214/.

Castillo:1997:AP

[CG97a] Jesús M. F. Castillo and Manuel González. Approximation properties. Lecture Notes in Mathematics, 1667:221–231, 1997. CO-DEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112518/.

Castillo:1997:CBS

[CG97b] Jesús M. F. Castillo and Manuel González. Classical Banach spaces. Lecture Notes in Mathematics, 1667:81–112, 1997. CO-DEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112514/.

Castillo:1997:GP

[CG97c] Jesús M. F. Castillo and Manuel González. Geometrical properties. Lecture Notes in Mathematics, 1667:156–195, 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112516/.

Castillo:1997:HP

[CG97d] Jesús M. F. Castillo and Manuel González. Homological properties. Lecture Notes in Mathematics, 1667:196–220, 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112517/.

Castillo:1997:MOI

[CG97e] Jesús M. F. Castillo and Manuel González. Methods to obtain 3 SP ideals. Lecture Notes in Mathematics, 1667:45–80, 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112513/.

Castillo:1997:TSC

[CG97f] Jesús M. F. Castillo and Manuel González. Three-space constructions. Lecture Notes in Mathematics, 1667:1–44, 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112512/.

Castillo:1997:TSP

[CG97g] Jesús M. F. Castillo and Manuel González. Three-space Problems in Banach Space Theory, volume 1667 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 267 pp. LCCN QA3 .L28 no. 1667.

URL http://link.springer.com/book/10.1007/BFb0112511;
 http://www.springerlink.com/content/978-3-540-69519-
6.

Castillo:1997:TPB

[CG97h] Jesús M. F. Castillo and Manuel González. Topological properties of Banach spaces. Lecture Notes in Mathematics, 1667:

113–155, 1997. CODEN LNMAA2. ISBN 3-540-63344-8 (print), 3-540-69519-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112515/.

Crawford:1991:BCS

[CGG^{+}91] J. D. Crawford, M. Golubitsky, M. G. M. Gomes, E. Knobloch, and I. N. Stewart. Boundary conditions as symmetry constraints. Lecture Notes in Mathematics, 1463:63–79, 1991. CODEN LN-MAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085426/.

Celeghini:1992:CQG

[CGST92] E. Celeghini, R. Giachetti, E. Sorace, and M. Tarlini. Contractions of quantum groups. Lecture Notes in Mathematics, 1510: 221–244, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101192/.

Chaatit:1991:TTU

[Cha91a] Fouad Chaatit. Twisted types and uniform stability. Lecture Notes in Mathematics, 1470:183–199, 1991. CODEN LN-MAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090222/.

Chabrowski:1991:ESD

[Cha91b] Jan Chabrowski. C_{n−1}-estimate of the solution of the Dirichlet problem with L^{2}-boundary data. Lecture Notes in Mathematics, 1482:142–167, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095761/.

Chabrowski:1991:AM

[Cha91c] Jan Chabrowski. Applications of the L^{2}-method. Lecture Notes in Mathematics, 1482:104–116, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095758/.

Chabrowski:1991:DPBa

[Cha91d] Jan Chabrowski. The Dirichlet problem in a bounded domain. Lecture Notes in Mathematics, 1482:46–66, 1991. CODEN LN-MAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095754/.

Chabrowski:1991:DPH

[Cha91e] Jan Chabrowski. The Dirichlet problem in a half-space. Lecture Notes in Mathematics, 1482:20–45, 1991. CODEN LN-MAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095753/.

Chabrowski:1991:DPBb

[Cha91f] Jan Chabrowski. The Dirichlet Problem with L^{2}-Boundary Data for Elliptic Linear Equations, volume 1482 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 173 pp. LCCN QA3 .L28 no. 1482. URL http:/

/link.springer.com/book/10.1007/BFb0095750;   http://
www.springerlink.com/content/978-3-540-38400-7.

Chabrowski:1991:DB

[Cha91g] Jan Chabrowski. Domains with C^{1,α}-boundary. Lecture Notes in Mathematics, 1482:117–130, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095759/.

Chabrowski:1991:ED

[Cha91h] Jan Chabrowski. Estimates of derivatives. Lecture Notes in Mathematics, 1482:67–77, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095755/.

Chabrowski:1991:ESB

[Cha91i] Jan Chabrowski. Exceptional sets on the boundary. Lecture Notes in Mathematics, 1482:90–103, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095757/.

Chabrowski:1991:HM

[Cha91j] Jan Chabrowski. Harmonic measure. Lecture Notes in Mathematics, 1482:78–89, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095756/.

Chabrowski:1991:I

[Cha91k] Jan Chabrowski. Introduction. Lecture Notes in Mathematics, 1482:1–4, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095751/.

Chabrowski:1991:S

[Cha91l] Jan Chabrowski. The space C_{n−1}(Q̄). Lecture Notes in Mathematics, 1482:131–141, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095760/.

Chabrowski:1991:WSS

[Cha91m] Jan Chabrowski. Weighted Sobolev space W̃^{1,2}. Lecture Notes in Mathematics, 1482:7–19, 1991. CODEN LNMAA2. ISBN 3-540-54486-0 (print), 3-540-38400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095752/.

Chang:1991:NLN

[Cha91n] Der-Chen Chang. Nankai lecture in ¯∂-Neumann problem. Lecture Notes in Mathematics, 1494:1–22, 1991. CODEN LN-MAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087752/.

Chang:1992:SNS

[Cha92] Huakang Chang. The steady Navier–Stokes problem for low Reynolds number viscous jets into a half space. Lecture Notes in Mathematics, 1530:85–96, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090335/.

Chatterji:1996:RIR

[Cha96] Professeur S. D. Chatterji. Remarques sur l’intégrale de Riemann généralisée. (French) []. Lecture Notes in Mathematics, 1626:1– 11, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094637/.

Cherry:1990:ENT

[Che90a] William Cherry. Equidimensional Nevanlinna theory on coverings of C^{n}. Lecture Notes in Mathematics, 1433:143–168, 1990. CODEN LNMAA2. ISBN 3-540-52785-0 (print), 3-540-47146-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093850/.

Cherry:1990:NTM

[Che90b] William Cherry. Nevanlinna theory for meromorphic functions on coverings of C. Lecture Notes in Mathematics, 1433:113– 142, 1990. CODEN LNMAA2. ISBN 3-540-52785-0 (print), 3-540-47146-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093849/.

Cheeger:1991:CPD

[Che91] Jeff Cheeger. Critical points of distance functions and applications to geometry. Lecture Notes in Mathematics, 1504:1–38, 1991. CODEN LNMAA2. ISBN 3-540-55017-8 (print), 3-540-46651-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094288/.

Chiarellotto:1990:DRA

[Chi90a] Bruno Chiarellotto. Duality in rigid analysis. Lecture Notes in Mathematics, 1454:142–172, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091137/.

Chicone:1990:BLC

[Chi90b] Carmen Chicone. On bifurcation of limit cycles from centers. Lecture Notes in Mathematics, 1455:20–43, 1990. CODEN LN-MAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085389/.

Chillingworth:1991:BM

[Chi91] David Chillingworth. Bifurcation from a manifold. Lecture Notes in Mathematics, 1463:22–37, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085424/.

Chiu:1995:EL

[Chi95] Yukuang Chiu. From an example of Lévy’s. Lecture Notes in Mathematics, 1613:162–165, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094208/.

Chiu:1997:MSP

[Chi97] Yukuang Chiu. The multiplicity of stochastic processes. Lecture Notes in Mathematics, 1655:207–215, 1997. CODEN LN-MAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119305/.

Christophersen:1991:CDV

[Chr91] Jan Arthur Christophersen. On the components and discriminant of the versal base space of cyclic quotient singularities. Lecture Notes in Mathematics, 1462:81–92, 1991. CODEN LN-MAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086376/.

Ciliberto:1992:EWG

[CHT92] Ciro Ciliberto, Joe Harris, and Montserrat Teixidor i Bigas. On the endomorphisms of Jac(W_{d}^{1}(C)) when p = 1 and C has general moduli. Lecture Notes in Mathematics, 1515:41–67, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098337/.

Chui:1993:WA

[Chu93] Charles K. Chui. On wavelet analysis. Lecture Notes in Mathematics, 1550:31–42, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0117472/.

Coquet:1990:CSA

[CJ90] François Coquet and Jean Jacod. Convergence des surmartingales — application aux vraisemblances partielles. (French) [Convergence of supermartingales — application to partial likelihoods]. Lecture Notes in Mathematics, 1426:282–299, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083771/.

Cecchini:1990:STP

[CK90] Carlo Cecchini and Burkhard Kümmerer. Stochastic transitions on preduals of von Neumann algebras. Lecture Notes in Mathematics, 1442:126–130, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085505/.

Connolly:1991:ELR

[CK91] Frank Connolly and Tadeusz Koźniewski. Examples of lack of rigidity in crystallographic groups. Lecture Notes in Mathematics, 1474:139–145, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084742/.

Cattiaux:1996:MKI

[CL96] P. Cattiaux and C. Léonard. Minimization of the Kullback information for some Markov processes. Lecture Notes in Mathematics, 1626:288–311, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094655/.

Croisille:1999:DIE

[CL99a] Jean-Pierre Croisille and Gilles Lebeau. Diffraction by an Immersed Elastic Wedge, volume 1723 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 134 pp. LCCN QA3 .L28 no. 1723.

URL http://link.springer.com/book/10.1007/BFb0092515;
 http://www.springerlink.com/content/978-3-540-46698-
7.

Croisille:1999:I

[CL99b] Jean-Pierre Croisille and Gilles Lebeau. Introduction. Lecture Notes in Mathematics, 1723:1–2, 1999. CODEN LNMAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092516.

Croisille:1999:NRa

[CL99c] Jean-Pierre Croisille and Gilles Lebeau. Notation and results. Lecture Notes in Mathematics, 1723:3–26, 1999. CODEN LN-MAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092517/.

Croisille:1999:NA

[CL99d] Jean-Pierre Croisille and Gilles Lebeau. Numerical algorithm. Lecture Notes in Mathematics, 1723:79–95, 1999. CODEN LN-MAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092520/.

Croisille:1999:NRb

[CL99e] Jean-Pierre Croisille and Gilles Lebeau. Numerical results. Lecture Notes in Mathematics, 1723:97–125, 1999. CODEN LN-MAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092521/.

Croisille:1999:PR

[CL99f] Jean-Pierre Croisille and Gilles Lebeau. Proofs of the results. Lecture Notes in Mathematics, 1723:57–78, 1999. CODEN LN-MAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092519/.

Croisille:1999:SF

[CL99g] Jean-Pierre Croisille and Gilles Lebeau. The spectral function. Lecture Notes in Mathematics, 1723:27–56, 1999. CODEN LN-MAA2. ISBN 3-540-66810-1 (print), 3-540-46698-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092518/.

Clemens:1990:OMM

[Cle90] Herbert Clemens. An obstruction to moving multiples of subvarieties. Lecture Notes in Mathematics, 1417:75–90, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083334/.

Clote:1990:MFO

[Clo90] P. Clote. The metamathematics of Fraïssé’s order type conjecture. Lecture Notes in Mathematics, 1432:41–56, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086113/.

Clunie:1992:CPR

[Clu92] J. G. Clunie. Convergence of polynomials with restricted zeros. Lecture Notes in Mathematics, 1511:100–105, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093802/.

Chong:1990:PSP

[CM90] C. T. Chong and K. J. Mourad. Positive solutions to Post’s problem. Lecture Notes in Mathematics, 1432:33–40, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086112/.

Cembranos:1997:Sa

[CM97a] Pilar Cembranos and José Mendoza. C(K, X) spaces. Lecture Notes in Mathematics, 1676:65–74, 1997. CODEN LN-MAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096769/.

Cembranos:1997:Sb

[CM97b] Pilar Cembranos and José Mendoza. L_{p}(μ, X) spaces. Lecture Notes in Mathematics, 1676:75–82, 1997. CODEN LN-MAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096770/.

Cembranos:1997:BSV

[CM97c] Pilar Cembranos and José Mendoza. Banach Spaces of Vector-Valued Functions, volume 1676 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 118 pp. LCCN QA3 .L28 no. 1676.

URL http://link.springer.com/book/10.1007/BFb0096765;
 http://www.springerlink.com/content/978-3-540-69639-
1.

Cembranos:1997:C

[CM97d] Pilar Cembranos and José Mendoza. Copies of c_{0} and ℓ_{1} in L_{p}(μ, X). Lecture Notes in Mathematics, 1676:41–63, 1997. CO-DEN LNMAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096768/.

Cembranos:1997:I

[CM97e] Pilar Cembranos and José Mendoza. Introduction. Lecture Notes in Mathematics, 1676:1–8, 1997. CODEN LNMAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096766/.

Cembranos:1997:P

[CM97f] Pilar Cembranos and José Mendoza. Preliminaries. Lecture Notes in Mathematics, 1676:9–40, 1997. CODEN LNMAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096767/.

Cembranos:1997:SRO

[CM97g] Pilar Cembranos and José Mendoza. Some related open problems. Lecture Notes in Mathematics, 1676:107–109, 1997. CODEN LNMAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096773/.

Cembranos:1997:Sc

[CM97h] Pilar Cembranos and José Mendoza. The space L_{∞}(μ, X). Lecture Notes in Mathematics, 1676:83–104, 1997. CODEN LN-MAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096771/.

Cembranos:1997:TR

[CM97i] Pilar Cembranos and José Mendoza. Tabulation of results. Lecture Notes in Mathematics, 1676:105–106, 1997. CODEN LN-MAA2. ISBN 3-540-63745-1 (print), 3-540-69639-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096772.

Chaleyat-Maurel:1990:PSH

[CML90] M. Chaleyat-Maurel and J.-F. Le Gall. On polar sets for hypoelliptic diffusion processes. Lecture Notes in Mathematics, 1444: 204–212, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083617/.

Cieslak:1991:ICS

[CMM91] Waldemar Cieślak, Andrzej Miernowski, and Witold Mozgawa. Isoptics of a closed strictly convex curve. Lecture Notes in Mathematics, 1481:28–35, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083625/.

Chaleyat-Maurel:1995:OMF

[CMN95] Mireille Chaleyat-Maurel and David Nualart. Onsager–Machlup functionals for solutions of stochastic boundary value problems. Lecture Notes in Mathematics, 1613:44–55, 1995. CODEN LN-MAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094199/.

Coste:1992:RAG

[CMR92] Michel Coste, Louis Mahé, and Marie-Françoise Roy, editors. Real Algebraic Geometry: Proceedings of the Conference held in Rennes, France, June 24–28, 1991, volume 1524 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1524.

URL http://link.springer.com/book/10.1007/BFb0084605;
 http://www.springerlink.com/content/978-3-540-47337-
4.

Capietto:1991:CDS

[CMZ91] Anna Capietto, Jean Mawhin, and Fabio Zanolin. The coincidence degree of some functional differential operators in spaces of periodic functions and related continuation theorems. Lecture Notes in Mathematics, 1475:76–87, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083481/.

Conti:1991:SSV

[CNZ91] G. Conti, P. Nistri, and P. Zecca. Systems of set-valued equations in Banach spaces. Lecture Notes in Mathematics, 1475: 98–109, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083483/.

Coates:1999:FIT

[Coa99] John Coates. Fragments of the GL_{2} Iwasawa theory of elliptic curves without complex multiplication. Lecture Notes in Mathematics, 1716:1–50, 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093452/.

Cockburn:1998:IDG

[Coc98] Bernardo Cockburn. An introduction to the Discontinuous Galerkin method for convection-dominated problems. Lecture Notes in Mathematics, 1697:151–268, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0096353/.

Cohen:1990:NCP

[Coh90] F. R. Cohen. A note concerning the v_{1}-periodic homotopy of odd spheres. Lecture Notes in Mathematics, 1418:152–155, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083700/.

Cohen:1992:RHT

[Coh92] F. R. Cohen. Remarks on the homotopy theory associated to perfect groups. Lecture Notes in Mathematics, 1509:95–103, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087503/.

Cohen:1995:SMP

[Coh95] S. Cohen. Some Markov properties of stochastic differential equations with jumps. Lecture Notes in Mathematics, 1613:181– 193, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094210/.

Coleman:1990:FMF

[Col90a] Robert F. Coleman. On the Frobenius matrices of Fermat curves. Lecture Notes in Mathematics, 1454:173–193, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091138/.

Colley:1990:SCF

[Col90b] Susan Jane Colley. Schubert’s coincidence formulas for line complexes and the contribution of embedded planar pencils. Lecture Notes in Mathematics, 1436:60–76, 1990. CODEN LN-MAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084041/.

Colombeau:1992:ATG

[Col92a] Jean François Colombeau. The abstract theory of generalized functions. Lecture Notes in Mathematics, 1532:158–171, 1992.

CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088960/.

Colombeau:1992:CHF

[Col92b] Jean François Colombeau. The canonical Hamiltonian formalism of interacting quantum fields. Lecture Notes in Mathematics, 1532:144–157, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088959/.

Colombeau:1992:CSC

[Col92c] Jean François Colombeau. The case of several constitutive equations. Lecture Notes in Mathematics, 1532:97–123, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088957/.

Colombeau:1992:EI

[Col92d] Jean François Colombeau. Elementary introduction. Lecture Notes in Mathematics, 1532:30–61, 1992. CODEN LN-MAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088955/.

Colombeau:1992:IGF

[Col92e] Jean François Colombeau. Introduction to generalized functions and distributions. Lecture Notes in Mathematics, 1532:1–12, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088953/.

Colombeau:1992:JFS

[Col92f] Jean François Colombeau. Jump formulas for systems in nonconservative form. New numerical methods. Lecture Notes in Mathematics, 1532:62–96, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088956/.

Colombeau:1992:LWP

[Col92g] Jean François Colombeau. Linear wave propagation in a medium with piecewise C^{∞} characteristics. Lecture Notes in Mathematics, 1532:124–143, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088958/.

Colombeau:1992:MDT

[Col92h] Jean-François Colombeau. Multiplication of Distributions: A tool in mathematics, numerical engineering and theoretical physics, volume 1532 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 184 pp. LCCN QA3 .L28 no. 1532. URL http:

//link.springer.com/book/10.1007/BFb0088952;   http:/
/www.springerlink.com/content/978-3-540-47510-1.

Colombeau:1992:MDC

[Col92i] Jean François Colombeau. Multiplications of distributions in classical physics. Lecture Notes in Mathematics, 1532:13–29, 1992. CODEN LNMAA2. ISBN 3-540-56288-5 (print), 3-540-47510-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088954/.

Connes:1992:NG

[Con92] Alain Connes. Noncommutative geometry. Lecture Notes in Mathematics, 1525:40–58, 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089204/.

Conway:1994:CNO

[Con94] John B. Conway. Close to normal operators. Lecture Notes in Mathematics, 1573:359–388, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100209/.

Cooper:1990:ERN

[Coo90] S. Barry Cooper. Enumeration reducibility, nondeterministic computations and relative computability of partial functions. Lecture Notes in Mathematics, 1432:57–110, 1990. CODEN LN-MAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086114/.

Corduneanu:1991:AVE

[Cor91] C. Corduneanu. Abstract Volterra equations and weak topologies. Lecture Notes in Mathematics, 1475:110–115, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083484/.

Coste:1990:SEA

[Cos90a] Michel Coste. Sous-ensembles algébriques réels de codimension 2. (French) []. Lecture Notes in Mathematics, 1420:111–120, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083914/.

Coster:1990:S

[Cos90b] Matthijs J. Coster. Supercongruences. Lecture Notes in Mathematics, 1454:194–204, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091139/.

Couveignes:1993:CSR

[Cou93] Jean-Marc Couveignes. Computing a square root for the number field sieve. Lecture Notes in Mathematics, 1554:95–102, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091540/.

Cacuci:1991:CPN

[CP91a] Dan G. Cacuci and V. Protopopescu. Canonical propagators for nonlinear systems: Theory and sample applications. Lecture Notes in Mathematics, 1460:43–56, 1991. CODEN LN-MAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091360/.

Courtieu:1991:NAA

[CP91b] Michel Courtieu and Alexei A. Panchishkin. 1. Non-Archimedean analytic functions, measures and distributions. Lecture Notes in Mathematics, 1471:13–44, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45178-5_2/.

Courtieu:1991:SMF

[CP91c] Michel Courtieu and Alexei A. Panchishkin. 2. Siegel modular forms and the holomorphic projection operator. Lecture Notes in Mathematics, 1471:45–93, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45178-5_3/.

Courtieu:1991:ADO

[CP91d] Michel Courtieu and Alexei A. Panchishkin. 3. Arithmetical differential operators on nearly holomorphic Siegel modular forms. Lecture Notes in Mathematics, 1471:95–125, 1991. CODEN LN-MAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45178-5_
4/.

Courtieu:1991:AMS

[CP91e] Michel Courtieu and Alexei A. Panchishkin. 4. Admissible measures for standard L-functions and nearly holomorphic Siegel modular forms. Lecture Notes in Mathematics, 1471:127–186, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45178-5_5/.

Courtieu:1991:I

[CP91f] Michel Courtieu and Alexei A. Panchishkin. Introduction. Lecture Notes in Mathematics, 1471:1–12, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45178-5_1/.

Courtieu:1991:NAF

[CP91g] Michel Courtieu and Alexei A. Panchishkin. Non-Archimedean L-Functions and Arithmetical Siegel Modular Forms, volume 1471 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second augmented edition, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 157 pp. LCCN QA3 .L28 no. 1471. URL

http://link.springer.com/book/10.1007/b13348;   http:
//www.springerlink.com/content/978-3-540-45178-5.

Courtieu:1991:R

[CP91h] Michel Courtieu and Alexei A. Panchishkin. References. Lecture Notes in Mathematics, 1471:187–193, 1991. CODEN LNMAA2. ISBN 3-540-40729-4 (print), 3-540-45178-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45178-5_6/.

Cenkl:1992:ACC

[CP92] B. Cenkl and R. Porter. Algorithm for the computation of the cohomology of ℑ-groups. Lecture Notes in Mathematics, 1509: 79–94, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087502/.

Cercignani:1993:NPMa

[CP93a] C. Cercignani and M. Pulvirenti. Nonequilibrium problems in many-particle systems. An introduction. Lecture Notes in Mathematics, 1551:1–13, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090927/.

Cercignani:1993:NPMb

[CP93b] Carlo Cercignani and Mario Pulvirenti, editors. Nonequilibrium Problems in Many-Particle Systems: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini, Italy, June 15–27, 1992, volume 1551 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1551. URL http:/

/link.springer.com/book/10.1007/BFb0090926;   http://
www.springerlink.com/content/978-3-540-47832-4.

Coornaert:1993:AFP

[CP93c] Michel Coornaert and Athanase Papadopoulos. Another finite presentation for the action of a hyperbolic group on its boundary. Lecture Notes in Mathematics, 1539:69–90, 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092582/.

Coornaert:1993:BHG

[CP93d] Michel Coornaert and Athanase Papadopoulos. The boundary of a hyperbolic group as a finitely presented dynamical system. Lecture Notes in Mathematics, 1539:43–68, 1993. CODEN LN-MAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092581/.

Coornaert:1993:BTF

[CP93e] Michel Coornaert and Athanase Papadopoulos. The boundary of a torsion-free hyperbolic group as a semi-Markovian space. Lecture Notes in Mathematics, 1539:118–134, 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092585/.

Coornaert:1993:I

[CP93f] Michel Coornaert and Athanase Papadopoulos. Introduction. Lecture Notes in Mathematics, 1539:1–4, 1993. CODEN LN-MAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092578/.

Coornaert:1993:QRG

[CP93g] Michel Coornaert and Athanase Papadopoulos. A quick review of Gromov hyperbolic spaces. Lecture Notes in Mathematics, 1539: 5–18, 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092579/.

Coornaert:1993:SMS

[CP93h] Michel Coornaert and Athanase Papadopoulos. Semi-Markovian spaces. Lecture Notes in Mathematics, 1539:107–117, 1993. CO-DEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092584/.

Coornaert:1993:SD

[CP93i] Michel Coornaert and Athanase Papadopoulos. Symbolic dynamics. Lecture Notes in Mathematics, 1539:19–42, 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092580/.

Coornaert:1993:SDH

[CP93j] Michel Coornaert and Athanase Papadopoulos. Symbolic Dynamics and Hyperbolic Groups, volume 1539 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 138 pp. LCCN QA3 .L28 no. 1539.

URL http://link.springer.com/book/10.1007/BFb0092577;
 http://www.springerlink.com/content/978-3-540-47573-
6.

Coornaert:1993:THB

[CP93k] Michel Coornaert and Athanase Papadopoulos. Trees and hyperbolic boundary. Lecture Notes in Mathematics, 1539:91–106, 1993. CODEN LNMAA2. ISBN 3-540-56499-3 (print), 3-540-47573-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092583/.

Carboni:1991:CTP

[CPR91] Aurelio Carboni, Maria Cristina Pedicchio, and Guiseppe Rosolini, editors. Category Theory: Proceedings of the International Conference held in Como, Italy, July 22–28, 1990, volume 1488 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1488. URL http:/

/link.springer.com/book/10.1007/BFb0084207;   http://
www.springerlink.com/content/978-3-540-46435-8.

Cogdell:1990:MCK

[CPS90] J. W. Cogdell and I. I. Piatetski-Shapiro. The meromorphic continuation of Kloosterman–Selberg zeta functions. Lecture Notes in Mathematics, 1422:23–35, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089402/.

Cattabriga:1991:MAA

[CR91] Lamberto Cattabriga and Luigi Rodino, editors. Microlocal Analysis and Applications: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Montecatini Terme, Italy, July 3–11, 1989, volume 1495 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1495.

URL http://link.springer.com/book/10.1007/BFb0085120;
 http://www.springerlink.com/content/978-3-540-46603-
1.

Coste:1992:TFF

[CR92] Michel Coste and Miloud Reguiat. Trivialités en famille. (French) []. Lecture Notes in Mathematics, 1524:193–204, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084620/.

Crauel:1991:LER

[Cra91] Hans Crauel. Lyapunov exponents of random dynamical systems on Grassmannians. Lecture Notes in Mathematics, 1486: 38–50, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086656/.

Crandall:1997:VSP

[Cra97] Michael G. Crandall. Viscosity solutions: a primer. Lecture Notes in Mathematics, 1660:1–43, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094294/.

Coates:1999:FM

[CRGR99] John H. Coates, Kenneth A. Ribet, Ralph Greenberg, and Karl Rubin. Front matter. Lecture Notes in Mathematics, 1716: ??, 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-48160-7/1.pdf.

Cristante:1990:WRA

[Cri90] Valentino Cristante. Witt realization of the p-adic Barsotti–Tate groups; some applications. Lecture Notes in Mathematics, 1454: 205–216, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091140/.

Cruz:1990:PKD

[Cru90] Ricardo N. Cruz. Periodic knots and desuspensions of free involutions on spheres. Lecture Notes in Mathematics, 1440:39–47, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084448/.

Cavaliere:1990:SGA

[CRV90] Maria Pia Cavaliere, Maria Evelina Rossi, and Giuseppe Valla. On short graded algebras. Lecture Notes in Mathematics, 1430: 21–31, 1990. CODEN LNMAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085534/.

Campbell:1990:ESD

[CRW90] C. M. Campbell, E. F. Robertson, and P. D. Williams. On the efficiency of some direct powers of groups. Lecture Notes in Mathematics, 1456:106–113, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100734/.

Cavaretta:1992:VDP

[CS92a] A. S. Cavaretta, Jr. and A. Sharma. Variation diminishing properties and convexity for the tensor product Bernstein operator.

Lecture Notes in Mathematics, 1511:18–32, 1992. CODEN LN-MAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093794/.

Cavicchioli:1992:CMS

[CS92b] Alberto Cavicchioli and Fulvia Spaggiari. The classification of 3manifolds with spines related to Fibonacci groups. Lecture Notes in Mathematics, 1509:50–78, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087501/.

Charzynski:1992:PPC

[CS92c] Zygmunt Charzyński and Przemysław Skibiński. Pseudoorthogonality of powers of the coordinates of a holomorphic mapping in two variables with the constant Jacobian. Lecture Notes in Mathematics, 1524:178–192, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084619/.

Choulli:1996:DAD

[CS96] Tahir Choulli and Christophe Stricker. Deux applications de la décomposition de Galtchouk–Kunita–Watanabe. (French) []. Lecture Notes in Mathematics, 1626:12–23, 1996. CODEN LN-MAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094638/.

Choulli:1998:SSP

[CS98] Tahir Choulli and Christophe Stricker. Séparation d’une suret d’une sousmartingale par une martingale. (French) []. Lecture Notes in Mathematics, 1686:67–72, 1998. CODEN LN-MAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101751/.

Cockburn:1998:FM

[CSJT98] Bernardo Cockburn, Chi-Wang Shu, Claes Johnson, and Eitan Tadmor. Front matter. Lecture Notes in Mathematics, 1697: ??, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-49804-9/1.pdf.

Sun:1991:HDC

[cSyW91] Dao chun Sun and Zhi ying Wen. The Hausdorff dimension of a class of lacunary trigonometric series. Lecture Notes in Mathematics, 1494:176–181, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087769/.

Catanese:1992:VBL

[CT92] F. Catanese and F. Tovena. Vector bundles, linear systems and extensions of π_{1}. Lecture Notes in Mathematics, 1507:51–70, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094510/.

Colliot-Thelene:1993:CAT

[CT93] Jean-Louis Colliot-Thélène. Cycles algébriques de torsion et Kthéorie algébrique. Cours au C.I.M.E., juin 1991. (French) []. Lecture Notes in Mathematics, 1553:1–49, 1993. CODEN LN-MAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084728/.

Colliot-Thelene:1993:AAG

[CTKV93a] Jean-Louis Colliot-Thélène, Kazuya Kato, and Paul Vojta. Arithmetic algebraic geometry, Trento, Italy 1991. Lecture Notes in Mathematics, 1553:e1–e2, 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0120521.

Colliot-Thelene:1993:BM

[CTKV93b] Jean-Louis Colliot-Thélène, Kazuya Kato, and Paul Vojta. Back matter. Lecture Notes in Mathematics, 1553:??, 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-47909-3/
1.pdf.

Colliot-Thelene:1993:FM

[CTKV93c] Jean-Louis Colliot-Thélène, Kazuya Kato, and Paul Vojta. Front matter. Lecture Notes in Mathematics, 1553:??, 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-47909-3/
1.pdf.

Cocozza-Thivent:1996:CLS

[CTR96] C. Cocozza-Thivent and M. Roussignol. Comparaison des lois stationnaires et quasi-stationnaires d’un processus de Markov et application à la fiabilité. (French) []. Lecture Notes in Mathematics, 1626:24–39, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094639/.

Carpenter:1993:SNR

[CV93] Amos J. Carpenter and Richard S. Varga. Some numerical results on best uniform polynomial approximation of X^{α} on [0, 1]. Lecture Notes in Mathematics, 1550:192–222, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117488/.

Ciliberto:1992:JHS

[CvdG92] Ciro Ciliberto and Gerard van der Geer. On the Jacobian of a hyperplane section of a surface. Lecture Notes in Mathematics, 1515:33–40, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098336/.

Cushman:1990:HHB

[CvdM90] R. Cushman and J.-C. van der Meer. The Hamiltonian Hopf bifurcation in the Lagrange top. Lecture Notes in Mathematics, 1416:26–38, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097463/.

Cvitanic:1997:OTU

[Cvi97] Jakša Cvitanić. Optimal trading under constraints. Lecture Notes in Mathematics, 1656:123–190, 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092000/.

Carmody:1991:CQA

[CW91a] S. Carmody and R. F. C. Walters. Computing quotients of actions of a free category. Lecture Notes in Mathematics, 1488: 63–78, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084213/.

Cooke:1991:SDE

[CW91b] Kenneth L. Cooke and Joseph Wiener. A survey of differential equations with piecewise continuous arguments. Lecture Notes in Mathematics, 1475:1–15, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083475/.

DaPrato:1999:SPK

[Da 99] Giueppe Da Prato, editor. Stochastic PDE’s and Kolmogorov Equations in Infinite Dimensions: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro, Italy, August 24–September 1, 1998, volume 1715 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66545-5 (print), 3-540-48161-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1715. URL http:/

/link.springer.com/book/10.1007/BFb0092416;   http://
www.springerlink.com/content/978-3-540-48161-4.

Dafermos:1996:ESC

[Daf96] C. M. Dafermos. Entropy and the stability of classical solutions of hyperbolic systems of conservation laws. Lecture Notes in Mathematics, 1640:48–69, 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093706/.

Damon:1991:EES

[Dam91a] James Damon. - equivalence and the equivalence of sections of images and discriminants. Lecture Notes in Mathematics, 1462: 93–121, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086377/.

Damon:1991:EBM

[Dam91b] James Damon. Equivariant bifurcations and morsifications for finite groups. Lecture Notes in Mathematics, 1463:80–106, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085427/.

Damailly:1992:SHM

[Dam92] Jean-Pierre Damailly. Singular Hermitian metrics on positive line bundles. Lecture Notes in Mathematics, 1507:87–104, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094512/.

Danilov:1991:RCT

[Dan91] V. I. Danilov. De Rham complex on toroidal variety. Lecture Notes in Mathematics, 1479:26–38, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086261/.

Darling:1991:LEP

[Dar91] R. W. R. Darling. The Lyapunov exponent for products of infinitedimensional random matrices. Lecture Notes in Mathematics, 1486:206–215, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086670/.

Datuashvili:1990:HDE

[Dat90] Tamar Datuashvili. Homological dimension of extensions of abelian categories and rings. Lecture Notes in Mathematics, 1437: 1–35, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086717/.

Dauns:1990:SMR

[Dau90] John Dauns. Semiprime modules and rings. Lecture Notes in Mathematics, 1448:41–62, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091250/.

David:1990:NDD

[Dav90] Guy David. Une nouvelle démonstration du théorème T (b), d’après Coifman et Semmes. (French) []. Lecture Notes in Mathematics, 1438:39–50, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083514/.

David:1991:SIO

[Dav91a] Guy David. Singular integral operators. Lecture Notes in Mathematics, 1465:26–54, 1991. CODEN LNMAA2. ISBN 3-540-53902-6 (print), 3-540-46377-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091546/.

David:1991:SIC

[Dav91b] Guy David. Singular integrals on curves and surfaces. Lecture Notes in Mathematics, 1465:55–92, 1991. CODEN LN-MAA2. ISBN 3-540-53902-6 (print), 3-540-46377-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091547/.

David:1991:W

[Dav91c] Guy David. Wavelets. Lecture Notes in Mathematics, 1465:1– 25, 1991. CODEN LNMAA2. ISBN 3-540-53902-6 (print), 3-540-46377-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091545/.

David:1991:WSI

[Dav91d] Guy David. Wavelets and Singular Integrals on Curves and Surfaces, volume 1465 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53902-6 (print), 3-540-46377-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 107 pp. LCCN QA3 .L28 no. 1465. URL http:

//link.springer.com/book/10.1007/BFb0091544;   http:/
/www.springerlink.com/content/978-3-540-46377-1.

Dawson:1993:MVM

[Daw93] Donald A. Dawson. Measure-valued Markov processes. Lecture Notes in Mathematics, 1541:1–260, 1993. CODEN LN-MAA2. ISBN 3-540-56622-8 (print), 3-540-47608-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084190/.

Dayal:1992:HFD

[Day92] Saheb Dayal. Higher Fréchet and discrete Gâteaux differentiability of n-convex functions on Banach spaces. Lecture Notes in Mathematics, 1511:154–171, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093809/.

Dazord:1990:GST

[Daz90] Pierre Dazord. Groupoídes symplectiques et troisième théorème de Lie ‘non linéaire’. (French) []. Lecture Notes in Mathematics, 1416:39–74, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097464/.

deBartolomeis:1991:GTR

[dBT91] Paolo de Bartolomeis and Franco Tricerri, editors. Geometric Topology: Recent Developments: Lectures given on the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Montecatini Terme, Italy, June 4–12, 1990, volume 1504 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-55017-8 (print), 3-540-46651-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1504. URL http:/

/link.springer.com/book/10.1007/BFb0094287;   http://
www.springerlink.com/content/978-3-540-46651-2.

Diener:1991:MD

[DD91] Francine Diener and Marc Diener. Maximal delay. Lecture Notes in Mathematics, 1493:71–86, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085025/.

Deslauriers:1990:SII

[DDD90] Gilles Deslauriers, Jacques Dubois, and Serge Dubuc. Schéma itératif d’interpolation. (French) [Iterative interpolation scheme]. Lecture Notes in Mathematics, 1438:93–101, 1990. CODEN LN-MAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083517/.

Diekmann:1999:FM

[DDH^{+}99] Odo Diekmann, Richard Durrett, Karl Peter Hadeler, Philip K. Maini, and Hal Smith. Front matter. Lecture Notes in Mathematics, 1714:??, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-48170-6/1.pdf.

DePierro:1991:MIM

[De 91] Alvaro R. De Pierro. Multiplicative iterative methods in computed tomography. Lecture Notes in Mathematics, 1497:167– 186, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084517/.

DeSamLazaro:1996:CET

[De 96] J. De Sam Lazaro. Un contre-exemple touchant à l’indépendance. (French) []. Lecture Notes in Mathematics, 1626:100–103, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094643/.

Debarre:1992:VSE

[Deb92] Olivier Debarre. Vers une stratification de l’espace des modules des variétés abeliennes principalement polarisées. (French) []. Lecture Notes in Mathematics, 1507:71–86, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094511/.

Degtyarev:1992:SOR

[Deg92] A. I. Degtyarev. Stiefel orientations on a real algebraic variety. Lecture Notes in Mathematics, 1524:205–220, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084621/.

Dekimpe:1996:ACA

[Dek96a] Karel Dekimpe. Algebraic characterizations of almost-crystallographic groups. Lecture Notes in Mathematics, 1639:31–46, 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094475/.

Dekimpe:1996:ABG

[Dek96b] Karel Dekimpe. Almost-Bieberbach Groups: Affine and Polynomial Structures, volume 1639 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 259 pp. LCCN QA3 .L28 no. 1639.

URL http://link.springer.com/book/10.1007/BFb0094472;
 http://www.springerlink.com/content/978-3-540-49564-
2.

Dekimpe:1996:CTR

[Dek96c] Karel Dekimpe. Canonical type representations. Lecture Notes in Mathematics, 1639:47–102, 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094476/.

Dekimpe:1996:CS

[Dek96d] Karel Dekimpe. Classification survey. Lecture Notes in Mathematics, 1639:159–230, 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094479/.

Dekimpe:1996:CVN

[Dek96e] Karel Dekimpe. The cohomology of virtually nilpotent groups. Lecture Notes in Mathematics, 1639:103–120, 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094477/.

Dekimpe:1996:INA

[Dek96f] Karel Dekimpe. Infra-nilmanifolds and almost-Bieberbach groups. Lecture Notes in Mathematics, 1639:13–30, 1996. CODEN LN-MAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094474/.

Dekimpe:1996:INT

[Dek96g] Karel Dekimpe. Infra-nilmanifolds and their topological invariants. Lecture Notes in Mathematics, 1639:121–157, 1996. CODEN LNMAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094478/.

Dekimpe:1996:PNC

[Dek96h] Karel Dekimpe. Preliminaries and notational conventions. Lecture Notes in Mathematics, 1639:1–11, 1996. CODEN LN-MAA2. ISBN 3-540-61899-6 (print), 3-540-49564-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094473/.

Dellacherie:1990:MST

[Del90a] C. Dellacherie. Modèles simples de la théorie du potentiel non linéaire. (French) []. Lecture Notes in Mathematics, 1426:63– 104, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083758/.

Dellacherie:1990:TPP

[Del90b] C. Dellacherie. Théorie des processus de production. (French) []. Lecture Notes in Mathematics, 1426:52–62, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083757/.

Delcroix:1991:TLS

[Del91a] Antoine Delcroix. A tool for the local study of slow-fast vector fields: The zoom. Lecture Notes in Mathematics, 1493:151– 167, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085029/.

Delfs:1991:ALS

[Del91b] Hans Delfs. Abstract locally semialgebraic spaces. Lecture Notes in Mathematics, 1484:1–16, 1991. CODEN LNMAA2. ISBN 3-540-54615-4 (print), 3-540-38494-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093940/.

Delfs:1991:HLS

[Del91c] Hans Delfs. Homology of Locally Semialgebraic Spaces, volume 1484 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54615-4 (print), 3-540-38494-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 136 pp. LCCN QA3 .L28 no. 1484. URL http:

//link.springer.com/book/10.1007/BFb0093939;   http:/
/www.springerlink.com/content/978-3-540-38494-6.

Delfs:1991:SBM

[Del91d] Hans Delfs. Semialgebraic Borel–Moore-homology. Lecture Notes in Mathematics, 1484:62–114, 1991. CODEN LNMAA2. ISBN 3-540-54615-4 (print), 3-540-38494-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093942/.

Delfs:1991:STL

[Del91e] Hans Delfs. Sheaf theory on locally semialgebraic spaces. Lecture Notes in Mathematics, 1484:17–61, 1991. CODEN LN-MAA2. ISBN 3-540-54615-4 (print), 3-540-38494-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093941/.

Delfs:1991:SIT

[Del91f] Hans Delfs. Some intersection theory. Lecture Notes in Mathematics, 1484:115–129, 1991. CODEN LNMAA2. ISBN 3-540-54615-4 (print), 3-540-38494-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093943/.

Dellacherie:1991:TNL

[Del91g] C. Dellacherie. Théorie non linéaire du potentiel: Un principe unifié de domination et du maximum et quelques applications. (French) []. Lecture Notes in Mathematics, 1485:1–9, 1991. CO-DEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100840/.

Delort:1992:BM

[Del92a] Jean-Marc Delort. Back matter. Lecture Notes in Mathematics, 1522:99–101, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21539-5/1.pdf.

Delort:1992:FBT

[Del92b] Jean-Marc Delort. F.B.I. Transformation: Second Microlocalization and Semilinear Caustics, volume 1522 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 101 pp. LCCN QA3 .L28 no.

1522. URL http://link.springer.com/book/10.1007/978-
3-662-21539-5;   http://www.springerlink.com/content/
978-3-662-21539-5.

Delort:1992:FBI

[Del92c] Jean-Marc Delort. Fourier–Bros–Iagolnitzer transformation and first microlocalization. Lecture Notes in Mathematics, 1522:7– 27, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21539-5_2/.

Delort:1992:FM

[Del92d] Jean-Marc Delort. Front matter. Lecture Notes in Mathematics, 1522:i–vi, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21539-5/1.pdf.

Delort:1992:GUB

[Del92e] Jean-Marc Delort. Geometric upper bounds. Lecture Notes in Mathematics, 1522:47–71, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21539-5_4/.

Delort:1992:I

[Del92f] Jean-Marc Delort. Introduction. Lecture Notes in Mathematics, 1522:1–6, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21539-5_1/.

Delort:1992:SM

[Del92g] Jean-Marc Delort. Second microlocalization. Lecture Notes in Mathematics, 1522:28–46, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21539-5_3/.

Delort:1992:SCP

[Del92h] Jean-Marc Delort. Semilinear Cauchy problem. Lecture Notes in Mathematics, 1522:72–98, 1992. CODEN LNMAA2. ISBN 3-540-55764-4 (print), 3-662-21539-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21539-5_5/.

deLaubenfels:1994:AMG

[deL94a] Ralph deLaubenfels. Arbitrary matrices of generators of bounded strongly continuous groups. Lecture Notes in Mathematics, 1570: 92–93, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073415.

deLaubenfels:1994:BVH

[deL94b] Ralph deLaubenfels. Boundary values of holomorphic semigroups. Lecture Notes in Mathematics, 1570:73–75, 1994. CODEN LN-MAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073411/.

deLaubenfels:1994:RHY

[deL94c] Ralph deLaubenfels. C-resolvents and Hille–Yosida type theorems. Lecture Notes in Mathematics, 1570:104–109, 1994. CO-DEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073418/.

deLaubenfels:1994:CPL

[deL94d] Ralph deLaubenfels. The Cauchy problem for the Laplace equation. Lecture Notes in Mathematics, 1570:71–72, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073410.

deLaubenfels:1994:EVE

[deL94e] Ralph deLaubenfels. Entire vectors and entire existence families. Lecture Notes in Mathematics, 1570:60–65, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073408/.

deLaubenfels:1994:EUF

[deL94f] Ralph deLaubenfels. Existence and uniqueness families. Lecture Notes in Mathematics, 1570:97–103, 1994. CODEN LN-MAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073417/.

deLaubenfels:1994:EF

[deL94g] Ralph deLaubenfels. Existence families. Lecture Notes in Mathematics, 1570:7–12, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073403/.

deLaubenfels:1994:EFF

[deL94h] Ralph deLaubenfels. Existence Families, Functional Calculi and Evolution Equations, volume 1570 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 234 pp. LCCN QA3 .L28 no. 1570.

URL http://link.springer.com/book/10.1007/BFb0073401;
 http://www.springerlink.com/content/978-3-540-48322-
9.

deLaubenfels:1994:EBB

[deL94i] Ralph deLaubenfels. Exponentially bounded (Banach) solution spaces. Lecture Notes in Mathematics, 1570:38–54, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073406/.

deLaubenfels:1994:FCC

[deL94j] Ralph deLaubenfels. Functional calculus for commuting generators of bounded strongly continuous groups. Lecture Notes in Mathematics, 1570:79–85, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073413/.

deLaubenfels:1994:IEE

[deL94k] Ralph deLaubenfels. Intuition and elementary examples. Lecture Notes in Mathematics, 1570:1–6, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073402/.

deLaubenfels:1994:MER

[deL94l] Ralph deLaubenfels. More examples of regularized semigroups. Lecture Notes in Mathematics, 1570:94–96, 1994. CODEN LN-MAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073416/.

deLaubenfels:1994:P

[deL94m] Ralph deLaubenfels. Perturbations. Lecture Notes in Mathematics, 1570:113–124, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073420/.

deLaubenfels:1994:PCM

[deL94n] Ralph deLaubenfels. Petrovsky correct matrices of generators of bounded strongly continuous groups. Lecture Notes in Mathematics, 1570:86–91, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073414/.

deLaubenfels:1994:RS

[deL94o] Ralph deLaubenfels. Regularized semigroups. Lecture Notes in Mathematics, 1570:13–23, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073404/.

deLaubenfels:1994:RIS

[deL94p] Ralph deLaubenfels. Relationship to integrated semigroups. Lecture Notes in Mathematics, 1570:110–112, 1994. CODEN LN-MAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073419/.

deLaubenfels:1994:RPP

[deL94q] Ralph deLaubenfels. Reversibility of parabolic problems. Lecture Notes in Mathematics, 1570:66–70, 1994. CODEN LN-MAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073409/.

deLaubenfels:1994:SE

[deL94r] Ralph deLaubenfels. The Schrödinger equation. Lecture Notes in Mathematics, 1570:76–78, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073412/.

deLaubenfels:1994:SSO

[deL94s] Ralph deLaubenfels. The solution space of an operator and automatic well-posedness. Lecture Notes in Mathematics, 1570:24– 37, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073405/.

deLaubenfels:1994:TO

[deL94t] Ralph deLaubenfels. Type of an operator. Lecture Notes in Mathematics, 1570:125–127, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073421/.

deLaubenfels:1994:WPL

[deL94u] Ralph deLaubenfels. Well-posedness on a larger space; generalized solutions. Lecture Notes in Mathematics, 1570:55–59, 1994. CODEN LNMAA2. ISBN 3-540-57703-3 (print), 3-540-48322-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073407/.

Delon:1998:SCF

[Del98] Françoise Delon. Separably closed fields. Lecture Notes in Mathematics, 1696:143–176, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-68521-0_9/.

Demailly:1996:VTP

[Dem96] Jean-Pierre Demailly. L^{2} vanishing theorems for positive line bundles and adjunction theory. Lecture Notes in Mathematics, 1646: 1–97, 1996. CODEN LNMAA2. ISBN 3-540-62038-9 (print), 3-540-49632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094302/.

Denkowska:1992:SSP

[Den92] Zofia Denkowska. Subanaliticity and the second part of Hilbert’s 16^{th} problem. Lecture Notes in Mathematics, 1524:221–234, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084622/.

Dermoune:1995:CSI

[Der95] A. Dermoune. Chaoticity on a stochastic interval [0, T]. Lecture Notes in Mathematics, 1613:117–124, 1995. CODEN LN-MAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094204/.

Dubins:1993:LTB

[DÉY93] L. E. Dubins, Michel Émery, and M. Yor. On the Lévy transformation of Brownian motions and continuous martingales. Lecture Notes in Mathematics, 1557:122–132, 1993. CODEN LN-MAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087970/.

deFoglio:1990:NOR

[dF90] Susana Fernandez Long de Foglio. New and old results concerning Henstock’s integrals. Lecture Notes in Mathematics, 1419: 25–37, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083097/.

Dethloff:1990:DCR

[DG90] Gerd Dethloff and Hans Grauert. Deformation of compact Riemann surfaces Y of genus p with distinguished points P_{1}, …, P_{m}∈Y. Lecture Notes in Mathematics, 1422:37–44, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089403/.

Degond:1991:SMB

[DGDMN91] P. Degond, F. Guyot-Delaurens, F. J. Mustieles, and F. Nier. Semiconductor modelling via the Boltzmann equation. Lecture Notes in Mathematics, 1460:87–106, 1991. CODEN LN-MAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091363/.

Diekmann:1991:SRE

[DGT91] Odo Diekmann, Mats Gyllenberg, and Horst R. Thieme. Semigroups and renewal equations on dual Banach spaces with applications to population dynamics. Lecture Notes in Mathematics, 1475:116–129, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083485/.

Dancer:1990:BSP

[DH90] E. N. Dancer and P. Hess. Behaviour of a semilinear periodicparabolic problem when a parameter is small. Lecture Notes in Mathematics, 1450:12–19, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084895/.

Dales:1994:BA

[DH94] H. G. Dales and A. Ya. Helemskii. Banach algebras. Lecture Notes in Mathematics, 1573:51–154, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100203/.

Dellacherie:1998:SMS

[DI98] C. Dellacherie and A. Iwanik. Sous-mesures symétriques sur un ensemble fini. (French) []. Lecture Notes in Mathematics, 1686: 1–5, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101745/.

Didierjean:1990:HEE

[Did90] G. Didierjean. Homotopie des espaces d’Équivalences. (French) [Homotopy of equivalent spaces]. Lecture Notes in Mathematics, 1425:32–39, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083828/.

Diekmann:1999:MAP

[Die99] Odo Diekmann. Modeling and analysing physiologically structured populations. Lecture Notes in Mathematics, 1714:1–37, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092374/.

Dimca:1991:DFH

[Dim91] Alexandru Dimca. Differential forms and hypersurface singularities. Lecture Notes in Mathematics, 1462:122–153, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086378/.

Dixmier:1991:IDG

[Dix91] Jacques Dixmier. Sur les invariants du groupe symétrique dans certaines représentations, II. (French) []. Lecture Notes in Mathematics, 1478:1–34, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083500/.

Dix:1997:A

[Dix97a] Daniel B. Dix. Applications. Lecture Notes in Mathematics, 1668: 155–193, 1997. CODEN LNMAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093373/.

Dix:1997:EIR

[Dix97b] Daniel B. Dix. Expansion in the inner region, matching. Lecture Notes in Mathematics, 1668:75–95, 1997. CODEN LN-MAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093370/.

Dix:1997:LEO

[Dix97c] Daniel B. Dix. Laplace expansions, outer regions. Lecture Notes in Mathematics, 1668:1–74, 1997. CODEN LNMAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093369/.

Dix:1997:LTB

[Dix97d] Daniel B. Dix. Large-time Behavior of Solutions of Linear Dispersive Equations, volume 1668 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 203 pp. LCCN QA3 .L28 no. 1668.

URL http://link.springer.com/book/10.1007/BFb0093368;
 http://www.springerlink.com/content/978-3-540-69545-
5.

Dix:1997:SRS

[Dix97e] Daniel B. Dix. Special results for special cases. Lecture Notes in Mathematics, 1668:114–154, 1997. CODEN LNMAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093372/.

Dix:1997:UVE

[Dix97f] Daniel B. Dix. Uniformly valid expansions as t → ∞. Lecture Notes in Mathematics, 1668:96–113, 1997. CODEN LN-MAA2. ISBN 3-540-63434-7 (print), 3-540-69545-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093371/.

deJong:1993:ALB

[dJ93] Johan de Jong. Ample line bundles and intersection theory. Lecture Notes in Mathematics, 1566:69–76, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48208-6_7/.

Downey:1990:ANS

[DJS90] Rod Downey, Carl Jockusch, and Michael Stob. Array nonrecursive sets and multiple permitting arguments. Lecture Notes in Mathematics, 1432:141–173, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086116/.

deJong:1991:D

[dJvS91] T. de Jong and D. van Straten. Disentanglements. Lecture Notes in Mathematics, 1462:199–211, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086383/.

Denker:1992:ULC

[DK92] M. Denker and K. F. Krämer. Upper and lower class results for subsequences of the champernowne number. Lecture Notes in Mathematics, 1514:83–89, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097529/.

Dimitrov:1993:SCE

[DK93a] B. Dimitrov and Z. Khalil. Some characterizations of the exponential distribution based on the service time properties of an unreliable server. Lecture Notes in Mathematics, 1546:17–25, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084478/.

Dobrushin:1993:FM

[DK93b] Roland Lvovich Dobrushin and Shigeo Kusuoka. Front matter. Lecture Notes in Mathematics, 1567:??, 1993. CODEN LNMAA2. ISBN 3-540-57516-2 (print), 3-540-48205-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-48205-5/1.pdf.

Dobrushin:1993:SMF

[DK93c] Roland Lvovich Dobrushin and Shigeo Kusuoka. Statistical Mechanics and Fractals, volume 1567 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57516-2 (print), 3-540-48205-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 98 pp. LCCN QA3 .L28 no. 1567.

URL http://link.springer.com/book/10.1007/BFb0074238;
 http://www.springerlink.com/content/978-3-540-48205-
5.

Dynkin:1994:SIB

[DK94] E. M. Dyn’kin and S. V. Kisliakov. Singular integrals, BMO, H^{p}. Lecture Notes in Mathematics, 1573:409–464, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100211/.

Denef:1990:PNP

[DL90] J. Denef and F. Loeser. Polyèdres de Newton et poids de sommes exponentielles. (French) []. Lecture Notes in Mathematics, 1454: 217–222, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091141/.

Dias:1996:CSH

[DL96a] Danielle Dias and Patrick Le Barz. Configuration Spaces over Hilbert Schemes and Applications, volume 1647 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62050-8 (print), 3-540-49634-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 143 pp. LCCN QA3 .L28 no. 1647. URL http:/

/link.springer.com/book/10.1007/BFb0093653;   http://
www.springerlink.com/content/978-3-540-49634-2.

Dias:1996:CCQ

[DL96b] Danielle Dias and Patrick Le Barz. Construction of a complete quadruples variety. Lecture Notes in Mathematics, 1647: 66–128, 1996. CODEN LNMAA2. ISBN 3-540-62050-8 (print), 3-540-49634-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093656/.

Dias:1996:DTP

[DL96c] Danielle Dias and Patrick Le Barz. Double and triple points formula. Lecture Notes in Mathematics, 1647:10–64, 1996. CODEN LNMAA2. ISBN 3-540-62050-8 (print), 3-540-49634-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093655/.

Dias:1996:I

[DL96d] Danielle Dias and Patrick Le Barz. Introduction. Lecture Notes in Mathematics, 1647:1–8, 1996. CODEN LNMAA2. ISBN 3-540-62050-8 (print), 3-540-49634-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093654/.

Dolcetta:1997:VSA

[DL97] Italo Capuzzo Dolcetta and Pierre Louis Lions, editors. Viscosity Solutions and Applications: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, June 12–20, 1995, volume 1660 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1660. URL http:/

/link.springer.com/book/10.1007/BFb0094293;   http://
www.springerlink.com/content/978-3-540-69043-6.

delaRue:1993:ELF

[dlR93] Thierry de la Rue. Espaces de Lebesgue. (French) [Lebesgue spaces]. Lecture Notes in Mathematics, 1557:15–21, 1993. CO-DEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087958/.

Droms:1990:TCF

[DLS90] Carl Droms, Jacques Lewin, and Herman Servatius. The Tits conjecture and the five string braid group. Lecture Notes in Mathematics, 1440:48–51, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084449/.

Davis:1990:PH

[DM90a] Donald M. Davis and Mark Mahowald. v_{1}-periodic homotopy of Sp(2), Sp(3), and S^{2n}. Lecture Notes in Mathematics, 1418: 219–237, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083706/.

Desolneux-Moulis:1990:DSH

[DM90b] N. Desolneux-Moulis. Dynamique des systèmes hamiltoniens complètement intégrables sur les variétés compactes. (French) []. Lecture Notes in Mathematics, 1416:75–83, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097465/.

Downey:1990:SD

[DM90c] Rod Downey and Joe Mourad. Superbranching degrees. Lecture Notes in Mathematics, 1432:175–186, 1990. CODEN LN-MAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086117/.

deMonvel:1993:ISD

[dM93a] L. Boutet de Monvel. Indice des systèmes différentiels. (French) []. Lecture Notes in Mathematics, 1565:1–30, 1993. CODEN LNMAA2. ISBN 3-540-57498-0 (print), 3-540-48195-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073465/.

DellAntonio:1993:DFL

[DM93b] Gianfausto Dell’Antonio and Umberto Mosco, editors. Dirichlet Forms: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Varenna, Italy, June 8–19, 1992, volume 1563 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1563. URL http:/

/link.springer.com/book/10.1007/BFb0074088;   http://
www.springerlink.com/content/978-3-540-48151-5.

Donagi:1996:SCA

[DM96] Ron Donagi and Eyal Markman. Spectral covers, algebraically completely integrable, Hamiltonian systems, and moduli of bundles. Lecture Notes in Mathematics, 1620:1–119, 1996. CODEN LNMAA2. ISBN 3-540-60542-8 (print), 3-540-47706-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094792/.

deMeyer:1999:SAT

[dM99] Bernard de Meyer. Une simplification de l’argument de Tsirelson sur le caractère non-brownien des processus de Walsh. (French) []. Lecture Notes in Mathematics, 1709:217–220, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096512/.

Dudley:1999:DCQ

[DN99a] R. M. Dudley and R. Norvaiša. Differentiability of the composition and quantile operators for regulated and A. E. continuous functions. Lecture Notes in Mathematics, 1703:209–240, 1999. CODEN LNMAA2. ISBN 3-540-65975-7 (print), 3-540-48814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100747/.

Dudley:1999:PIY

[DN99b] R. M. Dudley and R. Norvaiša. Product integrals, Young integrals and p-variation. Lecture Notes in Mathematics, 1703:73– 208, 1999. CODEN LNMAA2. ISBN 3-540-65975-7 (print), 3-540-48814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100746/.

Dudley:1999:SDS

[DN99c] R. M. Dudley and R. Norvaiša. A survey on differentiability of six operators in relation to probability and statistics. Lecture Notes in Mathematics, 1703:1–72, 1999. CODEN LNMAA2. ISBN 3-540-65975-7 (print), 3-540-48814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100745/.

Dudley:1999:DSO

[DN99d] Richard M. Dudley and Rimas Norvaiša. Differentiability of Six Operators on Nonsmooth Functions and p-Variation, volume 1703 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-65975-7 (print), 3-540-48814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 277 pp. LCCN QA3 .L28 no. 1703. URL http:/

/link.springer.com/book/10.1007/BFb0100744;   http://
www.springerlink.com/content/978-3-540-48814-9.

Dudley:1999:BVV

[DNQ99] R. M. Dudley, R. Norvaiša, and Jinghua Qian. Bibliographies on p-variation and φ-variation. Lecture Notes in Mathematics, 1703:241–272, 1999. CODEN LNMAA2. ISBN 3-540-65975-7 (print), 3-540-48814-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100748/.

Dobrushin:1993:WMF

[Dob93] R. L. Dobrushin. On the way to the mathematical foundations of statistical mechanics. Lecture Notes in Mathematics, 1567:1– 37, 1993. CODEN LNMAA2. ISBN 3-540-57516-2 (print), 3-540-48205-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074239/.

Dobrushin:1996:PMT

[Dob96] R. L. Dobrushin. Perturbation methods of the theory of Gibbsian fields. Lecture Notes in Mathematics, 1648:1–66, 1996. CODEN LNMAA2. ISBN 3-540-62055-9 (print), 3-540-49635-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095674/.

Donato:1990:GOC

[Don90] Paul Donato. Géométrie des orbites coadjointes des groupes de difféomorphismes. (French) []. Lecture Notes in Mathematics, 1416:84–104, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097466/.

Doney:1998:SCP

[Don98] R. A. Doney. Some calculations for perturbed Brownian motion. Lecture Notes in Mathematics, 1686:231–236, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101760/.

Dore:1993:RAD

[Dor93] Giovanni Dore. L^{p} regularity for abstract differential equations. Lecture Notes in Mathematics, 1540:25–38, 1993. CODEN LN-MAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085472/.

Downey:1990:NPM

[Dow90] Rod Downey. Notes on the O′′′ priority method with special attention to density results. Lecture Notes in Mathematics, 1432: 111–140, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086115/.

DeMasi:1991:EUE

[DP91a] Anna De Masi and Errico Presutti. Escape from an unstable equilibrium. Lecture Notes in Mathematics, 1501:147–166, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086466/.

DeMasi:1991:EF

[DP91b] Anna De Masi and Errico Presutti. Estimates on the V -functions. Lecture Notes in Mathematics, 1501:167–188, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086467/.

DeMasi:1991:GKP

[DP91c] Anna De Masi and Errico Presutti. The Glauber + Kawasaki process. Lecture Notes in Mathematics, 1501:97–111, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086463/.

DeMasi:1991:HLI

[DP91d] Anna De Masi and Errico Presutti. Hydrodynamic limits for independent particles. Lecture Notes in Mathematics, 1501:7–32, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086459/.

DeMasi:1991:HLK

[DP91e] Anna De Masi and Errico Presutti. Hydrodynamic limits in kinetic models. Lecture Notes in Mathematics, 1501:112–127, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086464/.

DeMasi:1991:HZR

[DP91f] Anna De Masi and Errico Presutti. Hydrodynamics of the zero range process. Lecture Notes in Mathematics, 1501:33–51, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086460/.

DeMasi:1991:I

[DP91g] Anna De Masi and Errico Presutti. Introduction. Lecture Notes in Mathematics, 1501:1–6, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086458/.

DeMasi:1991:MMH

[DP91h] Anna De Masi and Errico Presutti. Mathematical Methods for Hydrodynamic Limits, volume 1501 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 196 pp. LCCN QA3 .L28 no. 1501.

URL http://link.springer.com/book/10.1007/BFb0086457;
 http://www.springerlink.com/content/978-3-540-46636-
9.

DeMasi:1991:PMR

[DP91i] Anna De Masi and Errico Presutti. Particle models for reactiondiffusion equations. Lecture Notes in Mathematics, 1501:52– 66, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086461/.

DeMasi:1991:PMC

[DP91j] Anna De Masi and Errico Presutti. Particle models for the Carleman equation. Lecture Notes in Mathematics, 1501:67–96, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086462/.

DeMasi:1991:PSI

[DP91k] Anna De Masi and Errico Presutti. Phase separation and interface dynamics. Lecture Notes in Mathematics, 1501:128–146, 1991. CODEN LNMAA2. ISBN 3-540-55004-6 (print), 3-540-46636-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086465/.

DeConcini:1993:QG

[DP93] C. De Concini and C. Procesi. Quantum groups. Lecture Notes in Mathematics, 1565:31–140, 1993. CODEN LNMAA2. ISBN 3-540-57498-0 (print), 3-540-48195-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073466/.

Decker:1990:HCS

[DPlPS90] Wolfram Decker, Thomas Peternell, Joseph le Potier, and Michael Schneider. Half-canonical surfaces in IP_{4}. Lecture Notes in Mathematics, 1417:91–110, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083335/.

duPlessis:1991:RSM

[dPW91] Andrew du Plessis and Leslie Charles Wilson. Right-symmetry of mappings. Lecture Notes in Mathematics, 1462:258–275, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086387/.

Dumortier:1990:SLB

[DR90] Freddy Dumortier and Robert Roussarie. On the saddle loop bifurcation. Lecture Notes in Mathematics, 1455:44–73, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085390/.

Dolgachev:1991:RVB

[DR91] Igor Dolgachev and Igor Reider. On rank 2 vector bundles with c_{1}^{2} = 10 and c_{2} = 3 on Enriques surfaces. Lecture Notes in Mathematics, 1479:39–49, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086262/.

Drinfeld:1992:SUP

[Dri92] V. G. Drinfel’d. On some unsolved problems in quantum group theory. Lecture Notes in Mathematics, 1510:1–8, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101175/.

Drmota:1990:IDH

[Drm90] Michael Drmota. On irregularities of distribution on the hyperbolic plane. Lecture Notes in Mathematics, 1452:31–42, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096979/.

Dumortier:1991:AIU

[DRSŻ91a] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Abelian integrals in unfoldings of codimension 3 singular planar vector fields. Lecture Notes in Mathematics, 1480: 165–224, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098361/.

Dumortier:1991:BC

[DRSŻ91b] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Bifurcations of codimension 1 and 2. Lecture Notes in Mathematics, 1480:28–56, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098357/.

Dumortier:1991:CR

[DRSŻ91c] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. The central rescaling. Lecture Notes in Mathematics, 1480:85–134, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098359/.

Dumortier:1991:CDR

[DRSŻ91d] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Conclusions and discussion of remaining problems. Lecture Notes in Mathematics, 1480:135–164, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098360/.

Dumortier:1991:DN

[DRSŻ91e] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Definitions and notations. Lecture Notes in Mathematics, 1480:19–21, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098355/.

Dumortier:1991:EP

[DRSŻ91f] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Elementary properties. Lecture Notes in Mathematics, 1480:57–84, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098358/.

Dumortier:1991:I

[DRSŻ91g] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Introduction. Lecture Notes in Mathematics, 1480: 1–18, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098354/.

Dumortier:1991:TNF

[DRSŻ91h] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żaładek. Transformation into normal form. Lecture Notes in Mathematics, 1480:22–27, 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098356/.

Dumortier:1991:BPV

[DRSŻ91i] Freddy Dumortier, Robert Roussarie, Jorge Sotomayor, and Henryk Żoładek. Bifurcations of Planar Vector Fields: Nilpotent Singularities and Abelian Integrals, volume 1480 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54521-2 (print), 3-540-38433-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 226 pp. LCCN QA3 .L28 no. 1480. URL http:/

/link.springer.com/book/10.1007/BFb0098353;   http://
www.springerlink.com/content/978-3-540-38433-5.

Dovermann:1990:EST

[DS90a] Karl Heinz Dovermann and Reinhard Schultz. Equivariant Surgery Theories and Their Periodicity Properties, volume 1443 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 227 pp. LCCN QA3 .L28 no. 1443. URL http:/

/link.springer.com/book/10.1007/BFb0092354;   http://
www.springerlink.com/content/978-3-540-46394-8.

Dovermann:1990:IES

[DS90b] Karl Heinz Dovermann and Reinhard Schultz. Introduction to equivariant surgery. Lecture Notes in Mathematics, 1443:9– 36, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092356/.

Dovermann:1990:PTE

[DS90c] Karl Heinz Dovermann and Reinhard Schultz. Periodicity theorems in equivariant surgery. Lecture Notes in Mathematics, 1443: 80–114, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092358/.

Dovermann:1990:PPS

[DS90d] Karl Heinz Dovermann and Reinhard Schultz. Products and periodicity for surgery up to pseudoequivalence. Lecture Notes in Mathematics, 1443:141–191, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092360/.

Dovermann:1990:RBE

[DS90e] Karl Heinz Dovermann and Reinhard Schultz. Relations between equivariant surgery theories. Lecture Notes in Mathematics, 1443: 37–79, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092357/.

Dovermann:1990:SBM

[DS90f] Karl Heinz Dovermann and Reinhard Schultz. Summary: Background material and basic results. Lecture Notes in Mathematics, 1443:1–8, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092355/.

Dovermann:1990:TPF

[DS90g] Karl Heinz Dovermann and Reinhard Schultz. Twisted product formulas for surgery with coefficients. Lecture Notes in Mathematics, 1443:115–140, 1990. CODEN LNMAA2. ISBN 3-540-53042-8 (print), 3-540-46394-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092359/.

Dubins:1992:MDL

[DS92] Lester E. Dubins and Meir Smorodinsky. The modified, discrete, Lévy-transformation is Bernoulli. Lecture Notes in Mathematics, 1526:157–161, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084318/.

Delbaen:1995:IPP

[DS95] F. Delbaen and W. Schachermayer. An inequality for the predictable projection of an adapted process. Lecture Notes in Mathematics, 1613:17–24, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094195/.

Drmota:1997:A

[DT97a] Michael Drmota and Robert F. Tichy. Applications. Lecture Notes in Mathematics, 1651:368–432, 1997. CODEN LNMAA2. ISBN 3-540-62606-9 (print), 3-540-68333-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093407/.

Drmota:1997:DS

[DT97b] Michael Drmota and Robert F. Tichy. Discrepancy of sequences. Lecture Notes in Mathematics, 1651:1–203, 1997. CODEN LN-MAA2. ISBN 3-540-62606-9 (print), 3-540-68333-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093405/.

Drmota:1997:GCU

[DT97c] Michael Drmota and Robert F. Tichy. General concepts of uniform distribution. Lecture Notes in Mathematics, 1651:204– 367, 1997. CODEN LNMAA2. ISBN 3-540-62606-9 (print), 3-540-68333-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093406/.

Drmota:1997:SDA

[DT97d] Michael Drmota and Robert F. Tichy. Sequences, Discrepancies and Applications, volume 1651 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62606-9 (print), 3-540-68333-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 503 pp. LCCN QA3 .L28 no. 1651.

URL http://link.springer.com/book/10.1007/BFb0093404;
 http://www.springerlink.com/content/978-3-540-68333-
9.

Drmota:1990:CUD

[DTW90] M. Drmota, R. F. Tichy, and R. Winkler. Completely uniformly distributed sequences of matrices. Lecture Notes in Mathematics, 1452:43–57, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096980/.

Denker:1992:DHM

[DU92] Manfred Denker and Mariusz Urbański. The dichotomy of Hausdorff measures and equilibrium states for parabolic rational maps. Lecture Notes in Mathematics, 1514:90–113, 1992. CODEN LN-MAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097530/.

Dubrovin:1996:GTF

[Dub96] Boris Dubrovin. Geometry of 2D topological field theories. Lecture Notes in Mathematics, 1620:120–348, 1996. CODEN LN-MAA2. ISBN 3-540-60542-8 (print), 3-540-47706-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094793/.

Duggal:1992:NGC

[Dug92] B. P. Duggal. A note on generalised commutativity theorems in the Schatten norm. Lecture Notes in Mathematics, 1511:47– 54, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093796/.

Dula:1991:CS

[Dul91] Giora Dula. On conic spaces. Lecture Notes in Mathematics, 1474: 38–58, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084736/.

Dumas:1991:SCF

[Dum91] François Dumas. Sous-corps de fractions rationnelles des corps gauches de séries de Laurent. (French) []. Lecture Notes in Mathematics, 1478:192–214, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083504/.

Dung:1992:MCA

[Dũn92] Nguyẽn Viet Dũng. The modulo 2 cohomology algebra of the wreath product Σ_{∞} ∫ X. Lecture Notes in Mathematics, 1509: 115–119, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087505/.

Durand:1990:QATa

[Dur90a] A. Durand. Quelques aspects de la théorie analytique des polynômes I. (French) []. Lecture Notes in Mathematics, 1415: 1–42, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084877/.

Durand:1990:QATb

[Dur90b] A. Durand. Quelques aspects de la théorie analytique des polynômes II. (French) []. Lecture Notes in Mathematics, 1415:

43–85, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084878/.

Durand:1990:AAN

[Dur90c] Alain Durand. Approximations algébriques d’un nombre transcendant. (French) []. Lecture Notes in Mathematics, 1415:94– 96, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084880/.

Durand:1990:RSD

[Dur90d] Alain Durand. Relation de Szegő sur la derivée d’un polynôme. (French) []. Lecture Notes in Mathematics, 1415:86–93, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084879/.

Duren:1994:GFT

[Dur94] P. Duren. Geometric function theory. Lecture Notes in Mathematics, 1574:383–422, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101068/.

Durrett:1995:TLP

[Dur95] Rick Durrett. Ten lectures on particle systems. Lecture Notes in Mathematics, 1608:97–201, 1995. CODEN LNMAA2. ISBN 3-540-60015-9 (print), 3-540-49402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095747/.

Durrett:1999:SSM

[Dur99] Rick Durrett. Stochastic spatial models. Lecture Notes in Mathematics, 1714:39–94, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092375/.

Dillen:1991:GCS

[DV91] Franki Dillen and Luc Vrancken. Generalized Cayley surfaces. Lecture Notes in Mathematics, 1481:36–47, 1991. CODEN LN-MAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083626/.

Diekmann:1991:CMD

[DvG91] Odo Diekmann and Stephan A. van Gils. The center manifold for delay equations in the light of suns and stars. Lecture Notes in Mathematics, 1463:122–141, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085429/.

Dangelmayr:1991:CFB

[DW91] G. Dangelmayr and M. Wegelin. On a codimension-four bifurcation occurring in optical bistability. Lecture Notes in Mathematics, 1463:107–121, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085428/.

Deaconu:1997:CTA

[DW97] Mădălina Deaconu and Sophie Wantz. Comportement des temps d’atteinte d’une diffusion fortement rentrante. (French) []. Lecture Notes in Mathematics, 1655:168–175, 1997. CODEN LN-MAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119301/.

Dwork:1990:WPR

[Dwo90] B. Dwork. Work of Philippe Robba. Lecture Notes in Mathematics, 1454:1–10, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091131/.

Doney:1998:PBP

[DWY98] R. A. Doney, J. Warren, and M. Yor. Perturbed Bessel processes. Lecture Notes in Mathematics, 1686:237–249, 1998. CO-DEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101761/.

Dimaki:1993:CPD

[DX93] Caterina Dimaki and Evdokia Xekalaki. Characterizations of the Pareto distribution based on order statistics. Lecture Notes in Mathematics, 1546:1–16, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084477/.

Dontchev:1993:GWP

[DZ93a] Asen L. Dontchev and Tullio Zolezzi. Generic well-posedness. Lecture Notes in Mathematics, 1543:81–115, 1993. CODEN LN-MAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084198/.

Dontchev:1993:HTW

[DZ93b] Asen L. Dontchev and Tullio Zolezzi. Hadamard and Tykhonov well-posedness. Lecture Notes in Mathematics, 1543:38–80, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084197/.

Dontchev:1993:HWP

[DZ93c] Asen L. Dontchev and Tullio Zolezzi. Hadamard well-posedness in mathematical programming. Lecture Notes in Mathematics, 1543:335–380, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084204/.

Dontchev:1993:RVH

[DZ93d] Asen L. Dontchev and Tullio Zolezzi. Relaxation and value Hadamard well-posedness in optimal control. Lecture Notes in Mathematics, 1543:230–247, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084201/.

Dontchev:1993:SPO

[DZ93e] Asen L. Dontchev and Tullio Zolezzi. Singular perturbations in optimal control. Lecture Notes in Mathematics, 1543:248– 282, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084202/.

Dontchev:1993:TWP

[DZ93f] Asen L. Dontchev and Tullio Zolezzi. Tykhonov well-posedness. Lecture Notes in Mathematics, 1543:1–37, 1993. CODEN LN-MAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084196/.

Dontchev:1993:WPOb

[DZ93g] Asen L. Dontchev and Tullio Zolezzi. Well-Posed Optimization Problems, volume 1543 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 421 pp. LCCN QA3 .L28 no. 1543.

URL http://link.springer.com/book/10.1007/BFb0084195;
 http://www.springerlink.com/content/978-3-540-47644-
3.

Dontchev:1993:WPV

[DZ93h] Asen L. Dontchev and Tullio Zolezzi. Well-posedness and variational, epi- and Mosco convergences. Lecture Notes in Mathematics, 1543:116–175, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084199/.

Dontchev:1993:WPOa

[DZ93i] Asen L. Dontchev and Tullio Zolezzi. Well-posedness in optimal control. Lecture Notes in Mathematics, 1543:176–229, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084200/.

Dontchev:1993:WPC

[DZ93j] Asen L. Dontchev and Tullio Zolezzi. Well-posedness in the calculus of variations. Lecture Notes in Mathematics, 1543:283– 334, 1993. CODEN LNMAA2. ISBN 3-540-56737-2 (print), 3-540-47644-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084203/.

Ebenfelt:1998:HMB

[Ebe98] Peter Ebenfelt. Holomorphic mappings between real analytic submanifolds in complex space. Lecture Notes in Mathematics, 1684: 35–69, 1998. CODEN LNMAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096090/.

Eberle:1999:UFD

[Ebe99a] Andreas Eberle. L^{p} uniqueness in finite dimensions. Lecture Notes in Mathematics, 1718:41–87, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103048/.

Eberle:1999:FSI

[Ebe99b] Andreas Eberle. First steps in infinite dimensions. Lecture Notes in Mathematics, 1718:185–253, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103051/.

Eberle:1999:I

[Ebe99c] Andreas Eberle. Introduction. Lecture Notes in Mathematics, 1718:1–8, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103046/.

Eberle:1999:MU

[Ebe99d] Andreas Eberle. Markov uniqueness. Lecture Notes in Mathematics, 1718:89–167, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103049/.

Eberle:1999:MBD

[Ebe99e] Andreas Eberle. Motivation and basic definitions: Uniqueness problems in various contexts. Lecture Notes in Mathematics, 1718: 9–40, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103047/.

Eberle:1999:PAM

[Ebe99f] Andreas Eberle. Probabilistic aspects of L^{p} and Markov uniqueness. Lecture Notes in Mathematics, 1718:169–184, 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103050/.

Eberle:1999:UNU

[Ebe99g] Andreas Eberle. Uniqueness and Non-Uniqueness of Semigroups Generated by Singular Diffusion Operators, volume 1718 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66628-1 (print), 3-540-48076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 262 pp. LCCN QA3 .L28 no. 1718. URL http:/

/link.springer.com/book/10.1007/BFb0103045;   http://
www.springerlink.com/content/978-3-540-48076-1.

Ecalle:1990:FCL

[Eca90] J. Ecalle. Finitude des cycles-limites et accéléro-sommation de l’application de retour. (French) []. Lecture Notes in Mathematics, 1455:74–159, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085391/.

Echerbault:1996:MHF

[Ech96] Mireille Echerbault. Sur le modèle d’Heisenberg. (French) [on the Heisenberg model]. Lecture Notes in Mathematics, 1626:162– 177, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094647/.

Edixhoven:1993:APF

[Edi93] Bas Edixhoven. Arithmetic part of Faltings’s proof. Lecture Notes in Mathematics, 1566:97–110, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48208-6_11/.

Edixhoven:1993:DAA

[EE93] Bas Edixhoven and Jan-Hendrik Evertse. Diophantine Approximation and Abelian Varieties, volume 1566 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 127 pp. LCCN QA3 .L28 no.

1566. URL http://link.springer.com/book/10.1007/978-
3-540-48208-6;   http://www.springerlink.com/content/
978-3-540-48208-6.

Eggermont:1991:VLD

[Egg91] P. P. B. Eggermont. On Volterra–Lotka differential equations and multiplicative algorithms for monotone complementarity problems. Lecture Notes in Mathematics, 1497:141–152, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084515/.

Esposti:1994:EPO

[EGI94] Mirko Degli Esposti, Sandro Graffi, and Stefano Isola. Equidistribution of periodic orbits: an overview of classical vs quantum results. Lecture Notes in Mathematics, 1589:65–91, 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074075/.

Ehrhard:1992:DAG

[Ehr92] Dirk Ehrhard. Decoding algebraic-geometric codes by solving a key equation. Lecture Notes in Mathematics, 1518:18–25, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087989/.

Eisenbaum:1997:TLP

[Eis97] Nathalie Eisenbaum. Théorèmes limités pour les temps locaux d’un processus stable symétrique. (French) []. Lecture Notes in Mathematics, 1655:216–224, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0119306/.

Ekeland:1990:VPR

[Eke90] I. Ekeland. The ϵ-variational principle revisited. Lecture Notes in Mathematics, 1446:1–15, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084929/.

Elfving:1991:CRL

[Elf91] Tommy Elfving. Constrained regularized least squares problems. Lecture Notes in Mathematics, 1497:153–166, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084516/.

Elworthy:1999:A

[ELL99a] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Appendices. Lecture Notes in Mathematics, 1720:95–110, 1999. CODEN LN-MAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103071/.

Elworthy:1999:AAS

[ELL99b] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Application: Analysis on spaces of paths. Lecture Notes in Mathematics, 1720: 76–86, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103069/.

Elworthy:1999:CC

[ELL99c] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Construction of connections. Lecture Notes in Mathematics, 1720:7–29, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103066/.

Elworthy:1999:DNF

[ELL99d] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Decomposition of noise and filtering. Lecture Notes in Mathematics, 1720:57– 75, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103068/.

Elworthy:1999:IGA

[ELL99e] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. The infinitesimal generators and associated operators. Lecture Notes in Mathematics, 1720:30–56, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103067/.

Elworthy:1999:I

[ELL99f] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Introduction. Lecture Notes in Mathematics, 1720:3–6, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103065/.

Elworthy:1999:SSD

[ELL99g] K. D. Elworthy, J. Le Jan, and Xue-Mei Li. Stability of stochastic dynamical systems. Lecture Notes in Mathematics, 1720:87– 94, 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103070/.

Elworthy:1999:GDO

[ELL99h] K. David Elworthy, Yves Le Jan, and Xue-Mei Li. On the Geometry of Diffusion Operators and Stochastic Flows, volume 1720 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66708-3 (print), 3-540-47022-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 118 pp. LCCN QA3 .L28 no. 1720. URL http:

//link.springer.com/book/10.1007/BFb0103064;   http:/
/www.springerlink.com/content/978-3-540-47022-9.

Ewing:1990:PIS

[ELLV90] Richard E. Ewing, Raytcho D. Lazarov, Peng Lu, and Panayot S. Vassilevski. Preconditioning indefinite systems arising from mixed finite element discretization of second-order elliptic problems. Lecture Notes in Mathematics, 1457:28–43, 1990. CODEN LN-MAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090900/.

Elworthy:1997:TSQ

[ELY97] K. D. Elworthy, X. M. Li, and M. Yor. On the tails of the supremum and the quadratic variation of strictly local martingales. Lecture Notes in Mathematics, 1655:113–125, 1997. CO-DEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0119298/.

Emery:1991:BPV

[EM91] M. Emery and G. Mokobodzki. Sur le barycentre d’une probabilité dans une variété. (French) []. Lecture Notes in Mathematics, 1485:220–233, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100858/.

Emery:1991:QCR

[Eme91] M. Emery. Quelques cas de représentation chaotique. (French) []. Lecture Notes in Mathematics, 1485:10–23, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100841/.

Emery:1997:CSS

[Éme97] M. Émery. Closed sets supporting a continuous divergent martingale. Lecture Notes in Mathematics, 1655:176–189, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119302/.

Ekeland:1990:FM

[EMM^{+}90] Ivar Ekeland, Paolo Marcellini, Antonio Marino, Mario Tosques, Czesław Olech, Giulio Pianigiani, Tyrrell Rockafeller, and Michel Valadier. Front matter. Lecture Notes in Mathematics, 1446: ??, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46715-1/1.pdf.

Ene:1995:CF

[Ene95a] Vasile Ene. Classes of functions. Lecture Notes in Mathematics, 1603:25–125, 1995. CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095390/.

Ene:1995:E

[Ene95b] Vasile Ene. Examples. Lecture Notes in Mathematics, 1603:213– 292, 1995. CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095394/.

Ene:1995:FRC

[Ene95c] Vasile Ene. Finite representations for continuous functions. Lecture Notes in Mathematics, 1603:127–139, 1995. CODEN LN-MAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095391/.

Ene:1995:I

[Ene95d] Vasile Ene. Integrals. Lecture Notes in Mathematics, 1603:161– 211, 1995. CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095393/.

Ene:1995:M

[Ene95e] Vasile Ene. Monotonicity. Lecture Notes in Mathematics, 1603: 141–159, 1995. CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095392/.

Ene:1995:P

[Ene95f] Vasile Ene. Preliminaries. Lecture Notes in Mathematics, 1603: 1–23, 1995. CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095389/.

Ene:1995:RFC

[Ene95g] Vasile Ene. Real Functions — Current Topics, volume 1603 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995.

CODEN LNMAA2. ISBN 3-540-60008-6 (print), 3-540-49400-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 310 pp. LCCN QA3 .L28 no. 1603. URL http:/

/link.springer.com/book/10.1007/BFb0095388;   http://
www.springerlink.com/content/978-3-540-49400-3.

Ellia:1990:GPC

[EP90] Ph. Ellia and Ch. Peskine. Groupes de points de P^{2}: Caractère et position uniforme. (French) []. Lecture Notes in Mathematics, 1417:111–116, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083336/.

ElKaroui:1997:NLP

[EQ97] N. El Karoui and M. C. Quenez. Non-linear pricing theory and backward stochastic differential equations. Lecture Notes in Mathematics, 1656:191–246, 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092001/.

Erdmann:1990:ADT

[Erd90a] Karin Erdmann. Algebras of dihedral type. Lecture Notes in Mathematics, 1428:159–180, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084009/.

Erdmann:1990:AQT

[Erd90b] Karin Erdmann. Algebras of quaternion type. Lecture Notes in Mathematics, 1428:181–204, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084010/.

Erdmann:1990:AST

[Erd90c] Karin Erdmann. Algebras of semidihedral type. Lecture Notes in Mathematics, 1428:205–262, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084011/.

Erdmann:1990:AQR

[Erd90d] Karin Erdmann. Algebras, quivers, representation type, Auslander–Reiten theory, coverings. Lecture Notes in Mathematics, 1428:1–46, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084004/.

Erdmann:1990:BTR

[Erd90e] Karin Erdmann. Blocks of Tame Representation Type and Related Algebras, volume 1428 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 312 pp. LCCN QA3 .L28 no. 1428.

URL http://link.springer.com/book/10.1007/BFb0084003;
 http://www.springerlink.com/content/978-3-540-47115-
8.

Erdmann:1990:CBD

[Erd90f] Karin Erdmann. Centres, blocks, decomposition numbers. Lecture Notes in Mathematics, 1428:263–285, 1990. CODEN LN-MAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084012/.

Erdmann:1990:MMQ

[Erd90g] Karin Erdmann. More on modules, quivers, Auslander–Reiten sequences. Lecture Notes in Mathematics, 1428:93–120, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084007/.

Erdmann:1990:SA

[Erd90h] Karin Erdmann. Some applications. Lecture Notes in Mathematics, 1428:286–306, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084013/.

Erdmann:1990:SBA

[Erd90i] Karin Erdmann. Special biserial algebras and the local semidihedral algebra. Lecture Notes in Mathematics, 1428:47–79, 1990.

CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084005/.

Erdmann:1990:SAR

[Erd90j] Karin Erdmann. Stable Auslander–Reiten components for tame blocks. Lecture Notes in Mathematics, 1428:121–158, 1990. CO-DEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084008/.

Erdmann:1990:TSL

[Erd90k] Karin Erdmann. Tame symmetric local algebras. Lecture Notes in Mathematics, 1428:80–92, 1990. CODEN LNMAA2. ISBN 3-540-52709-5 (print), 3-540-47115-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084006/.

Erneux:1991:SPT

[ERHG91] T. Erneux, E. L. Reiss, L. J. Holden, and M. Georgiou. Slow passage through bifurcation and limit points. Asymptotic theory and applications. Lecture Notes in Mathematics, 1493:14–28, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085021/.

Emery:1999:BFS

[ÉS99a] M. Émery and W. Schachermayer. Brownian filtrations are not stable under equivalent time-changes. Lecture Notes in Mathematics, 1709:267–276, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096516/.

Emery:1999:RTS

[ÉS99b] M. Émery and W. Schachermayer. A remark on Tsirelson’s stochastic differential equation. Lecture Notes in Mathematics, 1709:291–303, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096518/.

Es-Sahib:1999:BCP

[ESH99] Aziz Es-Sahib and Heinich Henri. Barycentre canonique pour un espace métrique à courbure négative. (French) []. Lecture Notes in Mathematics, 1709:355–370, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096526/.

Essen:1996:PTP

[Ess96] Matts Essén. Potential theory part I. Lecture Notes in Mathematics, 1633:3–100, 1996. CODEN LNMAA2. ISBN 3-540-61583-0 (print), 3-540-69991-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093411/.

Evans:1992:ATN

[Eva92a] Anthony B. Evans. Automorphisms and translation nets. Lecture Notes in Mathematics, 1535:50–56, 1992. CODEN LN-MAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092367/.

Evans:1992:CO

[Eva92b] Anthony B. Evans. Cyclotomic orthomorphisms. Lecture Notes in Mathematics, 1535:35–49, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092366/.

Evans:1992:DSG

[Eva92c] Anthony B. Evans. Data for small groups. Lecture Notes in Mathematics, 1535:77–90, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092369/.

Evans:1992:EAG

[Eva92d] Anthony B. Evans. Elementary abelian groups. Lecture Notes in Mathematics, 1535:25–34, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092365/.

Evans:1992:FR

[Eva92e] Anthony B. Evans. Further results. Lecture Notes in Mathematics, 1535:57–76, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092368/.

Evans:1992:I

[Eva92f] Anthony B. Evans. Introduction. Lecture Notes in Mathematics, 1535:1–24, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092364/.

Evans:1992:OGG

[Eva92g] Anthony B. Evans. Orthomorphism Graphs of Groups, volume 1535 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 114 pp. LCCN QA3 .L28 no. 1535. URL http:/

/link.springer.com/book/10.1007/BFb0092363;   http://
www.springerlink.com/content/978-3-540-47541-5.

Evans:1992:RD

[Eva92h] Anthony B. Evans. Research directions. Lecture Notes in Mathematics, 1535:91–104, 1992. CODEN LNMAA2. ISBN 3-540-56351-2 (print), 3-540-47541-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092370/.

Evans:1997:RFN

[Eva97] Lawrence C. Evans. Regularity for fully nonlinear elliptic equations and motion by mean curvature. Lecture Notes in Mathematics, 1660:98–133, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094296/.

Evertse:1993:STW

[Eve93] Jan-Hendrik Evertse. The subspace theorem of W. M. Schmidt. Lecture Notes in Mathematics, 1566:31–50, 1993. CODEN LN-MAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48208-6_
4/.

Elworthy:1993:CED

[EY93] K. D. Elworthy and M. Yor. Conditional expectations for derivatives of certain stochastic flows. Lecture Notes in Mathematics, 1557:159–172, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087972/.

Emery:1998:TTR

[ÉY98] M. Émery and M. Yor. Sur un théorème de Tsirelson relatif à des mouvements browniens corrélés et à la nullité de certains temps locaux. (French) []. Lecture Notes in Mathematics, 1686: 306–312, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101764/.

Eyre:1998:GS

[Eyr98a] Timothy M. W. Eyre. Z_{2}-graded structures. Lecture Notes in Mathematics, 1692:23–31, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096853/.

Eyre:1998:CE

[Eyr98b] Timothy M. W. Eyre. Chaotic expansions. Lecture Notes in Mathematics, 1692:101–112, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096858/.

Eyre:1998:E

[Eyr98c] Timothy M. W. Eyre. Extensions. Lecture Notes in Mathematics, 1692:113–132, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096859/.

Eyre:1998:I

[Eyr98d] Timothy M. W. Eyre. Introduction. Lecture Notes in Mathematics, 1692:1–6, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096851/.

Eyre:1998:IS

[Eyr98e] Timothy M. W. Eyre. The Itô superalgebra. Lecture Notes in Mathematics, 1692:59–75, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096856/.

Eyre:1998:QSCa

[Eyr98f] Timothy M. W. Eyre. Quantum stochastic calculus. Lecture Notes in Mathematics, 1692:7–21, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096852/.

Eyre:1998:QSCb

[Eyr98g] Timothy M. W. Eyre. Quantum Stochastic Calculus and Representations of Lie Superalgebras, volume 1692 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CO-DEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 138 pp. LCCN QA3 .L28 no. 1692. URL http:/

/link.springer.com/book/10.1007/BFb0096850;   http://
www.springerlink.com/content/978-3-540-68385-8.

Eyre:1998:RLS

[Eyr98h] Timothy M. W. Eyre. Representations of Lie superalgebras in Z_{2}- graded quantum stochastic calculus. Lecture Notes in Mathematics, 1692:33–50, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096854/.

Eyre:1998:SRG

[Eyr98i] Timothy M. W. Eyre. Some results in Z_{2}-graded quantum stochastic calculus. Lecture Notes in Mathematics, 1692:77– 99, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096857/.

Eyre:1998:UHO

[Eyr98j] Timothy M. W. Eyre. The ungraded higher order Itô product formula. Lecture Notes in Mathematics, 1692:51–57, 1998. CODEN LNMAA2. ISBN 3-540-64897-6 (print), 3-540-68385-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096855/.

Faber:1993:GPF

[Fab93a] Carel Faber. Geometric part of Faltings’s proof. Lecture Notes in Mathematics, 1566:83–91, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48208-6_9/.

Fabes:1993:GUB

[Fab93b] E. B. Fabes. Gaussian upper bounds on fundamental solutions of parabolic equations; the method of Nash. Lecture Notes in Mathematics, 1563:1–20, 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074089/.

Fagnola:1990:QSC

[Fag90] F. Fagnola. Quantum stochastic calculus and a boson Lévy theorem. Lecture Notes in Mathematics, 1442:131–144, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085506/.

Faltings:1992:RPD

[Fal92] Gerd Faltings. Recent progress in Diophantine geometry. Lecture Notes in Mathematics, 1525:78–86, 1992. CODEN LN-MAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089207/.

Farjoun:1992:HLP

[Far92] Emmanuel Dror Farjoun. Homotopy localization and V_{1}-periodic spaces. Lecture Notes in Mathematics, 1509:104–113, 1992. CO-DEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087504/.

Farjoun:1996:PST

[Far96a] Emmanuel Dror Farjoun. v_{1}-periodic spaces and K-theory. Lecture Notes in Mathematics, 1622:144–154, 1996. CODEN LN-MAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094437/.

Farjoun:1996:AHI

[Far96b] Emmanuel Dror Farjoun. Augmented homotopy idempotent functors. Lecture Notes in Mathematics, 1622:39–58, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094431/.

Farjoun:1996:CI

[Far96c] Emmanuel Dror Farjoun. Cellular inequalities. Lecture Notes in Mathematics, 1622:155–175, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094438/.

Farjoun:1996:CSN

[Far96d] Emmanuel Dror Farjoun. Cellular Spaces, Null Spaces and Homotopy Localization, volume 1622 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 199 pp. LCCN QA3 .L28 no. 1622.

URL http://link.springer.com/book/10.1007/BFb0094429;
 http://www.springerlink.com/content/978-3-540-48449-
3.

Farjoun:1996:CNC

[Far96e] Emmanuel Dror Farjoun. Classification of nullity and cellular types of finite p-torsion suspension spaces. Lecture Notes in Mathematics, 1622:135–143, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094436/.

Farjoun:1996:CHI

[Far96f] Emmanuel Dror Farjoun. Coaugmented homotopy idempotent localization functors. Lecture Notes in Mathematics, 1622:1– 38, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094430/.

Farjoun:1996:CRP

[Far96g] Emmanuel Dror Farjoun. Commutation rules for Ω, L_{f} and CW_{A}, preservation of fibrations and cofibrations. Lecture Notes in Mathematics, 1622:59–78, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094432/.

Farjoun:1996:DTS

[Far96h] Emmanuel Dror Farjoun. Dold–Thom symmetric products and other colimits. Lecture Notes in Mathematics, 1622:79–99, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094433/.

Farjoun:1996:GTF

[Far96i] Emmanuel Dror Farjoun. General theory of fibrations, GEM error terms. Lecture Notes in Mathematics, 1622:100–126, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094434/.

Farjoun:1996:HLN

[Far96j] Emmanuel Dror Farjoun. Homological localization nearly preserves fibrations. Lecture Notes in Mathematics, 1622:127–134, 1996. CODEN LNMAA2. ISBN 3-540-60604-1 (print), 3-540-48449-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094435/.

Fernique:1991:CLF

[Fer91] Xavier Fernique. Convergence en loi de fonctions aléatoires continues ou càdlàg, propriétés de compacité des lois. (French) []. Lecture Notes in Mathematics, 1485:178–195, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100856/.

Feigin:1996:IMQ

[FF96] Boris Feigin and Edward Frenkel. Integrals of motion and quantum groups. Lecture Notes in Mathematics, 1620:349–418, 1996. CODEN LNMAA2. ISBN 3-540-60542-8 (print), 3-540-47706-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094794/.

Fang:1997:DIB

[FF97] Shizan Fang and Jacques Franchi. A differentiable isomorphism between Wiener space and path group. Lecture Notes in Mathematics, 1655:54–61, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0119291/.

Fabes:1993:FM

[FFG^{+}93] Eugene Fabes, Masatoshi Fukushima, Leonard Gross, Carlos Kenig, Michael Röckner, and Daniel W. Stroock. Front matter. Lecture Notes in Mathematics, 1563:??, 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-48151-5/1.pdf.

Francaviglia:1990:GGM

[FG90] Mauro Francaviglia and Francesco Gherardelli, editors. Global Geometry and Mathematical Physics: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Montecatini Terme, Italy, July 4–12, 1988, volume 1451 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1451. URL http:/

/link.springer.com/book/10.1007/BFb0085063;   http://
www.springerlink.com/content/978-3-540-46813-4.

Francaviglia:1996:ISQ

[FG96] Mauro Francaviglia and Silvio Greco, editors. Integrable Systems and Quantum Groups: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, June 14–22, 1993, volume 1620 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-60542-8 (print), 3-540-47706-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1620. URL http:/

/link.springer.com/book/10.1007/BFb0094791;   http://
www.springerlink.com/content/978-3-540-47706-8.

Ferrandez:1991:CCC

[FGL91] A. Ferrández, O. J. Garay, and P. Lucas. On a certain class of conformally flat Euclidean hypersurfaces. Lecture Notes in Mathematics, 1481:48–54, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083627/.

Frame:1990:FDL

[FH90] Michael Frame and James Hefferon. Fractal dimensions of limit sets of some Kleinian groups. Lecture Notes in Mathematics, 1440: 74–80, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084452/.

Field:1991:LSE

[Fie91] Mike Field. Local structure of equivariant dynamics. Lecture Notes in Mathematics, 1463:142–166, 1991. CODEN LN-MAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085430/.

Fujita:1990:FAM

[FIK90] Hiroshi Fujita, Teruo Ikebe, and Shige Toshi Kuroda, editors. Functional-Analytic Methods for Partial Differential Equations: Proceedings of a Conference and a Symposium held in Tokyo, Japan, July 3–9, 1989, volume 1450 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1450. URL http:/

/link.springer.com/book/10.1007/BFb0084893;   http://
www.springerlink.com/content/978-3-540-46818-9.

Fine:1990:SPC

[Fin90] Benjamin Fine. Subgroup presentations without coset representatives. Lecture Notes in Mathematics, 1440:59–73, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084451/.

Fischer:1990:GAL

[Fis90] T. M. Fischer. A Galerkin approximation for linear eigenvalue problems in two and three-dimensional boundary-layer flows. Lecture Notes in Mathematics, 1431:100–108, 1990. CODEN LN-MAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086060/.

Fitzpatrick:1993:PID

[Fit93] Patrick Fitzpatrick. The parity as an invariant for detecting bifurcation of the zeroes of one parameter families of nonlinear Fredholm maps. Lecture Notes in Mathematics, 1537:1–31, 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085074/.

Farrell:1991:CSP

[FJ91] Professor F. T. Farrell and Professor L. E. Jones. Computations of stable pseudoisotopy spaces for aspherical manifolds. Lecture Notes in Mathematics, 1474:59–74, 1991. CODEN LN-MAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084737/.

Frohlich:1993:DTL

[FK93a] Jürg Fröhlich and Thomas Kerler. Duality theory for local quantum theories, dimensions and balancing in quantum categories. Lecture Notes in Mathematics, 1542:176–283, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084251/.

Frohlich:1993:HAQ

[FK93b] Jürg Fröhlich and Thomas Kerler. Hopf algebras and quantum groups at roots of unity. Lecture Notes in Mathematics, 1542: 102–118, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084248/.

Frohlich:1993:ISR

[FK93c] Jürg Fröhlich and Thomas Kerler. Introduction and survey of results. Lecture Notes in Mathematics, 1542:1–16, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084245/.

Frohlich:1993:LQT

[FK93d] Jürg Fröhlich and Thomas Kerler. Local quantum theory with braid group statistics. Lecture Notes in Mathematics, 1542:17– 44, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084246/.

Frohlich:1993:PRB

[FK93e] Jürg Fröhlich and Thomas Kerler. Path representations of the braid groups for quantum groups at roots of unity. Lecture Notes in Mathematics, 1542:141–175, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084250/.

Frohlich:1993:QCG

[FK93f] Jürg Fröhlich and Thomas Kerler. The quantum categories with a generator of dimension less than two. Lecture Notes in Mathematics, 1542:284–411, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084252/.

Frohlich:1993:QGQ

[FK93g] Jürg Fröhlich and Thomas Kerler. Quantum Groups, Quantum Categories and Quantum Field Theory, volume 1542 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 431 pp. LCCN QA3 .L28 no. 1542. URL http:/

/link.springer.com/book/10.1007/BFb0084244;   http://
www.springerlink.com/content/978-3-540-47611-5.

Frohlich:1993:RT

[FK93h] Jürg Fröhlich and Thomas Kerler. Representation theory of U_{q}^{red}(sℓ_{2}). Lecture Notes in Mathematics, 1542:119–140, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084249/.

Frohlich:1993:SSS

[FK93i] Jürg Fröhlich and Thomas Kerler. Superselection sectors and the structure of fusion rule algebras. Lecture Notes in Mathematics, 1542:45–101, 1993. CODEN LNMAA2. ISBN 3-540-56623-6 (print), 3-540-47611-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084247/.

Flandrin:1990:QMT

[Fla90] Patrick Flandrin. Quelques méthodes temps-fréquence et tempséchelle en traitement du signal. (French) []. Lecture Notes in Mathematics, 1438:81–92, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083516/.

Flandoli:1991:SFL

[Fla91] Franco Flandoli. Stochastic flow and Lyapunov exponents for abstract stochastic PDEs of parabolic type. Lecture Notes in Mathematics, 1486:196–205, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086669/.

Froment:1990:AMV

[FM90] Jacques Froment and Jean-Michel Morel. Analyse multiéchelle, vision stéréo et ondelettes. (French) []. Lecture Notes in Mathematics, 1438:51–80, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083515/.

Fitzpatrick:1993:FM

[FMMN93] Patrick Fitzpatrick, Mario Martelli, Jean Mawhin, and Roger Nussbaum. Front matter. Lecture Notes in Mathematics, 1537: ??, 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-47563-7/1.pdf.

Fitzgibbon:1991:EST

[FMSW91] W. E. Fitzgibbon, J. J. Morgan, R. S. Sanders, and S. J. Waggoner. Estimates for spatio-temporally dependent reaction diffusion systems. Lecture Notes in Mathematics, 1475:130–146, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083486/.

Foias:1991:CLT

[Foi91] Ciprian Foiaş. Commutant lifting techniques for computing optimal H^{∞} controllers. Lecture Notes in Mathematics, 1496:1– 36, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084466/.

Fontenas:1998:MCS

[Fon98] Éric Fontenas. Sur les minorations des constantes de Sobolev et de Sobolev logarithmiques pour les opérateurs de Jacobi et de Laguerre. (French) []. Lecture Notes in Mathematics, 1686: 14–29, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101747/.

Forte:1992:IPM

[For92] Bruno Forte. Inverse problems in mathematics for industry. Lecture Notes in Mathematics, 1521:64–110, 1992. CODEN LN-MAA2. ISBN 3-540-55595-1 (print), 3-540-47247-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098366/.

Fouvry:1990:NPP

[Fou90] Etienne Fouvry. Nombres presque premiers dans les petits intervalles. (French) []. Lecture Notes in Mathematics, 1434:65– 85, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097125/.

Fourati:1995:PMP

[Fou95] S. Fourati. Une propriété de Markov pour les processus indexés par R. (French) []. Lecture Notes in Mathematics, 1613:133– 154, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094206/.

Furi:1991:FSP

[FP91] Massimo Furi and Maria Patrizia Pera. The forced spherical pendulum does have forced oscillations. Lecture Notes in Mathematics, 1475:176–182, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083489/.

Fulton:1998:PFO

[FP98a] William Fulton and Piotr Pragacz. P̃− and Q̃− polynomial formulas for other classical groups. Lecture Notes in Mathematics, 1689:79–91, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096387/.

Fulton:1998:AOP

[FP98b] William Fulton and Piotr Pragacz. Applications and open problems. Lecture Notes in Mathematics, 1689:97–103, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096389/.

Fulton:1998:CBN

[FP98c] William Fulton and Piotr Pragacz. The classes of Brill–Noether loci in Prym varieties. Lecture Notes in Mathematics, 1689:92– 96, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096388/.

Fulton:1998:ECD

[FP98d] William Fulton and Piotr Pragacz. The Euler characteristic of degeneracy loci. Lecture Notes in Mathematics, 1689:53–64, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096385/.

Fulton:1998:FBD

[FP98e] William Fulton and Piotr Pragacz. Flag bundles and determinantal formulas for the other classical groups. Lecture Notes in Mathematics, 1689:65–78, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096386/.

Fulton:1998:IDL

[FP98f] William Fulton and Piotr Pragacz. Introduction to degeneracy loci and Schubert polynomials. Lecture Notes in Mathematics, 1689:1–13, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096381/.

Fulton:1998:MFG

[FP98g] William Fulton and Piotr Pragacz. Modern formulation; grassmannians, flag varieties, Schubert varieties. Lecture Notes in Mathematics, 1689:14–25, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096382/.

Fulton:1998:PSD

[FP98h] William Fulton and Piotr Pragacz. Polynomials supported on degeneracy loci. Lecture Notes in Mathematics, 1689:40–52, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096384/.

Fulton:1998:SVD

[FP98i] William Fulton and Piotr Pragacz. Schubert Varieties and Degeneracy Loci, volume 1689 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 148 pp. LCCN QA3 .L28 no. 1689.

URL http://link.springer.com/book/10.1007/BFb0096380;
 http://www.springerlink.com/content/978-3-540-69804-
3.

Fulton:1998:SPU

[FP98j] William Fulton and Piotr Pragacz. Symmetric polynomials useful in geometry. Lecture Notes in Mathematics, 1689:26–39, 1998. CODEN LNMAA2. ISBN 3-540-64538-1 (print), 3-540-69804-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096383/.

Frey:1992:DAE

[FPS92] Gerhard Frey, Marc Perret, and Henning Stichtenoth. On the different of abelian extensions of global fields. Lecture Notes in Mathematics, 1518:26–32, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087990/.

Ferus:1991:GDG

[FPSW91] Dirk Ferus, Ulrich Pinkall, Udo Simon, and Berd Wegner, editors. Global Differential Geometry and Global Analysis: Proceedings of a Conference held in Berlin, 15–20 June, 1990, volume 1481 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1481. URL http:/

/link.springer.com/book/10.1007/BFb0083621;   http://
www.springerlink.com/content/978-3-540-46445-7.

Falqui:1990:BM

[FR90a] Gregorio Falqui and Cesare Reina. Back matter. Lecture Notes in Mathematics, 1451:??, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-46813-4/1.pdf.

Falqui:1990:SS

[FR90b] Gregorio Falqui and Cesare Reina. Supermoduli and superstrings. Lecture Notes in Mathematics, 1451:176–188, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085070/.

Francoise:1990:BPV

[FR90c] Jean-Pierre Françoise and Robert Roussarie, editors. Bifurcations of Planar Vector Fields: Proceedings of a Meeting held in Luminy, France, Sept. 18–22, 1989, volume 1455 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1455.

URL http://link.springer.com/book/10.1007/BFb0085387;
 http://www.springerlink.com/content/978-3-540-46722-
9.

Francoise:1992:DRA

[FR92] J.-P. Françoise and F. Ronga. The decidability of real algebraic sets by the index formula. Lecture Notes in Mathematics, 1524: 235–239, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084623/.

Francoise:1990:IPG

[Fra90] Jean-Pierre Françoise. Intégrales de périodes en geométries symplectique et isochore. (French) []. Lecture Notes in Mathematics, 1416:105–138, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097467/.

Francis:1991:LCS

[Fra91] Bruce Francis. Lectures on H_{∞} control and sampled-data systems. Lecture Notes in Mathematics, 1496:37–105, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084467/.

Franchi:1995:CMT

[Fra95] J. Franchi. Chaos multiplicatif: un traitement simple et complet de la fonction de partition. (French) []. Lecture Notes in Mathematics, 1613:194–201, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094211/.

Freund:1990:BTI

[Fre90] Roland Freund. On Bernstein type inequalities and a weighted Chebyshev approximation problem on ellipses. Lecture Notes in Mathematics, 1435:45–55, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087896/.

Freitag:1991:FL

[Fre91a] Eberhard Freitag. The fundamental lemma. Lecture Notes in Mathematics, 1487:111–152, 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093835/.

Freitag:1991:I

[Fre91b] Eberhard Freitag. Introduction. Lecture Notes in Mathematics, 1487:1–5, 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093830/.

Freitag:1991:R

[Fre91c] Eberhard Freitag. The results. Lecture Notes in Mathematics, 1487:153–169, 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093836/.

Freitag:1991:SMFa

[Fre91d] Eberhard Freitag. Siegel modular forms. Lecture Notes in Mathematics, 1487:8–37, 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093831/.

Freitag:1991:SMFc

[Fre91e] Eberhard Freitag. Singular Modular Forms and Theta Relations, volume 1487 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 172 pp. LCCN QA3 .L28 no. 1487. URL http:/

/link.springer.com/book/10.1007/BFb0093829;   http://
www.springerlink.com/content/978-3-540-46434-1.

Freitag:1991:SMFb

[Fre91f] Eberhard Freitag. Singular modular forms and theta series. Lecture Notes in Mathematics, 1487:89–110, 1991. CODEN LN-MAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093834/.

Freitag:1991:SW

[Fre91g] Eberhard Freitag. Singular weights. Lecture Notes in Mathematics, 1487:70–88, 1991. CODEN LNMAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093833/.

Freitag:1991:TSP

[Fre91h] Eberhard Freitag. Theta series with polynomial coefficients. Lecture Notes in Mathematics, 1487:38–69, 1991. CODEN LN-MAA2. ISBN 3-540-54704-5 (print), 3-540-46434-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093832/.

Freyd:1991:ACC

[Fre91i] Peter Freyd. Algebraically complete categories. Lecture Notes in Mathematics, 1488:95–104, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084215/.

Freidlin:1992:SLP

[Fre92] Mark I. Freidlin. Semi-linear pde’s and limit theorems for large deviations. Lecture Notes in Mathematics, 1527:1–109, 1992. CO-DEN LNMAA2. ISBN 3-540-56250-8 (print), 3-540-47490-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084699/.

Frey:1994:ACO

[Fre94] Gerhard Frey. On Artin’s Conjecture for Odd 2-dimensional Representations, volume 1585 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 121 (est.) pp. LCCN QA3 .L28 no. 1585.

URL http://link.springer.com/book/10.1007/BFb0074106;
 http://www.springerlink.com/content/978-3-540-48681-
7.

Frigerio:1990:QPP

[Fri90a] Alberto Frigerio. Quantum Poisson processes: Physical motivations and applications. Lecture Notes in Mathematics, 1442: 177, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085509.

Frigerio:1990:TIN

[Fri90b] Alberto Frigerio. Time-inhomogeneous and nonlinear quantum evolutions. Lecture Notes in Mathematics, 1442:162–176, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085508/.

Friedrich:1991:BPT

[Fri91] Volkmar Friedrich. Backscattered photons — are they useful for a surface-near tomography? Lecture Notes in Mathematics, 1497: 63–65, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084508/.

Frosali:1991:FAT

[Fro91] Giovanni Frosali. Functional-analytic techniques in the study of time-dependent electron swarms in weakly ionized gases. Lecture Notes in Mathematics, 1460:107–139, 1991. CODEN LN-MAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091364/.

Fruchard:1991:EBD

[Fru91] Augustin Fruchard. Existence of bifurcation delay: The discrete case. Lecture Notes in Mathematics, 1493:87–106, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085026/.

Fejer:1990:DCM

[FS90a] Peter A. Fejer and Richard A. Shore. A direct construction of a minimal recursively enumerable truth-table degree. Lecture Notes in Mathematics, 1432:187–204, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086118/.

Fichtner:1990:LIB

[FS90b] Karl-Heinz Fichtner and Uwe Schreiter. Locally independent boson systems. Lecture Notes in Mathematics, 1442:145–161, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085507/.

Fomenko:1990:ERM

[FS90c] A. T. Fomenko and V. V. Sharko. Exact round Morse functions, inequalities of Morse type and integrals of Hamiltonian systems. Lecture Notes in Mathematics, 1453:51–67, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085946/.

Francoise:1990:RAV

[FS90d] J.-P. Françoise and R. Silhol. Real abelian varieties and the singularities of an integrable Hamiltonian system. Lecture Notes in Mathematics, 1420:121–127, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083915/.

Franke:1991:EAA

[FS91] Jeffery M. Franke and Harlan W. Stech. Extensions of an algorithm for the analysis of nongeneric Hopf bifurcations, with applications to delay-difference equations. Lecture Notes in Mathematics, 1475:161–175, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083488/.

Farwig:1992:ARE

[FS92] Reinhard Farwig and Hermann Sohr. An approach to resolvent estimates for the Stokes equations in L^{q}-spaces. Lecture Notes in Mathematics, 1530:97–110, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090336/.

Fouvry:1990:DTE

[FT90] E. Fouvry and G. Tenenbaum. Diviseurs de Titchmarsh des entiers sans grand facteur premier. (French) []. Lecture Notes in Mathematics, 1434:86–102, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097126/.

Finkelstein:1993:DWR

[FT93] Mark Finkelstein and Howard G. Tucker. On the distribution of the Wilcoxon Rank-Sum statistic. Lecture Notes in Mathematics, 1546:26–32, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084479/.

Fuhrmann:1991:HSC

[Fuh91] P. A. Fuhrmann. On the Hamiltonian structure in the computation of singular values for a class of Hankel operators. Lecture Notes in Mathematics, 1496:250–276, 1991. CODEN LN-MAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084471/.

Fujii:1990:UDZ

[Fuj90a] Akio Fujii. Uniform distribution of the zeros of the Riemann zeta function and the mean value theorems of Dirichlet L-functions (II). Lecture Notes in Mathematics, 1434:103–125, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097127/.

Fujita:1990:SDP

[Fuj90b] Takao Fujita. On singular Del Pezzo varieties. Lecture Notes in Mathematics, 1417:117–128, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083337/.

Fujiki:1991:HSM

[Fuj91] Akira Fujiki. Hyperkähler structure on the moduli space of flat bundles. Lecture Notes in Mathematics, 1468:1–83, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086187/.

Fujita:1992:ABA

[Fuj92] Takao Fujita. On adjoint bundles of ample vector bundles. Lecture Notes in Mathematics, 1507:105–112, 1992. CODEN LN-MAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094513/.

Fujiwara:1993:SFP

[Fuj93] Daisuke Fujiwara. Some Feynman path integrals as oscillatory integrals over a Sobolev manifold. Lecture Notes in Mathematics, 1540:39–53, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085473/.

Fukushima:1993:TTR

[Fuk93] Masatoshi Fukushima. Two topics related to Dirichlet forms: quasi everywhere convergences and additive functionals. Lecture Notes in Mathematics, 1563:21–53, 1993. CODEN LN-MAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074090/.

Fursikov:1990:NSE

[Fur90a] A. V. Fursikov. Navier–Stokes equations from the point of view of the theory of ill-posed boundary value problems. Lecture Notes in Mathematics, 1431:23–39, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086052/.

Fursikov:1990:SAN

[Fur90b] A. V. Fursikov. On the statistical approach to the Navier–Stokes equations. Lecture Notes in Mathematics, 1431:40–48, 1990. CO-DEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086053/.

Furter:1991:BSR

[Fur91] J. E. Furter. On the bifurcations of subharmonics in reversible systems. Lecture Notes in Mathematics, 1463:167–192, 1991. CO-DEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085431/.

Fursikov:1992:CPC

[Fur92] A. V. Fursikov. The closure problem for the chain of the Friedman–Keller moment equations in the case of large Reynolds numbers. Lecture Notes in Mathematics, 1530:226–245, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090345/.

Fournier:1991:MCT

[FW91] G. Fournier and M. Willem. The mountain circle theorem. Lecture Notes in Mathematics, 1475:147–160, 1991. CODEN LN-MAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083487/.

Furi:1993:TMO

[FZ93] Massimo Furi and Pietro Zecca, editors. Topological Methods for Ordinary Differential Equations: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, June 24–July 2, 1991, volume 1537 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1537. URL http:/

/link.springer.com/book/10.1007/BFb0085073;   http://
www.springerlink.com/content/978-3-540-47563-7.

Fuller:1990:FCM

[FZH90] Kent R. Fuller and Birge Zimmermann-Huisgen. Filtered Cartan matrices for Artinian rings of low Loewy length. Lecture Notes in Mathematics, 1448:72–79, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091252/.

Galdi:1992:OBV

[Gal92] Giovanni P. Galdi. On the Oseen boundary-value problem in exterior domains. Lecture Notes in Mathematics, 1530:111–131, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090337/.

Gao:1997:NCT

[Gao97] Fuqing Gao. A note on Cramér’s theorem. Lecture Notes in Mathematics, 1655:77–79, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0119294/.

Gauduchon:1991:SDM

[Gau91] Paul Gauduchon. Self-dual manifolds with non-negative Ricci operator. Lecture Notes in Mathematics, 1481:55–61, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083628/.

Gondard-Cozette:1990:CFR

[GC90] Danielle Gondard-Cozette. Chainable fields and real algebraic geometry. Lecture Notes in Mathematics, 1420:128–148, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083916/.

Gondard-Cozette:1992:ONE

[GC92] Danielle Gondard-Cozette. Sur les ordres de niveau 2^{n} et sur une extension du 17ème problème de Hilbert. (French) []. Lecture Notes in Mathematics, 1524:257–266, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084625/.

Gonzalez-Corbalan:1990:SIL

[GCR90] A. González-Corbalan and T. Recio. Shape invariant lists and realization as plane real algebraic curves with doublepoints. Lecture Notes in Mathematics, 1420:149–169, 1990. CODEN LN-MAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083917/.

Deng:1991:GPD

[gDH91] Dong gao Deng and Yongsheng Han. On a generalized paraproduct defined by non-convolution. Lecture Notes in Mathematics, 1494:46–53, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087755/.

Grassin:1991:DTS

[GDT91] Patricia Grassin, Bernard Duchene, and Walid Tabbara. Diffraction tomography some applications and extension to 3-D ultrasound imaging. Lecture Notes in Mathematics, 1497:98–105, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084510/.

Ge:1992:NSY

[Ge92] Mo-Lin Ge. New solutions of Yang–Baxter equations and quantum group structures. Lecture Notes in Mathematics, 1510:245– 258, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101193/.

Grosse-Erdmann:1998:AMO

[GE98a] Karl-Goswin Grosse-Erdmann. Applications to matrix operators and inequalities. Lecture Notes in Mathematics, 1679:49–76, 1998. CODEN LNMAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093490/.

Grosse-Erdmann:1998:BT

[GE98b] Karl-Goswin Grosse-Erdmann. The blocking technique. Lecture Notes in Mathematics, 1679:7–22, 1998. CODEN LN-MAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093488/.

Grosse-Erdmann:1998:BTW

[GE98c] Karl-Goswin Grosse-Erdmann. The Blocking Technique, Weighted Mean Operators and Hardy’s Inequality, volume 1679 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 114 pp. LCCN QA3 .L28 no. 1679. URL http:/

/link.springer.com/book/10.1007/BFb0093486;   http://
www.springerlink.com/content/978-3-540-69664-3.

Grosse-Erdmann:1998:IA

[GE98d] Karl-Goswin Grosse-Erdmann. Integral analogues. Lecture Notes in Mathematics, 1679:77–92, 1998. CODEN LNMAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093491/.

Grosse-Erdmann:1998:I

[GE98e] Karl-Goswin Grosse-Erdmann. Introduction. Lecture Notes in Mathematics, 1679:1–6, 1998. CODEN LNMAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093487/.

Grosse-Erdmann:1998:SS

[GE98f] Karl-Goswin Grosse-Erdmann. The sequence spaces c(a, p, q) and d(a, p, q). Lecture Notes in Mathematics, 1679:23–47, 1998. CO-DEN LNMAA2. ISBN 3-540-63902-0 (print), 3-540-69664-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093489/.

Gehring:1992:TQM

[Geh92] F. W. Gehring. Topics in quasiconformal mappings. Lecture Notes in Mathematics, 1508:20–38, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094236/.

Gekeler:1990:RCG

[Gek90] Ernst-Ulrich Gekeler. De Rham cohomology and the Gauss– Manin connection for Drinfeld modules. Lecture Notes in Mathematics, 1454:223–255, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091142/.

Gelman:1990:SPT

[Gel90] B. D. Gel’man. On some problems from the theory of fixed points of multivalued mappings. Lecture Notes in Mathematics, 1453: 227–244, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085958/.

Gelman:1992:GDM

[Gel92] B. D. Gel’man. Generalized degree of multi-valued mappings. Lecture Notes in Mathematics, 1520:173–192, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084721/.

Geman:1990:RFI

[Gem90] D. Geman. Random fields and inverse problems in imaging. Lecture Notes in Mathematics, 1427:113–193, 1990. CODEN LN-MAA2. ISBN 3-540-53508-X (print), 3-540-46718-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103042/.

Gershkovich:1990:EBN

[Ger90] V. Ya. Gershkovich. Estimates for ϵ-balls of nonholonomic metrics. Lecture Notes in Mathematics, 1453:69–85, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085947/.

Gervais:1992:QGS

[Ger92] Jean-Loup Gervais. Quantum group symmetry of 2D gravity. Lecture Notes in Mathematics, 1510:259–276, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101194/.

Geronimo:1993:POU

[Ger93] J. S. Geronimo. Polynomials orthogonal on the unit circle with random recurrence coefficients. Lecture Notes in Mathematics, 1550:43–61, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0117473/.

Gerstenhaber:1992:QS

[GGS92] Murray Gerstenhaber, Anthony Giaquinto, and Samuel D. Schack. Quantum symmetry. Lecture Notes in Mathematics, 1510: 9–46, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101176/.

Giga:1993:ESS

[GGS93] Mariko Giga, Yoshikazu Giga, and Hermann Sohr. L^{p} estimates for the Stokes system. Lecture Notes in Mathematics, 1540:55– 67, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085474/.

Gine:1997:LPT

[GGSC97] Evarist Giné, Geoffrey R. Grimmett, and Laurent Saloff-Coste, editors. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXVI — 1996, volume 1665 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63190-9 (print), 3-540-69210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1665. URL http:/

/link.springer.com/book/10.1007/BFb0092617;   http://
www.springerlink.com/content/978-3-540-69210-2.

Gavrilov:1990:LCZ

[GH90] Ljubomir Gavrilov and Emil Horozov. Limit cycles and zeroes of Abelian integrals satisfying third order Picard–Fuchs equations. Lecture Notes in Mathematics, 1455:160–186, 1990. CODEN LN-MAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085392/.

Ghrist:1997:B

[GHS97a] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Bifurcations. Lecture Notes in Mathematics, 1654:107–142, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093392/.

Ghrist:1997:CR

[GHS97b] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Concluding remarks. Lecture Notes in Mathematics, 1654:167– 191, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093394/.

Ghrist:1997:Ia

[GHS97c] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Introduction. Lecture Notes in Mathematics, 1654:1–4, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093388/.

Ghrist:1997:Ib

[GHS97d] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Invariants. Lecture Notes in Mathematics, 1654:143–166, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093393/.

Ghrist:1997:KLT

[GHS97e] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Knots and Links in Three-Dimensional Flows, volume 1654 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 208 pp. LCCN QA3 .L28 no. 1654. URL http:/

/link.springer.com/book/10.1007/BFb0093387;   http://
www.springerlink.com/content/978-3-540-68347-6.

Ghrist:1997:P

[GHS97f] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Prerequisites. Lecture Notes in Mathematics, 1654:5–32, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093389/.

Ghrist:1997:TT

[GHS97g] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Template theory. Lecture Notes in Mathematics, 1654:69–106, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093391/.

Ghrist:1997:T

[GHS97h] Robert W. Ghrist, Philip J. Holmes, and Michael C. Sullivan. Templates. Lecture Notes in Mathematics, 1654:33–68, 1997. CODEN LNMAA2. ISBN 3-540-62628-X (print), 3-540-68347-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093390/.

Gautheron:1991:FDD

[GI91] Véronique Gautheron and Emmanuel Isambert. Finitely differentiable ducks and finite expansions. Lecture Notes in Mathematics, 1493:40–56, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085023/.

Gilles:1990:GBS

[Gil90] Christol Gilles. Globally bounded solutions of differential equations. Lecture Notes in Mathematics, 1434:45–64, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097124/.

Gillet:1992:CAT

[Gil92] Henri Gillet. Comparing algebraic and topological K-theory. Lecture Notes in Mathematics, 1491:55–99, 1992. CODEN LN-MAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088879/.

Gill:1994:LSA

[Gil94] Richard D. Gill. Lectures on survival analysis. Lecture Notes in Mathematics, 1581:115–241, 1994. CODEN LNMAA2. ISBN 3-540-58208-8 (print), 3-540-48568-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073873/.

Gine:1997:DLT

[Gin97a] Evarist Gine. Decoupling and limit theorems for u-statistics and u-processes. Lecture Notes in Mathematics, 1665:1–35, 1997. CO-DEN LNMAA2. ISBN 3-540-63190-9 (print), 3-540-69210-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092618/.

Gine:1997:LSA

[Gin97b] Evarist Gine. Lectures on some aspects of the bootstrap. Lecture Notes in Mathematics, 1665:37–151, 1997. CODEN LN-MAA2. ISBN 3-540-63190-9 (print), 3-540-69210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092619/.

Gindikin:1998:RIG

[Gin98] Simon Gindikin. Real integral geometry and complex analysis. Lecture Notes in Mathematics, 1684:70–98, 1998. CODEN LNMAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096091/.

Giroux:1990:FGI

[Gir90] Emmanuel Giroux. Formes generatrices d’immersions lagrangiennes dans un espace cotangent. (French) []. Lecture Notes in Mathematics, 1416:139–145, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097468/.

Git:1998:ASP

[Git98] Y. Git. Almost sure path properties of branching diffusion processes. Lecture Notes in Mathematics, 1686:108–127, 1998. CO-DEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101754/.

Ghoussoub:1991:OWF

[GJ91] N. Ghoussoub and W. B. Johnson. Operators which factor through Banach lattices not containing c_{0}. Lecture Notes in Mathematics, 1470:68–71, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090213/.

Goto:1990:SCU

[GK90] Kazuo Goto and Takeshi Kano. Some conditions on uniform distribution of monotone sequences. Lecture Notes in Mathematics, 1434:126–132, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097128/.

Gohberg:1994:HOF

[GK94] I. Gohberg and M. A. Kaashoek. Holomorphic operator functions. Lecture Notes in Mathematics, 1573:179–210, 1994. CO-DEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100205/.

Grandits:1998:CSS

[GK98] Peter Grandits and Leszek Krawczyk. Closedness of some spaces of stochastic integrals. Lecture Notes in Mathematics, 1686:73– 85, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101752/.

Gurarie:1991:RBS

[GKKL91] David Gurarie, Gerhard Kalisch, Mark Kon, and Edward Landesman. Radial bounds for Schrödinger operators in Euclidean domains. Lecture Notes in Mathematics, 1475:183–190, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083490/.

Graham:1996:FM

[GKM^{+}96] Carl Graham, Thomas G. Kurtz, Sylvie Méléard, Philip E. Protter, Mario Pulvirenti, and Denis Talay. Front matter. Lecture Notes in Mathematics, 1627:??, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-68513-5/1.pdf.

Gohberg:1991:TVE

[GKW91] I. Gohberg, M. A. Kaashoek, and H. J. Woerdeman. Time variant extension problems of Nehari type and the band method. Lecture Notes in Mathematics, 1496:309–323, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084473/.

Garcia:1992:GCW

[GL92] Arnaldo García and R. F. Lax. Goppa codes and Weierstrass gaps. Lecture Notes in Mathematics, 1518:33–42, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087991/.

Gliklikh:1992:SAG

[Gli92] Yu. E. Gliklikh. Stochastic analysis, groups of diffemorphisms and Lagrangian description of viscous incompressible fluid. Lecture Notes in Mathematics, 1520:1–18, 1992. CODEN LN-MAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084712/.

Goldberg:1994:ESF

[GLO94] A. A. Goldberg, B. Ya. Levin, and I. V. Ostrovskii. Entire and subharmonic functions. Lecture Notes in Mathematics, 1574:295– 338, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101066/.

Groszek:1990:ICH

[GM90] Marcia Groszek and Michael Mytilinaios. Σ_{2}-induction and the construction of a high degree. Lecture Notes in Mathematics, 1432:205–221, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086119/.

Galtchouk:1997:WED

[GN97] Leonid I. Galtchouk and Alexandre A. Novikov. On Wald’s equation. Discrete time case. Lecture Notes in Mathematics, 1655: 126–135, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0119299/.

Goldstein:1990:BCA

[Gol90] Richard Goldstein. Bounded cancellation of automorphisms of free products. Lecture Notes in Mathematics, 1440:81–89, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084453/.

Goldsheid:1991:LEA

[Gol91a] I. Ya. Goldsheid. Lyapunov exponents and asymptotic behaviour of the product of random matrices. Lecture Notes in Mathematics, 1486:23–37, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086655/.

Golse:1991:PTN

[Gol91b] François Golse. Particle transport in nonhomogeneous media. Lecture Notes in Mathematics, 1460:152–170, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091366/.

Goldfeld:1994:SIW

[Gol94] Dorian Goldfeld. A spectral interpretation of Weil’s explicit formula. Lecture Notes in Mathematics, 1593:135–152, 1994. CO-DEN LNMAA2. ISBN 3-540-58673-3 (print), 3-540-49041-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074041/.

Goryunov:1991:IFP

[Gor91] V. V. Goryunov. The intersection form of a plane isolated line singularity. Lecture Notes in Mathematics, 1462:172–184, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086380/.

Gottsche:1994:CRR

[Göt94a] Lothar Göttsche. The chow ring of relative Hilbert schemes of projective bundles. Lecture Notes in Mathematics, 1572:145– 183, 1994. CODEN LNMAA2. ISBN 3-540-57814-5 (print), 3-540-48338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073495/.

Gottsche:1994:CBN

[Göt94b] Lothar Göttsche. Computation of the Betti numbers of Hilbert schemes. Lecture Notes in Mathematics, 1572:12–80, 1994. CO-DEN LNMAA2. ISBN 3-540-57814-5 (print), 3-540-48338-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073493/.

Gottsche:1994:FF

[Göt94c] Lothar Göttsche. Fundamental facts. Lecture Notes in Mathematics, 1572:1–11, 1994. CODEN LNMAA2. ISBN 3-540-57814-5 (print), 3-540-48338-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073492/.

Gottsche:1994:HSZ

[Göt94d] Lothar Göttsche. Hilbert Schemes of Zero-Dimensional Subschemes of Smooth Varieties, volume 1572 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57814-5 (print), 3-540-48338-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 196 pp. LCCN QA3 .L28 no. 1572. URL http:/

/link.springer.com/book/10.1007/BFb0073491;   http://
www.springerlink.com/content/978-3-540-48338-0.

Gottsche:1994:VSH

[Göt94e] Lothar Göttsche. The varieties of second and higher order data. Lecture Notes in Mathematics, 1572:81–144, 1994. CODEN LN-MAA2. ISBN 3-540-57814-5 (print), 3-540-48338-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073494/.

Galdi:1991:FRN

[GP91a] G. P. Galdi and M. Padula. Further results in the nonlinear stability of the magnetic Bénard problem. Lecture Notes in Mathematics, 1460:140–151, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091365/.

Gromov:1991:RLI

[GP91b] M. Gromov and P. Pansu. Rigidity of lattices: an introduction. Lecture Notes in Mathematics, 1504:39–137, 1991. CODEN LNMAA2. ISBN 3-540-55017-8 (print), 3-540-46651-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094289/.

Gurevich:1992:YBE

[GR92] D. Gurevich and V. Rubtsov. Yang–Baxter equation and deformation of associative and Lie algebras. Lecture Notes in Mathematics, 1510:47–55, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101177/.

Grabner:1990:DER

[Gra90a] Peter J. Grabner. On digit expansions with respect to second order linear recurring sequences. Lecture Notes in Mathematics, 1452:58–64, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096981/.

Gramain:1990:FEO

[Gra90b] François Gramain. Fonctions entières d’une ou plusieurs variables complexes prenant des valeurs entières sur une progression géométrique. (French) []. Lecture Notes in Mathematics, 1415: 123–137, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084883/.

Grangeat:1991:MFC

[Gra91a] Pierre Grangeat. Mathematical framework of cone beam 3D reconstruction via the first derivative of the Radon transform. Lecture Notes in Mathematics, 1497:66–97, 1991. CODEN LN-MAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084509/.

Gray:1991:OES

[Gra91b] John W. Gray. Order-enriched sketches for typed lambda calculi. Lecture Notes in Mathematics, 1488:105–130, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084216/.

Graffi:1994:TCC

[Gra94] Sandro Graffi, editor. Transition to Chaos in Classical and Quantum Mechanics: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, July 6–13, 1991, volume 1589 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1589.

URL http://link.springer.com/book/10.1007/BFb0074073;
 http://www.springerlink.com/content/978-3-540-48782-
1.

Graham:1996:SPA

[Gra96] Carl Graham. A statistical physics approach to large networks. Lecture Notes in Mathematics, 1627:127–147, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093179/.

Grandits:1999:SRB

[Gra99] Peter Grandits. Some remarks on L^{∞}, H^{∞} and BMO. Lecture Notes in Mathematics, 1709:342–348, 1999. CODEN LN-MAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096524/.

Gorostiza:1990:PDP

[GRCW90] L. G. Gorostiza, S. Roelly-Coppoletta, and A. Wakolbinger. Sur la persistance du processus de Dawson–Watanabe stable. L’interversion de la limite en temps et de la renormalisation. (French) []. Lecture Notes in Mathematics, 1426:275–281, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083770/.

Greenlees:1990:PBH

[Gre90] J. P. C. Greenlees. The power of modP Borel homology. Lecture Notes in Mathematics, 1418:140–151, 1990. CODEN LN-MAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083699/.

Greither:1992:CU

[Gre92a] Cornelius Greither. Calculations with units. Lecture Notes in Mathematics, 1534:67–76, 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089169/.

Greither:1992:CHT

[Gre92b] Cornelius Greither. Corestriction and Hilbert’s Theorem 90. Lecture Notes in Mathematics, 1534:55–66, 1992. CODEN LN-MAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089168/.

Greither:1992:CGE

[Gre92c] Cornelius Greither. Cyclic Galois Extensions of Commutative Rings, volume 1534 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 145 pp. LCCN QA3 .L28 no. 1534. URL http:

//link.springer.com/book/10.1007/BFb0089165;   http:/
/www.springerlink.com/content/978-3-540-47539-2.

Greither:1992:CGT

[Gre92d] Cornelius Greither. Cyclic Galois theory without the condition “‘p^{−1} ≥ R”. Lecture Notes in Mathematics, 1534:109–139, 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089172/.

Greither:1992:CEI

[Gre92e] Cornelius Greither. Cyclic p-extensions and ie771--extensions of number fields. Lecture Notes in Mathematics, 1534:77–96, 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089170/.

Greither:1992:CD

[Gre92f] Cornelius Greither. Cyclotomic descent. Lecture Notes in Mathematics, 1534:32–54, 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089167/.

Greither:1992:GTCa

[Gre92g] Cornelius Greither. Galois theory of commutative rings. Lecture Notes in Mathematics, 1534:1–31, 1992. CODEN LN-MAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089166/.

Greither:1992:GTCb

[Gre92h] Cornelius Greither. Geometric theory: cyclic extensions of finitely generated fields. Lecture Notes in Mathematics, 1534: 97–108, 1992. CODEN LNMAA2. ISBN 3-540-56350-4 (print), 3-540-47539-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089171/.

Green:1994:IMH

[Gre94] Mark L. Green. Infinitesimal methods in Hodge theory. Lecture Notes in Mathematics, 1594:1–92, 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-49046-3_1/.

Greenberg:1999:ITE

[Gre99] Ralph Greenberg. Iwasawa theory for elliptic curves. Lecture Notes in Mathematics, 1716:51–144, 1999. CODEN LN-MAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093453/.

Grimmett:1997:PDS

[Gri97] Geoffrey Grimmett. Percolation and disordered systems. Lecture Notes in Mathematics, 1665:153–300, 1997. CODEN LN-MAA2. ISBN 3-540-63190-9 (print), 3-540-69210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092620/.

Grorud:1990:CNS

[Gro90a] Axel Grorud. Un crochet non-symétrique en calcul stochastique anticipatif. (French) []. Lecture Notes in Mathematics, 1444:183– 192, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083615/.

Groves:1990:RSH

[Gro90b] J. R. J. Groves. Rewriting systems and homology of groups. Lecture Notes in Mathematics, 1456:114–141, 1990. CODEN LN-MAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100735/.

Gross:1993:LSI

[Gro93] Leonard Gross. Logarithmic Sobolev inequalities and contractivity properties of semigroups. Lecture Notes in Mathematics, 1563:

54–88, 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074091/.

Groeneboom:1996:LIP

[Gro96] Piet Groeneboom. Lectures on inverse problems. Lecture Notes in Mathematics, 1648:67–164, 1996. CODEN LNMAA2. ISBN 3-540-62055-9 (print), 3-540-49635-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095675/.

Grosshans:1997:AHS

[Gro97a] Frank D. Grosshans. Algebraic Homogeneous Spaces and Invariant Theory, volume 1673 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 148 pp. LCCN QA3 .L28 no. 1673.

URL http://link.springer.com/book/10.1007/BFb0093525;
 http://www.springerlink.com/content/978-3-540-69617-
9.

Grosshans:1997:C

[Gro97b] Frank D. Grosshans. Complexity. Lecture Notes in Mathematics, 1673:106–137, 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093530/.

Grosshans:1997:E

[Gro97c] Frank D. Grosshans. Errata. Lecture Notes in Mathematics, 1673: e1–e2, 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093531.

Grosshans:1997:I

[Gro97d] Frank D. Grosshans. Introduction. Lecture Notes in Mathematics, 1673:1–4, 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093526/.

Grosshans:1997:IMU

[Gro97e] Frank D. Grosshans. Invariants of maximal unipotent subgroups. Lecture Notes in Mathematics, 1673:71–105, 1997. CODEN LN-MAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093529/.

Grosshans:1997:OS

[Gro97f] Frank D. Grosshans. Observable subgroups. Lecture Notes in Mathematics, 1673:5–32, 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093527/.

Grosshans:1997:TP

[Gro97g] Frank D. Grosshans. The transfer principle. Lecture Notes in Mathematics, 1673:33–70, 1997. CODEN LNMAA2. ISBN 3-540-63628-5 (print), 3-540-69617-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093528/.

Grubb:1991:PPD

[Gru91a] Gerd Grubb. Parabolic pseudo-differential boundary problems and applications. Lecture Notes in Mathematics, 1495:46–117, 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085122/.

Grunbaum:1991:DTR

[Grü91b] F. Alberto Grünbaum. Diffuse tomography: a refined model. Lecture Notes in Mathematics, 1497:106–111, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084511/.

Grubb:1992:IVP

[Gru92] Gerd Grubb. Initial value problems for the Navier–Stokes equations with Neumann conditions. Lecture Notes in Mathematics, 1530:262–283, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090347/.

Gasull:1990:BAD

[GS90a] A. Gasull and J. Sotomayor. On the basin of attraction of dissipative planar vector fields. Lecture Notes in Mathematics, 1455: 187–195, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085393/.

Gutierrez:1990:PLC

[GS90b] C. Gutiérrez and J. Sotomayor. Periodic lines of curvature bifurcating from Darbouxian umbilical connections. Lecture Notes in Mathematics, 1455:196–229, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085394/.

Gonchar:1993:MAT

[GS93] Andrei A. Gonchar and Edward B. Saff, editors. Methods of Approximation Theory in Complex Analysis and Mathematical Physics: Leningrad, May 13–24, 1991, volume 1550 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1550.

URL http://link.springer.com/book/10.1007/BFb0117469;
 http://www.springerlink.com/content/978-3-540-47792-
1.

Galbiati:1990:RAA

[GT90] Margherita Galbiati and Alberto Tognoli, editors. Real Analytic and Algebraic Geometry: Proceedings of the Conference held in Trento, Italy, October 3–7, 1988, volume 1420 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1420.

URL http://link.springer.com/book/10.1007/BFb0083906;
 http://www.springerlink.com/content/978-3-540-46952-
0.

Giblin:1991:LRR

[GT91] P. J. Giblin and F. Tari. Local reflexional and rotational symmetry in the plane. Lecture Notes in Mathematics, 1462:154– 171, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086379/.

Gamboa:1992:PPM

[GU92] J. M. Gamboa and C. Ueno. Proper polynomial maps: The real case. Lecture Notes in Mathematics, 1524:240–256, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084624/.

Gubeladze:1990:CAT

[Gub90] Joseph Gubeladze. Classical algebraic K-theory of monoid algebras. Lecture Notes in Mathematics, 1437:36–94, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086718/.

Gustafsson:1990:CPC

[Gus90] Ivar Gustafsson. A class of preconditioned conjugate gradient methods applied to finite element equations. Lecture Notes in Mathematics, 1457:44–57, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090901/.

Gorenflo:1991:AIE

[GV91a] Rudolf Gorenflo and Sergio Vessella. Abel Integral Equations: Analysis and Applications, volume 1461 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 215 pp. LCCN QA3 .L28 no. 1461.

URL http://link.springer.com/book/10.1007/BFb0084665;
 http://www.springerlink.com/content/978-3-540-46949-
0.

Gorenflo:1991:ATA

[GV91b] Rudolf Gorenflo and Sergio Vessella. Applications of a transformed Abel integral equation. Lecture Notes in Mathematics, 1461:35–63, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084669/.

Gorenflo:1991:AAO

[GV91c] Rudolf Gorenflo and Sergio Vessella. Applications of Abel’s original integral equation: Determination of potentials. Lecture Notes in Mathematics, 1461:26–34, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084668/.

Gorenflo:1991:BTR

[GV91d] Rudolf Gorenflo and Sergio Vessella. Basic theory and representation formulas. Lecture Notes in Mathematics, 1461:8–25, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084667/.

Gorenflo:1991:EUT

[GV91e] Rudolf Gorenflo and Sergio Vessella. Existence and uniqueness theorems. Lecture Notes in Mathematics, 1461:83–94, 1991. CO-DEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084671/.

Gorenflo:1991:ISL

[GV91f] Rudolf Gorenflo and Sergio Vessella. Illposedness and stabilization of linear Abel integral equations of first kind. Lecture Notes in Mathematics, 1461:154–181, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084674/.

Gorenflo:1991:I

[GV91g] Rudolf Gorenflo and Sergio Vessella. Introduction. Lecture Notes in Mathematics, 1461:1–7, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084666/.

Gorenflo:1991:NAI

[GV91h] Rudolf Gorenflo and Sergio Vessella. Nonlinear Abel integral equations of second kind. Lecture Notes in Mathematics, 1461: 129–153, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084673/.

Gorenflo:1991:NTF

[GV91i] Rudolf Gorenflo and Sergio Vessella. On numerical treatment of first kind Abel integral equations. Lecture Notes in Mathematics, 1461:182–194, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084675/.

Gorenflo:1991:RBA

[GV91j] Rudolf Gorenflo and Sergio Vessella. Relations between Abel transform and other integral transforms. Lecture Notes in Mathematics, 1461:95–128, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084672/.

Gorenflo:1991:SPA

[GV91k] Rudolf Gorenflo and Sergio Vessella. Smoothing properties of the Abel operators. Lecture Notes in Mathematics, 1461:64–82, 1991. CODEN LNMAA2. ISBN 3-540-53668-X (print), 3-540-46949-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084670/.

Goderis:1990:QCL

[GVV90] D. Goderis, A. Verbeure, and P. Vets. Quantum central limit and coarse graining. Lecture Notes in Mathematics, 1442:178– 193, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085510/.

Gusein-Zade:1991:DEM

[GZ91] S. M. Gusein-Zade. On the degree of an equivariant map. Lecture Notes in Mathematics, 1462:185–193, 1991. CODEN LN-MAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086381/.

Hog-Angeloni:1990:HTC

[HA90a] Cynthia Hog-Angeloni. On the homotopy type of 2-complexes with a free product of cyclic groups as fundamental group. Lecture Notes in Mathematics, 1440:96–108, 1990. CODEN LN-MAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084455/.

Hog-Angeloni:1990:STP

[HA90b] Cynthia Hog-Angeloni. A short topological proof of Cohn’s theorem. Lecture Notes in Mathematics, 1440:90–95, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084454/.

Hadeler:1999:RTS

[Had99] K. P. Hadeler. Reaction transport systems in biological modelling. Lecture Notes in Mathematics, 1714:95–150, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092376/.

Hajduk:1991:OGE

[Haj91] Bogusław Hajduk. On the obstruction group to existence of Riemannian metrics of positive scalar curvature. Lecture Notes in Mathematics, 1481:62–72, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083629/.

Hale:1991:DD

[Hal91] Jack K. Hale. Dynamics and delays. Lecture Notes in Mathematics, 1475:16–30, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083476/.

Hog-Angeloni:1990:BIO

[HALM90] Cynthia Hog-Angeloni, M. Paul Latiolais, and Wolfgang Metzler. Bias ideals and obstructions to simple-homotopy equivalence. Lecture Notes in Mathematics, 1440:109–121, 1990. CODEN LN-MAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084456/.

Hamenstadt:1991:CMP

[Ham91] Ursula Hamenstädt. Compact manifolds with 1/4-pinched negative curvature. Lecture Notes in Mathematics, 1481:73–78, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083630/.

Hansen:1990:SSM

[Han90] Vagn Lundsgaard Hansen. The space of self maps on the 2sphere. Lecture Notes in Mathematics, 1425:40–47, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083829/.

Han:1991:BCZ

[Han91] Yongsheng Han. H^{p} boundedness of Calderón–Zygmund operators for product domains. Lecture Notes in Mathematics, 1494: 54–67, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087756/.

Hansen:1992:DLV

[Han92] Johan P. Hansen. Deligne–Lusztig varieties and group codes. Lecture Notes in Mathematics, 1518:63–81, 1992. CODEN LN-MAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087993/.

Harase:1990:ADF

[Har90a] Takashi Harase. Algebraic dependence of formal power series. Lecture Notes in Mathematics, 1434:133–137, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097129/.

Harder:1990:SRE

[Har90b] Prof. Dr. Günter Harder. Some results on the Eisenstein cohomology of arithmetic subgroups of GL_{n}. Lecture Notes in Mathematics, 1447:85–153, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085728/.

Harris:1990:PIH

[Har90c] Michael Harris. Period invariants of Hilbert modular forms, I: Trilinear differential operators and L-functions. Lecture Notes in Mathematics, 1447:155–202, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085729/.

Harder:1993:AGM

[Har93a] Günter Harder. Andersons gemischte Motive. (German) [Anderson’s mixed motives]. Lecture Notes in Mathematics, 1562:103– 142, 1993. CODEN LNMAA2. ISBN 3-540-57408-5 (print), 3-540-48135-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090309/.

Harder:1993:BG

[Har93b] Günter Harder. Die Beispiele. (German) [The examples]. Lecture Notes in Mathematics, 1562:70–102, 1993. CODEN LN-MAA2. ISBN 3-540-57408-5 (print), 3-540-48135-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090308/.

Harder:1993:BDV

[Har93c] Günter Harder. Die Beĭlinson–Deligne-Vermutungen. (German) [The Beĭlinson–Deligne conjectures]. Lecture Notes in Mathematics, 1562:1–30, 1993. CODEN LNMAA2. ISBN 3-540-57408-5 (print), 3-540-48135-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090306/.

Harder:1993:KSV

[Har93d] Günter Harder. Die Kohomologie von Shimura-Varietäten. (German) [The cohomology of Shimura varieties]. Lecture Notes in Mathematics, 1562:31–69, 1993. CODEN LNMAA2. ISBN 3-540-57408-5 (print), 3-540-48135-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090307/.

Harder:1993:EKG

[Har93e] Günter Harder. Eisensteinkohomologie und die Konstruktion gemischter Motive. (German) [Eisenstein cohomology and the construction of mixed designs], volume 1562 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CO-DEN LNMAA2. ISBN 3-540-57408-5 (print), 3-540-48135-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 184 pp. LCCN QA3 .L28 no. 1562. URL http:/

/link.springer.com/book/10.1007/BFb0090305;   http://
www.springerlink.com/content/978-3-540-48135-5.

Hartfiel:1998:BMS

[Har98a] Darald J. Hartfiel. Behavior in Markov set-chains. Lecture Notes in Mathematics, 1695:91–113, 1998. CODEN LNMAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094591/.

Hartfiel:1998:CMS

[Har98b] Darald J. Hartfiel. Convergence of Markov set-chains. Lecture Notes in Mathematics, 1695:59–89, 1998. CODEN LN-MAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094590/.

Hartfiel:1998:I

[Har98c] Darald J. Hartfiel. Introduction. Lecture Notes in Mathematics, 1695:1–2, 1998. CODEN LNMAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094587.

Hartfiel:1998:IMS

[Har98d] Darald J. Hartfiel. Introduction to Markov set-chains. Lecture Notes in Mathematics, 1695:27–57, 1998. CODEN LN-MAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094589/.

Hartfiel:1998:MSC

[Har98e] Darald J. Hartfiel. Markov Set-Chains, volume 1695 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 131 pp. LCCN QA3 .L28 no. 1695. URL http:/

/link.springer.com/book/10.1007/BFb0094586;   http://
www.springerlink.com/content/978-3-540-68711-5.

Hartfiel:1998:SMT

[Har98f] Darald J. Hartfiel. Stochastic matrices and their variants. Lecture Notes in Mathematics, 1695:3–25, 1998. CODEN LN-MAA2. ISBN 3-540-64775-9 (print), 3-540-68711-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094588/.

Hausmann:1991:TBA

[Hau91] Jean-Claude Hausmann. Sur la topologie des bras articulés. (French) []. Lecture Notes in Mathematics, 1474:146–159, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084743/.

Hazod:1993:DSC

[Haz93] W. Hazod. On different stability-concepts for probabilities on groups. Lecture Notes in Mathematics, 1546:33–44, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084480/.

Hineva:1991:MHL

[HB91] Stefana Hineva and Evgeni Belchev. On the minimal hypersurfaces of a locally symmetric manifold. Lecture Notes in Mathematics, 1481:1–4, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083622/.

Hebey:1996:BCP

[Heb96a] Emmanuel Hebey. The best constants problems. Lecture Notes in Mathematics, 1635:58–89, 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092911/.

Hebey:1996:GP

[Heb96b] Emmanuel Hebey. Geometric preliminaries. Lecture Notes in Mathematics, 1635:1–9, 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092908/.

Hebey:1996:SE

[Heb96c] Emmanuel Hebey. Sobolev embeddings. Lecture Notes in Mathematics, 1635:17–57, 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092910/.

Hebey:1996:SS

[Heb96d] Emmanuel Hebey. Sobolev spaces. Lecture Notes in Mathematics, 1635:10–16, 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092909/.

Hebey:1996:SSP

[Heb96e] Emmanuel Hebey. Sobolev spaces in the presence of symmetries. Lecture Notes in Mathematics, 1635:90–105, 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092912/.

Hebey:1996:SSR

[Heb96f] Emmanuel Hebey. Sobolev Spaces on Riemannian Manifolds, volume 1635 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61722-1 (print), 3-540-69993-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 116 pp. LCCN QA3 .L28 no. 1635. URL http:

//link.springer.com/book/10.1007/BFb0092907;   http:/
/www.springerlink.com/content/978-3-540-69993-4.

Heck:1998:HDS

[Hec98] Matthias K. Heck. Homogeneous diffusions on the Sierpiński gasket. Lecture Notes in Mathematics, 1686:86–107, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101753/.

Hegerfeldt:1990:OPQ

[Heg90] Gerhard C. Hegerfeldt. An open problem in quantum shot noise. Lecture Notes in Mathematics, 1442:194–203, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085511/.

Hegenbarth:1992:NBL

[Heg92] Friedrich Hegenbarth. A note on the Brauer lift map. Lecture Notes in Mathematics, 1509:139–145, 1992. CODEN LN-MAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087508/.

Helton:1991:TTS

[Hel91] J. William Helton. Two topics in systems engineering: Frequency domain design and nonlinear systems. Lecture Notes in Mathematics, 1496:106–140, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084468/.

Helemskii:1992:THA

[Hel92] A. Ya. Helemskiĭ. From topological homology: algebras with different properties of homological triviality. Lecture Notes in Mathematics, 1520:19–40, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084713/.

Helgason:1998:RTW

[Hel98] Sigurdur Helgason. Radon transforms and wave equations. Lecture Notes in Mathematics, 1684:99–121, 1998. CODEN LN-MAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096092/.

Hennequin:1990:EEP

[Hen90a] Paul-Louis Hennequin, editor. École d’Été de Probabilités de Saint-Flour XVIII — 1988. (French) [], volume 1427 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53508-X (print), 3-540-46718-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1427.

URL http://link.springer.com/book/10.1007/BFb0103040;
 http://www.springerlink.com/content/978-3-540-46718-
2.

Henniart:1990:CTD

[Hen90b] Guy Henniart. Une conséquence de la théorie du changement de base pour GL(n). (French) []. Lecture Notes in Mathematics, 1434:138–142, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097130/.

Henstock:1990:IID

[Hen90c] Ralph Henstock. Integration in infinite-dimensional spaces. Lecture Notes in Mathematics, 1419:54–65, 1990. CODEN LN-MAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083099/.

Henstock:1990:INI

[Hen90d] Ralph Henstock. Introduction to the new integrals. Lecture Notes in Mathematics, 1419:7–9, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083094/.

Hensz:1990:MOE

[Hen90e] Ewa Hensz. A method of operator estimation and a strong law of large numbers in von Neumann algebras. Lecture Notes in Mathematics, 1442:204–210, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085512/.

Henri:1991:EET

[Hen91] Cornille Henri. Exact exponential type solutions to the discrete Boltzmann models. Lecture Notes in Mathematics, 1460: 70–86, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091362/.

Hennequin:1992:EEP

[Hen92] Paul Louis Hennequin, editor. École d’Été de Probabilités de Saint-Flour XX — 1990. (French) [], volume 1527 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56250-8 (print), 3-540-47490-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1527.

URL http://link.springer.com/book/10.1007/BFb0084696;
 http://www.springerlink.com/content/978-3-540-47490-
6.

Hennequin:1993:EEP

[Hen93] Paul-Louis Hennequin, editor. École d’Été de Probabilités de Saint-Flour XXI — 1991. (French) [], volume 1541 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56622-8 (print), 3-540-47608-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1541.

URL http://link.springer.com/book/10.1007/BFb0084189;
 http://www.springerlink.com/content/978-3-540-47608-
5.

Herrlich:1990:NET

[Her90a] Frank Herrlich. The nonarchimedean extended Teichmüller space. Lecture Notes in Mathematics, 1454:256–266, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091143/.

Herzog:1990:HAS

[Her90b] Jürgen Herzog. A homological approach to symbolic powers. Lecture Notes in Mathematics, 1430:32–46, 1990. CODEN LN-MAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085535/.

Herzog:1994:BL

[Her94a] Bernd Herzog. Basic lemmas. Lecture Notes in Mathematics, 1597:18–29, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074028/.

Herzog:1994:DB

[Her94b] Bernd Herzog. Distinguished bases. Lecture Notes in Mathematics, 1597:59–75, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074031/.

Herzog:1994:FF

[Her94c] Bernd Herzog. Flatifying filtrations. Lecture Notes in Mathematics, 1597:91–100, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074033/.

Herzog:1994:HS

[Her94d] Bernd Herzog. Hilbert series. Lecture Notes in Mathematics, 1597: 76–90, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074032/.

Herzog:1994:IRF

[Her94e] Bernd Herzog. Inequalities related with flat couples of local rings. Lecture Notes in Mathematics, 1597:127–142, 1994. CODEN LN-MAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074035/.

Herzog:1994:KSMa

[Her94f] Bernd Herzog. Kodaira–Spencer maps. Lecture Notes in Mathematics, 1597:101–126, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074034/.

Herzog:1994:KSMb

[Her94g] Bernd Herzog. Kodaira–Spencer Maps in Local Algebra, volume 1597 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 176 pp. LCCN QA3 .L28 no. 1597. URL http:/

/link.springer.com/book/10.1007/BFb0074026;   http://
www.springerlink.com/content/978-3-540-49103-3.

Herzog:1994:LRH

[Her94h] Bernd Herzog. On the local rings of the Hilbert scheme. Lecture Notes in Mathematics, 1597:143–170, 1994. CODEN LN-MAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074036/.

Herzog:1994:RFb

[Her94i] Bernd Herzog. Relation to flatness. Lecture Notes in Mathematics, 1597:48–58, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074030/.

Herzog:1994:RFa

[Her94j] Bernd Herzog. Ring filtrations. Lecture Notes in Mathematics, 1597:1–17, 1994. CODEN LNMAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074027/.

Herzog:1994:TFU

[Her94k] Bernd Herzog. Tangential flatness under base change. Lecture Notes in Mathematics, 1597:30–47, 1994. CODEN LN-MAA2. ISBN 3-540-58790-X (print), 3-540-49103-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074029/.

Heywood:1990:OPT

[Hey90] John G. Heywood. Open problems in the theory of the Navier– Stokes equations for viscous incompressible flow. Lecture Notes in Mathematics, 1431:1–22, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086051/.

Heyer:1993:FLT

[Hey93] Herbert Heyer. Functional limit theorems for random walks on one-dimensional hypergroups. Lecture Notes in Mathematics, 1546:45–57, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084481/.

Heindorf:1994:FBA

[HFS94a] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. σ-filtered Boolean algebras. Lecture Notes in Mathematics, 1596:97–109, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094108/.

Heindorf:1994:F

[HFS94b] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. Functors. Lecture Notes in Mathematics, 1596:69–96, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094107/.

Heindorf:1994:I

[HFS94c] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. Introduction. Lecture Notes in Mathematics, 1596:1–10, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094104/.

Heindorf:1994:RFB

[HFS94d] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. rcfiltered Boolean algebras. Lecture Notes in Mathematics, 1596: 37–68, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094106/.

Heindorf:1994:SS

[HFS94e] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. Setting the stage. Lecture Notes in Mathematics, 1596:11–35, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094105/.

Heindorf:1994:TEA

[HFS94f] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. The twisted embedding and its applications. Lecture Notes in Mathematics, 1596:147–163, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094110/.

Heindorf:1994:WPR

[HFS94g] Lutz Heindorf, Sakaé Fuchino, and Leonid B. Shapiro. Weakly projective and regularly filtered algebras. Lecture Notes in Mathematics, 1596:111–145, 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094109/.

Horie:1990:EIC

[HH90] Kuniaki Horie and Mitsuko Horie. On the exponents of ideal class groups of cm-fields. Lecture Notes in Mathematics, 1434: 143–148, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097131/.

Hamada:1992:CSM

[HH92] Noboru Hamada and Tor Helleseth. On a characterization of some minihypers in PG(t, q) (q = 3 or 4) and its applications to error-correcting codes. Lecture Notes in Mathematics, 1518: 43–62, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087992/.

Higman:1990:TCG

[Hig90] Graham Higman. Transversals and conjugacy in the group of recursive permutations. Lecture Notes in Mathematics, 1456:142– 160, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100736/.

Hindry:1998:IAV

[Hin98] Marc Hindry. Introduction to abelian varieties and the Mordell– Lang conjecture. Lecture Notes in Mathematics, 1696:85–100, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-68521-0_5/.

Hitchin:1990:GTM

[Hit90] N. J. Hitchin. The geometry and topology of moduli spaces. Lecture Notes in Mathematics, 1451:1–48, 1990. CODEN LN-MAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085064/.

Hitczenko:1991:RPM

[Hit91] Pawel Hitczenko. A remark on the paper “Martingale inequalities in rearrangement invariant function spaces” by W. B. Johnson and G. Schechtman. Lecture Notes in Mathematics, 1470: 177–182, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090221/.

Hill:1992:CGM

[HK92] Theodore P. Hill and Ulrich Krengel. On the construction of generalized measure preserving transformations with given marginals. Lecture Notes in Mathematics, 1514:114–123, 1992. CODEN LN-MAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097531/.

Hopkins:1992:MTC

[HKR92] Michael J. Hopkins, Nicholas J. Kuhn, and Douglas C. Ravenel. Morava K-theories of classifying spaces and generalized characters for finite groups. Lecture Notes in Mathematics, 1509:186– 209, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087510/.

Hayat-Legrand:1992:CHA

[HL92] Claude Hayat-Legrand. Classes homotopiques associées á une G-opération. (French) []. Lecture Notes in Mathematics, 1509: 134–138, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087507/.

Hlawka:1990:NBM

[Hla90] Edmund Hlawka. Näherungsformeln zur Berechnung von mehrfachen Integralen mit Anwendungen auf die Berechungen von Potentialen, Induktionskoeffizienten und Lösungen von Gleichungssystemen. (German) [Approximation formulas for the calculation of multiple integrals with applications to the calculation of potentials, induction coefficients, and solutions of equation systems]. Lecture Notes in Mathematics, 1452:65–111, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096982/.

Herman:1991:MMT

[HLN91] Gabor T. Herman, Alfred K. Louis, and Frank Natterer, editors. Mathematical Methods in Tomography: Proceedings of a Conference held in Oberwolfach, Germany, 5–11 June, 1990, volume 1497 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1497. URL http:/

/link.springer.com/book/10.1007/BFb0084502;   http://
www.springerlink.com/content/978-3-540-46615-4.

Hatcher:1990:FPM

[HM90] Allen Hatcher and Darryl McCullough. Finite presentation of 3-manifold mapping class groups. Lecture Notes in Mathematics, 1425:48–57, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083830/.

Hauser:1991:ADP

[HM91] Herwig Hauser and Gerd Müller. Automorphisms of direct products of algebroid spaces. Lecture Notes in Mathematics, 1462: 194–198, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086382/.

Helson:1992:CS

[HM92] Henry Helson and John E. McCarthy. Continuity of seminorms. Lecture Notes in Mathematics, 1511:88–90, 1992. CODEN LN-MAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093800/.

Hino:1991:EPS

[HMN91a] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Existence of periodic solutions and almost periodic solutions. Lecture Notes in Mathematics, 1473:254–276, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084441/.

Hino:1991:FMS

[HMN91b] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Fading memory spaces and functional differential equations. Lecture Notes in Mathematics, 1473:187–214, 1991. CODEN LN-MAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084439/.

Hino:1991:FDE

[HMN91c] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Functional Differential Equations with Infinite Delay, volume 1473 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 317 pp. LCCN QA3 .L28 no. 1473. URL http:/

/link.springer.com/book/10.1007/BFb0084432;   http://
www.springerlink.com/content/978-3-540-47388-6.

Hino:1991:FT

[HMN91d] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Fundamental theorems. Lecture Notes in Mathematics, 1473:35–52, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084434/.

Hino:1991:GLS

[HMN91e] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. General linear systems. Lecture Notes in Mathematics, 1473:99–129, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084436/.

Hino:1991:LAS

[HMN91f] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Linear autonomous systems. Lecture Notes in Mathematics, 1473:130– 176, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084437/.

Hino:1991:LPS

[HMN91g] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Linear periodic systems. Lecture Notes in Mathematics, 1473:177– 186, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084438/.

Hino:1991:PS

[HMN91h] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Phase spaces. Lecture Notes in Mathematics, 1473:1–34, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084433/.

Hino:1991:SPS

[HMN91i] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Stabilities in perturbed systems and limiting equations. Lecture Notes in Mathematics, 1473:215–253, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084440/.

Hino:1991:SIL

[HMN91j] Yoshiyuki Hino, Satoru Murakami, and Toshiki Naito. Stieltjes integrals and linear operators on B. Lecture Notes in Mathematics, 1473:53–98, 1991. CODEN LNMAA2. ISBN 3-540-54084-9 (print), 3-540-47388-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084435/.

Heywood:1990:NSE

[HMRS90] John G. Heywood, Kyûya Masuda, Reimund Rautmann, and Vsevolod A. Solonnikov, editors. The Navier–Stokes Equations Theory and Numerical Methods: Proceedings of a Conference held at Oberwolfach, FRG, Sept. 18–24, 1988, volume 1431 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 0-387-52770-2, 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1431; QA3 .L35 v.1431. URL http:/

/link.springer.com/book/10.1007/BFb0086050;   http://
www.springerlink.com/content/978-3-540-47141-7.

Heywood:1992:NSE

[HMRS92] John G. Heywood, Kyûya Masuda, Reimund Rautmann, and Vsevolod A. Solonnikov, editors. The Navier–Stokes Equations II — Theory and Numerical Methods: Proceedings of a Conference held in Oberwolfach, Germany, August 18–24, 1991, volume 1530 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 0-387-56261-3, 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1530; QA3 .L35 v.1530. URL http:

//link.springer.com/book/10.1007/BFb0090329;   http:/
/www.springerlink.com/content/978-3-540-47498-2.

Huo:1994:QPS

[HMS94] Y. Huo, I. Müller, and S. Seelecke. Quasiplasticity and pseudoelasticity in shape memory alloys. Lecture Notes in Mathematics, 1584:87–146, 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073396/.

Hilgert:1993:AOS

[HN93a] Joachim Hilgert and Karl-Hermann Neeb. Applications of ordered spaces to Lie semigroups. Lecture Notes in Mathematics, 1552:148–161, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084645/.

Hilgert:1993:CS

[HN93b] Joachim Hilgert and Karl-Hermann Neeb. Compression semigroups. Lecture Notes in Mathematics, 1552:202–253, 1993. CO-DEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084648/.

Hilgert:1993:E

[HN93c] Joachim Hilgert and Karl-Hermann Neeb. Examples. Lecture Notes in Mathematics, 1552:47–79, 1993. CODEN LN-MAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084642/.

Hilgert:1993:GTL

[HN93d] Joachim Hilgert and Karl-Hermann Neeb. Geometry and topology of Lie semigroups. Lecture Notes in Mathematics, 1552:80– 112, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084643/.

Hilgert:1993:ICO

[HN93e] Joachim Hilgert and Karl-Hermann Neeb. Invariant cones and Ol’shanskii semigroups. Lecture Notes in Mathematics, 1552:177– 201, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084647/.

Hilgert:1993:LSTb

[HN93f] Joachim Hilgert and Karl-Hermann Neeb. Lie Semigroups and their Applications, volume 1552 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 315 pp. LCCN QA3 .L28 no. 1552.

URL http://link.springer.com/book/10.1007/BFb0084640;
 http://www.springerlink.com/content/978-3-540-69987-
3.

Hilgert:1993:LSTa

[HN93g] Joachim Hilgert and Karl-Hermann Neeb. Lie semigroups and their tangent wedges. Lecture Notes in Mathematics, 1552:1– 46, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084641/.

Hilgert:1993:MSG

[HN93h] Joachim Hilgert and Karl-Hermann Neeb. Maximal semigroups in groups with cocompact radical. Lecture Notes in Mathematics, 1552:162–176, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084646/.

Hilgert:1993:OHS

[HN93i] Joachim Hilgert and Karl-Hermann Neeb. Ordered homogeneous spaces. Lecture Notes in Mathematics, 1552:113–147, 1993. CO-DEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084644/.

Hilgert:1993:RT

[HN93j] Joachim Hilgert and Karl-Hermann Neeb. Representation theory. Lecture Notes in Mathematics, 1552:254–296, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084649/.

Hilgert:1993:T

[HN93k] Joachim Hilgert and Karl-Hermann Neeb. The theory for Sl(2). Lecture Notes in Mathematics, 1552:297–302, 1993. CODEN LNMAA2. ISBN 3-540-56954-5 (print), 3-540-69987-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084650/.

Havin:1994:LCAa

[HN94a] Victor P. Havin and Nikolai K. Nikolski. Linear and Complex Analysis Problem Book 3: Part I, volume 1573 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 489 pp. LCCN QA3 .L28 no. 1573. URL http:/

/link.springer.com/book/10.1007/BFb0100201;   http://
www.springerlink.com/content/978-3-540-48367-0.

Havin:1994:LCAb

[HN94b] Victor P. Havin and Nikolai K. Nikolski. Linear and Complex Analysis Problem Book 3: Part II, volume 1574 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 507 pp. LCCN QA3 .L28 no. 1574. URL http:/

/link.springer.com/book/10.1007/BFb0101060;   http://
www.springerlink.com/content/978-3-540-48368-7.

Homer:1990:FM

[HNP^{+}90] Steven Homer, Anil Nerode, Richard A. Platek, Gerald E. Sacks, and Andre Scedrov. Front matter. Lecture Notes in Mathematics, 1429:??, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-47135-6/1.pdf.

Herman:1991:ERA

[HO91] Gabor T. Herman and Dewey Odhner. Evaluation of reconstruction algorithms. Lecture Notes in Mathematics, 1497:215– 228, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084520/.

Hobson:1998:MMM

[Hob98] David G. Hobson. The maximum maximum of a martingale. Lecture Notes in Mathematics, 1686:250–263, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101762/.

Hodges:1990:RTA

[Hod90] Timothy J. Hodges. Ring-theoretical aspects of the Bernstein– Beilinson theorem. Lecture Notes in Mathematics, 1448:155– 163, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091260/.

Hofbauer:1991:ILE

[Hof91] Franz Hofbauer. An inequality for the Ljapunov exponent of an ergodic invariant measure for a piecewise monotonic map of the interval. Lecture Notes in Mathematics, 1486:227–231, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086672/.

Holevo:1990:AID

[Hol90a] A. S. Holevo. An analog of the Itô decomposition for multiplicative processes with values in a Lie group. Lecture Notes in Mathematics, 1442:211–215, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085513/.

Holzapfel:1990:EFT

[Hol90b] R.-P. Holzapfel. An effective finiteness theorem for ball lattices. Lecture Notes in Mathematics, 1447:203–236, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085730/.

Homer:1990:ICG

[Hom90] Steven Homer. The isomorphism conjecture and its generalizations. Lecture Notes in Mathematics, 1429:1–11, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093922/.

Hong:1990:RCP

[Hon90] Le Van Hong. Relative calibrations and the problem of stability of minimal surfaces. Lecture Notes in Mathematics, 1453: 245–262, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085959/.

Horvathy:1990:DSM

[Hor90] P. A. Horváthy. Dynamical symmetries of monopole scattering. Lecture Notes in Mathematics, 1416:146–160, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097469/.

Haydon:1991:CCB

[HOR91a] R. Haydon, E. Odell, and H. Rosenthal. On certain classes of Baire-1 functions with applications to Banach space theory. Lecture Notes in Mathematics, 1470:1–35, 1991. CODEN LN-MAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090209/.

Hormander:1991:QHO

[Hör91b] Lars Hörmander. Quadratic hyperbolic operators. Lecture Notes in Mathematics, 1495:118–160, 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085123/.

Hough:1990:CMF

[Hou90] David M. Hough. Conformal mapping and Fourier–Jacobi approximations. Lecture Notes in Mathematics, 1435:57–70, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087897/.

Howland:1990:SCD

[How90] James S. Howland. Spectral concentration for dense point spectrum. Lecture Notes in Mathematics, 1450:1–11, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084894/.

Hernandez:1992:CMT

[HP92] L. J. Hernández and T. Porter. Categorical models of N-types for pro-crossed complexes and ℑ_{n}-prospaces. Lecture Notes in Mathematics, 1509:146–185, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087509/.

Huisken:1999:GEE

[HP99] Gerhard Huisken and Alexander Polden. Geometric evolution equations for hypersurfaces. Lecture Notes in Mathematics, 1713: 45–84, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092669/.

Hulek:1992:CAV

[HPSS92] Klaus Hulek, Thomas Peternell, Michael Schneider, and Frank-Olaf Schreyer, editors. Complex Algebraic Varieties: Proceedings of a Conference held in Bayreuth, Germany, April 2–6, 1990, volume 1507 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1507. URL http:/

/link.springer.com/book/10.1007/BFb0094506;   http://
www.springerlink.com/content/978-3-540-46786-1.

Habets:1991:JNO

[HRS91] P. Habets, M. Ramos, and L. Sanchez. Jumping nonlinearity for 2nd order ODE with positive forcing. Lecture Notes in Mathematics, 1475:191–203, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083491/.

Hrushovski:1998:BM

[Hru98a] Ehud Hrushovski. Back matter. Lecture Notes in Mathematics, 1696:207–216, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-68521-0/1.pdf.

Hrushovski:1998:PMT

[Hru98b] Ehud Hrushovski. Proof of Manin’s theorem by reduction to positive characteristic. Lecture Notes in Mathematics, 1696:197– 205, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-68521-0_11/.

Haught:1990:UIS

[HS90a] Christine Ann Haught and Richard A. Shore. Undecidability and initial segments of the wtt-degrees ≤ 0′. Lecture Notes in Mathematics, 1432:223–244, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086120/.

Hudson:1990:SDQ

[HS90b] R. L. Hudson and P. Shepperson. Stochastic dilations of quantum dynamical semigroups using one-dimensional quantum stochastic calculus. Lecture Notes in Mathematics, 1442:216–218, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085514/.

Harris:1991:ABS

[HS91] William A. Harris, Jr. and Yasutaka Sibuya. Asymptotic behaviors of solutions of a system of linear ordinary differential equations as t → ∞. Lecture Notes in Mathematics, 1475:210– 217, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083493/.

Harris:1992:LDF

[HS92] John C. Harris and R. James Shank. Lannes’ division functors on summands of H ∗ (B(Z/p)^{r}). Lecture Notes in Mathematics, 1509:120–133, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087506/.

Havin:1993:APH

[HS93] V. P. Havin and A. Presa Sagué. Approximation properties of harmonic vector fields and differential forms. Lecture Notes in Mathematics, 1550:149–156, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0117480/.

Heindorf:1994:NPB

[HS94] Lutz Heindorf and Leonid B. Shapiro. Nearly Projective Boolean Algebras, volume 1596 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58787-X (print), 3-540-49101-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 202 pp. LCCN QA3 .L28 no. 1596. URL http:

//link.springer.com/book/10.1007/BFb0094103;   http:/
/www.springerlink.com/content/978-3-540-49101-9. With

an Appendix by Sakaé Fuchino.

Hildebrandt:1999:CVG

[HS99] Stefan Hildebrandt and Michael Struwe, editors. Calculus of Variations and Geometric Evolution Problems: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro, Italy, June 15–22, 1996, volume 1713 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1713. URL http:/

/link.springer.com/book/10.1007/BFb0092667;   http://
www.springerlink.com/content/978-3-540-48813-2.

Hu:1991:CCM

[HSS91] Wei Hu, Xianliang Shi, and Qiyu Sun. A_{∞} condition characterized by maximal geometric mean operator. Lecture Notes in Mathematics, 1494:68–72, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087757/.

Hlawka:1990:NTA

[HT90a] Edmund Hlawka and Robert F. Tichy, editors. Number-Theoretic Analysis: Seminar, Vienna 1988–89, volume 1452 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1452.

URL http://link.springer.com/book/10.1007/BFb0096975;
 http://www.springerlink.com/content/978-3-540-46864-
6.

Holschneider:1990:RLF

[HT90b] Matthias Holschneider and Philippe Tchamitchian. Régularité locale de la fonction “non-différentiable” de Riemann. (French) []. Lecture Notes in Mathematics, 1438:102–124, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083518/.

Huneke:1990:GRI

[HU90] Graig Huneke and Bernd Ulrich. Generic residual intersections. Lecture Notes in Mathematics, 1430:47–60, 1990. CODEN LN-MAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085536/.

Hu:1991:WNI

[Hu91] Yue Hu. A weighted norm inequality for oscillatory singular integrals. Lecture Notes in Mathematics, 1494:73–81, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087758/.

Hu:1993:HPF

[Hu93] Yaozhong Hu. Hypercontractivité pour les fermions, d’après Carlen–Lieb. (French) []. Lecture Notes in Mathematics, 1557: 86–96, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087966/.

Hu:1995:DFO

[Hu95] Yaozhong Hu. On the differentiability of functions of an operator. Lecture Notes in Mathematics, 1613:218–219, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094213.

Hu:1996:SWO

[Hu96] Yaozhong Hu. Strong and weak order of time discretization schemes of stochastic differential equations. Lecture Notes in Mathematics, 1626:218–227, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094650/.

Hubbuck:1990:SSI

[Hub90] J. R. Hubbuck. Some stably indecomposable loop spaces. Lecture Notes in Mathematics, 1418:70–77, 1990. CODEN LN-MAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083693/.

Huber:1995:PIH

[Hub95a] Annette Huber. ⊗-product and internal Hom on D_{MR}. Lecture Notes in Mathematics, 1604:102–112, 1995. CODEN LN-MAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095516/.

Huber:1995:AC

[Hub95b] Annette Huber. 1-adic cohomology. Lecture Notes in Mathematics, 1604:73–80, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095512/.

Huber:1995:BN

[Hub95c] Annette Huber. Basic notions. Lecture Notes in Mathematics, 1604:2–10, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095504/.

Huber:1995:BOA

[Hub95d] Annette Huber. The Bloch–Ogus axioms. Lecture Notes in Mathematics, 1604:122–138, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095518/.

Huber:1995:CCL

[Hub95e] Annette Huber. The Chern class of a line bundle. Lecture Notes in Mathematics, 1604:140–146, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095519/.

Huber:1995:CS

[Hub95f] Annette Huber. Classifying spaces. Lecture Notes in Mathematics, 1604:146–154, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095520/.

Huber:1995:CFA

[Hub95g] Annette Huber. Comparison functors: 1-adic versus singular realization. Lecture Notes in Mathematics, 1604:80–86, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095513/.

Huber:1995:RC

[Hub95h] Annette Huber. De Rham cohomology. Lecture Notes in Mathematics, 1604:57–60, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095510/.

Huber:1995:DCE

[Hub95i] Annette Huber. Derived categories of exact categories. Lecture Notes in Mathematics, 1604:10–21, 1995. CODEN LN-MAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095505/.

Huber:1995:FDC

[Hub95j] Annette Huber. Filtered derived categories. Lecture Notes in Mathematics, 1604:22–27, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095506/.

Huber:1995:GC

[Hub95k] Annette Huber. Gluing of categories. Lecture Notes in Mathematics, 1604:28–45, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095507/.

Huber:1995:GR

[Hub95l] Annette Huber. Godement resolutions. Lecture Notes in Mathematics, 1604:45–48, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095508/.

Huber:1995:GM

[Hub95m] Annette Huber. Grothendieck motives. Lecture Notes in Mathematics, 1604:177–182, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095523/.

Huber:1995:HCC

[Hub95n] Annette Huber. Higher Chern classes. Lecture Notes in Mathematics, 1604:154–169, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095521/.

Huber:1995:HR

[Hub95o] Annette Huber. Hodge realization. Lecture Notes in Mathematics, 1604:60–73, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095511/.

Huber:1995:KM

[Hub95p] Annette Huber. The Künneth morphism. Lecture Notes in Mathematics, 1604:112–122, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095517/.

Huber:1995:MMT

[Hub95q] Annette Huber. Mixed Motives and Their Realization in Derived Categories, volume 1604 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 207 pp. LCCN QA3 .L28 no. 1604.

URL http://link.springer.com/book/10.1007/BFb0095503;
 http://www.springerlink.com/content/978-3-540-49274-
0.

Huber:1995:MR

[Hub95r] Annette Huber. The mixed realization. Lecture Notes in Mathematics, 1604:86–95, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095514/.

Huber:1995:OC

[Hub95s] Annette Huber. Operations of correspondences. Lecture Notes in Mathematics, 1604:172–177, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095522/.

Huber:1995:SC

[Hub95t] Annette Huber. Singular cohomology. Lecture Notes in Mathematics, 1604:50–56, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095509/.

Huber:1995:TT

[Hub95u] Annette Huber. The Tate twist. Lecture Notes in Mathematics, 1604:98–102, 1995. CODEN LNMAA2. ISBN 3-540-59475-2 (print), 3-540-49274-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0095515/.

Huck:1990:EAC

[Huc90] Günther Huck. Embeddings of acyclic 2-complexes in S^{4} with contractible complement. Lecture Notes in Mathematics, 1440: 122–129, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084457/.

Huisman:1993:HAV

[Hui93] Johan Huisman. Heights on Abelian varieties. Lecture Notes in Mathematics, 1566:51–61, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48208-6_5/.

Hulek:1990:ASP

[Hul90] Klaus Hulek. Abelian surfaces in products of projective spaces. Lecture Notes in Mathematics, 1417:129–137, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083338/.

Hummel:1990:NSS

[Hum90] J. A. Hummel. Numerical solutions of the Schiffer equation. Lecture Notes in Mathematics, 1435:71–79, 1990. CODEN LN-MAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087898/.

Hunt:1996:LCS

[Hun96a] Bruce Hunt. The 27 lines on a cubic surface. Lecture Notes in Mathematics, 1637:108–167, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094404/.

Hunt:1996:AQ

[Hun96b] Bruce Hunt. Arithmetic quotients. Lecture Notes in Mathematics, 1637:36–65, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094402/.

Hunt:1996:BQ

[Hun96c] Bruce Hunt. The Burkhardt quartic. Lecture Notes in Mathematics, 1637:168–221, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094405/.

Hunt:1996:GMU

[Hun96d] Bruce Hunt. A gem of the modular universe. Lecture Notes in Mathematics, 1637:222–254, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094406/.

Hunt:1996:GSS

[Hun96e] Bruce Hunt. The Geometry of some special Arithmetic Quotients, volume 1637 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 332 pp. LCCN QA3 .L28 no. 1637. URL http:/

/link.springer.com/book/10.1007/BFb0094399;   http://
www.springerlink.com/content/978-3-540-69997-2.

Hunt:1996:I

[Hun96f] Bruce Hunt. Introduction. Lecture Notes in Mathematics, 1637: 1–14, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094400/.

Hunt:1996:MSP

[Hun96g] Bruce Hunt. Moduli spaces of PEL structures. Lecture Notes in Mathematics, 1637:15–35, 1996. CODEN LNMAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094401/.

Hunt:1996:PEM

[Hun96h] Bruce Hunt. Projective embeddings of modular varieties. Lecture Notes in Mathematics, 1637:66–107, 1996. CODEN LN-MAA2. ISBN 3-540-61795-7 (print), 3-540-69997-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094403/.

Hannsgen:1991:MCV

[HW91] Kenneth B. Hannsgen and Robert L. Wheeler. Moment conditions for a Volterra integral equation in a Banach space. Lecture Notes in Mathematics, 1475:204–209, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083492/.

He:1997:HOU

[HW97] Sheng-Wu He and Jia-Gang Wang. The hypercontractivity of Ornstein–Uhlenbeck semigroups with drift, revisited. Lecture Notes in Mathematics, 1655:80–84, 1997. CODEN LN-MAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119295/.

Harmand:1993:BSWb

[HWW93a] Peter Harmand, Dirk Werner, and Wend Werner. Banach spaces which are L-summands in their biduals. Lecture Notes in Mathematics, 1547:157–214, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084359/.

Harmand:1993:BSWa

[HWW93b] Peter Harmand, Dirk Werner, and Wend Werner. Banach spaces which are M-ideals in their biduals. Lecture Notes in Mathematics, 1547:101–155, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084358/.

Harmand:1993:BTI

[HWW93c] Peter Harmand, Dirk Werner, and Wend Werner. Basic theory of M-ideals. Lecture Notes in Mathematics, 1547:1–47, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084356/.

Harmand:1993:GPI

[HWW93d] Peter Harmand, Dirk Werner, and Wend Werner. Geometric properties of M-ideals. Lecture Notes in Mathematics, 1547:49– 100, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084357/.

Harmand:1993:IBA

[HWW93e] Peter Harmand, Dirk Werner, and Wend Werner. M-ideals in Banach algebras. Lecture Notes in Mathematics, 1547:215–261, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084360/.

Harmand:1993:IBS

[HWW93f] Peter Harmand, Dirk Werner, and Wend Werner. M-Ideals in Banach Spaces and Banach Algebras, volume 1547 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CO-DEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 344 (est.) pp. LCCN QA3 .L28 no. 1547. URL http:/

/link.springer.com/book/10.1007/BFb0084355;   http://
www.springerlink.com/content/978-3-540-47753-2.

Harmand:1993:ISB

[HWW93g] Peter Harmand, Dirk Werner, and Wend Werner. M-ideals in spaces of bounded operators. Lecture Notes in Mathematics, 1547: 263–344, 1993. CODEN LNMAA2. ISBN 3-540-56814-X (print), 3-540-47753-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084361/.

Hawkins:1991:TNZ

[HYL91a] W. G. Hawkins, N.-C. Yang, and P. K. Leichner. Theorems for the number of zeros of the projection radial modulators of the 2D exponential Radon transform. Lecture Notes in Mathematics, 1497:194–214, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084519/.

Hyland:1991:FSS

[Hyl91b] J. M. E. Hyland. First steps in synthetic domain theory. Lecture Notes in Mathematics, 1488:131–156, 1991. CODEN LN-MAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084217/.

Iorio:1990:KBF

[I90] Rafael José Iório, Jr. KDV, BO and friends in weighted Sobolev spaces. Lecture Notes in Mathematics, 1450:104–121, 1990. CO-DEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084901/.

Ichinose:1990:FPI

[Ich90] Takashi Ichinose. Feynman path integral to relativistic quantum mechanics. Lecture Notes in Mathematics, 1450:196–209, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084907/.

Ingold:1990:SDP

[IG90] Gert-Ludwig Ingold and Hermann Grabert. Sluggish decay of preparation effects in low temperature quantum systems. Lecture Notes in Mathematics, 1442:219–230, 1990. CODEN LN-MAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085515/.

Ishibashi:1990:SAF

[IK90] M. Ishibashi and S. Kanemitsu. Some asymptotic formulas of Ramanujan. Lecture Notes in Mathematics, 1434:149–167, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097132/.

Iarrobino:1999:CVP

[IK99a] Anthony Iarrobino and Vassil Kanev. Closures of the variety Gor(T), and the parameter space G(T) of graded algebras. Lecture Notes in Mathematics, 1721:249–253, 1999. CODEN LN-MAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093434/.

Iarrobino:1999:CCD

[IK99b] Anthony Iarrobino and Vassil Kanev. Connectedness and components of the determinantal locus PV_{s}(u, v; r). Lecture Notes in Mathematics, 1721:237–247, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093433/.

Iarrobino:1999:FCM

[IK99c] Anthony Iarrobino and Vassil Kanev. Forms and catalecticant matrices. Lecture Notes in Mathematics, 1721:3–56, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093427/.

Iarrobino:1999:FZDa

[IK99d] Anthony Iarrobino and Vassil Kanev. Forms and zero-dimensional schemes I: Basic results, and the case r = 3. Lecture Notes in Mathematics, 1721:131–205, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093431/.

Iarrobino:1999:FZDb

[IK99e] Anthony Iarrobino and Vassil Kanev. Forms and zero-dimensional schemes, II: Annihilating schemes and reducible Gor(T). Lecture Notes in Mathematics, 1721:207–236, 1999. CODEN LN-MAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093432/.

Iarrobino:1999:LSP

[IK99f] Anthony Iarrobino and Vassil Kanev. The locus PS(s, j; r) of sums of powers, and determinantal loci of catalecticant matrices. Lecture Notes in Mathematics, 1721:91–127, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093430/.

Iarrobino:1999:PSG

[IK99g] Anthony Iarrobino and Vassil Kanev. Power Sums, Gorenstein Algebras, and Determinantal Loci, volume 1721 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxxii + 345 pp. LCCN QA3 .L28 no. 1721. URL http:/

/link.springer.com/book/10.1007/BFb0093426;   http://
www.springerlink.com/content/978-3-540-46707-6.

Iarrobino:1999:QP

[IK99h] Anthony Iarrobino and Vassil Kanev. Questions and problems. Lecture Notes in Mathematics, 1721:255–264, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093435/.

Iarrobino:1999:SPL

[IK99i] Anthony Iarrobino and Vassil Kanev. Sums of powers of linear forms, and Gorenstein algebras. Lecture Notes in Mathematics, 1721:57–72, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093428/.

Iarrobino:1999:TSC

[IK99j] Anthony Iarrobino and Vassil Kanev. Tangent spaces to catalecticant schemes. Lecture Notes in Mathematics, 1721:73–90, 1999. CODEN LNMAA2. ISBN 3-540-66766-0 (print), 3-540-46707-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093429/.

Ikawa:1990:DPS

[Ika90] Mitsuru Ikawa. On the distribution of poles of the scattering matrix for several convex bodies. Lecture Notes in Mathematics, 1450:210–225, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084908/.

Ikeda:1990:PMS

[Ike90] N. Ikeda. Probabilistic methods in the study of asymptotics. Lecture Notes in Mathematics, 1427:195–325, 1990. CODEN LN-MAA2. ISBN 3-540-53508-X (print), 3-540-46718-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103043/.

Maruyama:1990:L

[iM90a] Ken ichi Maruyama. Localizing ϵ_{#}(X). Lecture Notes in Mathematics, 1425:87–90, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083833/.

Mizumoto:1990:ICV

[iM90b] Shin ichiro Mizumoto. Integrality of critical values of triple product L-functions. Lecture Notes in Mathematics, 1434:188– 195, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097135/.

Inassaridze:1990:EHA

[Ina90a] Hvedri Inassaridze, editor. K-theory and Homological Algebra: A Seminar held at the Razmadze Mathematical Institute in Tbilisi, Georgia, USSR 1987–88, volume 1437 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1437. URL http:/

/link.springer.com/book/10.1007/BFb0086716;   http://
www.springerlink.com/content/978-3-540-47162-2.

Inassaridze:1990:TSN

[Ina90b] Hvedri Inassaridze. K-theory of special normed rings. Lecture Notes in Mathematics, 1437:95–156, 1990. CODEN LN-MAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086719/.

Inoue:1992:TST

[Ino92] Atsushi Inoue. A tiny step towards a theory of functional derivative equations — a strong solution of the space-time Hopf equation. Lecture Notes in Mathematics, 1530:246–261, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090346/.

Inoue:1998:FA

[Ino98a] Atsushi Inoue. Fundamentals of O^{∗}-algebras. Lecture Notes in Mathematics, 1699:7–40, 1998. CODEN LNMAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093331/.

Inoue:1998:I

[Ino98b] Atsushi Inoue. Introduction. Lecture Notes in Mathematics, 1699: 1–5, 1998. CODEN LNMAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093330/.

Inoue:1998:PA

[Ino98c] Atsushi Inoue. Physical applications. Lecture Notes in Mathematics, 1699:169–223, 1998. CODEN LNMAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093334/.

Inoue:1998:SSM

[Ino98d] Atsushi Inoue. Standard systems and modular systems. Lecture Notes in Mathematics, 1699:41–110, 1998. CODEN LN-MAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093332/.

Inoue:1998:SWA

[Ino98e] Atsushi Inoue. Standard weights on O^{∗}-algebras. Lecture Notes in Mathematics, 1699:111–168, 1998. CODEN LNMAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093333/.

Inoue:1998:TTT

[Ino98f] Atsushi Inoue. Tomita–Takesaki Theory in Algebras of Unbounded Operators, volume 1699 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-65194-2 (print), 3-540-49495-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 241 pp. LCCN QA3 .L28 no. 1699.

URL http://link.springer.com/book/10.1007/BFb0093329;
 http://www.springerlink.com/content/978-3-540-49495-
9.

Ionescu:1990:EPV

[Ion90] Paltin Ionescu. Embedded projective varieties of small invariants. III. Lecture Notes in Mathematics, 1417:138–154, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083339/.

Iluis-Puebla:1992:IAT

[IP92] Emilio Iluis-Puebla. Introduction to algebraic K-theory. Lecture Notes in Mathematics, 1491:1–30, 1992. CODEN LN-MAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088877/.

Izydorek:1992:PBU

[IR92] Marek Izydorek and Sławomir Rybicki. On parametrized Borsuk– Ulam theorem for free Z_{p}-action. Lecture Notes in Mathematics, 1509:227–234, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087512/.

Ivanov:1990:BLI

[IS90] K. G. Ivanov and E. B. Saff. Behavior of the Lagrange interpolants in the roots of unity. Lecture Notes in Mathematics, 1435: 81–87, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087899/.

Isac:1992:Ea

[Isa92a] G. Isac. Equivalences. Lecture Notes in Mathematics, 1528:52– 69, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084657/.

Isac:1992:Eb

[Isa92b] G. Isac. Errata. Lecture Notes in Mathematics, 1528:e1–e2, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0120443.

Isac:1992:ET

[Isa92c] G. Isac. Existence theorems. Lecture Notes in Mathematics, 1528: 70–138, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084658/.

Isac:1992:ICP

[Isa92d] G. Isac. The implicit complementarity problem. Lecture Notes in Mathematics, 1528:162–195, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084660/.

Isac:1992:I

[Isa92e] G. Isac. Introduction. Lecture Notes in Mathematics, 1528:1– 3, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084654/.

Isac:1992:IPC

[Isa92f] G. Isac. Isotone projection cones and complementarity. Lecture Notes in Mathematics, 1528:196–219, 1992. CODEN LN-MAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084661/.

Isac:1992:MA

[Isa92g] G. Isac. Models and applications. Lecture Notes in Mathematics, 1528:16–51, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084656/.

Isac:1992:OCP

[Isa92h] G. Isac. The order complementarity problem. Lecture Notes in Mathematics, 1528:139–161, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084659/.

Isac:1992:PDP

[Isa92i] G. Isac. Preliminaries and definitions of principal complementarity problems. Lecture Notes in Mathematics, 1528:4–15, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084655/.

Isac:1992:TCP

[Isa92j] G. Isac. Topics on complementarity problems. Lecture Notes in Mathematics, 1528:220–269, 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084662/.

Isac:1992:CP

[Isa92k] George Isac. Complementarity Problems, volume 1528 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56251-6 (print), 3-540-47491-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 297 pp. LCCN QA3 .L28 no. 1528. URL http:/

/link.springer.com/book/10.1007/BFb0084653;   http://
www.springerlink.com/content/978-3-540-47491-3.

Ishiguro:1992:CSC

[Ish92] Kenshi Ishiguro. Classifying spaces of compact simple Lie groups and p-tori. Lecture Notes in Mathematics, 1509:210–226, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087511/.

Iskovskih:1991:TPR

[Isk91] V. A. Iskovskih. Towards the problem of rationality of conic bundles. Lecture Notes in Mathematics, 1479:50–56, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086263/.

Imrich:1990:FSH

[IT90] W. Imrich and E. C. Turner. Fixed subsets of homomorphisms of free groups. Lecture Notes in Mathematics, 1440:130–147, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084458/.

Ito:1993:SPT

[Itō93a] Kiyosi Itō. Semigroups in probability theory. Lecture Notes in Mathematics, 1540:69–83, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085475/.

Iwamiya:1993:CNP

[ITO93b] Toshiyuki Iwamiya, Tadayasu Takahashi, and Shinnosuke Oharu. Characterization of nonlinearly perturbed semigroups. Lecture Notes in Mathematics, 1540:85–102, 1993. CODEN LN-MAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085476/.

Ivanov:1993:PAS

[Iva93] Oleg V. Ivanov. A problem of axler and shields on nontangential limits and maximal ideal space of some pseudonanalytic algebras. Lecture Notes in Mathematics, 1550:157–159, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117481/.

Iwatsuka:1990:SOM

[Iwa90] Akira Iwatsuka. On Schrödinger operators with magnetic fields. Lecture Notes in Mathematics, 1450:157–172, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084904/.

Iwaniec:1992:TQM

[Iwa92a] Tadeusz Iwaniec. L^{p}-theory of quasiregular mappings. Lecture Notes in Mathematics, 1508:39–64, 1992. CODEN LN-MAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094237/.

Iwanik:1992:PEI

[Iwa92b] Anzelm Iwanik. Positive entropy implies infinite L^{p}-multiplicity for p > 1. Lecture Notes in Mathematics, 1514:124–127, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097532/.

Yosimura:1990:QKH

[iY90] Zen ichi Yosimura. The quasi ko-homology types of the real projective spaces. Lecture Notes in Mathematics, 1418:156–174, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083701/.

Jacobsen:1990:OPC

[Jac90] Lisa Jacobsen. Orthogonal polynomials, chain sequences, threeterm recurrence relations and continued fractions. Lecture Notes in Mathematics, 1435:89–101, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087900/.

Jacod:1991:ROQ

[Jac91] Jean Jacod. Régularité d’ordre quelconque pour un modèle statistique filtré. (French) []. Lecture Notes in Mathematics, 1485:140– 161, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100854/.

Jagers:1993:SIP

[Jag93] Peter Jagers. Stabilities and instabilities in population dynamics. Lecture Notes in Mathematics, 1546:58–67, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084482/.

Jajte:1990:ASC

[Jaj90] Ryszard Jajte. Almost sure convergence of iterates of contractions in noncommutative L_{2}-spaces. Lecture Notes in Mathematics, 1442:231–246, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085516/.

Jajte:1991:ASC

[Jaj91a] Ryszard Jajte. Almost sure convergence in noncommutative L_{2}- spaces. Lecture Notes in Mathematics, 1477:1–9, 1991. CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098425/.

Jajte:1991:AF

[Jaj91b] Ryszard Jajte. Asymptotic formulae. Lecture Notes in Mathematics, 1477:37–51, 1991. CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098427/.

Jajte:1991:CCE

[Jaj91c] Ryszard Jajte. Convergence of conditional expectations and martingales. Lecture Notes in Mathematics, 1477:85–89, 1991. CO-DEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098430/.

Jajte:1991:CIC

[Jaj91d] Ryszard Jajte. Convergence of iterates of contractions. Lecture Notes in Mathematics, 1477:52–63, 1991. CODEN LN-MAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098428/.

Jajte:1991:COS

[Jaj91e] Ryszard Jajte. Convergence of orthogonal series and strong laws of large numbers. Lecture Notes in Mathematics, 1477:64– 84, 1991. CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098429/.

Jajte:1991:IET

[Jaj91f] Ryszard Jajte. Individual ergodic theorems in L_{2} over a von Neumann algebra. Lecture Notes in Mathematics, 1477:10–36, 1991.

CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098426/.

Jajte:1991:MR

[Jaj91g] Ryszard Jajte. Miscellaneous results. Lecture Notes in Mathematics, 1477:90–99, 1991. CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098431/.

Jajte:1991:SLT

[Jaj91h] Ryszard Jajte. Strong Limit Theorems in Noncommutative L_{2}- Spaces, volume 1477 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54214-0 (print), 3-540-47512-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 113 pp. LCCN QA3 .L28 no. 1477. URL http:

//link.springer.com/book/10.1007/BFb0098424;   http:/
/www.springerlink.com/content/978-3-540-47512-5.

James:1990:CFS

[Jam90] I. M. James. Continuous functions of several variable. Lecture Notes in Mathematics, 1418:53–56, 1990. CODEN LN-MAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083691/.

Janelidze:1990:CEI

[Jan90a] George Janelidze. Cohomology and extensions of internal modules. Lecture Notes in Mathematics, 1437:157–168, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086720/.

Jannsen:1990:ACT

[Jan90b] Uwe Jannsen. Algebraic cycles, K-theory, and extension classes. Lecture Notes in Mathematics, 1400:57–188, 1990. CODEN LN-MAA2. ISBN 3-540-52260-3 (print), 3-540-46941-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085082/.

Jannsen:1990:TAC

[Jan90c] Uwe Jannsen. K-theory and ℓ-adic cohomology. Lecture Notes in Mathematics, 1400:189–221, 1990. CODEN LNMAA2. ISBN 3-540-52260-3 (print), 3-540-46941-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085083/.

Jannsen:1990:MMAb

[Jan90d] Uwe Jannsen. Mixed Motives and Algebraic K-Theory, volume 1400 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52260-3 (print), 3-540-46941-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 246 pp. LCCN QA3 .L28 no. 1400. URL http:/

/link.springer.com/book/10.1007/BFb0085080;   http://
www.springerlink.com/content/978-3-540-46941-4.

Jannsen:1990:MMAa

[Jan90e] Uwe Jannsen. Mixed motives for absolute Hodge cycles. Lecture Notes in Mathematics, 1400:1–56, 1990. CODEN LN-MAA2. ISBN 3-540-52260-3 (print), 3-540-46941-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085081/.

Janelidze:1991:PGT

[Jan91] George Janelidze. Precategories and Galois theory. Lecture Notes in Mathematics, 1488:157–173, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084218/.

Jankovic:1993:SPR

[Jan93] Slobodanka Janković. Some properties of random variables which are stable with respect to the random sample size. Lecture Notes in Mathematics, 1546:68–75, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084483/.

Jaworowski:1991:ABC

[Jaw91] Jan Jaworowski. An additive basis for the cohomology of real grassmannians. Lecture Notes in Mathematics, 1474:231–234, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084749/.

Jay:1991:FLC

[Jay91] C. Barry Jay. Fixpoint and loop constructions as colimits. Lecture Notes in Mathematics, 1488:187–192, 1991. CODEN LN-MAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084220/.

Jie-cheng:1991:DBP

[JcC91] Chen Jie-cheng and Luo Cheng. Duality of H^{1} and BMO on positively curved manifolds and their characterizations. Lecture Notes in Mathematics, 1494:23–38, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087753/.

Jensen:1990:HEA

[Jen90] Arne Jensen. High energy asymptotics for the total scattering phase in potential scattering theory. Lecture Notes in Mathematics, 1450:187–195, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084906/.

Jibladze:1990:CCL

[Jib90] M. Jibladze. Coefficients for cohomology of ‘large’ categories. Lecture Notes in Mathematics, 1437:169–179, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086721/.

Jain:1992:BCB

[JJS92] P. K. Jain, A. M. Jarrah, and D. P. Sinha. Boundedly complete bases in various locally convex spaces. Lecture Notes in Mathematics, 1511:128–140, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093806/.

Jiang:1991:NLG

[JL91] Yaping Jiang and Xuebo Luo. The nilpotent Lie group G^{d+2} and a class of differential operators with multiple characteristics. Lecture Notes in Mathematics, 1494:82–83, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087759.

Jorgenson:1993:BAR

[JL93a] Jay A. Jorgenson and Serge Lang. Basic Analysis of Regularized Series and Products, volume 1564 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57488-3 (print), 3-540-48193-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 117 (est.) pp. LCCN QA3 .L28 no. 1564.

URL http://link.springer.com/book/10.1007/BFb0077194;
 http://www.springerlink.com/content/978-3-540-48193-
5.

Jorgenson:1993:PFF

[JL93b] Jay A. Jorgenson and Serge Lang. A Parseval formula for functions with a singular asymptotic expansion at the origin. Lecture Notes in Mathematics, 1564:89–117, 1993. CODEN LN-MAA2. ISBN 3-540-57488-3 (print), 3-540-48193-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0077196/.

Jorgenson:1993:SCA

[JL93c] Jay A. Jorgenson and Serge Lang. Some complex analytic properties of regularized products and series. Lecture Notes in Mathematics, 1564:ix–88, 1993. CODEN LNMAA2. ISBN 3-540-57488-3 (print), 3-540-48193-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0077195/.

Jorgenson:1994:EFRa

[JL94] Jay Jorgenson and Serge Lang. Explicit formulas for regularized products and series. Lecture Notes in Mathematics, 1593: 1–134, 1994. CODEN LNMAA2. ISBN 3-540-58673-3 (print), 3-540-49041-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074040/.

Jorgenson:1994:EFRb

[JLG94] Jay Jorgenson, Serge Lang, and Dorian Goldfeld. Explicit Formulas for Regularized Products and Series, volume 1593 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58673-3 (print), 3-540-49041-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 154 pp. LCCN QA3 .L28 no. 1593. URL http:/

/link.springer.com/book/10.1007/BFb0074039;   http://
www.springerlink.com/content/978-3-540-49041-8.

Jain:1990:NCR

[JLP90] Surender Kumar Jain and Sergio R. López-Permouth, editors. Non-Commutative Ring Theory: Proceedings of a Conference held in Athens, Ohio Sept. 29–30, 1989, volume 1448 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1448.

URL http://link.springer.com/book/10.1007/BFb0091244;
 http://www.springerlink.com/content/978-3-540-46745-
8.

Jagers:1996:ACS

[JN96] Peter Jagers and Olle Nerman. The asymptotic composition of supercritical, multi-type branching populations. Lecture Notes in Mathematics, 1626:40–54, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094640/.

Johnsen:1990:LMT

[Joh90] Trygve Johnsen. Local multiplicities of tangential trisecants to space curves. Lecture Notes in Mathematics, 1436:77–100, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084042/.

Johannson:1995:GOR

[Joh95a] Klaus Johannson. Generalized one-relator 3-manifolds. Lecture Notes in Mathematics, 1599:147–245, 1995. CODEN LN-MAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074008/.

Johannson:1995:H

[Joh95b] Klaus Johannson. Handlebodies. Lecture Notes in Mathematics, 1599:1–36, 1995. CODEN LNMAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074006/.

Johannson:1995:RM

[Joh95c] Klaus Johannson. N-relation 3-manifolds. Lecture Notes in Mathematics, 1599:246–282, 1995. CODEN LNMAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074009/.

Johannson:1995:RH

[Joh95d] Klaus Johannson. Relative handlebodies. Lecture Notes in Mathematics, 1599:37–146, 1995. CODEN LNMAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074007/.

Johannson:1995:SHG

[Joh95e] Klaus Johannson. The space of Heegaard graphs. Lecture Notes in Mathematics, 1599:283–426, 1995. CODEN LNMAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074010/.

Johannson:1995:TCM

[Joh95f] Klaus Johannson. Topology and Combinatorics of 3-Manifolds, volume 1599 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-59063-3 (print), 3-540-49181-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 446 pp. LCCN QA3 .L28 no. 1599. URL http:/

/link.springer.com/book/10.1007/BFb0074005;   http://
www.springerlink.com/content/978-3-540-49181-1.

Johnson:1995:DSL

[Joh95g] Russell Johnson, editor. Dynamical Systems: Lectures Given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, June 13–22, 1994, volume 1609 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1609. URL http:/

/link.springer.com/book/10.1007/BFb0095237;   http://
www.springerlink.com/content/978-3-540-49415-7.

Johnson:1998:AFE

[Joh98] Claes Johnson. Adaptive finite element methods for conservation laws. Lecture Notes in Mathematics, 1697:269–323, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096354/.

Jones:1992:KMP

[Jon92] Vaughan F. R. Jones. Knots in mathematics and physics. Lecture Notes in Mathematics, 1525:70–77, 1992. CODEN LN-MAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089206/.

Jones:1995:GSP

[Jon95] Christopher K. R. T. Jones. Geometric singular perturbation theory. Lecture Notes in Mathematics, 1609:44–118, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095239/.

Jackowski:1991:ATP

[JOP91] Stefan Jackowski, Bob Oliver, and Krzystof Pawałowski, editors. Algebraic Topology Poznań 1989: Proceedings of a Conference held in Poznań, Poland, June 22–27, 1989, volume 1474 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1474.

URL http://link.springer.com/book/10.1007/BFb0084733;

 http://www.springerlink.com/content/978-3-540-47403-
6.

Jouini:1997:MIE

[Jou97] Elyès Jouini. Market imperfections, equilibrium and arbitrage. Lecture Notes in Mathematics, 1656:247–307, 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092002/.

Jourdain:1998:PTD

[Jou98] B. Jourdain. Propagation trajectorielle du chaos pour les lois de conservation scalaire. (French) []. Lecture Notes in Mathematics, 1686:215–230, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101759/.

Jacod:1991:RED

[JP91a] J. Jacod and P. Protter. Une remarque sur les équations différentielles stochastiques à solutions markoviennes. (French) []. Lecture Notes in Mathematics, 1485:138–139, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100853.

Jost:1991:GMS

[JP91b] Jürgen Jost and Xiao-Wei Peng. The geometry of moduli spaces of stable vector bundles over Riemann surfaces. Lecture Notes in Mathematics, 1481:79–96, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083631/.

Jacod:1997:MWF

[JPA97] Jean Jacod and Víctor Pérez-Abreu. On martingales which are finite sums of independent random variables with time dependent coefficients. Lecture Notes in Mathematics, 1655:62–68, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0119292/.

Janeczko:1991:CSC

[JR91a] Staszek Janeczko and Mark Roberts. Classification of symmetric caustics I: symplectic equivalence. Lecture Notes in Mathematics, 1463:193–219, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085432/.

Johnson:1991:FGA

[JR91b] F. E. A. Johnson and E. G. Rees. The fundamental groups of algebraic varieties. Lecture Notes in Mathematics, 1474:75–82, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084738/.

Jipsen:1992:ALV

[JR92a] Peter Jipsen and Henry Rose. Amalgamation in lattice varieties. Lecture Notes in Mathematics, 1533:128–148, 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090230/.

Jipsen:1992:EB

[JR92b] Peter Jipsen and Henry Rose. Equational bases. Lecture Notes in Mathematics, 1533:115–127, 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090229/.

Jipsen:1992:GR

[JR92c] Peter Jipsen and Henry Rose. General results. Lecture Notes in Mathematics, 1533:13–45, 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090226/.

Jipsen:1992:MV

[JR92d] Peter Jipsen and Henry Rose. Modular varieties. Lecture Notes in Mathematics, 1533:46–76, 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090227/.

Jipsen:1992:NV

[JR92e] Peter Jipsen and Henry Rose. Nonmodular varieties. Lecture Notes in Mathematics, 1533:77–114, 1992. CODEN LN-MAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090228/.

Jipsen:1992:P

[JR92f] Peter Jipsen and Henry Rose. Preliminaries. Lecture Notes in Mathematics, 1533:1–12, 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090225/.

Jipsen:1992:VL

[JR92g] Peter Jipsen and Henry Rose. Varieties of Lattices, volume 1533 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56314-8 (print), 3-540-47514-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 162 pp. LCCN QA3 .L28 no. 1533. URL http:/

/link.springer.com/book/10.1007/BFb0090224;   http://
www.springerlink.com/content/978-3-540-47514-9.

Jolis:1990:GMS

[JS90] Maria Jolis and Marta Sanz. On generalized multiple stochastic integrals and multiparameter anticipative calculus. Lecture Notes in Mathematics, 1444:141–182, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083614/.

Janeczko:1991:SSO

[JS91a] S. Janeczko and Ian Stewart. Symplectic singularities and optical diffraction. Lecture Notes in Mathematics, 1463:220–255, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085433/.

Johnson:1991:RTD

[JS91b] William B. Johnson and Gideon Schechtman. Remarks on Talagrand’s deviation inequality for Rademacher functions. Lecture Notes in Mathematics, 1470:72–77, 1991. CODEN LN-MAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090214/.

Jeanneret:1992:RTC

[JS92] Alain Jeanneret and Ulrich Suter. Réalisation topologique de certaines algèbres associées aux algèbres de Dickson. (French) []. Lecture Notes in Mathematics, 1509:235–239, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087513/.

Jacod:1994:JFM

[JS94] J. Jacod and A. V. Skorohod. Jumping filtrations and martingales with finite variation. Lecture Notes in Mathematics, 1583: 21–35, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073831/.

Janelidze:1991:HAC

[JT91a] George Janelidze and Walter Tholen. How algebraic is the change-of-base functor? Lecture Notes in Mathematics, 1488: 174–186, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084219/.

Joyal:1991:SSC

[JT91b] André Joyal and Myles Tierney. Strong stacks and classifying spaces. Lecture Notes in Mathematics, 1488:213–236, 1991. CO-DEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084222/.

Johnstone:1991:PPP

[JV91] Peter Johnstone and Steven Vickers. Preframe presentations present. Lecture Notes in Mathematics, 1488:193–212, 1991. CO-DEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084221/.

Jeulin:1990:FPB

[JY90] T. Jeulin and M. Yor. Filtration des ponts browniens et équations différentielles stochastiques linéaires. (French) []. Lecture Notes in Mathematics, 1426:227–265, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083768/.

Jost:1991:HMK

[JY91] Jürgen Jost and Shing-Tung Yau. Harmonic maps and Kähler geometry. Lecture Notes in Mathematics, 1468:340–370, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086200/.

Jeulin:1993:MMS

[JY93] T. Jeulin and M. Yor. Moyennes mobiles et semimartingales. (French) []. Lecture Notes in Mathematics, 1557:53–77, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087964/.

Kaczynski:1991:IDE

[Kac91] Tomasz Kaczynski. Implicit differential equations which are not solvable for the highest derivative. Lecture Notes in Mathematics, 1475:218–224, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083494/.

Kahn:1990:LPR

[Kah90a] Donald W. Kahn. List of papers on or relevant to groups of self-homotopy equivalences. Lecture Notes in Mathematics, 1425: 208–214, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083842/.

Kahn:1990:RSG

[Kah90b] Donald W. Kahn. Representations of the stable group of selfequivalences. Lecture Notes in Mathematics, 1425:58–70, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083831/.

Kahn:1990:SRP

[Kah90c] Donald W. Kahn. Some research problems on homotopy-self-equivalences. Lecture Notes in Mathematics, 1425:204–207, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083841/.

Kahane:1992:TP

[Kah92] Jean-Pierre Kahane. On a theorem of Pólya. Lecture Notes in Mathematics, 1511:106–112, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093803/.

Kahane:1994:PP

[Kah94] Jean-Pierre Kahane. Probabilistic problems. Lecture Notes in Mathematics, 1573:155–178, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100204/.

Kaiser:1997:DBT

[Kai97a] Uwe Kaiser. 3-dimensional betti-trivial submanifolds. Lecture Notes in Mathematics, 1669:117–131, 1997. CODEN LN-MAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092693/.

Kaiser:1997:ELB

[Kai97b] Uwe Kaiser. Enumeration of link bordism in 3-manifolds. Lecture Notes in Mathematics, 1669:21–36, 1997. CODEN LN-MAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092688/.

Kaiser:1997:LBM

[Kai97c] Uwe Kaiser. Link bordism in manifolds. Lecture Notes in Mathematics, 1669:1–20, 1997. CODEN LNMAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092687/.

Kaiser:1997:LCB

[Kai97d] Uwe Kaiser. Link characteristic and band-operations in Bettitrivial 3-manifolds. Lecture Notes in Mathematics, 1669:99–116, 1997. CODEN LNMAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092692/.

Kaiser:1997:LIB

[Kai97e] Uwe Kaiser. Link invariants in betti-trivial 3-manifolds. Lecture Notes in Mathematics, 1669:85–98, 1997. CODEN LN-MAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092691/.

Kaiser:1997:LTM

[Kai97f] Uwe Kaiser. Link Theory in Manifolds, volume 1669 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 167 pp. LCCN QA3 .L28 no. 1669. URL http:/

/link.springer.com/book/10.1007/BFb0092686;   http://
www.springerlink.com/content/978-3-540-69546-2.

Kaiser:1997:LNM

[Kai97g] Uwe Kaiser. Linking number maps. Lecture Notes in Mathematics, 1669:37–68, 1997. CODEN LNMAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092689/.

Kaiser:1997:SSL

[Kai97h] Uwe Kaiser. Surface structures for links in 3-manifolds. Lecture Notes in Mathematics, 1669:69–84, 1997. CODEN LN-MAA2. ISBN 3-540-63435-5 (print), 3-540-69546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092690/.

Kajitani:1991:FIO

[Kaj91] Kunihiko Kajitani. Fourier integral operators with complexvalued phase function and the Cauchy problem for hyperbolic operators. Lecture Notes in Mathematics, 1505:1–70, 1991. CODEN LNMAA2. ISBN 3-540-55018-6 (print), 3-540-46655-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090883/.

Kalashnikov:1993:TSE

[Kal93] V. V. Kalashnikov. Two-side estimates of geometric convolutions. Lecture Notes in Mathematics, 1546:76–88, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084484/.

Kallsen:1999:SDE

[Kal99] Jan Kallsen. A stochastic differential equation with a unique (up to indistinguishability) but not strong solution. Lecture Notes in Mathematics, 1709:315–326, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096520/.

Kandelaki:1990:TGB

[Kan90] Tamazi Kandelaki. K-theory of Z_{2}-graded Banach categories. I. Lecture Notes in Mathematics, 1437:180–221, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086722/.

Kaneyuki:1991:CSS

[Kan91] Soji Kaneyuki. On the causal structures of the Silov boundaries of symmetric bounded domains. Lecture Notes in Mathematics, 1468:127–159, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086191/.

Kapranov:1991:DMR

[Kap91] M. M. Kapranov. On DG-modules over the de Rham complex and the vanishing cycles functor. Lecture Notes in Mathematics, 1479:57–86, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086264/.

Karandikar:1991:ASC

[Kar91] Rajeeva L. Karandikar. On almost sure convergence of modified Euler–Peano approximation of solution to an S.D.E. driven by a semimartingale. Lecture Notes in Mathematics, 1485:113– 120, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100851/.

Karpeshina:1997:IFW

[Kar97a] Yulia E. Karpeshina. The interaction of a free wave with a semi-bounded crystal. Lecture Notes in Mathematics, 1663:233– 338, 1997. CODEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094269/.

Karpeshina:1997:I

[Kar97b] Yulia E. Karpeshina. Introduction. Lecture Notes in Mathematics, 1663:1–22, 1997. CODEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094265/.

Karpeshina:1997:PTPa

[Kar97c] Yulia E. Karpeshina. Perturbation theory for a polyharmonic operator in the case of 2l > n. Lecture Notes in Mathematics, 1663:23–62, 1997. CODEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094266/.

Karpeshina:1997:PTSa

[Kar97d] Yulia E. Karpeshina. Perturbation theory for Schrödinger operator with a periodic potential. Lecture Notes in Mathematics, 1663:99–232, 1997. CODEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094268/.

Karpeshina:1997:PTPb

[Kar97e] Yulia E. Karpeshina. Perturbation theory for the polyharmonic operator in the case 4l > n + 1. Lecture Notes in Mathematics, 1663:63–97, 1997. CODEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094267/.

Karpeshina:1997:PTSb

[Kar97f] Yulia E. Karpeshina. Perturbation Theory for the Schrödinger Operator with a Periodic Potential, volume 1663 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CO-DEN LNMAA2. ISBN 3-540-63136-4 (print), 3-540-69156-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 352 pp. LCCN QA3 .L28 no. 1663. URL http:/

/link.springer.com/book/10.1007/BFb0094264;   http://
www.springerlink.com/content/978-3-540-69156-3.

Kato:1990:LFM

[Kat90] Tosio Kato. Liapunov functions and monotonicity in the Navier– Stokes equation. Lecture Notes in Mathematics, 1450:53–63, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084898/.

Kato:1993:LAI

[Kat93a] Kazuya Kato. Lectures on the approach to Iwasawa theory for Hasse–Weil L-functions via B_{dR}. Part I. Lecture Notes in Mathematics, 1553:50–163, 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084729/.

Kato:1993:AEE

[Kat93b] Tosio Kato. Abstract evolution equations, linear and quasilinear, revisited. Lecture Notes in Mathematics, 1540:103–125, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085477/.

Kawamata:1992:MDA

[Kaw92] Yujiro Kawamata. Moderate degenerations of algebraic surfaces. Lecture Notes in Mathematics, 1507:113–132, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094514/.

Kawahigashi:1993:ESO

[Kaw93] Yasuyuki Kawahigashi. Exactly solvable orbifold models and subfactors. Lecture Notes in Mathematics, 1540:127–147, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085478/.

Kazhikhov:1990:BVP

[Kaz90] A. V. Kazhikhov. Boundary-value problems for Navier–Stokes equations of viscous gas. Lecture Notes in Mathematics, 1431: 169–172, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086067/.

Kazamaki:1994:BM

[Kaz94a] Norihiko Kazamaki. BMO-martingales. Lecture Notes in Mathematics, 1579:25–52, 1994. CODEN LNMAA2. ISBN 3-540-58042-5 (print), 3-540-48421-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073587/.

Kazamaki:1994:CEM

[Kaz94b] Norihiko Kazamaki. Continuous Exponential Martingales and BMO, volume 1579 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58042-5 (print), 3-540-48421-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 91 pp. LCCN QA3 .L28 no. 1579. URL http:/

/link.springer.com/book/10.1007/BFb0073585;   http://
www.springerlink.com/content/978-3-540-48421-9.

Kazamaki:1994:EM

[Kaz94c] Norihiko Kazamaki. Exponential martingales. Lecture Notes in Mathematics, 1579:1–24, 1994. CODEN LNMAA2. ISBN 3-540-58042-5 (print), 3-540-48421-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073586/.

Kazamaki:1994:EB

[Kaz94d] Norihiko Kazamaki. Exponential of BMO. Lecture Notes in Mathematics, 1579:53–84, 1994. CODEN LNMAA2. ISBN 3-540-58042-5 (print), 3-540-48421-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073588/.

Kushkuley:1996:DTEb

[KB96a] Alexander Kushkuley and Zalman Balanov. Degree theory for equivariant maps of finite-dimensional manifolds: Smooth actions. Lecture Notes in Mathematics, 1632:43–73, 1996. CODEN LNMAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092826/.

Kushkuley:1996:DTEa

[KB96b] Alexander Kushkuley and Zalman Balanov. Degree theory for equivariant maps of finite-dimensional manifolds: Topological actions. Lecture Notes in Mathematics, 1632:31–42, 1996. CODEN LNMAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092825/.

Kushkuley:1996:FDE

[KB96c] Alexander Kushkuley and Zalman Balanov. Fundamental domains and extension of equivariant maps. Lecture Notes in Mathematics, 1632:13–30, 1996. CODEN LNMAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092824/.

Kushkuley:1996:GMD

[KB96d] Alexander Kushkuley and Zalman Balanov. Geometric Methods in Degree Theory for Equivariant Maps, volume 1632 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 136 pp. LCCN QA3 .L28 no. 1632. URL http:/

/link.springer.com/book/10.1007/BFb0092822;   http://
www.springerlink.com/content/978-3-540-68726-9.

Kushkuley:1996:I

[KB96e] Alexander Kushkuley and Zalman Balanov. Introduction. Lecture Notes in Mathematics, 1632:1–12, 1996. CODEN LN-MAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092823/.

Kushkuley:1996:SA

[KB96f] Alexander Kushkuley and Zalman Balanov. Some applications. Lecture Notes in Mathematics, 1632:86–125, 1996. CODEN LN-MAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092828/.

Kushkuley:1996:WNE

[KB96g] Alexander Kushkuley and Zalman Balanov. A winding number of equivariant vector fields in infinite dimensional Banach spaces. Lecture Notes in Mathematics, 1632:74–85, 1996. CODEN LNMAA2. ISBN 3-540-61529-6 (print), 3-540-68726-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092827/.

Keller:1991:LEC

[Kel91a] Gerhard Keller. Lyapunov exponents and complexity for interval maps. Lecture Notes in Mathematics, 1486:216–226, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086671/.

Kelly:1991:NRR

[Kel91b] G. M. Kelly. A note on relations relative to a factorization system. Lecture Notes in Mathematics, 1488:249–261, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084224/.

Keller:1998:CTE

[Kel98] Bernhard Keller. On the construction of triangle equivalences. Lecture Notes in Mathematics, 1685:155–176, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096374/.

Kempf:1991:EMI

[Kem91a] George R. Kempf. Effective methods in invariant theory. Lecture Notes in Mathematics, 1479:90–93, 1991. CODEN LN-MAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086266/.

Kempf:1991:MCI

[Kem91b] George R. Kempf. More on computing invariants. Lecture Notes in Mathematics, 1479:87–89, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086265/.

Kenig:1993:PTN

[Ken93] Carlos E. Kenig. Potential theory of non-divergence form elliptic equations. Lecture Notes in Mathematics, 1563:89–128, 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074092/.

Kenmochi:1994:SNP

[Ken94] Nobuyuki Kenmochi. Systems of nonlinear PDEs arising from dynamical phase transitions. Lecture Notes in Mathematics, 1584: 39–86, 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073395/.

Komura:1993:WEN

[KF93] Yukio Kōmura and Kiyoko Furuya. Wave equations in nonreflexive spaces. Lecture Notes in Mathematics, 1540:235–238, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085482/.

Khimshiashvili:1992:FAL

[Khi92] G. N. Khimshiashvili. On Fredholmian aspects of linear transmission problems. Lecture Notes in Mathematics, 1520:193–216, 1992.

CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084722/.

Khokhlov:1993:FLT

[Kho93] Yu. S. Khokhlov. The functional limit theorem on nilpotent Lie group. Lecture Notes in Mathematics, 1546:163–166, 1993. CO-DEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084491/.

Khoshnevisan:1997:SPS

[Kho97] Davar Khoshnevisan. Some polar sets for the Brownian sheet. Lecture Notes in Mathematics, 1655:190–197, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119303/.

Khrushchev:1993:POP

[Khr93] S. Khrushchëv. Parameters of orthogonal polynomials. Lecture Notes in Mathematics, 1550:185–191, 1993. CODEN LN-MAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117487/.

Khuri:1990:MRP

[Khu90] Soumaya Makdissi Khuri. Modules with regular, perfect, Noetherian or Artinian endomorphism rings. Lecture Notes in Mathematics, 1448:7–18, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091246/.

Kifer:1991:LDR

[Kif91] Yuri Kifer. Large deviations for random expanding maps. Lecture Notes in Mathematics, 1486:178–186, 1991. CODEN LN-MAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086667/.

Kiming:1994:EVA

[Kim94] Ian Kiming. On the experimental verification of the Artin conjecture for 2-dimensional odd Galois representations over Q liftings of 2-dimensional projective Galois representations over Q. Lecture Notes in Mathematics, 1585:1–36, 1994. CODEN LN-MAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074107/.

Kinzelbach:1994:GCD

[Kin94] Martin Kinzelbach. A. Geometrical construction of 2-dimensional Galois representations of A_{5}-type. B. On the realisation of the groups PSL_{2}(1) as Galois groups over number fields by means of l-torsion points of elliptic curves. Lecture Notes in Mathematics, 1585:47–58, 1994. CODEN LNMAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074109/.

Kirschenhofer:1990:VSD

[Kir90] Peter Kirschenhofer. On the variance of the sum of digits function. Lecture Notes in Mathematics, 1452:112–116, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096983/.

Kisliakov:1994:BS

[Kis94] S. V. Kisliakov. Banach spaces. Lecture Notes in Mathematics, 1573:1–50, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100202/.

Kitada:1993:ACB

[Kit93] Hitoshi Kitada. Asymptotic completeness of N-body wave operators II. A new proof for the short-range case and the asymptotic clustering for long-range systems. Lecture Notes in Mathematics, 1540:149–189, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085479/.

Kitahara:1994:ASF

[Kit94a] Kazuaki Kitahara. Approximation by step functions. Lecture Notes in Mathematics, 1576:78–89, 1994. CODEN LN-MAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091390/.

Kitahara:1994:AVV

[Kit94b] Kazuaki Kitahara. Approximation by vector-valued monotone increasing or convex functions. Lecture Notes in Mathematics, 1576: 58–77, 1994. CODEN LNMAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091389/.

Kitahara:1994:CAS

[Kit94c] Kazuaki Kitahara. Characterizations of approximating spaces of C[a, b] or C_{0}(Q). Lecture Notes in Mathematics, 1576:8–29, 1994. CODEN LNMAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091387/.

Kitahara:1994:P

[Kit94d] Kazuaki Kitahara. Preliminaries. Lecture Notes in Mathematics, 1576:1–7, 1994. CODEN LNMAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091386/.

Kitahara:1994:STH

[Kit94e] Kazuaki Kitahara. Some topics of Haar-like spaces of F [a, b]. Lecture Notes in Mathematics, 1576:30–57, 1994. CODEN LN-MAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091388/.

Kitahara:1994:SAF

[Kit94f] Kazuaki Kitahara. Spaces of Approximating Functions with Haarlike Conditions, volume 1576 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57974-5 (print), 3-540-48404-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 110 pp. LCCN QA3 .L28 no. 1576.

URL http://link.springer.com/book/10.1007/BFb0091385;
 http://www.springerlink.com/content/978-3-540-48404-

Kurzweil:1990:PID

[KJ90] Jaroslav Kurzweil and Jiří Jarník. The pu-integral: its definition and some basic properties. Lecture Notes in Mathematics, 1419: 66–81, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083100/.

Korolev:1993:LTR

[KK93] V. Yu. Korolev and V. M. Kruglov. Limit theorems for random sums of independent random variables. Lecture Notes in Mathematics, 1546:100–120, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084486/.

Kleiman:1990:MPF

[Kle90] Steven L. Kleiman. Multiple-point formulas II: The Hilbert scheme. Lecture Notes in Mathematics, 1436:101–138, 1990. CO-DEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084043/.

Klaas:1997:AFF

[KLGP97a] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Algorithm for finding a filtration and the obliquity. Lecture Notes in Mathematics, 1674:68–77, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094096/.

Klaas:1997:AF

[KLGP97b] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Algorithms on fields. Lecture Notes in Mathematics, 1674:59– 61, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094094/.

Klaas:1997:CG

[KLGP97c] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Chevalley groups. Lecture Notes in Mathematics, 1674:26–29, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094091/.

Klaas:1997:EPW

[KLGP97d] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Elementary properties of width. Lecture Notes in Mathematics, 1674:9–11, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094088/.

Klaas:1997:FSD

[KLGP97e] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Fields of small degree. Lecture Notes in Mathematics, 1674:62– 67, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094095/.

Klaas:1997:I

[KLGP97f] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Introduction. Lecture Notes in Mathematics, 1674:1–8, 1997. CO-DEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094087/.

Klaas:1997:LPG

[KLGP97g] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Linear Pro- p-Groups of Finite Width, volume 1674 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 115 pp. LCCN QA3 .L28 no. 1674. URL http:/

/link.springer.com/book/10.1007/BFb0094086;   http://
www.springerlink.com/content/978-3-540-69623-0.

Klaas:1997:ASG

[KLGP97h] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. p-adically simple groups (p̃-groups). Lecture Notes in Mathematics, 1674:12–20, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094089/.

Klaas:1997:P

[KLGP97i] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Periodicity. Lecture Notes in Mathematics, 1674:21–25, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094090/.

Klaas:1997:SCG

[KLGP97j] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Some classical groups. Lecture Notes in Mathematics, 1674:30– 54, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094092/.

Klaas:1997:SOP

[KLGP97k] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Some open problems. Lecture Notes in Mathematics, 1674: 108, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094100.

Klaas:1997:STG

[KLGP97l] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Some thin groups. Lecture Notes in Mathematics, 1674:55– 58, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094093/.

Klaas:1997:T

[KLGP97m] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Tables. Lecture Notes in Mathematics, 1674:92–105, 1997. CO-DEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094098/.

Klaas:1997:TBT

[KLGP97n] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. The theory behind the tables. Lecture Notes in Mathematics, 1674:78–91, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094097/.

Klaas:1997:UMJ

[KLGP97o] Gundel Klaas, Charles R. Leedham-Green, and Wilhelm Plesken. Uncountably many just infinite pro- p-groups of finite width. Lecture Notes in Mathematics, 1674:106–107, 1997. CODEN LNMAA2. ISBN 3-540-63643-9 (print), 3-540-69623-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094099.

Kumar:1997:FM

[KLS97] Shrawan Kumar, Gérard Laumon, and Ulrich Stuhler. Front matter. Lecture Notes in Mathematics, 1649:??, 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-49701-1/
1.pdf.

Kozlov:1997:AOD

[KM97a] Vladimir Kozlov and Vladimir Maz’ya. Application to ordinary differential equations with operator coefficients. Lecture Notes in Mathematics, 1659:127–136, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094708/.

Kozlov:1997:ASI

[KM97b] Vladimir Kozlov and Vladimir Maz’ya. Asymptotics of solutions at infinity. Lecture Notes in Mathematics, 1659:111–125, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094707/.

Kozlov:1997:BEC

[KM97c] Vladimir Kozlov and Vladimir Maz’ya. Basic equation with constant coefficients. Lecture Notes in Mathematics, 1659:1–14, 1997.

CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094701/.

Kozlov:1997:GFO

[KM97d] Vladimir Kozlov and Vladimir Maz’ya. Green’s function for the operator M(∂_{t}) − ω(t). Lecture Notes in Mathematics, 1659:37– 45, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094704/.

Kozlov:1997:OC

[KM97e] Vladimir Kozlov and Vladimir Maz’ya. The operator M(∂_{t}) − ω_{0} with constant ω_{0}. Lecture Notes in Mathematics, 1659:25–35, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094703/.

Kozlov:1997:OSI

[KM97f] Vladimir Kozlov and Vladimir Maz’ya. The operator M(∂_{t}) on a semiaxis and an interval. Lecture Notes in Mathematics, 1659: 15–24, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094702/.

Kozlov:1997:PUV

[KM97g] Vladimir Kozlov and Vladimir Maz’ya. Properties of M(∂_{t}−ω(t)) under various assumptions about ω(t). Lecture Notes in Mathematics, 1659:81–110, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094706/.

Kozlov:1997:THO

[KM97h] Vladimir Kozlov and Vladimir Maz’ya. Theory of a Higher-Order Sturm–Liouville Equation, volume 1659 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 140 pp. LCCN QA3 .L28 no. 1659.

URL http://link.springer.com/book/10.1007/BFb0094700;
 http://www.springerlink.com/content/978-3-540-69122-
8.

Kozlov:1997:USP

[KM97i] Vladimir Kozlov and Vladimir Maz’ya. Uniqueness and solvability properties of the operator M(∂_{t} − ω(t)). Lecture Notes in Mathematics, 1659:47–80, 1997. CODEN LNMAA2. ISBN 3-540-63065-1 (print), 3-540-69122-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094705/.

Kollar:1992:RCF

[KMM92] János Kollár, Yoichi Miyaoka, and Shigefumi Mori. Rational curves on Fano varieties. Lecture Notes in Mathematics, 1515: 100–105, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098339/.

Kajitani:1991:HCP

[KN91] Kunihiko Kajitani and Tatsuo Nishitani. The Hyperbolic Cauchy Problem, volume 1505 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-55018-6 (print), 3-540-46655-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 168 pp. LCCN QA3 .L28 no. 1505. URL http:/

/link.springer.com/book/10.1007/BFb0090882;   http://
www.springerlink.com/content/978-3-540-46655-0.

Knarr:1995:ES

[Kna95a] Norbert Knarr. Examples and supplements. Lecture Notes in Mathematics, 1611:48–60, 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096315/.

Knarr:1995:F

[Kna95b] Norbert Knarr. Foundations. Lecture Notes in Mathematics, 1611: 5–24, 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096312/.

Knarr:1995:I

[Kna95c] Norbert Knarr. Introduction. Lecture Notes in Mathematics, 1611:1–4, 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096311/.

Knarr:1995:KS

[Kna95d] Norbert Knarr. Kinematic spaces. Lecture Notes in Mathematics, 1611:40–47, 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096314/.

Knarr:1995:LCD

[Kna95e] Norbert Knarr. Locally compact 4-dimensional translation planes. Lecture Notes in Mathematics, 1611:61–77, 1995. CODEN LN-MAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096316/.

Knarr:1995:LCT

[Kna95f] Norbert Knarr. Locally compact translation planes of higher dimension. Lecture Notes in Mathematics, 1611:95–103, 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096318/.

Knarr:1995:PLT

[Kna95g] Norbert Knarr. Planes of Lenz type V with complex kernel. Lecture Notes in Mathematics, 1611:78–94, 1995. CODEN LN-MAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096317/.

Knarr:1995:SDP

[Kna95h] Norbert Knarr. Spreads of 3-dimensional projective spaces. Lecture Notes in Mathematics, 1611:25–39, 1995. CODEN LN-MAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096313/.

Knarr:1995:TPF

[Kna95i] Norbert Knarr. Translation Planes: Foundations and Construction Principles, volume 1611 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60208-9 (print), 3-540-44724-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 112 pp. LCCN QA3 .L28 no. 1611.

URL http://link.springer.com/book/10.1007/BFb0096310;
 http://www.springerlink.com/content/978-3-540-44724-
5.

Knebusch:1992:STL

[Kne92] Manfred Knebusch. Semialgebraic topology in the last ten years. Lecture Notes in Mathematics, 1524:1–36, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084606/.

Knill:1991:ULE

[Kni91] Oliver Knill. The upper Lyapunov exponent of Sl(2, R) cocycles: Discontinuity and the problem of positivity. Lecture Notes in Mathematics, 1486:86–97, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086660/.

Knight:1993:SRM

[Kni93] Frank Knight. Some remarks on mutual windings. Lecture Notes in Mathematics, 1557:36–43, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087962/.

Kincaid:1990:RVP

[KO90a] David R. Kincaid and Thomas C. Oppe. Recent vectorization and parallelization of ITPACKV. Lecture Notes in Mathematics, 1457:58–78, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090902/.

Knapp:1990:KCR

[KO90b] K. Knapp and E. Ossa. KO-codegree and real line bundles. Lecture Notes in Mathematics, 1418:130–139, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083698/.

Knaust:1991:WNS

[KO91] Helmut Knaust and Edward Odell. Weakly null sequences with upper ℓ_{p}-estimates. Lecture Notes in Mathematics, 1470:85– 107, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090216/.

Kobayashi:1993:SLL

[KO93] Yoshikazu Kobayashi and Shinnosuke Oharu. Semigroups of locally Lipschitzian operators and applications. Lecture Notes in Mathematics, 1540:191–211, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085480/.

Kobayashi:1990:MVB

[Kob90] Shoshichi Kobayashi. On moduli of vector bundles. Lecture Notes in Mathematics, 1422:45–57, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089404/.

Kochman:1990:CS

[Koc90a] Stanley O. Kochman. The Chicago stem (π_{N}^{S}, 32 ≤ N ≤ 45). Lecture Notes in Mathematics, 1423:139–211, 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083801/.

Kochman:1990:EAI

[Koc90b] Stanley O. Kochman. The elements of arf invariant one. Lecture Notes in Mathematics, 1423:284–293, 1990. CODEN LN-MAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083803/.

Kochman:1990:IJ

[Koc90c] Stanley O. Kochman. The image of J. Lecture Notes in Mathematics, 1423:72–98, 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083799/.

Kochman:1990:I

[Koc90d] Stanley O. Kochman. Introduction. Lecture Notes in Mathematics, 1423:1–11, 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083796/.

Kochman:1990:JS

[Koc90e] Stanley O. Kochman. The Japanese stems (π_{N}, 9 ≤ N ≤ 31). Lecture Notes in Mathematics, 1423:99–138, 1990. CODEN LN-MAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083800/.

Kochman:1990:LDC

[Koc90f] Stanley O. Kochman. Low dimensional computations. Lecture Notes in Mathematics, 1423:35–71, 1990. CODEN LN-MAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083798/.

Kochman:1990:NS

[Koc90g] Stanley O. Kochman. The new stems (π_{N}^{S}, 46 ≤ N ≤ 64). Lecture Notes in Mathematics, 1423:212–283, 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083802/.

Kochman:1990:SHG

[Koc90h] Stanley O. Kochman. Stable Homotopy Groups of Spheres: A Computer-Assisted Approach, volume 1423 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 330 pp. LCCN QA3 .L28 no. 1423. URL http:/

/link.springer.com/book/10.1007/BFb0083795;   http://
www.springerlink.com/content/978-3-540-46993-3.

Kochman:1990:TB

[Koc90i] Stanley O. Kochman. Toda brackets. Lecture Notes in Mathematics, 1423:12–34, 1990. CODEN LNMAA2. ISBN 3-540-52468-1 (print), 3-540-46993-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083797/.

Kock:1991:APM

[Koc91] Anders Kock. Algebras for the partial map classifier monad. Lecture Notes in Mathematics, 1488:262–278, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084225/.

Koecher:1999:CJA

[Koe99a] Max Koecher. Complex Jordan algebras. Lecture Notes in Mathematics, 1710:93–108, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096290/.

Koecher:1999:DP

[Koe99b] Max Koecher. Domains of positivity. Lecture Notes in Mathematics, 1710:1–33, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096286/.

Koecher:1999:HS

[Koe99c] Max Koecher. Half-spaces. Lecture Notes in Mathematics, 1710: 127–155, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096292/.

Koecher:1999:JA

[Koe99d] Max Koecher. Jordan algebras. Lecture Notes in Mathematics, 1710:53–72, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096288/.

Koecher:1999:JAO

[Koe99e] Max Koecher. Jordan algebras and omega domains. Lecture Notes in Mathematics, 1710:109–126, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096291/.

Koecher:1999:MNJ

[Koe99f] Max Koecher. The Minnesota Notes on Jordan Algebras and Their Applications, volume 1710 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 175 pp. LCCN QA3 .L28 no. 1710.

URL http://link.springer.com/book/10.1007/BFb0096285;
 http://www.springerlink.com/content/978-3-540-48402-

8. Edited and annotated by Aloys Krieg and Sebastian Walcher.

Koecher:1999:OD

[Koe99g] Max Koecher. Omega domains. Lecture Notes in Mathematics, 1710:35–51, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096287/.

Koecher:1999:RCJ

[Koe99h] Max Koecher. Real and complex Jordan algebras. Lecture Notes in Mathematics, 1710:73–92, 1999. CODEN LNMAA2. ISBN 3-540-66360-6 (print), 3-540-48402-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096289/.

Kolyvagin:1991:SST

[Kol91] V. A. Kolyvagin. On the structure of Shafarevich–Tate groups. Lecture Notes in Mathematics, 1479:94–121, 1991. CODEN LN-MAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086267/.

Komatsu:1991:MAG

[Kom91] Hikosaburo Komatsu. Microlocal analysis in Gevrey classes and in complex domains. Lecture Notes in Mathematics, 1495:161– 236, 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085124/.

Komatsu:1993:FAR

[Kom93a] Hikosaburo Komatsu, editor. Functional Analysis and Related Topics, 1991: Proceedings of the International Conference in Memory of Professor Kôsaku Yosida held at RIMS, Kyoto University, Japan, July 29–Aug. 2, 1991, volume 1540 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1540.

URL http://link.springer.com/book/10.1007/BFb0085469;
 http://www.springerlink.com/content/978-3-540-47565-
1.

Komatsu:1993:OCS

[Kom93b] Hikosaburo Komatsu. Operational calculus and semi-groups of operators. Lecture Notes in Mathematics, 1540:213–234, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085481/.

Konno:1990:URN

[Kon90] Yasuko Konno. Unitary representations with nonzero multiplicities in L^{2}(Γ \ G). Lecture Notes in Mathematics, 1447:237– 248, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085731/.

Konig:1998:BDS

[Kön98a] Steffen König. Basic definitions and some examples. Lecture Notes in Mathematics, 1685:5–32, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096368/.

Konig:1998:RFT

[Kön98b] Steffen König. Rickard’s fundamental theorem. Lecture Notes in Mathematics, 1685:33–50, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096369/.

Kooman:1993:FVS

[Koo93] Robert-Jan Kooman. Faltings’s version of Siegel’s lemma. Lecture Notes in Mathematics, 1566:93–96, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48208-6_10/.

Koschorke:1991:SLM

[Kos91] Ulrich Koschorke. Semicontractible link maps and their suspensions. Lecture Notes in Mathematics, 1474:160–169, 1991. CO-DEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084744/.

Koshelev:1995:DSS

[Kos95a] Alexander Koshelev. Differentiability of solutions for second order elliptic systems. Lecture Notes in Mathematics, 1614:108– 174, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094486/.

Koshelev:1995:NSS

[Kos95b] Alexander Koshelev. The Navier–Stokes system; strong solutions. Lecture Notes in Mathematics, 1614:216–247, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094488/.

Koshelev:1995:RSN

[Kos95c] Alexander Koshelev. Regularity of solutions for non degenerated quasilinear second order elliptic systems of the divergent form with bounded nonlinearities. Lecture Notes in Mathematics, 1614:23–71, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094484/.

Koshelev:1995:RSP

[Kos95d] Alexander Koshelev. Regularity of solutions for parabolic systems with some applications. Lecture Notes in Mathematics, 1614: 175–215, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094487/.

Koshelev:1995:RPQ

[Kos95e] Alexander Koshelev. Regularity Problem for Quasilinear Elliptic and Parabolic Systems, volume 1614 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 255 pp. LCCN QA3 .L28 no. 1614.

URL http://link.springer.com/book/10.1007/BFb0094482;
 http://www.springerlink.com/content/978-3-540-44772-
6.

Koshelev:1995:SPA

[Kos95f] Alexander Koshelev. Some properties and applications of regular solutions for quasilinear elliptic systems. Lecture Notes in Mathematics, 1614:72–107, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094485/.

Koshelev:1995:WSU

[Kos95g] Alexander Koshelev. Weak solutions and the universal iterative process. Lecture Notes in Mathematics, 1614:1–22, 1995. CODEN LNMAA2. ISBN 3-540-60251-8 (print), 3-540-44772-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094483/.

Kovacs:1990:GCA

[Kov90] L. G. Kovács, editor. Groups — Canberra 1989: Australian National University Group Theory Program 1989, volume 1456 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1456. URL http:/

/link.springer.com/book/10.1007/BFb0100725;   http://
www.springerlink.com/content/978-3-540-46900-1.

Kowalski:1992:MGS

[Kow92] Zbigniew S. Kowalski. On mixing generalized skew products. Lecture Notes in Mathematics, 1514:128–130, 1992. CODEN LN-MAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097533/.

Kozima:1990:GHC

[Koz90a] Kazumoto Kozima. On the generalized homology of the connective fibering of bu. Lecture Notes in Mathematics, 1418:205– 209, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083704/.

Kozono:1990:SSN

[Koz90b] Hideo Kozono. Strong solution for the Navier–Stokes flow in the half-space. Lecture Notes in Mathematics, 1431:84–86, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086057/.

Kozlowski:1991:SIA

[Koz91] Michael Kozlowski. Some improper affine spheres in A_{3}. Lecture Notes in Mathematics, 1481:104–107, 1991. CODEN LN-MAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083633/.

Kirschenhofer:1990:APC

[KP90] Peter Kirschenhofer and Helmut Prodinger. On the analysis of probabilistic counting. Lecture Notes in Mathematics, 1452:117– 120, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096984/.

Kurtz:1996:BM

[KP96a] Thomas G. Kurtz and Philip E. Protter. Back matter. Lecture Notes in Mathematics, 1627:??, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-68513-5/1.pdf.

Kurtz:1996:WCSa

[KP96b] Thomas G. Kurtz and Philip E. Protter. Weak convergence of stochastic integrals and differential equations. Lecture Notes in Mathematics, 1627:1–41, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093176/.

Kurtz:1996:WCSb

[KP96c] Thomas G. Kurtz and Philip E. Protter. Weak convergence of stochastic integrals and differential equations II: Infinite dimensional case. Lecture Notes in Mathematics, 1627:197–285, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093181/.

Kurdyka:1990:MFS

[KPR90] K. Kurdyka, J. B. Poly, and G. Raby. Moyennes des fonctions sous-analytiques, densité, cône tangent et tranches. (French) []. Lecture Notes in Mathematics, 1420:170–177, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083918/.

Kenig:1990:IVP

[KPV90] Carlos E. Kenig, Gustavo Ponce, and Luis Vega. The initial value problem for a class of nonlinear dispersive equations. Lecture Notes in Mathematics, 1450:141–156, 1990. CODEN LN-MAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084903/.

Koranyi:1990:QMC

[KR90] Adam Korányi and Hans Martin Reimann. Quasiconformal mappings on CR manifolds. Lecture Notes in Mathematics, 1422:

59–75, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089405/.

Keimel:1992:A

[KR92a] Klaus Keimel and Walter Roth. Approximation. Lecture Notes in Mathematics, 1517:68–80, 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089195/.

Keimel:1992:I

[KR92b] Klaus Keimel and Walter Roth. Introduction. Lecture Notes in Mathematics, 1517:1–7, 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089191/.

Keimel:1992:LCC

[KR92c] Klaus Keimel and Walter Roth. Locally convex cones. Lecture Notes in Mathematics, 1517:8–24, 1992. CODEN LN-MAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089192/.

Keimel:1992:NC

[KR92d] Klaus Keimel and Walter Roth. Nachbin cones. Lecture Notes in Mathematics, 1517:81–105, 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089196/.

Keimel:1992:OCA

[KR92e] Klaus Keimel and Walter Roth. Ordered Cones and Approximation, volume 1517 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 134 pp. LCCN QA3 .L28 no. 1517. URL http:/

/link.springer.com/book/10.1007/BFb0089190;   http://
www.springerlink.com/content/978-3-540-47079-3.

Keimel:1992:QE

[KR92f] Klaus Keimel and Walter Roth. Quantitative estimates. Lecture Notes in Mathematics, 1517:106–128, 1992. CODEN LN-MAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089197/.

Keimel:1992:S

[KR92g] Klaus Keimel and Walter Roth. Subcones. Lecture Notes in Mathematics, 1517:55–67, 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089194/.

Keimel:1992:UCO

[KR92h] Klaus Keimel and Walter Roth. Uniformly continuous operators and the dual cone. Lecture Notes in Mathematics, 1517: 25–54, 1992. CODEN LNMAA2. ISBN 3-540-55445-9 (print), 3-540-47079-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089193/.

Krattenthaler:1990:DES

[Kra90] Christian Krattenthaler. A determinant evaluation and some enumeration results for plane partitions. Lecture Notes in Mathematics, 1452:121–131, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096985/.

Kroner:1990:AEF

[Krö90] Dietmar Kröner. Asymptotic expansions for a flow with a dynamic contact angle. Lecture Notes in Mathematics, 1431:49–59, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086054/.

Krob:1991:ERA

[Kro91] Daniel Krob. Expressions rationnelles sur un anneau. (French) []. Lecture Notes in Mathematics, 1478:215–243, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083505/.

Krupkova:1997:BGT

[Kru97a] Olga Krupková. Basic geometric tools. Lecture Notes in Mathematics, 1678:20–40, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093440/.

Krupkova:1997:GIM

[Kru97b] Olga Krupková. Geometric intergration methods. Lecture Notes in Mathematics, 1678:174–207, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093447/.

Krupkova:1997:GOV

[Kru97c] Olga Krupková. The Geometry of Ordinary Variational Equations, volume 1678 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 251 pp. LCCN QA3 .L28 no. 1678. URL http:

//link.springer.com/book/10.1007/BFb0093438;   http:/
/www.springerlink.com/content/978-3-540-69657-5.

Krupkova:1997:HS

[Kru97d] Olga Krupková. Hamiltonian systems. Lecture Notes in Mathematics, 1678:80–96, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093443/.

Krupkova:1997:I

[Kru97e] Olga Krupková. Introduction. Lecture Notes in Mathematics, 1678:1–19, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093439/.

Krupkova:1997:LDF

[Kru97f] Olga Krupková. Lagrangean dynamics on fibered manifolds. Lecture Notes in Mathematics, 1678:41–51, 1997. CODEN LN-MAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093441/.

Krupkova:1997:LSF

[Kru97g] Olga Krupková. Lagrangean systems on π: R × M ^{≫}R. Lecture Notes in Mathematics, 1678:208–228, 1997. CODEN LN-MAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093448/.

Krupkova:1997:RLS

[Kru97h] Olga Krupková. Regular Lagrangean systems. Lecture Notes in Mathematics, 1678:97–128, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093444/.

Krupkova:1997:SLSa

[Kru97i] Olga Krupková. Singular Lagrangean systems. Lecture Notes in Mathematics, 1678:129–148, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093445/.

Krupkova:1997:SLSb

[Kru97j] Olga Krupková. Symmetries of Lagrangean systems. Lecture Notes in Mathematics, 1678:149–173, 1997. CODEN LN-MAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093446/.

Krupkova:1997:VE

[Kru97k] Olga Krupková. Variational equations. Lecture Notes in Mathematics, 1678:52–79, 1997. CODEN LNMAA2. ISBN 3-540-63832-6 (print), 3-540-69657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093442/.

Krengel:1992:ETR

[KRW92] Ulrich Krengel, Karin Richter, and Volker Warstat, editors. Ergodic Theory and Related Topics III: Proceedings of the International Conference held in Güstrow, Germany, October 22–27, 1990, volume 1514 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1514. URL http:/

/link.springer.com/book/10.1007/BFb0097523;   http://
www.springerlink.com/content/978-3-540-47076-2.

Krylov:1995:MPK

[Kry95] N. V. Krylov. A martingale proof of the Khinchin iterated logarithm law for Wiener processes. Lecture Notes in Mathematics, 1613:25–29, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094196/.

Krylov:1999:KEF

[Kry99] N. V. Krylov. On Kolmogorov’s equations for finite dimensional diffusions. Lecture Notes in Mathematics, 1715:1–63, 1999. CO-DEN LNMAA2. ISBN 3-540-66545-5 (print), 3-540-48161-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092417/.

Kosmann-Schwarzbach:1990:GLP

[KS90a] Yvette Kosmann-Schwarzbach. Groupes de Lie–Poisson quasitriangulaires. (French) []. Lecture Notes in Mathematics, 1416: 161–177, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097470/.

Krasilnikov:1990:FBL

[KS90b] A. N. Krasil’nikov and A. L. Shmel’kin. On finite bases for laws of triangular matrices. Lecture Notes in Mathematics, 1456:10– 13, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100727/.

Kiran:1992:RJL

[KS92] Shashi Kiran and Ajit Iqbal Singh. Role of James like spaces in multiplicative linear functionals on operator algebras. Lecture Notes in Mathematics, 1511:172–180, 1992. CODEN LN-MAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093810/.

Kipnis:1996:LER

[KS96] C. Kipnis and E. Saada. Un lien entre réseaux de neurones et systèmes de particules: Un modèle de rétinotopie. (French) []. Lecture Notes in Mathematics, 1626:55–67, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094641/.

Kowalski:1991:CCL

[KT91] O. Kowalski and F. Tricerri. A canonical connection for locally homogeneous Riemannian manifolds. Lecture Notes in Mathematics, 1481:97–103, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083632/.

Kawazu:1993:MDP

[KT93] Kiyoshi Kawazu and Hiroshi Tanaka. On the maximum of a diffusion process in a drifted Brownian environment. Lecture Notes in Mathematics, 1557:78–85, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087965/.

Katsman:1992:SLC

[KTV92] Gregory L. Katsman, Michael A. Tsfasman, and Serge G. Vladuţ. Spectra of linear codes and error probability of decoding. Lecture Notes in Mathematics, 1518:82–98, 1992. CODEN LN-MAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087994/.

Korezlioglu:1990:NCD

[KÜ90a] H. Körezlioğlu and A. S. Üstünel. A new class of distributions on Wiener spaces. Lecture Notes in Mathematics, 1444:106– 121, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083611/.

Korezlioglu:1990:SAR

[KÜ90b] Hayri Körezlioğlu and Ali Suleyman Üstünel, editors. Stochastic Analysis and Related Topics II: Proceedings of a Second Workshop held in Silivri, Turkey, July 18–30, 1988, volume 1444 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1444.

URL http://link.springer.com/book/10.1007/BFb0083608;
 http://www.springerlink.com/content/978-3-540-46596-
6.

Kucera:1990:RGF

[Kuč90] Antonín Kučera. Randomness and generalizations of fixed point free functions. Lecture Notes in Mathematics, 1432:245–254, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086121/.

Kuhnel:1995:CM

[Küh95a] Wolfgang Kühnel. (k − 1)-connected 2k-manifolds. Lecture Notes in Mathematics, 1612:56–77, 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096345/.

Kuhnel:1995:MTS

[Küh95b] Wolfgang Kühnel. 3-manifolds and twisted sphere bundles. Lecture Notes in Mathematics, 1612:78–84, 1995. CODEN LN-MAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096346/.

Kuhnel:1995:CSM

[Küh95c] Wolfgang Kühnel. Connected sums and manifolds with boundary. Lecture Notes in Mathematics, 1612:85–94, 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096347/.

Kuhnel:1995:IBN

[Küh95d] Wolfgang Kühnel. Introduction and basic notions. Lecture Notes in Mathematics, 1612:1–5, 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096342/.

Kuhnel:1995:MCP

[Küh95e] Wolfgang Kühnel. Miscellaneous cases and pseudomanifolds. Lecture Notes in Mathematics, 1612:95–109, 1995. CODEN LN-MAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096348/.

Kuhnel:1995:TPSb

[Küh95f] Wolfgang Kühnel. Tight Polyhedral Submanifolds and Tight Triangulations, volume 1612 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 122 pp. LCCN QA3 .L28 no. 1612.

URL http://link.springer.com/book/10.1007/BFb0096341;
 http://www.springerlink.com/content/978-3-540-49452-
2.

Kuhnel:1995:TPSa

[Küh95g] Wolfgang Kühnel. Tight polyhedral surfaces. Lecture Notes in Mathematics, 1612:6–40, 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096343/.

Kuhnel:1995:TT

[Küh95h] Wolfgang Kühnel. Tightness and k-tightness. Lecture Notes in Mathematics, 1612:41–55, 1995. CODEN LNMAA2. ISBN 3-540-60121-X (print), 3-540-49452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096344/.

Kuiken:1992:MMI

[Kui92] H. K. Kuiken. Mathematical modelling of industrial processes. Lecture Notes in Mathematics, 1521:1–63, 1992. CODEN LN-MAA2. ISBN 3-540-55595-1 (print), 3-540-47247-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098365/.

Kuksin:1993:NII

[Kuk93a] Sergej B. Kuksin. Nearly Integrable Infinite-Dimensional Hamiltonian Systems, volume 1556 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57161-2 (print), 3-540-47920-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxviii + 101 pp. LCCN QA3 .L28 no. 1556.

URL http://link.springer.com/book/10.1007/BFb0092243;
 http://www.springerlink.com/content/978-3-540-47920-
8.

Kuksin:1993:PMT

[Kuk93b] Sergej B. Kuksin. Proof of the main theorem. Lecture Notes in Mathematics, 1556:45–90, 1993. CODEN LNMAA2. ISBN 3-540-57161-2 (print), 3-540-47920-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092246/.

Kuksin:1993:SMT

[Kuk93c] Sergej B. Kuksin. Statement of the main theorem and its consequences. Lecture Notes in Mathematics, 1556:13–44, 1993. CO-DEN LNMAA2. ISBN 3-540-57161-2 (print), 3-540-47920-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092245/.

Kuksin:1993:SSH

[Kuk93d] Sergej B. Kuksin. Symplectic structures and Hamiltonian systems in scales of Hilbert spaces. Lecture Notes in Mathematics, 1556:1–12, 1993. CODEN LNMAA2. ISBN 3-540-57161-2 (print), 3-540-47920-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092244/.

Kulikov:1991:FGC

[Kul91] Vic. S. Kulikov. On the fundamental group of the complement of a hypersurface in C^{n}. Lecture Notes in Mathematics, 1479: 122–130, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086268/.

Kulish:1992:QGP

[Kul92] Petr P. Kulish, editor. Quantum Groups: Proceedings of Workshops held in the Euler International Mathematical Institute, Leningrad, Fall 1990, volume 1510 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1510. URL http:/

/link.springer.com/book/10.1007/BFb0101174;   http://
www.springerlink.com/content/978-3-540-47020-5.

Kummer:1990:RER

[Kum90] Martin Kummer. Recursive enumeration without repetition revisited. Lecture Notes in Mathematics, 1432:255–275, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086122/.

Kumar:1992:QUP

[Kum92] Ajay Kumar. A qualitative uncertainity principle for hypergroups. Lecture Notes in Mathematics, 1511:1–9, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093792/.

Kumar:1997:IGM

[Kum97] Shrawan Kumar. Infinite Grassmannians and moduli spaces of Gbundles. Lecture Notes in Mathematics, 1649:1–49, 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094424/.

Kunyavskii:1992:SPC

[Kun92] Boris È. Kunyavskii. Sphere packings centered at S-units of algebraic tori. Lecture Notes in Mathematics, 1518:108–121, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087996/.

Kurokawa:1990:ADS

[Kur90] Nobushige Kurokawa. Analyticity of Dirichlet series over prime powers. Lecture Notes in Mathematics, 1434:168–177, 1990. CO-DEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097133/.

Kusner:1991:MPI

[Kus91] Rob Kusner. A maximum principle at infinity and the topology of complete embedded surfaces with constant mean curvature. Lecture Notes in Mathematics, 1481:108–114, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083634/.

Kusuoka:1993:LDP

[Kus93] S. Kusuoka. Lecture on diffusion processes on nested fractals. Lecture Notes in Mathematics, 1567:39–98, 1993. CODEN LN-MAA2. ISBN 3-540-57516-2 (print), 3-540-48205-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074240/.

Kasangian:1991:TDC

[KV91] S. Kasangian and S. Vigna. Trees in distributive categories. Lecture Notes in Mathematics, 1488:237–248, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084223/.

Korogodsky:1992:QSH

[KV92] L. I. Korogodsky and L. L. Vaksman. Quantum G-spaces and Heisenberg algebra. Lecture Notes in Mathematics, 1510:56– 66, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101178/.

Kiming:1994:EDO

[KW94] Ian Kiming and Xiangdong Wang. Examples of 2-dimensional, odd Galois representations of A_{5}-type over Q satisfying the Artin conjecture. Lecture Notes in Mathematics, 1585:109–121, 1994. CODEN LNMAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074112/.

Kwakernaak:1991:PAO

[Kwa91] Huibert Kwakernaak. The polynomial approach to H_{∞}-optimal regulation. Lecture Notes in Mathematics, 1496:141–221, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084469/.

Kawabata:1991:NMS

[KY91] Shigetoku Kawabata and Toshio Yamada. On Newton’s method for stochastic differential equations. Lecture Notes in Mathematics, 1485:121–137, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100852/.

Klebanov:1993:SMR

[KY93] L. B. Klebanov and A. Yu. Yakovlev. A stochastic model of radiation carcinogenesis. Lecture Notes in Mathematics, 1546: 89–99, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084485/.

Kress:1991:TDR

[KZ91] Rainer Kress and Axel Zinn. Three dimensional reconstructions in inverse obstacle scattering. Lecture Notes in Mathematics, 1497: 112–125, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084512/.

Kwong:1992:DO

[KZ92a] Man Kam Kwong and Anton Zettl. The difference operator. Lecture Notes in Mathematics, 1536:117–143, 1992. CODEN LNMAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090869/.

Kwong:1992:I

[KZ92b] Man Kam Kwong and Anton Zettl. Introduction. Lecture Notes in Mathematics, 1536:1–2, 1992. CODEN LNMAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090865.

Kwong:1992:NID

[KZ92c] Man Kam Kwong and Anton Zettl. Norm Inequalities for Derivatives and Differences, volume 1536 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 150 pp. LCCN QA3 .L28 no. 1536.

URL http://link.springer.com/book/10.1007/BFb0090864;
 http://www.springerlink.com/content/978-3-540-47548-
4.

Kwong:1992:N

[KZ92d] Man Kam Kwong and Anton Zettl. The norms of y, y′, y′′. Lecture Notes in Mathematics, 1536:35–83, 1992. CODEN LN-MAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090867/.

Kwong:1992:UWF

[KZ92e] Man Kam Kwong and Anton Zettl. Unit weight functions. Lecture Notes in Mathematics, 1536:3–34, 1992. CODEN LN-MAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090866/.

Kwong:1992:W

[KZ92f] Man Kam Kwong and Anton Zettl. Weights. Lecture Notes in Mathematics, 1536:84–116, 1992. CODEN LNMAA2. ISBN 3-540-56387-3 (print), 3-540-47548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090868/.

Kalashnikov:1993:SPS

[KZ93] Vladimir V. Kalashnikov and Vladimir M. Zolatarev, editors. Stability Problems for Stochastic Models: Proceedings of the International Seminar, held in Suzdal, Russia, Jan. 27–Feb. 2, 1991, volume 1546 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1546. URL http:/

/link.springer.com/book/10.1007/BFb0084476;   http://
www.springerlink.com/content/978-3-540-47645-0.

Konig:1998:DEG

[KZ98] Steffen König and Alexander Zimmermann. Derived Equivalences for Group Rings, volume 1685 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 246 pp. LCCN QA3 .L28 no. 1685.

URL http://link.springer.com/book/10.1007/BFb0096366;
 http://www.springerlink.com/content/978-3-540-69748-
0.

Labesse:1990:SVM

[Lab90] J.-P. Labesse. Signature des variétés modulaires de Hilbert et représentations diédrales. (French) []. Lecture Notes in Mathematics, 1447:249–260, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085732/.

Lambert:1992:WSC

[Lam92] Alan Lambert. Weighted shifts and composition operators on L^{2}. Lecture Notes in Mathematics, 1511:10–17, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093793/.

Lang:1990:EHD

[Lan90a] Serge Lang. Equidimensional higher dimensional theory. Lecture Notes in Mathematics, 1433:57–107, 1990. CODEN LN-MAA2. ISBN 3-540-52785-0 (print), 3-540-47146-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093848/.

Lang:1990:NTO

[Lan90b] Serge Lang. Nevanlinna theory in one variable. Lecture Notes in Mathematics, 1433:9–55, 1990. CODEN LNMAA2. ISBN 3-540-52785-0 (print), 3-540-47146-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093847/.

Langevin:1990:SRG

[Lan90c] Michel Langevin. Sphère de Riemann et géométrie des polynômes. (French) []. Lecture Notes in Mathematics, 1415:138–159, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084884/.

Lang:1991:C

[Lan91a] Reinhard Lang. Conclusion. Lecture Notes in Mathematics, 1498: 113–118, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093936/.

Lang:1991:ETIa

[Lan91b] Reinhard Lang. Explanation of theorem 1 and introduction to an extended Boltzmann theory of entropy. Lecture Notes in Mathematics, 1498:43–70, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093934/.

Lang:1991:ETIb

[Lan91c] Reinhard Lang. Explanation of theorem 2 and introduction to an extended Floquet–Weyl theory. Lecture Notes in Mathematics, 1498:71–112, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093935/.

Lang:1991:GHP

[Lan91d] Reinhard Lang. The general heuristic picture. Lecture Notes in Mathematics, 1498:19–29, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093932/.

Lang:1991:I

[Lan91e] Reinhard Lang. Introduction. Lecture Notes in Mathematics, 1498:1–5, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093930/.

Lang:1991:SKR

[Lan91f] Reinhard Lang. Some known results and open problems. Lecture Notes in Mathematics, 1498:30–42, 1991. CODEN LN-MAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093933/.

Lang:1991:STR

[Lan91g] Reinhard Lang. Spectral Theory of Random Schrödinger Operators: A Genetic Introduction, volume 1498 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 125 pp. LCCN QA3 .L28 no. 1498. URL http:/

/link.springer.com/book/10.1007/BFb0093929;   http://
www.springerlink.com/content/978-3-540-46627-7.

Lang:1991:TSE

[Lan91h] Reinhard Lang. Two simple examples. Lecture Notes in Mathematics, 1498:6–18, 1991. CODEN LNMAA2. ISBN 3-540-54975-7 (print), 3-540-46627-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093931/.

Landvogt:1996:GSQ

[Lan96a] Erasmus Landvogt. The o_{K}-group schemes in the quasi-split case. Lecture Notes in Mathematics, 1619:31–66, 1996. CODEN LNMAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094597/.

Landvogt:1996:A

[Lan96b] Erasmus Landvogt. The apartment. Lecture Notes in Mathematics, 1619:14–30, 1996. CODEN LNMAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094596/.

Landvogt:1996:BC

[Lan96c] Erasmus Landvogt. The building and its compactification. Lecture Notes in Mathematics, 1619:98–149, 1996. CODEN LN-MAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094599/.

Landvogt:1996:BQS

[Lan96d] Erasmus Landvogt. The building in the quasi-split case. Lecture Notes in Mathematics, 1619:67–97, 1996. CODEN LN-MAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094598/.

Landvogt:1996:CBT

[Lan96e] Erasmus Landvogt. A Compactification of the Bruhat–Tits Building, volume 1619 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 152 pp. LCCN QA3 .L28 no. 1619. URL http:/

/link.springer.com/book/10.1007/BFb0094594;   http://
www.springerlink.com/content/978-3-540-45556-1.

Landvogt:1996:I

[Lan96f] Erasmus Landvogt. Introduction. Lecture Notes in Mathematics, 1619:1–13, 1996. CODEN LNMAA2. ISBN 3-540-60427-8 (print), 3-540-45556-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094595/.

Lang:1996:AP

[Lan96g] Serge Lang. Augmented products. Lecture Notes in Mathematics, 1625:109–115, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092629/.

Lang:1996:CF

[Lan96h] Serge Lang. Class formations. Lecture Notes in Mathematics, 1625:166–187, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092633/.

Lang:1996:CT

[Lan96i] Serge Lang. Cohomological triviality. Lecture Notes in Mathematics, 1625:62–72, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092627/.

Lang:1996:CP

[Lan96j] Serge Lang. Cup products. Lecture Notes in Mathematics, 1625: 73–108, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092628/.

Lang:1996:EU

[Lan96k] Serge Lang. Existence and uniqueness. Lecture Notes in Mathematics, 1625:3–36, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092625/.

Lang:1996:GE

[Lan96l] Serge Lang. Group extensions. Lecture Notes in Mathematics, 1625:156–165, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092632/.

Lang:1996:GGT

[Lan96m] Serge Lang. Groups of Galois type. Lecture Notes in Mathematics, 1625:123–155, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092631/.

Lang:1996:RS

[Lan96n] Serge Lang. Relations with subgroups. Lecture Notes in Mathematics, 1625:37–61, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092626/.

Lang:1996:SS

[Lan96o] Serge Lang. Spectral sequences. Lecture Notes in Mathematics, 1625:116–122, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092630/.

Lang:1996:TCG

[Lan96p] Serge Lang. Topics in Cohomology of Groups, volume 1625 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 226 pp. LCCN QA3 .L28 no. 1625. URL http:/

/link.springer.com/book/10.1007/BFb0092624;   http://
www.springerlink.com/content/978-3-540-68337-7.

Larcher:1990:IAD

[Lar90] Gerhard Larcher. An inequality with applications in Diophantine approximation. Lecture Notes in Mathematics, 1452:132– 138, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096986/.

Lascar:1998:OSG

[Las98] Daniel Lascar. Omega-stable groups. Lecture Notes in Mathematics, 1696:45–59, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-68521-0_3/.

Latiolais:1990:TCG

[Lat90] Paul Latiolais, editor. Topology and Combinatorial Group Theory: Proceedings of the Fall Foliage Topology Seminars held in New Hampshire 1986–1988, volume 1440 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1440. URL http:/

/link.springer.com/book/10.1007/BFb0084444;   http://
www.springerlink.com/content/978-3-540-46296-5.

Lauterbach:1991:DNS

[Lau91] Reiner Lauterbach. Dynamics near steady state bifurcations in problems with spherical symmetry. Lecture Notes in Mathematics, 1463:256–265, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085434/.

Laumon:1997:DS

[Lau97] G. Laumon. Drinfeld shtukas. Lecture Notes in Mathematics, 1649:50–109, 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094425/.

Lawvere:1991:ICH

[Law91a] F. William Lawvere. Intrinsic co-Heyting boundaries and the Leibniz rule in certain toposes. Lecture Notes in Mathematics, 1488:279–281, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084226/.

Lawvere:1991:STF

[Law91b] F. William Lawvere. Some thoughts on the future of category theory. Lecture Notes in Mathematics, 1488:1–13, 1991. CODEN LNMAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084208/.

Lax:1996:OTK

[Lax96] Peter D. Lax. Outline of a theory of the KdV equation. Lecture Notes in Mathematics, 1640:70–102, 1996. CODEN LN-MAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093707/.

Lang:1990:TNT

[LC90] Serge Lang and William Cherry. Topics in Nevanlinna Theory, volume 1433 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52785-0 (print), 3-540-47146-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 168 (est.) pp. LCCN QA3 .L28 no. 1433. URL http:/

/link.springer.com/book/10.1007/BFb0093846;   http://
www.springerlink.com/content/978-3-540-47146-2.

LeBarz:1990:QFM

[Le 90] Patrick Le Barz. Quelques formules multisécantes pour les surfaces. (French) []. Lecture Notes in Mathematics, 1436:151– 188, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084045/.

LeJan:1991:SOE

[Le 91] Y. Le Jan. A second order extension of Oseledets theorem. Lecture Notes in Mathematics, 1486:81–85, 1991. CODEN LN-MAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086659/.

LeGall:1992:SPP

[Le 92] Jean-François Le Gall. Some properties of planar Brownian motion. Lecture Notes in Mathematics, 1527:111–235, 1992. CODEN LNMAA2. ISBN 3-540-56250-8 (print), 3-540-47490-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084700/.

LeGall:1994:EMR

[Le 94] Jean-François Le Gall. Exponential moments for the renormalized self-intersection local time of planar Brownian motion. Lecture Notes in Mathematics, 1583:172–180, 1994. CODEN LN-MAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073845/.

LeGall:1997:MAA

[Le 97] Jean-François Le Gall. Marches aléatoires auto-évitantes et mesures de polymère. (French) []. Lecture Notes in Mathematics, 1655:103–112, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0119297/.

Leader:1990:DCF

[Lea90] Solomon Leader. 1-differentials on 1-cells: a further study. Lecture Notes in Mathematics, 1419:82–96, 1990. CODEN LN-MAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083101/.

Leandre:1996:CBN

[Léa96] R. Léandre. Cohomologie de Bismut–Nualart–Pardoux et cohomologie de Hochschild entière. (French) []. Lecture Notes in Mathematics, 1626:68–99, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094642/.

Ledoux:1990:NLD

[Led90] Michel Ledoux. A note on large deviations for Wiener chaos. Lecture Notes in Mathematics, 1426:1–14, 1990. CODEN LN-MAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083753/.

Ledrappier:1992:EPS

[Led92] François Ledrappier. Ergodic properties of the stable foliations. Lecture Notes in Mathematics, 1514:131–145, 1992. CODEN LN-MAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097534/.

Ledoux:1996:IGA

[Led96] Michel Ledoux. Isoperimetry and Gaussian analysis. Lecture Notes in Mathematics, 1648:165–294, 1996. CODEN LN-MAA2. ISBN 3-540-62055-9 (print), 3-540-49635-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095676/.

Ledoux:1999:CML

[Led99] Michel Ledoux. Concentration of measure and logarithmic Sobolev inequalities. Lecture Notes in Mathematics, 1709:120– 216, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096511/.

Lehrer:1990:GRG

[Leh90] G. I. Lehrer. Group representations, geometry and topology. Lecture Notes in Mathematics, 1456:14–31, 1990. CODEN LN-MAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100728/.

Leizarowitz:1991:ERL

[Lei91] Arie Leizarowitz. Eigenvalue representation for the Lyapunov exponents of certain Markov processes. Lecture Notes in Mathematics, 1486:51–63, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086657/.

Lemarie:1990:AME

[Lem90a] Pierre Gilles Lemarié. Analyse multi-échelles et ondelettes à support compact. (French) []. Lecture Notes in Mathematics, 1438: 26–38, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083513/.

Lemarie:1990:ITO

[Lem90b] Pierre Gilles Lemarié. Introduction à la théorie des ondelettes. (French) []. Lecture Notes in Mathematics, 1438:1–13, 1990. CO-DEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083511/.

Lemarie:1990:OSA

[Lem90c] Pierre Gilles Lemarié, editor. Les Ondelettes en 1989: Séminaire d’Analyse Harmonique, Université de Paris-Sud, Orsay. (French) [Wavelets in 1989: Seminar on Harmonic Analysis, University of Paris-Sud, Orsay], volume 1438 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1438. URL http:/

/link.springer.com/book/10.1007/BFb0083510;   http://
www.springerlink.com/content/978-3-540-47179-0.

Lemarie:1990:WES

[Lem90d] Pierre Gilles Lemarie. Wavelets in 1989: an extended summary. Lecture Notes in Mathematics, 1438:192–212, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083520/.

Lenstra:1993:NFSa

[Len93] H. W. Lenstra, Jr. The number field sieve: an annotated bibliography. Lecture Notes in Mathematics, 1554:1–3, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091535/.

Lepingle:1992:OIU

[Lép92] D. Lépingle. Orthogonalité et intégrabilité uniforme de martingales discrètes. (French) []. Lecture Notes in Mathematics, 1526: 167–169, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084320/.

Lesigne:1992:ETA

[Les92] Emmanuel Lesigne. Ergodic theorem along a return time sequence. Lecture Notes in Mathematics, 1514:146–152, 1992. CO-DEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097535/.

Leung:1991:ETP

[Leu91] Denny H. Leung. Embedding ℓ^{1} into tensor products of Banach spaces. Lecture Notes in Mathematics, 1470:171–176, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090220/.

Leuridan:1996:DEI

[Leu96] Christophe Leuridan. Une démonstration élémentaire d’une identité de Biane et Yor. (French) []. Lecture Notes in Mathematics, 1626:255–260, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094653/.

Li:1991:CBF

[Li91] Chun Li. Characterization of BMO_{p}^{sq}-functions by generalized Carleson measure. Lecture Notes in Mathematics, 1494:84–94, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087760/.

Lichnerowicz:1990:GLS

[Lic90] André Lichnerowicz. Groupes de Lie à structures symplectiques ou Kähleriennes invariantes. (French) []. Lecture Notes in Mathematics, 1416:192–209, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097472/.

Liess:1993:BSG

[Lie93a] Otto Liess. Bi-symplectic geometry and multihomogeneous maps. Lecture Notes in Mathematics, 1555:145–191, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084682/.

Liess:1993:CRHb

[Lie93b] Otto Liess. Conical Refraction and Higher Microlocalization, volume 1555 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 389 pp. LCCN QA3 .L28 no. 1555. URL http:

//link.springer.com/book/10.1007/BFb0084678;   http:/
/www.springerlink.com/content/978-3-540-47905-5.

Liess:1993:CRHa

[Lie93c] Otto Liess. Conical refraction, hyperbolicity and slowness surfaces. Lecture Notes in Mathematics, 1555:225–278, 1993. CO-DEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084684/.

Liess:1993:FIO

[Lie93d] Otto Liess. Fourier integral operators. Lecture Notes in Mathematics, 1555:193–223, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084683/.

Liess:1993:HOW

[Lie93e] Otto Liess. Higher order wave front sets. Lecture Notes in Mathematics, 1555:33–93, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084680/.

Liess:1993:I

[Lie93f] Otto Liess. Introduction. Lecture Notes in Mathematics, 1555: 1–32, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084679/.

Liess:1993:PAH

[Lie93g] Otto Liess. Partial analyticity, higher microlocalization and sheaves. Lecture Notes in Mathematics, 1555:347–379, 1993. CO-DEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084687/.

Liess:1993:PRB

[Lie93h] Otto Liess. Propagation of regularity up to the boundary. Lecture Notes in Mathematics, 1555:279–308, 1993. CODEN LN-MAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084685/.

Liess:1993:PO

[Lie93i] Otto Liess. Pseudodifferential operators. Lecture Notes in Mathematics, 1555:95–144, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084681/.

Liess:1993:SRT

[Lie93j] Otto Liess. Some results on transmission problems. Lecture Notes in Mathematics, 1555:309–345, 1993. CODEN LNMAA2. ISBN 3-540-57105-1 (print), 3-540-47905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084686/.

Liming:1990:TUR

[Lim90] Wu Liming. Un traitement unifié de la représentation des fonctionnelles de Wiener. (French) []. Lecture Notes in Mathematics, 1426:166–187, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083764/.

Lin:1990:CMT

[Lin90] Wen-Hsiung Lin. A conjecture of may on E_{2}-term of the may spectral sequence for the cohomology of the Steenrod algebra. Lecture Notes in Mathematics, 1418:15–52, 1990. CODEN LN-MAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083690/.

Linckelmann:1998:SEM

[Lin98] Markus Linckelmann. On stable equivalences of Morita type. Lecture Notes in Mathematics, 1685:221–232, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096377/.

Lions:1993:RSI

[Lio93a] J.-L. Lions. Remarks on systems with incompletely given initial data and incompletely given part of the boundary. Lecture Notes in Mathematics, 1540:239–250, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085483/.

Lions:1993:GSK

[Lio93b] P.-L. Lions. Global solutions of kinetic models and related questions. Lecture Notes in Mathematics, 1551:58–86, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090929/.

Liu:1996:BM

[Liu96a] Tai-Ping Liu. Back matter. Lecture Notes in Mathematics, 1640: ??, 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-49565-9/1.pdf.

Liu:1996:NHD

[Liu96b] Tai-Ping Liu. Nonlinear hyperbolic-dissipative partial differential equations. Lecture Notes in Mathematics, 1640:103–136, 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093708/.

Livorni:1990:ESS

[Liv90] Elvira Laura Livorni. On the existence of some surfaces. Lecture Notes in Mathematics, 1417:155–179, 1990. CODEN LN-MAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083340/.

Lempp:1990:PAU

[LL90] Steffen Lempp and Manuel Lerman. Priority arguments using iterated trees of strategies. Lecture Notes in Mathematics, 1432: 277–296, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086123/.

Liu:1991:AHM

[LL91] Zhixin Liu and Shanzhen Lu. Applications of Hörmander multiplier theorem to approximation in real Hardy spaces. Lecture Notes in Mathematics, 1494:119–129, 1991. CODEN LN-MAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087763/.

Lenstra:1993:DNF

[LL93] Arjen K. Lenstra and Hendrik W. Lenstra, Jr. The development of the number field sieve, volume 1554 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 126 (est.) pp. LCCN QA3 .L28 no. 1554.

URL http://link.springer.com/book/10.1007/BFb0091534;
 http://www.springerlink.com/content/978-3-540-47892-

Lenstra:1993:NFSb

[LLMP93] A. K. Lenstra, H. W. Lenstra, Jr., M. S. Manasse, and J. M. Pollard. The number field sieve. Lecture Notes in Mathematics, 1554:11–42, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091537/.

Lindsay:1990:DTA

[LM90] J. M. Lindsay and H. Maassen. Duality transform as *-algebraic isomorphism. Lecture Notes in Mathematics, 1442:247–250, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085517/.

Lindenstrauss:1991:GAF

[LM91] Joram Lindenstrauss and Vitali D. Milman, editors. Geometric Aspects of Functional Analysis: Israel Seminar (GAFA) 1989– 90, volume 1469 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1469. URL http:/

/link.springer.com/book/10.1007/BFb0089211;   http://
www.springerlink.com/content/978-3-540-47355-8.

Long:1991:WSI

[LN91] Ruilin Long and Fusheng Nie. Weighted Sobolev inequality and eigenvalue estimates of Schrödinger operators. Lecture Notes in Mathematics, 1494:131–141, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087765/.

Leandre:1997:IPC

[LN97] R. Léandre and J. R. Norris. Integration by parts and Cameron– Martin formulas for the free path space of a compact Riemannian manifold. Lecture Notes in Mathematics, 1655:16–23, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0119288/.

Li:1998:Ea

[LO98a] Ke-Zheng Li and Frans Oort. Examples on S_{g,1}. Lecture Notes in Mathematics, 1680:55–68, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095941/.

Li:1998:Eb

[LO98b] Ke-Zheng Li and Frans Oort. Examples on S_{g,d} (d > 1). Lecture Notes in Mathematics, 1680:84–86, 1998. CODEN LN-MAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095944/.

Li:1998:FTQ

[LO98c] Ke-Zheng Li and Frans Oort. Flag type quotients. Lecture Notes in Mathematics, 1680:19–23, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095935/.

Li:1998:I

[LO98d] Ke-Zheng Li and Frans Oort. Introduction. Lecture Notes in Mathematics, 1680:1–10, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095932/.

Li:1998:MRa

[LO98e] Ke-Zheng Li and Frans Oort. Main results on S_{g,1}. Lecture Notes in Mathematics, 1680:24–27, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095936/.

Li:1998:MRb

[LO98f] Ke-Zheng Li and Frans Oort. Main results on S_{g,d}. Lecture Notes in Mathematics, 1680:69–72, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095942/.

Li:1998:MRP

[LO98g] Ke-Zheng Li and Frans Oort. Moduli of rigid PFTQs of Dieudonné modules. Lecture Notes in Mathematics, 1680:39– 50, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095939/.

Li:1998:MSA

[LO98h] Ke-Zheng Li and Frans Oort. Moduli of Supersingular Abelian Varieties, volume 1680 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 116 pp. LCCN QA3 .L28 no. 1680. URL http:/

/link.springer.com/book/10.1007/BFb0095931;   http://
www.springerlink.com/content/978-3-540-69666-7.

Li:1998:PDM

[LO98i] Ke-Zheng Li and Frans Oort. PFTQs of Dieudonné modules over W. Lecture Notes in Mathematics, 1680:35–38, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095938/.

Li:1998:PAD

[LO98j] Ke-Zheng Li and Frans Oort. Prerequisites about Dieudonné modules. Lecture Notes in Mathematics, 1680:28–34, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095937/.

Li:1998:PPF

[LO98k] Ke-Zheng Li and Frans Oort. Proofs of the propositions on FTQs. Lecture Notes in Mathematics, 1680:73–83, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095943/.

Li:1998:STD

[LO98l] Ke-Zheng Li and Frans Oort. A scheme-theoretic definition of supersingularity. Lecture Notes in Mathematics, 1680:87–95, 1998.

CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095945/.

Li:1998:SCN

[LO98m] Ke-Zheng Li and Frans Oort. Some class numbers. Lecture Notes in Mathematics, 1680:51–54, 1998. CODEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095940/.

Li:1998:SPA

[LO98n] Ke-Zheng Li and Frans Oort. Some prerequisites about group schemes. Lecture Notes in Mathematics, 1680:16–18, 1998. CO-DEN LNMAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095934/.

Li:1998:SAV

[LO98o] Ke-Zheng Li and Frans Oort. Supersingular abelian varieties. Lecture Notes in Mathematics, 1680:11–15, 1998. CODEN LN-MAA2. ISBN 3-540-63923-3 (print), 3-540-69666-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095933/.

Lobry:1991:DB

[Lob91] Claude Lobry. Dynamic bifurcations. Lecture Notes in Mathematics, 1493:1–13, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085020/.

Loday:1992:IAT

[Lod92] Jean-Louis Loday. Introduction to algebraic K-theory and cyclic homology. Lecture Notes in Mathematics, 1491:31–54, 1992. CO-DEN LNMAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088878/.

Lomonaco:1992:NOC

[Lom92] Luciano Lomonaco. Normalized operations in cohomology. Lecture Notes in Mathematics, 1509:240–249, 1992. CODEN LN-MAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087514/.

Lorentz:1992:BPB

[Lor92a] Rudolph A. Lorentz. Basic properties of Birkhoff interpolation. Lecture Notes in Mathematics, 1516:9–22, 1992. CODEN LN-MAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088791/.

Lorentz:1992:DT

[Lor92b] Rudolph A. Lorentz. Decomposition theorems. Lecture Notes in Mathematics, 1516:62–71, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088794/.

Lorentz:1992:E

[Lor92c] Rudolph A. Lorentz. Examples. Lecture Notes in Mathematics, 1516:75–89, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088796/.

Lorentz:1992:I

[Lor92d] Rudolph A. Lorentz. Introduction. Lecture Notes in Mathematics, 1516:1–3, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088789/.

Lorentz:1992:KIB

[Lor92e] Rudolph A. Lorentz. Kergin interpolation via Birkhoff interpolation. Lecture Notes in Mathematics, 1516:162–170, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088801/.

Lorentz:1992:MBI

[Lor92f] Rudolph A. Lorentz. Multivariate Birkhoff Interpolation, volume 1516 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 192 pp. LCCN QA3 .L28 no. 1516. URL http:

//link.springer.com/book/10.1007/BFb0088788;   http:/
/www.springerlink.com/content/978-3-540-47300-8.

Lorentz:1992:R

[Lor92g] Rudolph A. Lorentz. Reduction. Lecture Notes in Mathematics, 1516:72–74, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088795/.

Lorentz:1992:SC

[Lor92h] Rudolph A. Lorentz. Shifts and coalescences. Lecture Notes in Mathematics, 1516:50–61, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088793/.

Lorentz:1992:SIS

[Lor92i] Rudolph A. Lorentz. Singular interpolation schemes. Lecture Notes in Mathematics, 1516:23–49, 1992. CODEN LN-MAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088792/.

Lorentz:1992:TS

[Lor92j] Rudolph A. Lorentz. A theorem of severi. Lecture Notes in Mathematics, 1516:156–161, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088800/.

Lorentz:1992:UHIa

[Lor92k] Rudolph A. Lorentz. Uniform Hermite interpolation of tensorproduct type. Lecture Notes in Mathematics, 1516:90–102, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088797/.

Lorentz:1992:UHIb

[Lor92l] Rudolph A. Lorentz. Uniform Hermite interpolation of type total degree. Lecture Notes in Mathematics, 1516:103–138, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088798/.

Lorentz:1992:UI

[Lor92m] Rudolph A. Lorentz. Univariate interpolation. Lecture Notes in Mathematics, 1516:4–8, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088790/.

Lorentz:1992:VD

[Lor92n] Rudolph A. Lorentz. Vandermonde determinants. Lecture Notes in Mathematics, 1516:139–155, 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088799/.

Lou:1991:WNI

[Lou91a] Hongwei Lou. Weighted norm inequalities for the restriction of Fourier transform to S^{n−1}. Lecture Notes in Mathematics, 1494: 130, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087764.

Louis:1991:MQB

[Lou91b] Alfred K. Louis. Mathematical questions of a biomagnetic imaging problem. Lecture Notes in Mathematics, 1497:126–132, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084513/.

Lacomba:1990:EES

[LP90a] Ernesto A. Lacomba and Felipe Peredo. Escape-equilibrium solutions in the repulsive Coulombian isosceles 3-body problem. Lecture Notes in Mathematics, 1416:178–191, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097471/.

Lindsay:1990:RPC

[LP90b] J. M. Lindsay and K. R. Parthasarathy. Rigidity of the Poisson convolution. Lecture Notes in Mathematics, 1442:251–262, 1990. CODEN LNMAA2. ISBN 3-540-53026-6 (print), 3-540-46311-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085518/.

Lloyd:1990:CCB

[LP90c] N. G. Lloyd and J. M. Pearson. Conditions for a centre and the bifurcation of limit cycles in a class of cubic systems. Lecture Notes in Mathematics, 1455:230–242, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085395/.

Lin:1991:BSP

[LP91] Peng Lin and Lizhong Peng. Besov spaces of Paley–Wiener type. Lecture Notes in Mathematics, 1494:95–112, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087761/.

Lluis-Puebla:1992:HAT

[LPLG^{+}92] Emilio Lluis-Puebla, Jean-Louis Loday, Henri Gillet, Christophe Soulé, and Victor Snaith. Higher Algebraic K-Theory: an overview, volume 1491 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 164 pp. LCCN QA3 .L28 no. 1491. URL http:

//link.springer.com/book/10.1007/BFb0088876;   http:/
/www.springerlink.com/content/978-3-540-46639-0.

Liu:1995:CMS

[LQ95a] Pei-Dong Liu and Min Qian. Characterization of measures satisfying entropy formula. Lecture Notes in Mathematics, 1606: 128–181, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094315/.

Liu:1995:ELE

[LQ95b] Pei-Dong Liu and Min Qian. Entropy and Lyapunov exponents of random diffeomorphisms. Lecture Notes in Mathematics, 1606: 22–44, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094310/.

Liu:1995:EEA

[LQ95c] Pei-Dong Liu and Min Qian. Estimation of entropy from above through Lyapunov exponents. Lecture Notes in Mathematics, 1606:45–54, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094311/.

Liu:1995:EEB

[LQ95d] Pei-Dong Liu and Min Qian. Estimation of entropy from below through Lyapunov exponents. Lecture Notes in Mathematics, 1606:91–108, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094313/.

Liu:1995:P

[LQ95e] Pei-Dong Liu and Min Qian. Preliminaries. Lecture Notes in Mathematics, 1606:1–21, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094309/.

Liu:1995:RPH

[LQ95f] Pei-Dong Liu and Min Qian. Random perturbations of hyperbolic attractors. Lecture Notes in Mathematics, 1606:182–206, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094316/.

Liu:1995:SET

[LQ95g] Pei-Dong Liu and Min Qian. Smooth Ergodic Theory of Random Dynamical Systems, volume 1606 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 221 pp. LCCN QA3 .L28 no. 1606.

URL http://link.springer.com/book/10.1007/BFb0094308;
 http://www.springerlink.com/content/978-3-540-49291-
7.

Liu:1995:SIM

[LQ95h] Pei-Dong Liu and Min Qian. Stable invariant manifolds of random diffeomorphisms. Lecture Notes in Mathematics, 1606:55– 90, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094312/.

Liu:1995:SFD

[LQ95i] Pei-Dong Liu and Min Qian. Stochastic flows of diffeomorphisms. Lecture Notes in Mathematics, 1606:109–127, 1995. CODEN LNMAA2. ISBN 3-540-60004-3 (print), 3-540-49291-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094314/.

Lugo:1990:HFS

[LR90] Adrian R. Lugo and Jorge G. Russo. Hamiltonian formulation of string theory and multiloop amplitudes in the operator context. Lecture Notes in Mathematics, 1451:148–162, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085068/.

Labesse:1990:CAG

[LS90a] Jean-Pierre Labesse and Joachim Schwermer, editors. Cohomology of Arithmetic Groups and Automorphic Forms: Proceedings of a Conference held in Luminy/Marseille, France, May 22–27 1989, volume 1447 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1447. URL http:/

/link.springer.com/book/10.1007/BFb0085723;   http://
www.springerlink.com/content/978-3-540-46876-9.

Laksov:1990:TCC

[LS90b] Dan Laksov and Robert Speiser. Transversality criteria in any characteristic. Lecture Notes in Mathematics, 1436:139–150, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084044/.

Loustaunau:1990:MC

[LS90c] Philippe Loustaunau and Jay Shapiro. Morita contexts. Lecture Notes in Mathematics, 1448:80–92, 1990. CODEN LN-MAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091253/.

Latushkin:1991:LSP

[LS91] Yu. D. Latushkin and A. M. Stepin. Linear skew-product flows and semigroups of weighted composition operators. Lecture Notes in Mathematics, 1486:98–111, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086661/.

Levin:1993:ECR

[LS93] A. L. Levin and E. B. Saff. Exact convergence rates for best L_{P} rational approximation to the signum function and for optimal quadrature in H^{P}. Lecture Notes in Mathematics, 1550: 98–109, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0117476/.

Lindenstrauss:1996:CBS

[LT96] Joram Lindenstrauss and Lior Tzafriri. Classical Banach Spaces I: Sequence Spaces, volume ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNMAA2. ISBN 3-540-60628-9 (paperback), 3-540-37732-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ????

pp. URL http://link.springer.com/book/10.1007/978-3-

540-37732-0. Reprint of Lecture Notes in Mathematics volume

Lumiste:1991:SPH

[Lum91] Ülo Lumiste. On submanifolds with parallel higher order fundamental form in Euclidean spaces. Lecture Notes in Mathematics, 1481:126–137, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083636/.

Lundell:1992:CTJ

[Lun92] Albert T. Lundell. Concise tables of James numbers and some homotopy of classical Lie groups and associated homogeneous spaces. Lecture Notes in Mathematics, 1509:250–272, 1992. CO-DEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087515/.

Lupton:1990:NCS

[Lup90] Gregory Lupton. Note on a conjecture of Stephen Halperin’s. Lecture Notes in Mathematics, 1440:148–163, 1990. CODEN LN-MAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084459/.

Lustig:1990:RDH

[Lus90] Martin Lustig. On the rank, the deficiency and the homological dimension of groups: The computation of a lower bound via Fox ideals. Lecture Notes in Mathematics, 1440:164–174, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084460/.

Laksov:1990:NCC

[LW90a] Dan Laksov and Karin Westin. Nullstellensätze; conjectures and counterexamples. Lecture Notes in Mathematics, 1420:178– 190, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083919/.

Langevin:1990:CAP

[LW90b] Michel Langevin and Michel Waldschmidt, editors. Cinquante Ans de Polynômes — Fifty Years of Polynomials: Proceedings of a Conference held in honour of Alain Durand at the Institut Henri Poincaré Paris, France, May 26–27, 1988, volume 1415 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1415.

URL http://link.springer.com/book/10.1007/BFb0084876;
 http://www.springerlink.com/content/978-3-540-46919-
3.

Leander:1990:RGI

[LW90c] Rémi Leander and Michel Weber. Une représentation gaussienne de l’indice d’un opérateur. (French) []. Lecture Notes in Mathematics, 1426:105–106, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083759.

Lyubashenko:1992:RIF

[Lyu92] Vladimir Lyubashenko. Real and imaginary forms of quantum groups. Lecture Notes in Mathematics, 1510:67–78, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101179/.

Lychagin:1992:DSS

[LZ92] V. V. Lychagin and L. V. Zil’bergleĭt. Duality in stable Spencer cohomologies. Lecture Notes in Mathematics, 1520:41–55, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084714/.

Maass:1991:SVD

[Maa91] Peter Maass. Singular value decompositions for Radon transforms. Lecture Notes in Mathematics, 1497:6–14, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084504/.

MacDonald:1991:CFP

[Mac91] John L. MacDonald. Concretely functorial programming. Lecture Notes in Mathematics, 1488:282–297, 1991. CODEN LN-MAA2. ISBN 3-540-54706-1 (print), 3-540-46435-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084227/.

Madych:1991:IRH

[Mad91] W. R. Madych. Image reconstruction in Hilbert space. Lecture Notes in Mathematics, 1497:15–45, 1991. CODEN LN-MAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084505/.

Maimeskul:1993:DAA

[Maĭ93a] V. V. Maĭmeskul. Degree of approximation of analytic functions by “near the best” polynomial approximants. Lecture Notes in Mathematics, 1550:160–163, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0117482/.

Maisonneuve:1993:PMN

[Mai93b] Bernard Maisonneuve. Processus de Markov: Naissance, retournement, régénération. (French) []. Lecture Notes in Mathematics, 1541:261–292, 1993. CODEN LNMAA2. ISBN 3-540-56622-8 (print), 3-540-47608-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084191/.

Maille:1994:UPM

[Mai94] Sophie Maille. Sur l’utilisation de processus de Markov dans le modèle d’Ising: attractivité et couplage. (French) []. Lecture Notes in Mathematics, 1583:195–235, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073848/.

Maini:1999:MMM

[Mai99] Philip K. Maini. Mathematical models in morphogenesis. Lecture Notes in Mathematics, 1714:151–189, 1999. CODEN LN-MAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092377/.

Majid:1992:RQG

[Maj92] Shahn Majid. Rank of quantum groups and braided groups in dual form. Lecture Notes in Mathematics, 1510:79–89, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101180/.

Malliavin:1991:TIT

[Mal91] Marie-Paule Malliavin, editor. Topics in Invariant Theory: Séminaire d’Algèbre P. Dubreil et M.-P. Malliavin 1989–1990 (40ème Année), volume 1478 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1478. URL http:/

/link.springer.com/book/10.1007/BFb0083499;   http://
www.springerlink.com/content/978-3-540-47592-7.

Malczak:1992:SLT

[Mal92] Jan Malczak. Some limit theorems for Markov operators and their applications. Lecture Notes in Mathematics, 1514:153–162, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097536/.

Manni:1992:MFF

[Man92] Riccardo Salvati Manni. Modular forms of the fourth degree. Lecture Notes in Mathematics, 1515:106–111, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098340/.

Mandal:1997:CI

[Man97a] Satya Mandal. Complete intersections. Lecture Notes in Mathematics, 1672:73–90, 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093563/.

Mandal:1997:EMP

[Man97b] Satya Mandal. Extended modules over polynomial rings. Lecture Notes in Mathematics, 1672:25–33, 1997. CODEN LN-MAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093560/.

Mandal:1997:I

[Man97c] Satya Mandal. Introduction. Lecture Notes in Mathematics, 1672: 1–2, 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093557.

Mandal:1997:MCR

[Man97d] Satya Mandal. Modules over commutative rings. Lecture Notes in Mathematics, 1672:35–62, 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093561/.

Mandal:1997:PMO

[Man97e] Satya Mandal. Patching modules and other preliminaries. Lecture Notes in Mathematics, 1672:13–23, 1997. CODEN LN-MAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093559/.

Mandal:1997:PFP

[Man97f] Satya Mandal. Préliminaires. (French) [Preliminaries]. Lecture Notes in Mathematics, 1672:3–12, 1997. CODEN LN-MAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093558/.

Mandal:1997:PMC

[Man97g] Satya Mandal. Projective Modules and Complete Intersections, volume 1672 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 114 pp. LCCN QA3 .L28 no. 1672. URL http:/

/link.springer.com/book/10.1007/BFb0093556;   http://
www.springerlink.com/content/978-3-540-69598-1.

Mandal:1997:TL

[Man97h] Satya Mandal. The techniques of Lindel. Lecture Notes in Mathematics, 1672:91–110, 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093564/.

Mandal:1997:TM

[Man97i] Satya Mandal. The theory of matrices. Lecture Notes in Mathematics, 1672:63–72, 1997. CODEN LNMAA2. ISBN 3-540-63564-5 (print), 3-540-69598-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093562/.

Mao:1991:LFA

[Mao91] Xuerong Mao. Lyapunov functions and almost sure exponential stability. Lecture Notes in Mathematics, 1486:165–177, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086666/.

Marcellini:1990:NCI

[Mar90a] Paolo Marcellini. Non convex integrals of the Calculus of Variations. Lecture Notes in Mathematics, 1446:16–57, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084930/.

Marcum:1990:HEC

[Mar90b] Howard J. Marcum. Homotopy equivalences in 2-categories. Lecture Notes in Mathematics, 1425:71–86, 1990. CODEN LN-MAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083832/.

Marin:1990:TCF

[Mar90c] Alexis Marin. Sur un théorème de Cheponkus. (French) []. Lecture Notes in Mathematics, 1420:191–193, 1990. CODEN LN-MAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083920/.

Maritz:1990:SAO

[Mar90d] P. Maritz. On some aspects of open multifunctions. Lecture Notes in Mathematics, 1419:110–130, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083103/.

Marar:1991:ECD

[Mar91] W. L. Marar. The Euler characteristic of the disentanglement of the image of a corank 1 map germ. Lecture Notes in Mathematics, 1462:212–220, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086384/.

Martio:1992:PDE

[Mar92] Olli Martio. Partial differential equations and quasiregular mappings. Lecture Notes in Mathematics, 1508:65–79, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094238/.

Markowich:1993:KMS

[Mar93a] Peter A. Markowich. Kinetic models for semiconductors. Lecture Notes in Mathematics, 1551:87–111, 1993. CODEN LN-MAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090930/.

Martelli:1993:CPB

[Mar93b] Mario Martelli. Continuation principles and boundary value problems. Lecture Notes in Mathematics, 1537:32–73, 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085075/.

Marshall:1996:ARS

[Mar96a] Murray A. Marshall. Abstract real spectra. Lecture Notes in Mathematics, 1636:99–131, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092703/.

Marshall:1996:FRT

[Mar96b] Murray A. Marshall. Fans and the representation theorem. Lecture Notes in Mathematics, 1636:37–59, 1996. CODEN LN-MAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092700/.

Marshall:1996:I

[Mar96c] Murray A. Marshall. Introduction. Lecture Notes in Mathematics, 1636:1–4, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092697/.

Marshall:1996:MGC

[Mar96d] Murray A. Marshall. Minimal generation of constructible sets. Lecture Notes in Mathematics, 1636:133–150, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092704/.

Marshall:1996:OF

[Mar96e] Murray A. Marshall. Orderings on fields. Lecture Notes in Mathematics, 1636:5–17, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092698/.

Marshall:1996:SCC

[Mar96f] Murray A. Marshall. P -structures, connected components and 0 the isotropy theorem. Lecture Notes in Mathematics, 1636: 61–82, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092701/.

Marshall:1996:RSR

[Mar96g] Murray A. Marshall. The real spectrum of a ring. Lecture Notes in Mathematics, 1636:83–97, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092702/.

Marshall:1996:SO

[Mar96h] Murray A. Marshall. Spaces of orderings. Lecture Notes in Mathematics, 1636:19–35, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092699/.

Marshall:1996:SOA

[Mar96i] Murray A. Marshall. Spaces of Orderings and Abstract Real Spectra, volume 1636 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 190 pp. LCCN QA3 .L28 no. 1636. URL http:/

/link.springer.com/book/10.1007/BFb0092696;   http://
www.springerlink.com/content/978-3-540-69996-5.

Marshall:1996:SRR

[Mar96j] Murray A. Marshall. Structure results and realization results. Lecture Notes in Mathematics, 1636:151–181, 1996. CODEN LNMAA2. ISBN 3-540-61729-9 (print), 3-540-69996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092705/.

Marker:1998:ZG

[Mar98] David Marker. Zariski geometries. Lecture Notes in Mathematics, 1696:107–128, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-68521-0_7/.

Martinelli:1999:LGD

[Mar99] Fabio Martinelli. Lectures on Glauber dynamics for discrete spin models. Lecture Notes in Mathematics, 1717:93–191, 1999. CODEN LNMAA2. ISBN 3-540-66593-5 (print), 3-540-48115-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48115-7_2/.

Massey:1995:AS

[Mas95a] David B. Massey. Aligned singularities. Lecture Notes in Mathematics, 1615:75–80, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094417/.

Massey:1995:CMF

[Mas95b] David B. Massey. Constancy of the Milnor fibrations. Lecture Notes in Mathematics, 1615:86–91, 1995. CODEN LN-MAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094419/.

Massey:1995:DBP

[Mas95c] David B. Massey. Definitions and basic properties. Lecture Notes in Mathematics, 1615:8–30, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094411/.

Massey:1995:EE

[Mas95d] David B. Massey. Elementary examples. Lecture Notes in Mathematics, 1615:31–36, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094412/.

Massey:1995:GIF

[Mas95e] David B. Massey. Generalized Lê–Iomdine formulas. Lecture Notes in Mathematics, 1615:42–60, 1995. CODEN LN-MAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094414/.

Massey:1995:HDM

[Mas95f] David B. Massey. A handle decomposition of the Milnor fibre. Lecture Notes in Mathematics, 1615:37–41, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094413/.

Massey:1995:I

[Mas95g] David B. Massey. Introduction. Lecture Notes in Mathematics, 1615:1–7, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094410/.

Massey:1995:CHS

[Mas95h] David B. Massey. Lê Cycles and Hypersurface Singularities, volume 1615 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 131 pp. LCCN QA3 .L28 no. 1615. URL http:/

/link.springer.com/book/10.1007/BFb0094409;   http://
www.springerlink.com/content/978-3-540-45521-9.

Massey:1995:NHA

[Mas95i] David B. Massey. Lê numbers and hyperplane arrangements. Lecture Notes in Mathematics, 1615:61–67, 1995. CODEN LN-MAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094415/.

Massey:1995:OCC

[Mas95j] David B. Massey. Other characterizations of the Lê cycles. Lecture Notes in Mathematics, 1615:92–104, 1995. CODEN LN-MAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094420/.

Massey:1995:SS

[Mas95k] David B. Massey. Suspending singularities. Lecture Notes in Mathematics, 1615:81–85, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094418/.

Massey:1995:TC

[Mas95l] David B. Massey. Thom’s a_{f} condition. Lecture Notes in Mathematics, 1615:68–74, 1995. CODEN LNMAA2. ISBN 3-540-60395-6 (print), 3-540-45521-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094416/.

Matano:1990:SSN

[Mat90a] Hiroshi Matano. Singular solutions of a nonlinear elliptic equation and an infinite dimensional dynamical system. Lecture Notes in Mathematics, 1450:64–87, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084899/.

Matone:1990:CFT

[Mat90b] Marco Matone. Conformal field theory, real weight differentials and KdV equation in higher genus. Lecture Notes in Mathematics, 1451:163–175, 1990. CODEN LNMAA2. ISBN 3-540-53286-2 (print), 3-540-46813-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085069/.

Matsumoto:1990:VDZ

[Mat90c] Kohji Matsumoto. Value-distribution of zeta-functions. Lecture Notes in Mathematics, 1434:178–187, 1990. CODEN LN-MAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097134/.

Matsuura:1993:NCC

[Mat93] Shigetake Matsuura. On non-convex curves of constant angle. Lecture Notes in Mathematics, 1540:251–268, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085484/.

Mathieu:1998:QIL

[Mat98] P. Mathieu. Quand l’inégalité log-Sobolev implique l’inégalité de trou spectral. (French) []. Lecture Notes in Mathematics, 1686: 30–35, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101748/.

Maubach:1990:SPH

[Mau90] J. Maubach. On the sparsity patterns of hierarchical finite element matrices. Lecture Notes in Mathematics, 1457:79–104, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090903/.

Mawhin:1993:TDB

[Maw93] J. Mawhin. Topological degree and boundary value problems for nonlinear differential equations. Lecture Notes in Mathematics, 1537:74–142, 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085076/.

May:1990:WEQ

[May90] J. P. May. Weak equivalences and quasifibrations. Lecture Notes in Mathematics, 1425:91–101, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083834/.

Mazet:1997:CSG

[Maz97] Olivier Mazet. Classification des semi-groupes de diffusion sur IR associés à une famille de polynômes orthogonaux. (French) []. Lecture Notes in Mathematics, 1655:40–53, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119290/.

McCleary:1990:HTC

[McC90a] John McCleary. Homotopy theory and closed geodesics. Lecture Notes in Mathematics, 1418:86–94, 1990. CODEN LN-MAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083695/.

McConnell:1990:QGF

[McC90b] J. C. McConnell. Quantum groups, filtered rings and Gelfand– Kirillov dimension. Lecture Notes in Mathematics, 1448:139– 147, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091258/.

McCullough:1990:TAA

[McC90c] Darryl McCullough. Topological and algebraic automorphisms of 3-manifolds. Lecture Notes in Mathematics, 1425:102–113, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083835/.

McCutcheon:1999:DRT

[McC99a] Randall McCutcheon. Density Ramsey theory. Lecture Notes in Mathematics, 1722:78–110, 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093963/.

McCutcheon:1999:EME

[McC99b] Randall McCutcheon. Elemental Methods in Ergodic Ramsey Theory, volume 1722 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 160 pp. LCCN QA3 .L28 no. 1722. URL http:/

/link.springer.com/book/10.1007/BFb0093959;   http://
www.springerlink.com/content/978-3-540-46699-4.

McCutcheon:1999:IRT

[McC99c] Randall McCutcheon. Infinitary Ramsey theory. Lecture Notes in Mathematics, 1722:40–77, 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093962/.

McCutcheon:1999:I

[McC99d] Randall McCutcheon. Introduction. Lecture Notes in Mathematics, 1722:1–4, 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093960/.

McCutcheon:1999:RTT

[McC99e] Randall McCutcheon. Ramsey theory and topological dynamics. Lecture Notes in Mathematics, 1722:5–39, 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093961/.

McCutcheon:1999:TER

[McC99f] Randall McCutcheon. Three ergodic Roth theorems. Lecture Notes in Mathematics, 1722:111–135, 1999. CODEN LN-MAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093964/.

McCutcheon:1999:TST

[McC99g] Randall McCutcheon. Two Szemerédi theorems. Lecture Notes in Mathematics, 1722:136–152, 1999. CODEN LNMAA2. ISBN 3-540-66809-8 (print), 3-540-46699-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093965/.

McGill:1992:GTQ

[McG92] Paul McGill. Generalised transforms, quasi-diffusions, and Désiré André’s equation. Lecture Notes in Mathematics, 1526:234– 247, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084325/.

McIntosh:1990:SRP

[McI90] Alan McIntosh. The square root problem for elliptic operators — a survey. Lecture Notes in Mathematics, 1450:122–140, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084902/.

McKenzie:1990:SIB

[McK90] Ralph McKenzie. Some interactions between group theory and the general theory of algebras. Lecture Notes in Mathematics, 1456: 32–48, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100729/.

Miret:1990:SDC

[MD90] J. M. Miret and S. Xambó Descamps. On Schubert’s degenerations of cuspidal plane cubics. Lecture Notes in Mathematics, 1436:189–214, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084046/.

Meleard:1996:ABS

[Mél96] Sylvie Méléard. Asymptotic behaviour of some interacting particle systems; McKean–Vlasov and Boltzmann models. Lecture Notes in Mathematics, 1627:42–95, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093177/.

Merel:1994:UFE

[Mer94] Loïc Merel. Universal Fourier expansions of modular forms. Lecture Notes in Mathematics, 1585:59–94, 1994. CODEN LN-MAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074110/.

Metsch:1991:BPR

[Met91a] Klaus Metsch. Basic properties and results of (n + 1, 1)-designs. Lecture Notes in Mathematics, 1490:15–20, 1991. CODEN LN-MAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083248/.

Metsch:1991:DBP

[Met91b] Klaus Metsch. Definition and basic properties of linear spaces. Lecture Notes in Mathematics, 1490:1–8, 1991. CODEN LN-MAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083246/.

Metsch:1991:DWC

[Met91c] Klaus Metsch. The Dowling Wilson conjecture. Lecture Notes in Mathematics, 1490:181–187, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083262/.

Metsch:1991:EDP

[Met91d] Klaus Metsch. Embedding (n+1, 1)-designs into projective planes. Lecture Notes in Mathematics, 1490:43–60, 1991. CODEN LN-MAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083251/.

Metsch:1991:HS

[Met91e] Klaus Metsch. A hypothetical structure. Lecture Notes in Mathematics, 1490:94–105, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083255/.

Metsch:1991:LSW

[Met91f] Klaus Metsch. Linear spaces in which every long line meets only few lines. Lecture Notes in Mathematics, 1490:150–160, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083260/.

Metsch:1991:LSP

[Met91g] Klaus Metsch. Linear spaces with n^{2} +n+1 points. Lecture Notes in Mathematics, 1490:86–93, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083254/.

Metsch:1991:LSL

[Met91h] Klaus Metsch. Linear spaces with n^{2} + n + 2 lines. Lecture Notes in Mathematics, 1490:106–117, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083256/.

Metsch:1991:LSFa

[Met91i] Klaus Metsch. Linear spaces with few lines. Lecture Notes in Mathematics, 1490:31–42, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083250/.

Metsch:1991:LSFb

[Met91j] Klaus Metsch. Linear Spaces with Few Lines, volume 1490 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 196 pp. LCCN QA3 .L28 no. 1490. URL http:/

/link.springer.com/book/10.1007/BFb0083245;   http://
www.springerlink.com/content/978-3-540-46444-0.

Metsch:1991:LBN

[Met91k] Klaus Metsch. Lower bounds for the number of lines. Lecture Notes in Mathematics, 1490:9–14, 1991. CODEN LN-MAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083247/.

Metsch:1991:NEC

[Met91l] Klaus Metsch. The non-existence of certain (7, 1)-designs and determination of A(5) and A(6). Lecture Notes in Mathematics, 1490:131–140, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083258/.

Metsch:1991:OBE

[Met91m] Klaus Metsch. An optimal bound for embedding linear spaces into projective planes. Lecture Notes in Mathematics, 1490:61– 73, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083252/.

Metsch:1991:PD

[Met91n] Klaus Metsch. Points of degree n. Lecture Notes in Mathematics, 1490:21–30, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083249/.

Metsch:1991:PDA

[Met91o] Klaus Metsch. Points of degree n and another characterization of the linear spaces L(n, d). Lecture Notes in Mathematics, 1490:

118–130, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083257/.

Metsch:1991:RGT

[Met91p] Klaus Metsch. A result on graph theory with an application to linear spaces. Lecture Notes in Mathematics, 1490:141–149, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083259/.

Metsch:1991:FIP

[Met91q] Klaus Metsch. s-Fold inflated projective planes. Lecture Notes in Mathematics, 1490:161–180, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083261/.

Metsch:1991:TT

[Met91r] Klaus Metsch. The theorem of totten. Lecture Notes in Mathematics, 1490:74–85, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083253/.

Metsch:1991:UE

[Met91s] Klaus Metsch. Uniqueness of embeddings. Lecture Notes in Mathematics, 1490:188–191, 1991. CODEN LNMAA2. ISBN 3-540-54720-7 (print), 3-540-46444-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083263/.

Meuser:1990:FEI

[Meu90] Diane Meuser. On a functional equation of Igusa’s local zeta function. Lecture Notes in Mathematics, 1454:309–313, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091145/.

Meyer:1990:OFM

[Mey90] Yves Meyer. Ondelettes, filtres miroirs en quadrature et traitement numérique de l’image. (French) []. Lecture Notes in Mathematics, 1438:14–25, 1990. CODEN LNMAA2. ISBN 3-540-52932-2 (print), 3-540-47179-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083512/.

Meyer:1991:ADB

[Mey91a] P.-A. Meyer. Application du “Bébé Fock” au modèle d’Ising. (French) []. Lecture Notes in Mathematics, 1485:52–60, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100846/.

Meyer:1991:MSP

[Mey91b] P.-A. Meyer. Sur la méthode de L. Schwartz pour les é.d.s.. (French) []. Lecture Notes in Mathematics, 1485:108–112, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100850/.

Meyer:1993:SPE

[Mey93a] P.-A. Meyer. Les systèmes-produits et l’espace de Fock d’après W. Arveson. (French) []. Lecture Notes in Mathematics, 1557: 106–113, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087968/.

Meyer:1993:RMO

[Mey93b] P.-A. Meyer. Représentation de martingales d’opérateurs d’après Parthasarathy–Sinha. (French) []. Lecture Notes in Mathematics, 1557:97–105, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087967/.

Meyer:1993:RFC

[Mey93c] P.-A. Meyer. Représentation des fonctions conditionnellement de type positif d’après V. P. Belavkin. (French) []. Lecture Notes in Mathematics, 1557:114–121, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087969/.

Meyer:1993:BM

[Mey93d] Paul-André Meyer. Back matter. Lecture Notes in Mathematics, 1538:201–293, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21558-6/1.pdf.

Meyer:1993:FS

[Mey93e] Paul-André Meyer. Fock space (1). Lecture Notes in Mathematics, 1538:55–96, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21558-6_4/.

Meyer:1993:FSM

[Mey93f] Paul-André Meyer. Fock space (2): Multiple Fock spaces. Lecture Notes in Mathematics, 1538:97–116, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21558-6_5/.

Meyer:1993:FM

[Mey93g] Paul-André Meyer. Front matter. Lecture Notes in Mathematics, 1538:N2–X, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21558-6/1.pdf.

Meyer:1993:HO

[Mey93h] Paul-André Meyer. The harmonic oscillator. Lecture Notes in Mathematics, 1538:41–54, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21558-6_3/.

Meyer:1993:II

[Mey93i] Paul-André Meyer. Independent increments. Lecture Notes in Mathematics, 1538:187–200, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21558-6_7/.

Meyer:1993:NCP

[Mey93j] Paul-André Meyer. Non-commutative probability. Lecture Notes in Mathematics, 1538:1–10, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21558-6_1/.

Meyer:1993:QPP

[Mey93k] Paul-André Meyer. Quantum Probability for Probabilists, volume 1538 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 287 pp. LCCN QA3 .L28 no. 1538. URL http://link.

springer.com/book/10.1007/978-3-662-21558-6;   http:/
/www.springerlink.com/content/978-3-662-21558-6.

Meyer:1993:S

[Mey93l] Paul-André Meyer. Spin. Lecture Notes in Mathematics, 1538: 11–40, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21558-6_2/.

Meyer:1993:SCF

[Mey93m] Paul-André Meyer. Stochastic calculus in Fock space. Lecture Notes in Mathematics, 1538:117–186, 1993. CODEN LNMAA2. ISBN 3-540-56476-4 (print), 3-662-21558-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21558-6_6/.

Meyer:1994:TDM

[Mey94] P.-A. Meyer. Sur une transformation du mouvement brownien due à Jeulin et Yor. (French) []. Lecture Notes in Mathematics, 1583:98–101, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073836/.

Meyer:1995:FS

[Mey95a] Paul-André Meyer. Fock space (1). Lecture Notes in Mathematics, 1538:57–102, 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084705/.

Meyer:1995:FSM

[Mey95b] Paul-André Meyer. Fock space (2): Multiple Fock spaces. Lecture Notes in Mathematics, 1538:103–124, 1995. CODEN LN-MAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084706/.

Meyer:1995:HO

[Mey95c] Paul-André Meyer. The harmonic oscillator. Lecture Notes in Mathematics, 1538:43–56, 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084704/.

Meyer:1995:II

[Mey95d] Paul-André Meyer. Independent increments. Lecture Notes in Mathematics, 1538:195–208, 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084708/.

Meyer:1995:NCP

[Mey95e] Paul-André Meyer. Non-commutative probability. Lecture Notes in Mathematics, 1538:1–11, 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084702/.

Meyer:1995:QPP

[Mey95f] Paul-André Meyer. Quantum Probability for Probabilists, volume 1538 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 208 (est.) pp. LCCN QA3 .L28 no. 1538. URL http:/

/link.springer.com/book/10.1007/BFb0084701;   http://
www.springerlink.com/content/978-3-540-36959-2.

Meyer:1995:S

[Mey95g] Paul-André Meyer. Spin. Lecture Notes in Mathematics, 1538: 13–42, 1995. CODEN LNMAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084703/.

Meyer:1995:SCF

[Mey95h] Paul-André Meyer. Stochastic calculus in Fock space. Lecture Notes in Mathematics, 1538:125–194, 1995. CODEN LN-MAA2. ISBN 3-540-60270-4 (print), 3-540-36959-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084707/.

Meyer:1999:CC

[Mey99a] Kenneth R. Meyer. Central configurations. Lecture Notes in Mathematics, 1719:39–49, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094681/.

Meyer:1999:CO

[Mey99b] Kenneth R. Meyer. Comet orbits. Lecture Notes in Mathematics, 1719:111–118, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094687/.

Meyer:1999:EP

[Mey99c] Kenneth R. Meyer. The elliptic problem. Lecture Notes in Mathematics, 1719:129–137, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094689/.

Meyer:1999:ECM

[Mey99d] Kenneth R. Meyer. Equations of celestial mechanics. Lecture Notes in Mathematics, 1719:9–18, 1999. CODEN LN-MAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094679/.

Meyer:1999:HS

[Mey99e] Kenneth R. Meyer. Hamiltonian systems. Lecture Notes in Mathematics, 1719:19–37, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094680/.

Meyer:1999:HLE

[Mey99f] Kenneth R. Meyer. Hill’s lunar equations. Lecture Notes in Mathematics, 1719:119–127, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094688/.

Meyer:1999:I

[Mey99g] Kenneth R. Meyer. Introduction. Lecture Notes in Mathematics, 1719:1–8, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094678/.

Meyer:1999:LO

[Mey99h] Kenneth R. Meyer. Lunar orbits. Lecture Notes in Mathematics, 1719:105–110, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094686/.

Meyer:1999:PSB

[Mey99i] Kenneth R. Meyer. Periodic Solutions of the N-Body Problem, volume 1719 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 144 pp. LCCN QA3 .L28 no. 1719. URL http:

//link.springer.com/book/10.1007/BFb0094677;   http:/
/www.springerlink.com/content/978-3-540-48073-0.

Meyer:1999:RP

[Mey99j] Kenneth R. Meyer. The restricted problem. Lecture Notes in Mathematics, 1719:91–103, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094685/.

Meyer:1999:SO

[Mey99k] Kenneth R. Meyer. Satellite orbits. Lecture Notes in Mathematics, 1719:87–90, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094684/.

Meyer:1999:SIR

[Mey99l] Kenneth R. Meyer. Symmetries, integrals, and reduction. Lecture Notes in Mathematics, 1719:51–70, 1999. CODEN LN-MAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094682/.

Meyer:1999:TPS

[Mey99m] Kenneth R. Meyer. Theory of periodic solutions. Lecture Notes in Mathematics, 1719:71–86, 1999. CODEN LNMAA2. ISBN 3-540-66630-3 (print), 3-540-48073-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094683/.

Mather:1994:AMO

[MF94a] John N. Mather and Giovanni Forni. Action minimizing orbits in Hamiltonian systems. Lecture Notes in Mathematics, 1589: 92–186, 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074076/.

Mather:1994:BM

[MF94b] John N. Mather and Giovanni Forni. Back matter. Lecture Notes in Mathematics, 1589:??, 1994. CODEN LNMAA2. ISBN 3-540-58416-1 (print), 3-540-48782-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-48782-1/1.pdf.

Miclo:1992:RSS

[Mic92] Laurent Miclo. Recuit simulé sans potentiel sur un ensemble fini. (French) []. Lecture Notes in Mathematics, 1526:47–60, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084309/.

Micchelli:1993:URE

[Mic93] Charles A. Micchelli. Using the refinement equation for the construction of pre-wavelets IV: Cube splines and elliptic splines united. Lecture Notes in Mathematics, 1550:62–70, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117474/.

Miclo:1997:RHE

[Mic97] Laurent Miclo. Remarques sur l’hypercontractivité et l’évolution de l’entropie pour des chaînes de Markov finies. (French) []. Lecture Notes in Mathematics, 1655:136–167, 1997. CODEN LN-MAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0119300/.

Miclo:1998:TSB

[Mic98] Laurent Miclo. Trous spectraux à basse température: un contreexemple à un comportement asymptotique escompté. (French) []. Lecture Notes in Mathematics, 1686:36–55, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101749/.

Mielke:1991:CSW

[Mie91a] Alexander Mielke. Capillarity surface waves. Lecture Notes in Mathematics, 1489:103–108, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097553/.

Mielke:1991:EVP

[Mie91b] Alexander Mielke. Elliptic variational problems on cylindrical domains. Lecture Notes in Mathematics, 1489:95–102, 1991. CO-DEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097552/.

Mielke:1991:HLF

[Mie91c] Alexander Mielke. Hamiltonian and Lagrangian Flows on Center Manifolds: with Applications to Elliptic Variational Problems, volume 1489 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 140 pp. LCCN QA3 .L28 no. 1489. URL http:

//link.springer.com/book/10.1007/BFb0097544;   http:/
/www.springerlink.com/content/978-3-540-46441-9.

Mielke:1991:HFC

[Mie91d] Alexander Mielke. Hamiltonian flows on center manifolds. Lecture Notes in Mathematics, 1489:27–40, 1991. CODEN LN-MAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097548/.

Mielke:1991:HSS

[Mie91e] Alexander Mielke. Hamiltonian systems with symmetries. Lecture Notes in Mathematics, 1489:41–59, 1991. CODEN LN-MAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097549/.

Mielke:1991:I

[Mie91f] Alexander Mielke. Introduction. Lecture Notes in Mathematics, 1489:1–6, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097545/.

Mielke:1991:LS

[Mie91g] Alexander Mielke. Lagrangian systems. Lecture Notes in Mathematics, 1489:61–83, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097550/.

Mielke:1991:LT

[Mie91h] Alexander Mielke. The linear theory. Lecture Notes in Mathematics, 1489:17–26, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097547/.

Mielke:1991:NSb

[Mie91i] Alexander Mielke. Necking of strips. Lecture Notes in Mathematics, 1489:109–119, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097554/.

Mielke:1991:NSa

[Mie91j] Alexander Mielke. Nonautonomous systems. Lecture Notes in Mathematics, 1489:85–92, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097551/.

Mielke:1991:NBF

[Mie91k] Alexander Mielke. Notations and basic facts on center manifolds. Lecture Notes in Mathematics, 1489:9–16, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097546/.

Mielke:1991:SVP

[Mie91l] Alexander Mielke. Saint-Venant’s problem. Lecture Notes in Mathematics, 1489:121–131, 1991. CODEN LNMAA2. ISBN 3-540-54710-X (print), 3-540-46441-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097555/.

Mignotte:1990:PLS

[Mig90] Maurice Mignotte. Polynômes et lemme de Siegel. (French) []. Lecture Notes in Mathematics, 1415:160–166, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084885/.

Miller:1990:PHC

[Mil90a] Andy Miller. Projecting homeomorphisms from covering spaces. Lecture Notes in Mathematics, 1425:114–132, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083836/.

Miller:1990:JKP

[Mil90b] Haynes Miller. On Jones’s Kahn–Priddy theorem. Lecture Notes in Mathematics, 1418:210–218, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083705/.

Milman:1991:SAD

[Mil91] V. Milman. Some applications of duality relations. Lecture Notes in Mathematics, 1469:13–40, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089213/.

Milman:1994:ILE

[Mil94a] Mario Milman. K/J inequalities and limiting embedding theorems. Lecture Notes in Mathematics, 1580:35–41, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073501/.

Milman:1994:BET

[Mil94b] Mario Milman. Background on extrapolation theory. Lecture Notes in Mathematics, 1580:7–34, 1994. CODEN LN-MAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073500/.

Milman:1994:BEL

[Mil94c] Mario Milman. Bilinear extrapolation and a limiting case of a theorem by Cwikel. Lecture Notes in Mathematics, 1580:59–73, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073503/.

Milman:1994:CMA

[Mil94d] Mario Milman. Calculations with the Δ method and applications. Lecture Notes in Mathematics, 1580:43–57, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073502/.

Milman:1994:ECR

[Mil94e] Mario Milman. Estimates for commutators in real interpolation. Lecture Notes in Mathematics, 1580:95–126, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073505/.

Milman:1994:EOD

[Mil94f] Mario Milman. Extrapolation and Optimal Decompositions: with Applications to Analysis, volume 1580 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 162 pp. LCCN QA3 .L28 no. 1580.

URL http://link.springer.com/book/10.1007/BFb0073498;
 http://www.springerlink.com/content/978-3-540-48439-
4.

Milman:1994:ERA

[Mil94g] Mario Milman. Extrapolation, reiteration, and applications. Lecture Notes in Mathematics, 1580:75–93, 1994. CODEN LN-MAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073504/.

Milman:1994:I

[Mil94h] Mario Milman. Introduction. Lecture Notes in Mathematics, 1580: 1–5, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073499/.

Milman:1994:ODS

[Mil94i] Mario Milman. Optimal decompositions, scales, and Nash–Moser iteration. Lecture Notes in Mathematics, 1580:139–147, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073508/.

Milman:1994:SIT

[Mil94j] Mario Milman. Sobolev imbedding theorems and extrapolation of infinitely many operators. Lecture Notes in Mathematics, 1580: 127–130, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073506/.

Milman:1994:SRE

[Mil94k] Mario Milman. Some remarks on extrapolation spaces and abstract parabolic equations. Lecture Notes in Mathematics, 1580: 131–137, 1994. CODEN LNMAA2. ISBN 3-540-58081-6 (print), 3-540-48439-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073507/.

Mimura:1990:HTR

[Mim90] Mamoru Mimura, editor. Homotopy Theory and Related Topics: Proceedings of the International Conference held at Kinosaki, Japan, August 19–24, 1988, volume 1418 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1418.

URL http://link.springer.com/book/10.1007/BFb0083687;
 http://www.springerlink.com/content/978-3-540-46938-
4.

Minda:1990:BMC

[Min90] David Minda. The Bloch and Marden constants. Lecture Notes in Mathematics, 1435:131–142, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087903/.

Mishchenko:1992:SSN

[Mis92] A. S. Mishchenko. Stationary solutions of nonlinear stochastic equations. Lecture Notes in Mathematics, 1520:217–236, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084723/.

Mischaikow:1995:CIT

[Mis95] Konstantin Mischaikow. Conley index theory. Lecture Notes in Mathematics, 1609:119–207, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095240/.

Mitrea:1994:BCA

[Mit94a] Marius Mitrea. The L^{2} boundedness of Clifford algebra valued singular integral operators. Lecture Notes in Mathematics, 1575: 42–59, 1994. CODEN LNMAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073559/.

Mitrea:1994:ATH

[Mit94b] Marius Mitrea. Applications to the theory of harmonic functions. Lecture Notes in Mathematics, 1575:87–105, 1994. CODEN LNMAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073561/.

Mitrea:1994:CA

[Mit94c] Marius Mitrea. Clifford algebras. Lecture Notes in Mathematics, 1575:1–15, 1994. CODEN LNMAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073557/.

Mitrea:1994:CWS

[Mit94d] Marius Mitrea. Clifford Wavelets, Singular Integrals, and Hardy Spaces, volume 1575 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 116 pp. LCCN QA3 .L28 no. 1575. URL http:/

/link.springer.com/book/10.1007/BFb0073556;   http://
www.springerlink.com/content/978-3-540-48379-3.

Mitrea:1994:CCW

[Mit94e] Marius Mitrea. Constructions of Clifford wavelets. Lecture Notes in Mathematics, 1575:16–41, 1994. CODEN LNMAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073558/.

Mitrea:1994:HSM

[Mit94f] Marius Mitrea. Hardy spaces of monogenic functions. Lecture Notes in Mathematics, 1575:60–86, 1994. CODEN LN-MAA2. ISBN 3-540-57884-6 (print), 3-540-48379-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073560/.

Mizera:1992:GPO

[Miz92] Ivan Mizera. Generic properties of one-dimensional dynamical systems. Lecture Notes in Mathematics, 1514:163–173, 1992. CO-DEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0097537/.

Mejia:1990:HGS

[MM90] Diego Mejía and David Minda. Hyperbolic geometry in spherically k-convex regions. Lecture Notes in Mathematics, 1435:117– 129, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087902/.

Martinez:1991:CAS

[MM91a] A. Martínez and F. Milán. Convex affine surfaces with constant affine mean curvature. Lecture Notes in Mathematics, 1481:138– 144, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083637/.

Mond:1991:STA

[MM91b] David Mond and James Montaldi, editors. Singularity Theory and its Applications: Warwick 1989, Part I: Geometric Aspects of Singularities, volume 1462 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1462. URL http:/

/link.springer.com/book/10.1007/BFb0086370;   http://
www.springerlink.com/content/978-3-540-47060-1.

Majer:1997:CEC

[MM97] Pietro Majer and Maria Elvira Mancino. A counter-example concerning a condition of Ogawa integrability. Lecture Notes in Mathematics, 1655:198–206, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0119304/.

Muller:1990:LPP

[MN90] Wolfgang Müller and Werner Georg Nowak. Lattice points in planar domains: Applications of Huxley’s ‘discrete Hardy– Littlewood method’. Lecture Notes in Mathematics, 1452:139– 164, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096987/.

Miyano:1991:MSH

[MN91] Toshiki Miyano and Junjiro Noguchi. Moduli spaces of harmonic and holomorphic mappings and Diophantine geometry. Lecture Notes in Mathematics, 1468:227–253, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086196/.

Mebkhout:1990:CRG

[MNM90] Z. Mebkhout and L. Narvaez-Macarro. Sur les coéfficients de de Rham–Grothendieck des variétés algébriques. (French) []. Lecture Notes in Mathematics, 1454:267–308, 1990. CODEN LN-MAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091144/.

Mikusinski:1990:SHI

[MO90] Piotr Mikusiński and Krzysztof Ostaszewski. The space of Henstock integrable functions II. Lecture Notes in Mathematics, 1419: 136–149, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083105/.

Morgan:1993:BF

[MO93a] John W. Morgan and Kieran G. O’Grady. The blow-up formula. Lecture Notes in Mathematics, 1545:112–166, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086771/.

Morgan:1993:CMS

[MO93b] John W. Morgan and Kieran G. O’Grady. Certain moduli spaces for bundles on elliptic surfaces with p_{g} = 1. Lecture Notes in Mathematics, 1545:57–98, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086769/.

Morgan:1993:DTC

[MO93c] John W. Morgan and Kieran G. O’Grady. Differential Topology of Complex Surfaces: Elliptic Surfaces with p_{g} = 1: Smooth Classification, volume 1545 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 224 pp. LCCN QA3 .L28 no. 1545. URL http:/

/link.springer.com/book/10.1007/BFb0086765;   http://
www.springerlink.com/content/978-3-540-47628-3.

Morgan:1993:Ib

[MO93d] John W. Morgan and Kieran G. O’Grady. Identification of δ_{3,r}(S, H) with γ_{3}(S). Lecture Notes in Mathematics, 1545:33– 56, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086768/.

Morgan:1993:Ia

[MO93e] John W. Morgan and Kieran G. O’Grady. Introduction. Lecture Notes in Mathematics, 1545:1–11, 1993. CODEN LN-MAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086766/.

Morgan:1993:PT

[MO93f] John W. Morgan and Kieran G. O’Grady. The proof of theorem 1.1.1. Lecture Notes in Mathematics, 1545:167–210, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086772/.

Morgan:1993:RCI

[MO93g] John W. Morgan and Kieran G. O’Grady. Representatives for classes in the image of the ν-map. Lecture Notes in Mathematics, 1545:99–111, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086770/.

Morgan:1993:UPA

[MO93h] John W. Morgan and Kieran G. O’Grady. Unstable polynomials of algebraic surfaces. Lecture Notes in Mathematics, 1545: 12–32, 1993. CODEN LNMAA2. ISBN 3-540-56674-0 (print), 3-540-47628-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086767/.

Moerdijk:1995:BTT

[Moe95a] Izak Moerdijk. Background in topos theory. Lecture Notes in Mathematics, 1616:5–19, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094443/.

Moerdijk:1995:CSCa

[Moe95b] Izak Moerdijk. Classifying spaces and classifying topoi. Lecture Notes in Mathematics, 1616:95, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094447.

Moerdijk:1995:CSCb

[Moe95c] Izak Moerdijk. Classifying Spaces and Classifying Topoi, volume 1616 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 94 pp. LCCN QA3 .L28 no. 1616. URL http:

//link.springer.com/book/10.1007/BFb0094441;   http:/
/www.springerlink.com/content/978-3-540-44912-6.

Moerdijk:1995:CTa

[Moe95d] Izak Moerdijk. Classifying topoi. Lecture Notes in Mathematics, 1616:21–55, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094444/.

Moerdijk:1995:CTb

[Moe95e] Izak Moerdijk. Comparison theorems. Lecture Notes in Mathematics, 1616:77–88, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094446/.

Moerdijk:1995:GR

[Moe95f] Izak Moerdijk. Geometric realization. Lecture Notes in Mathematics, 1616:57–76, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094445/.

Moerdijk:1995:I

[Moe95g] Izak Moerdijk. Introduction. Lecture Notes in Mathematics, 1616: 1–4, 1995. CODEN LNMAA2. ISBN 3-540-60319-0 (print), 3-540-44912-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094442/.

Mogilevskii:1992:ESB

[Mog92] Il’ia Mogilevskii. Estimates in C^{2l,l} for solution of a boundary value problem for the nonstationary Stokes system with a surface tension in boundary condition. Lecture Notes in Mathematics, 1530:284–290, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090348/.

Moller:1990:ESH

[Møl90] Jesper Michael Møller. Equivariant self-homotopy equivalences of 2-stage G-spaces. Lecture Notes in Mathematics, 1425:133– 146, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083837/.

Molchanov:1993:RVM

[Mol93a] I. S. Molchanov. On regularly varying multivalued functions. Lecture Notes in Mathematics, 1546:121–129, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084487/.

Molchanov:1993:ASC

[Mol93b] Ilya S. Molchanov. Almost sure convergence of unions of random closed sets. Lecture Notes in Mathematics, 1561:67–84, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073532/.

Molchanov:1993:ALT

[Mol93c] Ilya S. Molchanov. Applications of limit theorems. Lecture Notes in Mathematics, 1561:123–145, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073535/.

Molchanov:1993:DRC

[Mol93d] Ilya S. Molchanov. Distributions of random closed sets. Lecture Notes in Mathematics, 1561:1–14, 1993. CODEN LN-MAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073528/.

Molchanov:1993:IDS

[Mol93e] Ilya S. Molchanov. Infinite divisibility and stability of random sets with respect to unions. Lecture Notes in Mathematics, 1561: 29–44, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073530/.

Molchanov:1993:LTN

[Mol93f] Ilya S. Molchanov. Limit theorems for normalized unions of random closed sets. Lecture Notes in Mathematics, 1561:45–65, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073531/.

Molchanov:1993:LTU

[Mol93g] Ilya S. Molchanov. Limit Theorems for Unions of Random Closed Sets, volume 1561 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57393-3, 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 157 pp. LCCN QA3 .L28 no. 1561; QA3 .L28 no. 1561 QA273.5. URL http:/

/link.springer.com/book/10.1007/BFb0073527;   http://
www.springerlink.com/content/978-3-540-48111-9.

Molchanov:1993:MRV

[Mol93h] Ilya S. Molchanov. Multivalued regularly varying functions and their applications to limit theorems for unions of random sets. Lecture Notes in Mathematics, 1561:85–99, 1993. CODEN LN-MAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073533/.

Molchanov:1993:PMS

[Mol93i] Ilya S. Molchanov. Probability metrics in the space of random sets distributions. Lecture Notes in Mathematics, 1561:101–121, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073534/.

Molchanov:1993:SSR

[Mol93j] Ilya S. Molchanov. Survey on stability of random sets and limit theorems for Minkowski addition. Lecture Notes in Mathematics, 1561:15–27, 1993. CODEN LNMAA2. ISBN 3-540-57393-3 (print), 3-540-48111-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073529/.

Molchanov:1994:LRM

[Mol94] S. Molchanov. Lectures on random media. Lecture Notes in Mathematics, 1581:242–411, 1994. CODEN LNMAA2. ISBN 3-540-58208-8 (print), 3-540-48568-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073874/.

Mond:1991:VCA

[Mon91a] David Mond. Vanishing cycles for analytic maps. Lecture Notes in Mathematics, 1462:221–234, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086385/.

Montaldi:1991:CTR

[Mon91b] James Montaldi. Caustics in time reversible Hamiltonian systems. Lecture Notes in Mathematics, 1463:266–277, 1991. CO-DEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085435/.

Monat:1994:RIB

[Mon94] P. Monat. Remarques sur les inégalités de Burkholder–Davis– Gundy. (French) []. Lecture Notes in Mathematics, 1583:92– 97, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073835/.

Moore:1996:GAS

[Moo96a] John D. Moore. Global analysis of the Seiberg–Witten equations. Lecture Notes in Mathematics, 1629:65–100, 1996. CODEN LNMAA2. ISBN 3-540-61455-9 (print), 3-540-68592-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092951/.

Moore:1996:LSW

[Moo96b] John D. Moore. Lectures on Seiberg–Witten Invariants, volume 1629 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61455-9 (print), 3-540-68592-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 105 pp. LCCN QA3 .L28 no. 1629. URL http:/

/link.springer.com/book/10.1007/BFb0092948;   http://
www.springerlink.com/content/978-3-540-68592-0.

Moore:1996:P

[Moo96c] John D. Moore. Preliminaries. Lecture Notes in Mathematics, 1629:1–38, 1996. CODEN LNMAA2. ISBN 3-540-61455-9 (print), 3-540-68592-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092949/.

Moore:1996:SGF

[Moo96d] John D. Moore. Spin geometry on four-manifolds. Lecture Notes in Mathematics, 1629:39–64, 1996. CODEN LNMAA2. ISBN 3-540-61455-9 (print), 3-540-68592-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092950/.

Morava:1990:CCR

[Mor90a] Jack Morava. On the complex Cobordism ring as a Fock representation. Lecture Notes in Mathematics, 1418:184–204, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083703/.

Moreno:1990:PSC

[Mor90b] Carlos Moreno. Produits star sur certains G/K Kähleriens. Équation de Yang–Baxter et produits star sur G. (French) []. Lecture Notes in Mathematics, 1416:210–234, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097473/.

Morita:1990:VCR

[Mor90c] Yasuo Morita. On vanishing of cohomologies of rigid analytic spaces. Lecture Notes in Mathematics, 1454:314–318, 1990. CO-DEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091146/.

Moroz:1991:SCD

[Mor91] Irene M. Moroz. Some complex differential equations arising in telecommunications. Lecture Notes in Mathematics, 1463:278– 293, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085436/.

Morimoto:1993:ABW

[Mor93a] Hiroko Morimoto. Asymptotic behavior of weak solutions of the convection equation. Lecture Notes in Mathematics, 1540:269– 276, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085485/.

Morozov:1993:CTQ

[Mor93b] E. V. Morozov. A comparison theorem for queueing system with non-identical channels. Lecture Notes in Mathematics, 1546:130– 133, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084488/.

Min-Oo:1991:TCT

[MORT91] Maung Min-Oo, Ernst A. Ruh, and Philippe Tondeur. Transversal curvature and tautness for Riemannian foliations. Lecture Notes in Mathematics, 1481:145–146, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083638.

Motreanu:1991:GEM

[Mot91] D. Motreanu. Generic existence of Morse functions on infinite dimensional Riemannian manifolds and applications. Lecture Notes in Mathematics, 1481:175–184, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083640/.

Mourtada:1990:CFP

[Mou90] A. Mourtada. Cyclicité finie des polycycles hyperboliques de champs de vecteurs du plan mise sous forme normale. (French) []. Lecture Notes in Mathematics, 1455:272–314, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085397/.

Mountford:1992:CFP

[Mou92a] T. S. Mountford. A critical function for the planar Brownian convex hull. Lecture Notes in Mathematics, 1526:107–112, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084314/.

Mountford:1992:QEU

[Mou92b] T. S. Mountford. Quasi-everywhere upper functions. Lecture Notes in Mathematics, 1526:95–106, 1992. CODEN LN-MAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084313/.

Mosca:1991:CTL

[MP91] Edoardo Mosca and Luciano Pandolfi, editors. H_{∞}-Control Theory: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Como, Italy, June 18–26, 1990, volume 1496 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1496. URL http:/

/link.springer.com/book/10.1007/BFb0084465;   http://
www.springerlink.com/content/978-3-540-46604-8.

Marden:1990:TFP

[MR90] Al Marden and Burt Rodin. On Thurston’s formulation and proof of Andreev’s theorem. Lecture Notes in Mathematics, 1435:103– 115, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087901/.

Meyer:1991:CAR

[MR91a] M. Meyer and S. Reisner. Characterization of affinely-rotation-invariant log-concave measures by section-centroid location. Lecture Notes in Mathematics, 1469:145–152, 1991. CODEN LN-MAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089221/.

Montiel:1991:SOA

[MR91b] Sebastián Montiel and Antonio Ros. Schrödinger operators associated to a holomorphic map. Lecture Notes in Mathematics, 1481:147–174, 1991. CODEN LNMAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083639/.

Marcus:1994:ERC

[MR94] Michael B. Marcus and Jay Rosen. Exact rates of convergence to the local times of symmetric Lévy processes. Lecture Notes in Mathematics, 1583:102–109, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073837/.

Mikulevicius:1998:NSI

[MR98] R. Mikulevicius and B. L. Rozovskii. Normalized stochastic integrals in topological vector spaces. Lecture Notes in Mathematics, 1686:137–165, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101756/.

Maass:1990:RBC

[MS90a] Wolfgang Maass and Theodore A. Slaman. On the relationship between the complexity, the degree, and the extension of a computable set. Lecture Notes in Mathematics, 1432:297–322, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086124/.

Martinez:1990:BCB

[MS90b] Regina Martínez and Carles Simó. Blow up of collapsing binaries in the planar three body problem. Lecture Notes in Mathematics, 1416:255–267, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097476/.

Muller:1990:UMS

[MS90c] Bruno J. Müller and Surjeet Singh. Uniform modules over serial rings II. Lecture Notes in Mathematics, 1448:25–32, 1990.

CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091248/.

Memin:1991:CUS

[MS91a] Jean Mémin and Leszek Słominski. Condition UT et stabilité en loi des solutions d’équations différentielles stochastiques. (French) []. Lecture Notes in Mathematics, 1485:162–177, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100855/.

Sun:1991:HNL

[mS91b] Li min Sun. Hermitian nilpotent Lie groups: Harmonic analysis as spectral theory of Laplacians. Lecture Notes in Mathematics, 1494:182–184, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087770/.

Monat:1994:FF

[MS94a] P. Monat and C. Stricker. Fermeture de G_{T} (Θ) et de L^{2}(F_{′}) + G_{T} (×). (French) []. Lecture Notes in Mathematics, 1583:189– 194, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073847/.

Muller-Stach:1994:RHP

[MS94b] Stefan Müller-Stach. A remark on height pairings. Lecture Notes in Mathematics, 1594:253–259, 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-49046-3_6/.

Millet:1994:SPS

[MSS94] Annie Millet and Marta Sanz-Solé. A simple proof of the support theorem for diffusion processes. Lecture Notes in Mathematics, 1583:36–48, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073832/.

Marino:1990:SVP

[MT90] A. Marino and M. Tosques. Some variational problems with lack of convexity and some partial differential inequalities. Lecture Notes in Mathematics, 1446:58–83, 1990. CODEN LN-MAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084931/.

McIntosh:1991:CSI

[MT91a] Alan McIntosh and Qian Tao. Convolution singular integral operators on Lipschitz curves. Lecture Notes in Mathematics, 1494: 142–162, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087766/.

Moishezon:1991:BGT

[MT91b] Boris Moishezon and Mina Teicher. Braid group technique m complex geometry, II: From arrangements of lines and conics to cuspidal curves. Lecture Notes in Mathematics, 1479:131–180, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086269/.

Morayne:1999:SPD

[MT99] Michał Morayne and Krzysztof Tabisz. A short proof of decomposition of strongly reduced martingales. Lecture Notes in Mathematics, 1709:339–341, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096523/.

Mukai:1990:PTC

[Muk90] Juno Mukai. A proof of the theorem characterizing the generalized J-homomorphism. Lecture Notes in Mathematics, 1418: 95–104, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083696/.

Mukhometov:1991:PIG

[Muk91] R. G. Mukhometov. A problem of integral geometry for a family of rays with multiple reflections. Lecture Notes in Mathematics, 1497:46–52, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084506/.

Muldowney:1990:IDG

[Mul90a] P. Muldowney. Infinite-dimensional generalised Riemann integrals. Lecture Notes in Mathematics, 1419:131–135, 1990. CO-DEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083104/.

Muldowney:1990:ARH

[Mul90b] Pat Muldowney. About Ralph Henstock. Lecture Notes in Mathematics, 1419:1–6, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083093/.

Muller:1990:SLP

[Mul90c] Marie-Paule Muller. Une sphère Lagrangienne plongée dans une structure symplectique complète sur R^{6}. (French) []. Lecture Notes in Mathematics, 1416:235–241, 1990. CODEN LN-MAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097474/.

Muller:1991:PHS

[Mül91] Paul F. X. Müller. Permutations of the Haar system. Lecture Notes in Mathematics, 1469:125–126, 1991. CODEN LN-MAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089218.

Muller:1999:VMM

[Mül99] Stefan Müller. Variational models for microstructure and phase transitions. Lecture Notes in Mathematics, 1713:85–210, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092670/.

Mumford:1999:RBV

[Mum99a] David Mumford. The Red Book of Varieties and Schemes, volume 1358 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second expanded edition, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 306 pp. LCCN QA3 .L28 no. 1358. URL

http://link.springer.com/book/10.1007/b62130;   http:
//www.springerlink.com/content/978-3-540-46021-3. In-

cludes Michigan Lectures (1974) on Curves and their Jacobians.

Mumford:1999:ACT

[Mum99b] David B. Mumford. Appendix: Curves and their Jacobians. Lecture Notes in Mathematics, 1358:225–291, 1999. CODEN LN-MAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-46021-3_
4/.

Mumford:1999:GLR

[Mum99c] David B. Mumford. Guide to the literature and references: Curves and their Jacobians. Lecture Notes in Mathematics, 1358:294– 300, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46021-3_6/.

Mumford:1999:V

[Mum99d] David B. Mumford. I. Varieties. Lecture Notes in Mathematics, 1358:1–63, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46021-3_1/.

Mumford:1999:IP

[Mum99e] David B. Mumford. II. Preschemes. Lecture Notes in Mathematics, 1358:65–136, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46021-3_2/.

Mumford:1999:ILP

[Mum99f] David B. Mumford. III. Local properties of schemes. Lecture Notes in Mathematics, 1358:137–223, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46021-3_3/.

Mumford:1999:RRB

[Mum99g] David B. Mumford. References: The red book of varieties and schemes. Lecture Notes in Mathematics, 1358:293, 1999. CODEN LNMAA2. ISBN 3-540-63293-X (print), 3-540-46021-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-46021-3_5.

Murasugi:1991:IGT

[Mur91] Kunio Murasugi. Invariants of graphs and their applications to knot theory. Lecture Notes in Mathematics, 1474:83–97, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084739/.

Murata:1993:URP

[Mur93] Minoru Murata. Uniform restricted parabolic Harnack inequality, separation principle, and ultracontractivity for parabolic equations. Lecture Notes in Mathematics, 1540:277–288, 1993. CO-DEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085486/.

Murre:1994:ACA

[Mur94] Jacob P. Murre. Algebraic cycles and algebraic aspects of cohomology and K-theory. Lecture Notes in Mathematics, 1594: 93–152, 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49046-3_2/.

Musin:1992:SPC

[Mus92] O. R. Musin. On some problems of computational geometry and topology. Lecture Notes in Mathematics, 1520:57–80, 1992. CO-DEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084715/.

Mayer-Wolf:1992:LDM

[MWNPA92] Eduardo Mayer-Wolf, David Nualart, and Víctor Pérez-Abreu. Large deviations for multiple Wiener–Itô integral processes. Lecture Notes in Mathematics, 1526:11–31, 1992. CODEN LN-MAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084307/.

Meyer:1991:FCD

[MY91] P.-A. Meyer and J. A. Yan. Les “fonctions caractéristiques” des distributions sur l’espace de wiener. (French) []. Lecture Notes in Mathematics, 1485:61–78, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100847/.

Ma:1999:FBS

[MY99] Jin Ma and Jiongmin Yong. Forward-backward stochastic differential equations and their applications, volume 1702 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LN-MAA2. ISBN 3-540-65960-9. ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 270 pp. LCCN QA3 .L28 no. 1702; QA3 .L28 no. 1702 QA274.23. URL http://www.loc.gov/

catdir/enhancements/fy0815/99023640-d.html;   http://
www.loc.gov/catdir/enhancements/fy0815/99023640-t.html.

Ma:2007:FBS

[MY07] Jin Ma and Jiongmin Yong. Forward-backward stochastic differential equations and their applications, volume 1702 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., corrected 3rd printing edition, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 270 pp. LCCN QA3 .L28 no. 1702; QA274.23 .M3 2007 aa07 2007-12-11. URL

http://www.loc.gov/catdir/toc/fy0802/2007923869.html.

Muller:1990:OLF

[MZ90] Bruno J. Müller and Ying-Lan Zhang. Ore localization in the first Weyl algebra. Lecture Notes in Mathematics, 1448:148–154, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091259/.

Nabutovsky:1990:INR

[Nab90] Alexander Nabutovsky. Isotopies and non-recursive functions in real algebraic geometry. Lecture Notes in Mathematics, 1420: 194–205, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083921/.

Nagata:1990:SBE

[Nag90a] Wayne Nagata. Symmetry-breaking effects of distant sidewalls in Rayleigh–Bénard convection. Lecture Notes in Mathematics, 1431:109–116, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086061/.

Nagel:1990:SPS

[Nag90b] Rainer Nagel. On the stability of positive semigroups generated by operator matrices. Lecture Notes in Mathematics, 1422:77– 83, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089406/.

Nagasawa:1993:PSI

[Nag93] Masao Nagasawa. Principle of superposition and interference of diffusion processes. Lecture Notes in Mathematics, 1557:1– 14, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087957/.

Narasimhan:1990:LPA

[Nar90] Raghavan Narasimhan. The Levi problem on algebraic manifolds. Lecture Notes in Mathematics, 1422:85–91, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089407/.

Narayanaswami:1993:SQP

[Nar93] P. P. Narayanaswami. The separable quotient problem for barrelled spaces. Lecture Notes in Mathematics, 1540:289–308, 1993.

CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085487/.

Narkiewicz:1995:FIS

[Nar95a] Władysław Narkiewicz. Fully invariant sets for polynomial mappings. Lecture Notes in Mathematics, 1600:67–109, 1995. CODEN LNMAA2. ISBN 3-540-59435-3 (print), 3-540-49266-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076896/.

Narkiewicz:1995:PM

[Nar95b] Władysław Narkiewicz. Polynomial Mappings, volume 1600 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-59435-3 (print), 3-540-49266-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 130 pp. LCCN QA3 .L28 no. 1600. URL http:/

/link.springer.com/book/10.1007/BFb0076894;   http://
www.springerlink.com/content/978-3-540-49266-5.

Narkiewicz:1995:RIV

[Nar95c] Władysław Narkiewicz. Rings of integral-valued polynomials. Lecture Notes in Mathematics, 1600:1–66, 1995. CODEN LN-MAA2. ISBN 3-540-59435-3 (print), 3-540-49266-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0076895/.

Narasimhan:1997:VBC

[Nar97] M. S. Narasimhan, editor. Vector Bundles on Curves — New Directions: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro (Cosenza), Italy, June 19–27, 1995, volume 1649 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62401-5 (print), 3-540-49701-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1649.

URL http://link.springer.com/book/10.1007/BFb0094423;
 http://www.springerlink.com/content/978-3-540-49701-
1.

Nazarov:1992:YSL

[Naz92] M. L. Nazarov. Yangians of the “strange” Lie superalgebras. Lecture Notes in Mathematics, 1510:90–97, 1992. CODEN LN-MAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101181/.

Nerode:1990:SLI

[Ner90] Anil Nerode. Some lectures on intuitionistic logic. Lecture Notes in Mathematics, 1429:12–59, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093923/.

Neumann:1990:IFG

[Neu90] Walter D. Neumann. On intersections of finitely generated subgroups of free groups. Lecture Notes in Mathematics, 1456:161– 170, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100737/.

Neuenschwander:1996:BMH

[Neu96a] Daniel Neuenschwander. Brownian motions on H. Lecture Notes in Mathematics, 1630:29–84, 1996. CODEN LNMAA2. ISBN 3-540-61453-2 (print), 3-540-68590-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094032/.

Neuenschwander:1996:I

[Neu96b] Daniel Neuenschwander. Introduction. Lecture Notes in Mathematics, 1630:1–6, 1996. CODEN LNMAA2. ISBN 3-540-61453-2 (print), 3-540-68590-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094030/.

Neuenschwander:1996:OLT

[Neu96c] Daniel Neuenschwander. Other limit theorems on H. Lecture Notes in Mathematics, 1630:85–123, 1996. CODEN LN-MAA2. ISBN 3-540-61453-2 (print), 3-540-68590-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094033/.

Neuenschwander:1996:PHG

[Neu96d] Daniel Neuenschwander. Probabilities on the Heisenberg Group: Limit Theorems and Brownian Motion, volume 1630 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61453-2 (print), 3-540-68590-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 139 pp. LCCN QA3 .L28 no. 1630. URL http:/

/link.springer.com/book/10.1007/BFb0094029;   http://
www.springerlink.com/content/978-3-540-68590-6.

Neuenschwander:1996:PTS

[Neu96e] Daniel Neuenschwander. Probability theory on simply connected nilpotent Lie groups. Lecture Notes in Mathematics, 1630:7– 27, 1996. CODEN LNMAA2. ISBN 3-540-61453-2 (print), 3-540-68590-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094031/.

Neuberger:1997:CTG

[Neu97a] John William Neuberger. Comparison of two gradients. Lecture Notes in Mathematics, 1670:5–9, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092833/.

Neuberger:1997:CSDa

[Neu97b] John William Neuberger. Continuous steepest descent in Hilbert space: Linear case. Lecture Notes in Mathematics, 1670:11– 13, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092834/.

Neuberger:1997:CSDb

[Neu97c] John William Neuberger. Continuous steepest descent in Hilbert space: Nonlinear case. Lecture Notes in Mathematics, 1670:15– 31, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092835/.

Neuberger:1997:FDS

[Neu97d] John William Neuberger. Finite difference setting: the inner product case. Lecture Notes in Mathematics, 1670:59–68, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092839/.

Neuberger:1997:FPN

[Neu97e] John William Neuberger. Flow problems and non-inner product Sobolev spaces. Lecture Notes in Mathematics, 1670:107– 114, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092844/.

Neuberger:1997:FGB

[Neu97f] John William Neuberger. Foliations as a guide to boundary conditions. Lecture Notes in Mathematics, 1670:115–123, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092845/.

Neuberger:1997:IBC

[Neu97g] John William Neuberger. Introducing boundary conditions. Lecture Notes in Mathematics, 1670:43–52, 1997. CODEN LN-MAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092837/.

Neuberger:1997:MS

[Neu97h] John William Neuberger. Minimal surfaces. Lecture Notes in Mathematics, 1670:93–106, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092843/.

Neuberger:1997:NMC

[Neu97i] John William Neuberger. Newton’s method in the context of Sobolev gradients. Lecture Notes in Mathematics, 1670:53–58, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092838/.

Neuberger:1997:OPA

[Neu97j] John William Neuberger. Orthogonal projections, adjoints and Laplacians. Lecture Notes in Mathematics, 1670:33–42, 1997.

CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092836/.

Neuberger:1997:RAI

[Neu97k] John William Neuberger. A related analytic iteration method. Lecture Notes in Mathematics, 1670:135–138, 1997. CODEN LN-MAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092847/.

Neuberger:1997:SCC

[Neu97l] John William Neuberger. A sample computer code with notes. Lecture Notes in Mathematics, 1670:141–143, 1997. CODEN LN-MAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092849/.

Neuberger:1997:SG

[Neu97m] John William Neuberger. Several gradients. Lecture Notes in Mathematics, 1670:1–3, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092832/.

Neuberger:1997:SGD

[Neu97n] John William Neuberger. Sobolev Gradients and Differential Equations, volume 1670 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 150 pp. LCCN QA3 .L28 no. 1670.

URL http://link.springer.com/book/10.1007/BFb0092831;
 http://www.springerlink.com/content/978-3-540-69594-
3.

Neuberger:1997:SGW

[Neu97o] John William Neuberger. Sobolev gradients for weak solutions: Function space case. Lecture Notes in Mathematics, 1670:69– 73, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092840/.

Neuberger:1997:SGN

[Neu97p] John William Neuberger. Sobolev gradients in non-inner product spaces: Introduction. Lecture Notes in Mathematics, 1670: 75–78, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092841/.

Neuberger:1997:SRI

[Neu97q] John William Neuberger. Some related iterative methods for differential equations. Lecture Notes in Mathematics, 1670:125– 133, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092846/.

Neuberger:1997:SDC

[Neu97r] John William Neuberger. Steepest descent for conservation equations. Lecture Notes in Mathematics, 1670:139–140, 1997. CO-DEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092848.

Neuberger:1997:SEG

[Neu97s] John William Neuberger. The superconductivity equations of Ginzburg–Landau. Lecture Notes in Mathematics, 1670:79–91, 1997. CODEN LNMAA2. ISBN 3-540-63537-8 (print), 3-540-69594-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092842/.

Nevai:1994:OP

[Nev94] Paul Nevai. Orthogonal polynomials. Lecture Notes in Mathematics, 1574:177–206, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101063/.

Newman:1990:GPP

[New90] M. F. Newman. Groups of prime-power order. Lecture Notes in Mathematics, 1456:49–62, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100730/.

Nagasaka:1990:ANT

[NF90] Kenji Nagasaka and Etienne Fouvry, editors. Analytic Number Theory: Proceedings of the Japanese–French Symposium held in Tokyo, Japan, October 10–13, 1988, volume 1434 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1434.

URL http://link.springer.com/book/10.1007/BFb0097119;
 http://www.springerlink.com/content/978-3-540-47147-
9.

Niederreiter:1990:PNG

[Nie90] Prof. Dr. H. Niederreiter. Pseudorandom numbers generated from shift register sequences. Lecture Notes in Mathematics, 1452:165– 177, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096988/.

Nielsen:1991:BAL

[Nie91a] Torben T. Nielsen. The Bose algebra Γ_{0}H, ⟨, ⟩. Lecture Notes in Mathematics, 1472:4–22, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098305/.

Nielsen:1991:BAO

[Nie91b] Torben T. Nielsen. Bose algebras of operators. Lecture Notes in Mathematics, 1472:79–88, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098312/.

Nielsen:1991:BAC

[Nie91c] Torben T. Nielsen. Bose Algebras: The Complex and Real Wave Representations, volume 1472 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 132 pp. LCCN QA3 .L28 no. 1472.

URL http://link.springer.com/book/10.1007/BFb0098303;
 http://www.springerlink.com/content/978-3-540-47367-
1.

Nielsen:1991:CV

[Nie91d] Torben T. Nielsen. The coherent vectors in ΓH. Lecture Notes in Mathematics, 1472:33–44, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098307/.

Nielsen:1991:CWR

[Nie91e] Torben T. Nielsen. The complex wave representation. Lecture Notes in Mathematics, 1472:66–71, 1991. CODEN LN-MAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098310/.

Nielsen:1991:I

[Nie91f] Torben T. Nielsen. Introduction. Lecture Notes in Mathematics, 1472:1–3, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098304/.

Nielsen:1991:LO

[Nie91g] Torben T. Nielsen. Lifting operators to ΓH. Lecture Notes in Mathematics, 1472:23–32, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098306/.

Nielsen:1991:RWR

[Nie91h] Torben T. Nielsen. The real wave representation. Lecture Notes in Mathematics, 1472:72–78, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098311/.

Nielsen:1991:SSO

[Nie91i] Torben T. Nielsen. Some special operators. Lecture Notes in Mathematics, 1472:53–65, 1991. CODEN LNMAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098309/.

Nielsen:1991:WR

[Nie91j] Torben T. Nielsen. Wave representations of Γ(H+H*). Lecture Notes in Mathematics, 1472:89–93, 1991. CODEN LN-MAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098313/.

Nielsen:1991:WOW

[Nie91k] Torben T. Nielsen. The Wick ordering and the Weyl relations. Lecture Notes in Mathematics, 1472:45–52, 1991. CODEN LN-MAA2. ISBN 3-540-54041-5 (print), 3-540-47367-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098308/.

Nikolski:1994:IBM

[Nik94a] N. K. Nikolski. Interpolation, bases, multipliers. Lecture Notes in Mathematics, 1574:259–294, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101065/.

Nikolski:1994:SAS

[Nik94b] N. K. Nikolski. Spectral analysis and synthesis. Lecture Notes in Mathematics, 1574:1–72, 1994. CODEN LNMAA2. ISBN 3-540-57871-4 (print), 3-540-48368-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101061/.

Nillsen:1994:ADS

[Nil94a] Rodney Nillsen. Applications to differential and singular integral operators. Lecture Notes in Mathematics, 1586:118–151, 1994. CODEN LNMAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073515/.

Nillsen:1994:DSI

[Nil94b] Rodney Nillsen. Difference Spaces and Invariant Linear Forms, volume 1586 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 186 pp. LCCN QA3 .L28 no. 1586. URL http:/

/link.springer.com/book/10.1007/BFb0073511;   http://
www.springerlink.com/content/978-3-540-48652-7.

Nillsen:1994:GPR

[Nil94c] Rodney Nillsen. General and preparatory results. Lecture Notes in Mathematics, 1586:9–43, 1994. CODEN LNMAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073513/.

Nillsen:1994:I

[Nil94d] Rodney Nillsen. Introduction. Lecture Notes in Mathematics, 1586:1–8, 1994. CODEN LNMAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073512/.

Nillsen:1994:MDS

[Nil94e] Rodney Nillsen. Multiplication and difference spaces on R^{n}. Lecture Notes in Mathematics, 1586:44–117, 1994. CODEN LN-MAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073514/.

Nillsen:1994:RSG

[Nil94f] Rodney Nillsen. Results for L^{p} spaces on general groups. Lecture Notes in Mathematics, 1586:152–174, 1994. CODEN LN-MAA2. ISBN 3-540-58323-8 (print), 3-540-48652-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073516/.

Nishitani:1991:EHC

[Nis91] Tatsuo Nishitani. The effectively hyperbolic Cauchy problem. Lecture Notes in Mathematics, 1505:71–167, 1991. CODEN LN-MAA2. ISBN 3-540-55018-6 (print), 3-540-46655-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090884/.

Nishioka:1996:AIM

[Nis96a] Kumiko Nishioka. Algebraic independence of Mahler functions and their values. Lecture Notes in Mathematics, 1631:78–117, 1996. CODEN LNMAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093675/.

Nishioka:1996:AET

[Nis96b] Kumiko Nishioka. Applications of elimination theory. Lecture Notes in Mathematics, 1631:118–149, 1996. CODEN LN-MAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093676/.

Nishioka:1996:MFT

[Nis96c] Kumiko Nishioka. Mahler Functions and Transcendence, volume 1631 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 185 pp. LCCN QA3 .L28 no. 1631. URL http:/

/link.springer.com/book/10.1007/BFb0093672;   http://
www.springerlink.com/content/978-3-540-68598-2.

Nishioka:1996:RSM

[Nis96d] Kumiko Nishioka. Regular sequences and Mahler functions. Lecture Notes in Mathematics, 1631:150–175, 1996. CODEN LN-MAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093677/.

Nishioka:1996:TTMa

[Nis96e] Kumiko Nishioka. Transcendence theory of Mahler functions of one variable. Lecture Notes in Mathematics, 1631:1–32, 1996. CODEN LNMAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093673/.

Nishioka:1996:TTMb

[Nis96f] Kumiko Nishioka. Transcendence theory of Mahler functions of several variables. Lecture Notes in Mathematics, 1631:33– 77, 1996. CODEN LNMAA2. ISBN 3-540-61472-9 (print), 3-540-68598-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093674/.

Noumi:1992:AWP

[NM92] Masatoshi Noumi and Katsuhisa Mimachi. Askey–Wilson polynomials as spherical functions on SU_{q}(2). Lecture Notes in Mathematics, 1510:98–103, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101182/.

Noguchi:1991:PCG

[NO91] Junjiro Noguchi and Takeo Ohsawa, editors. Prospects in Complex Geometry: Proceedings of the 25th Taniguchi International Symposium held in Katata, and the Conference held in Kyoto, July 31–August 9, 1989, volume 1468 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1468. URL http:/

/link.springer.com/book/10.1007/BFb0086186;   http://
www.springerlink.com/content/978-3-540-47370-1.

Nogin:1991:NEV

[Nog91] D. Yu. Nogin. Notes on exceptional vector bundles and helices. Lecture Notes in Mathematics, 1479:181–195, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086270/.

Norris:1992:CDF

[Nor92] J. R. Norris. A complete differential formalism for stochastic calculus in manifolds. Lecture Notes in Mathematics, 1526:189– 209, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084322/.

Notay:1990:SPS

[Not90] Y. Notay. Solving positive (semi)definite linear systems by preconditioned iterative methods. Lecture Notes in Mathematics, 1457:105–125, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090904/.

Novikov:1992:RIM

[Nov92] Sergei Novikov. Rôle of integrable models in the development of mathematics. Lecture Notes in Mathematics, 1525:13–28, 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089202/.

Nowak:1990:DAP

[Now90] Werner Georg Nowak. Divisors in arithmetic progressions in imaginary quadratic number fields. Lecture Notes in Mathematics, 1452:178–192, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096989/.

Nerode:1990:PIS

[NR90] A. Nerode and J. B. Remmel. Polynomially isolated sets. Lecture Notes in Mathematics, 1432:323–362, 1990. CODEN LN-MAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086125/.

Nazaikinskii:1992:IMO

[NSS92] V. E. Nazaĭkinskiĭ, B. Yu. Sternin, and V. E. Shatalov. Introduction to Maslov’s operational method (non-commutative analysis and differential equations). Lecture Notes in Mathematics, 1520: 81–91, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084716/.

Nualart:1998:AWS

[Nua98] David Nualart. Analysis on Wiener space and anticipating stochastic calculus. Lecture Notes in Mathematics, 1690:123– 220, 1998. CODEN LNMAA2. ISBN 3-540-64620-5 (print), 3-540-69228-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092538/.

Nussbaum:1993:BM

[Nus93a] Roger D. Nussbaum. Back matter. Lecture Notes in Mathematics, 1537:??, 1993. CODEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47563-7/1.pdf.

Nussbaum:1993:FPI

[Nus93b] Roger D. Nussbaum. The fixed point index and fixed point theorems. Lecture Notes in Mathematics, 1537:143–205, 1993. CO-DEN LNMAA2. ISBN 3-540-56461-6 (print), 3-540-47563-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085077/.

Nualart:1990:SRI

[NÜZ90] D. Nualart, A. S. Üstünel, and M. Zakai. Some remarks on independence and conditioning on Wiener space. Lecture Notes in Mathematics, 1444:122–127, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083612/.

Nualart:1990:ACP

[NV90] David Nualart and Josep Vives. Anticipative calculus for the Poisson process based on the Fock space. Lecture Notes in Mathematics, 1426:155–165, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083763/.

Nakanishi:1991:IRT

[NV91] Toshihiro Nakanishi and John A. Velling. On inner radii of Teichmüller spaces. Lecture Notes in Mathematics, 1468:115– 126, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086190/.

Nikolskii:1994:FM

[NV94] N. K. Nikolskii and V. I. Vasyunin. Functional models. Lecture Notes in Mathematics, 1573:389–408, 1994. CODEN LN-MAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100210/.

Nikolaev:1999:ASF

[NZ99a] Igor Nikolaev and Evgeny Zhuzhoma. C^{∗}-algebras of surface flows. Lecture Notes in Mathematics, 1705:175–199, 1999. CO-DEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093607/.

Nikolaev:1999:AWP

[NZ99b] Igor Nikolaev and Evgeny Zhuzhoma. Anosov–Weil problem. Lecture Notes in Mathematics, 1705:209–237, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093609/.

Nikolaev:1999:DF

[NZ99c] Igor Nikolaev and Evgeny Zhuzhoma. Decomposition of flows. Lecture Notes in Mathematics, 1705:41–62, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093602/.

Nikolaev:1999:DE

[NZ99d] Igor Nikolaev and Evgeny Zhuzhoma. Definitions and examples. Lecture Notes in Mathematics, 1705:1–21, 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093600/.

Nikolaev:1999:ET

[NZ99e] Igor Nikolaev and Evgeny Zhuzhoma. Ergodic theory. Lecture Notes in Mathematics, 1705:95–114, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093605/.

Nikolaev:1999:FDM

[NZ99f] Igor Nikolaev and Evgeny Zhuzhoma. Flows on 2-dimensional Manifolds: An Overview, volume 1705 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 294 pp. LCCN QA3 .L28 no. 1705.

URL http://link.springer.com/book/10.1007/BFb0093599;
 http://www.springerlink.com/content/978-3-540-48759-
3.

Nikolaev:1999:ISF

[NZ99g] Igor Nikolaev and Evgeny Zhuzhoma. Invariants of surface flows. Lecture Notes in Mathematics, 1705:115–174, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093606/.

Nikolaev:1999:LT

[NZ99h] Igor Nikolaev and Evgeny Zhuzhoma. Local theory. Lecture Notes in Mathematics, 1705:63–71, 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093603/.

Nikolaev:1999:NCS

[NZ99i] Igor Nikolaev and Evgeny Zhuzhoma. Non-compact surfaces. Lecture Notes in Mathematics, 1705:239–255, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093610/.

Nikolaev:1999:PBT

[NZ99j] Igor Nikolaev and Evgeny Zhuzhoma. Poincaré–Bendixson’s theory. Lecture Notes in Mathematics, 1705:23–39, 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093601/.

Nikolaev:1999:SLT

[NZ99k] Igor Nikolaev and Evgeny Zhuzhoma. Semi-local theory. Lecture Notes in Mathematics, 1705:201–208, 1999. CODEN LN-MAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093608/.

Nikolaev:1999:SFV

[NZ99l] Igor Nikolaev and Evgeny Zhuzhoma. Space of flows and vector fields. Lecture Notes in Mathematics, 1705:73–93, 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093604/.

Nikolaev:1999:T

[NZ99m] Igor Nikolaev and Evgeny Zhuzhoma. Triptych. Lecture Notes in Mathematics, 1705:257–268, 1999. CODEN LNMAA2. ISBN 3-540-66080-1 (print), 3-540-48759-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093611/.

Obata:1994:A

[Oba94a] Nobuaki Obata. Addendum. Lecture Notes in Mathematics, 1577: 151–166, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073958/.

Obata:1994:OT

[Oba94b] Nobuaki Obata. Operator theory. Lecture Notes in Mathematics, 1577:71–108, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073956/.

Obata:1994:P

[Oba94c] Nobuaki Obata. Prerequisites. Lecture Notes in Mathematics, 1577:1–18, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073953/.

Obata:1994:THA

[Oba94d] Nobuaki Obata. Toward harmonic analysis. Lecture Notes in Mathematics, 1577:109–150, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073957/.

Obata:1994:WNC

[Oba94e] Nobuaki Obata. White Noise Calculus and Fock Space, volume 1577 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 183 pp. LCCN QA3 .L28 no. 1577. URL http:

//link.springer.com/book/10.1007/BFb0073952;   http:/
/www.springerlink.com/content/978-3-540-48411-0.

Obata:1994:WNF

[Oba94f] Nobuaki Obata. White noise functionals. Lecture Notes in Mathematics, 1577:33–70, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073955/.

Obata:1994:WNS

[Oba94g] Nobuaki Obata. White noise space. Lecture Notes in Mathematics, 1577:19–32, 1994. CODEN LNMAA2. ISBN 3-540-57985-0 (print), 3-540-48411-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073954/.

Oda:1990:MHS

[Oda90a] Takayuki Oda. Multiple Hecke series for class-1 Whittaker functions on GL(n) over p-adic fields. Lecture Notes in Mathematics, 1434:196–214, 1990. CODEN LNMAA2. ISBN 3-540-52787-7 (print), 3-540-47147-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097136/.

Oda:1990:RHP

[Oda90b] Takayuki Oda. The Riemann–Hodge period relation for Hilbert modular forms of weight 2. Lecture Notes in Mathematics, 1447: 261–286, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085733/.

Odifreddi:1990:LCS

[Odi90] Piergiorgio Odifreddi, editor. Logic and Computer Science: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held at Montecatini Terme, Italy, June 20–28, 1988, volume 1429 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1429. URL http:/

/link.springer.com/book/10.1007/BFb0093921;   http://
www.springerlink.com/content/978-3-540-47135-6.

Ogus:1990:AAC

[Ogu90] Arthur Ogus. A p-adic analogue of the Chowla–Selberg formula. Lecture Notes in Mathematics, 1454:319–341, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091147/.

Ohsawa:1991:SHR

[Ohs91] Takeo Ohsawa. A strong harmonic representation theorem on complex spaces with isolated singularities. Lecture Notes in Mathematics, 1468:170–176, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086193/.

Ogawa:1991:RTA

[OK91] Hidemitsu Ogawa and Itsuo Kumazawa. Radon transform and analog coding. Lecture Notes in Mathematics, 1497:229–241, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084521/.

Okamoto:1990:ADB

[Oka90] Hisashi Okamoto. Applications of degenerate bifurcation equations to the Taylor problem and the water wave problem. Lecture Notes in Mathematics, 1431:117–127, 1990. CODEN LN-MAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086062/.

Okonek:1991:IIA

[Oko91] Christian Okonek. Instanton invariants and algebraic surfaces. Lecture Notes in Mathematics, 1504:138–186, 1991. CODEN LNMAA2. ISBN 3-540-55017-8 (print), 3-540-46651-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094290/.

Odyniec:1990:I

[OL90a] Włodzimierz Odyniec and Grzegorz Lewicki. Introduction. Lecture Notes in Mathematics, 1449:1–17, 1990. CODEN LN-MAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094528/.

Odyniec:1990:IBS

[OL90b] Włodzimierz Odyniec and Grzegorz Lewicki. Isometries of Banach spaces and the problem of characterization of Hilbert spaces. Lecture Notes in Mathematics, 1449:131–152, 1990. CODEN LN-MAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094532/.

Odyniec:1990:KTC

[OL90c] Włodzimierz Odyniec and Grzegorz Lewicki. Kolmogorov’s type criteria for minimal projections. Lecture Notes in Mathematics, 1449:94–130, 1990. CODEN LNMAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094531/.

Odyniec:1990:MPB

[OL90d] Włodzimierz Odyniec and Grzegorz Lewicki. Minimal Projections in Banach Spaces: Problems of Existence and Uniqueness and their Application, volume 1449 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 168 pp. LCCN QA3 .L28 no. 1449.

URL http://link.springer.com/book/10.1007/BFb0094527;
 http://www.springerlink.com/content/978-3-540-46753-
3.

Odyniec:1990:MPC

[OL90e] Włodzimierz Odyniec and Grzegorz Lewicki. Minimal projections onto codimension one subspaces and a related mathematical programming problem. Lecture Notes in Mathematics, 1449:

52–93, 1990. CODEN LNMAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094530/.

Odyniec:1990:PUM

[OL90f] Włodzimierz Odyniec and Grzegorz Lewicki. Problem of uniqueness of minimal projections in Banach spaces. Lecture Notes in Mathematics, 1449:18–51, 1990. CODEN LNMAA2. ISBN 3-540-53197-1 (print), 3-540-46753-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094529/.

Olshanskii:1992:TYI

[Ol’92] G. I. Ol’shanskiĭ. Twisted Yangians and infinite-dimensional classical Lie algebras. Lecture Notes in Mathematics, 1510:104– 119, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101183/.

Olech:1990:LTE

[Ole90] Czesraw Olech. The Lyapunov Theorem: Its extensions and applications. Lecture Notes in Mathematics, 1446:84–103, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084932/.

Oliva:1990:PMP

[Oli90] Cristina Oliva. On the pluriadjoint maps of polarized normal Gorenstein surfaces. Lecture Notes in Mathematics, 1417:180– 183, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083341/.

Oller:1993:IBM

[Oll93] Josep M. Oller. On an intrinsic bias measure. Lecture Notes in Mathematics, 1546:134–158, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084489/.

Oort:1993:BM

[Oor93a] Frans Oort. Back matter. Lecture Notes in Mathematics, 1566: 123–127, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-48208-6/1.pdf.

Oort:1993:GCL

[Oor93b] Frans Oort. “The” general case of S. Lang’s conjecture (after Faltings). Lecture Notes in Mathematics, 1566:117–122, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48208-6_13/.

Opozda:1991:SER

[Opo91] Barbara Opozda. Some extensions of Radon’s theorem. Lecture Notes in Mathematics, 1481:185–191, 1991. CODEN LN-MAA2. ISBN 3-540-54728-2 (print), 3-540-46445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083641/.

Odwell:1991:FAP

[OR91] Edward E. Odwell, Jr. and Haskell P. Rosenthal, editors. Functional Analysis: Proceedings of the Seminar at the University of Texas at Austin 1987–89, volume 1470 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1470. URL http:/

/link.springer.com/book/10.1007/BFb0090208;   http://
www.springerlink.com/content/978-3-540-47493-7.

Orellana:1990:SAC

[Ore90] O. F. Orellana. On some analytic and computational aspects of two dimensional vortex sheet evolution. Lecture Notes in Mathematics, 1435:143–154, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087904/.

Ollagnier:1990:FIL

[OS90] Jean Moulin Ollagnier and Jean-Marie Strelcyn. On first integrals of linear systems, Frobenius integrability theorem and linear representations of Lie algebras. Lecture Notes in Mathematics, 1455:243–271, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085396/.

Okamoto:1993:CAA

[OSK93] H. Okamoto, M. Shōji, and M. Katsurada. A computer-assisted analysis of the two dimensional Navier–Stokes equations. Lecture Notes in Mathematics, 1540:309–318, 1993. CODEN LN-MAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085488/.

Ogawa:1990:BSN

[OT90] Takayoshi Ogawa and Yoshio Tsutsumi. Blow-up of solutions for the nonlinear Schrödinger equation with quartic potential and periodic boundary condition. Lecture Notes in Mathematics, 1450: 236–251, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084910/.

Otani:1993:PES

[Ôta93] Mitsuharu Ôtani. A priori estimates for some nonlinear parabolic equations via Lyapunov functions. Lecture Notes in Mathematics, 1540:319–327, 1993. CODEN LNMAA2. ISBN 3-540-56471-3 (print), 3-540-47565-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085489/.

Ohnita:1991:CAP

[OU91] Yoshihiro Ohnita and Seiichi Udagawa. Complex-analyticity of pluriharmonic maps and their constructions. Lecture Notes in Mathematics, 1468:371–407, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086201/.

Oudin:1991:DSD

[Oud91] Louis R. Oudin. Determination of the specific density of an aerosol through tomography. Lecture Notes in Mathematics, 1497:242– 260, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084522/.

Ouyang:1991:MRS

[Ouy91] Guangzhong Ouyang. Multipliers from L_{1}(G) to a reflexive Segal algebra. Lecture Notes in Mathematics, 1494:163–168, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087767/.

Overbeck:1995:PRP

[Ove95] L. Overbeck. On the predictable representation property for superprocesses. Lecture Notes in Mathematics, 1613:108–116, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094203/.

Ozawa:1990:SES

[Oza90] Tohru Ozawa. Smoothing effect for the Schrödinger evolution equations with electric fields. Lecture Notes in Mathematics, 1450: 226–235, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084909/.

Palleschi:1990:ALB

[Pal90] Marino Palleschi. On the adjoint line bundle to an ample and spanned one. Lecture Notes in Mathematics, 1417:184–190, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083342/.

Palamodov:1991:IFT

[Pal91] Victor P. Palamodov. Inversion formulas for the threedimensional ray transform. Lecture Notes in Mathematics, 1497: 53–62, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084507/.

Pan:1991:WNI

[Pan91a] Wenjie Pan. Weighted norm inequalities for certain maximal operators with approach regions. Lecture Notes in Mathematics, 1494:169–175, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087768/.

Panchishkin:1991:A

[Pan91b] Alexey A. Panchishkin. Acknowledgement. Lecture Notes in Mathematics, 1471:8, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21541-8_2.

Panchishkin:1991:BM

[Pan91c] Alexey A. Panchishkin. Back matter. Lecture Notes in Mathematics, 1471:146–161, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21541-8/1.pdf.

Panchishkin:1991:FM

[Pan91d] Alexey A. Panchishkin. Front matter. Lecture Notes in Mathematics, 1471:N2–vii, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21541-8/1.pdf.

Panchishkin:1991:I

[Pan91e] Alexey A. Panchishkin. Introduction. Lecture Notes in Mathematics, 1471:1–8, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21541-8_1/.

Panchishkin:1991:NAA

[Pan91f] Alexey A. Panchishkin. Non-Archimedean analytic functions, measures and distributions. Lecture Notes in Mathematics, 1471:

9–34, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21541-8_3/.

Panchishkin:1991:NAC

[Pan91g] Alexey A. Panchishkin. Non-Archimedean convolutions of Hilbert modular forms. Lecture Notes in Mathematics, 1471:117–145, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21541-8_6/.

Panchishkin:1991:NAF

[Pan91h] Alexey A. Panchishkin. Non-Archimedean L-Functions of Siegel and Hilbert Modular Forms, volume 1471 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 157 pp. LCCN QA3 .L28 no.

1471. URL http://link.springer.com/book/10.1007/978-
3-662-21541-8;   http://www.springerlink.com/content/
978-3-662-21541-8.

Panchishkin:1991:NAS

[Pan91i] Alexey A. Panchishkin. Non-Archimedean standard zeta functions of Siegel modular forms. Lecture Notes in Mathematics, 1471:81–116, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21541-8_5/.

Panchishkin:1991:SMF

[Pan91j] Alexey A. Panchishkin. Siegel modular forms and the holomorphic projection operator. Lecture Notes in Mathematics, 1471: 35–80, 1991. CODEN LNMAA2. ISBN 3-540-54137-3 (print), 3-662-21541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21541-8_4/.

Pardoux:1990:AAS

[Par90] Etienne Pardoux. Applications of anticipating stochastic calculus to stochastic differential equations. Lecture Notes in Mathematics, 1444:63–105, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083610/.

Pardoux:1991:FNL

[Par91a] Etienne Pardoux. Filtrage non linéaire et Équations aux derivées partielles stochastiques associées. (French) []. Lecture Notes in Mathematics, 1464:68–163, 1991. CODEN LNMAA2. ISBN 3-540-53841-0 (print), 3-540-46319-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085168/.

Parthasarathy:1991:ARU

[Par91b] K. R. Parthasarathy. An additional remark on unitary evolutions in Fock space. Lecture Notes in Mathematics, 1485:37–38, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100844.

Parthasarathy:1991:RCM

[Par91c] K. R. Parthasarathy. Realisation of a class of Markov processes through unitary evolutions in Fock space. Lecture Notes in Mathematics, 1485:31–36, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100843/.

Parimala:1992:AGM

[Par92] R. Parimala. Algebraic geometric methods in real algebraic geometry. Lecture Notes in Mathematics, 1524:37–51, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084607/.

Parfenov:1993:EPB

[Par93] O. G. Parfënov. Extremal problems for Blaschke products and widths. Lecture Notes in Mathematics, 1550:164–168, 1993. CO-DEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0117483/.

Pataraia:1990:QCP

[Pat90] P. Pataraia. On Quillen’s + construction of perfect groups. Lecture Notes in Mathematics, 1437:222–267, 1990. CODEN LN-MAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086723/.

Pazhitnov:1991:MTC

[Paz91] A. V. Pazhitnov. Morse theory of closed 1-forms. Lecture Notes in Mathematics, 1474:98–110, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084740/.

Pontier:1992:RHS

[PE92] Monique Pontier and Anne Estrade. Relévement horizontal d’une semi-martingale càdlàg. (French) []. Lecture Notes in Mathematics, 1526:127–145, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084316/.

Pearson:1991:NOC

[Pea91] J. B. Pearson. Notes on l^{1}-optimal control. Lecture Notes in Mathematics, 1496:222–249, 1991. CODEN LNMAA2. ISBN 3-540-54949-8 (print), 3-540-46604-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084470/.

Pedersen:1992:FFR

[Ped92] Jens Peter Pedersen. A function field related to the Ree group. Lecture Notes in Mathematics, 1518:122–131, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087997/.

Peetre:1994:HTO

[Pee94] Jaak Peetre. Hankel and Toeplitz operators. Lecture Notes in Mathematics, 1573:293–358, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100208/.

Pellikaan:1991:CCE

[Pel91] Ruud Pellikaan. On complete conditions in enumerative geometry. Lecture Notes in Mathematics, 1462:235–257, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086386/.

Pellikaan:1992:GCA

[Pel92] Ruud Pellikaan. On the gonality of curves, abundant codes and decoding. Lecture Notes in Mathematics, 1518:132–144, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087998/.

Perko:1990:BLC

[Per90] L. M. Perko. Bifurcation of limit cycles. Lecture Notes in Mathematics, 1455:315–333, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085398/.

Peres:1991:ADL

[Per91] Yuval Peres. Analytic dependence of Lyapunov exponents on transition probabilities. Lecture Notes in Mathematics, 1486:64– 80, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086658/.

Persson:1992:GTF

[Per92] Ulf Persson. Genus two fibrations revisited. Lecture Notes in Mathematics, 1507:133–144, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094515/.

Peres:1999:PTI

[Per99] Yuval Peres. Probability on trees: an introductory climb. Lecture Notes in Mathematics, 1717:193–280, 1999. CODEN LNMAA2. ISBN 3-540-66593-5 (print), 3-540-48115-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48115-7_3/.

Peters:1991:CTP

[Pet91] Martin Peters. Classification of two-parameter bifurcations. Lecture Notes in Mathematics, 1463:294–300, 1991. CODEN LN-MAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085437/.

Peters:1992:RNR

[Pet92] C. A. M. Peters. On the rank of non-rigid period maps in the weight one and two case. Lecture Notes in Mathematics, 1507: 157–165, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094517/.

Peternell:1996:MSP

[Pet96] Thomas Peternell. Manifolds of semi-positive curvature. Lecture Notes in Mathematics, 1646:98–142, 1996. CODEN LN-MAA2. ISBN 3-540-62038-9 (print), 3-540-49632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094303/.

Pfeffer:1990:DTV

[Pfe90] Washek F. Pfeffer. Divergence theorem for vector fields with singularities. Lecture Notes in Mathematics, 1419:150–166, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083106/.

Pierson-Gorez:1990:DIC

[PG90] Cecile Pierson-Gorez. Double integrals and convergence of double series. Lecture Notes in Mathematics, 1419:38–53, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083098/.

Phan:1991:TSR

[Phà91] Nguyen Huynh Phàn. On the topology of the space of reachable symmetric linear systems. Lecture Notes in Mathematics, 1474:235–253, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084750/.

Phelps:1993:ASR

[Phe93a] Robert R. Phelps. Asplund spaces, the RNP and perturbed optimization. Lecture Notes in Mathematics, 1364:79–94, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46077-0_5/.

Phelps:1993:BM

[Phe93b] Robert R. Phelps. Back matter. Lecture Notes in Mathematics, 1364:110–120, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-46077-0/1.pdf.

Phelps:1993:CFM

[Phe93c] Robert R. Phelps. Convex Functions, Monotone Operators and Differentiability, volume 1364 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 117 pp. LCCN QA3 .L28 no. 1364.

URL http://link.springer.com/book/10.1007/978-3-540-
46077-0;   http://www.springerlink.com/content/978-3-
540-46077-0.

Phelps:1993:CFR

[Phe93d] Robert R. Phelps. Convex functions on real Banach spaces. Lecture Notes in Mathematics, 1364:1–16, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46077-0_1/.

Phelps:1993:FM

[Phe93e] Robert R. Phelps. Front matter. Lecture Notes in Mathematics, 1364:I–XI, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-46077-0/1.pdf.

Phelps:1993:GDS

[Phe93f] Robert R. Phelps. Gâteaux differentiability spaces. Lecture Notes in Mathematics, 1364:95–101, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46077-0_6/.

Phelps:1993:GMO

[Phe93g] Robert R. Phelps. A generalization of monotone operators: Usco maps. Lecture Notes in Mathematics, 1364:102–109, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46077-0_7/.

Phelps:1993:LSC

[Phe93h] Robert R. Phelps. Lower semicontinuous convex functions. Lecture Notes in Mathematics, 1364:38–57, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46077-0_3/.

Phelps:1993:MOS

[Phe93i] Robert R. Phelps. Monotone operators, subdifferentials and Asplund spaces. Lecture Notes in Mathematics, 1364:17–37, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46077-0_2/.

Phelps:1993:SVP

[Phe93j] Robert R. Phelps. Smooth variational principles, Asplund spaces, weak Asplund spaces. Lecture Notes in Mathematics, 1364:58– 78, 1993. CODEN LNMAA2. ISBN 3-540-56715-1 (print), 3-540-46077-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46077-0_4/.

Philippon:1990:PIF

[Phi90] Patrice Philippon. Polynômes d’interpolation sur Z et Z[i]. (French) []. Lecture Notes in Mathematics, 1415:180–195, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084887/.

Pianigiani:1990:DIB

[Pia90] Giulio Pianigiani. Differential inclusions the Baire category method. Lecture Notes in Mathematics, 1446:104–136, 1990. CO-DEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084933/.

Piccinini:1990:GSE

[Pic90] Renzo A. Piccinini, editor. Groups of Self-Equivalences and Related Topics: Proceedings of a Conference held in Montreal, Canada, Aug. 8–12, 1988, volume 1425 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1425. URL http:/

/link.springer.com/book/10.1007/BFb0083825;   http://
www.springerlink.com/content/978-3-540-47091-5.

Picard:1991:CSA

[Pic91] Jean Picard. Calcul stochastique avec sauts sur une variété. (French) []. Lecture Notes in Mathematics, 1485:196–219, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100857/.

Pileckas:1990:NFB

[Pil90] Konstantin Pileckas. Noncompact free boundary problems for the Navier–Stokes equations. Lecture Notes in Mathematics, 1431: 60–72, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086055/.

Pilyugin:1994:DPR

[Pil94a] Sergei Yu. Pilyugin. Definitions and preliminary results. Lecture Notes in Mathematics, 1571:1–21, 1994. CODEN LN-MAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073520/.

Pilyugin:1994:GPD

[Pil94b] Sergei Yu. Pilyugin. Generic properties of dynamical systems. Lecture Notes in Mathematics, 1571:23–52, 1994. CODEN LN-MAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073521/.

Pilyugin:1994:LSD

[Pil94c] Sergei Yu. Pilyugin. Limit sets of domains. Lecture Notes in Mathematics, 1571:125–142, 1994. CODEN LNMAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073524/.

Pilyugin:1994:PA

[Pil94d] Sergei Yu. Pilyugin. Perturbations of attractors. Lecture Notes in Mathematics, 1571:87–123, 1994. CODEN LNMAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073523/.

Pilyugin:1994:SDS

[Pil94e] Sergei Yu. Pilyugin. The Space of Dynamical Systems with the C^{0}- Topology, volume 1571 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 188 pp. LCCN QA3 .L28 no. 1571. URL http:

//link.springer.com/book/10.1007/BFb0073519;   http:/
/www.springerlink.com/content/978-3-540-48314-4.

Pilyugin:1994:TS

[Pil94f] Sergei Yu. Pilyugin. Topological stability. Lecture Notes in Mathematics, 1571:53–85, 1994. CODEN LNMAA2. ISBN 3-540-57702-5 (print), 3-540-48314-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073522/.

Pillay:1998:MTC

[Pil98a] Anand Pillay. The model-theoretic content of Lang’s conjecture. Lecture Notes in Mathematics, 1696:101–106, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-68521-0_

Pillay:1998:MTA

[Pil98b] Anand Pillay. Model theory of algebraically closed fields. Lecture Notes in Mathematics, 1696:61–84, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-68521-0_4/.

Pilyugin:1999:NAS

[Pil99a] Sergei Yu. Pilyugin. Numerical applications of shadowing. Lecture Notes in Mathematics, 1706:219–257, 1999. CODEN LN-MAA2. ISBN 3-540-66299-5 (print), 3-540-48429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093188/.

Pilyugin:1999:SDS

[Pil99b] Sergei Yu. Pilyugin. Shadowing in Dynamical Systems, volume 1706 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66299-5 (print), 3-540-48429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 271 pp. LCCN QA3 .L28 no. 1706. URL http:/

/link.springer.com/book/10.1007/BFb0093184;   http://
www.springerlink.com/content/978-3-540-48429-5.

Pilyugin:1999:SNI

[Pil99c] Sergei Yu. Pilyugin. Shadowing near an invariant set. Lecture Notes in Mathematics, 1706:1–101, 1999. CODEN LN-MAA2. ISBN 3-540-66299-5 (print), 3-540-48429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093185/.

Pilyugin:1999:SSS

[Pil99d] Sergei Yu. Pilyugin. Systems with special structure. Lecture Notes in Mathematics, 1706:173–217, 1999. CODEN LNMAA2. ISBN 3-540-66299-5 (print), 3-540-48429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093187/.

Pilyugin:1999:TSS

[Pil99e] Sergei Yu. Pilyugin. Topologically stable, structurally stable, and generic systems. Lecture Notes in Mathematics, 1706:103– 172, 1999. CODEN LNMAA2. ISBN 3-540-66299-5 (print), 3-540-48429-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093186/.

Pirashvili:1990:CSC

[Pir90] Teimuraz Pirashvili. Cohomology of small categories in homotopical algebra. Lecture Notes in Mathematics, 1437:268–302, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086724/.

Pirola:1994:II

[Pir94] Gian Pietro Pirola. The infinitesimal invariant of C^{+} − C^{−}. Lecture Notes in Mathematics, 1594:223–232, 1994. CODEN LN-MAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-49046-3_
4/.

Pisier:1996:BM

[Pis96a] Gilles Pisier. Back matter. Lecture Notes in Mathematics, 1618: 143–161, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-662-21537-1/1.pdf.

Pisier:1996:CBH

[Pis96b] Gilles Pisier. Completely bounded homomorphisms and derivations. Lecture Notes in Mathematics, 1618:70–91, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21537-1_
5/.

Pisier:1996:CBMa

[Pis96c] Gilles Pisier. Completely bounded maps. Lecture Notes in Mathematics, 1618:53–69, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21537-1_4/.

Pisier:1996:CBMb

[Pis96d] Gilles Pisier. Completely bounded maps in the Banach space setting. Lecture Notes in Mathematics, 1618:133–142, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21537-1_
9/.

Pisier:1996:FM

[Pis96e] Gilles Pisier. Front matter. Lecture Notes in Mathematics, 1618: N2–vii, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21537-1/1.pdf.

Pisier:1996:HSM

[Pis96f] Gilles Pisier. Hankelian Schur multipliers. Herz–Schur multipliers. Lecture Notes in Mathematics, 1618:107–115, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21537-1_
7/.

Pisier:1996:IDC

[Pis96g] Gilles Pisier. Introduction. Description of contents. Lecture Notes in Mathematics, 1618:1–12, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21537-1_1/.

Pisier:1996:NUU

[Pis96h] Gilles Pisier. Non-unitarizable uniformly bounded group representations. Lecture Notes in Mathematics, 1618:30–52, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21537-1_3/.

Pisier:1996:SMG

[Pis96i] Gilles Pisier. Schur multipliers and Grothendieck’s inequality. Lecture Notes in Mathematics, 1618:92–106, 1996. CODEN LN-MAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21537-1_

Pisier:1996:SPCa

[Pis96j] Gilles Pisier. The similarity problem for cyclic homomorphisms on a C∗-algebra. Lecture Notes in Mathematics, 1618:116– 132, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21537-1_8/.

Pisier:1996:SPCb

[Pis96k] Gilles Pisier. Similarity Problems and Completely Bounded Maps, volume 1618 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 156 pp. LCCN QA3 .L28 no. 1618. URL http://link.

springer.com/book/10.1007/978-3-662-21537-1;   http:/
/www.springerlink.com/content/978-3-662-21537-1.

Pisier:1996:NIA

[Pis96l] Gilles Pisier. Von Neumann’s inequality and Ando’s generalization. Lecture Notes in Mathematics, 1618:13–29, 1996. CODEN LNMAA2. ISBN 3-540-60322-0 (print), 3-662-21537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21537-1_
2/.

Pitman:1999:DLT

[Pit99] Jim Pitman. The distribution of local times of a Brownian bridge. Lecture Notes in Mathematics, 1709:388–394, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096528/.

Liu:1991:WSH

[pL91] He ping Liu. The weak H^{p} spaces on homogeneous groups. Lecture Notes in Mathematics, 1494:113–118, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087762/.

Platek:1990:MCS

[Pla90] Richard A. Platek. Making computers safe for the world: an introduction to proofs of programs part I. Lecture Notes in Mathematics, 1429:60–89, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093924/.

Polotovskii:1992:CDP

[Pol92] G. M. Polotovskiĭ. On the classification of decomposing plane algebraic curves. Lecture Notes in Mathematics, 1524:52–74, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084608/.

Pollard:1993:FCI

[Pol93a] J. M. Pollard. Factoring with cubic integers. Lecture Notes in Mathematics, 1554:4–10, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091536/.

Pollard:1993:LS

[Pol93b] J. M. Pollard. The lattice sieve. Lecture Notes in Mathematics, 1554:43–49, 1993. CODEN LNMAA2. ISBN 3-540-57013-6 (print), 3-540-47892-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091538/.

Pommaret:1991:DTA

[Pom91] J. F. Pommaret. Deformation theory of algebraic and geometric structures. Lecture Notes in Mathematics, 1478:244–254, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083506/.

Pomp:1998:BDI

[Pom98a] Andreas Pomp. The Boundary-Domain Integral Method for Elliptic Systems, volume 1683 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 163 pp. LCCN QA3 .L28 no. 1683.

URL http://link.springer.com/book/10.1007/BFb0094576;
 http://www.springerlink.com/content/978-3-540-69697-

Pomp:1998:DED

[Pom98b] Andreas Pomp. The differential equations of the DV model. Lecture Notes in Mathematics, 1683:69–88, 1998. CODEN LN-MAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094580/.

Pomp:1998:EKS

[Pom98c] Andreas Pomp. An example: Katenoid shell under torsion. Lecture Notes in Mathematics, 1683:137–153, 1998. CODEN LN-MAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094583/.

Pomp:1998:LFE

[Pom98d] Andreas Pomp. Levi functions for elliptic systems of partial differential equations. Lecture Notes in Mathematics, 1683:29–45, 1998. CODEN LNMAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094578/.

Pomp:1998:LFS

[Pom98e] Andreas Pomp. Levi functions for the shell equations. Lecture Notes in Mathematics, 1683:89–108, 1998. CODEN LN-MAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094581/.

Pomp:1998:PD

[Pom98f] Andreas Pomp. Pseudohomogeneous distributions. Lecture Notes in Mathematics, 1683:1–27, 1998. CODEN LNMAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094577/.

Pomp:1998:SIEb

[Pom98g] Andreas Pomp. The system of integral equations and its numerical solution. Lecture Notes in Mathematics, 1683:109–135, 1998. CODEN LNMAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094582/.

Pomp:1998:SIEa

[Pom98h] Andreas Pomp. Systems of integral equations, generated by Levi functions. Lecture Notes in Mathematics, 1683:47–68, 1998. CO-DEN LNMAA2. ISBN 3-540-64163-7 (print), 3-540-69697-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094579/.

Pott:1995:CS

[Pot95a] Alexander Pott. Codes and sequences. Lecture Notes in Mathematics, 1601:149–168, 1995. CODEN LNMAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094455/.

Pott:1995:DSC

[Pot95b] Alexander Pott. Difference sets with classical parameters. Lecture Notes in Mathematics, 1601:69–102, 1995. CODEN LN-MAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094452/.

Pott:1995:EEN

[Pot95c] Alexander Pott. Examples: Existence and non-existence. Lecture Notes in Mathematics, 1601:35–68, 1995. CODEN LN-MAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094451/.

Pott:1995:FGC

[Pot95d] Alexander Pott. Finite Geometry and Character Theory, volume 1601 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 181 pp. LCCN QA3 .L28 no. 1601. URL http:/

/link.springer.com/book/10.1007/BFb0094449;   http://
www.springerlink.com/content/978-3-540-49182-8.

Pott:1995:PIS

[Pot95e] Alexander Pott. Preliminaries: Incidence structures with Singer groups. Lecture Notes in Mathematics, 1601:1–33, 1995. CODEN LNMAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094450/.

Pott:1995:PPQ

[Pot95f] Alexander Pott. Projective planes with quasiregular collineation groups. Lecture Notes in Mathematics, 1601:113–147, 1995. CO-DEN LNMAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094454/.

Pott:1995:SRD

[Pot95g] Alexander Pott. Semiregular relative difference sets. Lecture Notes in Mathematics, 1601:103–111, 1995. CODEN LN-MAA2. ISBN 3-540-59065-X (print), 3-540-49182-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094453/.

Pumarino:1997:BPb

[PR97a] Antonio Pumariño and J. Angel Rodríguez. The binding period. Lecture Notes in Mathematics, 1658:135–152, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093345/.

Pumarino:1997:BPa

[PR97b] Antonio Pumariño and J. Angel Rodríguez. The binding point. Lecture Notes in Mathematics, 1658:119–134, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093344/.

Pumarino:1997:CPS

[PR97c] Antonio Pumariño and J. Angel Rodríguez. Coexistence and Persistence of Strange Attractors, volume 1658 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 195 pp. LCCN QA3 .L28 no. 1658. URL http:/

/link.springer.com/book/10.1007/BFb0093337;   http://
www.springerlink.com/content/978-3-540-68496-1.

Pumarino:1997:CD

[PR97d] Antonio Pumariño and J. Angel Rodríguez. Contractive directions. Lecture Notes in Mathematics, 1658:53–72, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093341/.

Pumarino:1997:CPB

[PR97e] Antonio Pumariño and J. Angel Rodríguez. Critical points of the bidimensional map. Lecture Notes in Mathematics, 1658: 73–88, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093342/.

Pumarino:1997:EP

[PR97f] Antonio Pumariño and J. Angel Rodríguez. The exclusion of parameters. Lecture Notes in Mathematics, 1658:153–190, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093346/.

Pumarino:1997:IP

[PR97g] Antonio Pumariño and J. Angel Rodríguez. The inductive process. Lecture Notes in Mathematics, 1658:89–118, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093343/.

Pumarino:1997:I

[PR97h] Antonio Pumariño and J. Angel Rodríguez. Introduction. Lecture Notes in Mathematics, 1658:1–10, 1997. CODEN LN-MAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093338/.

Pumarino:1997:SFC

[PR97i] Antonio Pumariño and J. Angel Rodríguez. Saddle-focus connections. Lecture Notes in Mathematics, 1658:11–20, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093339/.

Pumarino:1997:UF

[PR97j] Antonio Pumariño and J. Angel Rodríguez. The unimodal family. Lecture Notes in Mathematics, 1658:21–52, 1997. CODEN LNMAA2. ISBN 3-540-62731-6 (print), 3-540-68496-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093340/.

Praeger:1990:FPP

[Pra90a] Cheryl E. Praeger. Finite primitive permutation groups: a survey. Lecture Notes in Mathematics, 1456:63–84, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100731/.

Pratelli:1990:LMD

[Pra90b] Luca Pratelli. Sur le lemme de mesurabilité de Doob. (French) []. Lecture Notes in Mathematics, 1426:46–51, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083756/.

Pragacz:1991:AGA

[Pra91] Piotr Pragacz. Algebro-geometric applications of Schur s- and qpolynomials. Lecture Notes in Mathematics, 1478:130–191, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083503/.

Pratelli:1992:CCA

[Pra92] Luca Pratelli. Une caractérisation de la convergence dans L^{1}. Application aux quasimartingales. (French) []. Lecture Notes in Mathematics, 1526:61–69, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084310/.

Pratelli:1994:DCE

[Pra94] Luca Pratelli. Deux contre-exemples sur la convergence d’intégrales anticipatives. (French) []. Lecture Notes in Mathematics, 1583:110–112, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073838/.

Pratelli:1999:APT

[Pra99] Maurizio Pratelli. An alternative proof of a theorem of aldous concerning convergence in distribution for martingales. Lecture Notes in Mathematics, 1709:334–338, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096522/.

Previato:1996:SYS

[Pre96] Emma Previato. Seventy years of spectral curves: 1923–1993. Lecture Notes in Mathematics, 1620:419–481, 1996. CODEN LNMAA2. ISBN 3-540-60542-8 (print), 3-540-47706-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094795/.

Priddy:1990:CSC

[Pri90] Stewart B. Priddy. On characterizing summands in the classifying space of a group, II. Lecture Notes in Mathematics, 1418: 175–183, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083702/.

Pritsker:1993:CBP

[Pri93] I. E. Pritsker. On the convergence of Bieberbach polynomials in domains with interior zero angles. Lecture Notes in Mathematics, 1550:169–172, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0117484/.

Prodinger:1990:FRP

[Pro90a] Helmut Prodinger. Further results on a problem of Knödel concerning the analysis of bin-packing. Lecture Notes in Mathematics, 1452:193–198, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096990/.

Prouse:1990:UCS

[Pro90b] Giovanni Prouse. A uniqueness criterion for the solution of the stationary Navier–Stokes equations. Lecture Notes in Mathematics, 1431:128–133, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086063/.

Proskurin:1998:CMF

[Pro98a] Nikolai Proskurin. Cubic Metaplectic Forms and Theta Functions, volume 1677 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-63751-6 (print), 3-540-69641-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 196 pp. LCCN QA3 .L28 no. 1677. URL http:/

/link.springer.com/book/10.1007/BFb0093533;   http://
www.springerlink.com/content/978-3-540-69641-4.

Proskurin:1998:Pa

[Pro98b] Nikolai Proskurin. Part 0. Lecture Notes in Mathematics, 1677: 1–62, 1998. CODEN LNMAA2. ISBN 3-540-63751-6 (print), 3-540-69641-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093534/.

Proskurin:1998:Pb

[Pro98c] Nikolai Proskurin. Part 1. Lecture Notes in Mathematics, 1677: 63–125, 1998. CODEN LNMAA2. ISBN 3-540-63751-6 (print), 3-540-69641-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093535/.

Proskurin:1998:Pc

[Pro98d] Nikolai Proskurin. Part 2. Lecture Notes in Mathematics, 1677: 126–187, 1998. CODEN LNMAA2. ISBN 3-540-63751-6 (print), 3-540-69641-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093536/.

Papamichael:1990:NPD

[PS90] N. Papamichael and N. S. Stylianopoulos. On the numerical performance of a domain decomposition method for conformal mapping. Lecture Notes in Mathematics, 1435:155–169, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087905/.

Piene:1990:CBR

[PsT90] Ragni Piene and Hsin sheng Tai. A characterization of balanced rational normal scrolls in terms of their osculating spaces. Lecture Notes in Mathematics, 1436:215–224, 1990. CODEN LN-MAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084047/.

Peternell:1992:NEV

[PSW92] Th. Peternell, M. Szurek, and J. A. Wiśniewski. Numerically effective vector bundles with small Chern classes. Lecture Notes in Mathematics, 1507:145–156, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094516/.

Pukhnachov:1990:PMF

[Puk90] V. V. Pukhnachov. The problem of momentumless flow for the Navier–Stokes equations. Lecture Notes in Mathematics, 1431: 87–94, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086058/.

Pulvirenti:1996:KLS

[Pul96] Mario Pulvirenti. Kinetic limits for stochastic particle systems. Lecture Notes in Mathematics, 1627:96–126, 1996. CODEN LN-MAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093178/.

Pusz:1993:CED

[Pus93] Jerzy Pusz. Characterization of exponential distributions by conditional moments. Lecture Notes in Mathematics, 1546:159– 162, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084490/.

Puschnigg:1996:TAC

[Pus96a] Michael Puschnigg. KK-theory and asymptotic cohomology. Lecture Notes in Mathematics, 1642:182–201, 1996. CODEN LN-MAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094468/.

Puschnigg:1996:ARC

[Pus96b] Michael Puschnigg. Algebraic de Rham complexes. Lecture Notes in Mathematics, 1642:19–26, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094460/.

Puschnigg:1996:ACa

[Pus96c] Michael Puschnigg. The analytic X-complex. Lecture Notes in Mathematics, 1642:59–96, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094463/.

Puschnigg:1996:ACCb

[Pus96d] Michael Puschnigg. Asymptotic Cyclic Cohomology, volume 1642 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 238 pp. LCCN QA3 .L28 no. 1642. URL http:/

/link.springer.com/book/10.1007/BFb0094458;   http://
www.springerlink.com/content/978-3-540-49579-6.

Puschnigg:1996:ACCa

[Pus96e] Michael Puschnigg. Asymptotic cyclic cohomology of dense subalgebras. Lecture Notes in Mathematics, 1642:118–126, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094465/.

Puschnigg:1996:AHC

[Pus96f] Michael Puschnigg. The asymptotic homotopy category. Lecture Notes in Mathematics, 1642:1–18, 1996. CODEN LN-MAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094459/.

Puschnigg:1996:ACb

[Pus96g] Michael Puschnigg. The asymptotic X-complex. Lecture Notes in Mathematics, 1642:97–117, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094464/.

Puschnigg:1996:CC

[Pus96h] Michael Puschnigg. Cyclic cohomology. Lecture Notes in Mathematics, 1642:27–39, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094461/.

Puschnigg:1996:ES

[Pus96i] Michael Puschnigg. Exact sequences. Lecture Notes in Mathematics, 1642:158–181, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094467/.

Puschnigg:1996:E

[Pus96j] Michael Puschnigg. Examples. Lecture Notes in Mathematics, 1642:202–231, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094469/.

Puschnigg:1996:HPC

[Pus96k] Michael Puschnigg. Homotopy properties of X-complexes. Lecture Notes in Mathematics, 1642:40–58, 1996. CODEN LN-MAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094462/.

Puschnigg:1996:P

[Pus96l] Michael Puschnigg. Products. Lecture Notes in Mathematics, 1642:127–157, 1996. CODEN LNMAA2. ISBN 3-540-61986-0 (print), 3-540-49579-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094466/.

Pytlak:1999:ESD

[Pyt99a] Radosław Pytlak. Estimates on solutions to differential equations and their approximations. Lecture Notes in Mathematics, 1707: 13–26, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097246/.

Pytlak:1999:FOM

[Pyt99b] Radosław Pytlak. First order method. Lecture Notes in Mathematics, 1707:27–53, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097247/.

Pytlak:1999:Ib

[Pyt99c] Radosław Pytlak. Implementation. Lecture Notes in Mathematics, 1707:55–79, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097248/.

Pytlak:1999:Ia

[Pyt99d] Radosław Pytlak. Introduction. Lecture Notes in Mathematics, 1707:1–12, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097245/.

Pytlak:1999:NMO

[Pyt99e] Radosław Pytlak. Numerical Methods for Optimal Control Problems with State Constraints, volume 1707 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LN-MAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 215 pp. LCCN QA3 .L28 no. 1707; QA3 .L35 v.1707. URL http:/

/link.springer.com/book/10.1007/BFb0097244;   http://
www.springerlink.com/content/978-3-540-48662-6.

Pytlak:1999:RKB

[Pyt99f] Radosław Pytlak. Runge–Kutta based procedure for optimal control of differential–algebraic equations. Lecture Notes in Mathematics, 1707:129–168, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097250/.

Pytlak:1999:SOM

[Pyt99g] Radosław Pytlak. Second order method. Lecture Notes in Mathematics, 1707:81–128, 1999. CODEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097249/.

Qian:1995:HOU

[QH95] Zhongmin Qian and Sheng-Wu He. On the hypercontractivity of Ornstein–Uhlenbeck semigroups with drift. Lecture Notes in Mathematics, 1613:202–217, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094212/.

Quarteroni:1998:ANA

[Qua98] Alfio Quarteroni, editor. Advanced Numerical Approximation of Nonlinear Hyperbolic Equations: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro, Italy, June 23–28, 1997, volume 1697 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1697.

URL http://link.springer.com/book/10.1007/BFb0096351;
 http://www.springerlink.com/content/978-3-540-49804-
9.

Rabiger:1991:LUE

[Räb91] Frank Räbiger. Lower and upper 2-estimates for order bounded sequences and Dunford–Pettis operators between certain classes of Banach lattices. Lecture Notes in Mathematics, 1470:159– 170, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090219/.

Rainer:1996:PDS

[Rai96] Catherine Rainer. Projection d’une diffusion sur sa filtration lente. (French) []. Lecture Notes in Mathematics, 1626:228– 242, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094651/.

Rajeev:1990:SMA

[Raj90] B. Rajeev. On semi-martingales associated with crossings. Lecture Notes in Mathematics, 1426:107–116, 1990. CODEN LN-MAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083760/.

Rajagopal:1991:SRR

[Raj91] K. R. Rajagopal. Some recent results on swirling flows of Newtonian and non-Newtonian fluids. Lecture Notes in Mathematics, 1460:171–185, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091367/.

Rajeev:1996:FOC

[Raj96] B. Rajeev. First order calculus and last entrance times. Lecture Notes in Mathematics, 1626:261–287, 1996. CODEN LN-MAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094654/.

Rakhmanov:1993:SAO

[Rak93] E. A. Rakhmanov. Strong asymptotics for orthogonal polynomials. Lecture Notes in Mathematics, 1550:71–97, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117475/.

Rannacher:1990:NAN

[Ran90a] Rolf Rannacher. On the numerical analysis of the nonstationary Navier–Stokes equations. Lecture Notes in Mathematics, 1431: 180–193, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086069/.

Ransford:1990:PSS

[Ran90b] T. J. Ransford. Predictable sets and set-valued processes. Lecture Notes in Mathematics, 1426:41–45, 1990. CODEN LN-MAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083755/.

Rannacher:1992:CPM

[Ran92] Rolf Rannacher. On Chorin’s projection method for the incompressible Navier–Stokes equations. Lecture Notes in Mathematics, 1530:167–183, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090341/.

Raugel:1995:BM

[Rau95a] Geneviève Raugel. Back matter. Lecture Notes in Mathematics, 1609:??, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-49415-7/1.pdf.

Raugel:1995:DPD

[Rau95b] Geneviève Raugel. Dynamics of partial differential equations on thin domains. Lecture Notes in Mathematics, 1609:208–315, 1995. CODEN LNMAA2. ISBN 3-540-60047-7 (print), 3-540-49415-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095241/.

Rao:1992:NMP

[RB92] Geetha S. Rao and T. L. Bhaskaramurthi. Nonexpansive mappings and proximinality in normed almost linear spaces. Lecture Notes in Mathematics, 1511:80–87, 1992. CODEN LN-MAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093799/.

Reeder:1990:MSS

[Ree90] Mark Reeder. Modular symbols and the Steinberg representation. Lecture Notes in Mathematics, 1447:287–302, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085734/.

Reid:1990:IVE

[Rei90a] Miles Reid. Infinitesimal view of extending a hyperplane sectiondeformation theory and computer algebra. Lecture Notes in Mathematics, 1417:214–286, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083344/.

Reid:1990:QTC

[Rei90b] Miles Reid. Quadrics through a canonical surface. Lecture Notes in Mathematics, 1417:191–213, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083343/.

Reider:1990:TAJ

[Rei90c] Igor Reider. Toward Abel–Jacobi theory for higher dimensional varieties. Lecture Notes in Mathematics, 1417:287–300, 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083345/.

Reithmeier:1991:DDSa

[Rei91a] Eduard Reithmeier. Differentiable dynamical systems. Lecture Notes in Mathematics, 1483:9–109, 1991. CODEN LN-MAA2. ISBN 3-540-54512-3 (print), 3-540-38427-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094523/.

Reithmeier:1991:DDSb

[Rei91b] Eduard Reithmeier. Differentiable dynamical systems with discontinuities. Lecture Notes in Mathematics, 1483:110–151, 1991. CODEN LNMAA2. ISBN 3-540-54512-3 (print), 3-540-38427-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094524/.

Reithmeier:1991:I

[Rei91c] Eduard Reithmeier. Introduction. Lecture Notes in Mathematics, 1483:3–8, 1991. CODEN LNMAA2. ISBN 3-540-54512-3 (print), 3-540-38427-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094522/.

Reithmeier:1991:PSN

[Rei91d] Eduard Reithmeier. Periodic Solutions of Nonlinear Dynamical Systems: Numerical Computation, Stability, Bifurcation and Transition to Chaos, volume 1483 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54512-3 (print), 3-540-38427-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 171 pp. LCCN QA3 .L28 no. 1483.

URL http://link.springer.com/book/10.1007/BFb0094521;
 http://www.springerlink.com/content/978-3-540-38427-
4.

Reimann:1997:I

[Rei97a] Harry Reimann. Introduction. Lecture Notes in Mathematics, 1657:1–8, 1997. CODEN LNMAA2. ISBN 3-540-62645-X (print), 3-540-68414-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093996/.

Reimann:1997:P

[Rei97b] Harry Reimann. Part I. Lecture Notes in Mathematics, 1657: 9–65, 1997. CODEN LNMAA2. ISBN 3-540-62645-X (print), 3-540-68414-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093997/.

Reimann:1997:PI

[Rei97c] Harry Reimann. Part II. Lecture Notes in Mathematics, 1657: 66–88, 1997. CODEN LNMAA2. ISBN 3-540-62645-X (print), 3-540-68414-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093998/.

Reimann:1997:SSZ

[Rei97d] Harry Reimann. The semi-simple zeta function of quaternionic Shimura varieties, volume 1657 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62645-X (print), 3-540-68414-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 143 pp. LCCN QA3 .L28 no. 1657.

URL http://link.springer.com/book/10.1007/BFb0093995;
 http://www.springerlink.com/content/978-3-540-68414-
5.

Reshetnyak:1992:FCI

[Res92] Yu. G. Reshetnyak. On functional classes invariant relative to homotheties. Lecture Notes in Mathematics, 1508:80–92, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094239/.

Reznikov:1991:SII

[Rez91] Alexander G. Reznikov. A strengthened isoperimetric inequality for simplices. Lecture Notes in Mathematics, 1469:90–93, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089216/.

Richter-Gebert:1996:ACT

[RG96a] Jürgen Richter-Gebert. Alternative construction techniques. Lecture Notes in Mathematics, 1643:149–172, 1996. CODEN LN-MAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093767/.

Richter-Gebert:1996:AU

[RG96b] Jürgen Richter-Gebert. Applications of university. Lecture Notes in Mathematics, 1643:77–115, 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093765/.

Richter-Gebert:1996:I

[RG96c] Jürgen Richter-Gebert. Introduction. Lecture Notes in Mathematics, 1643:1–11, 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093762/.

Richter-Gebert:1996:OT

[RG96d] Jürgen Richter-Gebert. The objects and the tools. Lecture Notes in Mathematics, 1643:13–40, 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093763/.

Richter-Gebert:1996:P

[RG96e] Jürgen Richter-Gebert. Problems. Lecture Notes in Mathematics, 1643:173–177, 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093768/.

Richter-Gebert:1996:RSP

[RG96f] Jürgen Richter-Gebert. Realization Spaces of Polytopes, volume 1643 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 187 pp. LCCN QA3 .L28 no. 1643. URL http:/

/link.springer.com/book/10.1007/BFb0093761;   http://
www.springerlink.com/content/978-3-540-49640-3.

Richter-Gebert:1996:TDP

[RG96g] Jürgen Richter-Gebert. Three-dimensional polytopes. Lecture Notes in Mathematics, 1643:117–147, 1996. CODEN LN-MAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093766/.

Richter-Gebert:1996:UT

[RG96h] Jürgen Richter-Gebert. The universality theorem. Lecture Notes in Mathematics, 1643:41–76, 1996. CODEN LNMAA2. ISBN 3-540-62084-2 (print), 3-540-49640-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093764/.

Ribet:1999:TPG

[Rib99] Kenneth A. Ribet. Torsion points on J_{0}(N) and Galois representations. Lecture Notes in Mathematics, 1716:145–166, 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093454/.

Rickman:1992:PTD

[Ric92] Seppo Rickman. Picard’s theorem and defect relation for quasiregular mappings. Lecture Notes in Mathematics, 1508:93–103, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094240/.

Rickard:1998:TCM

[Ric98] Jeremy Rickard. Triangulated categories in the modular representation theory of finite groups. Lecture Notes in Mathematics, 1685:177–198, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096375/.

Ricker:1999:ABA

[Ric99a] Werner Ricker. Abstract Boolean algebras and stone spaces. Lecture Notes in Mathematics, 1711:25–40, 1999. CODEN LN-MAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096186/.

Ricker:1999:BFA

[Ric99b] Werner Ricker. Bade functionals: an application to scalartype spectral operators. Lecture Notes in Mathematics, 1711: 91–104, 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096190/.

Ricker:1999:BAP

[Ric99c] Werner Ricker. Boolean algebras of projections and uniformly closed operator algebras. Lecture Notes in Mathematics, 1711: 41–56, 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096187/.

Ricker:1999:IRS

[Ric99d] Werner Ricker. Integral representation of the strongly closed algebra generated by a Boolean algebra of projections. Lecture Notes in Mathematics, 1711:67–90, 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096189/.

Ricker:1999:OAG

[Ric99e] Werner Ricker. Operator Algebras Generated by Commuting Projections: a Vector Measure Approach, volume 1711 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 159 pp. LCCN QA3 .L28 no. 1711. URL http:/

/link.springer.com/book/10.1007/BFb0096184;   http://
www.springerlink.com/content/978-3-540-48279-6.

Ricker:1999:RSM

[Ric99f] Werner Ricker. Ranges of spectral measures and Boolean algebras of projections. Lecture Notes in Mathematics, 1711:57–66, 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096188/.

Ricker:1999:RTB

[Ric99g] Werner Ricker. The reflexivity theorem and bicommutant algebras. Lecture Notes in Mathematics, 1711:105–119, 1999. CODEN LNMAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096191/.

Ricker:1999:VMB

[Ric99h] Werner Ricker. Vector measures and Banach spaces. Lecture Notes in Mathematics, 1711:1–24, 1999. CODEN LN-MAA2. ISBN 3-540-66461-0 (print), 3-540-48279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096185/.

Rivkind:1990:NMN

[Riv90] Valeriy Ya. Rivkind. Numerical methods for the Navier–Stokes equations with an unknown boundary between two viscous incompressible fluids. Lecture Notes in Mathematics, 1431:194– 200, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086070/.

Rossello-Llompart:1990:CR

[RL90] Francesc Rosselló-Llompart. The Chow ring of Hilb^{3}P^{3}. Lecture Notes in Mathematics, 1436:225–255, 1990. CODEN LN-MAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084048/.

Robson:1990:SFA

[Rob90] Robert O. Robson. Slices: Functions for abstract real analysis. Lecture Notes in Mathematics, 1420:206–222, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083922/.

Rockafellar:1990:NAP

[Roc90] R. T. Rockafellar. Nonsmooth analysis and parametric optimization. Lecture Notes in Mathematics, 1446:137–151, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084934/.

Rockner:1993:GTD

[Röc93] Michael Röckner. General theory of Dirichlet forms and applications. Lecture Notes in Mathematics, 1563:129–193, 1993. CO-DEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074093/.

Rockner:1999:AFI

[Röc99] Michael Röckner. L^{p}-analysis of finite and infinite dimensional diffusion operators. Lecture Notes in Mathematics, 1715:65– 116, 1999. CODEN LNMAA2. ISBN 3-540-66545-5 (print), 3-540-48161-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092418/.

Rodier:1992:MCS

[Rod92] François Rodier. Minoration de certaines sommes exponentielles binaires. (French) []. Lecture Notes in Mathematics, 1518:199– 209, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088003/.

Rodrigues:1994:VMS

[Rod94] José-Francisco Rodrigues. Variational methods in the Stefan problem. Lecture Notes in Mathematics, 1584:147–212, 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073397/.

Roger:1990:DUC

[Rog90] Claude Roger. Déformations universelles des crochets de Poissonxo. (French) []. Lecture Notes in Mathematics, 1416:242– 254, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097475/.

Rohlfs:1990:LNA

[Roh90] Jürgen Rohlfs. Lefschetz numbers for arithmetic groups. Lecture Notes in Mathematics, 1447:303–313, 1990. CODEN LN-MAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085735/.

Rosebrock:1990:RSD

[Ros90] Stephan Rosebrock. A reduced spherical diagram into a ribbondisk complement and related examples. Lecture Notes in Mathematics, 1440:175–185, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084461/.

Rosenberg:1991:KAM

[Ros91] Jonathan Rosenberg. The ko-assembly map and positive scalar curvature. Lecture Notes in Mathematics, 1474:170–182, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084745/.

Rosenthal:1994:GOT

[Ros94] Peter Rosenthal. General operator theory. Lecture Notes in Mathematics, 1573:211–258, 1994. CODEN LNMAA2. ISBN 3-540-57870-6 (print), 3-540-48367-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100206/.

Rousseau:1990:UUS

[Rou90] Christiane Rousseau. Universal unfolding of a singularity of a symmetric vector field with 7-jet C^{∞}-equivalent to y∂/∂x+(±x^{3}± x^{6}y)∂/∂y. Lecture Notes in Mathematics, 1455:334–355, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085399/.

Rouquier:1998:DCB

[Rou98] Raphaël Rouquier. The derived category of blocks with cyclic defect groups. Lecture Notes in Mathematics, 1685:199–220, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096376/.

Rees:1990:CTE

[RS90a] Elmer Rees and Christopher Smyth. On the constant in the Tarry–Escott problem. Lecture Notes in Mathematics, 1415:196– 208, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084888/.

Robbiano:1990:SB

[RS90b] Lorenzo Robbiano and Moss Sweedler. Subalgebra bases. Lecture Notes in Mathematics, 1430:61–87, 1990. CODEN LN-MAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085537/.

Rohlfs:1990:BCL

[RS90c] Jürgen Rohlfs and Birgit Speh. Boundary contributions to Lefschetz numbers for arithmetic groups I. Lecture Notes in Mathematics, 1447:315–332, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085736/.

Rothe:1990:BQP

[RS90d] Franz Rothe and Douglas S. Shafer. Bifurcation in a quartic polynomial system arising in biology. Lecture Notes in Mathematics, 1455:356–368, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085400/.

Roy:1990:CCC

[RS90e] M.-F. Roy and A. Szpirglas. Complexity of the computation of cylindrical decomposition and topology of real algebraic curves using Thom’s lemma. Lecture Notes in Mathematics, 1420:223– 236, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083923/.

Roberts:1991:STA

[RS91a] Mark Roberts and Ian Stewart, editors. Singularity Theory and its Applications: Warwick 1989, Part II: Singularities, Bifurcations and Dynamics, volume 1463 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1463. URL http:/

/link.springer.com/book/10.1007/BFb0085422;   http://
www.springerlink.com/content/978-3-540-47047-2.

Rosenthal:1991:TPO

[RS91b] H. P. Rosenthal and S. J. Szarek. On tensor products of operators from L^{p} to L^{q}. Lecture Notes in Mathematics, 1470:108–132, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090217/.

Rahman:1993:CBR

[RS93] Mizan Rahman and S. K. Suslov. Classical biorthogonal rational functions. Lecture Notes in Mathematics, 1550:131–146, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0117478/.

Rabeherimanana:1994:PPS

[RS94] T. J. Rabeherimanana and S. N. Smirnov. Petites perturbations de systèmes dynamiques et algèbres de Lie nilpotentes. Une extension des estimations de Doss & Stroock. (French) []. Lecture Notes in Mathematics, 1583:49–72, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073833/.

Ruscheweyh:1990:CMF

[RSSV90] Stephan Ruscheweyh, Edward B. Saff, Luis C. Salinas, and Richard S. Varga, editors. Computational Methods and Function Theory: Proceedings of a Conference, held in Valparaíso, Chile, March 13–18, 1989, volume 1435 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1435. URL http:/

/link.springer.com/book/10.1007/BFb0087892;   http://
www.springerlink.com/content/978-3-540-47139-4.

Rubin:1999:BM

[Rub99a] Karl Rubin. Back matter. Lecture Notes in Mathematics, 1716: ??, 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-48160-7/1.pdf.

Rubin:1999:ECC

[Rub99b] Karl Rubin. Elliptic curves with complex multiplication and the conjecture of Birch and Swinnerton-Dyer. Lecture Notes in Mathematics, 1716:167–234, 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093455/.

Rudyak:1992:PRH

[Rud92] Yu. B. Rudyak. The problem of realization of homology classes from Poincaré up to the present. Lecture Notes in Mathematics, 1520:93–110, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084717/.

Ruggeri:1996:RMM

[Rug96] Tommaso Ruggeri, editor. Recent Mathematical Methods in Nonlinear Wave Propagation: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.), held in Montecatini Terme, Italy, May 23–31, 1994, volume 1640 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61907-0 (print), 3-540-49565-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1640. URL http:/

/link.springer.com/book/10.1007/BFb0093704;   http://
www.springerlink.com/content/978-3-540-49565-9.

Ruiz:1990:TGS

[Ruí90] Jesús M. Ruíz. On the topology of global semianalytic sets. Lecture Notes in Mathematics, 1420:237–246, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083924/.

Runggaldier:1997:FML

[Run97] Wolfgang J. Runggaldier, editor. Financial Mathematics: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Bressanone, Italy, July 8–13, 1996, volume 1656 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62642-5 (print), 3-540-68356-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1656. URL http:/

/link.springer.com/book/10.1007/BFb0091997;   http://
www.springerlink.com/content/978-3-540-68356-8.

Russo:1990:LEC

[Rus90] Francesco Russo. Linear extrapolation concerning Hilbert valued planar functions. Lecture Notes in Mathematics, 1444:244– 268, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083619/.

Rutter:1990:SPH

[Rut90] John W. Rutter. On skeleton preserving homotopy selfequivalences of CW complexes. Lecture Notes in Mathematics, 1425:147–156, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083838/.

Rutter:1997:FPN

[Rut97a] John W. Rutter. ϵ^{∗}(X) finitely presented, nilpotent. Lecture Notes in Mathematics, 1662:40–43, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093745/.

Rutter:1997:FG

[Rut97b] John W. Rutter. ϵ^{∗}(X) not finitely generated. Lecture Notes in Mathematics, 1662:34–35, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093743.

Rutter:1997:BB

[Rut97c] John W. Rutter. Building blocks. Lecture Notes in Mathematics, 1662:4–6, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093738/.

Rutter:1997:CHCb

[Rut97d] John W. Rutter. Cellular, homology complexes: calculations. Lecture Notes in Mathematics, 1662:54–62, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093748/.

Rutter:1997:CHCa

[Rut97e] John W. Rutter. Cellular/homology complexes: methods. Lecture Notes in Mathematics, 1662:45–53, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093747/.

Rutter:1997:GSM

[Rut97f] John W. Rutter. Generators: surface, modular groups. Lecture Notes in Mathematics, 1662:19–27, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093741/.

Rutter:1997:GTP

[Rut97g] John W. Rutter. Group theoretic properties. Lecture Notes in Mathematics, 1662:108–112, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093754/.

Rutter:1997:HAS

[Rut97h] John W. Rutter. Homotopy automorphisms of H-spaces. Lecture Notes in Mathematics, 1662:121–123, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093756/.

Rutter:1997:HSC

[Rut97i] John W. Rutter. Homotopy systems, chain complexes. Lecture Notes in Mathematics, 1662:74–79, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093750/.

Rutter:1997:HTH

[Rut97j] John W. Rutter. Homotopy type, homotopy groups. Lecture Notes in Mathematics, 1662:113–120, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093755/.

Rutter:1997:RD

[Rut97k] John W. Rutter. L-R duality. Lecture Notes in Mathematics, 1662:44, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093746.

Rutter:1997:L

[Rut97l] John W. Rutter. Localization. Lecture Notes in Mathematics, 1662:36–39, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093744/.

Rutter:1997:MDT

[Rut97m] John W. Rutter. Manifolds of dimension three or more. Lecture Notes in Mathematics, 1662:28–33, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093742/.

Rutter:1997:NCP

[Rut97n] John W. Rutter. Non-1-connected Postnikov: methods. Lecture Notes in Mathematics, 1662:63–73, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093749/.

Rutter:1997:NCS

[Rut97o] John W. Rutter. Non-1-connected spaces: calculations. Lecture Notes in Mathematics, 1662:80–93, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093751/.

Rutter:1997:P

[Rut97p] John W. Rutter. Preliminaries. Lecture Notes in Mathematics, 1662:1–3, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093737/.

Rutter:1997:RHH

[Rut97q] John W. Rutter. Representations: homology and homotopy. Lecture Notes in Mathematics, 1662:7–10, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093739/.

Rutter:1997:SHS

[Rut97r] John W. Rutter. Spaces of Homotopy Self-Equivalences: A Survey, volume 1662 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 170 pp. LCCN QA3 .L28 no. 1662. URL http:

//link.springer.com/book/10.1007/BFb0093736;   http:/
/www.springerlink.com/content/978-3-540-69135-8.

Rutter:1997:S

[Rut97s] John W. Rutter. Surfaces. Lecture Notes in Mathematics, 1662: 11–18, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093740/.

Rutter:1997:UP

[Rut97t] John W. Rutter. Unions and products. Lecture Notes in Mathematics, 1662:98–107, 1997. CODEN LNMAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093753/.

Rutter:1997:WTS

[Rut97u] John W. Rutter. Whitehead torsion, simple homotopy. Lecture Notes in Mathematics, 1662:94–97, 1997. CODEN LN-MAA2. ISBN 3-540-63103-8 (print), 3-540-69135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093752/.

Rizvi:1990:CSM

[RY90] S. Tariq Rizvi and Mohamed F. Yousif. On continuous and singular modules. Lecture Notes in Mathematics, 1448:116–124, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091256/.

Sabatier:1991:RIC

[Sab91] P. C. Sabatier. Remark on the informative content of few measurements. Lecture Notes in Mathematics, 1497:187–193, 1991. CODEN LNMAA2. ISBN 3-540-54970-6 (print), 3-540-46615-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084518/.

Sacks:1990:PP

[Sac90] Gerald E. Sacks. Prolog programming. Lecture Notes in Mathematics, 1429:90–110, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093925/.

Sadowski:1991:ESA

[Sad91] Michał Sadowski. Equivariant splittings associated with smooth toral actions. Lecture Notes in Mathematics, 1474:183–192, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084746/.

Saffari:1990:EPP

[Saf90] Bahman Saffari. Extremal problems on polynomials. Lecture Notes in Mathematics, 1415:209–211, 1990. CODEN LN-MAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084889/.

Saito:1991:HEI

[Sai91a] Kyoji Saito. Higher Eichler integrals and vector bundles over the moduli of spinned Riemann surfaces. Lecture Notes in Mathematics, 1468:408–421, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086202/.

Saito:1991:HCM

[Sai91b] Morihiko Saito. Hodge conjecture and mixed motives II. Lecture Notes in Mathematics, 1479:196–215, 1991. CODEN LN-MAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086271/.

Sakai:1990:RSM

[Sak90] Fumio Sakai. Reider–Serrano’s method on normal surfaces. Lecture Notes in Mathematics, 1417:301–319, 1990. CODEN LN-MAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083346/.

Salvi:1992:EPS

[Sal92] Rodolfo Salvi. The exterior problem for the stationary Navier– Stokes equations: On the existence and regularity. Lecture Notes in Mathematics, 1530:132–145, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090338/.

Samborski:1991:RPV

[Sam91] S. N. Samborski. Rivers from the point of view of the qualitative theory. Lecture Notes in Mathematics, 1493:168–180, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085030/.

Sapronov:1990:BSR

[Sap90] Yu. I. Sapronov. The bifurcation of stationary rotations of a multidimensional asymmetric rigid body from the sleeping top regime. Lecture Notes in Mathematics, 1453:87–100, 1990. CO-DEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085948/.

Sarason:1992:WCH

[Sar92] Donald Sarason. Weak compactness of holomorphic composition operators on H^{1}. Lecture Notes in Mathematics, 1511:75– 79, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093798/.

Sommese:1990:AGP

[SBL90] Andrew John Sommese, Aldo Biancofiore, and Elvira Laura Livorni, editors. Algebraic Geometry: Proceedings of the International Conference held in L’Aquila, Italy, May 30–June 4, 1988, volume 1417 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52217-4 (print), 3-540-46934-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1417. URL http:/

/link.springer.com/book/10.1007/BFb0083328;   http://
www.springerlink.com/content/978-3-540-46934-6.

Saloff-Coste:1997:LFM

[SC97] Laurent Saloff-Coste. Lectures on finite Markov chains. Lecture Notes in Mathematics, 1665:301–413, 1997. CODEN LN-MAA2. ISBN 3-540-63190-9 (print), 3-540-69210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092621/.

Scedrov:1990:BM

[Sce90a] Andre Scedrov. Back matter. Lecture Notes in Mathematics, 1429:??, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47135-6/1.pdf.

Scedrov:1990:GPT

[Sce90b] Andre Scedrov. A guide to polymorphic types. Lecture Notes in Mathematics, 1429:111–150, 1990. CODEN LNMAA2. ISBN 3-540-52734-6 (print), 3-540-47135-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093926/.

Schaaf:1990:DBB

[Sch90a] Renate Schaaf. Dirichlet branches bifurcating from zero. Lecture Notes in Mathematics, 1458:1–44, 1990. CODEN LN-MAA2. ISBN 3-540-53514-4 (print), 3-540-46742-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098347/.

Schaaf:1990:GPT

[Sch90b] Renate Schaaf. General properties of time maps. Lecture Notes in Mathematics, 1458:110–136, 1990. CODEN LNMAA2. ISBN 3-540-53514-4 (print), 3-540-46742-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098350/.

Schaaf:1990:G

[Sch90c] Renate Schaaf. Generalizations. Lecture Notes in Mathematics, 1458:69–109, 1990. CODEN LNMAA2. ISBN 3-540-53514-4 (print), 3-540-46742-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098349/.

Schaaf:1990:GSB

[Sch90d] Renate Schaaf. Global Solution Branches of Two Point Boundary Value Problems, volume 1458 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53514-4 (print), 3-540-46742-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 141 pp. LCCN QA3 .L28 no. 1458.

URL http://link.springer.com/book/10.1007/BFb0098346;
 http://www.springerlink.com/content/978-3-540-46742-
7.

Schaaf:1990:NPP

[Sch90e] Renate Schaaf. Neumann problems, period maps and semilinear Dirichlet problems. Lecture Notes in Mathematics, 1458:45– 68, 1990. CODEN LNMAA2. ISBN 3-540-53514-4 (print), 3-540-46742-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0098348/.

Schaefer:1990:BST

[Sch90f] H. H. Schaefer. A Banach–Steinhaus theorem for weak and order continuous operators. Lecture Notes in Mathematics, 1422: 93–100, 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089408/.

Schenzel:1990:FIT

[Sch90g] Peter Schenzel. Flatness and ideal-transforms of finite type. Lecture Notes in Mathematics, 1430:88–97, 1990. CODEN LN-MAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085538/.

Schikhof:1990:CPA

[Sch90h] Wim H. Schikhof. The complementation property of ℓ^{∞} in padic Banach spaces. Lecture Notes in Mathematics, 1454:342– 350, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0091148/.

Schinzel:1990:CIP

[Sch90i] A. Schinzel. Un critère d’irreductibilité de polynômes. (French) []. Lecture Notes in Mathematics, 1415:212–224, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084890/.

Schlomiuk:1990:AIQ

[Sch90j] Dana Schlomiuk. Algebraic integrals of quadratic systems with a weak focus. Lecture Notes in Mathematics, 1455:373–384, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085402/.

Schober:1990:PHM

[Sch90k] Glenn Schober. Planar harmonic mappings. Lecture Notes in Mathematics, 1435:171–176, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087906/.

Schoissengeier:1990:AE

[Sch90l]  Johannes Schoißengeier. An asymptotic expansion for ∑_{n≤N{nα+β}}.

Lecture Notes in Mathematics, 1452:199–205, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096991/.

Schulz:1990:CMR

[Sch90m] Friedmar Schulz. Conformal mappings with respect to a Riemannian metric. Lecture Notes in Mathematics, 1445:61–71, 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098283/.

Schulz:1990:FTE

[Sch90n] Friedmar Schulz. Function theory of elliptic equations. Lecture Notes in Mathematics, 1445:39–52, 1990. CODEN LN-MAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098281/.

Schulz:1990:ICH

[Sch90o] Friedmar Schulz. Integral criteria for Hölder continuity. Lecture Notes in Mathematics, 1445:1–14, 1990. CODEN LN-MAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098278/.

Schulz:1990:LBS

[Sch90p] Friedmar Schulz. Local behavior of solutions of differential inequalities. Lecture Notes in Mathematics, 1445:72–84, 1990. CO-DEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098284/.

Schulz:1990:PEM

[Sch90q] Friedmar Schulz. A priori estimates for Monge–Ampère equations. Lecture Notes in Mathematics, 1445:94–105, 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098286/.

Schulz:1990:RPE

[Sch90r] Friedmar Schulz. Regularity and a priori estimates for locally convex surfaces. Lecture Notes in Mathematics, 1445:106–114, 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098287/.

Schulz:1990:RLE

[Sch90s] Friedmar Schulz. Regularity for linear elliptic equations and quasilinear systems. Lecture Notes in Mathematics, 1445:15–27, 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098279/.

Schulz:1990:RMA

[Sch90t] Friedmar Schulz. Regularity for Monge–Ampère equations. Lecture Notes in Mathematics, 1445:28–38, 1990. CODEN LN-MAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098280/.

Schulz:1990:RTQ

[Sch90u] Friedmar Schulz. Regularity Theory for Quasilinear Elliptic Systems and Monge–Ampère Equations in Two Dimensions, volume 1445 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 123 pp. LCCN QA3 .L28 no. 1445. URL http:/

/link.springer.com/book/10.1007/BFb0098277;   http://
www.springerlink.com/content/978-3-540-46678-9.

Schulz:1990:USB

[Sch90v] Friedmar Schulz. Univalent solutions of binary elliptic systems. Lecture Notes in Mathematics, 1445:53–60, 1990. CODEN LN-MAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098282/.

Schulz:1990:USH

[Sch90w] Friedmar Schulz. Univalent solutions of Heinz–Lewy type systems. Lecture Notes in Mathematics, 1445:85–93, 1990. CODEN LNMAA2. ISBN 3-540-53103-3 (print), 3-540-46678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098285/.

Schwermer:1990:CAG

[Sch90x] Joachim Schwermer. Cohomology of arithmetic groups, automorphic forms and L-functions. Lecture Notes in Mathematics, 1447: 1–29, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085724/.

Schaumloffel:1991:MET

[Sch91a] Kay-Uwe Schaumlöffel. Multiplicative ergodic theorems in infinite dimensions. Lecture Notes in Mathematics, 1486:187–195, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086668/.

Schlumprecht:1991:LSI

[Sch91b] Thomas Schlumprecht. Limited sets in injective tensor products. Lecture Notes in Mathematics, 1470:133–158, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090218/.

Schmid:1991:FGI

[Sch91c] Barbara J. Schmid. Finite groups and invariant theory. Lecture Notes in Mathematics, 1478:35–66, 1991. CODEN LN-MAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083501/.

Schmidt:1991:DA

[Sch91d] Wolfgang M. Schmidt. Diophantine approximation. Lecture Notes in Mathematics, 1467:34–72, 1991. CODEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098248/.

Schmidt:1991:DAD

[Sch91e] Wolfgang M. Schmidt. Diophantine Approximations and Diophantine Equations, volume 1467 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 217 pp. LCCN QA3 .L28 no. 1467.

URL http://link.springer.com/book/10.1007/BFb0098246;
 http://www.springerlink.com/content/978-3-540-47374-
9.

Schmidt:1991:DEM

[Sch91f] Wolfgang M. Schmidt. Diophantine equations in more than two variables. Lecture Notes in Mathematics, 1467:176–204, 1991. CODEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098251/.

Schmidt:1991:UEH

[Sch91g] Wolfgang M. Schmidt. S-unit equations and hyperelliptic equations. Lecture Notes in Mathematics, 1467:127–175, 1991. CO-DEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098250/.

Schmidt:1991:SLH

[Sch91h] Wolfgang M. Schmidt. Siegel’s lemma and heights. Lecture Notes in Mathematics, 1467:1–33, 1991. CODEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098247/.

Schmidt:1991:TE

[Sch91i] Wolfgang M. Schmidt. The Thue equation. Lecture Notes in Mathematics, 1467:73–126, 1991. CODEN LNMAA2. ISBN 3-540-54058-X (print), 3-540-47374-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0098249/.

Schmuckenschlager:1991:DTJ

[Sch91j] M. Schmuckenschläger. On the dependence on ϵ in a theorem of J. Bourgain, J. Lindenstrauss and V. D. Milman. Lecture Notes in Mathematics, 1469:166–173, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089223/.

Schrauwen:1991:DMN

[Sch91k] Rob Schrauwen. Deformations and the Milnor number of nonisolated plane curve singularities. Lecture Notes in Mathematics, 1462:276–291, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086388/.

Schurmann:1991:AMC

[Sch91l] Michael Schürmann. The Azéma martingales as components of quantum independent increment processes. Lecture Notes in Mathematics, 1485:24–30, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0100842/.

Scheiderer:1992:RAA

[Sch92a] Claus Scheiderer. Real algebra and its applications to geometry in the last ten years: Some major developments and results. Lecture Notes in Mathematics, 1524:75–96, 1992. CODEN LNMAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084609/.

Schonbek:1992:SRA

[Sch92b] Maria E. Schonbek. Some results on the asymptotic behaviour of solutions to the Navier–Stokes equations. Lecture Notes in Mathematics, 1530:146–160, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090339/.

Schurmann:1993:BCF

[Sch93a] Michael Schürmann. Basic concepts and first results. Lecture Notes in Mathematics, 1544:12–40, 1993. CODEN LN-MAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089239/.

Schurmann:1993:I

[Sch93b] Michael Schürmann. Introduction. Lecture Notes in Mathematics, 1544:1–11, 1993. CODEN LNMAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089238/.

Schurmann:1993:LT

[Sch93c] Michael Schürmann. Limit theorems. Lecture Notes in Mathematics, 1544:128–137, 1993. CODEN LNMAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089244/.

Schurmann:1993:QCC

[Sch93d] Michael Schürmann. Quadratic components of conditionally positive linear functionals. Lecture Notes in Mathematics, 1544:114– 127, 1993. CODEN LNMAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089243/.

Schurmann:1993:SWN

[Sch93e] Michael Schürmann. Symmetric white noise on Bose Fock space. Lecture Notes in Mathematics, 1544:41–68, 1993. CODEN LN-MAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089240/.

Schurmann:1993:S

[Sch93f] Michael Schürmann. Symmetrization. Lecture Notes in Mathematics, 1544:69–80, 1993. CODEN LNMAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089241/.

Schurmann:1993:WNBb

[Sch93g] Michael Schürmann. White Noise on Bialgebras, volume 1544 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 146 pp. LCCN QA3 .L28 no. 1544. URL http:/

/link.springer.com/book/10.1007/BFb0089237;   http://
www.springerlink.com/content/978-3-540-47614-6.

Schurmann:1993:WNBa

[Sch93h] Michael Schürmann. White noise on Bose Fock space. Lecture Notes in Mathematics, 1544:81–113, 1993. CODEN LN-MAA2. ISBN 3-540-56627-9 (print), 3-540-47614-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089242/.

Scheiderer:1994:BCT

[Sch94a] Claus Scheiderer. Base change theorems. Lecture Notes in Mathematics, 1588:180–190, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074284/.

Scheiderer:1994:CDR

[Sch94b] Claus Scheiderer. Cohomological dimension of X_{b}, I: Reduction to the field case. Lecture Notes in Mathematics, 1588:68–86, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074276/.

Scheiderer:1994:CDI

[Sch94c] Claus Scheiderer. Cohomological dimension of X_{b}, II: The field case. Lecture Notes in Mathematics, 1588:96–106, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074278/.

Scheiderer:1994:CAV

[Sch94d] Claus Scheiderer. Cohomology of affine varieties. Lecture Notes in Mathematics, 1588:205–211, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074286/.

Scheiderer:1994:CT

[Sch94e] Claus Scheiderer. Comparison theorems. Lecture Notes in Mathematics, 1588:173–179, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074283/.

Scheiderer:1994:CSF

[Sch94f] Claus Scheiderer. Constructible sheaves and finiteness theorems. Lecture Notes in Mathematics, 1588:191–204, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074285/.

Scheiderer:1994:ESA

[Sch94g] Claus Scheiderer. Equivariant sheaves for actions of topological groups. Lecture Notes in Mathematics, 1588:87–95, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074277/.

Scheiderer:1994:EC

[Sch94h] Claus Scheiderer. Examples and complements. Lecture Notes in Mathematics, 1588:219–243, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074288/.

Scheiderer:1994:FLE

[Sch94i] Claus Scheiderer. The fundamental long exact sequence. Lecture Notes in Mathematics, 1588:56–67, 1994. CODEN LN-MAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074275/.

Scheiderer:1994:T

[Sch94j] Claus Scheiderer. G-toposes. Lecture Notes in Mathematics, 1588: 107–127, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074279/.

Scheiderer:1994:GER

[Sch94k] Claus Scheiderer. Glueing étale and real étale site. Lecture Notes in Mathematics, 1588:9–17, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074271/.

Scheiderer:1994:GAS

[Sch94l] Claus Scheiderer. Group actions on spaces: Topological versus topos-theoretic constructions. Lecture Notes in Mathematics, 1588:161–165, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074281/.

Scheiderer:1994:ILT

[Sch94m] Claus Scheiderer. Inverse limits of G-toposes: Two examples. Lecture Notes in Mathematics, 1588:128–160, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074280/.

Scheiderer:1994:LTS

[Sch94n] Claus Scheiderer. Limit theorems, stalks, and other basic facts. Lecture Notes in Mathematics, 1588:18–29, 1994. CODEN LN-MAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074272/.

Scheiderer:1994:QTT

[Sch94o] Claus Scheiderer. Quotient topos of a G-topos, for G of prime order. Lecture Notes in Mathematics, 1588:166–172, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074282/.

Scheiderer:1994:REC

[Sch94p] Claus Scheiderer. Real and Étale Cohomology, volume 1588 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 273 pp. LCCN QA3 .L28 no. 1588. URL http:/

/link.springer.com/book/10.1007/BFb0074269;   http://
www.springerlink.com/content/978-3-540-48797-5.

Scheiderer:1994:RSR

[Sch94q] Claus Scheiderer. Real spectrum and real étale site. Lecture Notes in Mathematics, 1588:1–8, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074270/.

Scheiderer:1994:RSE

[Sch94r] Claus Scheiderer. Real spectrum of X and étale site of X[\sqrt{−1}]. Lecture Notes in Mathematics, 1588:42–55, 1994. CODEN LN-MAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074274/.

Scheiderer:1994:RZT

[Sch94s] Claus Scheiderer. Relations to the Zariski topology. Lecture Notes in Mathematics, 1588:212–218, 1994. CODEN LNMAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074287/.

Scheiderer:1994:SRW

[Sch94t] Claus Scheiderer. Some reminders on Weil restrictions. Lecture Notes in Mathematics, 1588:30–41, 1994. CODEN LN-MAA2. ISBN 3-540-58436-6 (print), 3-540-48797-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074273/.

Schwartz:1994:SMB

[Sch94u] Laurent Schwartz. Semi-martingales banachiques: Le théorème des trois opérateurs. (French) []. Lecture Notes in Mathematics, 1583:1–20, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073830/.

Schwarz:1995:ADF

[Sch95a] Günter Schwarz. Analysis of differential forms. Lecture Notes in Mathematics, 1607:9–58, 1995. CODEN LNMAA2. ISBN 3-540-60016-7 (print), 3-540-49403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095980/.

Schwarz:1995:BVP

[Sch95b] Günter Schwarz. Boundary value problems for differential forms. Lecture Notes in Mathematics, 1607:113–145, 1995. CODEN LNMAA2. ISBN 3-540-60016-7 (print), 3-540-49403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095982/.

Schwarz:1995:HD

[Sch95c] Günter Schwarz. The Hodge decomposition. Lecture Notes in Mathematics, 1607:59–112, 1995. CODEN LNMAA2. ISBN 3-540-60016-7 (print), 3-540-49403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095981/.

Schwarz:1995:HDM

[Sch95d] Günter Schwarz. Hodge Decomposition — a Method for Solving Boundary Value Problems, volume 1607 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60016-7 (print), 3-540-49403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 155 pp. LCCN QA3 .L28 no. 1607.

URL http://link.springer.com/book/10.1007/BFb0095978;
 http://www.springerlink.com/content/978-3-540-49403-
4.

Schwarz:1995:I

[Sch95e] Günter Schwarz. Introduction. Lecture Notes in Mathematics, 1607:1–8, 1995. CODEN LNMAA2. ISBN 3-540-60016-7 (print), 3-540-49403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095979/.

Schachermayer:1999:CPE

[Sch99] W. Schachermayer. On certain probabilities equivalent to Wiener measure, d’après Dubins, Feldman, Smorodinsky and Tsirelson. Lecture Notes in Mathematics, 1709:221–239, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096513/.

Servatius:1990:FBE

[SDS90] Herman Servatius, Carl Droms, and Brigitte Servatius. The finite basis extension property and graph groups. Lecture Notes in Mathematics, 1440:52–58, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084450/.

Segal:1990:RFG

[Seg90] Dan Segal. Residually finite groups. Lecture Notes in Mathematics, 1456:85–95, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100732/.

Seng:1990:SOS

[Sen90] Chew Tuan Seng. The superposition operators in the space of Henstock–Kurzweil integrable functions. Lecture Notes in Mathematics, 1419:19–24, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083096/.

Serre:1992:AF

[Ser92a] Jean-Pierre Serre. Analytic functions. Lecture Notes in Mathematics, 1500:67–75, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70634-2_9/.

Serre:1992:AG

[Ser92b] Jean-Pierre Serre. Analytic groups. Lecture Notes in Mathematics, 1500:102–128, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70634-2_11/.

Serre:1992:AM

[Ser92c] Jean-Pierre Serre. Analytic manifolds. Lecture Notes in Mathematics, 1500:76–101, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70634-2_10/.

Serre:1992:BM

[Ser92d] Jean-Pierre Serre. Back matter. Lecture Notes in Mathematics, 1500:161–172, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-70634-2/1.pdf.

Serre:1992:CF

[Ser92e] Jean-Pierre Serre. Complete fields. Lecture Notes in Mathematics, 1500:64–66, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70634-2_8/.

Serre:1992:FGL

[Ser92f] Jean-Pierre Serre. Filtered groups and Lie algebras. Lecture Notes in Mathematics, 1500:6–10, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70634-2_2/.

Serre:1992:FLA

[Ser92g] Jean-Pierre Serre. Free Lie algebras. Lecture Notes in Mathematics, 1500:18–30, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70634-2_4/.

Serre:1992:FMa

[Ser92h] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 1500:I–VII, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-70634-2/1.pdf.

Serre:1992:FMb

[Ser92i] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 1500:1, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-70634-2/1/1.pdf.

Serre:1992:FMc

[Ser92j] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 1500:63, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-70634-2/2/1.pdf.

Serre:1992:LAL

[Ser92k] Jean-Pierre Serre, editor. Lie Algebras and Lie Groups: 1964 Lectures given at Harvard University, volume 1500 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1500.

URL http://link.springer.com/book/10.1007/978-3-540-
70634-2;   http://www.springerlink.com/content/978-3-
540-70634-2.

Serre:1992:LAD

[Ser92l] Jean-Pierre Serre. Lie algebras: Definition and examples. Lecture Notes in Mathematics, 1500:2–5, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70634-2_1/.

Serre:1992:LT

[Ser92m] Jean-Pierre Serre. Lie theory. Lecture Notes in Mathematics, 1500:129–160, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70634-2_12/.

Serre:1992:NSL

[Ser92n] Jean-Pierre Serre. Nilpotent and solvable Lie algebras. Lecture Notes in Mathematics, 1500:31–43, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-70634-2_5/.

Serre:1992:R

[Ser92o] Jean-Pierre Serre. Representations of ∫↕_{\}. Lecture Notes in Mathematics, 1500:56–62, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70634-2_7/.

Serre:1992:SLA

[Ser92p] Jean-Pierre Serre. Semisimple Lie algebras. Lecture Notes in Mathematics, 1500:44–55, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70634-2_6/.

Serre:1992:UAL

[Ser92q] Jean-Pierre Serre. Universal algebra of a Lie algebra. Lecture Notes in Mathematics, 1500:11–17, 1992. CODEN LNMAA2. ISBN 3-540-55008-9 (print), 3-540-70634-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-70634-2_3/.

Serre:1994:CGP

[Ser94a] Jean-Pierre Serre. Cohomologie des groupes profinis. (French) []. Lecture Notes in Mathematics, 5:1–79, 1994. CODEN LN-MAA2. ISBN 3-540-58002-6 (print), 3-540-24927-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0108759/.

Serre:1994:CGCa

[Ser94b] Jean-Pierre Serre. Cohomologie galoisième — cas commutatif. (French) [Galois cohomology — commutative case]. Lecture Notes in Mathematics, 5:81–126, 1994. CODEN LNMAA2. ISBN 3-540-58002-6 (print), 3-540-24927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0108760/.

Serre:1994:CGF

[Ser94c] Jean-Pierre Serre. Cohomologie Galoisienne. (French) [Galois cohomology], volume 5 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth revised and supplemented edition, 1994. CODEN LNMAA2. ISBN 3-540-58002-6 (print), 3-540-24927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 181 pp. LCCN QA3 .L28 no. 5. URL http:/

/link.springer.com/book/10.1007/BFb0108758;   http://
www.springerlink.com/content/978-3-540-24927-6.

Serre:1994:CGN

[Ser94d] Jean-Pierre Serre. Cohomologie galoisienne non commutative. (French) [Noncommutative Galois cohomology]. Lecture Notes in Mathematics, 5:127–170, 1994. CODEN LNMAA2. ISBN 3-540-58002-6 (print), 3-540-24927-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0108761/.

Serre:1997:ALM

[SG97] Jean-Pierre Serre and Peter Gabriel, editors. Algèbre locale, multiplicités; cours au Collège de France, 1957–1958. (French) [Local algebra, multiplicities; courses at the Collège of France, 1957– 1958], volume 11 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-07028-1 (Berlin), 0-387-07028-1 (New York). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 11. URL http://link.springer.

com/book/10.1007/BFb0087610.

Shekhtman:1993:DPL

[She93] Boris Shekhtman. Duality principle in linearized rational approximation. Lecture Notes in Mathematics, 1550:173–177, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0117485/.

Shiota:1990:PLS

[Shi90] Masahiro Shiota. Piecewise linearization of subanalytic functions II. Lecture Notes in Mathematics, 1420:247–307, 1990. CODEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083925/.

Shiga:1991:HSB

[Shi91a] Hiroshige Shiga. Hardy spaces and BMO on Riemann surfaces. Lecture Notes in Mathematics, 1468:84–93, 1991. CODEN LN-MAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086188/.

Shioda:1991:MWL

[Shi91b] Tetsuji Shioda. Mordell–Weil lattices of type E_{8} and deformation of singularities. Lecture Notes in Mathematics, 1468:177– 202, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086194/.

Shi:1994:LBW

[Shi94] Z. Shi. Liminf behaviours of the windings and Lévy’s stochastic areas of planar Brownian motion. Lecture Notes in Mathematics, 1583:122–137, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073841/.

Shi:1996:HLD

[Shi96] Zhan Shi. How long does it take a transient Bessel process to reach its future infimum? Lecture Notes in Mathematics, 1626: 207–217, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094649/.

Shkolnik:1993:SPM

[Shk93] S. M. Shkol’nik. Some properties of the median of the stable distributions close to the symmetric ones. Lecture Notes in Mathematics, 1546:170–173, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084493/.

Shokranian:1992:AT

[Sho92a] Salahoddin Shokranian. The Adèlic theory. Lecture Notes in Mathematics, 1503:41–44, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092309/.

Shokranian:1992:GET

[Sho92b] Salahoddin Shokranian. The geometric expansion of the trace formula. Lecture Notes in Mathematics, 1503:60–68, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092311/.

Shokranian:1992:GT

[Sho92c] Salahoddin Shokranian. The geometric theory. Lecture Notes in Mathematics, 1503:45–59, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092310/.

Shokranian:1992:ITF

[Sho92d] Salahoddin Shokranian. The invariant trace formula and its applications. Lecture Notes in Mathematics, 1503:79–86, 1992. CO-DEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092313/.

Shokranian:1992:KFC

[Sho92e] Salahoddin Shokranian. Kernel functions and the convergence theorem. Lecture Notes in Mathematics, 1503:24–40, 1992. CO-DEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092308/.

Shokranian:1992:NTA

[Sho92f] Salahoddin Shokranian. Number theory and automorphic representations. Lecture Notes in Mathematics, 1503:1–10, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092306/.

Shokranian:1992:SAT

[Sho92g] Salahoddin Shokranian. The Selberg–Arthur Trace Formula: Based on Lectures by James Arthur, volume 1503 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 97 pp. LCCN QA3 .L28 no. 1503. URL http:/

/link.springer.com/book/10.1007/BFb0092305;   http://
www.springerlink.com/content/978-3-540-46659-8.

Shokranian:1992:STF

[Sho92h] Salahoddin Shokranian. Selberg’s trace formula. Lecture Notes in Mathematics, 1503:11–23, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092307/.

Shokranian:1992:ST

[Sho92i] Salahoddin Shokranian. The spectral theory. Lecture Notes in Mathematics, 1503:69–78, 1992. CODEN LNMAA2. ISBN 3-540-55021-6 (print), 3-540-46659-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092312/.

Short:1992:BT

[Sho92j] Mark W. Short. Background theory. Lecture Notes in Mathematics, 1519:10–42, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090197/.

Short:1992:C

[Sho92k] Mark W. Short. Conclusion. Lecture Notes in Mathematics, 1519: 114–120, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090205/.

Short:1992:ISSa

[Sho92l] Mark W. Short. The imprimitive soluble subgroups of GL(2, p^{k}). Lecture Notes in Mathematics, 1519:43–54, 1992. CODEN LN-MAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090198/.

Short:1992:ISSc

[Sho92m] Mark W. Short. The imprimitive soluble subgroups of GL(4, 2) and GL(4, 3). Lecture Notes in Mathematics, 1519:84–92, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0090202/.

Short:1992:I

[Sho92n] Mark W. Short. Introduction. Lecture Notes in Mathematics, 1519:1–9, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090196/.

Short:1992:ISSb

[Sho92o] Mark W. Short. The irreducible soluble subgroups of GL(2, p^{k}). Lecture Notes in Mathematics, 1519:62–74, 1992. CODEN LN-MAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090200/.

Short:1992:ISSd

[Sho92p] Mark W. Short. The irreducible soluble subgroups of GL(6, 2). Lecture Notes in Mathematics, 1519:108–113, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090204/.

Short:1992:NSC

[Sho92q] Mark W. Short. The normaliser of a Singer cycle of prime degree. Lecture Notes in Mathematics, 1519:55–61, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090199/.

Short:1992:PSPa

[Sho92r] Mark W. Short. The primitive soluble permutation groups of degree less than 256. Lecture Notes in Mathematics, 1519:146, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0120507.

Short:1992:PSPb

[Sho92s] Mark W. Short. The Primitive Soluble Permutation Groups of Degree less than 256, volume 1519 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 145 pp. LCCN QA3 .L28 no. 1519.

URL http://link.springer.com/book/10.1007/BFb0090195;
 http://www.springerlink.com/content/978-3-540-47120-
2.

Short:1992:PSS

[Sho92t] Mark W. Short. The primitive soluble subgroups of GL(4, p^{k}). Lecture Notes in Mathematics, 1519:93–107, 1992. CODEN LN-MAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090203/.

Short:1992:SIS

[Sho92u] Mark W. Short. Some irreducible soluble subgroups of GL(q, p^{k}), q > 2. Lecture Notes in Mathematics, 1519:75–83, 1992. CODEN LNMAA2. ISBN 3-540-55501-3 (print), 3-540-47120-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090201/.

Shpilrain:1990:CFC

[Shp90] V. E. Shpilrain. On the centers of free central extensions of some groups. Lecture Notes in Mathematics, 1456:181–184, 1990.

CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100739/.

Shustin:1992:TRP

[Shu92] E. I. Shustin. Topology of real plane algebraic curves. Lecture Notes in Mathematics, 1524:97–109, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084610/.

Shu:1998:BM

[Shu98a] Chi-Wang Shu. Back matter. Lecture Notes in Mathematics, 1697:??, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-49804-9/1.pdf.

Shu:1998:ENO

[Shu98b] Chi-Wang Shu. Essentially non-oscillatory and weighted essentially non-oscillatory schemes for hyperbolic conservation laws. Lecture Notes in Mathematics, 1697:325–432, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096355/.

Siddiqi:1992:MIL

[Sid92] Jamil A. Siddiqi. Maximal ideals in local Carleman algebras. Lecture Notes in Mathematics, 1511:91–99, 1992. CODEN LN-MAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093801/.

Siersma:1991:VCS

[Sie91] Dirk Siersma. Vanishing cycles and special fibres. Lecture Notes in Mathematics, 1462:292–301, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086389/.

Silhol:1990:CBS

[Sil90] R. Silhol. Classification birationnelle des surfaces rationnelles réelles. Lecture Notes in Mathematics, 1420:308–324, 1990. CO-DEN LNMAA2. ISBN 3-540-52313-8 (print), 3-540-46952-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083926/.

Silhol:1992:MPR

[Sil92] R. Silhol. Moduli problems in real algebraic geometry. Lecture Notes in Mathematics, 1524:110–119, 1992. CODEN LN-MAA2. ISBN 3-540-55992-2 (print), 3-540-47337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084611/.

Simis:1990:TRA

[Sim90] Aron Simis. Topics in Rees algebras of special ideals. Lecture Notes in Mathematics, 1430:98–114, 1990. CODEN LN-MAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085539/.

Simpson:1991:ABM

[Sim91a] Carlos Simpson. Asymptotic Behavior of Monodromy: Singularly Perturbed Differential Equations on a Riemann Surface, volume 1502 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 139 pp. LCCN QA3 .L28 no. 1502. URL http:/

/link.springer.com/book/10.1007/BFb0094551;   http://
www.springerlink.com/content/978-3-540-46641-3.

Simpson:1991:BM

[Sim91b] Carlos Simpson. Bounds on multiplicities. Lecture Notes in Mathematics, 1502:93–100, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094561/.

Simpson:1991:CE

[Sim91c] Carlos Simpson. Complements and examples. Lecture Notes in Mathematics, 1502:111–126, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094563/.

Simpson:1991:CF

[Sim91d] Carlos Simpson. Construction of flows. Lecture Notes in Mathematics, 1502:31–40, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094555/.

Simpson:1991:FL

[Sim91e] Carlos Simpson. Finiteness lemmas. Lecture Notes in Mathematics, 1502:60–67, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094558/.

Simpson:1991:I

[Sim91f] Carlos Simpson. Introduction. Lecture Notes in Mathematics, 1502:1–11, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094552/.

Simpson:1991:LTA

[Sim91g] Carlos Simpson. Laplace transform, asymptotic expansions, and the method of stationary phase. Lecture Notes in Mathematics, 1502:17–30, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094554/.

Simpson:1991:ML

[Sim91h] Carlos Simpson. The main lemma. Lecture Notes in Mathematics, 1502:54–59, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094557/.

Simpson:1991:MRH

[Sim91i] Carlos Simpson. Moving relative homology chains. Lecture Notes in Mathematics, 1502:41–53, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094556/.

Simpson:1991:MCI

[Sim91j] Carlos Simpson. Moving the cycle of integration. Lecture Notes in Mathematics, 1502:84–92, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094560/.

Simpson:1991:ODE

[Sim91k] Carlos Simpson. Ordinary differential equations on a Riemann surface. Lecture Notes in Mathematics, 1502:12–16, 1991. CO-DEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094553/.

Simpson:1991:RIT

[Sim91l] Carlos Simpson. Regularity of individual terms. Lecture Notes in Mathematics, 1502:101–110, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094562/.

Simpson:1991:SC

[Sim91m] Carlos Simpson. Sizes of cells. Lecture Notes in Mathematics, 1502:68–83, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094559/.

Simpson:1991:SLP

[Sim91n] Carlos Simpson. The Sturm–Liouville problem. Lecture Notes in Mathematics, 1502:127–134, 1991. CODEN LNMAA2. ISBN 3-540-55009-7 (print), 3-540-46641-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094564/.

Simons:1998:DCA

[Sim98a] Stephen Simons. A digression into convex analysis. Lecture Notes in Mathematics, 1693:43–51, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093637/.

Simons:1998:DPL

[Sim98b] Stephen Simons. Discontinuous positive linear operators. Lecture Notes in Mathematics, 1693:141–151, 1998. CODEN LN-MAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093642/.

Simons:1998:FAP

[Sim98c] Stephen Simons. Functional analytic preliminaries. Lecture Notes in Mathematics, 1693:13–28, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093635/.

Simons:1998:GMM

[Sim98d] Stephen Simons. General monotone multifunctions. Lecture Notes in Mathematics, 1693:53–73, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093638/.

Simons:1998:I

[Sim98e] Stephen Simons. Introduction. Lecture Notes in Mathematics, 1693:1–11, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093634/.

Simons:1998:MM

[Sim98f] Stephen Simons. Minimax and Monotonicity, volume 1693 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 172 pp. LCCN QA3 .L28 no. 1693. URL http:/

/link.springer.com/book/10.1007/BFb0093633;   http://
www.springerlink.com/content/978-3-540-68931-7.

Simons:1998:M

[Sim98g] Stephen Simons. Multifunctions. Lecture Notes in Mathematics, 1693:29–41, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093636/.

Simons:1998:OP

[Sim98h] Stephen Simons. Open problems. Lecture Notes in Mathematics, 1693:163–164, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093644.

Simons:1998:SMM

[Sim98i] Stephen Simons. Special maximal monotone multifunctions. Lecture Notes in Mathematics, 1693:97–109, 1998. CODEN LN-MAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093640/.

Simons:1998:S

[Sim98j] Stephen Simons. Subdifferentials. Lecture Notes in Mathematics, 1693:111–139, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093641/.

Simons:1998:SPG

[Sim98k] Stephen Simons. The sum problem for general Banach spaces. Lecture Notes in Mathematics, 1693:153–161, 1998. CODEN LNMAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093643/.

Simons:1998:SPR

[Sim98l] Stephen Simons. The sum problem for reflexive spaces. Lecture Notes in Mathematics, 1693:75–95, 1998. CODEN LN-MAA2. ISBN 3-540-64755-4 (print), 3-540-68931-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093639/.

Sinai:1991:MPT

[Sin91] Ya. G. Sinai. Mathematical problems in the theory of quantum chaos. Lecture Notes in Mathematics, 1469:41–59, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089214/.

Singh:1992:CFT

[Sin92] U. N. Singh. The Carleman–Fourier transform and its applications. Lecture Notes in Mathematics, 1511:181–214, 1992. CO-DEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093811/.

Sipos:1992:PSF

[Šip92] Ján Šipoš. Peculiar submeasures on finite algebras. Lecture Notes in Mathematics, 1514:192–197, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097539/.

Siu:1991:GNC

[Siu91] Yum-Tong Siu. Global nondeformability of the complex projective space. Lecture Notes in Mathematics, 1468:254–280, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086197/.

Sjostrand:1991:MAP

[Sjö91] Johannes Sjöstrand. Microlocal analysis for the periodic magnetic Schrödinger equation and related questions. Lecture Notes in Mathematics, 1495:237–332, 1991. CODEN LNMAA2. ISBN 3-540-54948-X (print), 3-540-46603-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085125/.

Skorobogatov:1992:LCS

[Sko92] Alexei N. Skorobogatov. Linear codes, strata of grassmannians, and the problems of Segre. Lecture Notes in Mathematics, 1518: 210–223, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088004/.

Skvortsov:1990:SPD

[Skv90] V. A. Skvortsov. Some properties of dyadic primitives. Lecture Notes in Mathematics, 1419:167–179, 1990. CODEN LN-MAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083107/.

Slominska:1991:HCC

[Sło91] Jolanta Słominska. Homotopy colimits on E-I-categories. Lecture Notes in Mathematics, 1474:273–294, 1991. CODEN LN-MAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084752/.

Schwartz:1999:ARA

[SM99a] Niels Schwartz and James J. Madden. Arities of reflectors and approximations by H-closed reflectors. Lecture Notes in Mathematics, 1712:161–166, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093982/.

Schwartz:1999:CRa

[SM99b] Niels Schwartz and James J. Madden. Comparing reflectors. Lecture Notes in Mathematics, 1712:79–92, 1999. CODEN LN-MAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093977/.

Schwartz:1999:CRb

[SM99c] Niels Schwartz and James J. Madden. Constructing reflectors. Lecture Notes in Mathematics, 1712:93–105, 1999. CODEN LN-MAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093978/.

Schwartz:1999:EER

[SM99d] Niels Schwartz and James J. Madden. Epimorphic extensions of reduced porings. Lecture Notes in Mathematics, 1712:167– 181, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093983/.

Schwartz:1999:ERP

[SM99e] Niels Schwartz and James J. Madden. Epimorphisms of reduced porings. Lecture Notes in Mathematics, 1712:51–53, 1999. CO-DEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093974/.

Schwartz:1999:EM

[SM99f] Niels Schwartz and James J. Madden. Essential monoreflectors. Lecture Notes in Mathematics, 1712:183–188, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093984/.

Schwartz:1999:FRP

[SM99g] Niels Schwartz and James J. Madden. Functions and representable porings. Lecture Notes in Mathematics, 1712:55–61, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093975/.

Schwartz:1999:CE

[SM99h] Niels Schwartz and James J. Madden. H-closed epireflectors. Lecture Notes in Mathematics, 1712:107–123, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093979/.

Schwartz:1999:I

[SM99i] Niels Schwartz and James J. Madden. Introduction. Lecture Notes in Mathematics, 1712:1–19, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093969/.

Schwartz:1999:PPO

[SM99j] Niels Schwartz and James J. Madden. Preordered and partially ordered rings. Lecture Notes in Mathematics, 1712:21–33, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093970/.

Schwartz:1999:QCR

[SM99k] Niels Schwartz and James J. Madden. Quotient-closed reflectors. Lecture Notes in Mathematics, 1712:125–132, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093980/.

Schwartz:1999:RCR

[SM99l] Niels Schwartz and James J. Madden. The real closure reflector. Lecture Notes in Mathematics, 1712:133–160, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093981/.

Schwartz:1999:RSP

[SM99m] Niels Schwartz and James J. Madden. Real spectra of preordered rings. Lecture Notes in Mathematics, 1712:45–49, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093973/.

Schwartz:1999:RR

[SM99n] Niels Schwartz and James J. Madden. Reduced f-rings. Lecture Notes in Mathematics, 1712:217–227, 1999. CODEN LN-MAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093988/.

Schwartz:1999:RTO

[SM99o] Niels Schwartz and James J. Madden. Reflections of totally ordered fields. Lecture Notes in Mathematics, 1712:189–199, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093985/.

Schwartz:1999:RS

[SM99p] Niels Schwartz and James J. Madden. Reflective subcategories. Lecture Notes in Mathematics, 1712:35–42, 1999. CODEN LN-MAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093971/.

Schwartz:1999:SAFb

[SM99q] Niels Schwartz and James J. Madden. Semi-algebraic Function Rings and Reflectors of Partially Ordered Rings, volume 1712 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 279 pp. LCCN QA3 .L28 no. 1712. URL http:/

/link.springer.com/book/10.1007/BFb0093968;   http://
www.springerlink.com/content/978-3-540-48284-0.

Schwartz:1999:SAFa

[SM99r] Niels Schwartz and James J. Madden. Semi-algebraic functions. Lecture Notes in Mathematics, 1712:63–77, 1999. CODEN LN-MAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093976/.

Schwartz:1999:TOR

[SM99s] Niels Schwartz and James J. Madden. Totally ordered and real closed fields. Lecture Notes in Mathematics, 1712:43–44, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093972.

Schwartz:1999:TOD

[SM99t] Niels Schwartz and James J. Madden. Totally ordered domains. Lecture Notes in Mathematics, 1712:209–215, 1999. CODEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093987/.

Schwartz:1999:NRR

[SM99u] Niels Schwartz and James J. Madden. von Neumann regular frings. Lecture Notes in Mathematics, 1712:201–208, 1999. CO-DEN LNMAA2. ISBN 3-540-66460-2 (print), 3-540-48284-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093986/.

Smale:1992:TC

[Sma92] Stephen Smale. Theory of computation. Lecture Notes in Mathematics, 1525:59–69, 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089205/.

Smith:1990:CMW

[Smi90] Patrick F. Smith. Cs-modules and weak cs-modules. Lecture Notes in Mathematics, 1448:99–115, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091255/.

Smith:1999:BM

[Smi99a] H. L. Smith. Back matter. Lecture Notes in Mathematics, 1714: ??, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-48170-6/1.pdf.

Smith:1999:DC

[Smi99b] H. L. Smith. Dynamics of competition. Lecture Notes in Mathematics, 1714:191–240, 1999. CODEN LNMAA2. ISBN 3-540-66522-6 (print), 3-540-48170-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092378/.

Schinzel:1990:LZP

[SN90] A. Schinzel and J. L. Nicolas. Localisation des zéros de polynômes intervenant en théorie du signal. (French) []. Lecture Notes in Mathematics, 1415:167–179, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084886/.

Snaith:1992:AGC

[Sna92] Victor Snaith. Applications of group cohomology to bilinear forms. Lecture Notes in Mathematics, 1491:133–164, 1992. CO-DEN LNMAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088881/.

Sone:1991:ECR

[SOA91] Yoshio Sone, Taku Ohwada, and Kazuo Aoki. Evaporation and condensation of a rarefied gas between its two parallel plane condensed phases with different temperatures and negative temperature-gradient phenomenon — numerical analysis of the Boltzmann equation for hard-sphere molecules. Lecture Notes in Mathematics, 1460:186–202, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0091368/.

Soardi:1994:CPF

[Soa94a] Paolo M. Soardi. Currents and potentials with finite energy. Lecture Notes in Mathematics, 1590:32–71, 1994. CODEN LN-MAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073998/.

Soardi:1994:FN

[Soa94b] Paolo M. Soardi. Finite networks. Lecture Notes in Mathematics, 1590:22–31, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073997/.

Soardi:1994:KL

[Soa94c] Paolo M. Soardi. Kirchhoff’s laws. Lecture Notes in Mathematics, 1590:1–21, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073996/.

Soardi:1994:PTI

[Soa94d] Paolo M. Soardi. Potential Theory on Infinite Networks, volume 1590 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 187 pp. LCCN QA3 .L28 no. 1590. URL http:/

/link.springer.com/book/10.1007/BFb0073995;   http://
www.springerlink.com/content/978-3-540-48798-2.

Soardi:1994:RI

[Soa94e] Paolo M. Soardi. Rough isometries. Lecture Notes in Mathematics, 1590:160–172, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074002/.

Soardi:1994:RC

[Soa94f] Paolo M. Soardi. Royden’s compactification. Lecture Notes in Mathematics, 1590:131–159, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074001/.

Soardi:1994:SEC

[Soa94g] Paolo M. Soardi. Some examples and computations. Lecture Notes in Mathematics, 1590:100–130, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074000/.

Soardi:1994:URT

[Soa94h] Paolo M. Soardi. Uniqueness and related topics. Lecture Notes in Mathematics, 1590:72–99, 1994. CODEN LNMAA2. ISBN 3-540-58448-X (print), 3-540-48798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073999/.

Sobolev:1990:NIM

[Sob90] V. A. Sobolev. Nonlocal integral manifolds and decoupling of nonlinear parabolic systems. Lecture Notes in Mathematics, 1453: 101–108, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085949/.

Socolowsky:1992:FBP

[Soc92] Jürgen Socolowsky. On a free boundary problem for the stationary Navier–Stokes equations with a dynamic contact line. Lecture Notes in Mathematics, 1530:17–29, 1992. CODEN LN-MAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090331/.

Songling:1990:FCB

[Son90] Shi Songling. On the finiteness of certain boundary cycles for N th degree polynomial vector fields. Lecture Notes in Mathematics, 1455:369–372, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085401/.

Soner:1997:CMP

[Son97] Halil Mete Soner. Controlled Markov processes, viscosity solutions and applications to mathematical finance. Lecture Notes in Mathematics, 1660:134–185, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094297/.

Soule:1992:ATI

[Sou92] Christophe Soulé. Algebraic K-theory of the integers. Lecture Notes in Mathematics, 1491:100–132, 1992. CODEN LN-MAA2. ISBN 3-540-55007-0 (print), 3-540-46639-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088880/.

Souganidis:1997:BM

[Sou97a] Panagiotis E. Souganidis. Back matter. Lecture Notes in Mathematics, 1660:??, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-69043-6/1.pdf.

Souganidis:1997:FPT

[Sou97b] Panagiotis E. Souganidis. Front propagation: Theory and applications. Lecture Notes in Mathematics, 1660:186–242, 1997. CODEN LNMAA2. ISBN 3-540-62910-6 (print), 3-540-69043-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094298/.

Spencer:1993:NLR

[Spe93] Joel Spencer. Nine lectures on random graphs. Lecture Notes in Mathematics, 1541:293–347, 1993. CODEN LNMAA2. ISBN 3-540-56622-8 (print), 3-540-47608-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084192/.

Spiga:1991:RSE

[Spi91] Giampiero Spiga. Rigorous solution to the extended kinetic equations for homogeneous gas mixtures. Lecture Notes in Mathematics, 1460:203–221, 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091369/.

Spreen:1990:CET

[Spr90] Dieter Spreen. A characterization of effective topological spaces. Lecture Notes in Mathematics, 1432:363–387, 1990. CODEN LNMAA2. ISBN 3-540-52772-9 (print), 3-540-47142-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086126/.

Sprindzuk:1993:AF

[Spr93a] Vladimir G. Sprindžuk. Algebraic foundations. Lecture Notes in Mathematics, 1559:14–29, 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073788/.

Sprindzuk:1993:CNV

[Spr93b] Vladimir G. Sprindžuk. The class number value problem. Lecture Notes in Mathematics, 1559:155–187, 1993. CODEN LN-MAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073794/.

Sprindzuk:1993:CDE

[Spr93c] Vladimir G. Sprindžuk. Classical Diophantine Equations, volume 1559 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 228 pp. LCCN QA3 .L28 no. 1559. URL http:/

/link.springer.com/book/10.1007/BFb0073786;   http://
www.springerlink.com/content/978-3-540-48083-9.

Sprindzuk:1993:EHE

[Spr93d] Vladimir G. Sprindžuk. Elliptic and hyperelliptic equations. Lecture Notes in Mathematics, 1559:111–137, 1993. CODEN LN-MAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073792/.

Sprindzuk:1993:EHT

[Spr93e] Vladimir G. Sprindžuk. Equations of hyperelliptic type. Lecture Notes in Mathematics, 1559:138–154, 1993. CODEN LN-MAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073793/.

Sprindzuk:1993:LFL

[Spr93f] Vladimir G. Sprindžuk. Linear forms in the logarithms of algebraic numbers. Lecture Notes in Mathematics, 1559:30–60, 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073789/.

Sprindzuk:1993:O

[Spr93g] Vladimir G. Sprindžuk. Origins. Lecture Notes in Mathematics, 1559:1–13, 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073787/.

Sprindzuk:1993:RPD

[Spr93h] Vladimir G. Sprindžuk. Reducibility of polynomials and Diophantine equations. Lecture Notes in Mathematics, 1559:188– 218, 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073795/.

Sprindzuk:1993:TME

[Spr93i] Vladimir G. Sprindžuk. The Thue–Mahler equation. Lecture Notes in Mathematics, 1559:85–110, 1993. CODEN LN-MAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073791/.

Sprindzuk:1993:TE

[Spr93j] Vladimir G. Sprindžuk. The Thue equation. Lecture Notes in Mathematics, 1559:61–84, 1993. CODEN LNMAA2. ISBN 3-540-57359-3 (print), 3-540-48083-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073790/.

Srebro:1992:TPQ

[Sre92] Uri Srebro. Topological properties of quasiregular mappings. Lecture Notes in Mathematics, 1508:104–118, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094241/.

Srinivasan:1990:MRF

[Sri90] Bhama Srinivasan. Modular representations of finite groups of Lie type in a non-defining characteristic. Lecture Notes in Mathematics, 1456:96–105, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100733/.

Sternin:1990:LRT

[SS90] B. Yu. Sternin and V. E. Shatalov. On Leary’s residue theory. Lecture Notes in Mathematics, 1453:109–119, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085950/.

Schechtman:1991:ARV

[SS91] G. Schechtman and M. Schmuckenschläger. Another remark on the volume of the intersection of two L_{p}^{n} balls. Lecture Notes in Mathematics, 1469:174–178, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089224/.

Sternin:1992:CSE

[SS92] B. Yu. Sternin and V. E. Shatalov. Continuation of solutions to elliptic equations and localization of singularities. Lecture Notes in Mathematics, 1520:237–259, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084724/.

Schapira:1993:ITC

[SS93] Pierre Schapira and Jean-Pierre Schneiders. Index theorems for R-constructible sheaves and for D-modules. Lecture Notes in Mathematics, 1565:141–156, 1993. CODEN LNMAA2. ISBN 3-540-57498-0 (print), 3-540-48195-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073467/.

Schmeling:1992:HCH

[SSS92] J. Schmeling and Ra. Siegmund-Schultze. Hölder continuity of the holonomy maps for hyperbolic basic sets I. Lecture Notes in Mathematics, 1514:174–191, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097538/.

Solonnikov:1992:EFB

[ST92a] V. A. Solonnikov and A. Tani. Evolution free boundary problem for equations of motion of viscous compressible barotropic liquid. Lecture Notes in Mathematics, 1530:30–55, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090332/.

Stichtenoth:1992:AGC

[ST92b] Henning Stichtenoth and Michael A. Tsfasman. Algebraic geometry and coding theory an introduction. Lecture Notes in Mathematics, 1518:1–3, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087987/.

Stichtenoth:1992:CTA

[ST92c] Henning Stichtenoth and Michael A. Tsfasman, editors. Coding Theory and Algebraic Geometry: Proceedings of the International Workshop held in Luminy, France, June 17–21, 1991, volume 1518 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1518. URL http:/

/link.springer.com/book/10.1007/BFb0087986;   http://
www.springerlink.com/content/978-3-540-47267-4.

Stasheff:1992:DGL

[Sta92] Jim Stasheff. Differential graded Lie algebras, quasi-Hopf algebras and higher homotopy algebras. Lecture Notes in Mathematics, 1510:120–137, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0101184/.

Stahl:1993:URA

[Sta93] Herbert Stahl. Uniform rational approximation of |X|. Lecture Notes in Mathematics, 1550:110–130, 1993. CODEN LN-MAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117477/.

Stegeman:1990:APM

[Ste90a] J. D. Stegeman. Analysis of P. Malliavin’s proof of non spectral synthesis. Lecture Notes in Mathematics, 1419:180–200, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083108/.

Stegeman:1990:PGC

[Ste90b] J. D. Stegeman. Papers of G. Cross, Y. Kubota, J. L. Mawhin, M. Morayne, W. F. Pfeffer and W.-C. Yang, and C. A. Rogers. Lecture Notes in Mathematics, 1419:201, 1990. CODEN LN-MAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083109.

Stegeman:1990:P

[Ste90c] J. D. Stegeman. Problems. Lecture Notes in Mathematics, 1419: 202, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083110.

Stevens:1991:VDC

[Ste91] Jan Stevens. On the versal deformation of cyclic quotient singularities. Lecture Notes in Mathematics, 1462:302–319, 1991. CODEN LNMAA2. ISBN 3-540-53737-6 (print), 3-540-47060-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086390/.

Steffen:1999:BM

[Ste99a] Klaus Steffen. Back matter. Lecture Notes in Mathematics, 1713: ??, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-48813-2/1.pdf.

Steffen:1999:PSP

[Ste99b] Klaus Steffen. Parametric surfaces of prescribed mean curvature. Lecture Notes in Mathematics, 1713:211–265, 1999. CODEN LNMAA2. ISBN 3-540-65977-3 (print), 3-540-48813-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092671/.

Stoer:1992:BI

[Sto92a] Mechthild Stoer. Basic inequalities. Lecture Notes in Mathematics, 1531:49–68, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088968/.

Stoer:1992:CI

[Sto92b] Mechthild Stoer. Comb inequalities. Lecture Notes in Mathematics, 1531:113–123, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088973/.

Stoer:1992:CR

[Sto92c] Mechthild Stoer. Computational results. Lecture Notes in Mathematics, 1531:175–194, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088976/.

Stoer:1992:D

[Sto92d] Mechthild Stoer. Decomposition. Lecture Notes in Mathematics, 1531:33–47, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088967/.

Stoer:1992:DSN

[Sto92e] Mechthild Stoer. Design of Survivable Networks, volume 1531 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 205 pp. LCCN QA3 .L28 no. 1531. URL http:/

/link.springer.com/book/10.1007/BFb0088963;   http://
www.springerlink.com/content/978-3-540-47500-2.

Stoer:1992:HFV

[Sto92f] Mechthild Stoer. How to find valid inequalities. Lecture Notes in Mathematics, 1531:125–154, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088974/.

Stoer:1992:ICP

[Sto92g] Mechthild Stoer. Implementation of the cutting plane algorithm. Lecture Notes in Mathematics, 1531:155–173, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088975/.

Stoer:1992:LCI

[Sto92h] Mechthild Stoer. Lifted r-cover inequalities. Lecture Notes in Mathematics, 1531:101–112, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088972/.

Stoer:1992:LT

[Sto92i] Mechthild Stoer. Lifting theorems. Lecture Notes in Mathematics, 1531:69–76, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088969/.

Stoer:1992:M

[Sto92j] Mechthild Stoer. Motivation. Lecture Notes in Mathematics, 1531: 5–6, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088964.

Stoer:1992:NSM

[Sto92k] Mechthild Stoer. Network survivability models using node types. Lecture Notes in Mathematics, 1531:7–18, 1992. CODEN LN-MAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088965/.

Stoer:1992:NPI

[Sto92l] Mechthild Stoer. Node partition inequalities. Lecture Notes in Mathematics, 1531:91–99, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088971/.

Stoer:1992:PI

[Sto92m] Mechthild Stoer. Partition inequalities. Lecture Notes in Mathematics, 1531:77–90, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088970/.

Stoer:1992:SND

[Sto92n] Mechthild Stoer. Survivable network design under connectivity constraints — a survey. Lecture Notes in Mathematics, 1531: 19–32, 1992. CODEN LNMAA2. ISBN 3-540-56271-0 (print), 3-540-47500-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0088966/.

Stokes:1992:DCC

[Sto92o] Philip Stokes. The domain of covering codes. Lecture Notes in Mathematics, 1518:170–177, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088000/.

Stroock:1993:BM

[Str93a] Daniel W. Stroock. Back matter. Lecture Notes in Mathematics, 1563:??, 1993. CODEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-48151-5/1.pdf.

Stroock:1993:LSI

[Str93b] Daniel W. Stroock. Logarithmic Sobolev inequalities for Gibbs states. Lecture Notes in Mathematics, 1563:194–228, 1993. CO-DEN LNMAA2. ISBN 3-540-57421-2 (print), 3-540-48151-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074094/.

Shparlinski:1992:CMP

[STV92] Igor E. Shparlinski, Michael A. Tsfasman, and Serge G. Vladut. Curves with many points and multiplication in finite fields. Lecture Notes in Mathematics, 1518:145–169, 1992. CODEN LN-MAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0087999/.

Suffridge:1990:EPN

[Suf90] T. J. Suffridge. Extremal problems for non-vanishing H^{p} functions. Lecture Notes in Mathematics, 1435:177–190, 1990. CO-DEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087907/.

Suzuki:1990:IGP

[Suz90] Takashi Suzuki. Introduction to geometric potential theory. Lecture Notes in Mathematics, 1450:88–103, 1990. CODEN LN-MAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084900/.

Sohr:1990:DPS

[SV90] Hermann Sohr and Werner Varnhorn. On decay properties of the Stokes equations in exterior domains. Lecture Notes in Mathematics, 1431:134–151, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086064/.

Schwanzl:1991:HRS

[SV91a] R. Schwänzl and R. M. Vogt. Homotopy ring spaces and their matrix rings. Lecture Notes in Mathematics, 1474:254–272, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084751/.

Smith:1991:STA

[SV91b] Roy Smith and Robert Varley. Singularity theory applied to Θ-divisors. Lecture Notes in Mathematics, 1479:238–257, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086273/.

Svertchkov:1993:WSR

[Sve93] M. Yu. Svertchkov. On wide-sense regeneration. Lecture Notes in Mathematics, 1546:167–169, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084492/.

Schmitt:1992:DSF

[SvW92] Burkhard J. Schmitt and Wolf von Wahl. Decomposition of solenoidal fields into poloidal fields, toroidal fields and the mean flow. Applications to the Boussinesq-equations. Lecture Notes in Mathematics, 1530:291–305, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090349/.

Suli:1992:ASL

[SW92] Endre Süli and Antony F. Ware. Analysis of the spectral Lagrange–Galerkin method for the Navier–Stokes equations. Lecture Notes in Mathematics, 1530:184–195, 1992. CODEN LN-MAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090342/.

Schneider:1993:RTB

[SW93] Dominique Schneider and Michel Weber. Une remarque sur un théorème de bourgain. (French) []. Lecture Notes in Mathematics, 1557:202–206, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087976/.

Seeley:1991:AMS

[SY91] Craig Seeley and Stephen S.-T. Yau. Algebraic methods in the study of simple-elliptic singularities. Lecture Notes in Mathematics, 1479:216–237, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086272/.

Stricker:1998:SRO

[SY98] C. Stricker and J. A. Yan. Some remarks on the optional decomposition theorem. Lecture Notes in Mathematics, 1686:56–66, 1998. CODEN LNMAA2. ISBN 3-540-64376-1 (print), 3-540-69762-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101750/.

Schaufele:1990:GGB

[SZ90] Christopher Schaufele and Nancy Zumoff. ∗-groups, graphs, and bases. Lecture Notes in Mathematics, 1440:186–191, 1990. CO-DEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084462/.

Szarek:1991:GBM

[Sza91] S. J. Szarek. On the geometry of the Banach–Mazur compactum. Lecture Notes in Mathematics, 1470:48–59, 1991. CODEN LN-MAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090211/.

Sznitman:1991:TPC

[Szn91] Alain-Sol Sznitman. Topics in propagation of chaos. Lecture Notes in Mathematics, 1464:165–251, 1991. CODEN LNMAA2. ISBN 3-540-53841-0 (print), 3-540-46319-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085169/.

Tadmor:1998:ASN

[Tad98] Eitan Tadmor. Approximate solutions of nonlinear conservation laws. Lecture Notes in Mathematics, 1697:1–149, 1998. CODEN LNMAA2. ISBN 3-540-64977-8 (print), 3-540-49804-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096352/.

Taft:1992:QDF

[Taf92] Earl J. Taft. Quantum deformation of the flag variety. Lecture Notes in Mathematics, 1510:138–141, 1992. CODEN LN-MAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101185/.

Taira:1991:AEB

[Tai91a] Kazuaki Taira. L^{p} approach to elliptic boundary value problems. Lecture Notes in Mathematics, 1499:41–49, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092033/.

Taira:1991:TPD

[Tai91b] Kazuaki Taira. L^{p} theory of pseudo-differential operators. Lecture Notes in Mathematics, 1499:23–40, 1991. CODEN LN-MAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092032/.

Taira:1991:ASI

[Tai91c] Kazuaki Taira. Application to semilinear initial-boundary value problems. Lecture Notes in Mathematics, 1499:105–111, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092039/.

Taira:1991:BVP

[Tai91d] Kazuaki Taira. Boundary Value Problems and Markov Processes, volume 1499 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 132 pp. LCCN QA3 .L28 no. 1499. URL http:/

/link.springer.com/book/10.1007/BFb0092029;   http://
www.springerlink.com/content/978-3-540-46635-2.

Taira:1991:IR

[Tai91e] Kazuaki Taira. Introduction and results. Lecture Notes in Mathematics, 1499:1–9, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092030/.

Taira:1991:PE

[Tai91f] Kazuaki Taira. A priori estimates. Lecture Notes in Mathematics, 1499:55–60, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092035/.

Taira:1991:PTa

[Tai91g] Kazuaki Taira. Proof of theorem 1. Lecture Notes in Mathematics, 1499:50–54, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092034/.

Taira:1991:PTb

[Tai91h] Kazuaki Taira. Proof of theorem 2. Lecture Notes in Mathematics, 1499:61–69, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092036/.

Taira:1991:PTPa

[Tai91i] Kazuaki Taira. Proof of Theorem 3 — Part (I). Lecture Notes in Mathematics, 1499:70–80, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092037/.

Taira:1991:PTPb

[Tai91j] Kazuaki Taira. Proof of Theorem 3 — Part (II). Lecture Notes in Mathematics, 1499:81–104, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092038/.

Taira:1991:ST

[Tai91k] Kazuaki Taira. Semigroup theory. Lecture Notes in Mathematics, 1499:10–22, 1991. CODEN LNMAA2. ISBN 3-540-54996-X (print), 3-540-46635-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092031/.

Takegoshi:1991:ACI

[Tak91] Kensho Takegoshi. Application of a certain integral formula to complex analysis. Lecture Notes in Mathematics, 1468:94– 114, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086189/.

Takaoka:1999:SRU

[Tak99] Koichiro Takaoka. Some remarks on the uniform integrability of continuous martingales. Lecture Notes in Mathematics, 1709: 327–333, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096521/.

Talagrand:1991:NII

[Tal91] Michel Talagrand. A new isoperimetric inequality and the concentration of measure phenomenon. Lecture Notes in Mathematics, 1469:94–124, 1991. CODEN LNMAA2. ISBN 3-540-54024-5 (print), 3-540-47355-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089217/.

Talay:1996:PNM

[Tal96] Denis Talay. Probabilistic numerical methods for partial differential equations: Elements of analysis. Lecture Notes in Mathematics, 1627:148–196, 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093180/.

Tamura:1990:EBS

[Tam90] Hideo Tamura. Existence of bound states for double well potentials and the Efimov effect. Lecture Notes in Mathematics, 1450:173–186, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084905/.

Tamanoi:1999:EG

[Tam99a] Hirotaka Tamanoi. Elliptic genera. Lecture Notes in Mathematics, 1704:22–48, 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092543/.

Tamanoi:1999:EGV

[Tam99b] Hirotaka Tamanoi. Elliptic Genera and Vertex Operator Super-Algebras, volume 1704 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 390 pp. LCCN QA3 .L28 no. 1704. URL http:/

/link.springer.com/book/10.1007/BFb0092541;   http://
www.springerlink.com/content/978-3-540-48788-3.

Tamanoi:1999:IVO

[Tam99c] Hirotaka Tamanoi. G-invariant vertex operator super subalgebras. Lecture Notes in Mathematics, 1704:93–221, 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092545/.

Tamanoi:1999:GSV

[Tam99d] Hirotaka Tamanoi. Geometric structure in vector spaces and reduction of structure groups on manifolds. Lecture Notes in Mathematics, 1704:222–302, 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092546/.

Tamanoi:1999:IDS

[Tam99e] Hirotaka Tamanoi. Infinite dimensional symmetries in elliptic genera for Kähler manifolds. Lecture Notes in Mathematics, 1704: 303–378, 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092547/.

Tamanoi:1999:ISR

[Tam99f] Hirotaka Tamanoi. Introduction and summary of results. Lecture Notes in Mathematics, 1704:1–21, 1999. CODEN LN-MAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092542/.

Tamanoi:1999:VOS

[Tam99g] Hirotaka Tamanoi. Vertex operator super algebras. Lecture Notes in Mathematics, 1704:49–92, 1999. CODEN LNMAA2. ISBN 3-540-66006-2 (print), 3-540-48788-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092544/.

Taniguchi:1991:BEL

[Tan91] Masahiko Taniguchi. The behavior of the extremal length function on arbitrary Riemann surface. Lecture Notes in Mathematics, 1468:160–169, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086192/.

Tate:1996:AGC

[Tat96] John Tate. Applications of Galois cohomology in algebraic geometry. Lecture Notes in Mathematics, 1625:188–215, 1996. CODEN LNMAA2. ISBN 3-540-61181-9 (print), 3-540-68337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092634/.

Taylor:1992:SPR

[Tay92] J. C. Taylor. Skew products, regular conditional probabilities and stochastic differential equations: a technical remark. Lecture Notes in Mathematics, 1526:113–126, 1992. CODEN LN-MAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084315/.

Toscani:1991:MAF

[TBR91] Giuseppe Toscani, Vinicio Boffi, and Salvatore Rionero, editors. Mathematical Aspects of Fluid and Plasma Dynamics: Proceedings of an International Workshop held in Salice Terme, Italy, 26–30 September 1988, volume 1460 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-53545-4 (print), 3-540-46779-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1460. URL http:/

/link.springer.com/book/10.1007/BFb0091356;   http://
www.springerlink.com/content/978-3-540-46779-3.

Temlyakov:1993:UFC

[Tem93] V. N. Temlyakov. Universality of the Fibonacci cubature formulas. Lecture Notes in Mathematics, 1550:178–184, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117486/.

Thakur:1990:GKF

[Tha90] Dinesh S. Thakur. Gross–Koblitz formula for function fields. Lecture Notes in Mathematics, 1454:351–355, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091149/.

Thieullen:1991:GDT

[Thi91] P. Thieullen. Généralisation du théorème de Pesin pour l’αentropie. (French) []. Lecture Notes in Mathematics, 1486:232– 242, 1991. CODEN LNMAA2. ISBN 3-540-54662-6 (print), 3-540-46431-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086673/.

Thompson:1990:GGZ

[Tho90a] John G. Thompson. Groups of genus zero and certain rational functions. Lecture Notes in Mathematics, 1456:185–190, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100740/.

Thorup:1990:RET

[Tho90b] Anders Thorup. Rational equivalence theory on arbitrary Noetherian schemes. Lecture Notes in Mathematics, 1436:256–297, 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084049/.

Thom:1992:LMP

[Tho92] René Thom. Leaving mathematics for philosophy. Lecture Notes in Mathematics, 1525:1–12, 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089201/.

Thorisson:1993:RSS

[Tho93] Hermann Thorisson. Regeneration, stationarity and simulation. Lecture Notes in Mathematics, 1546:174–179, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084494/.

Thron:1990:SRS

[Thr90] W. J. Thron. Some results on separate convergence of continued fractions. Lecture Notes in Mathematics, 1435:191–200, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087908/.

Tian:1996:KEM

[Tia96] Gang Tian. Kähler–Einstein metrics on algebraic manifolds. Lecture Notes in Mathematics, 1646:143–185, 1996. CODEN LN-MAA2. ISBN 3-540-62038-9 (print), 3-540-49632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094304/.

Tiba:1990:ENA

[Tib90a] Dan Tiba. Elements of nonlinear analysis. Lecture Notes in Mathematics, 1459:1–28, 1990. CODEN LNMAA2. ISBN 3-540-53524-1 (print), 3-540-46755-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085565/.

Tiba:1990:FBP

[Tib90b] Dan Tiba. Free boundary problems. Lecture Notes in Mathematics, 1459:122–151, 1990. CODEN LNMAA2. ISBN 3-540-53524-1 (print), 3-540-46755-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085568/.

Tiba:1990:OCN

[Tib90c] Dan Tiba. Optimal Control of Nonsmooth Distributed Parameter Systems, volume 1459 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-53524-1 (print), 3-540-46755-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 159 pp. LCCN QA3 .L28 no. 1459. URL http:/

/link.springer.com/book/10.1007/BFb0085564;   http://
www.springerlink.com/content/978-3-540-46755-7.

Tiba:1990:SE

[Tib90d] Dan Tiba. Semilinear equations. Lecture Notes in Mathematics, 1459:29–76, 1990. CODEN LNMAA2. ISBN 3-540-53524-1 (print), 3-540-46755-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085566/.

Tiba:1990:VI

[Tib90e] Dan Tiba. Variational inequalities. Lecture Notes in Mathematics, 1459:77–121, 1990. CODEN LNMAA2. ISBN 3-540-53524-1 (print), 3-540-46755-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085567/.

Tijdeman:1993:DAA

[Tij93a] Rob Tijdeman. Diophantine approximation and its applications. Lecture Notes in Mathematics, 1566:13–20, 1993. CODEN LN-MAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48208-6_
2/.

Tijdeman:1993:RT

[Tij93b] Rob Tijdeman. Roth’s theorem. Lecture Notes in Mathematics, 1566:21–30, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48208-6_3/.

Tikhomirov:1990:KPM

[Tik90] V. M. Tikhomirov. A. N. Kolmogorov and the progress of mathematics. Lecture Notes in Mathematics, 1453:161–169, 1990. CO-DEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085953/.

Tralle:1997:EM

[TO97a] Aleksy Tralle and John Oprea. The examples of McDuff. Lecture Notes in Mathematics, 1661:120–136, 1997. CODEN LN-MAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092612/.

Tralle:1997:N

[TO97b] Aleksy Tralle and John Oprea. Nilmanifolds. Lecture Notes in Mathematics, 1661:45–69, 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092610/.

Tralle:1997:Sa

[TO97c] Aleksy Tralle and John Oprea. Solvmanifolds. Lecture Notes in Mathematics, 1661:70–119, 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092611/.

Tralle:1997:SPH

[TO97d] Aleksy Tralle and John Oprea. The starting point: Homotopy properties of Kähler manifolds. Lecture Notes in Mathematics, 1661:1–44, 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092609/.

Tralle:1997:Sb

[TO97e] Aleksy Tralle and John Oprea. Survey. Lecture Notes in Mathematics, 1661:173–199, 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092614/.

Tralle:1997:SMN

[TO97f] Aleksy Tralle and John Oprea. Symplectic Manifolds with no Kähler Structure, volume 1661 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 207 pp. LCCN QA3 .L28 no. 1661.

URL http://link.springer.com/book/10.1007/BFb0092608;
 http://www.springerlink.com/content/978-3-540-69145-
7.

Tralle:1997:SST

[TO97g] Aleksy Tralle and John Oprea. Symplectic structures in total spaces of bundles. Lecture Notes in Mathematics, 1661:137– 172, 1997. CODEN LNMAA2. ISBN 3-540-63105-4 (print), 3-540-69145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092613/.

Todorcevic:1997:CCG

[Tod97a] Stevo Todorcevic. Compact and compactly generated groups. Lecture Notes in Mathematics, 1652:79–120, 1997. CODEN LN-MAA2. ISBN 3-540-62611-5 (print), 3-540-68338-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096298/.

Todorcevic:1997:CSF

[Tod97b] Stevo Todorcevic. Compact sets in function spaces. Lecture Notes in Mathematics, 1652:1–60, 1997. CODEN LNMAA2. ISBN 3-540-62611-5 (print), 3-540-68338-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096296/.

Todorcevic:1997:H

[Tod97c] Stevo Todorcevic. Hyperspaces. Lecture Notes in Mathematics, 1652:121–147, 1997. CODEN LNMAA2. ISBN 3-540-62611-5 (print), 3-540-68338-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096299/.

Todorcevic:1997:S

[Tod97d] Stevo Todorcevic. The semigroup βN. Lecture Notes in Mathematics, 1652:61–78, 1997. CODEN LNMAA2. ISBN 3-540-62611-5 (print), 3-540-68338-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096297/.

Todorcevic:1997:TT

[Tod97e] Stevo Todorcevic. Topics in Topology, volume 1652 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62611-5 (print), 3-540-68338-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 153 pp. LCCN QA3 .L28 no. 1652. URL http:/

/link.springer.com/book/10.1007/BFb0096295;   http://
www.springerlink.com/content/978-3-540-68338-4.

Top:1993:DMR

[Top93] Jaap Top. D. Mumford’s “A Remark on Mordell’s Conjecture”. Lecture Notes in Mathematics, 1566:63–67, 1993. CODEN LN-MAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48208-6_
6/.

Totik:1994:A

[Tot94a] Vilmos Totik. Applications. Lecture Notes in Mathematics, 1569: 79–110, 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076138/.

Totik:1994:AGW

[Tot94b] Vilmos Totik. Approximation with general weights. Lecture Notes in Mathematics, 1569:21–48, 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0076136/.

Totik:1994:FW

[Tot94c] Vilmos Totik. Freud weights. Lecture Notes in Mathematics, 1569: 7–20, 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076135/.

Totik:1994:I

[Tot94d] Vilmos Totik. Introduction. Lecture Notes in Mathematics, 1569: 1–5, 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0076134/.

Totik:1994:VW

[Tot94e] Vilmos Totik. Varying weights. Lecture Notes in Mathematics, 1569:49–77, 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0076137/.

Totik:1994:WAV

[Tot94f] Vilmos Totik. Weighted Approximation with Varying Weight, volume 1569 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57705-X (print), 3-540-48323-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 115 pp. LCCN QA3 .L28 no. 1569. URL http:/

/link.springer.com/book/10.1007/BFb0076133;   http://
www.springerlink.com/content/978-3-540-48323-6.

Tian-Ping:1991:OIP

[TPDZ91] Chen Tian-Ping and Zhang De-Zhi. Oscillatory integral with polynomial phase. Lecture Notes in Mathematics, 1494:39–45, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087754/.

Tarabusi:1998:IGR

[TPZ98] Enrico Casadio Tarabusi, Massimo A. Picardello, and Giuseppe Zampieri, editors. Integral Geometry, Radon Transforms and Complex Analysis: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Venice, Italy, June 3–12, 1996, volume 1684 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1684. URL http:/

/link.springer.com/book/10.1007/BFb0096088;   http://
www.springerlink.com/content/978-3-540-69702-2.

Troitsky:1991:LNC

[Tro91] E. V. Troitsky. Lefschetz numbers of C∗-complexes. Lecture Notes in Mathematics, 1474:193–206, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084747/.

Tsfasman:1992:SRA

[Tsf92] Michael A. Tsfasman. Some remarks on the asymptotic number of points. Lecture Notes in Mathematics, 1518:178–192, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088001/.

Teply:1990:PIN

[TT90] Mark L. Teply and Blas Torrecillas. Primitive ideals of nice Ore localizations. Lecture Notes in Mathematics, 1448:63–71, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091251/.

Talay:1996:PMN

[TT96] Denis Talay and Luciano Tubaro, editors. Probabilistic Models for Nonlinear Partial Differential Equations: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, May 22–30, 1995, volume 1627 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61397-8 (print), 3-540-68513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1627. URL http:/

/link.springer.com/book/10.1007/BFb0093175;   http://
www.springerlink.com/content/978-3-540-68513-5.

Tucker:1990:STG

[Tuc90] Thomas W. Tucker. Some topological graph theory for topologists: a sampler of covering space constructions. Lecture Notes in Mathematics, 1440:192–207, 1990. CODEN LNMAA2. ISBN 3-540-52990-X (print), 3-540-46296-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084463/.

Tumanov:1998:ADE

[Tum98] Alexander Tumanov. Analytic discs and the extendibility of CR functions. Lecture Notes in Mathematics, 1684:123–141, 1998. CODEN LNMAA2. ISBN 3-540-64207-2 (print), 3-540-69702-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096093/.

Turiel:1990:DMO

[Tur90a] F. J. Turiel. Dimension minimale des orbites d’une action symplectique de R^{n}. (French) []. Lecture Notes in Mathematics, 1416: 268–289, 1990. CODEN LNMAA2. ISBN 3-540-52191-7 (print), 3-540-46920-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0097477/.

Turnwald:1990:NRN

[Tur90b] Gerhard Turnwald. A note on the Ramanujan–Nagell equation. Lecture Notes in Mathematics, 1452:206–207, 1990. CODEN LNMAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096992.

Tezuka:1990:CFG

[TY90] M. Tezuka and N. Yagita. Cohomology of finite groups and Brown–Peterson cohomology II. Lecture Notes in Mathematics, 1418:57–69, 1990. CODEN LNMAA2. ISBN 3-540-52246-8 (print), 3-540-46938-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0083692/.

Tyurin:1991:SGM

[Tyu91] A. N. Tyurin. A slight generalization of the Mehta–Ramanathan theorem. Lecture Notes in Mathematics, 1479:258–272, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086274/.

Tyurin:1992:GSC

[Tyu92] A. N. Tyurin. The geometry of the special components of moduli space of vector bundles over algebraic surfaces of general type. Lecture Notes in Mathematics, 1507:166–175, 1992. CODEN LNMAA2. ISBN 3-540-55235-9 (print), 3-540-46786-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094518/.

Tyurin:1996:SLF

[Tyu96] Andrei Tyurin. Six lectures on four manifolds. Lecture Notes in Mathematics, 1646:186–246, 1996. CODEN LNMAA2. ISBN 3-540-62038-9 (print), 3-540-49632-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094305/.

Urabe:1993:CCR

[Ura93a] Tohsuke Urabe. Concept of co-root modules. Lecture Notes in Mathematics, 1548:185–226, 1993. CODEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084374/.

Urabe:1993:DGQ

[Ura93b] Tohsuke Urabe. Dynkin Graphs and Quadrilateral Singularities, volume 1548 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 233 pp. LCCN QA3 .L28 no. 1548. URL http:/

/link.springer.com/book/10.1007/BFb0084369;   http://
www.springerlink.com/content/978-3-540-47769-3.

Urabe:1993:I

[Ura93c] Tohsuke Urabe. Introduction. Lecture Notes in Mathematics, 1548:1–16, 1993. CODEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084370/.

Urabe:1993:OC

[Ura93d] Tohsuke Urabe. Obstruction components. Lecture Notes in Mathematics, 1548:98–184, 1993. CODEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084373/.

Urabe:1993:QSE

[Ura93e] Tohsuke Urabe. Quadrilateral singularities and elliptic K3 surfaces. Lecture Notes in Mathematics, 1548:17–59, 1993. CODEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084371/.

Urabe:1993:TIC

[Ura93f] Tohsuke Urabe. Theorems with the Ik-conditions for J_{3,0}, Z_{1,0} and Q_{2,0}. Lecture Notes in Mathematics, 1548:60–97, 1993. CO-DEN LNMAA2. ISBN 3-540-56877-8 (print), 3-540-47769-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084372/.

Uridia:1990:TEC

[Uri90] M. Uridia. U-theory of exact categories. Lecture Notes in Mathematics, 1437:303–313, 1990. CODEN LNMAA2. ISBN 3-540-52836-9 (print), 3-540-47162-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0086725/.

Ustunel:1995:MTM

[Üst95a] Ali Süleyman Üstünel. L^{p}-multipliers theorem, Meyer inequalities and distributions. Lecture Notes in Mathematics, 1610:31– 39, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096333/.

Ustunel:1995:CIS

[Üst95b] Ali Süleyman Üstünel. Characterization of independence of some Wiener functionals. Lecture Notes in Mathematics, 1610:61– 67, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096336/.

Ustunel:1995:GSD

[Üst95c] Ali Süleyman Üstünel. Gross–Sobolev derivative, divergence and Ornstein–Uhlenbeck operator. Lecture Notes in Mathematics, 1610:9–18, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096330/.

Ustunel:1995:H

[Üst95d] Ali Süleyman Üstünel. Hypercontractivity. Lecture Notes in Mathematics, 1610:27–30, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096332/.

Ustunel:1995:IAW

[Üst95e] Ali Süleyman Üstünel. An Introduction to Analysis on Wiener Space, volume 1610 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 95 pp. LCCN QA3 .L28 no. 1610. URL http:/

/link.springer.com/book/10.1007/BFb0096328;   http://
www.springerlink.com/content/978-3-540-44662-0.

Ustunel:1995:ITR

[Üst95f] Ali Süleyman Üstünel. Introduction to the theorem of Ramer. Lecture Notes in Mathematics, 1610:81–90, 1995. CODEN LN-MAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096338/.

Ustunel:1995:MI

[Üst95g] Ali Süleyman Üstünel. Meyer inequalities. Lecture Notes in Mathematics, 1610:19–25, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096331/.

Ustunel:1995:MIW

[Üst95h] Ali Süleyman Üstünel. Moment inequalities for Wiener functional. Lecture Notes in Mathematics, 1610:69–79, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096337/.

Ustunel:1995:PDA

[Üst95i] Ali Süleyman Üstünel. Positive distributions and applications. Lecture Notes in Mathematics, 1610:53–60, 1995. CODEN LN-MAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096335/.

Ustunel:1995:P

[Üst95j] Ali Süleyman Üstünel. Preliminaries. Lecture Notes in Mathematics, 1610:1–7, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096329/.

Ustunel:1995:SAD

[Üst95k] Ali Süleyman Üstünel. Some applications of the distributions. Lecture Notes in Mathematics, 1610:41–51, 1995. CODEN LN-MAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096334/.

Ueno:1992:ZSF

[UT92] Kimio Ueno and Tadayoshi Takebayashi. Zonal spherical functions on quantum symmetric spaces and MacDonald’s symmetric polynomials. Lecture Notes in Mathematics, 1510:142–147, 1992. CODEN LNMAA2. ISBN 3-540-55305-3 (print), 3-540-47020-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0101186/.

Ustunel:1997:CVF

[ÜZ97] A. S. Üstünel and M. Zakai. The change of variables formula on Wiener space. Lecture Notes in Mathematics, 1655:24–39, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0119289/.

Vaisala:1992:DM

[Väi92] Jussi Väisälä. Domains and maps. Lecture Notes in Mathematics, 1508:119–131, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094242/.

Vainikko:1993:CCA

[Vai93a] Gennadi Vainikko. Composite cubature algorithms. Lecture Notes in Mathematics, 1549:94–111, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088985/.

Vainikko:1993:HOM

[Vai93b] Gennadi Vainikko. Higher order methods. Lecture Notes in Mathematics, 1549:112–136, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088986/.

Vainikko:1993:MWS

[Vai93c] Gennadi Vainikko. Multidimensional Weakly Singular Integral Equations, volume 1549 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 159 pp. LCCN QA3 .L28 no. 1549.

URL http://link.springer.com/book/10.1007/BFb0088979;
 http://www.springerlink.com/content/978-3-540-47773-
0.

Vainikko:1993:NIE

[Vai93d] Gennadi Vainikko. Nonlinear integral equation. Lecture Notes in Mathematics, 1549:137–144, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088987/.

Vainikko:1993:ODC

[Vai93e] Gennadi Vainikko. Outlines of the discrete convergence theory. Lecture Notes in Mathematics, 1549:51–59, 1993. CODEN LN-MAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0088983/.

Vainikko:1993:PCC

[Vai93f] Gennadi Vainikko. Piecewise constant collocation and related methods. Lecture Notes in Mathematics, 1549:60–93, 1993. CO-DEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0088984/.

Vainikko:1993:P

[Vai93g] Gennadi Vainikko. Preliminaries. Lecture Notes in Mathematics, 1549:10–23, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088981/.

Vainikko:1993:SS

[Vai93h] Gennadi Vainikko. Smoothness of the solution. Lecture Notes in Mathematics, 1549:24–50, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088982/.

Vainikko:1993:SPL

[Vai93i] Gennadi Vainikko. Some problems leading to multidimensional weakly singular integral equations. Lecture Notes in Mathematics, 1549:1–9, 1993. CODEN LNMAA2. ISBN 3-540-56878-6 (print), 3-540-47773-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088980/.

Valadier:1990:BM

[Val90a] Michel Valadier. Back matter. Lecture Notes in Mathematics, 1446:??, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46715-1/1.pdf.

Valadier:1990:YM

[Val90b] Michel Valadier. Young measures. Lecture Notes in Mathematics, 1446:152–188, 1990. CODEN LNMAA2. ISBN 3-540-53120-3 (print), 3-540-46715-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084935/.

Valli:1990:ODN

[Val90c] Alberto Valli. On the one-dimensional Navier–Stokes equations for compressible fluids. Lecture Notes in Mathematics, 1431:173– 179, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086068/.

Vallois:1994:OUI

[Val94] Pierre Vallois. Orthogonalité et uniforme intégrabilité de martingales. Étude d’une classe d’exemples. (French) []. Lecture Notes in Mathematics, 1583:73–91, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073834/.

vanderVorst:1990:CBP

[van90] H. A. van der Vorst. The convergence behaviour of preconditioned CG and CG-S in the presence of rounding errors. Lecture Notes in Mathematics, 1457:126–136, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090905/.

VandenBergh:1991:DOS

[Van91] Michel Van den Bergh. Differential operators on semi-invariants for tori and weighted projective spaces. Lecture Notes in Mathematics, 1478:255–272, 1991. CODEN LNMAA2. ISBN 3-540-54377-5 (print), 3-540-47592-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0083507/.

Vanhaecke:1996:ACI

[Van96a] Pol Vanhaecke. Algebraic completely integrable Hamiltonian systems. Lecture Notes in Mathematics, 1638:123–138, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21535-7_5/.

Vanhaecke:1996:BM

[Van96b] Pol Vanhaecke. Back matter. Lecture Notes in Mathematics, 1638:209–221, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21535-7/1.pdf.

Vanhaecke:1996:FM

[Van96c] Pol Vanhaecke. Front matter. Lecture Notes in Mathematics, 1638:N2–viii, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-21535-7/1.pdf.

Vanhaecke:1996:GHH

[Van96d] Pol Vanhaecke. The Garnier and Hénon–Heiles potentials and the Toda lattice. Lecture Notes in Mathematics, 1638:171– 208, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21535-7_7/.

Vanhaecke:1996:IHSb

[Van96e] Pol Vanhaecke. Integrable Hamiltonian systems and symmetric products of curves. Lecture Notes in Mathematics, 1638:67– 93, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21535-7_3/.

Vanhaecke:1996:IHSa

[Van96f] Pol Vanhaecke. Integrable Hamiltonian systems on affine Poisson varieties. Lecture Notes in Mathematics, 1638:17–65, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21535-7_2/.

Vanhaecke:1996:ISR

[Van96g] Pol Vanhaecke. Integrable Systems in the realm of Algebraic Geometry, volume 1638 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 218 pp. LCCN QA3 .L28 no. 1638. URL http:

//link.springer.com/book/10.1007/978-3-662-21535-7;
http://www.springerlink.com/content/978-3-662-21535-7.

Vanhaecke:1996:IGA

[Van96h] Pol Vanhaecke. Interludium: the geometry of Abelian varieties. Lecture Notes in Mathematics, 1638:95–122, 1996. CODEN LN-MAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21535-7_
4/.

Vanhaecke:1996:I

[Van96i] Pol Vanhaecke. Introduction. Lecture Notes in Mathematics, 1638:1–15, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21535-7_1/.

Vanhaecke:1996:MS

[Van96j] Pol Vanhaecke. The master systems. Lecture Notes in Mathematics, 1638:139–169, 1996. CODEN LNMAA2. ISBN 3-540-61886-4 (print), 3-662-21535-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21535-7_6/.

Varnhorn:1992:FSM

[Var92] Werner Varnhorn. A fractional step method for regularized Navier–Stokes equations. Lecture Notes in Mathematics, 1530: 196–209, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0090343/.

Varadhan:1993:BM

[Var93a] S. R. S. Varadhan. Back matter. Lecture Notes in Mathematics, 1551:??, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47832-4/1.pdf.

Varadhan:1993:EMH

[Var93b] S. R. S. Varadhan. Entropy methods in hydrodynamic scaling. Lecture Notes in Mathematics, 1551:112–145, 1993. CODEN LNMAA2. ISBN 3-540-56945-6 (print), 3-540-47832-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090931/.

Vasconcelos:1990:SA

[Vas90] Wolmer V. Vasconcelos. Symmetric algebras. Lecture Notes in Mathematics, 1430:115–160, 1990. CODEN LNMAA2. ISBN 3-540-52745-1 (print), 3-540-47136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085540/.

Vasudevan:1992:RMM

[Vas92] R. Vasudevan. A representation of the multipler module hom_{A}(A, W). Lecture Notes in Mathematics, 1511:141–146, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093807/.

Vath:1997:BDP

[Vät97a] Martin Väth. Basic definitions and properties. Lecture Notes in Mathematics, 1664:7–27, 1997. CODEN LNMAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093550/.

Vath:1997:IS

[Vät97b] Martin Väth. Ideal Spaces, volume 1664 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 146 pp. LCCN QA3 .L28 no. 1664.

URL http://link.springer.com/book/10.1007/BFb0093548;
 http://www.springerlink.com/content/978-3-540-69192-
1.

Vath:1997:ISP

[Vät97c] Martin Väth. Ideal spaces on product measures and calculus. Lecture Notes in Mathematics, 1664:75–104, 1997. CODEN LN-MAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093552/.

Vath:1997:ISA

[Vät97d] Martin Väth. Ideal spaces with additional properties. Lecture Notes in Mathematics, 1664:29–74, 1997. CODEN LN-MAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093551/.

Vath:1997:I

[Vät97e] Martin Väth. Introduction. Lecture Notes in Mathematics, 1664: 1–6, 1997. CODEN LNMAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093549/.

Vath:1997:OA

[Vät97f] Martin Väth. Operators and applications. Lecture Notes in Mathematics, 1664:105–126, 1997. CODEN LNMAA2. ISBN 3-540-63160-7 (print), 3-540-69192-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093553/.

Varga:1990:AZP

[VC90] R. S. Varga and A. J. Carpenter. Asymptotics for the zeros of the partial sums of e^{z}. II. Lecture Notes in Mathematics, 1435: 201–207, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087909/.

vandenBerg:1991:MR

[vdB91] Imme P. van den Berg. Macroscopic rivers. Lecture Notes in Mathematics, 1493:190–209, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085032/.

vanderGeer:1993:PDC

[vdG93] Gerard van der Geer. Points of degree d on curves over number fields. Lecture Notes in Mathematics, 1566:111–116, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48208-6_12/.

vanderPut:1993:PT

[vdP93] Marius van der Put. The product theorem. Lecture Notes in Mathematics, 1566:77–82, 1993. CODEN LNMAA2. ISBN 3-540-57528-6 (print), 3-540-48208-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48208-6_8/.

vanderPut:1997:ADE

[vdPS97a] Marius van der Put and Michael F. Singer. Algorithms for difference equations. Lecture Notes in Mathematics, 1666:28–34, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096120/.

vanderPut:1997:CCF

[vdPS97b] Marius van der Put and Michael F. Singer. Classification and canonical forms. Lecture Notes in Mathematics, 1666:71–76, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096125/.

vanderPut:1997:DM

[vdPS97c] Marius van der Put and Michael F. Singer. Difference modules over P. Lecture Notes in Mathematics, 1666:60–67, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096124/.

vanderPut:1997:EEG

[vdPS97d] Marius van der Put and Michael F. Singer. Examples of equations and Galois groups. Lecture Notes in Mathematics, 1666: 111–126, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096128/.

vanderPut:1997:EPC

[vdPS97e] Marius van der Put and Michael F. Singer. An excursion in positive characteristic. Lecture Notes in Mathematics, 1666:52– 59, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096123/.

vanderPut:1997:GTD

[vdPS97f] Marius van der Put and Michael F. Singer. Galois Theory of Difference Equations, volume 1666 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 180 pp. LCCN QA3 .L28 no. 1666.

URL http://link.springer.com/book/10.1007/BFb0096118;
 http://www.springerlink.com/content/978-3-540-69241-
6.

vanderPut:1997:IPD

[vdPS97g] Marius van der Put and Michael F. Singer. The inverse problem for difference equations. Lecture Notes in Mathematics, 1666: 35–44, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096121/.

vanderPut:1997:MDE

[vdPS97h] Marius van der Put and Michael F. Singer. Mild difference equations. Lecture Notes in Mathematics, 1666:95–110, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096127/.

vanderPut:1997:PVR

[vdPS97i] Marius van der Put and Michael F. Singer. Picard–Vessiot rings. Lecture Notes in Mathematics, 1666:4–27, 1997. CODEN LN-MAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096119/.

vanderPut:1997:DE

[vdPS97j] Marius van der Put and Michael F. Singer. q-Difference equations. Lecture Notes in Mathematics, 1666:149–174, 1997. CO-DEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096130/.

vanderPut:1997:RS

[vdPS97k] Marius van der Put and Michael F. Singer. The ring S of sequences. Lecture Notes in Mathematics, 1666:45–51, 1997. CO-DEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096122/.

vanderPut:1997:SRD

[vdPS97l] Marius van der Put and Michael F. Singer. Semi-regular difference equations. Lecture Notes in Mathematics, 1666:77–94, 1997. CODEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096126/.

vanderPut:1997:WDE

[vdPS97m] Marius van der Put and Michael F. Singer. Wild difference equations. Lecture Notes in Mathematics, 1666:127–148, 1997. CO-DEN LNMAA2. ISBN 3-540-63243-3 (print), 3-540-69241-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096129/.

Velte:1990:OCS

[Vel90] Waldemar Velte. On optimal constants in some inequalities. Lecture Notes in Mathematics, 1431:158–168, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086066/.

Vershik:1990:TQR

[Ver90] A. M. Vershik. On topological questions of real complexity theory and combinatorial optimization. Lecture Notes in Mathematics, 1453:263–270, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085960/.

Vershinin:1991:BRP

[Ver91] Vladimir V. Vershinin. On bordism rings with principal torsion ideal. Lecture Notes in Mathematics, 1474:295–309, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084753/.

Verdi:1994:NAP

[Ver94] Claudio Verdi. Numerical aspects of parabolic free boundary and hysteresis problems. Lecture Notes in Mathematics, 1584:213– 284, 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073398/.

vanGeemen:1992:PMP

[vG92] Bert van Geemen. Projective models of Picard modular varieties. Lecture Notes in Mathematics, 1515:68–99, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0098338/.

vanGeemen:1994:IHC

[vG94] Bert van Geemen. An introduction to the Hodge conjecture for Abelian varieties. Lecture Notes in Mathematics, 1594:233– 252, 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49046-3_5/.

vanHamme:1990:TGM

[vH90] Lucien van Hamme. Three generalizations of Mahler’s expansion for continuous functions on Z_{p}. Lecture Notes in Mathematics, 1454:356–361, 1990. CODEN LNMAA2. ISBN 3-540-53477-6 (print), 3-540-46906-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0091150/.

Vigue:1990:FPH

[Vig90] Jean-Pierre Vigué. Fixed points of holomorphic mappings. Lecture Notes in Mathematics, 1422:101–106, 1990. CODEN LN-MAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089409/.

Villani:1990:CGA

[Vil90] Vinicio Villani, editor. Complex Geometry and Analysis: Proceedings of the International Symposium in honour of Edoardo Vesentini held in Pisa (Italy), May 23–27, 1988, volume 1422 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52434-7 (print), 3-540-46988-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1422.

URL http://link.springer.com/book/10.1007/BFb0089399;
 http://www.springerlink.com/content/978-3-540-46988-

Viola:1999:ATE

[Vio99] Carlo Viola, editor. Arithmetic Theory of Elliptic Curves: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro, Italy, July 12–19, 1997, volume 1716 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66546-3 (print), 3-540-48160-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1716. URL http:/

/link.springer.com/book/10.1007/BFb0093451;   http://
www.springerlink.com/content/978-3-540-48160-7.

Vistoli:1992:EGG

[Vis92a] Angelo Vistoli. Equivariant Grothendieck groups and equivariant Chow groups. Lecture Notes in Mathematics, 1515:112–133, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0098341/.

Vistoli:1992:TE

[Vis92b] Angelo Vistoli. Trento examples. Lecture Notes in Mathematics, 1515:134–139, 1992. CODEN LNMAA2. ISBN 3-540-55295-2 (print), 3-540-47016-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0098342/.

Visintin:1994:PTH

[Vis94] Augusto Visintin, editor. Phase Transitions and Hysteresis: Lectures given at the 3rd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Montecatini Terme, Italy, July 13–21, 1993, volume 1584 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-58386-6 (print), 3-540-48678-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1584. URL http:/

/link.springer.com/book/10.1007/BFb0073393;   http://
www.springerlink.com/content/978-3-540-48678-7.

Vorobev:1990:DCP

[VK90] Yu. M. Vorob’ev and M. V. Karasev. Deformation and cohomologies of Poisson brackets. Lecture Notes in Mathematics, 1453: 271–289, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085961/.

Vasavada:1992:ADS

[VM92] M. H. Vasavada and R. D. Mehta. Algebra direct sum decomposition of C_{R}(X). Lecture Notes in Mathematics, 1511:113– 119, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093804/.

vanNeerven:1992:T

[vN92a] Jan van Neerven. The σ(X, X^{⊙})-topology. Lecture Notes in Mathematics, 1529:19–39, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085010/.

vanNeerven:1992:APS

[vN92b] Jan van Neerven. The adjoint of a positive semigroup. Lecture Notes in Mathematics, 1529:144–175, 1992. CODEN LN-MAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085016/.

vanNeerven:1992:ASL

[vN92c] Jan van Neerven. The Adjoint of a Semigroup of Linear Operators, volume 1529 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 195 pp. LCCN QA3 .L28 no. 1529. URL http:

//link.springer.com/book/10.1007/BFb0085008;   http:/
/www.springerlink.com/content/978-3-540-47497-5.

vanNeerven:1992:AS

[vN92d] Jan van Neerven. The adjoint semigroup. Lecture Notes in Mathematics, 1529:1–18, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085009/.

vanNeerven:1992:ASR

[vN92e] Jan van Neerven. Adjoint semigroups and the RNP. Lecture Notes in Mathematics, 1529:111–121, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085014/.

vanNeerven:1992:DT

[vN92f] Jan van Neerven. Dichotomy theorems. Lecture Notes in Mathematics, 1529:96–110, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085013/.

vanNeerven:1992:IED

[vN92g] Jan van Neerven. Interpolation, extrapolation and duality. Lecture Notes in Mathematics, 1529:40–68, 1992. CODEN LN-MAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085011/.

vanNeerven:1992:PT

[vN92h] Jan van Neerven. Perturbation theory. Lecture Notes in Mathematics, 1529:69–95, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085012/.

vanNeerven:1992:TP

[vN92i] Jan van Neerven. Tensor products. Lecture Notes in Mathematics, 1529:122–143, 1992. CODEN LNMAA2. ISBN 3-540-56260-5 (print), 3-540-47497-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085015/.

Voisin:1994:TMS

[Voi94] Claire Voisin. Transcendental methods in the study of algebraic cycles. Lecture Notes in Mathematics, 1594:153–222, 1994. CODEN LNMAA2. ISBN 3-540-58692-X (print), 3-540-49046-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49046-3_3/.

Vojta:1993:AAA

[Voj93] Paul Vojta. Applications of arithmetic algebraic geometry to Diophantine approximations. Lecture Notes in Mathematics, 1553: 164–208, 1993. CODEN LNMAA2. ISBN 3-540-57110-8 (print), 3-540-47909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084730/.

Voss:1992:WTC

[Vos92] Conny Voss. On the weights of trace codes. Lecture Notes in Mathematics, 1518:193–198, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0088002/.

Vuorinen:1992:QSM

[Vuo92] Matti Vuorinen, editor. Quasiconformal Space Mappings: A collection of surveys 1960–1990, volume 1508 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1508.

URL http://link.springer.com/book/10.1007/BFb0094234;
 http://www.springerlink.com/content/978-3-540-47061-
8.

vonWahl:1990:NSC

[vW90] Wolf von Wahl. On necessary and sufficient conditions for the solvability of the equations rotμ = γ and divμ = ϵ with μ vanishing on the boundary. Lecture Notes in Mathematics, 1431: 152–157, 1990. CODEN LNMAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0086065/.

vonWaldenfels:1996:CMK

[vW96] Wilhelm von Waldenfels. Continuous Maassen kernels and the inverse oscillator. Lecture Notes in Mathematics, 1626:117– 161, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094646/.

Wacker:1992:IPC

[Wac92] Ulrich Wacker. Invariance principles and central limit theorems for nonadditive, stationary processes. Lecture Notes in Mathematics, 1514:198–228, 1992. CODEN LNMAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097540/.

Wahlbin:1995:CSI

[Wah95a] Lars B. Wahlbin. Chapter 10. Superconvergence in isoparametric mappings of translation invariant meshes: an example. Lecture Notes in Mathematics, 1605:98–106, 1995. CODEN LN-MAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096845/.

Wahlbin:1995:CIS

[Wah95b] Lars B. Wahlbin. A computational investigation of superconvergence for first derivatives in the plane. Lecture Notes in Mathematics, 1605:125–135, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096847/.

Wahlbin:1995:ISS

[Wah95c] Lars B. Wahlbin. Introduction to several space dimensions: some results about superconvergence in L_{2}-projections. Lecture Notes in Mathematics, 1605:42–47, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096839/.

Wahlbin:1995:LGP

[Wah95d] Lars B. Wahlbin. Local and global properties of L_{2}-projections. Lecture Notes in Mathematics, 1605:36–41, 1995. CODEN LN-MAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096838/.

Wahlbin:1995:SNP

[Wah95e] Lars B. Wahlbin. On superconvergence in nonlinear problems. Lecture Notes in Mathematics, 1605:93–97, 1995. CODEN LN-MAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096844/.

Wahlbin:1995:RAS

[Wah95f] Lars B. Wahlbin. Remarks about some of the tools used in chapter 1. Lecture Notes in Mathematics, 1605:28–35, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096837/.

Wahlbin:1995:SOE

[Wah95g] Lars B. Wahlbin. Second order elliptic boundary value problems in any number of space dimensions: preliminary considerations on local and global estimates and presentation of the main technical tools for showing superconvergence. Lecture Notes in Mathematics, 1605:48–64, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096840/.

Wahlbin:1995:SOD

[Wah95h] Lars B. Wahlbin. Some one-dimensional superconvergence results. Lecture Notes in Mathematics, 1605:1–27, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096836/.

Wahlbin:1995:SAM

[Wah95i] Lars B. Wahlbin. Superconvergence by averaging: mainly, the K-operator. Lecture Notes in Mathematics, 1605:107–124, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0096846/.

Wahlbin:1995:SLS

[Wah95j] Lars B. Wahlbin. Superconvergence by local symmetry. Lecture Notes in Mathematics, 1605:74–83, 1995. CODEN LN-MAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096842/.

Wahlbin:1995:SDQ

[Wah95k] Lars B. Wahlbin. Superconvergence for difference quotients on translation invariant meshes. Lecture Notes in Mathematics, 1605: 84–92, 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096843/.

Wahlbin:1995:SGF

[Wah95l] Lars B. Wahlbin. Superconvergence in Galerkin Finite Element Methods, volume 1605 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 166 pp. LCCN QA3 .L28 no. 1605. URL http:/

/link.springer.com/book/10.1007/BFb0096835;   http://
www.springerlink.com/content/978-3-540-49401-0.

Wahlbin:1995:STP

[Wah95m] Lars B. Wahlbin. Superconvergence in tensor-product elements. Lecture Notes in Mathematics, 1605:65–73, 1995. CODEN LN-MAA2. ISBN 3-540-60011-6 (print), 3-540-49401-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096841/.

Waldschmidt:1990:IAN

[Wal90a] Michel Waldschmidt. Indépendance algébrique de nombres de Liouville. (French) []. Lecture Notes in Mathematics, 1415:225– 235, 1990. CODEN LNMAA2. ISBN 3-540-52190-9 (print), 3-540-46919-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084891/.

Wall:1990:DLR

[Wal90b] G. E. Wall. Dependence of Lie relators for Burnside varieties. Lecture Notes in Mathematics, 1456:191–197, 1990. CODEN LN-MAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100741/.

Wallach:1990:LM

[Wal90c] Nolan R. Wallach. Limit multiplicities in L^{2}(Γ\G). Lecture Notes in Mathematics, 1447:31–56, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085725/.

Wallet:1991:OAD

[Wal91a] Guy Wallet. Overstability in arbitrary dimension. Lecture Notes in Mathematics, 1493:57–70, 1991. CODEN LNMAA2. ISBN 3-540-54900-5 (print), 3-540-46471-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085024/.

Waltman:1991:BSP

[Wal91b] Paul Waltman. A brief survey of persistence in dynamical systems. Lecture Notes in Mathematics, 1475:31–40, 1991. CODEN LNMAA2. ISBN 3-540-54120-9 (print), 3-540-47418-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0083477/.

Walsh:1993:SR

[Wal93] John B. Walsh. Some remarks on A(t, B_{t}). Lecture Notes in Mathematics, 1557:173–176, 1993. CODEN LNMAA2. ISBN 3-540-57282-1 (print), 3-540-48034-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087973/.

Wang:1990:EFJ

[Wan90] S. P. Wang. Embedding of Flensted–Jensen modules in L^{2}(Γ \ G) in the noncompact case. Lecture Notes in Mathematics, 1447: 333–356, 1990. CODEN LNMAA2. ISBN 3-540-53422-9 (print), 3-540-46876-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085737/.

Wan:1991:VDI

[Wan91a] Yieh-Hei Wan. Versal deformations of infinitesimally symplectic transformations with antisymplectic involutions. Lecture Notes in Mathematics, 1463:301–320, 1991. CODEN LNMAA2. ISBN 3-540-53736-8 (print), 3-540-47047-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0085438/.

Wang:1991:WCA

[Wan91b] Xue-Ping Wang. Weak coupling asymptotics of Schrödinger operators with Stark effect. Lecture Notes in Mathematics, 1494: 185–195, 1991. CODEN LNMAA2. ISBN 3-540-54901-3 (print), 3-540-46474-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087771/.

Wang:1994:HOC

[Wan94] Xiangdong Wang. The Hecke operators on the cusp forms of Γ_{0}(N) with nebentype. Lecture Notes in Mathematics, 1585:95– 108, 1994. CODEN LNMAA2. ISBN 3-540-58387-4 (print), 3-540-48681-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074111/.

Warren:1997:BPR

[War97] Jonathan Warren. Branching processes, the Ray–Knight theorem, and sticky Brownian motion. Lecture Notes in Mathematics, 1655:1–15, 1997. CODEN LNMAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0119287/.

Warren:1999:JSB

[War99] J. Warren. On the joining of sticky Brownian motion. Lecture Notes in Mathematics, 1709:257–266, 1999. CODEN LN-MAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096515/.

Watanabe:1990:STA

[Wat90] Shinzo Watanabe. Short time asymptotic problems in Wiener functional integration theory. Applications to heat kernels and index theorems. Lecture Notes in Mathematics, 1444:1–62, 1990. CODEN LNMAA2. ISBN 3-540-53064-9 (print), 3-540-46596-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0083609/.

Watanabe:1999:EMS

[Wat99] Shinzo Watanabe. The existence of a multiple spider martingale in the natural filtration of a certain diffusion in the plane.

Lecture Notes in Mathematics, 1709:277–290, 1999. CODEN LNMAA2. ISBN 3-540-66342-8 (print), 3-540-48407-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096517/.

Wehlen:1990:ARM

[Weh90] Joseph A. Wehlen. Azumaya rings and Maschke’s Theorem. Lecture Notes in Mathematics, 1448:19–24, 1990. CODEN LN-MAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0091247/.

Weisz:1994:IVF

[Wei94a] Ferenc Weisz. Inequalities for Vilenkin–Fourier coefficients. Lecture Notes in Mathematics, 1568:183–203, 1994. CODEN LN-MAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073454/.

Weisz:1994:MHS

[Wei94b] Ferenc Weisz. Martingale Hardy Spaces and their Applications in Fourier Analysis, volume 1568 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNMAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 218 pp. LCCN QA3 .L28 no. 1568.

URL http://link.springer.com/book/10.1007/BFb0073448;
 http://www.springerlink.com/content/978-3-540-48295-
6.

Weisz:1994:OPM

[Wei94c] Ferenc Weisz. One-parameter martingale Hardy spaces. Lecture Notes in Mathematics, 1568:13–79, 1994. CODEN LN-MAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073450/.

Weisz:1994:PN

[Wei94d] Ferenc Weisz. Preliminaries and notations. Lecture Notes in Mathematics, 1568:1–12, 1994. CODEN LNMAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073449/.

Weisz:1994:RI

[Wei94e] Ferenc Weisz. Real interpolation. Lecture Notes in Mathematics, 1568:164–182, 1994. CODEN LNMAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073453/.

Weisz:1994:TM

[Wei94f] Ferenc Weisz. Tree martingales. Lecture Notes in Mathematics, 1568:141–163, 1994. CODEN LNMAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073452/.

Weisz:1994:TPM

[Wei94g] Ferenc Weisz. Two-parameter martingale Hardy spaces. Lecture Notes in Mathematics, 1568:80–140, 1994. CODEN LN-MAA2. ISBN 3-540-57623-1 (print), 3-540-48295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073451/.

Werner:1994:REA

[Wer94] Wendelin Werner. Rate of explosion of the Ampèrean area of the planar Brownian loop. Lecture Notes in Mathematics, 1583: 153–163, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073843/.

Werner:1995:SRP

[Wer95] Wendelin Werner. Some remarks on perturbed reflecting Brownian motion. Lecture Notes in Mathematics, 1613:37–43, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0094198/.

Wesolowski:1993:MID

[Wes93] Jacek Wesołowski. Multivariate infinitely divisible distributions with the Gaussian second order conditional structure. Lecture Notes in Mathematics, 1546:180–183, 1993. CODEN LN-MAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084495/.

Wetton:1992:FDV

[Wet92] Brian T. R. Wetton. Finite difference vorticity methods. Lecture Notes in Mathematics, 1530:210–225, 1992. CODEN LN-MAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090344/.

Wicks:1991:FH

[Wic91a] Keith R. Wicks. Fractals and Hyperspaces, volume 1492 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 168 pp. LCCN QA3 .L28 no. 1492. URL http:/

/link.springer.com/book/10.1007/BFb0089156;   http://
www.springerlink.com/content/978-3-540-46610-9.

Wicks:1991:HIS

[Wic91b] Keith R. Wicks. Hutchinson’s invariant sets. Lecture Notes in Mathematics, 1492:44–85, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089161/.

Wicks:1991:I

[Wic91c] Keith R. Wicks. Introduction. Lecture Notes in Mathematics, 1492:1–2, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0089157.

Wicks:1991:NDH

[Wic91d] Keith R. Wicks. Nonstandard development of the Hausdorff metric. Lecture Notes in Mathematics, 1492:29–43, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089160/.

Wicks:1991:NDV

[Wic91e] Keith R. Wicks. Nonstandard development of the Vietoris topology. Lecture Notes in Mathematics, 1492:13–28, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089159/.

Wicks:1991:P

[Wic91f] Keith R. Wicks. Preliminaries. Lecture Notes in Mathematics, 1492:3–12, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089158/.

Wicks:1991:VFN

[Wic91g] Keith R. Wicks. Views and fractal notions. Lecture Notes in Mathematics, 1492:86–130, 1991. CODEN LNMAA2. ISBN 3-540-54965-X (print), 3-540-46610-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089162/.

Wiegner:1990:DSS

[Wie90] Michael Wiegner. Decay and stability in L_{p} for strong solutions of the Cauchy problem for the Navier–Stokes equations. Lecture Notes in Mathematics, 1431:95–99, 1990. CODEN LN-MAA2. ISBN 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086059/.

Wiegner:1992:AWS

[Wie92] Michael Wiegner. Approximation of weak solutions of the Navier– Stokes equations in unbounded domains. Lecture Notes in Mathematics, 1530:161–166, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090340/.

Williams:1995:NLW

[Wil95] David Williams. Non-linear Wiener–Hopf theory, 1: an appetizer. Lecture Notes in Mathematics, 1613:155–161, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094207/.

Wildeshaus:1997:CCM

[Wil97a] Jörg Wildeshaus. The canonical construction of mixed sheaves on mixed Shimura varieties. Lecture Notes in Mathematics, 1650:

77–140, 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093054/.

Wildeshaus:1997:I

[Wil97b] Jörg Wildeshaus. Introduction. Lecture Notes in Mathematics, 1650:1–21, 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093052/.

Wildeshaus:1997:MSF

[Wil97c] Jörg Wildeshaus. Mixed structures on fundamental groups. Lecture Notes in Mathematics, 1650:23–76, 1997. CODEN LN-MAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093053/.

Wildeshaus:1997:PEMc

[Wil97d] Jörg Wildeshaus. Polygogarithmic extensions on mixed Shimura varieties. Part III: The elliptic polygogarithm. Lecture Notes in Mathematics, 1650:249–335, 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093057/.

Wildeshaus:1997:PEMa

[Wil97e] Jörg Wildeshaus. Polylogarithmic extensions on mixed Shimura varieties. Part I: Construction and basic properties. Lecture Notes in Mathematics, 1650:141–197, 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093055/.

Wildeshaus:1997:PEMb

[Wil97f] Jörg Wildeshaus. Polylogarithmic extensions on mixed Shimura varieties. Part II: The classical polylogarithm. Lecture Notes in Mathematics, 1650:199–248, 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093056/.

Wildeshaus:1997:RP

[Wil97g] Jörg Wildeshaus. Realizations of Polylogarithms, volume 1650 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNMAA2. ISBN 3-540-62460-0 (print), 3-540-49728-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 343 pp. LCCN QA3 .L28 no. 1650. URL http:/

/link.springer.com/book/10.1007/BFb0093051;   http://
www.springerlink.com/content/978-3-540-49728-8.

Winkler:1990:SWP

[Win90] R. Winkler. Strong Weyl property in uniform distribution. Lecture Notes in Mathematics, 1452:208–220, 1990. CODEN LN-MAA2. ISBN 3-540-53408-3 (print), 3-540-46864-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096993/.

Winkelmann:1995:CTDc

[Win95a] Jörg Winkelmann. The Classification of Three-Dimensional Homogeneous Complex Manifolds, volume 1602 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CO-DEN LNMAA2. ISBN 3-540-59072-2 (print), 3-540-49185-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 230 pp. LCCN QA3 .L28 no. 1602. URL http:/

/link.springer.com/book/10.1007/BFb0095837;   http://
www.springerlink.com/content/978-3-540-49185-9.

Winkelmann:1995:CTDa

[Win95b] Jörg Winkelmann. The classification of three-dimensional homogeneous complex manifolds X = G/H where G is a complex Lie group. Lecture Notes in Mathematics, 1602:20–84, 1995. CODEN LNMAA2. ISBN 3-540-59072-2 (print), 3-540-49185-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095839/.

Winkelmann:1995:CTDb

[Win95c] Jörg Winkelmann. The classification of three-dimensional homogeneous complex manifolds X = G/H where G is a real Lie group. Lecture Notes in Mathematics, 1602:85–224, 1995. CO-DEN LNMAA2. ISBN 3-540-59072-2 (print), 3-540-49185-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0095840/.

Winkelmann:1995:S

[Win95d] Jörg Winkelmann. Survey. Lecture Notes in Mathematics, 1602: 1–19, 1995. CODEN LNMAA2. ISBN 3-540-59072-2 (print), 3-540-49185-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095838/.

Wirsching:1998:AAC

[Wir98a] G. J. Wirsching. 3-adic averages of counting functions. Lecture Notes in Mathematics, 1681:76–95, 1998. CODEN LN-MAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095989/.

Wirsching:1998:ACG

[Wir98b] G. J. Wirsching. Analysis of the Collatz graph. Lecture Notes in Mathematics, 1681:31–75, 1998. CODEN LNMAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095988/.

Wirsching:1998:AHM

[Wir98c] G. J. Wirsching. An asymptotically homogeneous Markov chain. Lecture Notes in Mathematics, 1681:96–122, 1998. CODEN LN-MAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095990/.

Wirsching:1998:I

[Wir98d] G. J. Wirsching. Introduction. Lecture Notes in Mathematics, 1681:1–9, 1998. CODEN LNMAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0095986/.

Wirsching:1998:MPD

[Wir98e] G. J. Wirsching. Mixing and predecessor density. Lecture Notes in Mathematics, 1681:123–140, 1998. CODEN LNMAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0095991/.

Wirsching:1998:SIA

[Wir98f] G. J. Wirsching. Some ideas around 3n + 1 iterations. Lecture Notes in Mathematics, 1681:10–30, 1998. CODEN LN-MAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0095987/.

Wirsching:1998:DSG

[Wir98g] Günther J. Wirsching. The Dynamical System Generated by the 3n + 1 Function, volume 1681 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-63970-5 (print), 3-540-69677-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 158 pp. LCCN QA3 .L28 no. 1681.

URL http://link.springer.com/book/10.1007/BFb0095985;
 http://www.springerlink.com/content/978-3-540-69677-
3.

Wittmann:1992:FPR

[Wit92] Rainer Wittmann. Fixed point rays of nonexpansive mappings. Lecture Notes in Mathematics, 1514:229–233, 1992. CODEN LN-MAA2. ISBN 3-540-55444-0 (print), 3-540-47076-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0097541/.

Wujie:1990:CPE

[WJ90] Shi Wujie and Bi Jianxing. A characteristic property for each finite projective special linear group. Lecture Notes in Mathematics, 1456:171–180, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100738/.

Wojtaszczyk:1991:SRA

[Woj91] P. Wojtaszczyk. Some remarks about the space of measures with uniformly bounded partial sums and Banach–Mazur distances between some spaces of polynomials. Lecture Notes in Mathematics, 1470:60–67, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090212/.

Wolpert:1991:SRS

[Wol91] Scott A. Wolpert. The spectrum of a Riemann surface with a cusp. Lecture Notes in Mathematics, 1468:203–226, 1991. CODEN LNMAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086195/.

Wolff:1992:HCF

[Wol92] Michael Wolff. Heat-conducting fluids with free surface in the case of slip-condition on the walls. Lecture Notes in Mathematics, 1530:56–70, 1992. CODEN LNMAA2. ISBN 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0090333/.

Wood:1998:DCF

[Woo98] Carol Wood. Differentially closed fields. Lecture Notes in Mathematics, 1696:129–141, 1998. CODEN LNMAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-68521-0_8/.

Woyczynski:1998:BKT

[Woy98a] Wojbor A. Woyczyński. Burgers–KPZ Turbulence: Göttingen Lectures, volume 1700 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 318 pp. LCCN QA3 .L28 no. 1700. URL http:/

/link.springer.com/book/10.1007/BFb0093107;   http://
www.springerlink.com/content/978-3-540-49480-5.

Woyczynski:1998:FBT

[Woy98b] Wojbor A. Woyczyński. Forced Burgers turbulence. Lecture Notes in Mathematics, 1700:135–201, 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093113/.

Woyczynski:1998:FBK

[Woy98c] Wojbor A. Woyczyński. Fractal Burgers–KPZ models. Lecture Notes in Mathematics, 1700:271–298, 1998. CODEN LN-MAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093115/.

Woyczynski:1998:HCF

[Woy98d] Wojbor A. Woyczyński. Hopf–Cole formula and its asymptotic analysis. Lecture Notes in Mathematics, 1700:25–42, 1998. CO-DEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093110/.

Woyczynski:1998:HLN

[Woy98e] Wojbor A. Woyczyński. Hydrodynamic limits, nonlinear diffusions, and propagation of chaos. Lecture Notes in Mathematics, 1700:13–24, 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093109/.

Woyczynski:1998:HAI

[Woy98f] Wojbor A. Woyczyński. Hyperbolic approximation and inviscid limit. Lecture Notes in Mathematics, 1700:97–133, 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093112/.

Woyczynski:1998:PTT

[Woy98g] Wojbor A. Woyczyński. Passive tracer transport in Burgers’ and related flows. Lecture Notes in Mathematics, 1700:203–270, 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093114/.

Woyczynski:1998:SWL

[Woy98h] Wojbor A. Woyczyński. Shock waves and the large scale structure (LSS) of the universe. Lecture Notes in Mathematics, 1700: 1–11, 1998. CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093108/.

Woyczynski:1998:SDP

[Woy98i] Wojbor A. Woyczyński. Statistical description, parabolic approximation. Lecture Notes in Mathematics, 1700:43–95, 1998.

CODEN LNMAA2. ISBN 3-540-65237-X (print), 3-540-49480-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093111/.

Wright:1990:LBP

[Wri90] Mary H. Wright. Links between prime ideals of a serial ring with Krull dimension. Lecture Notes in Mathematics, 1448:33–40, 1990. CODEN LNMAA2. ISBN 3-540-53164-5 (print), 3-540-46745-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0091249/.

Weiss:1990:DRD

[WS90] R. Weiss and W. Schönauer. Data reduction (dare) preconditioning for generalized conjugate gradient methods. Lecture Notes in Mathematics, 1457:137–153, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0090906/.

Wurgler:1991:MTS

[Wür91] Urs Würgler. Morava K-theories: a survey. Lecture Notes in Mathematics, 1474:111–138, 1991. CODEN LNMAA2. ISBN 3-540-54098-9 (print), 3-540-47403-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084741/.

Wadhwa:1992:AWB

[WY92] B. L. Wadhwa and B. S. Yadav. Approximation in weighted Banach spaces of power series. Lecture Notes in Mathematics, 1511: 33–46, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0093795/.

Xambo-Descamps:1990:EGP

[XD90] Sebastian Xambó-Descamps, editor. Enumerative Geometry: Proceedings of a Conference held in Sitges, Spain, June 1–6, 1987, volume 1436 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CODEN LNMAA2. ISBN 3-540-52811-3 (print), 3-540-47154-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1436. URL http:/

/link.springer.com/book/10.1007/BFb0084038;   http://
www.springerlink.com/content/978-3-540-47154-7.

Xi:1994:AWM

[Xi94a] Nanhua Xi. q_{s}-Analogue of weight multiplicity. Lecture Notes in Mathematics, 1587:44–47, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074134/.

Xi:1994:AWG

[Xi94b] Nanhua Xi. Affine Weyl groups and affine Hecke algebras. Lecture Notes in Mathematics, 1587:18–26, 1994. CODEN LN-MAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074132/.

Xi:1994:BRC

[Xi94c] Nanhua Xi. The based rings of cells in affine Weyl groups of type G̃_{2}, B̃_{2}, Ã_{2}. Lecture Notes in Mathematics, 1587:102–115, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0074141/.

Xi:1994:ER

[Xi94d] Nanhua Xi. An equivalence relation in T × C^{∗}. Lecture Notes in Mathematics, 1587:63–79, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074136/.

Xi:1994:GTS

[Xi94e] Nanhua Xi. A generalized two-sided cell of an affine Weyl group. Lecture Notes in Mathematics, 1587:27–43, 1994. CODEN LN-MAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074133/.

Xi:1994:HA

[Xi94f] Nanhua Xi. Hecke algebras. Lecture Notes in Mathematics, 1587: 1–17, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074131/.

Xi:1994:IAH

[Xi94g] Nanhua Xi. Isogenous affine Hecke algebras. Lecture Notes in Mathematics, 1587:93–98, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074139/.

Xi:1994:KLC

[Xi94h] Nanhua Xi. Kazhdan–Lusztig classification on simple modules of affine Hecke algebras. Lecture Notes in Mathematics, 1587: 48–62, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0074135/.

Xi:1994:LTS

[Xi94i] Nanhua Xi. The lowest two-sided cell. Lecture Notes in Mathematics, 1587:80–84, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074137/.

Xi:1994:PSR

[Xi94j] Nanhua Xi. Principal series representations and induced modules. Lecture Notes in Mathematics, 1587:85–92, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0074138/.

Xi:1994:QA

[Xi94k] Nanhua Xi. Quotient algebras. Lecture Notes in Mathematics, 1587:99–101, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0074140/.

Xi:1994:RAH

[Xi94l] Nanhua Xi. Representations of Affine Hecke Algebras, volume 1587 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 137 pp. LCCN QA3 .L28 no. 1587. URL http:/

/link.springer.com/book/10.1007/BFb0074130;   http://
www.springerlink.com/content/978-3-540-48682-4.

Xi:1994:SMA

[Xi94m] Nanhua Xi. Simple modules attached to c_{1}. Lecture Notes in Mathematics, 1587:116–128, 1994. CODEN LNMAA2. ISBN 3-540-58389-0 (print), 3-540-48682-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0074142/.

Xia:1992:WCJ

[Xia92] Aihua Xia. Weak convergence of jump processes. Lecture Notes in Mathematics, 1526:32–46, 1992. CODEN LNMAA2. ISBN 3-540-56021-1 (print), 3-540-47342-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084308/.

Xiao:1998:APb

[XL98a] Ti-Jun Xiao and Jin Liang. Almost periodicity. Lecture Notes in Mathematics, 1701:239–261, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-49479-9_7/.

Xiao:1998:APa

[XL98b] Ti-Jun Xiao and Jin Liang. Analyticity and parabolicity. Lecture Notes in Mathematics, 1701:141–176, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-49479-9_4/.

Xiao:1998:BM

[XL98c] Ti-Jun Xiao and Jin Liang. Back matter. Lecture Notes in Mathematics, 1701:263–309, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-49479-9/1.pdf.

Xiao:1998:CPH

[XL98d] Ti-Jun Xiao and Jin Liang. The Cauchy Problem for Higher Order Abstract Differential Equations, volume 1701 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 301 pp. LCCN QA3 .L28 no.

1701. URL http://link.springer.com/book/10.1007/978-
3-540-49479-9;   http://www.springerlink.com/content/
978-3-540-49479-9.

Xiao:1998:DNC

[XL98e] Ti-Jun Xiao and Jin Liang. Differentiability and norm continuity. Lecture Notes in Mathematics, 1701:199–238, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-49479-9_
6/.

Xiao:1998:EGB

[XL98f] Ti-Jun Xiao and Jin Liang. Exponential growth bound and exponential stability. Lecture Notes in Mathematics, 1701:177– 197, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-49479-9_5/.

Xiao:1998:FM

[XL98g] Ti-Jun Xiao and Jin Liang. Front matter. Lecture Notes in Mathematics, 1701:N2–XII, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-49479-9/1.pdf.

Xiao:1998:GW

[XL98h] Ti-Jun Xiao and Jin Liang. Generalized wellposedness. Lecture Notes in Mathematics, 1701:85–140, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-49479-9_3/.

Xiao:1998:LTO

[XL98i] Ti-Jun Xiao and Jin Liang. Laplace transforms and operator families in locally convex spaces. Lecture Notes in Mathematics, 1701:1–44, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-49479-9_1/.

Xiao:1998:WS

[XL98j] Ti-Jun Xiao and Jin Liang. Wellposedness and solvability. Lecture Notes in Mathematics, 1701:45–83, 1998. CODEN LNMAA2. ISBN 3-540-65238-8 (print), 3-540-49479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-49479-9_2/.

Xu:1996:ACR

[Xu96a] Jinzhong Xu. Applications in commutative rings. Lecture Notes in Mathematics, 1634:107–151, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094179/.

Xu:1996:EC

[Xu96b] Jinzhong Xu. Envelopes and covers. Lecture Notes in Mathematics, 1634:5–25, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094175/.

Xu:1996:FCCa

[Xu96c] Jinzhong Xu. Flat covers and cotorsion envelopes. Lecture Notes in Mathematics, 1634:51–79, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094177/.

Xu:1996:FCM

[Xu96d] Jinzhong Xu. Flat Covers of Modules, volume 1634 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 161 pp. LCCN QA3 .L28 no. 1634. URL http:/

/link.springer.com/book/10.1007/BFb0094173;   http://
www.springerlink.com/content/978-3-540-69992-7.

Xu:1996:FCCb

[Xu96e] Jinzhong Xu. Flat covers over commutative rings. Lecture Notes in Mathematics, 1634:81–106, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094178/.

Xu:1996:FT

[Xu96f] Jinzhong Xu. Fundamental theorems. Lecture Notes in Mathematics, 1634:27–50, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0094176/.

Xu:1996:I

[Xu96g] Jinzhong Xu. Introduction. Lecture Notes in Mathematics, 1634: 1–3, 1996. CODEN LNMAA2. ISBN 3-540-61640-3 (print), 3-540-69992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094174/.

Xue:1990:ZOL

[Xue90] Xing-Xiong Xue. A zero–one law for integral functionals of the Bessel process. Lecture Notes in Mathematics, 1426:137– 153, 1990. CODEN LNMAA2. ISBN 3-540-52694-3 (print), 3-540-47098-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083762/.

Xue:1992:ARI

[Xue92a] Weimin Xue. Artinian rings (II) — Azumaya’s exact rings. Lecture Notes in Mathematics, 1523:111–148, 1992. CODEN LN-MAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0089075/.

Xue:1992:ARM

[Xue92b] Weimin Xue. Artinian rings with Morita duality (I). Lecture Notes in Mathematics, 1523:84–110, 1992. CODEN LNMAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089074/.

Xue:1992:IMD

[Xue92c] Weimin Xue. Introduction to Morita duality. Lecture Notes in Mathematics, 1523:1–53, 1992. CODEN LNMAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089072/.

Xue:1992:MDR

[Xue92d] Weimin Xue. Morita duality and ring extensions. Lecture Notes in Mathematics, 1523:54–83, 1992. CODEN LNMAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089073/.

Xue:1992:OTR

[Xue92e] Weimin Xue. Other types of rings with duality. Lecture Notes in Mathematics, 1523:149–159, 1992. CODEN LNMAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0089076/.

Xue:1992:RMD

[Xue92f] Weimin Xue. Rings with Morita Duality, volume 1523 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55770-9 (print), 3-540-47283-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 167 pp. LCCN QA3 .L28 no. 1523. URL http:/

/link.springer.com/book/10.1007/BFb0089071;   http://
www.springerlink.com/content/978-3-540-47283-4.

Yajima:1990:SPS

[Yaj90] Kenji Yajima. On smoothing property of Schrödinger propagators. Lecture Notes in Mathematics, 1450:20–35, 1990. CODEN LNMAA2. ISBN 3-540-53393-1 (print), 3-540-46818-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0084896/.

Yamaguchi:1990:SHE

[Yam90] Kohhei Yamaguchi. Self-homotopy equivalences and highly connected Poincaré complexes. Lecture Notes in Mathematics, 1425: 157–169, 1990. CODEN LNMAA2. ISBN 3-540-52658-7 (print), 3-540-47091-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083839/.

Yanqian:1990:RVF

[Yan90] Ye Yanqian. Rotated vector fields decomposition method and its application. Lecture Notes in Mathematics, 1455:385–392, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085403/.

Yan:1991:NWS

[Yan91a] J. A. Yan. Notes on the Wiener semigroup and renormalization. Lecture Notes in Mathematics, 1485:79–94, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100848/.

Yan:1991:SRT

[Yan91b] J. A. Yan. Some remarks on the theory of stochastic integration. Lecture Notes in Mathematics, 1485:95–107, 1991. CODEN LNMAA2. ISBN 3-540-54616-2 (print), 3-540-38496-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100849/.

Yanushkevichiene:1993:CRS

[Yan93] O. L. Yanushkevichiene. On the convergence of random symmetric polynomials. Lecture Notes in Mathematics, 1546:184–188, 1993. CODEN LNMAA2. ISBN 3-540-56744-5 (print), 3-540-47645-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0084496/.

Yan:1996:AEH

[Yan96] J. A. Yan. An asymptotic evaluation of heat kernel for short time. Lecture Notes in Mathematics, 1626:104–107, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094644/.

Yap:1996:BTI

[Yap96a] Hian-Poh Yap. Basic terminology and introduction. Lecture Notes in Mathematics, 1623:1–6, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092896/.

Yap:1996:CTT

[Yap96b] Hian-Poh Yap. Classification of type 1 and type 2 graphs. Lecture Notes in Mathematics, 1623:53–95, 1996. CODEN LN-MAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092901/.

Yap:1996:CPG

[Yap96c] Hian-Poh Yap. Complete r-partite graphs. Lecture Notes in Mathematics, 1623:15–24, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092898/.

Yap:1996:CR

[Yap96d] Hian-Poh Yap. Concluding remarks. Lecture Notes in Mathematics, 1623:114–120, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092904/.

Yap:1996:GHD

[Yap96e] Hian-Poh Yap. Graphs of high degree. Lecture Notes in Mathematics, 1623:35–52, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092900/.

Yap:1996:GLD

[Yap96f] Hian-Poh Yap. Graphs of low degree. Lecture Notes in Mathematics, 1623:25–34, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092899/.

Yap:1996:SBR

[Yap96g] Hian-Poh Yap. Some basic results. Lecture Notes in Mathematics, 1623:7–14, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092897/.

Yap:1996:SUB

[Yap96h] Hian-Poh Yap. Some upper bounds for the total chromatic number of graphs. Lecture Notes in Mathematics, 1623:104–113, 1996. CODEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092903/.

Yap:1996:TCN

[Yap96i] Hian-Poh Yap. Total chromatic number of planar graphs. Lecture Notes in Mathematics, 1623:96–103, 1996. CODEN LN-MAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092902/.

Yap:1996:TCG

[Yap96j] Hian-Poh Yap. Total Colourings of Graphs, volume 1623 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNMAA2. ISBN 3-540-60717-X (print), 3-540-49301-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 131 pp. LCCN QA3 .L28 no. 1623. URL http:/

/link.springer.com/book/10.1007/BFb0092895;   http://
www.springerlink.com/content/978-3-540-49301-3.

Yau:1992:CSP

[Yau92] Shing-Tung Yau. The current state and prospects of geometry and nonlinear differential equations. Lecture Notes in Mathematics, 1525:29–39, 1992. CODEN LNMAA2. ISBN 3-540-56011-4 (print), 3-540-47341-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0089203/.

Yee:1990:GCT

[Yee90] Lee Peng Yee. Generalized convergence theorems for Denjoy– Perron integrals. Lecture Notes in Mathematics, 1419:97–109, 1990. CODEN LNMAA2. ISBN 3-540-52322-7 (print), 3-540-46955-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0083102/.

Yang:1992:TMD

[YK92] Kyeongcheol Yang and P. Vijay Kumar. On the true minimum distance of Hermitian codes. Lecture Notes in Mathematics, 1518: 99–107, 1992. CODEN LNMAA2. ISBN 3-540-55651-6 (print), 3-540-47267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087995/.

Yevenes:1990:MRP

[YR90] Antonio Córdova Yévenes and Stephan Ruscheweyh. On the maximal range problem for slit domains. Lecture Notes in Mathematics, 1435:33–44, 1990. CODEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0087895/.

Yadav:1992:FAO

[YS92] B. S. Yadav and D. Singh, editors. Functional Analysis and Operator Theory: Proceedings of a Conference held in Memory of U. N. Singh New Delhi, India, 2–6 August, 1990, volume 1511 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1511.

URL http://link.springer.com/book/10.1007/BFb0093791;
 http://www.springerlink.com/content/978-3-540-47041-
0.

Yadav:1992:BMT

[YSA92] B. S. Yadav, Dinesh Singh, and Sanjeev Agrawal. De Branges modules in H^{2}(C^{K}) of the torus. Lecture Notes in Mathematics, 1511:55–74, 1992. CODEN LNMAA2. ISBN 3-540-55365-7 (print), 3-540-47041-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093797/.

Yukich:1998:AE

[Yuk98a] Joseph E. Yukich. Applications and examples. Lecture Notes in Mathematics, 1675:97–109, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093480/.

Yukich:1998:AEF

[Yuk98b] Joseph E. Yukich. Asymptotics for Euclidean functionals: The uniform case. Lecture Notes in Mathematics, 1675:32–52, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0093476/.

Yukich:1998:GLP

[Yuk98c] Joseph E. Yukich. Geometric location problems. Lecture Notes in Mathematics, 1675:126–130, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093482/.

Yukich:1998:I

[Yuk98d] Joseph E. Yukich. Introduction. Lecture Notes in Mathematics, 1675:1–8, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0093473/.

Yukich:1998:ICI

[Yuk98e] Joseph E. Yukich. Isoperimetry and concentration inequalities. Lecture Notes in Mathematics, 1675:64–77, 1998. CODEN LN-MAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093478/.

Yukich:1998:MT

[Yuk98f] Joseph E. Yukich. Minimal triangulations. Lecture Notes in Mathematics, 1675:110–125, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093481/.

Yukich:1998:PTC

[Yuk98g] Joseph E. Yukich. Probability Theory of Classical Euclidean Optimization Problems, volume 1675 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 152 pp. LCCN QA3 .L28 no. 1675; QA3 L28 no. 1675

QA273.45. URL http://link.springer.com/book/10.1007/
BFb0093472;   http://www.loc.gov/catdir/enhancements/
fy0812/98010342-d.html;   http://www.loc.gov/catdir/enhancements/
fy0812/98010342-t.html;   http://www.springerlink.com/
content/978-3-540-69627-8.

Yukich:1998:RCH

[Yuk98h] Joseph E. Yukich. Rates of convergence and heuristics. Lecture Notes in Mathematics, 1675:53–63, 1998. CODEN LN-MAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093477/.

Yukich:1998:SSE

[Yuk98i] Joseph E. Yukich. Subadditive and superadditive Euclidean functionals. Lecture Notes in Mathematics, 1675:18–31, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093475/.

Yukich:1998:SS

[Yuk98j] Joseph E. Yukich. Subadditivity and superadditivity. Lecture Notes in Mathematics, 1675:9–17, 1998. CODEN LN-MAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093474/.

Yukich:1998:UTE

[Yuk98k] Joseph E. Yukich. Umbrella theorems for Euclidean functionals. Lecture Notes in Mathematics, 1675:78–96, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0093479/.

Yukich:1998:WCG

[Yuk98l] Joseph E. Yukich. Worst case growth rates. Lecture Notes in Mathematics, 1675:131–137, 1998. CODEN LNMAA2. ISBN 3-540-63666-8 (print), 3-540-69627-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0093483/.

Yurinsky:1995:FAM

[Yur95a] Vadim Vladimirovich Yurinsky. Fine asymptotics of moderate deviations. Lecture Notes in Mathematics, 1617:217–254, 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092605/.

Yurinsky:1995:GRA

[Yur95b] Vadim Vladimirovich Yurinsky. Gaussian and related approximations for distributions of sums. Lecture Notes in Mathematics, 1617:163–216, 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092604/.

Yurinsky:1995:GME

[Yur95c] Vadim Vladimirovich Yurinsky. Gaussian measures in Euclidean space. Lecture Notes in Mathematics, 1617:1–42, 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0092600/.

Yurinsky:1995:ISS

[Yur95d] Vadim Vladimirovich Yurinsky. Inequalities for seminorms: Sums of independent random vectors. Lecture Notes in Mathematics, 1617:79–122, 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0092602/.

Yurinsky:1995:RAL

[Yur95e] Vadim Vladimirovich Yurinsky. Rough asymptotics of large deviations. Lecture Notes in Mathematics, 1617:123–162, 1995. CO-DEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092603/.

Yurinsky:1995:SGVa

[Yur95f] Vadim Vladimirovich Yurinsky. Seminorms of Gaussian vectors in infinite dimensions. Lecture Notes in Mathematics, 1617:43– 78, 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092601/.

Yurinsky:1995:SGVb

[Yur95g] Vadim Vladimirovich Yurinsky. Sums and Gaussian Vectors, volume 1617 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNMAA2. ISBN 3-540-60311-5 (print), 3-540-44791-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 254 (est.) pp. LCCN QA3 .L28 no. 1617. URL http:/

/link.springer.com/book/10.1007/BFb0092599;   http://
www.springerlink.com/content/978-3-540-44791-7.

Zabczyk:1999:PEH

[Zab99] J. Zabczyk. Parabolic equations on Hilbert spaces. Lecture Notes in Mathematics, 1715:117–213, 1999. CODEN LNMAA2. ISBN 3-540-66545-5 (print), 3-540-48161-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092419/.

Zaidenberg:1990:HRP

[Zai90] M. G. Zaidenberg. Holomorphic rigidity of polynomial polyhedrons and quasihomogeneity. Lecture Notes in Mathematics, 1453: 291–307, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085962/.

Zak:1991:SPD

[Zak91] F. L. Zak. Some properties of dual varieties and their applications in projective geometry. Lecture Notes in Mathematics, 1479:273–280, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086275/.

Zarhin:1991:LIL

[Zar91a] Yuri G. Zarhin. Linear irreducible Lie algebras and Hodge structures. Lecture Notes in Mathematics, 1479:281–297, 1991. CO-DEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0086276/.

Zarhin:1991:UP

[Zar91b] Yuri G. Zarhin. Ussr participants. Lecture Notes in Mathematics, 1479:298–300, 1991. CODEN LNMAA2. ISBN 3-540-54456-9 (print), 3-540-38388-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0086277/.

Zvyagin:1992:PNE

[ZD92] V. G. Zvyagin and V. T. Dmitrienko. Properness of nonlinear elliptic differential operators in Hölder spaces. Lecture Notes in Mathematics, 1520:261–284, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0084725/.

Zampieri:1993:DMR

[ZD93] Giuseppe Zampieri and Andrea D’Agnolo, editors. D-modules, Representation Theory, and Quantum Groups: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Venezia, Italy, June 12–20, 1992, volume 1565 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNMAA2. ISBN 3-540-57498-0 (print), 3-540-48195-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1565. URL http:/

/link.springer.com/book/10.1007/BFb0073464;   http://
www.springerlink.com/content/978-3-540-48195-9.

Zheng:1996:MTB

[Zhe96] Weian Zheng. Meyer’s topology and Brownian motion in a composite medium. Lecture Notes in Mathematics, 1626:108– 116, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094645/.

Zieschang:1996:AAA

[Zie96a] Paul-Hermann Zieschang. An Algebraic Approach to Association Schemes, volume 1628 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 189 pp. LCCN QA3 .L28 no. 1628. URL http:/

/link.springer.com/book/10.1007/BFb0097032;   http://
www.springerlink.com/content/978-3-540-68523-4.

Zieschang:1996:AP

[Zie96b] Paul-Hermann Zieschang. Algebraic prerequisites. Lecture Notes in Mathematics, 1628:65–96, 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097035/.

Zieschang:1996:BR

[Zie96c] Paul-Hermann Zieschang. Basic results. Lecture Notes in Mathematics, 1628:1–32, 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0097033/.

Zieschang:1996:DT

[Zie96d] Paul-Hermann Zieschang. Decomposition theory. Lecture Notes in Mathematics, 1628:33–64, 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097034/.

Zieschang:1996:RT

[Zie96e] Paul-Hermann Zieschang. Representation theory. Lecture Notes in Mathematics, 1628:97–121, 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097036/.

Zieschang:1996:TG

[Zie96f] Paul-Hermann Zieschang. Theory of generators. Lecture Notes in Mathematics, 1628:123–176, 1996. CODEN LNMAA2. ISBN 3-540-61400-1 (print), 3-540-68523-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0097037/.

Ziegler:1998:IST

[Zie98] Martin Ziegler. Introduction to stability theory and Morley rank. Lecture Notes in Mathematics, 1696:19–44, 1998. CODEN LN-MAA2. ISBN 3-540-64863-1 (print), 3-540-68521-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-68521-0_
2/.

Zimmermann:1998:HR

[Zim98a] Alexander Zimmermann. Historical remarks. Lecture Notes in Mathematics, 1685:151–154, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0096373/.

Zimmermann:1998:I

[Zim98b] Alexander Zimmermann. Introduction. Lecture Notes in Mathematics, 1685:1–4, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0096367/.

Zimmermann:1998:OTC

[Zim98c] Alexander Zimmermann. Onesided tilting complexes for group rings. Lecture Notes in Mathematics, 1685:81–104, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096371/.

Zimmermann:1998:SML

[Zim98d] Alexander Zimmermann. Some modular and local representation theory. Lecture Notes in Mathematics, 1685:51–80, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0096370/.

Zimmermann:1998:TAS

[Zim98e] Alexander Zimmermann. Tilting with additional structure: twosided tilting complexes. Lecture Notes in Mathematics, 1685: 105–149, 1998. CODEN LNMAA2. ISBN 3-540-64311-7 (print), 3-540-69748-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096372/.

Zippin:1991:GAC

[Zip91] M. Zippin. A global approach to certain operator extension problems. Lecture Notes in Mathematics, 1470:78–84, 1991. CODEN LNMAA2. ISBN 3-540-54206-X (print), 3-540-47493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0090215/.

Zilbergleit:1990:SCD

[ZL90] L. V. Ziĺbergleit and V. V. Lychagin. Spencer cohomology of differential equations. Lecture Notes in Mathematics, 1453:121– 136, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0085951/.

Zoladek:1990:RDL

[Żoł90] Henryk Żoładek. Remarks on the delay of the loss of stability of systems with changing parameter. Lecture Notes in Mathematics, 1455:393–396, 1990. CODEN LNMAA2. ISBN 3-540-53509-8 (print), 3-540-46722-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0085404/.

Zorich:1992:GHT

[Zor92] V. A. Zorich. The global homeomorphism theorem for space quasiconformal mappings, its development and related open problems. Lecture Notes in Mathematics, 1508:132–148, 1992. CODEN LNMAA2. ISBN 3-540-55418-1 (print), 3-540-47061-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0094243/.

Zvyagin:1992:ODF

[ZR92] V. G. Zvyagin and N. M. Ratiner. Oriented degree of Fredholm maps of non-negative index and its application to global bifurcation of solutions. Lecture Notes in Mathematics, 1520:111– 137, 1992. CODEN LNMAA2. ISBN 3-540-55583-8 (print), 3-540-47223-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0084718/.

Zucker:1991:CSC

[Zuc91] Steven Zucker. L^{p}-cohomology and Satake compactifications. Lecture Notes in Mathematics, 1468:317–339, 1991. CODEN LN-MAA2. ISBN 3-540-54053-9 (print), 3-540-47370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086199/.

Zuo:1999:HMF

[Zuo99a] Kang Zuo. Harmonic metrics on flat vector bundles. Lecture Notes in Mathematics, 1708:25–51, 1999. CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092572/.

Zuo:1999:I

[Zuo99b] Kang Zuo. Introduction. Lecture Notes in Mathematics, 1708: 1–9, 1999. CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092570/.

Zuo:1999:NAH

[Zuo99c] Kang Zuo. Non-abelian Hodge theory, factorization theorems for non rigid or p-adic unbounded representations. Lecture Notes in Mathematics, 1708:52–103, 1999. CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0092573/.

Zuo:1999:P

[Zuo99d] Kang Zuo. Preliminaries. Lecture Notes in Mathematics, 1708: 10–24, 1999. CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092571/.

Zuo:1999:RFG

[Zuo99e] Kang Zuo. Representations of Fundamental Groups of Algebraic Varieties, volume 1708 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 135 pp. LCCN QA3 .L28 no. 1708. URL http:/

/link.springer.com/book/10.1007/BFb0092569;   http://
www.springerlink.com/content/978-3-540-48424-0.

Zuo:1999:SMR

[Zuo99f] Kang Zuo. Shafarevich maps for representations of fundamental groups, Kodaira dimension and Chern hyperbolicity of Shafarevich varieties. Lecture Notes in Mathematics, 1708:104–124, 1999.

CODEN LNMAA2. ISBN 3-540-66312-6 (print), 3-540-48424-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0092574/.

Zvyagin:1990:SSS

[Zvy90a] V. G. Zvyagin. On the structure of the set of solutions of a non-linear elliptic problem with fixed boundary conditions. Lecture Notes in Mathematics, 1453:309–320, 1990. CODEN LN-MAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0085963/.

Zvyagin:1990:PEP

[Zvy90b] V. G. Zvyagin. The properness of elliptic and parabolic differential operators. Lecture Notes in Mathematics, 1453:137–159, 1990. CODEN LNMAA2. ISBN 3-540-53407-5 (print), 3-540-46861-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0085952/.